A Complete Bibliography of Publications in Limnology and Oceanography: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.06

Title word cross-reference

+ [SK96]. ^{−} [LBL^{+}98]. ^{13} [GSC^{+}96, KHRS98, LTPB98, MBK^{+}99, QWR^{+}92, dGF96]. ^{137} [CDP96]. ^{14} [GSC^{+}96, Jön91b, KV99, LID93, Lig92, PSS91, RLL93, SLA93, VV98]. ^{1}5 [BG94b, AR96, HDR^{+}98, KMWK92, PVL^{+}96, SRG98, SRG00, WH94b]. ^{16} [LK94a, QWR^{+}95, RKG97]. ^{1}8 [MZ90, HB94, QWR^{+}95, RA98]. ^{2} [HB94]. ^{2+} [SH98b]. ^{21} [NWS94]. ^{210} [BPS94, LSS^{+}97]. ^{222} [CBBC96]. ^{226} [BPS94]. ^{234} [LSS^{+}97, WM92]. ^{3} [AHKH^{+}96, Bel90, KV99, TB91]. ^{31} [HGR95]. ^{32} [BP97, IH95, LBL91, TLvB^{+}91]. ^{34} [NWS94]. ^{60} [AGZ99]. ^{63} [BW94a]. ^{7} [LBL91]. _{12} [Hof90]. _{2} [MZ90, MZ90, ASS^{+}99, BFZED94, BZR99, CR93a, CI94, CP97b, CPSA90, DHSS95, DHWW97, FCG94, FBW96, GRG98, HA92, HAAM99, JRT96, KMWK92, LBP97, LB95, LBL^{+}98, ME98a, McC91, Mil91, NJLW97, NKT^{+}93, ND91, PB91, SO91, SHK95, WK96, dBGEK97]. _{3} [TM95a]. _{3}^{−} [CP97b, KMWK92]. _{4} [CBBC96, dARB91]. _{4}^{+} [HNS95, HA92]. _{6} [CC98]. _{CO2} [CP97b]. a [Ban90b]. b/a [Ban90b]. β [VM94]. : [BZR99, CGVL95, DM92, FZP92, FVL98, FVL99, FSSI97, HRD^{+}95, LS92, MDE97, MB90a, QWR^{+}95, QP91, dGP94]. δ [HDR^{+}98, HB94, KHRS98, LTPB98, MBK^{+}99, NWS94, PVL^{+}96, RA98, dGF96]. f [NJLW97, QP91, TKL^{+}97, Véz94]. μ [HBH97]. × [DW99].

-fixing [CPSA90]. -II [LBL^{+}98]. -N-acetylglucosaminidase [VM94]. -ratio [QP91, Véz94]. -ratios [TKL^{+}97]. -segregating [HN94].

1-nl [MNAH96]. 12-yr [KFEC91, VL91]. 14^{o} [RDK96]. 150^{o} [CLKJ95]. 15^{o} [FFNS99]. 187-196 [Ban90c]. 1989 [EK90]. 1992 [Ano92a, Ano92b, Ano92c]. 1993 [Ano93a, Ano93b, Sut93a]. 1994 [Ano94, Ano95b].

2-2 [HBH97]. 2-2nm [PW93]. 234Th [HP95, VJM96]. 238U [HP95]. 238U-234Th [HP95]. 27^{o} [FFNS99]. 2nm [PW93].

3-yr [NB97]. 38’S [FFNS99].

5 [AA91a, AA91b, THS^{+}91, YK95]. 5’-nucleotidase [AA91a, AA91b]. 53706 [Cot99]. 53rd [Ano90i]. 57’S [FFNS99]. 57th [Ano94]. 59Fe [Cur93].

75^{o} [SML96]. 7Be [VJM96].

9-m [FNWK96].

a-like [HHR94]. a-specific [DW95a]. A. [TJK^{+}97]. Aarhus [JMA^{+}95]. Abdullah [JD96b]. Abiotic [dHJS90]. above [DMM95]. abraded [Ber99b]. absence [CSS96]. absolute [DW95a, MAN91a, Som90]. absorbed [RF94b]. absorbing [JLO97]. Absorption [BDR99, Gor91, HS92, SM91, AP92, ABB^{+}95, Ano93y, BTLC93, BSMK95, BPCL97, BS90b, CW93b, CH94, CP99, DSV^{+}99, DASV98, FT91, GCP90, GSG94, GB94, HVB93, LL98, MSK^{+}97, NT94, RSK97, Roe98, Sos99, Str90, TF95, VFAA98, VHS^{+}95, SCKF99]. Abundance [CTSB98, PB94, YFF^{+}98, ADC90, Ano93p, DD92b, Li98, LDL91, MHSP90, Pic91, PPS^{+}97, SDJK98, SD95, SD94, WSB98, dGBPP96]. abundances [BFKL91, GAA93, Kam92, SPB90, SBGBC96b, SSS93]. abyssal [LSS^{+}97, PRS94, Smi92, SKB94, SGB^{+}97]. Acartia [Ano93m, DDC92, DD92b, DD92a, EMH95, KBH98, PWR94, STJ^{+}93, Sai94, SB90, Tis92, TJK^{+}97, WLY93]. acclimation [GMK98, LW93, MR92b, MM99, PF91]. Accounting [DC96a]. accretion [CDP96, LS93, PD90a]. Accumulation [HHR94, MAG95, THR97, WF98, And90, BK99, HCHD94, HBI95, JAMH96, MTLR98, SDS95, SRLRL^{+}99, WTH98]. accuracy [FT91, Roe98]. Accurate [NJLW97, Ver93]. Acetate [BDV99, Hin94, WGS97, SW96b]. Acetogenesis [HAAM99]. acetylene [SNB92]. acetylglucosamine [NVH^{+}94]. acetylglucosaminidase [VM94]. Achievement [Ano95a]. acid [BLK95, CH91, GDG99, GMR99, HB95a, HCHD94, NML99, PM90a, Sta90a, TMB96, TK90, dGBPP96]. acid-volatile [HCHD94]. acidic [GAAG96a, GAAG96b, TBM95, Pse92]. Acidification [Ege95, FMK^{+}99, GLM92, KHS^{+}98, LFHS99, RKST90, THS^{+}91, Web95]. acidified [BVSBB93, WHOL99]. acidity [YH92]. acidophilic [Gen91]. acids [BL98, CH92a, DM98, DMHH99, EMH95, Fie97, HCR^{+}94, KK93, KHS97, LLZ99, MSS^{+}95, MAS91, MASA94, NK91, PL94, RS93, RKST90, SRLRL^{+}99, SR92, WG97, WJRB97, WMD90]. Acoustic [GRG99, JJGS96, TT97a, FHF99]. acoustical [Hol99]. acquisition [CP97b, LD99, PRWG97, RJ91, RB91]. Acrocalanus [McK96]. across [BP91, Dua92, GV93, HCS95, Hur90, KMU95, OGH^{+}96, PL95, WMD97]. actinometry [MJC93]. action [SCKF99]. activated [HSSA94, WS96]. Activation [FR99]. active [GW99, JLO97, Jun92, KF93, LPA96, Wal92, ZMP98, dGGV^{+}96]. activities [Mar94, NVH^{+}94, TC93]. Activity [CK95, ASMB98, AA91a, AA91b, Ano93m, BH95, CHA91, CSS96, DCM90, GWY^{+}92, HB95b, JM93a, JRH90, JJGS96, LHWZ92, PH92a, PRS94, PRT^{+}93, RDK96, SLSLBD97, WM92, Wil90, YD99, dARB91]. acutifrons [GMR99]. acutus [LRvE94]. Adam [Man94]. Adams [SW93]. Adaptation [Lam94, DBS99, LB96]. adapted [NS93, OCP92]. Adaptive [AB99, LGW91, Pla98]. added [SSM^{+}95]. Addenda [Ano96a]. addition [CC98, LB95, MT99, RKST90, RB97]. additional [JF99]. Adhesion [JAS95]. adhesive [AWL94]. adhesive-filament [AWL94]. Adirondack [MHZ98, YAF^{+}92]. adjacent [HK95, RJC^{+}97, Sti91]. Adriatic [MMSKH94]. adsorption [DBT91, RS99, SMV98, ZMP98]. adult [SCFB98]. Aerobactin [HHB93]. Aerobic [KAD95, JA92, SCK93]. aeruginosa [FSSI97, OJ98, RKW94, WH99]. affect [Ano93l, LWS^{+}96, Lei90, MEM99, Paw98]. affected [NML99]. affecting [BKT96, GNL97]. affects [HJ94, LK99, Rot95]. affinis [Ano93n, BK90, HQCG92, JW92]. Africa [Mei90, HBO^{+}94, Kil90, LB94]. African [Hec99]. after [Cul95, HKvS^{+}92, KHS^{+}98, MHIHH99, RG92a]. against [FL95, GLS90, Jun92]. Agatti [MAN91a]. Age [MLBC94, DD92a]. aged [Pro90]. agent [CHC97, FMK^{+}99]. agents [SCPS98]. ages [CBD^{+}99]. aggregate [DFA93, KTMI^{+}98, KW98b]. aggregate-attached [DFA93]. aggregates [AV94a, AV94b, GS93, GSL97, LW90, SCPS98, TK98]. aggregation [DDG94, OC97, WJRB97, WMML98]. aggregations [MMSKH94, PRCV98]. aggression [PHMH96]. aging [SGP96]. agreement [FDLD93]. agriculture [HLQ^{+}99]. Aiptasia [CMPD92]. air [AR95, Bal96, Con90, HCS95, Liv99, PL95, RERB91, SHK95]. air-water [SHK95]. airborne [VHS^{+}95]. al [Jør93, Sta91, UGU^{+}90]. al. [CK92a, CA93, DW95b, Gib92, HSM^{+}95, Lap99, MSH96b, NRPRB96]. Alaska [Wet98]. Alaskan [KE91]. albedo [MMG94b]. Alberta [Ano94, MHSP90, SSFP90]. alcohol [EMH95]. aldehyde [NSK95]. Aldoses [SB97]. Alexandrium [And97, TOKA97]. alga [GAAG96a, GAAG96b, Hof90, Joh94, LRvE94, PSJ99, SH98b, WH92]. algae [AECP95, AKM95, AM95, CHA91, Edm96a, FK98, HK98b, HHR94, HW94, SJP94, Ken93, Leh95a, MHM^{+}99, NT93b, PSEJ99, RHT92, VBAO95, WVP97, WC91]. Algal [ES90, Han96, HSS^{+}99, JM93a, LFHS99, Rot95, ADR93, AS94, AMP98, Ban90a, BLDP96, BMAW97, CW93a, CCDL97, DO97, FZD93, FPL^{+}91, FPC99, FG93, Fra97a, Fra97b, Han92, HGP97, JL92, LB91, MASA94, MSK^{+}97, Pae97, PW93, PFDH93, Plu97, RLL93, RBH99, SWS^{+}96, SLC^{+}99, Sma97a, Som98, SG93, TBM95, TMB96, VV98, VCT^{+}99]. algal-rich [MASA94]. algorithm [VFAA98]. alimentary [Bre97]. alkaline [BW90, JCM^{+}99, SMH94]. Alkalinities [Sta90b]. alkalinity [CG91, GLWH90, KGFP91, Sta91, WMD90]. Allan [RW95]. Allanson [Mei90]. allelopathic [vEJ97]. Allen [Hur96]. allochthonous [VL98]. Allometric [NSJ90, WVP97]. allometry [Gla91]. allozyme [ČH99, SY95]. ALOHA [DK96, Ven97]. along [CLKJ95, CCT93, GA96, JD96a, JD96b, Leb90, LM91, RDK96, SML96, SR95, SM99b, War90, WAC^{+}98]. Alosa [TJC^{+}98]. alpine [Ano93i, GLM92, MSP^{+}99, MWI92, Sta90a, VL96, VL98]. also [LPA96]. Altamaha [CW98]. alteration [CH96, DMHH99, JCH92]. Altered [VLHL97]. alternative [KCC98, RKW94, TF95]. alterniflora [BFS91, LKD99, WH94c]. Alteromonas [RB91]. alters [HB91]. altivelis [WS90]. aluminum [Alt90, Gen91]. alvinellid [CJVD97]. Amazon [BOCL^{+}95, ES90, EM93, GW94, HCR^{+}94, Les95, QWR^{+}92, QWR^{+}95, RHD^{+}90]. ambient [HDMB97, PD90b, RB97]. ambigua [Han91, SP90a]. America [Fri96, HH93]. American [Ano94, Ano90a, Ano90i, Ano91f, BG94a, TJC^{+}98, WSM^{+}96]. americana [Gen91]. amides [PH97]. Amino [DM98, HB95a, PM90a, TK90, CH91, CH92a, DMHH99, Fie97, FR99, GMR99, HCR^{+}94, KK93, LLZ99, MSS^{+}95, NK91, NVH^{+}94, PL94, RS93, SRLRL^{+}99, SR92]. ammonia [JCM^{+}99]. ammonification [SK90]. Ammonium [CMPD92, HK95, KSC^{+}90, TA97, WK90, Arm99, EW99, HHI96, HFK92, HS90, Jon91a, KKW90, PVL^{+}96, Roc98, SK90, SFC90, TK90, WHN^{+}98]. ammonium-rich [EW99]. among [BG94b, ČH99, Cur90, FCMB99, FFF98, FDLD93, KK91, LMH98, MW98b, MC99, Som94, SH98b, TB91, dHK98]. among-lake [dHK98]. amorphous [MMSKH94]. amperometric [GK97]. amphipod [MW98b]. amphipods [HQCG92, TT97a]. amurensis [WH93a]. Anabaena [Gil94, Gil96, HRF91, SWJO98]. anaerobic [AR94, KAD95]. Analog [PSH^{+}96]. Analyses [MHZ98, CBFC98, CIM92, Joh91, UMS^{+}99, YMT^{+}99]. Analysis [Cot99, DŠ99, KSH96, Sch92, THK^{+}99, Ano93r, Ben94, BBH^{+}92, BdBC99, ES90, ECA99, Gae93, HA92, HRD^{+}95, Kap94, Kna99, KW98b, LBH^{+}93, Li90, Log93, LVFD93, MS93b, MSH96a, Mil95, NJLW97, Pol95, Pse90, SH91, SY95, SMC98, SFFF97, TB96, VSBZ^{+}94, Ver97, VPBD97, WW95, WGSP91, Wel94]. Analytical [Swa99]. analyze [WDGS92]. analyzer [YK95]. analyzing [KW98b, TB96]. Ancient [MBHM99, Cha96]. Anderson [CGSD93]. anemone [AMP98, CMPD92]. Angel [Cha96]. Angeletti [Mat94]. angle [Ban90b, HM99]. Angola [FFNS99]. Angular [GVK93, VdHH^{+}98]. anhydrase [BFZED94, YM96]. animal [FPZFW99, Som98, TTS93, Ano93g]. animal-to-animal [TTS93]. animals [CHC97]. annandalei [TT97a]. Announcement [Ano90b, Ano90c, Ano90d, Ano90e, Ano91a, Ano91b, Ano91c]. Annual [Ano90i, Ano92e, Ano94, Li98, PSH^{+}96, WH91, BFZED94, BN90, HY97, KKK^{+}93, Nür98]. Anodonta [CCT93, TCCA93]. anomalies [MKD90]. Anomalous [BA92, Rah90]. anomaly [CG91]. anophagefferens [BL97]. Anostraca [BW99a]. anoxia [Nür95]. Anoxic [SLA93, AEL^{+}99, Ano93z, BMP92a, CL91, CPCS93, CP95a, HTDM96, HSCM97, HAAM99, JUSI91, JSL92, LLB^{+}96, MMS93, MLBC94, OBL^{+}91, PPK95, RT92, RC98, SDS95]. answers [SC93b]. antagonism [FCMB99]. Antagonisms [SH96b]. Antarctic [Pri94, CL96, Dal90, FAE99, HK98a, HVHH91, HG99, KWS98, MASA94, MA96, Mol98, MWW97, PDM96, RGR^{+}98, TCB^{+}98, VRL^{+}98, WMM^{+}94, Amm96]. Antarctica [FP99, HHJ98, KL90, MTM^{+}98, MAS91, MM99]. Anthopleura [AMP98]. anthracene [GDG99]. anthropogenic [AGZ99, GNL97, Kli92, MBCB97]. antiqua [AWKH98, WKK91, WKK^{+}95]. AO [SSS93]. Apache [KSH^{+}99]. Aphanizomenon [HSI95]. Aphotic [LC90b]. aplastidic [Ver91]. Appalachian [Hur90]. Apparatus [BS90a]. apparent [RNR98]. appendage [LB96]. appendages [HSI95]. appendicularian [BDR99]. Applicability [Ano93c]. Application [CBBC96, EF95, LVFD93, SHB99, TBK98, Alt94, Ano93u, HTSDZ99, Liv90, Sti91]. applications [SS99d, Bur94]. applied [BW94a]. applying [WW98]. appraisal [Hur96]. approach [AO96, Ano92d, CE96, JJGS96, KRW99, LL91, MSS^{+}95, SSFP90, SHB99, TF95, WW98, Wil98, YD99, Fen94]. approaches [CP94, Roe98]. Approximation [JF96, JF99]. April [EK90]. Aqaba [LP95]. aquae [Gil94, Gil96, HSI95]. aquagenic [WJRB97]. aquaria [Ovi99]. Aquatic [GG97, MAS91, MASA94, Mor90, SW93, Ant99, CWW98, CW93b, CPSA90, Dua92, Duc94, EW99, GV93, GC97, Hai90, HT98, HA90, KL90, KT92, KMK95, KJSJ98, MB98a, NLSG99, RPJ97, PL95, Pse92, RJC^{+}97, TF95, VPBD97, Wet91a, YSW90, Fen94, Man95]. aqueous [KMWK92]. aquifers [Mat93]. Arabian [BDY^{+}98]. Arcady [Edm96b]. Arcady-Meyer [Edm96b]. archaea [MTM^{+}98]. Arctic [Arn98, DMA^{+}96, Eic94, HD97, MVT99, MHM96, OAF^{+}99, OBA99, PSS^{+}96, YD99]. Area [SBP^{+}96, BC97, BL94a, BS90b, Ven97]. areal [TKAD95]. areas [DWM90, Fro91, Vos92]. arenaria [EG91b]. arenicolid [KW93]. Argentina [CNP94, CP94, CV99]. armatus [PRS94]. array [KJ92]. arrays [NK99]. Arsenic [KS93, DBT91]. Artifacts [Str90]. Artificial [KLR94, BLD99, BW90, PTJ96, SF95]. artificially [MGW96]. ascending [CCK98]. ascent [MV96]. Asiatic [FRLS90, FFJ92]. ASLO [Ano93a, Ano95a, Ano95c, Wet93]. ASLO/ONR [Ano93a, Ano95c]. Asparagopsis [MHMD99]. aspects [VPBD97]. Assay [Cam94, BH95, PA95a, Tra97]. assays [OCS96]. assemblage [BUW96]. assemblages [BTLC93, DFA93, FCC99, FR91, FGN93, GHT^{+}96, LF91, PG97, SLD97, TM95b]. assessed [PTJ96, SB94]. Assessing [HPTC97, RFHG99, TKAD95]. Assessment [JOG96, TBM90, CBBC96, CCDL97, HTW90, Sch90]. assimilated [BDR99, CBD^{+}99, KHB96]. Assimilation [AMF98, WRLF96, WF96, Ano93x, CLMH93, DL96, Leh93, McC91, RF94a, SK90, WHN^{+}98, WH93a]. associated [HFK92, Kim92b, LCG94, ŠKN^{+}99, SHF^{+}96, THS^{+}91]. Association [LDL91, MDWS97, MHSW99, SE90]. assumptions [CL90]. Astaxanthin [JOG96]. Asterionella [Gen91]. astreoides [Gle93]. asymmetry [CP97b]. Asymptotic [McC92, FM94]. Atlantic [GSC^{+}96, HC96, Lap95, Liv99, MJC^{+}98, OGH^{+}96, VBD^{+}97, WBP99, WL94, BKT96, BKH^{+}96, BUW96, BL97, CCK98, EK90, FBS99, FFNS99, HWP95, HS92, KDKS90, MBKG95, OCZA90, PAD^{+}99, PRCV98, PRT^{+}93, RYC99, TJC97, TV99, VDA99, Was90, WG94]. Atlas [Ohm96]. atmosphere [ASS^{+}99, Mat94, MGW96, RKR91]. atmospheres [GC90]. Atmospheric [CC98, DT91, FBW96, HKPM99, SS99b, CCL90, FCG94, KPSB95, LWS^{+}96, Pae97, PP97, PB91, SO91]. Atoll [MAN91a, LJBTB99]. atomic [SBW^{+}98]. ATP [ZF93]. attached [DFA93]. Attenuation [WSM^{+}96, AMLH97, BSMK98, FM94, GCP90, KJSJ98, McC92, McC95, MZW^{+}95, Vos92, VBK98]. attractant [ZF93]. attributes [SR95]. Auhcoseira [SWJO98]. Aureococcus [BL97]. aureolum [Ano93v, NE93]. austral [NCD98, SN90]. Australian [LPPSR97]. australis [SMB^{+}97]. Authors [Ano90f, Ano91d]. Automated [Car98]. autonomous [SGB^{+}97]. autotrophic [DŠ98, DC96b, JBJB96, PSH^{+}96, SPB90]. autotrophs [FBS99]. autumn [SN90]. Availability [RM90, AML93, BFZED94, BR92, CV99, FRY95, FFF98, Gil96, IM98, IKT99, JMMH98, LM91, PASC95, PY99, PL95, SZA92, Wel99b]. available [MR99]. average [BSMK95, BSMK98, Kir91, Li98, SSS93]. averaged [And93]. avoid [GD96]. Avoidance [LL96]. Avrainvillea [HDP^{+}90]. AVS [HCHD94]. Award [Ano95a, Wet93, Ano90g, Ano91e, Ano92a, Ano92b, Ano92c, Ano93b, Ano95d, Sut93a]. Awards [Ano92e, Ano93a, Ano95c]. Azimuthal [BC90]. azooxanthellate [FYG98]. Azotobacter [WH95]. aztecus [CBS99].

B [Cha96, Duc94, Edm96c, McC97, Mei90, MDdM^{+}99, Tho90, Thr90, Wet91a, RR92, Wel94, Hof90, HHL98, NE93, RHH97, SRLRL^{+}99, Wil95]. back [Muc91]. back-titration [Muc91]. backwater [KV99]. Bacteria [Bre97, Duc98, THK^{+}99, PAS90, Ano93t, AR94, BP91, BS90a, Bre91, CTF98, CBD^{+}99, EMH95, GV93, GLS90, GP99, HB95a, HK95, JC95, KK93, KSSP91, KKW90, KJLA98, LK94a, LID93, LJD95, Li98, PC96, PH92a, Pse90, RS93, SFRB98, SS99b, SR92, SCK93, SBGBC96b, SCEG95, TZR98, Tra94, TK90, VORJ93, VBAO95, WC91, dGGV^{+}96]. Bacterial [AA91a, AA91b, AB96, BOCL^{+}95, CK95, CPCS93, CP95a, CNL90, DM90, Kir99, Kro93, MH90, PLTL96, SBB99, ZFdSY96, Ano93k, Bel90, BB98, BFKL91, BSC93, CPS90, CP95b, DFA93, FPL^{+}91, FN95, HPM96, JRH90, KCC98, KV99, LPA91, LDL91, LGT95, MHZ98, MH95, MMSKH94, MBKG95, OBH^{+}91, PMF90, PS92, RDK96, SDJK98, ŠKN^{+}99, SBGBC96a, SH98a, SSS93, TB91, Tra97, VHH96, dGBPP96]. bacterially [MB98b]. bacteriochlorophylls [HW91]. Bacterioplankton [BS99, MHSW99, dGGV^{+}96, BL98, BOB94, CSS96, CW92, CW95, Cur90, FFF98, JBJB96, LF91, LPA91, MD90, ML92, MNAH96, MBKG95, ORH99, RA97, SD95, SD94, SFFF97, SLT99, SH98a, TH98, VVIK99, WH93a, WH91]. bacterium [HFK92, OCP92, RB91]. bacterivore [JRH90]. Bacterivorous [SP90a, JD95, NK91, Rot92]. Bacterivory [CM91, NT93b, SBGBC96a, SPB90]. Baetis [Wil90]. Bahama [SH90]. Baikal [SS98, Ano93q, CG97, FMH^{+}91, FCM^{+}97, Gra97, HKP^{+}97, KCWM96, Liv99, PCFS97, PPS^{+}97, SGZ93, YMT^{+}99]. bakeri [MP92a, RB94]. balance [CPMW99, CE98, CCP92, DGS98, Leh91, MM91, MB99]. balances [CG97, JRH90]. Baldeggersee [TMLS99]. Baltic [KKM^{+}96, KHK^{+}98, BL94a, BR92, HB95b, KKK^{+}93, SH96a, Sti91, VRHK99]. bands [Ken93]. bank [MLBC94, MMPS98, RYC99]. banks [Hai96b, PH92b]. Bannister [Kie90]. Bannister-Laws [Ban90a]. Banse [MBFG90]. Banyoles [CR93b]. Barbara [ALB90, Cot99, OC97, RLTB90]. Barber [War90]. Barents [KFW94]. barite [FRLS90, FFJ92]. barnacles [CR99]. Barrier [TM95b, HSM^{+}98]. bars [SGZ93]. based [Arm99, BF97b, BL94b, DHWW97, ER95, Fri90, Gae93, JOG96, LBH^{+}93, LDC99, MSS^{+}95, MAS91, MLB99, MBSW98, OGH^{+}96, SWS93, TKL^{+}97, WPBS99]. Basic [TZF96, DBZF98]. Basin [RLTB90, And90, FNWK96, GW94, Hur90, LFGM96, MLBC94, SDS95, TGSB96, WBM93, ALB90, MBK^{+}99, MQ96, OC97, QWR^{+}95]. basins [AHKH^{+}96, Ber91, CE96, RWH^{+}98]. basis [BH97, IPP97, MUGK92, TNG^{+}92]. basket [MC95]. batch [LTPB98, NJLW97]. Bathymodiolus [TV99]. Bay [BW99b, CT96, DFN97, JMA^{+}95, SSP^{+}94, DASV98, Ker95, Kit96, BKB97, BN90, BN95, CCR^{+}95, CCB^{+}99, DDC92, DD92a, FZP92, GGFR91, HCP90, KSC^{+}90, LL98, LvGLC98, MD90, MPG95, MHSW99, PD90a, PBL99b, PWR94, RGRW93, SD94, WM92, WBKF95]. Be [LBL91]. beach [LS93, MJG91]. beach-stranded [MJG91]. beam [AMLH97, MVM98, Vos92]. bearing [GAA98, YK95]. beat [LB96]. beaver [YAF^{+}92]. bed [Dad93, FHF99, NSBB98, OMk93]. bedforms [MCED94]. bedload [EG91b]. bedrock [BL94a]. beds [BK99, JMMH98]. been [FRL94]. before [MHIHH99]. Behavior [BG97a, Tis92, Ano93z, Ber99a, BFM^{+}97, BPA96, CHC97, DFN97, DW99, DC91a, NFGP^{+}93, GPMF96, GWY^{+}92, HG99, KDB91, LHS90b, MK95, PB93, Pla98, RS99, Sai94, SB93, TJK^{+}97, Tjo90, TTS93, vLM91]. Behavioral [JD95, LLZ99, YW93]. Belostoma [RB94]. belowground [BFS91]. beneath [ESG91, MJ94]. Benguela [RV99]. Benthic [AD92, KV99, Smi92, AML93, AK92, Ant99, BR91, CFP97, CWSX96, CNSR90, CTP^{+}95, FG93, GGRJ94, GJ96, GG90, GW95, GGJ97, HCHD94, HLL98, HCJ96, HJ94, HQCG92, HM99, HS99, JW92, JJGS96, KAB92, MB98a, MSH96a, MWW97, MB90b, OMK95, PPP97, RPRNR94, PZ93, RG92a, SA94, SKB94, Som97, SB90, TG92, WTH98, WO98, Wil90, WLBQ95, YK94]. benthic-pelagic [MB98a, YK94]. benthos [Ano93g, Str95]. Bering [GCD90, Ree90]. Bermuda [BCM92, CPLD99, Goe98, JMMH98, NB97, PPK^{+}99]. Bermudian [MSSM96]. beryllina [RF94b]. betaine [KW98a]. Better [GG92]. between [AHKH^{+}96, Ano93x, BTL^{+}95, BPK94, BPS94, BS90b, CACM98, CSS96, CGVL95, CP97b, EMH95, FFNS99, FRY95, FGN93, Fra95, GV93, Gei90, HBH94, Hay96, HDR^{+}98, JB91a, KHP91, LPA91, MMCL94, MS93a, MDWS97, MHZ98, Mof94, OJ98, PASC95, PRT^{+}93, QP91, Ree90, RV99, RA98, RG92b, RJC^{+}97, Sch90, SZW95, SWJO98, Smi91, SH92, SH96b, TZGS98, VRT^{+}92, Was90, Wic95, dGF96]. beyond [Thr92]. biases [McL97]. Bibliography [Ano93d]. Bicarbonate [MT99, McC91]. Biel [AGZ99]. Bight [CM94, GSC^{+}96, HC96, LBL91, MJC^{+}98, RG92a, RYC99, VBD^{+}97]. binding [NLSG99, PA95a]. Bio [HS92, SCKF99, SCF99, Ano93x, AS94, FAE99]. Bio-optical [HS92, SCKF99, SCF99, Ano93x, AS94, FAE99]. bioaccumulation [CBS99, WB92]. bioactive [BDH91]. bioassay [TKAD95]. bioassessment [CWW98]. Bioavailability [PP97, SLT99, GHF^{+}97, ORH99, SPMW97]. Bioavailable [MSS^{+}95]. Biochemical [AV94a, MN95, BDR99, CH96, EA99, ER95]. Biochemistry [Plu97]. biodegradable [VVK97]. Biodegradation [WH95]. biodiversity [TSD97, SW93]. Bioerosion [LL91]. biofilm [FL95]. biofilms [NLSG99, SAG98, dBGEK97]. Biogenic [SLSLBD97, Con97, Kli92, TLvB^{+}91]. biogenous [BS90b]. Biogeochemical [BE94, JEP^{+}99, BMP92a, Bra91, BZ98, FC91, HTDM96, HSCM97, HL94, HWF95, KHK97, MDC95, MFH^{+}93, Pri94, Sut93b, WAC^{+}98, ZFdSY96, MJOC93]. Biogeochemistry [RHD^{+}90, HP95, Kir99, LBLO91, Mel99, WCD^{+}94]. biogeography [Man91b]. bioirrigators [MW98a]. Biological [BHAV91, ES93, FAC^{+}99, LM91, NSK95, Sal92b, SM98a, SH95b, BDH91, But95, CJVD97, DO97, HDB93, KL90, KSYK96, LSB92, LW90, Pla93, YFF^{+}98, Ban90c, Jum94]. Biological-physical [Sal92b, DO97]. Biologically [HSFR98, KL90, MZ97]. Biology [Whe91, Ano93h, Ban90e, HPB^{+}99, KMK95, Lew98, Pai96, Plu97, Smi96, Str95, Men90]. bioluminescence [BS90c, LCG94]. bioluminescent [GWY^{+}92]. Biomanipulation [DFM92, CK92a]. biomarker [CFW97]. biomarkers [BLS^{+}95, BdBC99]. Biomass [CW95, GdGD97, AC92, BFS91, BdBC99, BP96, CCKP98, CAS93, Coa91, CPCS93, EP97, FZD93, Gae92, Gae93, GG90, HK98b, Han92, LL98, Lew90, LID93, MD90, MC92, MBHH^{+}91, NK94, PS92, RG92a, TN94, TL94]. biomimetic [MSS^{+}95]. Biomineralization [FRLS90, WBG^{+}91]. biomonitor [HT98]. biopolymers [WJRB97]. biosphere [Lon91]. biota [Dum98, SBB^{+}93b]. biotic [Hai96b]. Bioturbation [MBS98, SS99c, Swa99, Whe92]. birds [Ano93l]. birth [Pol95]. bisphosphate [PPK^{+}99]. bivalve [CCT93, HBI95, MKT^{+}90, OMK95, RF94a, TCCA93, WMTB93b, WSNM98, WWW98, YK93]. bivalves [CH94, DL96, Jør93, LL98, OMk93, WMTB93a, Wil98]. Bivalvia [FRLS90]. Bivariate [Li90]. Biwa [RWH^{+}98, Spi96, TT97a, UNK95]. Black [OCP92]. Blackburn [Ano93f]. blackwater [CM91]. blast [WMD90]. Bloom [And97, ZPHH98, BKT96, BKH^{+}96, BFZED94, BFE96, BN95, BMAW97, Cra95, DFN97, DO97, DDC92, DD92a, HHR94, JL92, KOWH99, KTMI^{+}98, LvGLC98, MHIHH99, Mol98, Mor97, Nie91, NZHB95, PVL^{+}96, Sma97b, SP90b, TSV^{+}91, Véz94, WKK^{+}95, WMPC^{+}90, WGS97]. bloom-forming [BFE96]. Blooms [PBL99b, BHAV91, CDHS98, CP97b, CCDL97, Fra97a, Fra97b, FHN^{+}96, GJC92, HGP97, HSS^{+}99, HvOW99, IM98, KKM^{+}96, Lap97, Lap99, MBHM99, MSK^{+}97, Pae97, SdBdJL98, Sel94, Sel97, Sma97a, TS97, VMH^{+}97]. blue [BFGS94, HW94, MBHM99]. blue-green [HW94, MBHM99]. bluefish [HC96]. bluegill [EEL97]. bodies [AMLH97, Kim92b]. Bodo [DŠ98]. Body [DDC92, GL94, AML93, PAS90, Bra98, FL90, GU94, HBH97, ML90, RV99, TYL92]. Body-size [GL94, TYL92]. Bolivia [GW94]. bon [SHM95]. Bone [Ano93h]. Bonney [FP99]. Book [Abb90, Alt94, Amm96, Ano93h, Ano93e, Ano93f, Ano93g, Arc93, Ban90c, BCBC93, Boo94, Bur94, But95, CT96, Cha96, Con90, Cur94, Dol91, Dua95, Duc94, Edm96a, Edm96c, Edm96b, Fen94, Hai96a, Har95, HL94, Joh91, Jón95, Jum94, Kim92a, Leh95a, Leh95b, Lew95, Log95, Man94, Man95, Mat94, May92, McC97, Mei90, Men90, Mil95, Ohm96, Paa95, Pai96, PA95b, Pla93, Pri94, Pse92, Rav95, RW95, Sal92a, Sal92b, Sel94, Smi90a, Spi96, SS93, Str94, Str95, SW93, Sut93b, Thi94, Tho90, Thr90, Ver90, Wal96, War90, Web95, Wet91a, Wet95, Wet91d, Whe91, Won91, Wri90]. borderland [Ber91]. bore [Pin99]. boreal [Boy94, SBP^{+}96, Was90]. bores [LWMD96]. Bosmina [Ker95, SS99a, BL91]. Bosque [KSH^{+}99]. Boston [TGG94, GG90]. Bosumtwi [TGSB96]. bottleneck [Hen98]. bottom [ADC92, CJVD97, DMF^{+}98, GCD90, Lap97, Lap99, Leh91, MMG94b, MMC94, PS92, RT92, SKB94, SGB^{+}97, Ver93]. bottom-transecting [SGB^{+}97]. bottom-up [ADC92, Lap97, Lap99, Leh91, MMC94, PS92]. bottoms [Eck90]. Boudreaús [Swa99]. bound [UMS^{+}99]. boundaries [War90]. Boundary [MFJR99, BFGS94, GGRJ94, JD90, MB90b, OMk93, OMK95, PSEJ99, PSJ99, Smi92]. boundary-layer [MB90b, OMk93]. boundary-layerflow [BFGS94]. Box [WL99, BR91, MM91]. Boxshall [Ano93e]. bphosphorus [CW92]. Brachionus [Rot90a, Som98, Ste90]. brackish [Kit96, ND91, YK93]. Branchinecta [BW99a]. branickii [Ano93q]. Bransfield [FAE99]. Brazil [MBK^{+}99]. break [Liv99]. break-up [Liv99]. breakdown [JvROG99, JM93b]. breakup [ARM96, WMC^{+}96]. Brett [Hes93, UW93]. breve [MSK^{+}97, TS97]. Brevoortia [CCK98, KDKS90]. briefly [SRPG92]. Brine [FR99, FH92, PD90b, WDGS92]. brine-filled [PD90b]. bromide [MGN92, SM98a]. bromine [GNL97]. Bromley [Whe91]. bromocarbons [MHMD99]. bromoform [GNL97, MGN92]. Bronk [SRG00]. brooding [Gla91]. brown [BL97, BMAW97, CBS99, Gle93]. bryozoan [Grü97, PHMH96]. bubble [BS90a, ZMP98]. bubbles [RC98]. budget [Con97, DASV98, FG93, Gra97, KCKC90, LD99, MLR^{+}94, SM98b]. budgets [BL94b, CPS90, DCML96]. buffer [FL95]. buffering [CV99]. building [But95, VSBZ^{+}94, Ovi99]. bulk [DMHH99, DO98, SH95a]. Bults [Ano93h]. buoyancy [VC90, WH99]. buoyant [MV96, PPK^{+}96, SWH92, SAG98]. burdens [BK99]. Burial [AZS^{+}98, San92]. burrow [Mar94]. burrows [Web92]. bursting [BS90a]. Bythotrephes [BL94b, HY97, BG94a, LC93, SY95].

C [Ano93h, Boo94, Cha96, Jum94, Leh95a, May92, Mei90, MMH^{+}95, Pai96, Str95, SW93, Wal96, War90, Web95, Alt90, BLK95, BZR99, FZP92, FVL98, FVL99, GHF^{+}97, GSC^{+}96, HTDM96, Jön91b, KHRS98, KV99, LTPB98, Leh93, LID93, Lig92, LLB^{+}96, MBK^{+}99, MB90a, PSS91, QWR^{+}92, RLL93, SLA93, VV98, dGF96]. C-incubation [Jön91b]. C-labeled [SLA93]. C-leucine [KV99]. Cadagno [SFRB98]. Cadmium [LRM95, MHT99, Bru92, CCT93, SH96b, WH92, WTH98]. caging [HQ90]. Calanoid [AE91, KBH98, LHS90a, PK94]. Calanoida [vLM91]. calanoids [Har96]. Calanus [BOFW93, CKR93, CR96, CH96, DFN97, GH95, JOG96, MHIHH99, MMPS98, OR94, OC97, PF91, TTS93]. Calcification [McC91, MLR^{+}94, ASH99, CG91, FCG94]. calcite [HSFR98]. calcium [ASH99]. calculated [BH96, EKT99]. Calculating [AMLH97]. Calculation [MFLR96, McB92]. calculations [Pol90]. Calibration [SJHS98, TMLS99]. California [CM94, Cot99, JGRR99, LBL91, Ber91, BW99b, BPC^{+}97, FBR91, HDZB98, JPG90, JM93a, JM93b, JMMD93, JRM98, LKSA97, MFJR99, MGN92, MJOC93, RLTB90, SSP^{+}94, WBM93, WBRL98, WB96]. calm [MK95]. calyciflorus [Rot90a, Ste90]. camera [KW98b]. Cameroon [BGC94, BGN98, NKT^{+}93]. Can [CCP92, WW98, GSG94, GD96, Hai96a, Lei90]. Canadian [FHKC96, RKR91, SSM^{+}95, SCEG95]. Canary [FNWK96]. canopies [LL95]. canopy [CW93a]. Canter [Edm96a]. Canter-Lund [Edm96a]. canyon [HLL98]. capacity [IMB^{+}99]. Cape [CT96, DHWW97, PPK^{+}99, RKG97]. capillary [DSV^{+}99]. Capitella [FL90]. Capone [Cur94]. Capriulo [Dol91]. capsular [SH98a]. capture [ADM96, Lou90, SC90, Swi92, WSNM98]. car [SHM95]. car-bon [SHM95]. carbohydrate [BB97, KHB96, SGA98, UPP95]. carbohydrates [GHT^{+}96, GK97, HCR^{+}94, PB94, ZMP98]. Carbon [AH91, BBS97, BB97, CPS90, CGD90, DK93b, HSM^{+}98, HAL90, KKW90, MVT99, MLR^{+}94, OAF^{+}99, PL95, QWR^{+}92, SR92, TC94, TC95a, WL92, YKM95, AEL^{+}99, Ano93x, Ano95e, BTL^{+}95, BOCL^{+}95, BFE96, BLS^{+}95, BOB94, BGC94, BGN98, BLDP96, BdBC99, BR92, CW98, CPMW99, Car98, CDHS98, CCKP98, CKCP99, CWSX96, CBD^{+}99, CCC^{+}96, CE96, CGVL95, CCP92, CC98, CMPD92, CNL90, Cur93, DGS98, FBS99, FSGG93, FPC99, FG93, Fra95, FCG94, GJ90, GLT96, GGJ97, GSW95, GSC^{+}96, GW94, HCH93, HG99, HCLD95, IM98, JC95, KL96, Kap92, Kap94, KMC90, KCC98, KZM90, KV99, KRB98, KHB96, KCKC90, KAB92, LS90, Lan93, LBP97, LTPB98, LFHS99, LM91, LDL91, Lon91, LAC^{+}96b, MM93, MDC95, MB90a, MA94, MMH^{+}95, MBHT94, MH90, MZW^{+}95, MH97, MB99, NZHB95, PBL99a]. carbon [PPS^{+}97, PHX92, RJ91, RHD^{+}90, RLL93, RC95, RW96, Rom91, SH95a, SC93a, SML96, SBB^{+}93a, SB97, Smi92, SSM^{+}95, SM98b, SBB^{+}93b, SM98c, SBZD97, SFC90, TC96, Thi94, VV98, VRT^{+}92, VVIK99, WF96, WMPO99, YK95, ZEG^{+}94, ZWH^{+}95, dGP94]. carbon-14 [SSM^{+}95]. Carbon-isotope [TC94, BGC94, BGN98]. Carbonate [BFW92, GAA93, WBG^{+}91, CR93a, DMM95, GAA98, JMMH98, JL94, RA98, YK95]. carbonate-bearing [YK95]. carbonic [BFZED94, YM96]. carbonyl [KZM90, UA97]. carboxylase [PPK^{+}99]. carboxylic [BL98]. Carder [Str94]. Caribbean [AB99, Gle93, PG97]. carinata [HSI95]. Carlo [Gor91, Kir94b]. carnivory [ZF93]. Carolina [DHWW97, KMC90, MB90a, RKG97]. carotenoid [DB93, MEM99]. Carpenter [Cur94]. carrion [PL91, Sau92]. cascade [Ano93l, CK92a, Ham94, JPG90]. cascades [ME98b]. Case [Ant99, Wet91d, Gae93, LM98b, MGF91, MSK^{+}97, Mor97, Pse92, SS93]. Caspian [Dum98]. cast [WDGS92]. Castle [JPG90]. catchment [HCLD95, WLCG96]. catechol [WMT98]. catechol-type [WMT98]. cause [FN95, Les96]. caused [BC97, CM94, CA91, CGSD93, PS92, THR97]. Causes [BL97, FBR91, HSS^{+}99, RG92b]. Caution [SK96]. Cayuga [WPAA95]. Cd [LL98, TCCA93]. CDOM [Vod92, VBD^{+}97]. cedar [SM90]. cedar-dominated [SM90]. cederstroemi [SY95]. Cell [LV98, PL94, Ano93r, Ban91, Bel90, BF98, CPCS93, DO98, DOC99, LK94a, LDL91, MFLR96, OJ98, PRWG97, SSS93, TOKA97, VLHL97, VRT^{+}92]. cell-specific [MFLR96]. Cell-surface [PL94]. cells [Ano93j, BSC93, CSS96, CPCS93, OCS96, OP92, VBK98, Web90, WH94a, WWT^{+}93b]. cellular [BH97, Hof90, HM90, PSJ99, SH98b, SH98c]. cellulose [GCD90]. Center [Cot99]. centers [SCPS98]. central [BKT96, BKH^{+}96, BDY^{+}98, BR92, Bru92, CNV94, Eic94, FBR91, HB95b, LBB^{+}97, LLLB97, Smi92, SHF^{+}96, SM98c, SHKI95, TC96, TM95b, Ven93, VDA99, KHS^{+}98, PAD^{+}99]. centric [Gra91]. centrifugation [SAK91]. Century [LFGM96, SLC^{+}99, Fri90]. Ceratium [Nie91]. Cercopagidae [BL94b]. Ceriodaphnia [AB94]. cerium [Mof94]. cf [Ano93v]. CH [CBBC96, TM95a, dARB91]. Chad [IM90]. chaetognaths [TC93, Ano93h]. chain [PRWG97, TNG^{+}92, WT93]. chains [Arm94, FDG99, HLGA98, KBJ98, Thi94]. challenges [GG97]. chamber [AD92, CHA91]. Champ [Edm96b]. change [BMC^{+}96, BM99, DHK^{+}96, DD92b, EP96, FP99, Fri96, Hec99, JRT96, KTF96, LSB92, LKSA97, LGW91, MD97, ML96, MQ96, PTJ96, PJSW96, PRS94, Sch99, SLC^{+}99, May92]. Changes [AR95, KG95, LWS^{+}96, MCB90, Rot90a, THS^{+}91, WBP99, WLCG96, vLM91, BSMK95, BM96, CKCP99, CCC^{+}96, CE96, DD92a, Edm91a, FCC99, NFGP^{+}93, FAC^{+}99, GREGG97, HGS94, HQCG92, HCP90, JD96a, JD96b, KSYK96, LV97, MC95, MR98, MD97, MBK^{+}99, PS92, SH91, SLC^{+}99, WG97, WHOL99, WB96, YK94]. changing [FL95, Hai96b, JAMH96, Sta90b]. channel [RJC^{+}97, WDGS92]. Chaoborus [BO96, Bla93, DPC90, Han91, Sve97, Ber99a, HCKW90, HT98, MHT99, Swi92, Tjo90, TvE94]. Chaoborus-released [Han91]. Chaotic [SWH92]. Chara [McC91]. character [JAS95]. Characteristics [Ano93i, Kir94b, Bal96, DHP^{+}96, FCI91, GS93, HKvS^{+}92, HS92, OZ95, OCS96, SHF^{+}96, SM99b, VSBZ^{+}94]. Characterization [AA91a, AR94, DGW92, Dod92a, IPP97, OBH^{+}91, RB91, SIO98, BH95, CPSA90, GRG99, KY97b, PBJD93, PSS^{+}96, TvE94, VBvG91]. Charcoal [Ver97]. charge [FK98]. Charles [Pse92]. Charlotte [MM91]. Chatham [Nod97]. Chattonella [AWKH98, WKK91, WKK^{+}95]. Chavez [DW95b]. chelating [HFBB99]. chelators [KNM96, MB96]. chelipes [Som97]. Chemical [CPT^{+}94, DBZF98, KK93, KKS^{+}99, LRvE94, LL96, PPP97, VVK97, WK96, ZFT94, Gra97, Han91, HJ94, Kil90, LSB92, LL91, Mac93a, Mor90, Pse90, SBB^{+}93a, SM98b, Tjo90, TvE94, TZFP^{+}94, WKB^{+}96, Arc93, Log95]. chemically [HHB^{+}92, HDP^{+}90]. chemicals [BO96]. chemistry [BKB97, CR93a, CW98, ESC^{+}93, EBH92, EP96, FCM^{+}97, LMH98, MZ90, RB97, SO91, UG98, WB96]. chemoautotrophic [Hay96, RC95]. chemocline [OBKH96]. Chemoreception [LLZ99, MFY99]. chemosynthetic [BKB97]. chemotaxis [ZFdSY96]. chemotaxonomic [WGSP91]. Cheng [Won91]. Chernobyl [CDP96, HRD^{+}95]. Chesapeake [GGFR91, KSC^{+}90, MD90, MPG95, PWR94, RGRW93, SD94]. Chile [TC96]. chitin [ER95]. chitin-degrading [ER95]. chitobiase [ER95]. Chl [BPCL97, DW95a]. Chl-a-specific [BPCL97]. Chlamydomonas [SH98b, WG97]. Chlorella [WWT^{+}93b]. chloride [TM95a]. chlorinated [SM99a]. chloroform [NML95]. Chlorophyceae [GBC95]. Chlorophyll [DW93, FM94, PF91, BTL^{+}95, BF97b, CKCP99, CBB^{+}95, CGVL95, DW91, GDG99, GW93, JCMW91, KHK^{+}98, LM98b, MEM99, MUGK92, MBHT94, NS91, PASC95, RLL93, RYC99, TKAD95, Wel94, WPBS99]. chlorophyll-a [BTL^{+}95, CBB^{+}95, GDG99, GW93, MUGK92, MBHT94, RLL93, TKAD95, Wel94]. chlorophyll-b [Wel94]. chlorophyll-labeling [GW93]. chlorophyll/nutrient [WPBS99]. Chlorophyta [Hof90]. chlorophyte [DO98]. chloropigments [SLA93]. CHN [YK95]. choice [HM99]. chromatographic [KMWK92]. chromatography [ECA99]. Chromium [PM90b, JSL92]. chromophoric [GB94, HVB93, VGB94]. chronically [CACM98]. Chróst [Dua95, PA95b]. Chrysophyceae [CPS90]. chrysophyte [GHF^{+}97, Leh95a]. chrysophytes [Siv93]. Chukchi [GCD90]. Ciliata [RERB91]. ciliate [CDPR98, DC91b, DŠ97, EMH95, FPAG^{+}98, LPPSR97, MM97a, OS91, SJPGH97, WBG94]. ciliated [Wic95]. Ciliategrazing [ŠBM^{+}95]. ciliates [CDPR98, CJD^{+}99, Ver91, WLY93, NK94]. cinematographic [TTS93].

Circulation [Pin99, JB91a, JB91b, Kit96, LWS^{+}96, TS97, TCB^{+}98, WBRL98]. circulations [LJBTB99]. Cisó [GGA91]. citrate [Ano93m]. Cl [TM95a]. clade [PPK^{+}99]. clade-specific [PPK^{+}99]. Cladocera [BL94b, GB93, ML90]. Cladoceran [GU94, AB94, BG94a, HD95, LB91, SY95, TW93]. cladocerans [Bre91, Bre93, KK91, UW92]. Cladophora [Hof90]. Clair [HL92]. clam [Ano93u, EG91b, FRY95, FRLS90, FFJ92]. clams [BKB97]. Clare [HM99]. Clarification [PPK95, MFLR96]. clarity [CM94, FHKC96, LCR99, MT94]. classes [AB96]. classic [Wic95]. classifying [IPP97]. clausi [TJK^{+}97]. Clay [LDV^{+}92, CGB90, CH92b]. clean [HGFC96]. clear [LVM^{+}90]. clear-water [LVM^{+}90]. clearance [WH93a]. Climate [DMA^{+}96, MQ96, PJSW96, BMC^{+}96, CE96, DHK^{+}96, ES96, EP96, HLQ^{+}99, JRT96, KTF96, LWS^{+}96, Mac93b, ML96, Nic98, PTJ96, RKA96, SH91, Sch99, VWF96, WLCG96, Mat94]. Climatic [SHKP96, FHKC96, Fri96, HGS94, LKSA97, MF99, PSH^{+}96, SBP^{+}96]. Climatically [ALB90]. climatology [RSZ97]. Climax [Ven97]. Clinocottus [AMP98]. cloned [RKG97]. clones [DW99, WH93b, WS96]. closed [LFGM96, TGSB96]. closed-basin [LFGM96]. closely [Rot90b]. cloudless [GC90]. Clupeidae [KDKS90]. clustering [BSC93]. Clutch [GL94]. Clutch-size [GL94]. cnidarians [TKAD95]. CO [NJLW97, ASS^{+}99, BFZED94, BZR99, CP97b, DHWW97, FCG94, FBW96, HAAM99, JRT96, LBP97, McC91, NKT^{+}93, ND91, PB91, SO91, SHK95, WK96, dBGEK97, AGZ99]. coagulation [JL92]. coarse [SSE90]. coast [GA96, HGP97, JD96a, JD96b, VTCB96]. Coastal [Mal98, Pae97, AK97, AMM97, AA91a, AA91b, AK92, BL94a, BL97, CBBC96, CDP96, Clo98, CH92a, DW95a, DWM90, Eli98, FZD93, FSS94, FGN93, FAC^{+}99, FN95, GPMF96, HKvS^{+}92, HK95, HS92, HSM^{+}95, HFBB99, JLMV98, JB94, JMA^{+}95, JRT96, KOWH99, KSH96, KHRS98, Kit96, Kli92, KJ94, LPJ92, Leb90, LD96, LDC99, LO97, MD90, MB98a, Mat93, MVM97, MKWD97, MS99, MM97b, MH94a, MR99, MH94b, MB99, NK94, PP97, PHX92, RKG97, SM92, SdBdJL98, SK90, SH90, SH96b, SH98b, SH98c, SGSA92, SK95, SIO98, TN94, TM95b, TNG^{+}92, Véz94, VRHK99, Vos92, Was90, WBRL98, YFF^{+}98, ZWH^{+}95, dGGV^{+}96]. Cobalt [SH95b, LF93, YM96]. cobble [NSBB98]. coccolith [BKH^{+}96]. coccolithophore [BHAV91, BKT96, BKH^{+}96]. Coccolithophores [Paa95]. coccolithophorids [And99, Boo94]. coccoliths [VBK98]. Cod [CT96]. coefficient [AMLH97, BTLC93, BSMK98, HVB93, McC95, Vos92]. coefficients [ABB^{+}95, BS90b, CP99, GCP90, HCS95, Roe98, SKR^{+}91, VHS^{+}95]. coexistence [SH97]. coexisting [Two94]. Cohen [Hai96a]. coherence [WMC^{+}96]. coherent [Ken93]. cohesive [PWM^{+}98, WWW98]. cohort [PK94]. cohorts [McL97]. cold [BKB97, BDR99, PY99, RKA96]. cold-seep [PY99]. Cole [Duc94, Joh91]. Colimitation [PM91, Arm99]. colleagues [Hai90]. collect [Kap94]. collected [Alt90, RS91, Str90]. collection [HGFC96, LHWZ92, Liv90, Set92, SRG98]. collections [Coa90]. colloid [WJRB97]. colloidal [SGWS96, Tra94]. colloidally [CBS99]. colloids [GG97, HP95, MDC95, RM90, RS99, WG94, YSW90]. colonial [AP92, SJPGH97]. colonies [BC90]. Colonization [TW93, TK98, SGP92]. colonized [RPRNR94]. colony [Grü97, LRvE94, PSEJ99, PSJ99]. colony-forming [PSEJ99, PSJ99]. color [CdG92, GSG94, HPJW94, KE91, RCCdG99]. coloradensis [BW99a]. Colorado [ML92]. colored [DSV^{+}99]. colorimetric [MS93b, SMC98]. colour [San99]. column [BMP92a, BMP92b, Ben94, CL91, ESC^{+}93, HB95b, KYM97, MAG95, Mol98, SJDJ94, SBB99, WM92, WL94, WGS97]. columns [HTSDZ99]. combined [BW94a, KK93, NK91, RS93, VBvG91]. Combining [KRW99]. combustion [SBB^{+}93a]. Comment [Bre93, CGSD93, ADC92, CG93, CK92a, CKCP99, CP95b, Gib92, GLH91, Hes93, HM99, HR90, HSM^{+}95, JD96b, Jør93, Kie90, Kim90, Kir94a, Lap99, MBFG90, MSH96b, NT93a, NKMO91, Pol90, Pra90, Rot91, Sau92, SRG00, Som90, Sta91, TMB96, UW93, WMTB93b, Wet91b]. commentary [Sma97b]. comments [CA93]. commercial [GDG99]. common [DW91, MSSM96, SVF91, WT94]. communities [ADC92, AB92, BA97, BH91, BdBNvB91, CBFC98, CG91, DCM90, DY99, EP96, FS99, GdGD97, HLL98, HB95b, HT96, KJSJ98, KJ92, KJ94, LV98, Leh91, MM97a, MHZ98, MBKG95, PBJD93, PZ93, RFHG99, Sam91, Smi92, SSP^{+}94, SG93, SFC90, TW93, VBAO95, VCT^{+}99, WT93, WWT93a, WSM^{+}96, SW93]. Community [HB91, AS94, BKB97, BA95, CNV94, CWW98, CMT93, CRKS93, CL96, CGB90, De 90, DC91a, DAC90, DAC92, Gae93, GJ96, Goe98, GW95, GGJ97, Han96, KB92, KKK^{+}93, LBH^{+}93, LVFD93, MT90, MCB90, MDWS97, MHZ98, MDdM^{+}99, PMF90, PG97, ŠBM^{+}95, ŠKN^{+}99, SGB^{+}97, TZGS98, Ura90, WBG94, WH91, WHOL99, dGGV^{+}96]. Comparative [CBS99, Two94, RGH94, Joh91]. compare [DWM90]. compared [CP94, SSS93]. Comparing [HY97]. Comparison [Ban90a, BBH^{+}92, CH91, HCSE97, Kap92, Mac93a, NT94, Ven97, WH93a, AKM95, Ami90, AD92, BS90b, BZR99, DW95a, GNL97, Lap95, RA98, RKST90, SWS93, WPBS99]. Comparisons [Lew90, CFT92, FSS94, MBKG95, Sti92]. compartment [LS93]. compartmental [WVP97]. compensating [SGC^{+}97]. competing [HJ94]. Competition [SZW95, Smi91, XKS95, FSSI97, PHMH96, Pra90, Sch90, Sie98, Som89, Som92, Som93, Som94, Som98, Ura90]. competitive [GKM94]. complete [MVM98]. Complex [Wic95, CPMW99, SH97, TH92]. Complexation [Bru92, RB97]. complexed [CBS99]. complexes [CNL90]. complexing [JL94, LVvdB99, SDB97]. complications [EF95]. component [Coa90]. components [MH90]. composition [AV94a, Alt90, Ano93u, BGC94, BGN98, BDY^{+}98, Bra98, CAS93, DHB93, GREGG97, GCD90, GMR99, HQCG92, HSFR98, HS98, HC97, KHRS98, Kil90, LV98, MS93a, MBK^{+}99, MHZ98, PK91, Pic91, PRT^{+}93, SY95, Sha94, ŠKN^{+}99, SR92, SPMW97, Ven97, VVK97, WF96, WMD97, WBM93]. compositions [BUW96, CFW97, QWR^{+}92]. compounds [BDR99, CFW97, CLP^{+}92, CW92, JLO97, Jun92, KZM90, MZ97, vEJ97]. compression [CA91, CGSD93, Wri93]. Computations [Sti91]. Computerized [VSBZ^{+}94]. concentrated [BFG94]. concentrating [PW93]. Concentration [NVH^{+}94, WMHB95, ASH99, BKH^{+}96, BF97b, BRPR98, BZR99, Dua92, FL90, FM94, GW99, GW90, GRG98, HD95, Hof90, KSYK96, LBP97, LM98b, MEM99, MDWS97, ML90, OMk93, RRK94, RG92b, Sai94, SAK91, Str97, SH95c, SFC90, WR95, XS93]. concentration-dependent [SFC90]. Concentrations [RDK96, And90, CCT93, DW93, DMF^{+}98, HHI96, Hin94, JD96a, JD96b, Jon91a, LHS90b, MM93, Mul92, Nür98, OP92, PPB91, RKW94, RM92, SSKW91, Som90, TCCA93, UGU^{+}90, VHS^{+}95, WTH98, WL94, Wym92, YAF^{+}92]. Concepts [GG97, Pla93]. conceptual [CKR93]. Conch [LWMD96]. concinnus [Gra91]. concretions [DHP^{+}96, SVC^{+}98]. condition [Edm91a]. Conditions [AWKH98, BMP92a, Cul91, GAAG96a, GAAG96b, ND91, PVL^{+}96, SZA92, THK^{+}99]. conductivity [MCCD92, Siv93]. conductivity- [MCCD92]. Confluence [NCD98]. confocal [HC97]. conjunction [BM92b]. Conkling [CT96]. Connecticut [SLC^{+}99]. Consequences [CF96, EM93, FP99, MR98, Mor97, ADR93, BL97, CP97a, FHF99, FC91, JRM98, Mac93b, TS97, VMH^{+}97]. Consequnces [Joh90]. conservation [But95]. conservative [SK96, TMB96]. considerations [CP97b, LPA91, PZ93]. Constance [GS93, RS93, SC93a, SR92, Sti92, VJM96]. Constancy [SDJK98]. constant [DA92]. constituents [CH96, GHT^{+}96]. Constraints [FPZFW99, Ham94, Kam92, LBB^{+}97, MFY99]. consumer [BF97a]. consumers [LAC^{+}96b, PPP97, Smi91]. consuming [SBGCC91]. Consumption [CDPR98, BL94b, EKT99, HPM96, KG95, RHH97, SGB^{+}97]. contamination [CFP97]. content [AH91, DB93, GLT96, GH95, HQCG92, LK94a, MMSKH94, SCFB98, THH98, TSH99, VRT^{+}92]. contents [DŠ99]. continental [AZS^{+}98, BLDP96, FFNS99, HLL98, MJC^{+}98, MPS^{+}91, RKG97, Sam91, SMV98, TC96, WS90]. continental-shelf [HLL98]. continents [Kim92b]. Continuous [SHK95, HD97, SZW95, SGC^{+}97, HWP95]. continuously [Pol90]. Contrasting [ORH99, Sar92, CHC97, EG91a]. Contrasts [MHZ98, Har96]. contribute [NHJ91]. contributing [BMAW97, SGP96]. Contribution [BN95, Nie91, RDK96, UNK95, WMD90, BdBC99, Con97, SCK93, VFAA98]. contributions [FBS99, WTH98]. contributor [LPPSR97]. Control [Ive90, AMM97, ADC92, AC92, Ano93z, BFKL91, Bra91, BE94, Dad93, DO97, EG91b, HCKW90, KKK^{+}93, Lap97, Lap99, MC92, NS91, PPK95, PS92, WTH92, WW98, dGGV^{+}96]. controlled [ALB90, PBL99a, PAM94]. controlling [Kil90, PK94, REL97, SH98b, SH98c, TZF96, UGU^{+}90, YW93]. Controls [KNM96, VMH^{+}97, BG97a, DHB93, Leh91, MHSP90, PWR94, WBM93, vEJ97]. controversies [Jør93, WMTB93a]. convection [MJ94, WFG99]. Convective [MIM90, BSO92, JB91a, JB91b]. conventional [TJ98]. conversion [Bel90, CGVL95, KWF90]. cool [LG99]. cool-water [LG99]. coordination [HM90]. Copepod [HWF95, RG97, ZCGP98, BPA96, DA92, DZC91, DM90, DDC92, DD92b, DD92a, EMH95, GU94, GH95, GMR99, HWP95, IP93, IMB^{+}99, KBH98, LHS90a, MR92b, McK96, MEM99, McL97, MMPS98, OAR96, PD90b, PK94, PWR94, SB90, Sve97, TJK^{+}97, TB96, VGP^{+}92, VRHK99, WF98, WH93b, YSSO91, Ano93e]. Copepoda [TJK^{+}97, vLM91]. copepodite [Har96]. copepods [AE91, BOFW93, GL94, HPJW94, Hay95, KK91, KHP91, MFY99, OW96, PF90, RV99, Rom91, SCFB98, THH98, TSH99, Two94, WRLF96, Wic95, vdBG94, TN94, Ohm96]. Copper [LVvdB99, Coa91, KLR94, SDB97, SH95c, SH98c, XKS95]. copper-complexing [SDB97]. Copper-induced [LVvdB99]. Coral [Paw98, Ant99, AB92, BA97, BA92, BA95, CHA91, CW93a, CG91, DY99, FYG98, FPAG^{+}98, GJ90, Gle93, HDR^{+}98, HSS^{+}99, HC99, JLO97, KSYK96, Lap97, Lap99, LL91, LJBTB99, MT99, MHM^{+}99, PLP99, Sam91, SRLRL^{+}99, TA97, YPF^{+}98, YKM95, Sal92a]. coral-reef [BA95]. corallina [McC91]. corals [FGB95, GKM94, PG97, PSO91, RCG98, VDG^{+}97, WWD90]. Corbicula [FRLS90, FFJ92, YK93]. Core [CA91, CGSD93, Wri93, ALB90, BPS94, CC91, Liv90, NCRS90]. core-to-core [ALB90]. corer [BR91, RH93]. cores [Ege95, Jön91b, NWS94, SNB92, Ver93]. coring [CA91, CGSD93, Liv90]. Correction [Ano96b, CW93b]. Corrections [JF99]. correlation [ALB90]. correlations [Ber97]. correspondence [Ker95]. Coryphaenoides [PRS94]. Coscinodiscus [Gra91]. cosine [Ban92, BSMK95, BSMK98, McC95]. cosines [Kir91]. cosmogenic [LBL91]. cosmopolitan [WSB98]. Cost [DL92a, Gla91, Grü97, SC93b]. Costa [PRT^{+}93]. costatum [NS93, PVL^{+}96, SZA92]. costs [DL92b]. Couette [LCG94]. Council [Ano90h]. counter [SJHS98]. countervailing [MC96]. counting [SSS93, XB97]. countries [McC97]. counts [HS95]. coupled [CR96, KJ92]. Coupling [Ker93, SKB94, CPLD99, CW95, FPL^{+}91, FG93, GV93, GGFR91, GJ96, JW92, MB98a, YK94]. couplings [CFSW91]. course [RLL93]. coustal [Mol98]. Cove [GG90, TGG94]. cover [Eic94, ESG91, FCC99, FP99, Nic98]. covered [Gli98, HHJ98, LP94, MBT^{+}98, MA96, RLH95, TCB^{+}98, VRL^{+}98]. covers [DMA^{+}96]. Cr [LL98]. Crassostrea [CNL90, TZFP^{+}94]. Crater [CC97, DCML96, MCCD92, MCD^{+}96]. Craven [CP95b]. creek [SD95]. Crescenti [Ohm96]. crevice [Ber99b]. Crevices [Ber99b]. Critical [HvOW99, NS91, SP90b]. critical-depth [SP90b]. critique [BCBC93]. croaker [TJC97]. crops [MMC94]. Cross [FBR91, MB90b, YSW90, LF91, Sam91]. cross-continental [Sam91]. Cross-flow [YSW90]. cross-hybridization [LF91]. Cross-shelf [FBR91]. cruise [MK95]. Crusius [CGSD93]. Crustacea [CBS99, ER95, Som97]. crustacean [AMF98, Dod92b, MDE97, MCC^{+}96, SP90a, SZW95, VM94, VAH99, WT93, WC91]. crustaceans [ASC94]. cryptic [TH92]. cryptophyte [LLZ99]. crystallinus [Swi92]. Cs [CDP96, HRD^{+}95]. CTD [Old94]. Cu [HTDM96, MB96, MBCB97, RHT92, SH98b, XS93]. cue [TZFP^{+}94, ZFT94]. cues [DBZF98, LL96, Tjo90]. Culex [BK90]. culture [Ban90e, CP99, LTPB98]. cultured [CWSX96, MC99, SMB^{+}97]. cultures [CTF98, GW93, HS95, LVvdB99, LDC99, LB91, MWB^{+}98, MdlC97, OJ98, RLL93, SZW95, SGC^{+}97, SM98a, Som95, TM95a]. Cumming [CA93]. Cummins [Wal96]. cupric [KLB^{+}99, XS93]. Current [Gli98, GG97, HSSA94, Jør93, PB93, SG93, WMTB93a]. current-activated [HSSA94]. currents [CJVD97, JCH92, MBT^{+}98, MKT^{+}90, Ver94, Won91, YSSO91]. curvilineata [SGA98]. Cushing [Wal96]. cyanobacteria [AP92, CLMH93, Cur94, DZC91, Epp96, HR90, IM98, Li94, ME98a, MHSP90, MB96, PFB96, SWH92, Sel97, SCKF99, SCF99, WT94, Wym92, vEJ97]. cyanobacterial [KKM^{+}96, MBCB97, SLSLBD97]. cyanobacterium [Gil96, MC99, PPP97, PBL99b, VC90, WPB94, WMT98]. cycle [Ano93r, Ano93u, BN90, BGN98, CL96, ER95, FGN93, JC95, JUSI91, JSL92, Jør90, KV99, Lon91, LAC^{+}96a, LAC^{+}96b, OBKH96, PSH^{+}96, SGSA92, TOKA97, VGP^{+}92, Thi94]. cycles [BDY^{+}98, Els98, MB98a, SH90, SH95c, Sut93b, SRPG92]. Cycling [GSC^{+}96, Hon97, BMP92a, BMP92b, BE94, BAO97, BZ98, CK92b, GCCJ97, GCB^{+}98, HTDM96, HAL90, JMA^{+}95, JEP^{+}99, Kan93, Kim92b, Kir99, KSH^{+}99, KFW94, KS93, LBLO91, LD96, LvGLC98, MDC95, MFH^{+}93, PL91, PBL99a, QWR^{+}92, RHT92, Sau92, SdBdJL98, SW96b, SB97, Smi92, SM90, SFC90, UNK95, VORJ93, WGS97, ZFdSY96]. cyclopoid [GL94, Wic95]. Cyclosalpa [MP92a]. Cylindrical [YGB91]. Cymodocea [MCED94]. cysts [SAK91]. cytometer [CBFC98, DF94, FBCS90]. Cytometric [Li97, Li94, dGBPP96]. cytometrically [UC98]. cytometry [BC90, Li90, OZ95].

D [Alt94, Ano93g, Cur94, Leh95a, Man95, Mei90, Pse92, RW95, Sut93b, Tho90, Ver90, Wet91d]. D.C [RJC^{+}97]. d18 [LKSA97]. Dabob [BN95, DFN97, WM92]. daily [McB92, WHM97]. Dakota [Fri90]. damicornis [JLO97]. dammed [Ano93i]. Danish [Lew98]. Daphnia [ADR93, ASH99, BH97, Bla93, BV95, ČH99, DL92a, DL92b, DB93, DW99, DBS99, DZC91, DGS98, DG99, Den97, DC91a, Ebe92, Epp96, Gla91, GL94, GD96, HPB^{+}99, Han91, HD95, HSI95, HH93, Jun92, KM95, LRvE94, Lam94, LB96, LCR99, LL96, LW93, Loo93, LVM^{+}90, ME98a, Mac93a, MN95, NRR96, PV94, Pol95, RKW94, SP90a, Sar92, Sch90, SC93b, Som92, SH97, SHSS93, SR94, SM98b, Sti92, TH92, TYL92, UCS97, VKB96, WG97, WH93b, YW93]. DAPI [SSS93]. Dark [LID93, CMPD92, PSO91, PSS91]. darkness [HD97, MC97]. Data [Kil90, Ano95e, AS94, DC91b, LTPB98, MMTJ94, TJ98, Ben94]. database [SM97]. date [DK93a]. dated [NWS94]. dates [WMC^{+}96]. day [BG97b]. daylength [Som94, TC94]. Dead [AS91]. death [BF98, PG97]. debris [BGF92]. decades [ASS^{+}98]. decapod [MCC^{+}96]. Decapoda [CBS99]. decision [But95]. decline [KFEC91, LFHS99, NHJ91, Nie91, VL91]. Declines [JD96a, TC93]. decomposing [GC97, WH94c]. Decomposition [KLB^{+}99, AK92, FHN^{+}96, KAD95, PL91, PBL99a, Sau92]. Decrease [VKB96]. decreases [SRLRL^{+}99]. Deep [BR91, SGZ93, Ano93n, BB98, Ber91, CT93, CC97, GREGG97, HBO^{+}94, HKP^{+}97, JM98, KCWM96, LP95, MGWB97, ML95, MB90b, OC97, PG97, RBMH91, SGP92, SGP96, SHM96, Str95, TCB^{+}98, VF94, Whe92, dABL91, Thi94]. deep-mixing [CC97]. Deep-sea [BR91, Ano93n, Ber91, CT93, GREGG97, RBMH91, SGP92, SGP96, SHM96, Str95, VF94, dABL91, Thi94]. deep-water [HKP^{+}97, KCWM96, MB90b, PG97]. defended [HDP^{+}90]. defense [Grü97]. defenses [Dod99, PPP97]. deficiency [CP97a, DMM95, IPPD92]. deficient [DGS98, VLHL97]. defined [RM90]. defining [IPP97]. definitions [GG97]. degradable [THR97]. Degradation [BGF92, MR92b, AR94, DM98, HB95a, KW98a, KAB92, LC90b, MM97b, MH97, PF91, SLA93, Tay95, THK^{+}99]. degrading [ER95]. Degree [KK91]. del [KSH^{+}99]. Delaware [HK95]. delicate [ADS94]. delivery [LWMD96]. delta [PSS^{+}96, LMH98, ML96]. DeMelo [CK92a]. demersal [Lev96, MAN91a, PRS94]. demise [Ban90e]. Demographic [Har96, LGW91]. demography [Two94]. demonstration [CCK98]. Dendraster [App94]. Denitrification [CNSR90, KSCG98, NCRS90, EW99, KSC^{+}90, MSH96a, RPRNR94, RPJ97, RMT96, RRPP^{+}94]. Denmark [JMA^{+}95, LPJ92]. Denny [Men90]. dense [Fra97b, OBH^{+}91, WMML98]. densities [MAN91a]. Density [AHKH^{+}96, FR91, LJBTB99, Ano93q, Dod92a, FPAG^{+}98, MAG95, SAK91, SG93, Tra97]. Density-dependent [FR91]. Density-driven [AHKH^{+}96]. density-gradient [SAK91]. Density-induced [LJBTB99]. dentrification [BR92]. Deoxygenation [HBO^{+}94, WBRJ91]. Dependence [Hof90, MKD90, PF91, FK98, JGRR99, Str97]. Dependent [GGFR91, Ano93y, FGB95, FR91, FVL98, FVL99, HC94, MCCD92, NTH^{+}91, PPB91, Rot90b, SRLRL^{+}99, SFC90, Whe92]. depletion [FYG98, SO91]. depletion/dilution [EF95]. depletions [HG99]. deposit [BM99, FL90, Hen98, KM94, MT90, WSB98]. deposit-feeding [BM99, FL90, Hen98, WSB98]. deposited [TG92]. deposition [DT91, GJ96, Pae97, PP97, Pse92, Sta90a]. deposits [WBG^{+}91]. Depth [DW99, Bou98, FNWK96, HvOW99, MMG94b, SH96a, SLD97, SP90b, TC93, VWF96, Ver99]. depth-related [TC93]. depths [FHKC96, NS91, Smi92, SKB94, SGB^{+}97, WS90]. derivative [CIM92]. derived [AV94a, AV94b, BF97b, BdBC99, HWP95, MV98, MPS^{+}91, Roe98, SLT99, WMC^{+}96, ZKM93]. deriving [Bel90]. Description [PPK^{+}96]. descriptive [KRW99]. desert [MAS91, RB94]. desiccation [MB98b]. desiccation/oxidation [MB98b]. Design [Boy94, HN94, WLD91, DDG94, Liv90, Mil95]. designed [KL90, MMG94a]. despite [SRLRL^{+}99]. destratification [HLGA98, SF95]. destruction [MEM99]. detached [VBK98]. detect [GDG99, KHK97]. detected [DC99, HDMB97, SH91]. Detection [MSK^{+}97, SMB^{+}97, Ano93o, CCDL97, CPSA90, GK97, MS93b, SMC98]. detector [KJ92]. Determinants [WMC^{+}96, LV98]. Determination [MS93b, MUGK92, MMC94, SMC98, SMH94, BBH^{+}92, Car98, HHB^{+}92, Jon91a, Kap92, LMSP99, LB95, MCCD92, RS91, RNR98, RCCdG99, TLB99, WTB98, YK95, dGBPP96]. determinations [TBK98]. determine [BS90a, CDP96, MPA94, SCK93, VHS^{+}95]. determined [App94, AR95, GKS^{+}94, GK97, JB94, KLB^{+}99, KRB98, KSCG98, Log93, MM93, MBOCK95, UMS^{+}99]. determines [PMF90]. Determining [DSV^{+}99, MMH^{+}95, Nür99, QWR^{+}95, VFAA98, AML93, FT91, Muc91, PV94, SCFB98, WHK^{+}93]. deterrent [MSSM96]. Detrital [BL94a, CNL90, LO97]. detritivorous [SVW^{+}97]. detritivory [KK91]. detritus [KHRS98, KAB92, SSP^{+}94, TNG^{+}92]. developing [McC97]. Development [CI94, HC97, KHK^{+}98, LD99, Rut92, Ano93n, CR93a, CKR93, Dal90, Eli98, FSS94, HB91, HvOW99, Lew98, McK96, Mol98, RLL93, TB96, vdBG94, Leh95a]. developmental [Rom91]. developments [MaB98]. device [ADS94, Coa90, HTSDZ99, KC92, WO98]. Devils [Fri90]. DGT [HTSDZ99]. DGT-sediment [HTSDZ99]. di-methylsulfide [Ano93s]. Diagenesis [BFS91, BCM92, MR92a, NWS94, OBL^{+}91]. Diagenetic [GPMF96, DMHH99, DBT91]. Diagnostic [WBKF95]. diapausing [LHS90a]. Diaptomus [LHS90a, vLM91]. diapycnal [TJ98]. Diatom [NB97, And90, Ano93s, BFG94, BN95, CH96, Cra95, DDC92, DD92a, Fri90, FHN^{+}96, Gen91, GJ96, Gra91, GGF91, GGJ97, HJ94, HWF95, JW92, KBJ98, KTMI^{+}98, LBP97, LRM95, Mol98, MC97, NHJ91, NZHB95, RSK97, RM90, SLD97, SR93, SLA93, SH96b, SH98c, SGA98, TBM90, TC94, TK98, WHN^{+}98, YM96]. diatom-based [Fri90]. diatoms [Ber99b, GB93, GBC95, GG90, KTF96, KHP91, LPA94, LG99, MV96, PRWG97, SVF91, SZA92, SU98, SWS93, WTH92]. diazotrophic [SCKF99, SCF99, VC90]. diazotrophs [Cur94]. dibromomethane [GNL97]. dictionary [Kim92a]. did [VRHK99]. Diel [CHD91, DŠ99, DO98, Gra91, KM95, SR93, SH90, SRPG92, VC90, WKK91, YK94, PAS90, Ano93u, DL92a, Dod90, FH92, FR99, HPJW94, Hay95, HWP95, KY97a, KSH^{+}99, Loo93, RF94c, WSMS96]. dielvertical [DL92b]. Diet [SY95, BM99, Bra98, PPS^{+}97, Som98]. dietary [RBMH91]. diets [CLP^{+}92, DGS98, SR94, Wil92]. differ [BPK94]. Differences [Rot90b, BC97, Hay96, HPJW94, MC99]. Different [WJRB97, AB96, GLT96, HvOW99, NT94, NS93, ND91, RG92b, Rut92, ŠKN^{+}99]. Differential [CBMC^{+}99, Gle93, KSSP91, LB91, PH92a, WBG94]. Differentiation [Fra95, BG94a, ČH99]. differing [Eck90, SSM^{+}95, WMD97]. differs [BG94b]. Diffuse [MMG94b, BSMK98, GCP90, McC95]. diffusion [BH96, CSPG94, NSZ97, SKR^{+}91, UBBS94]. Diffusional [Ano93j, Dad93]. Diffusive [PSEJ99, PSJ99, AS91, GGRJ94, JD90, PAD^{+}99, RC98]. diffusivity [RNR98]. Digestion [CH96, DL96, DC91b, DŠ98, TLB99]. digestive [Bre97, MR92b, PF91, Wil92]. Digging [SC90]. digital [AE91]. Dillon [ML92]. dilute [EBH92]. dilution [SR92]. dilution-gradient [EF95].

Dimension [BGC99]. dimensional [CR96, SO91]. dimensions [GRG98]. dimethyl [WS96]. dimethylsulfide [BKR^{+}90, MWB^{+}98, ZFdSY96]. dimethylsulfonium [BKR^{+}90]. Dimethylsulphide [Mat94]. dimictic [PSH^{+}96, SFW^{+}95, UNN^{+}98a]. dinoflagellate [BFZED94, BFE96, LR99, MSK^{+}97, SAK91, SW96a, SB93, TOKA97, TK98]. dinoflagellates [BG97a, IMB^{+}99, JPJ97, KY97a, LGI94, Les96, SRLRL^{+}99]. diode [KJ92]. Dioithona [BPA96]. dioxide [CW98, CPMW99, Car98, CC98, FCG94, GJ90, PL95, SM98c]. Diploptene [PHX92]. Direct [CL99, HS95, JPG90, Pol95, Sch90, VL98, AC92, Ano93o, CI94, KLB^{+}99, LPA96, RA98, SSS93, YK95]. disaggregation [AGGD90]. discern [KHB96]. discharge [CBBC96, CE96, CCB^{+}99, Guy91]. discharged [Fie97]. discharges [WCW^{+}91]. discontinuities [MAG95]. Discrepancies [TB91]. Discrete [TIM96, NJLW97, Sie98]. discrimination [KK91]. Discussion [BW91]. disequilibria [FGN93, HP95]. Disequilibrium [BPS94]. disk [KE91]. Dispersal [BW99a, CTP^{+}95, CJVD97, LHM^{+}93]. dispersion [LAYE95]. disproportionating [CTF98]. Disruption [BMAW97]. dissipation [YSSO91]. dissolution [BFG94, GAA93, GAA98]. Dissolved [ASMB98, GK97, HHJ98, HCLD95, KK93, KRB98, MM93, MR99, MH94b, RM92, SSKW91, VV98, VDA99, WMPO99, AV94a, AV94b, Ami90, AB96, BP91, BB98, BTM92, BLS^{+}95, BB97, BS99, BZ98, CL91, CCKP98, CKCP99, CTSB98, CCC^{+}96, CE96, CH90, CW92, Cur93, DSV^{+}99, ECA99, FSGG93, FRY95, GLT96, GB94, GSW95, GSC^{+}96, GW94, HTDM96, HVB93, HT96, Kap92, Kap94, KT92, KY97b, KHBL98, KSYK96, Kro93, LS90, LFHS99, LM91, LGT95, LB95, MS93b, MCD^{+}96, MM97b, MMH^{+}95, MH90, MPS^{+}91, MZ97, MZW^{+}95, MH97, NK91, NZHB95, ORH99, OB98, OBA99, PBJD93, QWR^{+}95, RRK94, RCCdG99, RS93, SML96, SS99b, SBB^{+}93a, Sha94, SMC98, SBB99, SRG98, SRG00, SHM95, SFW^{+}95, SHF^{+}96, SBZD97, SPMW97, SIO98, TIM96, Tay95, UNN^{+}98a, UNN^{+}98b, VHH96, VL98, VGB94, VVK97]. dissolved [VM94, WHB95, ZWH^{+}95, dGP94]. distinct [Den97, HSL^{+}97, RWH^{+}98]. Distinguished [Wet93]. distinguishing [LHS90a]. Distribution [Ano93k, LS90, MBOCK95, MPS^{+}91, SHKI95, AGZ99, AE91, BKT96, BKB97, BLD99, CJVD97, DD92a, Gae92, GdGD97, GVK93, HVHH91, HS92, HW94, HCP90, JF96, JF99, KJS98, KJ94, LJH91, MMS93, MGN92, MLBC94, MTM^{+}98, MCD^{+}96, MBCB97, MNAH96, PB94, PL99, Paw98, Pin99, SM92, SA94, SSM^{+}95, SK95, TSV^{+}91, TZGS98, VV98, VAH99, Web90, WH94a, WG94, WHM97, WH93b, WBRL98, YPF^{+}98, YFF^{+}98]. distributional [KM95]. distributions [BOFW93, CPCS93, DOC99, Gae93, GCJ97, Lei90, LJD95, Log93, MKWD97, MA96, OCZA90, Sie98, VdHH^{+}98]. District [NTH^{+}91]. disturbance [FS99, KW93, MPT99, Som95, TSD97]. Diurnal [CNSR90, RPRNR94, RS98]. divergent [BH97, MW98b]. Diverse [FMK^{+}99, PPP97, BFM^{+}97]. diversions [JRM98]. Diversity [FS99, But95, DFA93, HB95b, Li97, RKG97, SGP92, SGP96, UC98]. divinylchlorophyll [RR92]. divinylchlorophyll-a [RR92]. division [Ano92d, Ban91, CCC^{+}96, OJ98]. DMS [BKR^{+}90, LD96, WS96]. DMSP [BKR^{+}90, LD96]. DNA [LF91, LJD95, TB91]. Do [LGI94, Paw98, SC93b]. DOC [CE98, FBS99, FHN^{+}96, GSW95, THR97, VHS^{+}95, VBD^{+}97, WSM^{+}96]. Does [Ban90d, DW93, LPA96, Lev96, WTH92, WHOL99, BPK94, HSM^{+}95, MBFG90, TMB96, Wil95]. doliolum [vEJ97]. dollar [App94]. DOM [PC97, RDK96]. DOM-synchronous [PC97]. dome [MMCL94]. dome-shaped [MMCL94]. dominance [GGA91, GW90, Han96, SWJO98]. dominant [FZP92, PCC^{+}97]. dominated [KV99, Kit96, SM90, Wet95]. DON [BW99b, SLT99]. Donahue [HM99]. doom [Coa90]. Doppler [FHF99]. dormancy [De 90]. dortmanna [RPJ97]. dosimeter [KL90]. double [AS91, dHJS90]. double-diffusive [AS91]. double-isotope [dHJS90]. doubled [JRT96]. down [ADC92, Lap99, Leh91, MMC94, PS92]. drainage [GBC95, GW94, NML99, RKA96, TMB96]. draining [PRT^{+}93]. Drastic [GREGG97]. drift [HLL98, Wil90]. drifter [KYM97]. driven [AHKH^{+}96, CTP^{+}95, MJ94, Ver94, AOTR97]. drivepoint [DMF^{+}98]. droplets [LGI94]. drops [BS90a]. drought [WKB^{+}96]. drying [Bal96]. dual [CBFC98, KHB96, Leh93]. dual-labeled [KHB96]. dubia [AB94]. Dubinsky [Sal92a]. Duce [Con90]. due [JD96a, RWH^{+}98, SLSLBD97, SE90, SGZ93]. Dunaliella [SGC^{+}97, SM91]. Duphnia [Pla98]. duration [ER95]. during [AB99, Ano93u, BB97, BN95, CDHS98, CPS90, CR99, CP97b, DFN97, DL92a, DL92b, DDC92, DD92a, EA99, FH92, HHR94, HLGA98, JB91a, JB91b, JM93b, JAMH96, JRM98, KTMI^{+}98, KKK^{+}93, LFHS99, LKSA97, LPA94, LvGLC98, MS93a, MR98, MTM^{+}98, MPT99, MR92b, MEM99, MKWD97, MHIHH99, Mol98, NTH^{+}91, NCD98, NZHB95, NWS94, NT93b, PVL^{+}96, RYC99, SLD97, Sch99, ŠBM^{+}95, TOKA97, UNN^{+}98a, Véz94, VM94, WHN^{+}98, WBRJ91, WWMI98, dAL94]. Dutch [DG99]. dye [PA95a]. dye-binding [PA95a]. dying [HHR94]. Dynamic [DGHD98, LHS90b, Ovi99, SHM96, GMK96, GMK98, HPTC97, Sch92, WLD93, WWT^{+}93b]. Dynamics [AOTR97, CS95, GSW95, ML95, SHM95, UNN^{+}98a, ZWH^{+}95, And97, Ano93k, BV95, CCL92, CR96, CTP^{+}95, CFT92, CMN93, De 90, DK93b, DO97, ESG91, EKT99, FCI91, FFF98, FP99, GS93, HTW90, HHJ98, JLMV98, JB91b, KMU95, KLB^{+}99, KGFP91, KFW94, LB94, LMH98, MJC^{+}98, MT90, MFW^{+}91, MHSW99, NTH^{+}91, NVH^{+}94, OMK95, PFB96, PLTL96, PVL^{+}96, PJSW96, RBH99, RGH94, SC93a, SBWB94, SM98c, TBM95, TYL92, Two94, VTCB96, VCT^{+}99, WPBS99, WH91, ZPHH98, Sal92b].

Early [BCM92, MR92a, OBL^{+}91, Ban90c, MTM^{+}98, NWS94]. earth [Hai96a, JL94, LHM^{+}93, Abb90]. earthen [SRPG92]. East [HBO^{+}94, LB94, Hec99]. eastern [BDY^{+}98, CJVD97, EK90, FPC91, GKS^{+}94, KKBG91, PRS94, SKB94, War90, ZR98]. Ebre [IPP97]. ebullition [FLH96]. echinoid [Ano93n]. echinoplutei [FSS94]. Echinus [Ano93n].

Eckelbarger [Str95]. Ecological [Ban98, Cot99, FVL98, FVL99, HCSE97, MFW^{+}91, Thr90, Wet98, Ano93l, BGC99, BL97, Fen94, HK98a, Mei90, Mil95, PTJ96, PRT^{+}93, Som98, ZFT94]. Ecology [Leh95a, Wil98, Ano93c, CWW98, Dod99, Dua95, Edm96c, FHS98, GB93, GAAG96b, LDC99, Man95, PA95b, PRCV98, Rav95, RW95, RB94, Sel97, Thr92, TTS93, UW93, VPBD97, Wet91d, BCBC93, Duc94, Man94, Dol91, Wet91a]. ecophysiological [VMH^{+}97]. ecophysiology [Kir99, Man91b, Sma97a]. Ecosystem [BMC^{+}96, Gae93, IKT99, dGF96, Arm94, CL96, FPL^{+}91, Fur91, HCPR99, JD96a, KHRS98, KTF96, MFW^{+}91, MC92, MQ96, NK94, PFDH93, PLP99, PAM94, QP91, RGH^{+}93, RGH94, SVW^{+}97, SHB99, TH96, TN94, VMH^{+}97, Wil98, WMML98, ZWH^{+}95]. Ecosystem-specific [dGF96]. Ecosystems [Sal92a, Arm94, Clo98, DW91, Fin95, HCSE97, HM99, Joh91, Kim92b, MAS91, Ovi99, Sal92b, Smi90b, Web95, Wel99a, Wil95, WMPO99, YKM95, Pse92, Wal96]. Ecotypic [MC99]. ectoenzyme [FFF98]. ectoenzymes [CK95]. ed [Amm96, Boo94, But95, CT96, Con90, Dol91, Dua95, Fen94, Leh95b, PA95b, Pse92, Sal92a, Smi90a, SS93, Wet91a, Won91, Wri90]. eddy [BH96]. edge [DCML96]. Editorial [Kir94a]. Edmonton [Ano94]. eds [Ano93h, Ano93f, Ano93g, Bur94, Cha96, Cur94, Duc94, Edm96c, Edm96b, Har95, HL94, Joh91, Leh95a, Man95, Mat94, May92, McC97, Mil95, Paa95, Pri94, Rav95, Sel94, Spi96, Str94, Str95, SW93, Sut93b, Thi94, Tho90, Thr90, Wal96, War90, Web95, Wet95, Wet91d]. education [Wet91b]. educational [Wel99a]. edulis [BFGS94, WF96]. Eel [MR99]. eelgrass [PBL99a, SBK95]. Effect [Ano93m, Ber99b, BSMK98, BGC94, BZR99, Cur93, GJ90, Gil96, GGF91, GH95, HVHH91, JL92, LBP97, ML90, NE93, PW93, RC98, THH98, TSH99, Web90, WH94a, AK97, ASH99, Ano93l, DC91a, Eic94, FL90, HPM96, JCH92, LW93, LPA96, MDdM^{+}99, OMK95, PBL99a, Pol95, PLP99, RSK97, Sha94, SAG98, SGA98, WH99, WLY93]. effective [RC93]. Effectiveness [LHWZ92]. Effects [ASC94, AB92, Bal96, BB98, BA95, BW99a, BdBNvB91, BLK95, CSPG94, CW93a, CT93, CPT^{+}94, CDPR98, Coa91, CMPD92, DA92, DZC91, DGS98, DC99, DD92b, ES96, FHKC96, FPAG^{+}98, FPC99, FDG99, FH92, GCCJ97, GKM94, GGRJ94, GAAG96a, HLQ^{+}99, Han91, HSI95, HQ90, HD97, HSRR96, HS90, JM98, JRT96, KDKS90, KE91, KAB92, MP92b, Mar94, MA96, MHM^{+}99, MB98b, MS97, MB99, MC97, RKW94, Rom91, RGRW93, Sai94, SdBdJL98, SG93, SM97, TA97, TJK^{+}97, Tjo90, Tra94, TZFP^{+}94, VL96, VVIK99, WCW^{+}91, WF96, AC92, BGC99, BK90, BSG^{+}99, BL94b, CWSX96, CH90, DHK^{+}96, DL93, DCML96, EG91a, EW99, EP96, FLH96, FR91, FM94, GCB^{+}98, GW99, Gen91, GDG99, Gle93, GL94, GAAG96b, Guy91, HSCM97, HD95, HDMB97, HM99, HHL98, JPG90]. effects [JLO97, LHWZ92, MLLR99, NSBB98, ORH99, PSO91, PCFS97, PB93, PPS^{+}97, PSH^{+}96, PCC^{+}97, STJ^{+}93, Sar92, Sar99, SBP^{+}96, Sch90, SW96b, SBK95, SH98c, VL98, WBG94, WWT93a]. Efficiencies [Leh93, BDR99, PY99, Rot90b, RKST90, Str97, WH93a]. efficiency [AZS^{+}98, BS90a, Bre91, KWF90, Kro93, LL98, NT94, OGH^{+}96, PPS^{+}97, Som98]. Egg [IP93, PK94, DDC92, GL94, Hai96b, KBH98, MLBC94, STJ^{+}93, VKB96]. eggs [BV95, CCK98, GH95, LHS90a, MLBC94]. Eh [ND91]. Eilat [LP95]. Eilers [Gib92]. Elbe [KG95, Ker93]. electrode [VBvG91]. electrodes [MBSW98]. electron [SBW^{+}98]. electronic [Kir98]. electronics [Fil91]. elegantissima [AMP98]. element [Bre93, FMH^{+}91, FCM^{+}97, HTSDZ99, HRD^{+}95, JL94, RF99, TJC^{+}98]. elemental [BGF92, CTF98, Lan93, SPMW97, UW93, WTB98]. elements [AMF98, BL94a, BSG^{+}99, CG97, MLLR99, RF94a, RF94b, WRLF96, WF96, WF98]. Elevated [Les96, MWB^{+}98, KOWH99, VKB96, dARB91]. elevation [WBM93, WLCG96]. Elimination [LB95, MGWB97]. elucidated [HP95]. embayment [MH94a]. Embryos [Ano93n, Gla91]. Emeis [Cha96]. Emerald [WB96]. emergence [De 90, MAN91a]. emergent [BSO92, NSZ97]. Emiliania [CP97b, LTPB98, LVvdB99, PH97, TC95a, VBK98, WS96]. emission [DGW92, MSM96, ND91]. emissions [KHS^{+}98, MGW96]. emphasis [HCSE97]. Empirical [Ban90b, Lou90, WP94, CGVL95]. emplacement [WFG99]. employing [LDC99]. empty [Web92]. Ems [dJvB95]. enclosure [THK^{+}99]. enclosures [BGC99]. Encounter [MK95, Ano93j, BS90c, KB92]. encrusting [Grü97, MB90b]. Encyclopedia [Fee99a]. endless [Ven93]. endogenous [Loo93]. Endomycorrhizae [WAC^{+}98]. endoscopy [Jør93, WMTB93a]. endosymbionts [KCKC90]. energy [AE91, KWF90, WTH92, YSSO91]. Engel [Bur94]. engineering [Kna99, Tho90, WMML98]. England [Sta91, GLWH90, WH94b]. English [MMC94, NTH^{+}91, Kim92a]. enhance [Thr91]. Enhanced [Joh94, LGT95, RYC99, Som98, dABL91, BC97, RF94c, XB97]. Enhancement [WWMI98, CMPD92, WK96, Web92]. enriched [Bac99, SB90, TK90]. Enrichment [Cul95, TvE94, BEB92, Ban91, BdBNvB91, CTF98, EA99, HLL98, LDC99]. enrichments [AR96, Coa91]. enteromorphoides [PPP97]. entrance [KKM^{+}96, KHK^{+}98]. entry [HLGA98]. enumeration [Ano93o]. environment [Ano93i, CW93a, CH92a, DCM90, JRH90, Lam94, Man91b, MM97b, Mof94, MH94b, NSK95, Sie98, Log95, Men90]. Environmental [DBS99, Hec99, BG97a, CCT93, FL90, Kna99, LK99, LMSP99, LGW91, MD97, Rav95, TJC97]. environments [CHD91, CPSA90, DBZF98, Duc98, GSW95, HNS95, Hai96b, HCJ96, HDB93, KL90, KT92, Ker95, MK95, Mel99, NVH^{+}94, PPB91, PZ93, Tis92, YFF^{+}98, YSW90]. Enzymatic [SFFF97, TC93, Arn98, SMH94]. enzyme [AA91a, BH95, ER95, NVH^{+}94]. enzymes [Bre97, SIO98]. enzymic [HHB^{+}92]. epibenthic [HK98b, HBI95, LSS^{+}97]. epibiont [ADR93, TW93, WWT93a]. epibionts [CMT93]. Epibiotic [WCT90]. epifauna [TR98]. epifluorescence [Ano93o, HS95, XB97]. epilimnetic [MM97a, Sha94]. epilithic [PFDH93]. epilithon [THS^{+}91]. epipelic [SU98, VV98]. epiphytes [BW90]. Epiphytic [BW90]. epiphyton [Ano93w, THS^{+}91]. Episodic [HCP90, Bur92, GREGG97, LHS90b]. epizoan [WT93]. epizoic [GB93]. EqPac [LBB^{+}97]. equation [Sch92, Swa99]. equations [Ban90b, GG92, McB92]. equatorial [BC91, CLKJ95, CBMC^{+}99, CBC^{+}91, GCJ97, GKS^{+}94, LBB^{+}97, LLLB97, OHMK91, PAM94, RDK96, RG97, RB97, WPB94]. equilibration [NJLW97]. Erie [BM96, SH95a, TCA96]. Erken [GJ96, IPPD92, Wet99]. erodibility [SAG98, SGA98]. erosion [LS93, WWW98]. Errata [Ano96c, Ano96d, JF99]. errors [BM92a]. essential [GMR99]. esterase [ASMB98]. Estimate [Ano92d, CLKJ95, DC91b, KF93, NKT^{+}93]. estimated [DD92a]. Estimates [GG99, MAN91a, BTLC93, Gor91, HY97, KE91, PH92b, VV98, WSM^{+}96]. Estimating [ASS^{+}99, BLR92, CG95, MM91, MBHT94, TJ98, TB91, FDLD93, McL97, SHB99]. Estimation [IH95, JB91a, SK90, TKL^{+}97, VBAO95, AO96, Ano93r, Ban92, OW96, PJ90, SHK95, Wym92]. estuaries [CE98, HR90, HSM^{+}95, IPP97, LvGLC98, MP92b, MB90a, PJ90, SGWS96, SHB99, VMH^{+}97]. estuarine [AEL^{+}99, AA91a, AA91b, Ano93w, BSG^{+}99, BZ98, CW98, CPMW99, CFW97, CK92b, CP99, DC91b, Ege95, HPTC97, HS90, KSCG98, LK94b, MVM97, MV98, MLLR99, MM91, MKD90, RT92, RFHG99, SDS95, SS99b, SDB97, Smi90b, SK90, SM99b, SK95, YK94]. estuarine-intertidal [CPMW99]. estuary [AC92, CCP92, ES93, FPL^{+}91, FBW96, GCP90, GGFR91, HK95, JCMW91, KMC90, KG95, Ker93, Kim92a, KBB98, KBMSP99, Leb90, LKM98, LVFD93, MFH^{+}93, PLTL96, Roc98, SS99c, SAG98, WHM94, dJvB95]. eucaryotic [FK98, Li94]. Euchaeta [YSSO91]. euglenoid [WCT90]. Euphausia [Dal90]. euphausiid [HK98a]. euphotic [BW99b, DK96, Gar91a]. Eurasian [OAF^{+}99]. Europe [KHS^{+}98]. European [BG94a, CDP96, HKvS^{+}92]. Euterpina [GMR99]. eutrophic [Bal96, CRKS93, DCM90, JA92, KSSP91, KS93, Mac93b, MGWB97, PH92a, Per97, SPB90, Sar92, ŠBM^{+}95, ŠKN^{+}99, SHM95, SBGCC91, UNN^{+}98b, VT90, XS93, XKS95]. Eutrophication [Clo98, ASS^{+}98, BC97, JD96a, KDKS90, Pae97, SH95a, SBK95, TMLS99, Wet95]. Evaluating [FHF99]. Evaluation [Boy94, BP97, HB95a, HCS95, IPPD92, KL90, MSH96a, RCCdG99, BW90, HN94, NT94, PF90, PHT^{+}93, Sch92, SBB^{+}93a]. Evans [HL94]. evasion [JM98]. Evelyn [Ano90g, Ano91e, Ano92a, Ano93b, Ano95b, Edm91b]. evening [HCKW90]. Event [GJC92, CC97, HSCM97, RG92a]. Event-scale [GJC92]. events [LHS90b, SLSLBD97, WWMI98]. Everglades [CGHK99]. Evidence [ADC90, ASMB98, ASS^{+}98, ARM96, CBD^{+}99, DG99, DMM95, FZP92, GLS90, HDB93, LFHS99, MMCL94, MLBC94, MMW^{+}97, MMW^{+}99, PPK^{+}99, PPB91, SRG98, TCB^{+}98, WMT98, YK94, Ano93r, FAE99, HBI95, IMB^{+}99, Kam92, KCC98, LSS^{+}97, RG92b, SGP96]. evoking [ZF93]. evolution [Dod99, FK98, Ano93e]. examine [CCR^{+}95, HC97, WHM94, WDGS92]. examining [Ant99, VSBZ^{+}94]. example [MS97, WMML98]. Examples [HDR^{+}98]. excentricus [App94]. exchange [AHKH^{+}96, ASS^{+}99, BH96, CH91, CC98, Con90, Fin95, GS91, JB91a, Jan93, Les95, MHT99, Ree90, SMV98, VMDC96, WK96, Web92, WHM94]. Excitation [LCG94, JPJ97, MC97]. exclusion [ECA99]. exclusive [HRF91]. excretion [CG93, Jan93, Lan93, SVW^{+}97, WP94]. Excretory [TR98]. excurrent [OMK95]. exilis [HPB^{+}99]. exist [DW93]. exogenous [GLS90, LK94b]. Exopolymer [DL93, SU98, DM90, GSL97, PA95a, WH95, ZMP98]. exopolymers [UPP95]. exotic [HY97, HH93]. expatriate [TSV^{+}91]. Expectations [Smi90a]. experienced [TJC97]. experiences [Smi90a]. experiencing [Wil92]. Experiment [Cul95, HBL^{+}93, JEP^{+}99, MB90b, PBL99a, RB97, Sam91, SB94, THK^{+}99, WLD91]. Experimental [BFG94, CKR93, HBI95, LW93, MDdM^{+}99, STJ^{+}93, SGP96, Whe92, ZR98, Ano93r, AR94, BA97, Bre93, CCKP98, EW99, FS99, FMK^{+}99, FHN^{+}96, KRW99, LPA96, LMSP99, NSBB98, NZHB95, Nür99, Roe98, Sar99, Som95, TJK^{+}97, THS^{+}91, UW92, VGB94, SBP^{+}96]. Experiments [HHL98, AR96, Ban91, BdBNvB91, BFGS94, DDG94, EEL97, HW94, LF93, PC96, Rot92, RKST90, SJL95, SBK95, SGP92, Mil95]. explain [Cul91]. explaining [SH97]. explanation [Bel90, Vod92, WSNM98]. explicit [BAH97]. exploitative [Sch90]. exploration [Wal92]. explored [Edm96a]. Export [MA94, WCW^{+}91, AW93, BG97b, CE96, KL96, Was90]. exposed [GHT^{+}96, HSL^{+}97]. exposure [Bal96, FH92]. Expression [RC95, Hay96, MCCD92, PPK^{+}99]. extensive [MMG94a]. external [Per97, SH98c]. extracellular [Arn98, MB96, NVH^{+}94, VM94, vEJ97]. extraction [Ben94, Rut92, TB91]. extrapolations [Smi90a]. Extreme [PDM96, WCW^{+}91]. extremely [OCP92]. extruder [Ver93]. exuded [SM98b]. Exuvium [TB96].

F [Arc93, Duc94, Log95, May92, Pse92, SS93, Thr90, Ver90, Web95, Wet91d, CBB^{+}95]. facilitates [SC90]. facilitation [Som92]. FACScan [DF94]. factor [Bel90, BC97, BMAW97, CGVL95, SGP96, SH97]. Factors [BP96, GNL97, Han92, ME98b, PB91, REL97, AGZ99, BKT96, WMHB95]. failed [HPB^{+}99]. fair [Ami90]. fairy [BW99a]. Falkenbergia [MHMD99]. Falkowski [Sut93b]. fall [CF96, HLGA98]. Fasham [HL94]. fast [Old94, dBGEK97]. fast-responding [dBGEK97]. fast-response [Old94]. fastest [KAD95]. Fate [KZM90, SSM^{+}95, BLDP96, BG94b, DC96b, GGJ97, HWF95, KAB92, Log95, LAC^{+}96b, PMF90, RMT96, UMS^{+}99]. fates [KHB96]. fatty [EMH95, KHS97, WG97]. fauna [MPT99]. faunal [Pal92]. Fe [GHF^{+}97, HTDM96, HTSDZ99, LBL^{+}98, OHMK91, RW96, UGU^{+}90, WHK^{+}93, YH92]. feature [DW91]. Fecal [BN95, WMML98, VRHK99]. Fechner [Log95]. fecundity [OR94, OAR96]. fed [MLR^{+}94, SWS^{+}96, Ver91]. Feedback [SH92, DMA^{+}96, FCG94, Swa99]. feeder [GW99, KM94]. feeders [Ano93j, Joh90, LPA96, MT90, NT94, TG92, Wil92, WMML98]. Feeding [JWRS96, MP92a, MSSM96, MLB99, SW96a, SB93, YSSO91, App94, BM99, CMPD92, CH94, DFN97, Dal90, DZC91, FL90, FH92, HSI95, HDP^{+}90, Hen98, Jac93, Jør93, KK91, KHB96, KDB91, MC95, MR92b, MEM99, MHIHH99, PCFS97, PF90, PPS^{+}97, Pla98, RCG98, Rot90b, SJL95, TZF96, TJK^{+}97, TTS93, VBAO95, WMTB93a, WMTB93b, WSNM98, WSB98, YK93]. Feeding-deterrent [MSSM96]. females [DFN97, Gla91, vLM91]. femorata [HQCG92]. fermentation [Ker93]. Ferredoxin [EA99]. fertile [SRPG92]. fertilization [CBMC^{+}99, FC91, IMB^{+}99, PB91, WLD91]. fertilized [PFDH93]. Fiber [KMK95, Alt90, CBB^{+}95, CW93b, Dod92a, KJ92]. Fiber-optic [KMK95, Dod92a, KJ92]. fiberoptic [KHK97]. Fibrillar [SBW^{+}98]. Field [BMT^{+}96, EEL97, GG90, MBT^{+}98, PHT^{+}93, SGP92, Wal92, Ban91, BSMK95, HCHD94, HCJ96, Kap94, Kir91, Kir94b, KWF90, KJ94, LW93, MMTJ94, RF94c, SCFB98, Som93, TV99, Ver93, dABL91]. fields [SJL95]. filament [AWL94]. filamentous [Epp96, GAAG96a, GAAG96b, HBI95]. filled [PD90b, SL98]. Filter [Bre91, LB96, Lam94, Roe98, Str90, Wil92, YK93]. filter-feeding [YK93]. filtering [Lou90, ML90]. filters [Alt90, CBB^{+}95, CW93b, Log93, RS91, TF95]. filtration [Lig92, YSW90, dHJS90]. Financial [Ano90a]. Findlay [Joh91]. Fine [HCJ96, ALB90, CMN93, GWY^{+}92, GWMB91, KW98b, Old94, WCW^{+}91]. fine-grained [KW98b]. Fine-scale [HCJ96, ALB90, Old94]. Finland [KKM^{+}96, KHK^{+}98]. finmarchicus [CKR93, CR96, MHIHH99, MMPS98, OR94, TTS93]. fiords [RGH94]. First [IMB^{+}99]. Fish [EABK99, NFGP^{+}93, SH97, Ber99a, BLD99, CRKS93, DPC90, DW99, DHK^{+}96, DC91a, ES96, EEL97, FR99, Ive90, JB94, LL96, Lev96, LJH91, LVM^{+}90, MMCL94, MK95, Mac93a, MCB90, MT95, MSP^{+}99, PL91, Per97, RKA96, RF94b, RBH99, SM92, Sau92, SHF^{+}96, SM98b, TLB99, TM95b, Tjo90, WBG^{+}91, WCT90]. fish-exuded [SM98b]. fish-produced [Mac93a]. fishes [Fee99a, Paw98]. fishless [RB94]. Fissidens [NSBB98]. Fissidentaceae [NSBB98]. fitting [GG92]. Five [FSS94, CPMW99, MS93a, McB92, RRK94]. fixation [BTL^{+}95, CMPD92, HSM^{+}95, HBL^{+}93, LID93, ME98a, MVT99, PPK95, PFB96, PPB91, Smi90b, WL92]. fixing [CPSA90, HR90, MHSP90]. fjord [Ano93z, OW96, RGH^{+}93, Mil91]. flagellate [CGD90]. flagellated [PAS90, CŠ90]. flagellates [And99, Boo94, NK91, Rot92, SP90a, WCT90]. Flathead [LS93]. flats [BA92, dJvB95]. flavodoxin [EA99]. Fleeger [Liv90]. flexibility [JD95]. Flexible [DL96, DGHD98]. floating [EM93, KYM97, MJ94]. flocculator [DDG94]. flocs [PW93]. flooded [KMU95]. flooding [EM93]. floodplain [AB94, CNP94, CP94, CV99, ES90, EM93, Les95, Two94]. floor [Dad93]. flora [DL93]. Florida [LWMD96, ADC90, BPS94, CGHK99, CCB^{+}99, DAC92, FZP92, HCS95, Lap97, MM91, MSK^{+}97, PBL99b, PLP99, WBRJ91]. flos [Gil94, Gil96, HSI95]. flos-aquae [Gil94, Gil96, HSI95]. Flow [AML93, FGB95, HZF96, Joh90, LL95, SC90, dGBPP96, BSO92, BC90, CHA91, CW93a, CBFC98, DF94, ECA99, FBS99, FBCS90, GKM94, GW99, HK98b, HA92, HCJ96, HSL^{+}97, IH95, JLO97, JM98, KBJ98, LR99, LPA96, Li90, Li94, LB95, LCG94, MS93b, MH94a, MHSW99, NSBB98, OZ95, PSO91, SMC98, SMH94, TZFP^{+}94, UC98, VFW^{+}92, YSW90]. Flow-dependent [FGB95]. Flow-induced [HZF96]. flows [FHF99, FPZFW99, MB90b, NK99, War90]. fluctuating [Ver99, VCT^{+}99, ZCGP98]. fluctuations [Fri90, JPG90, WLD93]. fluid [LCG94]. Flume [BFGS94, NSBB98, WWW98]. fluminea [FRLS90, FFJ92]. fluorescamine [CH91]. Fluorescence [JPJ97, BTL^{+}95, DOC99, FT91, FNWK96, GMM96, GDG99, GVK93, GB94, GKS^{+}94, HVB93, HV98, Jon91a, KF93, Li90, LP94, MC97, OGH^{+}96, OZ95, OCS96, PF90, PC97, SM91, Vod92, VGB94, VHS^{+}95]. Fluorescent [LJA91, LK94a]. Fluorometric [Wel94, Tra97]. flushing [Ber91, MM91, Roc98]. fluvial [CV99]. Flux [Jac93, AD92, BPK94, BH96, BOB94, BUW96, BN95, Dod92a, DCML96, HSM^{+}98, KTMI^{+}98, Mat93, NKT^{+}93, OBA99, PJ90, PPS^{+}97, RDK96, RC98, RJC^{+}97, Sar99, SE90, SCPS98, SBB^{+}93b, WHM97]. Fluxes [HA90, SDB97, AD92, BB97, Boy94, CW98, FBW96, GJ90, HKPM99, JMMD93, KRB98, MFJR99, MKBN99, Nod97, PD90b, PL95, SVF91, SHK95, Sti91, TBM90, VDA99, WVP97, WB96, YAF^{+}92]. fmax [Ano92d]. Focke [MC95]. focused [CR93b]. focusing [BK95, DCML96]. following [CCC^{+}96, GHF^{+}97, GD96, MSP^{+}99, MSM96, WGS97]. Food [LC93, ME98b, MEM99, Ste90, SB90, AML93, Arm94, BVSBB93, BKR^{+}90, BC97, BAH97, BSG^{+}99, BLK95, BFGS94, CL99, CCR^{+}95, CFP97, CFT92, DA92, DŠ99, DD92b, Ebe92, ES93, FCI91, FSS94, FPC99, FL90, FDG99, Fra95, Gil96, Har96, Hen98, HLGA98, KSH96, Lam94, LB96, MR98, MVM97, MLLR99, MO91, MSP^{+}99, MF90, ML90, PF91, PCC^{+}97, Pro90, RKW94, RV99, Rom91, STJ^{+}93, Sai94, SP90a, SWS^{+}96, SS99a, SGP92, SGP96, SHSS93, Str97, Swa99, Thi94, TNG^{+}92, Tis92, TT97b, TCA96, VF94, VL98, WF96, Wic95, WSMS96, WC91, YMT^{+}99, ZEG^{+}94, BC97, SS93]. food-feedback [Swa99]. food-independent [Ebe92]. food-limited [FSS94]. food-related [Hen98]. Food-web [LC93, BLK95, FCI91, MR98, TCA96]. foraging [NFGP^{+}93, KSH^{+}99, MJG91, MFY99, TJK^{+}97]. Foraminifera [CWSX96, GAA93, UMS^{+}99]. Foraminifera-symbiont [UMS^{+}99]. force [IH95, JAS95, SBW^{+}98]. force-flow [IH95]. forces [DG90, DGHD98, Lap99]. forcing [DMA^{+}96, Goe98, SHKP96]. Ford [Fen94]. forereef [CW93a]. forest [BGC94, Boy94, JM98, Mul92, MMW^{+}97, MMW^{+}99, WLD93]. forested [AB94]. forests [HLL98]. forgotten [MB98a]. form [AP92, FSS94, KK93, NSK95, RBMH91, RC95]. Formation [GSL97, KZM90, AOTR97, KHS97, KTMI^{+}98, LRvE94, MZ97, OMk93, PRWG97, PW93, RLTB90, Rie92, SML96, SVC^{+}98, ZMP98]. formed [KG95]. forming [BFE96, PSEJ99, PSJ99]. Forms [JMMH98, Hay96, HSL^{+}97, Rut92, SGWS96]. forward [BC90]. fossil [KFEC91, VL91]. fossils [Whe91]. Foundation [Thr91]. four [ASS^{+}98, ČH99, HB95a, JA92, RGH94, Swi92]. Fractal [LW90]. fraction [PW93]. fractionated [HNS95, Pro90, TL94, WL94]. fractionation [BDR99, CTF98, HCS95, HC94, HFK92, LBP97, Lig92, PVL^{+}96, TC94, TC95a, WHN^{+}98]. fractions [RG92b, Sha94, ŠKN^{+}99, SB97, VFAA98]. fragmentation [RKA96]. Framvaren [Mil91]. France [AEL^{+}99]. Francisco [CCR^{+}95, HCP90, LL98, LvGLC98, MHSW99, PD90a, WBKF95]. Free [GD96, KYM97, San99, XS93, BS99, CSPG94, CBS99, CP97b, DFA93, FGB95, NK91, RS93, Sai94]. Free-floating [KYM97]. free-ionic [CBS99]. Free-living [San99, DFA93]. Free-swimming [GD96, Sai94]. freeze [RH93]. freezing [MA96]. Frequency [DOC99, DD92a, JB94]. frequent [ADC90]. Fresh [New99, BLDP96, Kap94, ND91, SJHS98]. fresh-water [Kap94]. Freshwater [Lew98, MH94a, Thr91, Wel99a, AH91, BP91, BFKL91, BL98, BSO92, BFM^{+}97, CTSB98, CŠ90, CCP92, CLMH93, CCT93, CdG92, De 90, DAC90, FRLS90, GGJ97, Har96, HM99, HHL98, SJP94, JX94, Jør90, KHRS98, MW98a, MHZ98, MHSW99, NVH^{+}94, PL91, PPK^{+}96, Pic91, RW96, SP90a, San99, Sau92, SCK93, SM90, SBGBC96a, SBGBC96b, SLT99, SKR^{+}91, TCCA93, TT97a, Vod92, WL92, Web95, WVP97, Wil95, Edm96a]. Freshwaters [Wet98, CP95a, HA92, Kap92, QWR^{+}95, WJRB97]. Friedmann [Pri94]. front [Ano93v, GW90, RGR^{+}98]. frontal [KKM^{+}96]. frontier [Kir98]. frontiers [HCPR99]. fronts [Pin99]. Fryxell [HHJ98]. fueling [PH92a]. fullness [PF90]. fulvic [MAS91, MASA94, WJRB97]. function [BFGS94, Dum98, FVL98, FVL99, Fri90, Kir91, MVM98, PB93, RW95]. Functional [GW95, KM94, Hay96, PPK95, Rot90a, SR95]. functions [JF96, JF99]. Fundamentals [Hol99]. funding [Kal91]. Fundy [CT96]. fundyense [TOKA97]. fungal [BLR92]. fungilike [Ano93o]. furrows [VFW^{+}92]. Further [FAE99, Kam92, RG92b, TvE94]. future [Web95].

G [Ano92a, Ano93e, Ano93b, Ano95b, Cur94, Dol91, HL94, Joh91, Jón95, Man95, Mat94, McC97, Ohm96, Paa95, Sut93b, Thi94, Wal96, Whe91]. G. [Ano90g, Ano91e, Edm91b, Smi96]. G.Lund [Edm96a]. galeata [TH92, DW99, Mac93a, WG97]. Galilee [KKS^{+}99]. Gambusia [BK90]. Gas [GS91, PAS90, BSO92, HHJ98, RC98]. gas-vacuolate [PAS90]. Gastropod [Ano93p]. Gastropoda [Bre97]. gatunense [BFE96, ZPHH98]. GC [MMH^{+}95]. GC/C/IIRMS [MMH^{+}95]. geese [KSH^{+}99]. gel [KDZD94, MKBN99, dHJS90]. gelatinous [CG95, CT93, PWR94]. gene [PPK^{+}99]. general [GV93, HTW90, Sma97a]. generale [Edm96b]. generalizing [HM99]. generates [McC91]. generating [MOF90]. Generation [WKK^{+}95, GLWH90, Sta91]. genes [RKG97]. Genetic [BG94a, KM95, MW98b, MBKG95, UC98, MBOCK95]. genetics [Plu97]. Geneva [ML95]. genomic [LF91]. genus [LHS90a]. Geochemical [CG97, SBWB94, UGU^{+}90, WBM93, BGF92, DHP^{+}96, SH95b]. Geochemistry [YH92, AZS^{+}98, Bur94, FMH^{+}91, JL94, TGSB96]. Geographic [PSS91]. Geographically [Sam91]. geography [Ban98]. geolipid [BM96]. geometry [LW90]. George [MHZ98]. Georges [MMPS98, RYC99]. Georgia [CW98, SPMW97, WPBS99]. geothermal [PRT^{+}93]. Gerlache [MTM^{+}98]. German [MS93a]. germanium [CPT^{+}94]. Germany [TZGS98, WL99]. GF [CBB^{+}95]. GF/F [CBB^{+}95]. Ghana [TGSB96]. Giant [SCPS98, Ano93u, FRY95, HSL^{+}97, SSP^{+}94, GNL97]. gibba [GDG99]. gibber [McK96]. Gibbs [EBH92]. Giblin [Sta91]. Giller [Man95]. Gironde [AEL^{+}99]. gizzard [SVW^{+}97]. glacial [BW91, CPT^{+}94, MA96]. glaciated [Ano93i]. glandiforme [Joh94]. glass [Alt90, CBB^{+}95, CW93b]. glass-fiber [Alt90, CBB^{+}95, CW93b]. Global [MHM96, DHK^{+}96, Har95, Lon91, May92, Sch99, Thi94, But95]. globiceps [AMP98]. glomerata [Hof90]. glucose [Ano93u, LK94b, RDK96, SBB99]. glycerol [Ano93u]. Glycine [KW98a]. glycogen [HHB^{+}92]. go [VRHK99]. GOMEX [BR91]. Good [Wet91d]. Gopal [McC97, Wet91a]. Gordon [FAC^{+}99]. gorgonian [RCG98]. governing [Hay96]. gradient [BH96, CCT93, DW93, EF95, SAK91, SML96, SR95, SM99b, WAC^{+}98]. gradients [LVFD93, MBS98, Rah90, RG92b]. Graduate [Ano90h]. grained [KW98b]. Grand [PH92b]. grandis [CCT93, TCCA93]. Granite [Sta90a]. gravel [SA94]. gravid [vLM91]. gravity [CC91, CA91, CGSD93]. Gray [JD96b]. Grazer [MM97a, PAM94, BMAW97]. Grazer-controlled [PAM94]. grazers [HDMB97, LRvE94, LV97, SBZD97, TT97b, VLHL97]. Grazing [Arm94, Epp96, WS96, AW93, Ano93r, CCKP98, CKCP99, CBC^{+}91, CJD^{+}99, DŠ99, DD92b, EF95, Fro91, HBH97, Han96, HWF95, Jac93, KKK^{+}93, LBB^{+}97, LLLB97, LCR99, LKM98, MT95, MPG95, Nie91, Pet94, QLBAP96, RG97, Som97, SW91, TC95b, VT90, WW95, YPF^{+}98, dAL94]. Grazing-activated [WS96]. Great [MF99, SWS93, HLQ^{+}99, LFGM96, MQ96, Nic98, TM95b]. Greater [IKT99]. Greece [SLD97]. Green [Pri94, Gle93, GAAG96a, GAAG96b, HW94, LRvE94, MBHM99, dGBPP96]. Greenland [CSM94]. Greifen [KS93]. Greigite [Hil90]. grenadiers [PRS94]. gross [SRG98, BW99b, Str97]. grounding [Kim92a]. grounds [HC96]. Groundwater [HK98b, IM90, MH94a, Boy94, CBBC96, CR93b, CCB^{+}99, Fie97, Guy91, MLR^{+}94, MKBN99, Pae97, SSFP90, VMDC96, WPAA95]. groundwater-fed [MLR^{+}94]. group [Den97, MMC94, Str97]. growing [DMM95]. grown [TK90]. Growth [AB94, CKR93, CBC^{+}91, CJD^{+}99, CFT92, GC97, Hur90, KDKS90, McK96, OS91, RV99, SGC^{+}97, SZA92, VDG^{+}97, WG97, WMT96, AP92, Ano93r, Ant99, Arm99, AS94, BB98, CL90, CGVL95, CW95, Dal90, DC96a, DHB93, EF95, FGB95, FSS94, FR91, FL90, FFF98, FSSI97, Fur91, Gei90, GW99, Gen91, Goe98, GG90, GAAG96a, Gra91, HY97, HBH97, Har96, HD97, HJ94, HSRR96, HS99, JRH90, JCH92, KMM90, Kro93, LLLB97, LC90a, LBP97, Leh91, LPA96, LGT95, LB91, MP92a, MDE97, MD97, MT99, MFLR96, ML92, NB97, NSJ90, NE93, PH92a, PSJ99, PM91, Rot95, SCFB98, SD95, SD94, SN90, SM91, SR94, SCEG95, Str97, SW91, SB93, SH98b, SH98c, Tra94, TB96, UCS97, VHH96, WW95, WGSP91, WT94, WMT98, WWT^{+}93b, vdBG94]. Growth-compensating [SGC^{+}97]. Guam [Mat93]. Guglielmo [Ohm96]. guide [And99, Bac99, BF97a, Boo94, San99]. guidelines [Ano95e]. Gulf [Ano93x, BTL^{+}95, BHAV91, JCH92, JRT96, KCC98, KWF90, KKM^{+}96, KHK^{+}98, KBMSP99, LP95, MMPS98, OR94, SBB99, SS99c, UC98]. gulls [MJG91]. Gurevitch [Mil95]. gut [BM99, MR92b, PF90, THH98, TSH99]. gut-fullness [PF90]. Gymnodinium [MSK^{+}97, TS97]. gyre [NB97]. Gyrodinium [Ano93v, NE93]. Gyrotaxis [MOF90].

H [Abb90, AHKH^{+}96, Ano93h, Ano93f, Ano93g, BCBC93, Bur94, Edm96a, Lew95, Log95, Mei90, Rav95, Sal92b, Smi90a, SW93, Jum90, Bel90, HB94, KV99, Mil91, MZ90, TB91]. H-thymidine [KV99]. habitat [ES96, LL96, RKA96, SGG^{+}99, SHF^{+}96]. habitats [Ano93p, Den97, HC96]. Hackney [SW93]. halocarbons [Kli92, NML95]. halophytes [KLB^{+}99]. Halosaccion [Joh94]. Handbook [Duc94]. Hanes [Tho90]. Harbor [MJG91, MM91, TGG94]. harbors [MBCB97]. hard [Sar92, Wal92]. hard-water [Sar92]. hardwood [JM98]. Harmful [HGP97, Sma97a, DO97, Fra97a, KL90, MSK^{+}97, Pae97]. harpacticoid [DM90]. Hart [Mei90]. hatching [GH95]. Hatteras [DHWW97, RKG97, PPK^{+}99]. Hawkins [Ano93g]. HCO [CP97b]. head [BH97]. headspace [NJLW97]. headwater [VMDC96, WCW^{+}91]. heartbeat [ASC94]. heat [BH96, ESG91]. heavy [CBS99]. Hebridean [RS98]. height [CW93a]. Helens [WMD90]. helgolandicus [GH95]. hemolymph [Ano93u]. Hemond [Log95]. Herbivores [Els98, HDP^{+}90, MHM^{+}99]. herbivorous [CH96, SWS^{+}96]. herbivory [FGB95, KK91, WGSP91]. Heritable [DB93]. hermatypic [TKAD95]. Hessen [Bre93]. heterocysts [PPK95]. heterogeneity [Ker95, MOF90, Sti92]. Heteroptera [RB94]. Heterotrophic [RCG98, SPB90, CP97a, DŠ98, DŠ99, FPL^{+}91, GV93, GP99, HK95, JC95, JBJB96, JWRS96, KKW90, Li98, MLB99, MMW^{+}97, MMW^{+}99, PSH^{+}96, RDK96, SW96a, SD94, SBGBC96a, SBGBC96b, SB93, TZR98, TK90]. heterotrophs [FBS99, GGFR91]. hexaphosphate [SK95]. hexosamines [DM98]. Hg [SS99c]. High [ECA99, FBCS90, KDZD94, MBSW98, ASMB98, Ano93n, AS94, BLS^{+}95, BPCL97, CA91, Cul91, CGSD93, DMA^{+}96, DW91, EP97, FCC99, GSC^{+}96, Joh94, Kap92, KLR94, MAN91a, MVT99, MM93, MKBN99, NHJ91, PPB91, Rah90, SBB^{+}93a, SGP92, SGP96, SR94, SS99d, SKR^{+}91, TMLS99, WR95, WBM93, WLCG96]. high-elevation [WBM93, WLCG96]. high-light [BPCL97]. high-molecular-weight [BLS^{+}95, GSC^{+}96]. high-nutrient [Cul91, DW91]. high-plains [Rah90]. High-pressure [ECA99, SS99d]. High-resolution [KDZD94, MKBN99, TMLS99]. High-sensitivity [FBCS90]. high-temperature [Kap92, MM93]. highly [FBW96, Kir94b, MaB98]. Hildrew [Man95]. Hill [TGG94, GG90]. Historical [JD96b, BM96, KRW99, RA98]. histories [DW99]. History [Dum98, Bla93, DA92, Kir94a, MR92b, SH95a, Ban90c]. Hit [DFM92]. holarctic [ČH99]. holoplankton [MSSM96, SE90]. holothurian [MBOCK95]. holothurians [RBMH91]. Homeostasis [GMR99]. Hootsmans [Wet95]. horizontal [Ura90]. Hormothamnion [PPP97]. Host [CMT93]. hosts [SC93b]. hot [FCM^{+}97]. houses [SSP^{+}94]. HPLC [GK97, HW91, SHKI95]. Hsi [KL96]. Hudson [FPL^{+}91, SHB99]. hudsonica [DDC92, DD92b, DD92a, SB90, WLY93]. Hughes [Lap99]. human [CM94]. human-caused [CM94]. humic [CH90, De 93, GDG99, GLT96, HAL90, KZM90, KHS97, MH90, MH94b, NE93, RHH97, SJDJ94, Sha94, dHJS90]. humic-rich [RHH97]. humid [TGSB96]. Hure [Ohm96]. Huron [FCI91]. Hurricane [FAC^{+}99]. Hutchinson [Ano90g, Ano91e, Ano92a, Ano93b, Ano95b, Jón95, Edm91b]. huxleyi [CP97b, LTPB98, LVvdB99, PH97, TC95a, VBK98, WS96]. Huys [Ano93e]. hyacinth [CNP94]. hyalina [DW99]. hybrid [DW99, SH97]. hybridization [LF91, LK94a]. hybrids [TH92]. Hydra [EEL97]. Hydroacoustic [Ano93q, BLD99]. hydrocarbon [CFP97]. hydrocarbons [SM99a]. Hydrochemical [WB96, WLCG96]. Hydrodynamic [AWL94, Dad93, ES93, MC92]. hydrodynamically [SGP92]. hydrodynamics [BMT^{+}96, JB94, LL95, PB93]. Hydrogen [SC93a, SML96, AOTR97, Ano95e, Ben94, BW94b, HB94, KMM90, MZ90, PM90b, SS98, WCD^{+}94]. Hydrologic [SM98c, ML96, WBM93]. hydrological [ND91, PTJ96, SK96]. hydrologically [TGSB96]. hydrolysis [Arn98]. hydrolyzable [SMH94]. Hydromechanical [Grü97]. hydroperiod [BW99a]. hydrophobicity [MLB99]. hydrostatic [CT93, dABL91]. Hydrothermal [CJVD97, PRCV98, TV99, VF94, dABL91, dARB91]. Hydrothermal-vent [CJVD97]. hydroxide [KNM96]. hydroxylation [McC91]. hypereutrophic [DG99]. hypersaline [JM93a, JM93b, JMMD93, JAMH96, Mel99, MJOC93, SSKW91]. hypertrophic [Pra90, Som89]. hyphomycetes [GC97]. Hypolimnetic [HLGA98, VGP^{+}92]. hypolimnion [LC90b, MMS93, ML95]. hyporheic [BDV99, DMF^{+}98, Fie97, FSGG93, Fin95, MMW^{+}97, MMW^{+}99]. Hypotheses [Cul91]. hypothesis [BW91, Cul95, FS99, Ham94, LG99, MDE97, PLP99, Som93, Som95, TSD97]. hypothesized [CACM98]. hypotrichous [RERB91]. hypoxia [JRT96].

I.Friedmann [Amm96]. Ianora [Ohm96]. Ice [Liv99, ARM96, AS94, AR95, Dal90, DMA^{+}96, Eic94, ESG91, FCC99, FP99, Gli98, HHJ98, KFW94, LP94, MaB98, MVM98, MBT^{+}98, MSM96, MA96, Nic98, Rah90, RLH95, SN90, TCB^{+}98, VWF96, VRL^{+}98, WDGS92, WMC^{+}96]. ice-covered [Gli98, HHJ98, LP94, MBT^{+}98, MA96, RLH95, TCB^{+}98, VRL^{+}98]. Idaho [CMN93]. Identification [CM95, HW91, HGR95, LDC99]. identity [ZFT94]. Idothea [Som97]. II [LBL^{+}98, OHMK91, WHK^{+}93, XS93, RC95, SCF99]. III [KNM96, RW96, WHK^{+}93]. IIRMS [MMH^{+}95]. image [Pse90, VSBZ^{+}94]. imaged [SBW^{+}98]. imager [GSG94]. images [CHC97]. imagines [Ber99a]. Imberger [Spi96]. immersed [JAS95]. immobilization [RMT96]. immobilized [Fie97]. Immunofluorescence [CPSA90]. immunoprobe [OP92]. Impact [BTLC93, CCKP98, MWW97, PRWG97, WH93b, BLR92, CG95, HKPM99, MS99, MPG95, OYWK96, PL95, RBH99, SO91, Som94, VLHL97, WSB98]. Impacts [MT95, BG97a, DO97, KOWH99, LK99, MM97a, MHM96, RFHG99]. impediments [AWL94]. imperforata [CGD90, LDC99]. Implication [KHS97, KV99]. Implications [AP92, BKB97, BLK95, DAC90, DCML96, ES90, FAE99, FPC99, Gla91, GCJ97, Ham94, KK91, LC90b, LVM^{+}90, LvGLC98, MVT99, Nic98, PVL^{+}96, Pro90, PC97, QWR^{+}92, RC95, SS98, SB97, SFC90, UW93, WB92, Web95, WSMS96, AGGD90, BFG94, Bra91, BGF92, FVL98, FVL99, GP94, HK98a, HCHD94, HPJW94, Jac93, PSJ99, RJ91, SH95b, WSM^{+}96, ZFT94]. Importance [AGZ99, BG97b, Fin95, JA92, MR92b, Pae97, SCPS98, SGWS96, AK92, Ano93p, BP91, BTM92, CNV94, CL91, CFSW91, CS97, CFT92, EP97, FRL94, FBS99, GP99, JHM96, JL94, JW92, MB98a, PV94, WMT98, WC91]. important [CWW98, SBZD97]. imported [HLL98]. imposed [GGRJ94]. imposing [WKK^{+}95]. impoundments [YAF^{+}92]. improve [Roe98]. improved [DDG94, Jon91a, VFAA98]. Improving [Fil91]. in-water [GD92]. inapplicability [EBH92]. Inc. [Ano90a]. Incorporating [Pal92, CHA91]. incorporation [Bel90, BH95, CCR^{+}95, GLS90, KV99, Rom91, TB91, WH94c]. Increase [CP97b, RG92a, MMW^{+}97, MMW^{+}99, TIOS99, WCT90]. Increased [CR99, RF94c]. Increasing [DF94, FCG94, MVM97]. incubated [SRPG92]. incubation [Jön91b]. incubator [LSB92, MMG94a]. independent [Ebe92, Sch90]. index [Ano92d, ER95, WWMI98]. Indian [Wet91a, RGR^{+}98]. indicated [HCR^{+}94]. Indicator [PHX92, BC90]. indicators [DM98, EA99, GCD90, GLM92, IPPD92]. indigenous [RFHG99]. Indirect [BK90, CL99]. Individual [ADR93, App94, DO98, OCS96, OP92, WB92]. individual-cell [DO98]. induced [AML93, Bla93, DGW92, DPC90, DL92b, FT91, FAC^{+}99, GREGG97, HSFR98, HZF96, KTMI^{+}98, KMM90, LVvdB99, LJBTB99, LV97, MC92, NRR96, OYWK96, RWH^{+}98, WFG99, WLD93, ZCGP98, dJvB95, KZM90]. Inducible [PHMH96, Dod99]. Inducing [HFBB99]. Induction [BF98, FFJ92, LRvE94, OCS96, AM95]. inexpensive [KC92]. infaunal [MW98a, WWW98]. infecting [CS95]. infection [MWB^{+}98]. inferred [SLC^{+}99]. Inferring [Siv93, PF90]. inflows [MA96]. Influence [Ber99a, BR92, CV99, CK92b, GAA98, HGS94, KPSB95, KS93, LL98, MMG94b, MH95, Ver99, WH92, WWW98, dGP94, Ano93r, Ano93s, BKB97, BK95, CKR93, KNM96, LMSP99, PD90b, RG92b, SM98b, SM98c, TC94, Ver94, WKB^{+}96, CJVD97]. influenced [KY97a]. Influences [BSMK95, CAS93, NML99, SJL95, BDH91, CL99, CGB90, CH92b, HK98b, Sam91]. infochemicals [LV97]. information [Ben94, KJLA98, KJS98]. infrared [Ano93y]. ingested [CH96, DŠ97, KSSP91, RF94a]. Ingestion [DŠ98, CPS90, CH94, DC91b, LSS^{+}97, Leh93, MMCL94, MO91, PF90, PF91, Rot95, TZF96]. inhabited [HCJ96]. Inherent [HVB93, BRZ95]. inhibit [Les96]. Inhibition [NCD98, HJ94, LK94b, SNB92]. inhomogeneity [BC90]. Initial [BW94a]. Initiation [KKM^{+}96, TS97, BMAW97, SP90b]. Initiative [Thr91, Wet91c]. injection [HA92, LB95, MS93b, SMC98, SMH94]. inlakes [Ver94]. inland [MUGK92, Mei90]. inlet [KSCG98]. inner [PBL99b]. inner-shelf [PBL99b]. Inorganic [BFE96, LD99, RA97, CW92, DY99, FRY95, GGFR91, GLT96, IM98, KRB98, LTPB98, MMH^{+}95, RJ91, RMT96, VV98, VL98]. inorganic-carbon [RJ91]. inositol [SK95]. input [GREGG97, HKPM99, JM98, KPSB95]. inputs [GJC92, MS99, MH94a, MBCB97, SS99b, VBD^{+}97]. insect [HT98]. Insights [CCL92, NWS94]. instars [Swi92]. instead [Som92]. Institute [WL99]. instream [Mul92]. instruments [GD92]. intact [Jön91b]. integral [WWT^{+}93b]. integrated [Dod92a, KC92, MPA94, SSFP90, WMPO99]. integration [ZKM93]. integrifolia [HSL^{+}97]. intensity [Ano93s, Som94]. Intensive [KTMI^{+}98]. Interacting [Leh91, PB93]. Interaction [DHB93, MM97b, Ham94, HCSE97, ME98b, SLD97, SML96, TW93]. Interactions [SH98b, TBM95, AC92, Ano93g, CL99, CMT93, Cur90, DC99, FDG99, HHI96, KHP91, LL90, MPG95, RFHG99, Sal92b, SH92, THK^{+}99, Wic95, YD99]. Interactive [BDH91, GW99, JLO97, SH98c, DL93, VL98]. Interannual [JPG90, MD97, TH96, JW92]. interceptor [HN94]. Intercontinental [ČH99]. interdisciplinary [Pla93]. Interface [CK92a, CR93a, HCS95, JCMW91, KG95, PL95, RERB91, SBGCC91, TG92]. interference [LB95, RCCdG99, Sch90]. intermediate [FS99, Som95, TSD97]. intermittent [DA92, WLBQ95]. Internal [BSO92, RSZ97, SI98, GP94, LCR99, LWMD96, LJH91, MWI92, Per97, Pin99, TJ98]. interplay [RV99]. Interpretation [BRPR98, RC95]. interpreted [MA96]. interpreting [HRD^{+}95]. interrelationship [OAR96]. interreplacement [SH95b]. Intersegmental [AE91]. interspecies [AKM95]. Interspecific [HPJW94, Sti92, TH92]. intertidal [AOTR97, Ano93l, CPMW99, DG90, EG91b, FBR91, HQ90, Joh94, Ker93, Meh99, MDWS97, PWM^{+}98, PZ93, Roc98, Set92, SU98, SM99b, UPP95, WF90, WWW98]. intestines [WBG^{+}91]. intra [Sti92]. intra- [Sti92]. intracellular [SR92]. Intraspecific [Som92, PHMH96]. introduced [MT95]. introducing [CLP^{+}92]. introduction [Jum94]. Intrusion [WPAA95]. inundation [HKvS^{+}92]. invader [HY97]. invasion [HPB^{+}99, LC93, RFHG99]. Invertebrate [BL91, BL94b, CM95, HPM96, Leh93, LC93, ME98b, Pol95, VCT^{+}99, WCW^{+}91, WO98]. invertebrates [HCHD94, HCJ96, HHL98, LHM^{+}93, MB90b, SE90, SGG^{+}99, WTH98, WLBQ95, dARB91]. investigating [CG91]. Investigation [MMTJ94, RB91, BMT^{+}96, MZ90, WWT^{+}93b]. investment [Wil92]. involved [FR99, Wal92]. involving [TBM95]. iodide [MGN92, MdlC97, SM99a]. ion [EP96, MBSW98, SG93, XS93]. ion-selective [MBSW98]. ionic [CBS99, Sha94]. ions [MA94]. Irish [ASS^{+}98]. Iron [CH90, GCJ97, HDB93, LBB^{+}97, WPB94, WT94, AZS^{+}98, Arm99, BMP92b, Ban90d, Ban91, BW91, Bra91, BdBNvB91, CBMC^{+}99, CP97a, CBC^{+}91, Coa91, Cul95, DGW92, DBT91, DHB93, DMM95, DT91, EA99, EP97, GP99, GGF91, HVHH91, Hon97, HWF95, HDZB98, HFBB99, LBLO91, MGF91, MFW^{+}91, PB91, PAM94, RB91, RM90, RB97, SdBdJL98, Sha94, TBM95, WLD91, Wel99b, WMT96, WMT98, WL94, dHJS90, KNM96]. Iron-limited [WT94, MFW^{+}91, PAM94, WMT96, WMT98]. iron-replete [HFBB99]. Irradiance [NTH^{+}91, AMLH97, Ban90b, Ban92, BSMK98, CP99, Dod92a, FVL98, FVL99, Gor91, GC90, HB91, LPJ92, LM98a, MP92b, McC95, NHJ91, NS91, SWH92, SZA92, TC94, WH99, WHB95]. Irradiance-dependent [NTH^{+}91]. irradiance-mixing [NS91]. irradiance-regulated [SWH92]. ISAAC [KW98b]. Island [GAA93, PL99, PK94, SLA93, WLY93, WGS97]. Islands [Mat93, SH90]. isoetids [WAC^{+}98]. isolated [ECA99, Jun92]. isolates [MC99]. isolation [MW98b]. Isopoda [Som97]. Isotope [CTF98, HFK92, Ano93f, Ano95e, BBS97, BGC94, BGN98, BdBC99, BLK95, CE98, FPC99, KCC98, KSH96, Leh93, LKSA97, MVM97, MSH96a, MA96, MMH^{+}95, RC95, SS98, SR92, TMLS99, TC94, TC95a, UMS^{+}99, WHN^{+}98, YMT^{+}99, dHJS90]. isotope-balance [CE98]. isotope-pairing [MSH96a]. isotope-ratio [Ano95e]. isotopes [Ben94, BW94b, BLK95, Fra95, HB94, MB90a, Rav95, SH95a, WHM94, YKM95, ZEG^{+}94]. Isotopic [BDY^{+}98, CFW97, PVL^{+}96, Alt90, CCR^{+}95, CWSX96, FGN93, FAC^{+}99, GCD90, GSC^{+}96, HSFR98, HS98, LBP97, SH91]. isotropic [HCS95]. Israel [KKS^{+}99, MKBN99]. Italy [MDC95]. ither [Jum90].

J [Ano93g, Arc93, Cur94, Dua95, Duc94, Edm96a, Edm96b, Hai96a, HL94, Joh91, Kim92a, Leh95a, Log95, May92, Mei90, Mil95, Ohm96, Pai96, PA95b, Pri94, RW95, Sal92b, Sel94, Spi96, SS93, Str94, Str95, War90, Wet95, Wet91d, Wri90]. J.-M [May92]. Jackson [NT93a]. Jahnke [CP95b]. Jamaica [HDR^{+}98, Lap97]. Japan [RWH^{+}98, TT97a, UNN^{+}98b, YK93, Spi96]. japonica [YK93]. jellies [EABK99]. Jellyfish [LBLO91, LLB^{+}96, OBL^{+}91]. Jersey [RS99]. Jesogammarus [TT97a]. jet [BS90a, OMK95]. JGOFS [FBS99]. Joe [Ste94]. John [Ano93g]. Jorgensen [WMTB93b]. Jumars [Pla93]. June [Ano94]. juvenile [CBS99, HC96, KDKS90, RF94b, RBH99, TJC97, TJC^{+}98]. juveniles [Hen98].

K.-C [Cha96]. K.-H [Abb90]. K7 [Jan93]. kairomone [HD95, TvE94]. kairomones [Ber99a, DW99]. Kalff [Wet91b]. Kamenir [ADC92]. Kapp [Ano93h]. Karelian [BMT^{+}96]. Karsten [MM99]. Kasumigaura [UNN^{+}98b]. kelp [HLL98, HD97, HSL^{+}97, WLD93, GNL97]. Kemp [Duc94]. Kenya [Kit96, RA98]. key [HLGA98]. Keys [LWMD96, PLP99]. Kiefer [Ban90a]. Kiefer-Mitchell [Ban90a]. Kiel [RG92a]. Kilham [Ano93d].

Kimmel [Thr90]. Kimmerer [Pol90]. kindti [Bra98]. kindtii [BL91, MC95]. kinetic [WWT^{+}93b, YD99]. Kinetics [HM90, FFF98, Gar91b, HHI96, IH95, MSS^{+}95, Mor90, OS91, VHH96]. Kinneret [MKBN99, BFZED94, BSY^{+}95, KKS^{+}99, SBB^{+}93b, ZEG^{+}94, ZPHH98]. Kitchell [SS93]. Kitching [Pai96]. Knowles [Ano93f]. Knox [Smi96]. kowalewkii [KM94]. krill [GWY^{+}92]. Kristiansen [Leh95a]. Kszos [TMB96]. Kumagai [Spi96]. Kvet [Wet91d].

L [Arc93, Ban90c, Cha96, Ohm96, Str94, Thr90, War90]. L. [BFGS94, RPJ97]. labeled [KHB96, SRG98, SLA93]. labeling [GW93, Leh93, dHJS90]. labile [LS90, MZ97]. lability [SHM95]. Laboratory [MGN92, VdHH^{+}98, BPA96, EEL97, NFGP^{+}93, HS95, Kap94, MK95, PBL99a, Rot92, WF90, Liv90]. Labyrinth [Coa90]. Laccadives [MAN91a]. Lack [FDLD93, GV93, Loo93, TC93, ADC90]. lacustrine [AGZ99, DGW92, Smi90b, TMLS99]. lacustris [SJP94]. Ladoga [MMTJ94]. Lago [MC95]. Lagoon [MDC95, RMT96, Ano93k, MC92, PBL99b, TM95b, VTCB96]. lagoons [FZD93]. Lagrangian [FBS99, WW98]. Lake [BEB92, Ben94, BCM92, CC97, Dua95, DCML96, Fri90, FMK^{+}99, HB94, JPG90, JM93a, JM93b, JMMD93, JRM98, LBLO91, LS93, LLB^{+}96, MFJR99, MCCD92, MCD^{+}96, MTLR98, MJOC93, NTH^{+}91, OBL^{+}91, PA95b, SSFP90, SBWB94, Ste94, SS93, UBBS94, And90, ASS^{+}99, AR95, AR96, BVSBB93, BMT^{+}96, BK95, BBH^{+}92, Boy94, BPS94, BLD99, BFW92, BS99, BW90, CCL92, CL91, CSPG94, CAS93, CACM98, CCC^{+}96, CRKS93, CCL90, CC98, CW95, CA91, CGB90, CGSD93, De 93, DC91a, DMA^{+}96, ESG91, ES90, EM93, FHKC96, FCC99, FP99, Gae92, Gae93, GLWH90, GS93, GSL97, HK98b, HGS94, HTW90, HTDM96, Han96, HLGA98, HAL90, Hon97, HCLD95, HGR95, JEP^{+}99, JUSI91, JSL92, JCM^{+}99, KSSP91, KM95, KGFP91, Kna99, KJLA98, LFGM96, LAYE95, LFHS99, Les95, LJH91, LKSA97, LDL91, LDV^{+}92]. lake [LHS90b, LAC^{+}96a, LAC^{+}96b, Mac93b, MMS93, MBT^{+}98, MR98, MLR^{+}94, MHM96, MSP^{+}99, MGW96, ML95, MA96, MMH^{+}95, MWW97, MBSW98, MWI92, OYWK96, OBH^{+}91, OBKH96, PMF90, PC96, Per97, PASC95, PSH^{+}96, Pra90, PDM96, QLBAP96, Rah90, REL97, RHH97, RHT92, RA98, RLH95, RSZ97, RM96, RKST90, SI98, SJDJ94, SPB90, Sar92, SML96, Sha94, Siv93, Som89, SHM95, Spi96, Sta91, SFW^{+}95, SCEG95, SBGCC91, TMLS99, TCCA93, TJ98, Tho99, Tra94, Tra97, THS^{+}91, Two94, UNN^{+}98b, VT90, VWF96, Ver93, Ver99, VCT^{+}99, VL96, WKB^{+}96, WMPC^{+}90, WMM^{+}94, WBRJ91, WMPO99, XS93, XKS95, YK93, YK94, YH92, dHJS90, dGP94, dGBPP96, dHK98, AHKH^{+}96, AGZ99, AR95, BMP92a, BMP92b, BW94b, BFZED94, BSY^{+}95, BBS97, BL91, BE94, CG97, CFSW91, CR93b]. Lake [CFT92, Dum98, ESC^{+}93, Edm91a, FMH^{+}91, FCM^{+}97, FG93, FP99, GCB^{+}98, GGA91, GJ96, Gra97, GS93, HL92, HBO^{+}94, HSFR98, HS98, HKP^{+}97, HHJ98, HB94, HA90, IM90, IPPD92, JBJB96, JGRR99, JAMH96, JL94, KSH96, KLR94, Ker95, KCWM96, KKS^{+}99, KS93, LS90, LFGM96, LCR99, LC93, LB94, LKSA97, Liv99, LL90, MMTJ94, MCB90, MKWD97, MHZ98, ML95, ML92, MKBN99, NKT^{+}93, PCFS97, PPS^{+}97, PJ91, PSH^{+}96, RA98, RWH^{+}98, RS93, SFRB98, SH91, SH95a, SC93a, SS98, SGZ93, SR92, Sti92, SWSC90, SBB^{+}93b, SBGCC91, TH98, Thr92, TT97a, TGSB96, TCA96, UNK95, UNN^{+}98a, UNN^{+}98b, VGP^{+}92, VFW^{+}92, VJM96, WH95, Wet99, WB96, YMT^{+}99, ZEG^{+}94, ZPHH98, Smi90a, Wet95]. lake-atmosphere [ASS^{+}99]. lake-ice [AR95, DMA^{+}96]. Lakes [BM96, MQ96, WPAA95, ADC90, ARM96, Ano93i, Bac99, BPK94, Boy94, CNP94, CV99, CCKP98, CPCS93, DG99, DHK^{+}96, DK93a, DM92, DAC92, DCML96, EBH92, EG91a, EP96, EP97, FHKC96, Fra95, Gli98, GLT96, GLM92, Guy91, HNS95, HT98, HCSE97, HCH93, Hec99, HCLD95, HR90, HW91, HW94, IKT99, JA92, JEP^{+}99, KWS98, Kil90, KTF96, KE91, KPSB95, KHS^{+}98, KFEC91, LPPSR97, LMH98, LS92, Lew90, LP94, MS93a, MHSP90, MVT99, ML96, MT94, MHM96, MAS91, MGWB97, MHZ98, MSM96, MZW^{+}95, MH97, MMC94, Nür95, Nür98, Nür99, PV94, PPK^{+}96, PK91, PSH^{+}96, QWR^{+}95, RKR91, RRK94, SBP^{+}96, Sch99, SF95, SHK95, SLC^{+}99, Sta90a, Sta90b, SHF^{+}96, SHKP96, SSM^{+}95, SM98c, SB94, TZGS98, TCB^{+}98, UG98, UGU^{+}90, VORJ93, VL91, VRL^{+}98, WL92]. lakes [WR95, WMHB95, WMD97, Web90, WH94a, WHM97, WSM^{+}96, WHOL99, WMD90, WMC^{+}96, dGP94, FCI91, MF99, Nic98, SWS93, Pri94]. lakes-Evaluation [Boy94]. Lalli [Jum94]. Lamarra [Sau92]. laminar [LR99, LCG94, SJL95]. Laminaria [HD97]. land [MV98, SLC^{+}99, WPBS99]. land-based [WPBS99]. land-derived [MV98]. land-use [SLC^{+}99]. Landcover [MBK^{+}99]. Landry [CG93]. Landscape [PRT^{+}93, KSH^{+}99, PLP99, WKB^{+}96]. landslide [Ano93i]. landslide-dammed [Ano93i]. Lanyang [KL96]. Large [Cur90, ADS94, Gae92, Gae93, GGFR91, GSL97, IKT99, KTF96, LDL91, LDV^{+}92, MCED94, PASC95, Rie92, RM96, Spi96, SB94, WMPC^{+}90]. Large-scale [Cur90, MCED94, SB94]. larger [CPCS93]. larvacean [SSP^{+}94]. larvae [AMF98, App94, CM95, DPC90, Eck90, FSS94, FH92, HCKW90, HBI95, Hur90, MBOCK95, MC96, MHT99, RF94a, SGG^{+}99, TZF96, TZFP^{+}94]. larval [Dal90, EEL97, FR99, HC96, LHM^{+}93, Lev96, LJA91, MMCL94, MK95, Pin99, RBH99, SM92, SE90, Str95, Swi92, TM95b, TJC97, Wal92, YGB91, ZFT94]. laser [HC97]. last [LKSA97]. late [HHR94]. lateral [HCP90]. Lathja [Rav95]. latitudinal [DW93]. Laurentian [SWS93, WMC^{+}96]. law [GD96]. Lawrence [MF99, BTL^{+}95, KBMSP99, OR94, PLTL96, SS99c]. Laws [Ban90a]. layer [GGRJ94, HHI96, JX94, JD90, Ken93, MJ94, MB90b, OMk93, OMK95, OBH^{+}91, Smi92]. layered [KZ92]. layerflow [BFGS94]. layers [ALB90, CP94, GWY^{+}92, GWMB91, PSEJ99, PSJ99]. Lazier [Sal92b]. Lead [FGN93, MF90, NLSG99]. leads [RF94c]. leaf [GC97]. leakage [IH95]. legacy [Ver90]. LeMahaute [Tho90]. Lemna [GDG99]. length [Bra98, ML90]. Lepomis [EEL97]. Leptodora [LL90, BL91, Bra98, MC95]. leptopus [vLM91]. leucine [KV99]. level [ADR93, Fri90, MKD90, PD90a, RA98, ŠBM^{+}95]. levels [Ano93u, Ant99, Cam94, Har96, NS91, REL97, SZA92, ZCGP98]. Li [Har95, SK96, TMB96]. Life [Bla93, AO96, DA92, DW99, FGN93, LH91, MB98a, VGP^{+}92]. life-cycle [VGP^{+}92]. life-history [DA92]. lifestyle [Duc98]. Lifetime [Ano95a]. ligand [HFBB99]. ligands [Bru92, LVvdB99, SDB97, XKS95]. Light [AP92, FK98, GAAG96b, HCKW90, KJSJ98, LM98b, MBHH^{+}91, And93, Ano93s, Arm99, BH95, BF98, BM92b, BSMK95, BPCL97, BH91, BC90, CP94, De 93, Dod92a, DASV98, FT91, FPAG^{+}98, FAE99, GCCJ97, GCB^{+}98, GJ90, GMK98, HDR^{+}98, HS90, JL92, JPJ97, Kir91, Kir94b, KE91, KJ92, KJ94, LW93, Mac93b, McB92, MM99, MWW97, NS93, OCP92, PASC95, Pic91, PZ93, SGC^{+}97, Som94, Sos99, SH98c, VdHH^{+}98, WLD93, WWT^{+}93b, WKZK98, vEJ97]. light-absorption [Sos99]. light-photosynthesis [BM92b]. light-scattering [WKZK98]. lightweight [Ver93]. lignin [OB98]. like [BG97a, HHR94, NK99, SRLRL^{+}99]. Limfjorden [LPJ92]. limit [Ban90d]. Limitation [CNP94, MHP91, ADC90, Arm94, BH95, BF98, Bre93, CP94, CBC^{+}91, CMPD92, DG99, EA99, FZD93, FAE99, FZP92, GGF91, GKS^{+}94, HDZB98, HFBB99, JL92, JBJB96, KKW90, KKK^{+}93, KMM90, KKBG91, MBHH^{+}91, MM99, ML92, MN95, NT93b, PWR94, RSK97, RA97, RWH^{+}98, Rot95, SS99a, Ste90, SHSS93, Ste94, SB90, SH96b, TZR98, TIOS99, UW92, UCS97, VT90, WR95, WPB94, ZR98, vEJ97]. limitations [HCPR99]. limited [FSS94, FSSI97, Lap95, MFW^{+}91, OJ98, PPB91, PAM94, Rot92, SGC^{+}97, Som92, SM91, SB90, THK^{+}99, WT94, WMT96, WMT98]. limiting [BC97, EF95, PB91]. limits [SHF^{+}96, WK90]. Limnetic [GS93, BTM92]. Limnological [Kna99, Wet91b, LPA91, Wet99]. Limnologie [Edm96b]. Limnology [Ano90a, Ano90i, Ano91f, Ano94, Cot99, WL99, Jón95, Jum90, Kal91, Thr90, Thr91, Wel99a, Wet91c, Ano90f, Ano91d, Lew95, Leh95b, McC97]. limpets [HQ90]. Lindeman [Ano92b, Ano92c, Sut93a, Ano95d]. lineages [MBKG95]. linear [IH95]. Lingulodinium [LR99]. link [CACM98, VORJ93, Wic95]. linkages [PRT^{+}93]. Linking [DMHH99, KTF96, MV98]. links [Els98]. Linuche [KCKC90]. Lipid [HK98a, CFW97, HQCG92, LGI94, WL92]. Lipids [VGP^{+}92, Hen98, MMPS98]. liquid [BW94a, DSV^{+}99, MBSW98]. lithium [SK96, TMB96]. lithology [VMDC96]. litter [GC97, WH94c]. Little [Pai96, BEB92, FMK^{+}99, SBWB94, UBBS94]. Littoral [JB91b, Fra95, JB91a, MD97, PCFS97, PK91, PPS^{+}97, SBGBC96a, SBGBC96b, SBGCC91, VL98]. Littoral-pelagic [JB91b]. living [BS90b, DFA93, Ovi99, San99, SAK91]. load [Per97]. loading [BA95, Bur92, CGB90, CH92b, DY99, MB99, SH95a, Smi90a]. Lobelia [RPJ97]. lobster [AB99]. Local [BC97, DBS99, Liv99]. locally [BC97]. location [GJC92]. locations [Wal92]. lochs [BFM^{+}97]. Locus [HC94]. log [JF96, JF99]. log-normal [JF96, JF99]. Long [FHN^{+}96, NTH^{+}91, SH96a, VCT^{+}99, WH94b, BSY^{+}95, DG99, HGS94, HC99, JCH92, MR98, PJSW96, Pse90, Ree90, SGB^{+}97, SHF^{+}96, SB94, Won91, GAA93, PL99, PK94, SLA93, WLY93, WGS97]. Long-term [FHN^{+}96, SH96a, VCT^{+}99, WH94b, BSY^{+}95, HGS94, HC99, JCH92, MR98, PJSW96, Pse90, SB94, Won91]. longa [Hay95]. longicaulis [HDP^{+}90]. longicornis [Ano93m, PK94]. longirostris [BL91]. longispina [KM95]. loop [PH92a, WMPC^{+}90]. Loosdrecht [SB94]. loss [CA91, CGSD93, LFHS99, Leh91, RKA96, RKR91]. losses [GG99, SRG98]. lotic [Hur90, Pal92]. Lough [Edm96c]. Louisiana [BOB94]. Lovett [Joh91]. Low [DW91, CC98, Gli98, HD95, JB94, KTMI^{+}98, Lam94, LB96, MVT99, NML95, OCP92, RGRW93, SSE90, SRG98, SR94, SKR^{+}91]. low-chlorophyll [DW91]. low-food [Lam94, LB96]. low-frequency [JB94]. low-light [OCP92]. low-speed [Gli98]. low-wind [CC98]. lron [HM90]. LSU [SMB^{+}97]. lucens [BOFW93, Hay95, HWP95]. Lucerne [AHKH^{+}96]. Lucky [TV99]. Lugano [BBS97]. lumholtzi [HH93]. Luminescent [LR99]. Lund [Edm96a]. luteoviolaceus [RB91]. lutheri [CL90]. Lymnaeidae [Bre97]. lyngbyaceus [DCM90]. lysis [ASMB98, GHF^{+}97]. lysogeny [TH98]. lytic [CS95].

M [Ano93g, Arc93, Bur94, Cha96, Dol91, Edm96b, HL94, Jum94, May92, Men90, Mil95, Ohm96, Spi96, Str94, Str95, SW93, Tho90, Wet95, BUW96, FNWK96, HBH97]. Mackenzie [LMH98, ML96]. Macko [Bur94]. macro [JWRS96, Ste94]. macro- [JWRS96, Ste94]. macroalgae [MGN92, NML95]. Macroalgal [VMH^{+}97, HLL98, KAB92, Lap97, Lap99]. macrobenthos [HPTC97]. macrochirus [EEL97]. Macrocystis [HSL^{+}97, WH92]. macrofauna [AK92, KAB92]. macrofaunal [SGP96]. Macrohectopus [Ano93q]. macromolecular [SBW^{+}98, TB91]. macronutrient [MBHH^{+}91]. macronutrients [CL96]. macrophyte [AV94a, AV94b, BK99, BdBC99, KV99, LL96, RPJ97, Wet95]. macrophyte-derived [AV94a, AV94b, BdBC99]. macrophyte-dominated [KV99, Wet95]. macrophytes [BSO92, BW90, EW99, HC94, KMM90, LL96]. macroscopic [GS93, GSL97]. macrozoobenthic [BC97]. macrozooplankter [Ano93q]. macrura [HK98a]. Madison [Cot99]. Maestrini [Har95]. Magellan [Ohm96]. magellanicus [CH94]. Maggiore [MC95]. MAGIC [KT92]. magna [LL96, ASH99, DB93, DBS99, DGS98, Ebe92, Gla91, Pla98, SR94]. magnetic [Hil90]. magnitude [GJC92, KOWH99, RRK94]. Maine [Ano93x, BHAV91, KWF90, MMPS98, Sta90b]. maintaina [CCP92]. maintained [Ura90]. maintaining [SGP92]. maintenance [Gar91a, KBB98]. Major [OBA99, CG97, DHB93, FMH^{+}91, LPPSR97, LMH98, MA94]. making [SGB^{+}97, But95]. Malawi [PJ91]. male [IMB^{+}99, vLM91]. malfunctioning [THR97]. malhamensis [CPS90]. Management [Sch99, CK92a, Kna99, SS93, Wet91a, Wet91d]. Manganese [BFM^{+}97, BMP92b, BdBNvB91, Coa91, Hon97, JUSI91, LBLO91, MS93b, Mof94, SdBdJL98, SH90, SH96b, SH98c]. mangrove [Kit96, RMT96, BCM92]. mangrove-dominated [Kit96]. Manipulating [Wel99b, MHM^{+}99]. manipulation [SC90, WCW^{+}91]. manipulations [DC91a]. Mann [Sal92b]. Mantoura [May92]. manuscript [Ano95e, Kir94a]. many [Hai96a]. Maracaibo [GCB^{+}98]. March [EK90]. Margalef [Leh95b]. margin [AZS^{+}98, BG97b, FFNS99, KMU95, May92, SMV98, TC96]. marginal [Dal90, KFW94, SN90]. Mariana [Mat93]. mariculture [Bie99]. marina [SBK95]. Marine [AECP95, And99, FCG94, JHM96, Plu97, AWL94, AKM95, AM95, AV94a, ABB^{+}95, AGGD90, AA92, AA91a, AA91b, AK92, Ano93g, Ano93o, Ano93s, Arm94, AR94, BP91, BB98, BH95, BF98, BB97, BOFW93, BCM92, Bra91, BG94b, BP97, BAO97, BZR99, But95, CL90, CP97a, CBD^{+}99, CDPR98, CJD^{+}99, CK95, CL96, CH91, Clo98, CHD91, CM95, CS95, CH92a, Cra95, DA92, DFA93, DL96, DBZF98, DCM90, Dol91, DŠ97, DC96b, DW91, DDC92, DD92b, DD92a, GdGD97, GHF^{+}97, GHT^{+}96, GP99, GGF91, GMR99, HA92, HBI95, HHB93, HPJW94, HFK92, HAAM99, HC97, HM90, HDB93, Ive90, JMA^{+}95, JRH90, JF96, JF99, KY97a, KK93, KHS97, KAD95, LPJ92, LBP97, LTPB98, LK94a, LRM95, LLZ99, LHM^{+}93, Li97, LH91, LW90, Lon91, LBL^{+}98, LLB^{+}96, MdlC97, MM93]. marine [MTM^{+}98, MHZ98, MF90, MM97b, MB96, Mof94, MLB99, MBHT94, MV96, MC99, MH94b, MC96, Muc91, MC97, NK91, NCD98, NSK95, ORH99, Ohm96, OBL^{+}91, PHMH96, PBJD93, PFB96, PRWG97, PM90a, PPB91, PB93, PW93, PPP97, PK94, RR92, RKG97, RJ91, RB91, RF94a, RSK97, RM90, Rie92, Rom91, RFHG99, Rut92, Sal92b, SLD97, SBW^{+}98, SDJK98, SAK91, Sel97, SE90, SW96a, SB97, SRG98, SRG00, SGG^{+}99, SM98a, Som94, Som95, Sos99, SLT99, SH98a, SR93, SBZD97, SCKF99, SCF99, SH95b, SGSA92, TC94, TL94, TK90, TT97b, Ver97, Ver91, VRT^{+}92, VC90, WTH92, WBG^{+}91, WH92, Wal92, WRLF96, WF98, WHN^{+}98, Wil98, WKZK98, Wym92, YM96, ZFdSY96, dAL94, dGGV^{+}96, Boo94, Cur94]. marines [MAG95]. marinus [RR92]. maritime [GC90]. marker [ALB90]. markers [CM95, QLBAP96]. marl [BFW92]. marsh [BFS91, BdBC99, CPMW99, HKvS^{+}92, Hin94, KLB^{+}99, LL95, LKM98, ND91, PL91, Sau92, SD95, SM99b, SVC^{+}98, TH96, WH94b]. marshes [PD90a]. Martin [May92, SW93]. Mass [PJ90, BA97, BA92, BA95, CPMW99, CG97, CE98, Dad93, HCS95, HC94, JHM96, KLB^{+}99, KSCG98, OMK95, TA97]. mass- [CE98]. mass-dependent [HC94]. mass-transfer [BA95, HCS95]. masses [GCD90]. Massive [SVF91, PRCV98]. mat [DCM90, EKT99, JD90, PBJD93, PFB96]. material [AV94b, CSPG94, HHB^{+}92, HSRR96, KW98b, LHWZ92, PL99, SCPS98, SH98a, VRHK99]. materials [CH90]. Mathematical [McC95, DC96a, JUSI91]. matrix [FL95, WKZK98]. mats [MWW97, SBPV99, dBGEK97]. matter [AV94a, Alt90, AB96, BP91, BB98, BBS97, BB97, BL94a, BdBC99, BS90b, BG97b, CCR^{+}95, CTSB98, CMN93, DSV^{+}99, DM98, DMHH99, DŠ97, ECA99, GREGG97, GB94, GSC^{+}96, HCH93, HCR^{+}94, HS98, HVB93, HZF96, JAMH96, KHBL98, KAD95, Kro93, Lev96, LGT95, LO97, MBK^{+}99, MTLR98, MM97b, MPS^{+}91, MZ97, ORH99, OBA99, OBL^{+}91, PBJD93, PSS^{+}96, RGR^{+}98, SBW^{+}98, SH95a, SB97, SRG98, Sti91, SPMW97, TG92, THK^{+}99, Tra94, UMS^{+}99, VHH96, VL98, VGB94, VVK97, WS90, WHB95, dHK98]. maturation [Ebe92]. maturity [Ebe92]. Max [WL99]. maxima [LM98a]. maximal [DG90, NSJ90, OP92]. maximum [AEL^{+}99, BTL^{+}95, KHK^{+}98, NTH^{+}91, NHJ91, NS91, OGH^{+}96, ŠBM^{+}95]. mayfly [MT95, SC90]. Mazzocchi [Ohm96]. meadows [EM93]. Mean [Bou98, Ban92, McC95, MKD90]. means [CM95, DWM90, Jør93, KFEC91, McB92, VL91, WMTB93a]. measure [CCK98, CP95b, Dod92a, MO91, MKBN99, OP92, SRG00, TG94]. measured [AMLH97, BRPR98, CC98, GJ90, KV99, LK94a, MCED94, SNB92, SHKI95, TLvB^{+}91]. Measurement [BA97, HV98, OHMK91, Ver91, Ano93q, Ano93u, CI94, CH94, Gli98, GG90, GG99, HTSDZ99, HCJ96, LBL^{+}98, PA95a, SHK95, UPP95, Har95]. Measurements [CHA91, Li94, VBvG91, Ami90, AD92, AE91, DHSS95, DHWW97, DO98, FT91, Gor91, HVB93, JC95, Kim90, KJ92, LSB92, MVM98, MMG94a, MJC93, MBSW98, OZ95, PSO91, Pro90, SBB^{+}93a, SGB^{+}97, Sos99, TF95, TT97a, VHS^{+}95, VdHH^{+}98, WSM^{+}96]. measures [Ano93x, JOG96]. Measuring [GW93, HT96, SRG98, VWF96, ABB^{+}95, Boy94, BRZ95, BP97, CH91, CdG92, DMF^{+}98, FHF99, Jön91b, KT92, LW93, MSH96a, NT94, Str90]. mechanics [Men90]. Mechanism [KMM90, SVC^{+}98, ADM96, Jac93, MOF90, PM90a, PJ90, RB91, WS90]. Mechanisms [Jør93, Kil90, RJ91, WMTB93a, Bur92, Grü97, HC96, HvOW99, Joh91, MMC94, TZF96, YW93]. Medal [Ano90g, Ano91e, Ano92a, Ano93b, Ano95b]. mediate [KDB91]. mediated [LL96, LF93, MB98b]. mediation [CNL90, PWM^{+}98]. Mediterranean [ASMB98, Ano93k, BG97b, DASV98, GW90, HKPM99, JLMV98, KKBG91, MD97, MCC^{+}96, SM92, TZR98, VTCB96, VPN^{+}90, ZR98]. medium [HM90]. Meeting [Ano90i, Ano92e, Ano94]. megafauna [LSS^{+}97]. meiobenthos [GAA98, Liv90]. Meiofauna [AA92, HJ94, Pal92, SE90]. melt [Eic94, MSM96]. meltwater [CPT^{+}94, MA96]. membranacea [Grü97, PHMH96]. membrane [KSCG98, MBSW98]. Membranipora [Grü97, PHMH96]. Memoriam [Edm91b]. menace [DGHD98]. Mendota [HA90, LCR99, LL90, SS93]. mendotae [TH92]. menhaden [CCK98, KDKS90]. Menidia [RF94b]. menu [Rot91]. Mercenaria [JCH92]. Mercury [CGHK99, MFH^{+}93, SGWS96, WB92, HLGA98, REL97, WMHB95]. meres [MMC94]. MERL [KDKS90]. Mermaids [Ban90e]. Mermaids- [Ban90e]. meromictic [Hon97, OBH^{+}91, OBKH96]. Meromixis [JM93b, JMMD93, MJOC93, JM93a, JRM98, SL98]. Meroplanktonic [WBRL98]. mesh [Bre91, LB96, Lou90]. meso [SBGCC91]. meso-eutrophic [SBGCC91]. Mesocosm [LAC^{+}96a, LAC^{+}96b, SBK95, AR96, HBL^{+}93, KOWH99, WKK^{+}95]. mesocosm- [AR96]. mesocosms [KDKS90]. mesohaline [MT90, SHB99]. mesopelagic [SCPS98]. Mesoscale [Véz94, BHAV91, RB97]. mesotrophic [BLD99, BS99, GJ96, HNS95, LAC^{+}96a, LAC^{+}96b, QLBAP96]. Mesozooplankton [CL99, GG99]. Metabolic [CP97a, DL92b, CT93, CSS96, GJ90, Gla91, HB95b, TC93]. Metabolism [FSGG93, GHT^{+}96, ASC94, BDV99, BOCL^{+}95, CTF98, CNSR90, HK98a, HAL90, KY97a, Lan93, MP92a, MMW^{+}97, MMW^{+}99, MMSKH94, WHB95, YK93]. metabolism-influenced [KY97a]. metabolites [BBH^{+}92]. Metal [BdBNvB91, LvGLC98, SVC^{+}98, AMM97, BW94a, DL96, HTSDZ99, HCHD94, HGFC96, NLSG99, WBKF95]. Metal-rich [SVC^{+}98]. metallothionein [CCT93]. metals [AM95, BMP92a, BDH91, CBS99, GAAG96a, HTW90, HT98, MDC95, MHP91, SH98c, TC95b, TCA96]. metazoan [Str97, UNK95]. metazoans [SE90]. meter [BM92a, Boy94]. meters [AK97, Gli98]. Methane [HCS95, JM98, KMU95, YAF^{+}92, dAL94, FLH96, HC94, JCM^{+}99, MSM96, RW96, RC98, SC93a, SDS95, SW96b, SM98c, UNN^{+}98a, UNN^{+}98b, dABL91]. Methanogenic [Ano93t]. method [AHKH^{+}96, ABB^{+}95, Ano92d, Ano93o, Ano93u, BH96, CCK98, CI94, CdG92, GW93, HHB^{+}92, Jon91a, KT92, KMWK92, KJLA98, KJS98, LHS90a, LVFD93, MPA94, MCCD92, MSH96a, MHM^{+}99, Muc91, PF90, Pol95, RS91, RC93, Rut92, SCK93, TLB99, TBK98, TB96, WWT^{+}93b, Wym92]. Methodological [PZ93]. Methodology [MS97, DM90]. Methods [FHS98, Ami90, BBH^{+}92, BRZ95, CH91, Duc94, HB95a, Kap92, LW93, LHM^{+}93, MR99]. Methyl [MdlC97, TM95a, MGN92, REL97, SM99a, SM98a]. methylation [CGHK99]. methylene [MGN92]. methylmercury [GPMF96, HLGA98, WB92]. methylsulfide [Ano93s]. Metridia [BOFW93, Hay95, HWP95]. Mexican [GMM96]. Mexico [JCH92, JRT96, KCC98, KSH^{+}99, RMT96, SBB99]. Meybeck [Edm96b]. Meyer [Edm96b]. Michener [Rav95]. Michigan [AR95, BL91, BE94, CFSW91, CFT92, FCI91, FG93, LS90, LC93, MCB90, SW96b, SWSC90, VGP^{+}92]. microalga [SM99a]. microalgae [AK92, Som97]. Microalgal [LPJ92, AS94, BC90, CL90, HS99, LC90a, LTPB98, LL98, PZ93]. Microautoradiography [BW94a]. microbenthic [KJ92, KJ94]. microbes [WH95]. Microbial [Ano93z, JvROG99, LD96, LVFD93, MD90, Tay95, WR95, AV94b, Ano93c, BVSBB93, BKR^{+}90, BAH97, BBH^{+}92, BdBC99, CL99, CL96, DL93, Duc94, ES93, EKT99, FCC99, GS93, HB95b, JD90, Ker93, KB92, LHWZ92, MM97b, MS97, MWW97, NVH^{+}94, PBJD93, PFB96, PH92a, RHH97, RG92a, RW96, SCPS98, SM97, SH90, TCA96, UPP95, WMPC^{+}90, Wic95, YD99, ZR98, dABL91, dARB91, dBGEK97, Dua95, PA95b]. microbialloop [THR97]. Microbially [LF93, MAS91]. Microbiological [Fie97, PWM^{+}98, DHP^{+}96]. Microbiology [Mel99, Fen94, Amm96]. microcapsules [KHB96]. microchemistry [TJC97]. Microcoleus [DCM90]. Microconsumer [EF95]. microcosm [CWSX96, GGJ97]. microcosm-cultured [CWSX96]. microcosms [NFGP^{+}93, WF90]. microcystin [OJ98]. microcystins [Jun92]. Microcystis [Jun92, FSSI97, OJ98, RKW94, WH99]. Microdistribution [PAS90]. microelectrode [AD92, LBL^{+}98, dBGEK97]. microelectrodes [CR93a, GGRJ94]. microencapsulation [CLP^{+}92]. microenvironments [DL93]. microfauna [SA94]. microfossil [SWSC90]. microgradients [JD90]. microhabitat [CWSX96, RERB91]. microheterotrophs [PFB96]. Microheterotrophy [FPAG^{+}98]. Micromonas [CS95]. micronutrient [Ste94]. Microorganisms [VF94, CPSA90, KW98a, KJS98]. microparticles [LJA91]. microphytes [RPRNR94]. microphytobenthic [DY99, Jön91b]. microphytobenthos [dJvB95]. microplankton [BUW96, KHP91, MOF90, dGF96]. Micropogonias [TJC97]. microprobe [Dod92a, KJ92]. microprofiles [EKT99]. Microscale [MNAH96, PBJD93, CW93a, KJ94]. microscopic [Edm96a, HC97]. microscopy [Ano93o, HS95, SBW^{+}98, XB97]. microsensor [NCRS90, RNR98]. microsensors [GRG98, KMK95, KHK97]. microstructure [Old94]. microtidal [SAG98]. microtopographical [Joh94]. Microwave [TLB99, XB97, Bar99]. microzones [PPK95]. microzooplankton [BN95, JWRS96, Lan93, ŠKN^{+}99, SW91, WW95]. mid [BL97, PRCV98, RYC99, TV99]. mid-Atlantic [BL97, PRCV98, RYC99, TV99]. midcontinent [LWS^{+}96]. Middelburg [NRPRB96]. Middle [GSC^{+}96, HC96, VBD^{+}97]. midge [Hur90]. Midland [MMC94]. midlatitude [RLH95]. midocean [NB97]. Midwater [SSP^{+}94]. Migration [MCED94, PAS90, DFN97, DPC90, DL92a, DL92b, Dod90, FH92, FR99, GWY^{+}92, HPJW94, Hay95, HWP95, KY97a, KBB98, KM95, LW93, Loo93, MOF90, NRR96, RF94c, Tjo90, WS90, WKK91, WSMS96, YW93]. migrations [HCKW90, KSH^{+}99]. Migratory [Ano93l, DC91a]. mild [BC97]. Millero [Arc93]. Mills [Ban90c]. mine [NML99, SL98, TMB96]. mineral [Hes93, Kir99, MN95]. mineralization [AEL^{+}99, FFJ92, dHK98]. minerals [WHK^{+}93]. mini [DMF^{+}98]. minimize [Coa90]. Minimum [Bra91, BTL^{+}95]. mining [KLR94]. Minireview [MB98a]. Minnesota [SM98c]. Minor [FCM^{+}97, FMH^{+}91]. Minshall [Wal96]. missing [BG94b]. Mississippi [GCCJ97]. Mitchell [Ban90a]. mitigating [GDG99]. mitotic [Ano92d]. mixed [Bou98, CP94, CP99, ES93, HHI96, LB91, Som98]. mixed-algal [LB91]. mixing [AS91, BDY^{+}98, CE98, CC97, DOC99, FHKC96, JM93b, LCR99, LBL91, LP95, Mac93b, MFJR99, ML95, MA96, NS91, OYWK96, PC97, RWH^{+}98, RM96, TJ98, Ver99, WBRJ91, ZCGP98]. mixotrophic [CPS90, JBJB96, LPPSR97, SPB90]. Mn [HTDM96, HTSDZ99, LBL^{+}98, SH98b, UGU^{+}90, YH92]. moat [MA96]. mobile [LSS^{+}97, TR98]. mode [MWI92]. mode1 [SBB^{+}93b]. Model [Ban92, And93, Arm94, Arm99, AS94, BM92b, BS90c, CKR93, CR96, CE98, CGVL95, DK93a, Eck90, EBH92, GP94, GMK96, GMK98, GC90, HCHD94, HL94, HB94, IH95, JUSI91, KB92, KZ92, LC90a, MM91, MKT^{+}90, NSZ97, OMk93, RGH^{+}93, SP90b, SHM96, SCF99, THK^{+}99, TL94, VWF96, VGB94, Vos92, WHM97, Wil92, WLD93, WB96, vdBG94]. Modeling [AW93, Ano93r, FBS99, GCP90, HRD^{+}95, KCWM96, McL97, TCCA93, VHH96, Ver94, FAE99, GBC95, Guy91, MS97, MWW97, WWW98, WB96, WBKF95]. Modelling [WTH98]. Models [Fra97a, Ano93p, Ban90a, BF97a, CL90, DC96a, FDLD93, McC95, PH92b, Sar99, Sch92, WP94]. Modern [Cha96, DHP^{+}96]. modes [RSZ97]. Modification [LS92]. modified [Dod92a, FR99]. Molecular [CCR^{+}95, PL99, PSS^{+}96, SZA92, BLS^{+}95, GSC^{+}96, Har95, NML95, Plu97, SKR^{+}91]. molecular-weight [NML95]. molecules [DBZF98]. molluscs [RC93, WSNM98, FRLS90]. molt [ER95]. molybdate [LB95]. Molybdenum [CLMH93, MHSP90, EP97, MMS93, PPB91]. molybdenum-dependent [PPB91]. momentum [DGHD98]. monitoring [Abb90]. mono [LB91, JM93a, JM93b, JMMD93, JAMH96, JRM98, JL94, LKSA97, MFJR99, MJOC93]. mono- [LB91]. Monod [VHH96]. monomictic [HGS94, PSH^{+}96]. Monoporeia [HQCG92, JW92]. monosaccharide [GHT^{+}96]. monosaccharides [BS99, RDK96]. monoxide [SC93a]. Montana [LS93]. montane [MA94, NML99]. Monte [Gor91, Kir94b]. Monterey [BKB97, BW99b, BPC^{+}97, SSP^{+}94]. Mooers [War90]. Moon [LFGM96]. mooring [DHWW97, TJ98]. mooring-based [DHWW97]. Morphological [Ham94, LV97, Bla93, BC90]. morphologically [HSL^{+}97, RWH^{+}98]. morphologies [AML93, BH97]. morphology [HPJW94, VLHL97]. morphometry [BK95]. morphs [Gle93]. Mortality [OW96, AO96, FN95, GG99, OAR96, PV94, PK94, Sam91, Sve97, ZCGP98]. mosaic [HDZB98]. mosquito [BK90]. mosquitofish [BK90]. moss [NSBB98]. mosshead [AMP98]. motility [NE93]. Motion [KBJ98, Sai94, SWH92, TG94]. motions [MIM90, RS98]. Mougeotia [GAAG96a, GAAG96b]. moulting [VM94]. Mount [WMD90]. mountain [BDV99, FCC99, HPM96, KPSB95, KHS^{+}98, TBM95, Hur90]. mountainous [KL96]. Mountains [WLCG96, RKA96]. Movement [BO96, MPT99]. movements [HSI95]. MT [YK95]. MT-5 [YK95]. mucopolysaccharides [JvROG99]. mucus [FR99, WMTB93b]. mudflat [Ker93, TGG94]. muds [AA92]. multi [MC97]. multi-excitation [MC97]. multicellular [Hof90]. multilevel [WF90]. Multiple [DL93, FMK^{+}99, HC99, WT94, Arm94, BLK95, CW93b, FCMB99, LK94a, LMSP99, PLP99]. multipurpose [Edm96c]. Multiscale [SD95]. Musci [NSBB98]. mussel [BFGS94, WF96]. mussels [PY99, RF99, TV99, WW98]. mutually [HRF91]. Mya [EG91b]. mycosporine [SRLRL^{+}99]. mycosporine-like [SRLRL^{+}99]. Mysis [GRG99, MCB90]. myth [DFM92]. Mytilus [BFGS94, WF96].

N [Bre93, BZR99, FZP92, Ohm96, Sal92b, War90, Alt90, AR96, BG94b, BLK95, CIM92, CPSA90, FPR94, FVL98, FVL99, FSSI97, GHF^{+}97, HCSE97, HDR^{+}98, LS92, LLB^{+}96, ME98a, MDE97, MB90a, NVH^{+}94, PVL^{+}96, SML96, SRG98, SRG00, UW92, VM94, WH94b]. N-acetylglucosamine [NVH^{+}94]. N-labeled [SRG98]. N-nitrogen [WH94b]. naked [And99, Boo94]. Namibia [FFNS99]. Nannochloris [DO98]. nanoflagellate [DŠ98, SBGBC96b, TC95b]. nanoflagellates [DŠ99, GV93, JBJB96, JD95, JWRS96, MLB99, SPB90, SBGBC96a]. nanomolar [Cam94, HHI96, Jon91a]. nanoplankton [CPLD99, Ver91, VRT^{+}92]. nanoplanktonic [FR91]. Nanoscale [KJLA98, KJS98]. Narragansett [BN90, DDC92, DD92a]. Narrow [SSFP90]. natal [TJC^{+}98]. natans [Lap95]. National [KSH^{+}99, Thr91, Wet91c, Ano90h, Cot99]. Natural [CM94, Kir94a, KWF90, LAYE95, LH91, LP94, WHB95, BTLC93, Bru92, BW90, BS90c, CI94, CRKS93, GMM96, GCCJ97, GHT^{+}96, GB94, HS95, HSRR96, HV98, Jon91a, KZM90, Kro93, KNM96, LV98, LF91, LK94a, MHIHH99, MJ94, Mor90, NLSG99, SMB^{+}97, SMC98, SMH94, VBAO95, WK96, ZCGP98]. nature [MHT99]. Naupliar [Har96, GH95]. navigation [FPZFW99]. Neagh [Edm96c]. Near [Dad93, Ano93y, AR95, CR93a, FHF99, GSG94, LM98a, NB97, NSBB98, RJC^{+}97, SKB94, WPBS99, YPF^{+}98]. Near-bed [Dad93, FHF99, NSBB98]. near-bottom [SKB94]. near-infrared [Ano93y]. near-surface [GSG94, LM98a]. nearshore [CL96, GAA98, HLL98, MGN92, MF99, MDWS97, Wel99b, Wri90]. necessary [AWKH98]. Necessity [LPA91]. Need [Wet91c, JC95]. needle [VBvG91]. needs [Har96]. negative [GA96, WWMI98]. neighbor [Joh90]. Nematonums [PRS94]. neritic [Lap95]. Neshyba [War90]. Net [Fur91, BW99b, IH95, LSB92, Sti91, dGGV^{+}96]. Netherlands [SBGCC91]. nets [Lou90]. network [CE96, GBC95, PTJ96]. networks [BLD99]. neural [BLD99, CE96, PTJ96]. Neuse [MB90a]. neutral [GHT^{+}96, Hen98, RDK96]. neutrally [PPK^{+}96]. Nevada [WBM93, WB96, JGRR99, MTLR98]. newmani [OW96]. NH [HNS95, HA92]. Ni [BW94a]. nickel [PM91]. Nielsen [MSH96b]. nighttime [JB91a, JB91b]. Niño [ARM96, CR99, Nic98]. Niño-related [ARM96]. Nitrate [AR96, BH95, CIM92, DW95a, EK90, HHI96, PAD^{+}99, Arm99, BDY^{+}98, CNSR90, CHD91, DK96, DW91, Gar91b, JA92, LLZ99, LG99, PVL^{+}96, SGC^{+}97, Som94, VMDC96, WHN^{+}98, WK90]. nitrate-rich [CNSR90, CHD91]. nitric [RKST90]. Nitrification [DK96, RPJ97, SNB92, WWD90, BN90, EW99, KMWK92, RPRNR94, RRPP^{+}94]. nitrite [DK96, WHN^{+}98]. nitrobenzaldehyde [MJC93]. Nitrogen [Ano93v, FZD93, GCB^{+}98, HR90, KFW94, MVM97, MGWB97, SVW^{+}97, Smi90b, WHN^{+}98, WH94c, YK93, AH91, BF98, BBS97, BB97, BDY^{+}98, BW99b, BZ98, CNP94, Car98, CPS90, CE96, CHD91, CMPD92, CNL90, CK92b, DM92, FRY95, GCCJ97, GGFR91, HHI96, HCH93, HD97, HKPM99, HG99, HSM^{+}95, HBL^{+}93, KPSB95, KHS^{+}98, KHB96, LD99, LKM98, MHSP90, MV98, MA94, MSH96a, MPG95, MBHT94, OJ98, PPK95, PFB96, Pae97, PH97, PPB91, PBL99a, PP97, PM91, Pro90, RS91, RSK97, RMT96, Sar92, SS99b, SR92, SRG98, SRG00, SZA92, SM91, SCEG95, TR98, TC94, UMS^{+}99, UNK95, VRT^{+}92, Véz94, VDA99, VVIK99, WG97, WP94, WH94b, WPBS99, YK94, YKM95, YK95, JMMD93, Ano93f]. Nitrogen-fixing [HR90, MHSP90]. nitrogen-limited [OJ98, SM91]. Nitrogen-stable [MVM97]. Nitrogenase [DCM90, CHA91]. nitrogenfixation [HRF91]. Nitrous [Ano93w, MGW96, DK96, PDM96]. Nitzschia [SGA98, SMB^{+}97]. nl [MNAH96]. NMR [HGR95]. NO [KMWK92, HDMB97, FAE99, KMWK92]. Noctiluca [SW96a, TK98]. Nocturnal [SDS95, WKK91]. nodosa [MCED94]. Nojiri [UNN^{+}98a]. Non [RFHG99]. Non-indigenous [RFHG99]. nonalgal [BS90b, SWS^{+}96, VFAA98]. Nonconcordance [Wil90]. nonconsumptive [GG99]. nondestructive [TKAD95, TB96]. nonhumic [MH90]. nonlittoral [Boy94]. nonnutrient [Kam92]. nonsaline [Cam94]. nonviable [BV95]. normal [JF96, JF99]. Norse [But95]. North [DHWW97, Fri90, MB90a, RKG97, DHK^{+}96, HK98b, MS93a, MSM96, SHF^{+}96, SM98c, BKT96, BKH^{+}96, BG94a, BDY^{+}98, Bru92, CNV94, CR96, DM98, DHB93, DW93, EK90, Fri96, HH93, HS92, KY97b, KHBL98, KMC90, Lap95, Liv99, OGH^{+}96, OCZA90, PRS94, RKA96, Ree90, SMV98, Smi92, SKB94, TBK98, UA97, Ven93, Was90, WG94, WBP99, WSM^{+}96, WWMI98, WL94]. north-central [SHF^{+}96, SM98c]. north-temperate [DHK^{+}96, MSM96]. northeast [CR99, FGN93, HWP95, HP95, RM92, WWMI98, FBS99]. northeastern [And97, BUW96, SHKP96, TBM90]. northern [CDP96, FLH96, GCD90, HLQ^{+}99, JCH92, JRT96, LFGM96, MMSKH94, MHSW99, PJSW96, TZGS98, VRHK99, WMHB95, WKB^{+}96, WBRL98]. northwest [TZR98, VTCB96, YFF^{+}98]. northwestern [ASMB98, JLMV98, MCC^{+}96, SBP^{+}96, VPN^{+}90]. norvegica [TJK^{+}97]. Norwegian [GA96, JD96a, JD96b, MHIHH99]. note [DW95b, NRPRB96]. Notonecta [Bla93]. Nova [UGU^{+}90]. novel [HTSDZ99, LG99, PM90a, RNR98, RBMH91, SHB99, TKAD95, TBK98, VBvG91]. nucleic [dGBPP96]. nucleoid [CSS96]. nucleotidase [AA91a, AA91b]. null [Ano93p]. Numerical [Guy91, LJBTB99, RLH95, VWF96]. Nürnberg [CKCP99]. nursery [HC96]. nutricline [SH95c, WKK^{+}95]. Nutrient [Bre93, Dua92, ESC^{+}93, FSSI97, JBJB96, KSH^{+}99, KKK^{+}93, LBLO91, Lap97, LB94, ME98a, Mat93, ML92, PL91, Pra90, Som89, ADC90, Arm94, Bac99, BA95, CP94, Cul91, DC96a, DY99, DW91, EF95, Els98, NFGP^{+}93, Fro91, Gar91a, HK98b, HS99, JL92, Lap95, LRM95, LDV^{+}92, MFJR99, MM97a, MJC^{+}98, MB99, Mul92, NT93b, Pae97, PRWG97, PSJ99, RA97, Rot95, SO91, SH95a, Smi90a, SHSS93, SB90, TIOS99, VLHL97, VT90, WR95, WMD97, WPBS99, Sau92]. nutrient-deficient [VLHL97]. nutrient-depletion [EF95]. nutrient-depletion/dilution-gradient [EF95]. nutrient-enriched [Bac99, SB90]. Nutrient-limited [FSSI97, Lap95, SB90]. nutrient-rich [Fro91, Gar91a, LDV^{+}92]. Nutrients [JCMW91, BSG^{+}99, DHB93, EF95, FDG99, GMK98, Hes93, MLLR99, MHM^{+}99, TIOS99, VL98, VVIK99, WKK91, WMPO99]. Nutrition [KBH98, GP99, Hes93, MT90, TV99]. nutritional [Ano93m]. NW [BG97b]. Nyos [NKT^{+}93].

O [MZ90, QWR^{+}95, CH91, MJC93, DHWW97, GRG98, HB94, LBL^{+}98, QWR^{+}95, RA98]. o-nitrobenzaldehyde [MJC93]. o-phthaldialdehyde [CH91]. Oak [KMC90]. Oblea [SB93]. Observation [GBC95, Ree90]. Observations [CC97, Ken93, MJG91, MWI92, OC97, RLH95, RJC^{+}97, Sai94, Tho99, App94, BM92b, ZKM93]. observed [AK97, RSZ97]. obtaining [KC92]. obtusa [Mac93a, SC93b]. Occurrence [GS93, SVF91]. occurring [CL91]. Ocean [Cul95, LM98a, Tho90, AGGD90, BW91, BSMK98, BSC93, BRZ95, CBBC96, CL99, Cha96, DT91, DOC99, EA99, FC91, GSG94, GW93, GWMB91, HL94, HVB93, HP95, Li98, MS99, MR99, OB98, PB94, SO91, TIOS99, Vos92, WCD^{+}94, War90, Whe92, CNV94, DHB93, HP95, KY97b, KHBL98, Lap95, MBHH^{+}91, NS91, Nod97, OGH^{+}96, OAF^{+}99, OBA99, PAM94, PRS94, RGR^{+}98, RM92, RB97, SO91, SKB94, Smi96, SW91, SHKI95, TBK98, VDA99, WWMI98, WL94, May92, Str94]. Oceanic [FPC91, HS90, TBM90, BPCL97, BDH91, Cla94, Con97, DF94, GSW95, HHI96, Lap95, MMG94b, MSSM96, MS97, Mor97, PSS91, RA97, SM97, SH95c, TIM96, WK90]. oceanica [MD97]. oceanographic [CK95]. Oceanography [Ano90a, Ano90h, Ano90i, Ano91f, Ano94, Ban90c, Hol99, Jum94, MDWS97, Pla93, Sch92, Ano90f, Ano91d, Arc93]. Oceans [Mat94, BPK94, Bar99, CBFC98, HCSE97, HCH93, TKL^{+}97, OCZA90, WG94, Sal92b]. oculata [BPA96]. Odor [FPZFW99]. off [BG97b, DHWW97, Goe98, PPK^{+}99, RKG97, TC96]. off-shelf [BG97b]. offshore [Ban90d]. Ogeechee [LM91]. Oglethorpe [PSH^{+}96]. Oikopleura [ADM96, BDR99]. O’Keefe [Mei90]. Okuda [Spi96]. oligo [NB97, QLBAP96]. oligo-mesotrophic [QLBAP96]. oligo-trophic [NB97]. oligohaline [SHB99]. oligonucleotide [LDC99]. Oligosaccharide [AR94]. oligotrich [OS91]. Oligotrichida [DŠ97]. oligotrophic [AR96, BVSBB93, BW90, CL99, CCL92, CSPG94, CBFC98, CCL90, CC98, CW95, Duc98, Fur91, KY97b, KHBL98, LD96, PH92a, SCEG95, TBK98]. oligotrophication [MR98]. omnivorous [CGD90, OS91]. Omnivory [OR94, JOG96]. one [CR96]. one-dimensional [CR96]. only [BC97]. Onondaga [Kna99]. ONR [Ano93a, Ano95c]. onset [GDG99, JM93b, JRM98]. onshore [AB99]. Ontario [BM96, HSFR98, HS98, SH91, SBP^{+}96]. onto [Eck90]. Ontogcnetic [WS90]. Ontogenetic [Hay95, TV99]. ontogeny [MDE97]. Open [Cul95, CL99, Cul91, EA99, FDG99, Fro91, GW93, SBPV99, SW91]. Open-Ocean [Cul95, EA99]. Ophrydium [SJPGH97]. opportunistic [MT90]. opportunity [Thr91]. Opposing [NRR96]. optic [Dod92a, KMK95, KJ92]. Optical [AE91, CCDL97, GB94, Ano93x, AS94, BHAV91, BRZ95, DSV^{+}99, DO98, FAE99, FM94, FP99, GD92, HS92, HVB93, KZ92, MaB98, McC92, Mor97, SJHS98, SS99d, SR93, SM97, SCKF99, SCF99, BKT96]. Optical-digital [AE91]. optics [BKH^{+}96, KJ94, MS97, SM97, Str94]. Optimal [Wil92, HS99]. Optimization [VHH96, Arm99]. optimization-based [Arm99]. Orbulina [UMS^{+}99]. orders [RRK94]. Oregon [CC97, DCML96, MCCD92, MCD^{+}96]. Organic [Alt90, CDHS98, HLL98, JAMH96, MMSKH94, RKR91, RRK94, RW96, dHK98, AEL^{+}99, AV94a, AV94b, AB96, BP91, BB98, BTM92, BBS97, BLS^{+}95, BB97, BGN98, BdBC99, Bru92, BZ98, BG97b, CCR^{+}95, CCKP98, CKCP99, CTSB98, CCC^{+}96, CE96, CW92, CIM92, Cur93, CMN93, DSV^{+}99, DM98, DMHH99, ECA99, FSGG93, FL95, GPMF96, GREGG97, GGFR91, GB94, GS93, GSL97, GH95, GSW95, GSC^{+}96, GW94, HTDM96, HCR^{+}94, HS98, HVB93, HCLD95, KL96, Kap92, Kap94, KY97b, KHBL98, KAB92, KAD95, Kro93, KNM96, LS90, LFHS99, LM91, LDL91, LGT95, MM93, MDC95, MBK^{+}99, MTLR98, MM97b, MR99, MH90, MPS^{+}91, MZ97, MZW^{+}95, MH97, NLSG99, NZHB95, ORH99, OBA99, OBL^{+}91, PBJD93, PH97, PP97, PSS^{+}96, PHX92, QLBAP96, RGR^{+}98, RS91, RM92, RB97, SBW^{+}98, SH95a, SML96]. organic [SS99b, SBB^{+}93a, SB97, SRG98, SRG00, SHM95, Sti91, SBZD97, SPMW97, SIO98, THK^{+}99, Tra94, UMS^{+}99, VHH96, VDA99, VL98, VGB94, VVK97, WS90, WCW^{+}91, WHB95, WMPO99, WMD90, YK95, ZWH^{+}95, dGP94, Bur94]. organic-rich [GPMF96]. Organism [CACM98, WTH98]. organisms [AML93, ADS94, Ant99, CH91, DGHD98, GP94, KHB96, MF90, TCA96, VSBZ^{+}94]. Organization [WT93, BKB97, Grü97, TW93]. organized [VFW^{+}92]. orientation [KY97a]. oriented [KBB98]. Origin [SK95, LO97, MH94b, SLSLBD97, WHM97]. Origins [HCR^{+}94, JGRR99, YFF^{+}98]. orthophosphate [IH95, Jan93]. Örträsket [JBJB96]. Oscillation [Liv99, WBP99]. oscillations [AEL^{+}99, JW92]. oscillatory [CHA91]. Otago [MJG91]. other [BG97a, Cur94, HSCM97, KHK97, LW90, MHZ98, NML95, Pae97, PH97, RJ91, WMHB95]. otolith [TJC97]. otoliths [TJC^{+}98]. otters [KDB91]. our [Pal92]. outbreaks [AWKH98]. outburst [WBP99]. Overbeck [Dua95, PA95b]. overestimated [FRL94]. overlooked [CS97, SBPV99]. overlying [BW94b, WTH98, YK94]. overturn [MJOC93]. Overwintering [Dal90]. ovipositing [Ber99a]. Oxic [AEL^{+}99, CP95a, SLA93, TCCA93]. Oxic/anoxic [AEL^{+}99]. Oxidation [JCM^{+}99, LKD99, UNN^{+}98b, CPT^{+}94, Hin94, HS90, Kap92, KLB^{+}99, LF93, MM93, Mil91, MB98b, Mof94, PL94, RS91, RW96, SBB^{+}93a, SH90, TC96, UNN^{+}98a, dABL91, dARB91]. oxidative [GLT96, Les96]. oxide [Alt90, Ano93w, DK96, KLB^{+}99, MGW96, NLSG99, PDM96, RW96, TBM95]. oxides [HTSDZ99]. Oxygen [CPMW99, GG92, GCD90, JD90, RRPP^{+}94, SS98, SBGCC91, Ami90, Ano95e, AD92, Ben94, BW94b, CSPG94, Car98, CNSR90, EKT99, FK98, FL90, GJ90, GA96, HD95, HB94, JD96a, JD96b, Kan93, KG95, KMK95, KSYK96, MCD^{+}96, Old94, QWR^{+}95, Rah90, RHH97, RCCdG99, RGRW93, SSKW91, SGB^{+}97, SFW^{+}95, SHF^{+}96, Sti91, SS99d, VBvG91, WMM^{+}94].

Oxygen-consuming [SBGCC91]. oxygenated [MGW96, WMM^{+}94]. oxyhydroxides [DBT91]. oyster [TZFP^{+}94]. oysters [LPA96, LMSP99].

P [Bre93, CT96, Cha96, Duc94, Edm96b, Leh95a, Man94, Man95, Pla93, Sut93b, BZR99, CP97b, FZP92, FSSI97, GHF^{+}97, HCSE97, HGR95, IH95, LBL91, Leh93, LS92, LLB^{+}96, MDE97, MB98b, QP91, TZR98, THK^{+}99, UW92, ZR98, dGP94]. P-limited [THK^{+}99]. P-uptake [IH95]. P.O [WL99]. Pacific [CNV94, DHB93, KY97b, KHBL98, PRS94, SKB94, TBK98, Ban90d, BC91, BDY^{+}98, Bru92, CLKJ95, CBMC^{+}99, CBC^{+}91, CJVD97, Coa91, CR99, DW93, FPC91, FGN93, GMM96, GCJ97, GKS^{+}94, HSM^{+}98, HP95, KKW90, LBB^{+}97, LLLB97, LSS^{+}97, MP92a, MFW^{+}91, MBKG95, Nod97, OHMK91, OCZA90, PAM94, Ree90, RDK96, RM92, RG97, RB97, Smi92, SW91, SHKI95, TBM90, Ven93, WPB94, WGSP91, WK90, WWMI98, YFF^{+}98]. pacificus [BOFW93, CH96, DFN97, JOG96, OC97, PF91]. packaging [BPCL97]. paddy [RC98]. Paerl [HSM^{+}95]. paired [JEP^{+}99]. paired-lake [JEP^{+}99]. pairing [MSH96a]. Palau [HSM^{+}98, LBLO91, LLB^{+}96, OBL^{+}91]. paleoclimate [SS98]. paleoclimatic [Ben94, LFGM96]. paleoenvironmental [BFG94]. Paleolimnological [Fri96, KLR94, SWS93]. Paleolimnology [HPB^{+}99, KRW99, LC90b, LFHS99]. pallida [CMPD92]. Pamlico [MB90a]. Pantanal [HSCM97]. paradigm [Leh95b, WMPO99]. paradox [Hur96]. paradoxical [Sie98]. Paraguay [HSCM97]. Paralytic [KDB91]. parameter [HM99]. parameters [DA92, FDLD93, MMG94a, WB96]. Paraná [CNP94, CV99]. Parani [CP94, DK93b]. Paraphysomonas [CGD90, LDC99]. parasites [BLR92, LK99, LMSP99, SC93b]. Parent [VMDC96]. Pareuchaeta [TJK^{+}97]. Pariente [Thi94]. Parmenter [Sau92]. Parsons [Jum94]. Partial [KY97b, Car98]. Particle [ADM96, BM99, Lou90, WM92, AGGD90, App94, CDPR98, DDG94, DCML96, GW99, HN94, MHSW99, Rot90a, Rot90b, SM97, WSNM98, Whe92]. particles [AML93, ABB^{+}95, Ano93j, BW94a, CCK98, CDPR98, CW93b, DGW92, GSL97, HTSDZ99, HSM^{+}98, JF96, JF99, LSS^{+}97, MO91, MS97, PA95a, QWR^{+}92, SM97, TF95, WCW^{+}91, WHM97, YFF^{+}98, ZMP98]. Particulate [KL96, Alt90, BB97, BL94a, CMN93, DASV98, GSG94, HCH93, HZF96, Jac93, MBK^{+}99, RS91, Roe98, Sar99, Sha94, SRG98, Sos99, TG92, TIM96, WS90]. Partitioning [DASV98, MS99, Bur92, CDHS98, RG92b, SFC90, TCCA93, WBKF95]. pass [TIOS99]. Passive [Bar99, Eck90, HBI95]. past [MTLR98]. patches [SGP92, SGP96]. patchiness [KJLA98, STJ^{+}93, Ver94]. Patchy [MPT99, Tis92]. pathway [JM98, SRLRL^{+}99]. pathway-dependent [SRLRL^{+}99]. Pathways [Rom91, TC96]. pattern [GW94, Man95, SM92]. Patterns [BP91, CCB^{+}99, Dua92, DAC92, Hur90, Joh91, WMD97, dGP94, AML93, AE91, BKB97, CJVD97, De 90, Dod92a, DO98, Fra97b, Han96, HWP95, HPTC97, LWS^{+}96, MWW97, Ste94, VT90, WH93b, dGF96]. pause [MK95]. pause-travel [MK95]. Pavlova [CL90]. Payne [Thr90]. Pb [BK99, BPS94, LSS^{+}97, NWS94]. Pb-dated [NWS94]. PCBs [JEP^{+}99]. PCC [WMT96]. PCC786 [Jun92]. pCO [CR93a, CI94, DHSS95]. PCR [CM95]. PCR-RAPD [CM95]. peat [NWS94]. peatland [FLH96]. peatlands [SW96b]. peaty [De 93]. pedogenic [WJRB97]. peeper [WFG99]. Pelagic [CCC^{+}96, Els98, GJ96, JW92, YD99, ADM96, BGC99, Cur94, FG93, Fra95, HNS95, HC96, HCSE97, JB91a, JB91b, KC92, LAC^{+}96a, LAC^{+}96b, MR98, MB98a, MF90, PPK95, PCC^{+}97, Pro90, QP91, SKB94, SSP^{+}94, TC93, TT97b, TCA96, YK94, YSSO91, YMT^{+}99, ZEG^{+}94]. Pelagic-benthic [GJ96, FG93]. pellet [BN95]. pellets [VRHK99, WMML98]. penaeid [BLK95]. Penaeidae [CBS99]. Penaeus [CBS99]. Penetration [KBMSP99, And93, AMLH97, De 93, KHK97, Pic91]. people [Hai96a]. peptides [DBZF98]. perception [Hai90]. perched [ML96]. perchloroethylene [AECP95]. peregra [Bre97]. perennially [HHJ98, LP94, TCB^{+}98]. perfoliatus [CK92b]. performance [DWM90]. Peridinium [BFE96, ZPHH98]. period [JAMH96]. Periodic [JRH90]. periodically [ES93]. periodicity [CHD91, DŠ99, Gra91]. periphytic [Han92]. periphyton [AR96, BH91, CGHK99, Dod92a, HB91, HDMB97, NML99, SG93, VL96]. peritrich [WLY93]. permanent [FP99, HJ94]. permanently [Ano93z]. permeable [HZF96]. peroxide [AOTR97, MZ90, PM90b, SML96]. Perry [Str94]. persistence [JM93b, KW93, TSV^{+}91]. perspective [HC99, Mei90]. perspectives [Thr90]. persulfate [Kap92]. perturbations [Ano93k]. perturbed [LHS90b]. Peru [BS90b]. pesticide [HD95]. Peter [Ano93d]. Peters [BCBC93, Lew95]. Pfiesreria [BG97a]. Pfiesreria-like [BG97a]. Pfiesteria [BG97a]. pH [CR93a, CV99, FRY95, Gen91, GAAG96a, JA92, KRB98, KPSB95, KNM96, LFHS99, Sha94, TC94, THS^{+}91, VKB96]. Phaeocystis [PSEJ99, PSJ99, CSM94, JvROG99, MWB^{+}98, MM99]. Phaeodactylum [LBP97]. Phagotrophy [Rav97]. phase [HW91, LVM^{+}90, SGWS96]. phases [NLSG99]. phenolic [HSRR96]. phenomena [SGC^{+}97]. phenomenon [GV93]. Phenotypic [Lam94, Bla93, LB96]. phenotypically [MW98b]. pheophorbide [HHR94]. pheopigments [Wel94]. Phormidium [FSSI97]. phoronids [Joh90]. phosphatase [BW90, SMH94]. phosphatase-hydrolyzable [SMH94]. Phosphate [Bur92, Leb90, TBM95, AB92, Bal96, BA92, CV99, CPT^{+}94, Hon97, Jan93, JA92, LB94, RG92b, Sha94, SCK93, SMV98, SFC90, TMB96, dHJS90]. phosphate-starved [Jan93]. phosphohydrolytic [SIO98]. phosphorous [VVIK99]. Phosphorus [CV99, DG99, FZP92, JMA^{+}95, KKBG91, MS99, Per97, Som92, SM99b, UCS97, VAH99, vEJ97, AA91b, AH91, AZS^{+}98, BTM92, BE94, BL94b, BW90, Cam94, CCL92, CPS90, CCKP98, CKCP99, CCL90, CH90, CGB90, CH92b, Cur90, DGS98, DM92, ES90, EP97, FZD93, HCH93, HKPM99, HG99, HT96, IPPD92, JB91a, JB91b, JMMH98, KT92, KY97b, KPSB95, LCR99, LAC^{+}96a, MR99, MMW^{+}97, MMW^{+}99, Nic98, Nür98, RM92, RWH^{+}98, Rot92, Rut92, SJDJ94, Sar92, SVW^{+}97, SS99a, SSFP90, SMH94, Smi90b, SR94, SCEG95, SGSA92, SIO98, TLB99, TBK98, UNK95, VORJ93, VBAO95, VDA99, WH92, WG97, WP94, WVP97]. phosphorus-deficient [DGS98]. Phosphorus-limited [Som92, Rot92]. Photo [GLT96]. Photo-oxidative [GLT96]. photoadaptation [GMK96, SFRB98]. Photoadaptive [Mol98]. photoautotrophic [FRL94]. Photobiological [NS93]. Photochemical [BB98, OB98, UA97, KYM97, LGT95, MM97b, MH97, RHH97]. Photochemically [BL98]. photodamage [MC96]. photodegradation [De 93, KZM90]. photoinduced [GDG99]. Photoinhibition [IM98]. photolysis [WHB95]. photolytic [SB94]. photometer [KE91]. photon [Dod92a]. Photooxidation [KHS97, VBD^{+}97]. Photoperiodic [Den97]. Photophysiological [MM99, PZ93]. Photophysiology [MC99]. photopigments [MSK^{+}97]. photoreactions [MZ97]. photoreduction [TBM95]. photoresponses [FH92, FR99]. photosensitive [CH90]. photosynethetic [OCS96]. Photosynthesis [BH91, CSM94, EKT99, SJP94, And93, BM92b, CG91, Dod92a, FAE99, Gei90, GAAG96b, GGF91, GKS^{+}94, HD97, HB91, Hof90, IM98, KF93, KJSJ98, LPJ92, LTPB98, LPPSR97, Les96, LK94b, MVT99, McB92, MLR^{+}94, NCD98, NCRS90, OGH^{+}96, PSEJ99, QWR^{+}95, SJPGH97]. photosynthesis-irradiance [Dod92a, HB91]. photosynthesis-light [McB92]. Photosynthetic [BF97b, SFRB98, CS95, CP99, Dod92a, FDLD93, JM93a, KWF90, KCKC90, MMG94a, NSJ90, OZ95, OP92, SLSLBD97, SCKF99, VRT^{+}92]. photosynthetically [JLO97]. phototaxis [RF94c]. phototrophic [KJ92, OCP92, SFRB98]. phototrophs [Rav97]. phthaldialdehyde [CH91]. phycoerythrin [Wym92]. Phylogenetic [DFA93, RKG97]. phylogeny [Leh95a]. Physical [Gar91a, MCCD92, MFY99, AGZ99, AGGD90, AC92, CLKJ95, DO97, Goe98, GRG98, JPG90, KSYK96, Sal92b, SF95, Pri94, Spi96]. Physicochemical [PC97]. Physics [DHK^{+}96, UG98]. Physiological [BF98, FCI91, GAAG96b, GKS^{+}94, CDPR98, KHB96, PLP99, SZA92]. Physiology [Sel97, BKB97]. Phytochelatin [AKM95, AM95, AMM97]. phytoflagellate [CPS90]. phytoplankter [CL90]. Phytoplankton [ADC90, Bur92, Li90, MD90, OCS96, SS99a, SR95, SN90, Som93, SHSS93, Ven93, WCD^{+}94, WGSP91, YPF^{+}98, ADC92, ASMB98, AC92, And99, Ano92d, Ano93r, Arm99, AR96, BTLC93, BP91, Ban90d, Ban91, BF97a, BH95, BF98, BSY^{+}95, BB97, Boo94, BSC93, Bra91, BS90b, BLR92, BP96, BZR99, BGF92, CR96, CDHS98, CCKP98, CAS93, CS97, CBFC98, CBMC^{+}99, CBC^{+}91, CL96, Cla94, CGVL95, CHD91, CCP92, CLMH93, CW92, CW95, CGB90, Cur90, DFN97, DC96a, DHB93, DAC90, DAC92, DOC99, EG91a, EF95, EP96, EP97, FYG98, FAE99, FVL98, FVL99, FP99, Fur91, GJC92, Gar91b, Gei90, GMK96, Goe98, GHT^{+}96, GVK93, GW90, GKS^{+}94, Han96, HCSE97, HVHH91, HSRR96, HS92, HM90, HvOW99, HFBB99, HCP90, IKT99, IPPD92, JBJB96, Kam92, KOWH99]. phytoplankton [KZ92, KF93, LLLB97, LV98, LBH^{+}93, Lew90, LK94b, LKM98, LID93, LDV^{+}92, LP94, LvGLC98, Mac93b, MS93a, MP92b, MdlC97, MMG94a, MVT99, MDWS97, MHIHH99, MC92, MBHH^{+}91, MBHT94, MMC94, MNAH96, NTH^{+}91, NCD98, OR94, OZ95, OP92, PM90a, PH97, PPK^{+}99, PP97, Pra90, PH92b, PM91, PAM94, RJ91, Rot92, SdBdJL98, SS99b, Sel94, ŠBM^{+}95, SRG98, SRG00, Sma97a, SM98a, Som89, Som94, Som95, SCEG95, Str90, SW91, SH92, SH95b, SH95c, TM95a, TZR98, TL94, TKL^{+}97, TT97b, VLHL97, VT90, Ven97, Ver94, Véz94, VdHH^{+}98, VVIK99, WW95, WMD97, Web90, WH94a, WMPC^{+}90, WH93a, WKZK98, dAL94]. Phytoplankton-bacterioplankton [MD90]. phytoplanktonic [GMK98, QLBAP96]. pico [FR91]. pico- [FR91]. picocyanobacteria [Pic91]. picoflagellate [CS95]. picophytoplankton [VTCB96]. picoplankter [DŠ98]. Picoplankton [CPLD99, JLMV98, CDPR98, DF94, FRL94, FBCS90, MLB99, RKG97, SLSLBD97, ŠBM^{+}95, TC95b]. picoplankton-size [CDPR98]. picoplanktonic [PBL99b]. Piedmont [GW95, PSH^{+}96]. Pierrot [Ano93h]. Pierrot-Bults [Ano93h]. Pigment [BPCL97, LLLB97, SW91, BM92b, Cla94, GLM92, HHR94, HS92, JOG96, LC90b, LBH^{+}93, MR92b, PD90b]. pigment-based [JOG96]. Pigment-specific [SW91]. pigmentation [WWT93a]. Pigments [OCZA90, BLS^{+}95, HA90, QLBAP96, RR92, VFAA98]. Pinelands [RS99]. Pinhey [SB90]. Piracicaba [MBK^{+}99]. Pisces [KDKS90]. piscicida [BG97a]. piscivore [Ham94, PCC^{+}97]. piscivore-planktivore [Ham94]. pistillata [FPAG^{+}98, SRLRL^{+}99]. pit [SL98]. Placopecten [CH94]. plain [Leb90]. plains [Rah90, HLQ^{+}99, LFGM96]. Planck [WL99]. planetary [Leh95b]. Planktivore [WWT93a, Ham94]. planktivorous [LVM^{+}90]. planktological [WW98]. Plankton [BOB94, HWP95, PY99, BTM92, BdBNvB91, CRKS93, DHK^{+}96, EK90, NFGP^{+}93, FGN93, Gae92, Gae93, GSL97, HY97, HLGA98, HR90, HSM^{+}95, KKK^{+}93, LPPSR97, Leh91, LGW91, MJG91, PSEJ99, PSJ99, RG92b, SJDJ94, Set92, Sie98, SJHS98, TZGS98, UNK95, WW98, WBP99, WSM^{+}96, dGBPP96]. plankton-nutrient [NFGP^{+}93]. Planktonic [PMF90, SBZD97, ADS94, ASC94, BPK94, Bre93, BS90c, CFT92, Eck90, ER95, FPR94, FS99, GP94, GdGD97, Gil94, Gil96, HBH94, HHB93, HBL^{+}93, HT96, KB92, KTF96, LK94a, LJD95, MTM^{+}98, MF90, MDdM^{+}99, OW96, Pet94, RF94a, RHT92, RS93, SJL95, SR92, SBB^{+}93b, SFC90, SRPG92, TNG^{+}92, UW92, VORJ93, VBAO95, Ver91, WPBS99, dGP94]. Plant [BLS^{+}95, GLM92, CK92b, KJSJ98, LD99, MJ94, MH94b, NK99, RJC^{+}97, Ano93g]. Plant-animal [Ano93g]. plant-like [NK99]. plants [BW90, CFW97, Dua92, SVC^{+}98]. Plaster [TG94]. plasticity [Bla93, DAC90, Lam94, LLZ99]. plastidic [Ver91]. Plateau [MH97]. Platte [RKA96]. play [Wil95, WHOL99]. plicatilis [Som98]. Plön [WL99]. plume [GCCJ97, GGFR91, MD90, TM95b]. Plußee [Dua95, PA95b]. Plußsee [Som93]. PN [DW95a]. PN-specific [DW95a]. Pocillopora [JLO97]. pockets [WDGS92]. Pocono [MH97]. point [HLGA98]. points [Joh94]. poisoning [KDB91]. polar [Hen98, RGR^{+}98]. Poleward [War90]. Polishchuk [Kim90]. polluted [CACM98, FBW96, Kna99]. polychaete [CJVD97, Hen98, KW93, Mar94, WSB98]. polychaetes [BM99]. polycheate [FL90]. polyedrum [LR99]. polymer [AR94]. polymictic [Nür98]. polynya [CDHS98, YD99]. Polyphosphate [HGR95]. polysaccharide [FL95]. polysaccharides [SBW^{+}98]. POM [dGF96]. Pomatomus [HC96]. pond [SRPG92]. ponds [Eic94, MASA94, WT93]. Pontoporeia [HQCG92, MCB90]. Pool [HAL90, Ste94, KY97b, SWJO98, SRG98]. pools [BFE96, VDA99]. poorly [PDM96]. Population [HPB^{+}99, TYL92, ADR93, BV95, HY97, Hen98, JLMV98, JW92, MT90, RKA96, SRLRL^{+}99, Two94, WSB98]. populations [BG94a, BLR92, BMAW97, CRKS93, DŠ99, FBR91, FGN93, LVM^{+}90, LGW91, MCC^{+}96, MCB90, MW98b, PPK^{+}99, Pol90, PAM94, PWR94, SMB^{+}97, SB90, UC98]. Por [Ver90]. Pore [WFG99, BKB97, BRPR98, CR93a, CTSB98, DMF^{+}98, KDZD94, LBL^{+}98, Mar94, MM93, MKBN99, Muc91, SSE90, SBWB94, WTB98, WF90, WHM94]. Pore-water [WFG99, BKB97, KDZD94, MKBN99, SSE90, SBWB94, WF90, WHM94]. porewater [GREGG97]. Porites [Gle93]. porosity [CA91, CGSD93, MBS98, SSE90, SKR^{+}91]. Porphyridium [SM99a]. Portage [KLR94]. portions [Ano93y]. Posidonia [MD97]. position [KBB98, WKB^{+}96]. Positive [FCG94, CCP92, Pol95]. positively [MV96]. Possibility [Bre93, UW92]. Possible [FC91, Mac93b, HKPM99, RG92b]. possibly [MFW^{+}91]. post [AS91, HRD^{+}95]. post- [AS91]. post-Chernobyl [HRD^{+}95]. Poster [Ano93a, Ano95c]. postlarvae [AB99]. postsettlement [Sam91]. Potamocorbula [WH93a]. Potamogeton [CK92b]. Potential [CRKS93, DHK^{+}96, EP96, ML96, MSM96, OZ95, PFB96, RKA96, Arn98, BMC^{+}96, FLH96, Hay96, MPT99, MB98b, PWR94, WPBS99, ZF93, ZFdSY96]. potentially [ME98b, SBZD97]. Potentials [SB94]. potentiometric [Muc91]. Poterioochromonas [CPS90]. Potomac [RJC^{+}97]. pouchetii [MWB^{+}98]. Pourriot [Edm96b]. prairie [EP96, PJSW96, WR95, Som90]. prealpine [WMPC^{+}90]. precise [KT92, RCCdG99]. Precision [Ami90]. predacious [DC91b]. Predation [EEL97, PWR94, AB99, AMP98, BOFW93, BL91, CRKS93, GU94, GL94, HDP^{+}90, ME98b, MPA94, MC96, NRR96, PC96, STJ^{+}93, SH97, Sve97, TJK^{+}97, THH98, TSH99, Ura90, WCT90, WHOL99, WH93b]. Predator [Bla93, BO96, Bra98, BL94b, DL92b, FH92, HD95, LL96, Leh93, Str97]. Predator-induced [Bla93, DL92b]. predator-mediated [LL96]. predator-prey [Str97]. predators [FDG99, HBH94, Lei90, MK95, Pol95, WPBS99]. predatory [BL94b, CG95, LC93, MT95, Paw98, SY95, SC90, TJK^{+}97]. predict [BS90c, SPMW97]. predictable [Ant99]. Predicted [KOWH99]. Predicting [Bra98, BLD99, CG93, CKCP99, Dod90, Dod92b, GRG98, Lan93, DK93a]. Prediction [Nür98, Pet94, BA97, DG90]. Predictions [PK91, CL90]. predictive [HB94]. predicts [BH95, Bre91]. preference [CMT93]. Preferential [TIOS99]. Preliminary [DC91b, DHP^{+}96, SC93b]. Prelude [Kna99]. presence [CSS96, DPC90, RR92, VPN^{+}90, Wel94]. presence-absence [CSS96]. Present [BG97b]. Present-day [BG97b]. Preservation [PG97, KJLA98, KJS98]. preserve [Kap94]. preserved [HHB^{+}92, LHS90a]. pressure [Ano93n, CT93, ECA99, SS99d]. pressures [Car98, MC96, NRR96, dABL91]. pressurization [BSO92]. prevent [HSM^{+}95, TMB96]. preventing [ME98b]. Prey [BL94b, CŠ90, Swi92, BO96, DC91b, HBH94, JD95, MC95, MLB99, RF94b, SC90, Str97, Ver91, ZF93]. Prey-size [CŠ90]. Price [Ano93g]. Primary [BSY^{+}95, Li94, Ano93x, BTLC93, BPK94, BC91, BF97a, BE94, CCL92, CHA91, CKCP99, FZP92, GMM96, GG90, Har95, JPG90, Jön91b, KHRS98, KKBG91, LBB^{+}97, Lig92, LAC^{+}96b, PSO91, SLD97, VV98, Was90, YK93, YKM95, Sut93b]. primer [Pla93]. Principles [Bur94]. priorities [Lew98]. PRO [LJD95, XB97]. probable [PM90b]. probe [LDC99, OZ95]. probes [LK94a, MBOCK95, SMB^{+}97]. Probing [GWMB91]. Problems [Hai90, Lig92, HGP97, Leh95b]. procaryote [RR92]. Procaryotes [RBMH91]. procaryotic [GGA91]. procedure [Kap94, WHK^{+}93]. procedures [TB91]. Proceedings [Ano90i, Ano91f, Ano94]. process [Kap94, MR98, UA97, Man95]. Processes [HKP^{+}97, PK94, BMC^{+}96, BGC99, CL91, Cha96, Gar91a, Ker93, KS93, LS92, May92, Meh99, MM97b, Mor90, Mul92, Pal92, Pri94, SBWB94, SKB94, Spi96, SBGCC91, UGU^{+}90, HL94]. Processing [DŠ97, DC96a, HCR^{+}94]. Prochlorococcus [SHKI95, CNV94, CJD^{+}99, MC99, RR92, UC98]. prochlorophyte [SHKI95]. prochlorophytes [Li94, VPN^{+}90]. produced [AV94a, BL98, KMWK92, Mac93a, MHMD99]. producers [KHRS98, MV98, YKM95]. product [HHR94]. Production [AV94b, BKR^{+}90, GNL97, HS98, MB96, NML95, NZHB95, SM99a, SH98a, TLvB^{+}91, AMM97, AKM95, AM95, AW93, AC92, Ano93k, Ano93s, Ano93w, Ano93x, BTLC93, BP91, BPK94, Ban90d, Bel90, BSY^{+}95, BB97, Bra91, Bre93, BE94, BFW92, BDH91, BP97, BPC^{+}97, CCL92, CCKP98, CKCP99, CLKJ95, CP95a, CP95b, CW95, CFT92, DC96b, DWM90, DDC92, FPL^{+}91, FZP92, GC97, GCJ97, GG90, GLT96, GW95, Har95, HHB93, HRF91, HDB93, Ive90, JPG90, Ker95, Kim90, KV99, KBH98, KFW94, LBB^{+}97, LPA91, Li94, LDV^{+}92, LAC^{+}96b, MWB^{+}98, MGN92, MdlC97, McL97, MH95, MH90, NK94, NSK95, OJ98, PMF90, PBJD93, PSO91, PFDH93, PK94, Pol90, PSH^{+}96, PH92b, RB91, RW96, STJ^{+}93, SFRB98, SD95, SBPV99, SD94, SFFF97, SN90, SU98, SM98a, Som98, SBGBC96a]. production [Sti91, SGA98, SHB99, TM95a, TN94, TIOS99, TL94, UA97, VV98, WCD^{+}94, Was90, WLY93, WT94, WMT96, WS96, WWMI98, YK93, YAF^{+}92, ZFdSY96, dAL94, dGGV^{+}96]. productive [CAS93, CS97, RHT92]. productivities [BM96]. Productivity [BW91, BC91, BF97a, CHA91, CGVL95, Coa91, CGB90, FPC91, FP99, GMM96, HVHH91, Jön91b, KHBL98, KKBG91, Lap95, LID93, Lig92, MP92b, MWW97, MHP91, MKD90, NS91, NB97, OYWK96, SH91, Sut93b, ZKM93]. products [HSI95, LH91, TR98]. profiler [MKBN99]. profilers [Old94]. profiles [BRPR98, CW93a, CC91, KRB98, MR99, NWS94, SFW^{+}95, VBvG91, WMM^{+}94]. profiling [dBGEK97]. Profundal [GJ96, SBGCC91]. program [Con90]. programs [Wel99a]. promotes [MT99]. propagules [AWL94]. properties [BHAV91, BKT96, BRZ95, CBB^{+}95, CTSB98, DO98, FP99, GB94, Hil90, HVB93, HCLD95, KZ92, MaB98, MSSM96, MCCD92, Mor97, RSK97, SBP^{+}96, Sel97, SR93, SM97, SCKF99, SCF99, VBK98]. propionate [BKR^{+}90]. proportion [dGGV^{+}96]. Proportionality [Ano93x]. proportions [HS99]. proteases [BF98]. protein [KHB96, MR92a, MBHT94, Tay95]. proteinaceous [PL99]. proteins [RLL93, TIM96]. protist [THK^{+}99]. Protista [RERB91]. protistan [MO91, Pet94]. Protists [dGGV^{+}96, Ano93o, CM91, FN95, Tra94, Wic95]. proton [DGW92]. proton-induced [DGW92]. protons [McC91]. protozoa [CG93, CP97a, CŠ90, CFT92, FPR94, San99, SJL95, Dol91]. Protozoan [BFKL91, CFSW91, LVFD93, Str97]. protozoans [FR91, PV94]. provide [KLR94]. province [TBM90]. provinces [CK95, Cla94]. proxy [LKSA97]. Pseudo [SMB^{+}97]. Pseudo-nitzschia [SMB^{+}97]. Pseudocalanus [OW96]. Pseudomonas [Jan93]. pseudonana [LTPB98, RSK97]. PSII [FK98]. publishing [Kir98]. pulex [Bla93, Den97, GD96, HD95, RKW94, WH93b]. pulicaria [Epp96]. Pulsed [LWMD96, GK97]. pulses [FNWK96]. pump [OZ95, RH93]. pumping [PCFS97]. punctata [JB94]. punctipennis [HCKW90]. pure [CTF98, CLP^{+}92]. purification [TvE94]. purified [DZC91, HSI95]. purple [PAS90, OBH^{+}91]. purpureum [SM99a]. pusilla [CS95]. PVC [MBSW98]. pycnocline [SDS95]. Pyramid [Ben94, BW94b, HB94, BW94b, HB94, MTLR98]. pyrifera [WH92]. pyrite [CPT^{+}94].

Q [Ano93h]. quadricauda [Jan93]. Qualitative [Ker95, QLBAP96]. quality [CDPR98, HLQ^{+}99, ME98b, NML99, SS99a, SHSS93]. quantification [Hen98, SMB^{+}97]. quantify [SSFP90, SBK95]. Quantifying [CW93b, JC95, Nür95]. quantitative [HTW90, Ker95, KC92, KW98b, Roe98]. quantitatively [BH95]. quantity [Rom91]. Quantum [PH92b, BTL^{+}95, FK98, GKS^{+}94, LP94, OGH^{+}96, SM91]. Quebec [HCLD95]. question [EABK99, KHP91]. quotients [GJ90].

R [Ano93e, Ano93f, BCBC93, Boo94, Cha96, Con90, Dua95, Edm96c, Edm96b, HL94, Jum94, Kim92a, Leh95b, Lew95, May92, Mei90, PA95b, Rav95, Sal92b, Str94, War90, Web95, Wet91d, Whe91, Won91, Wri90, QP91, dGP94]. Ra [BPS94]. Radiance [KJ94, AK97, Ban92]. radiation [DC99, Gle93, GHT^{+}96, HDMB97, HHL98, JLO97, KL90, KBMSP99, Les96, MJC93, MZW^{+}95, MDdM^{+}99, NCD98, NE93, ORH99, RHH97, SLD97, SML96, SRLRL^{+}99, VRL^{+}98, VL96, VL98, Wil95, WSM^{+}96, ZCGP98]. radiative [MS97]. radiatively [MJ94]. Radiocarbon [CBD^{+}99]. radiolarian [BUW96]. radionuclides [LBL91]. Radiotracer [SJDJ94, LF93]. radium [WHM94]. Radix [Bre97]. Raffaelli [Man95]. rainwater [MS99]. ralfsii [Gen91]. Raman [BSMK98]. random [KB92]. range [CCL90, HBH97]. ranging [RRK94]. RAPD [CM95]. Rapid [Arn98, HA92, Kan93, RCCdG99, SFC90, Tra97, YK95, KMWK92, LSS^{+}97, PS92]. rapidbacterial [WHB95]. Rare [LHM^{+}93, CWW98, JL94]. rate [Ano93r, BC91, BH95, CGVL95, CW95, DW91, Gla91, GG90, JC95, LSB92, LBP97, MDE97, MV96, NSJ90, PF91, Pol95, Rot95, SD95, SD94, SH98b, SVC^{+}98, WSB98]. Rates [Ban91, RF99, ASMB98, ASC94, And90, Ano92d, Arn98, BF97b, BA95, BP97, BGF92, CDP96, CG93, CBC^{+}91, CT93, CH94, DW95a, FSS94, FR91, Fur91, GCCJ97, GCB^{+}98, GW99, Gla91, GW93, Goe98, Gra91, IH95, JHM96, JWRS96, KMC90, LB96, Lan93, LLLB97, Leh91, LB94, LB91, MMCL94, MK95, MAN91a, MFLR96, MPG95, Mor90, ML90, MHT99, OP92, OJ98, PV94, PCFS97, PF90, Pet94, PK94, QWR^{+}95, RV99, SCFB98, SW91, SRPG92, TC93, VBAO95, VDG^{+}97, Véz94, WTH92, WW95, WP94, WH93a, dHK98]. rather [Sch92]. ratio [Ano93c, Ano95e, BZR99, CGVL95, HBH94, LS92, MB90a, MMH^{+}95, NSJ90, NWS94, QWR^{+}95, QP91, Som93, Str97, UW93, Véz94]. Ratios [Ber97, BLK95, CK95, FPC99, FZP92, FSSI97, GLM92, HK98b, HRD^{+}95, Lan93, MS93a, Sar92, Som90, TKL^{+}97, dGP94]. ray [CHC97, DGW92]. Re [BW90, PF90, SBB^{+}93a]. Re-evaluation [BW90, PF90, SBB^{+}93a]. Reaction [Mor90, JHM96, SS99d]. reactions [HM90]. reactive [Cam94, WHK^{+}93]. reactivities [CH92a]. reactivity [DMHH99, OB98]. real [UCS97]. Realization [DWM90]. really [Ban90d, Sta90b]. Reassessment [MR99, LTPB98]. recalcitrant [WHB95]. recharge [IM90]. Recognition [CP94, TZF96]. reconsidered [Ven93]. reconstruct [SH95a]. Reconstruction [KFEC91, VL91, BFG94, KRW99, LFGM96]. record [BSY^{+}95, Edm91a, KLR94, MVM97, MTLR98, NB97, RA98, TJC^{+}98]. recorded [FNWK96]. Recorder [HWP95]. records [AR95, BM96, Fri96, HY97, STJ^{+}93, Cha96]. recovery [KMWK92]. Recruitment [Lev96, MB90b, CR99, EG91b, FBR91, Han96, HJ94, Hen98, PK94, PCC^{+}97, Sam91, Str95, WO98]. rectangular [Lou90]. rectangular-mesh [Lou90]. recycled [CCL92, WPBS99]. Recycling [RGR^{+}98, HDB93, KSC^{+}90, ME98a, TIOS99, TCA96, WK90]. Red [LP95, AWKH98, Ano93y, HGP97, JPJ97, Joh94, LR99, MSK^{+}97, SM99a, TSV^{+}91, TS97, VDG^{+}97]. Redfield [HS99, HG99]. redistribution [SJDJ94]. redox [CH90, KS93, SH90]. reduce [AB99, HDP^{+}90]. reduces [ME98a]. reducing [GLS90, HB95a, IMB^{+}99, JRM98]. reductase [BH95]. Reduction [Han91, FFNS99, KHS^{+}98, LG99, PB91, RW96, Smi90a, UBBS94]. Reef [LWMD96, AB92, BA97, BA92, BA95, CHA91, CG91, DY99, HSM^{+}98, KSYK96, MHM^{+}99, PG97, PSO91, Paw98, PLP99, VSBZ^{+}94, WWD90, YKM95, TM95b]. reef-building [VSBZ^{+}94]. reef-hat [AB92]. reefs [HSS^{+}99, HC99, Lap97, Lap99, LL91, LJBTB99, MHM^{+}99, Sal92a, TA97, YPF^{+}98]. reexamination [LM98b]. reflect [FGN93]. reflectance [MMG94b, MUGK92, PWM^{+}98, SCF99]. refractory [MM97b]. Refuge [KSH^{+}99]. refugia [Ber99b, TSD97]. regenerated [HDB93, KFW94]. Regeneration [CL91, LLB^{+}96, TC95b, TCA96, AA91b, GGFR91, HNS95, HT96, MPG95, Roc98, SH95c, WRLF96]. regime [DW95a, FGN93, HDZB98, MMTJ94, ML96, NS91, Ver99]. regimes [KE91, MWW97, NS93]. region [KKM^{+}96, MBK^{+}99, WLBQ95]. Regional [GW94, Pse92, AR95, KFEC91, Liv99, VL91]. regression [Ber97]. regulated [SWH92]. regulating [BP96, Han92, MT90, MH97]. Regulation [BC91, BL91, LDV^{+}92, Mul92, NK94, PC96, SJPGH97, SH95c, TN94, ADC92, ES93, GGFR91, KBH98, LKM98, LG99, LVM^{+}90, RRPP^{+}94, SD94, SB90, VC90, WH99]. regulators [WMPO99]. regulatory [GMK98]. reinhardtii [WG97]. related [CKR93, CNSR90, GL94, Hen98, LID93, MC95, Rot90b, TC93, WBP99, ARM96]. Relating [DO98, Ban90b, BKH^{+}96]. Relation [CSS96, Ano93k, Han96, MNAH96, Per97, Pic91, YH92]. relations [BH91, NFGP^{+}93, HB91, Wic95]. Relationship [BTL^{+}95, FRY95, HDR^{+}98, OJ98, Was90, BPK94, DM92, GMM96, Gei90, KPSB95, Liv99, MMCL94, Mof94, Mol98, QP91, WMHB95, dGF96]. Relationships [FFF98, HCLD95, LPA91, MS93a, PASC95, TZGS98, VRT^{+}92, DL93, MD90, MV96, SLC^{+}99]. Relative [EP97, CL91, DWM90, FBS99, JHM96, RMT96, WTH98, WC91]. Relaxation [RWH^{+}98]. Release [GHF^{+}97, NK91, VM94, BW99b, BGF92, CH90, GNL97, HTSDZ99, HC94, JA92, KHBL98, LVvdB99, MB98b, Per97, RT92, SRG00, SH98a, VBAO95, vEJ97]. released [Han91, HSI95, LV97]. relevance [Liv90, Sma97a]. relicta [GRG99]. remains [Ker95, SLC^{+}99]. remineralization [FPR94, KMC90]. remobilization [HTSDZ99]. remote [Bar99, JJGS96, SCF99]. remotely [BM92b, ZKM93]. removal [Mar94, WWSJ98]. renewal [HKP^{+}97, KCWM96]. repair [WWSJ98]. replaces [VHH96]. replete [HFBB99]. Reply [CG93, CKCP99, CA93, DW95b, Hes93, HM99, JD96b, Lap99, MSH96b, NT93a, NRPRB96, TMB96, UW93, ADC92, CK92a, CP95b, Gib92, GLH91, HR90, HSM^{+}95, Kie90, Kim90, MBFG90, NKMO91, Rot91, SRG00, Som90, WMTB93b, Wet91b]. reporting [Ano95e]. reproducing [Pol90]. reproduction [AB94, Den97, Ste90, WG97, WSMS96, Str95]. Reproductive [SWS^{+}96, MW98b]. require [Ano93n]. requirements [Bra91, WMT96]. research [CK92a, Hai90, Lew98, Wet99]. reserves [AE91]. reservoir [Bal96, JX94, MIM90, SVW^{+}97, ŠBM^{+}95, ŠKN^{+}99, Ura90, Thr90]. reservoirs [Hai96b, Sch99, WTH98]. residence [MM91, Rie92]. residual [DMA^{+}96, Won91]. Resolution [Jør93, WMTB93a, AS94, FHF99, KLR94, KDZD94, MKBN99, MBSW98, TMLS99]. resolved [KLB^{+}99]. Resource [Sie98, Som90, Ano93c, IKT99, KTF96, MS93a, PWR94, Som93, Edm96c]. resource-ratio [Ano93c, Som93]. Resources [Lei90, OR94, PC96, PF91, RJ91, SP90a, SWS^{+}96, VF94, WSMS96]. respiration [BOB94, CP97b, Gla91, GAAG96b, JC95, Ker93, LK94b, PSO91, PSEJ99, QWR^{+}95, SHB99, SRPG92]. responding [dBGEK97]. Response [BMP92a, CFP97, CCT93, Goe98, RKA96, TM95b, TJC97, WMPC^{+}90, BO96, BF98, BC97, BM99, CACM98, CBMC^{+}99, CGB90, CH92b, CP99, Den97, GJC92, Gil96, GJ96, Hec99, JM93a, KM94, LR99, LLZ99, LJH91, LGT95, LB91, Mac93a, MCB90, MJOC93, MB96, Mol98, Old94, PH92a, Pla98, PRT^{+}93, RB97, SF95, SHM96, TMLS99, WG97, WT94, vLM91]. Responses [IKT99, LK94b, NHJ91, Bla93, BSG^{+}99, CKCP99, CCC^{+}96, CK95, DY99, FMK^{+}99, LCR99, LFHS99, LC93, LGW91, ML96, MLLR99, NRR96, ND91, PZ93, PTJ96, Rot90a, SZA92, WKB^{+}96, WLCG96]. Restelli [Mat94]. resting [MLBC94, SAK91]. restless [Fee99b]. Restoration [MSP^{+}99, Bac99, JRM98, Smi90a]. restrict [IM98]. restricting [EP97]. resulting [MWB^{+}98]. Results [MAS91, MNAH96, BW94a, CKR93, DM90, JF99, Sar99]. Resurrection [WMPO99]. resuspended [BL94a]. Resuspension [CR93b, HL92, RG92a, San92, dJvB95]. retained [TF95]. retarding [LHWZ92]. Retention [RF94b, AMF98, BDV99, Bre91, CBB^{+}95, KW98a, LPA94, PY99, VMDC96, WH94b, WBRL98]. Retrieval [HVB93]. reveal [MBKG95]. revealed [CR96, Cla94, GSC^{+}96, LF93, dHJS90]. reverse [HW91]. reverse-phase [HW91]. reversed [Den97]. Reversibility [KHS^{+}98]. Review [Abb90, Alt94, Amm96, Ano93h, Ano93e, Ano93f, Ano93g, Arc93, Ban90c, BCBC93, Boo94, Bur94, But95, CT96, Cha96, Con90, Cur94, Dol91, Dua95, Duc94, Edm96a, Edm96c, Edm96b, Fen94, Hai96a, Har95, HL94, Joh91, Jón95, Jum94, Kim92a, Leh95a, Leh95b, Lew95, Log95, Man94, Man95, Mat94, May92, McC97, Mei90, Men90, Mil95, Ohm96, Paa95, Pai96, PA95b, Pla93, Pri94, Pse92, Rav95, RW95, Sal92a, Sal92b, Sel94, Smi90a, Spi96, SS93, Str94, Str95, SW93, Sut93b, Thi94, Tho90, Thr90, Ver90, Wal96, War90, Web95, Wet91a, Wet95, Wet91d, Whe91, Won91, Wri90, Gra97]. revised [BH95]. revisited [Ano92d, Kir94a, NS91, WH94a]. Revitalizing [Wel99a]. Rhizosolenia [SBPV99]. rhizosolenid [SVF91]. rhizosphere [RPJ97]. rhodamine [SB94]. Rhodophyta [MHMD99]. Rhone [IPP97]. Rhythmic [Roc98]. rhythmicity [Loo93]. ribulose [PPK^{+}99]. Rica [PRT^{+}93]. RICE [MR92a]. rich [CNSR90, CHD91, EW99, Fro91, GPMF96, Gar91a, LDV^{+}92, MASA94, PW93, RHH97, SVC^{+}98]. richness [Dod92b, SHKP96]. Ridge [PRCV98, TV99]. rigidulus [NSBB98]. Rigler [Lew95]. rimana [YSSO91]. riparian [Mul92]. Rise [Nod97, PD90a]. risk [TJK^{+}97]. River [Ben94, BW94b, BGC94, BGN98, HB94, MBK^{+}99, MR99, CM91, FPC99, GW95, KL96, OB98, PFDH93, SWJO98, SPMW97, TJC^{+}98, BOCL^{+}95, DK93b, FPL^{+}91, GBC95, GCCJ97, HSCM97, HCR^{+}94, KMC90, LM91, QWR^{+}92, RKA96, RHD^{+}90, RJC^{+}97, SHB99, Wal96]. riverbank [KMU95]. Riverine [Con97, EM93, FFF98, GW94, KZM90, MBK^{+}99, SFFF97, TM95b]. Rivers [CW98, MB90a, BP96, CPMW99, Kil90, QWR^{+}95, WW98]. Rn [CBBC96]. RNA [LK94a, SCFB98]. Robarts [Mei90]. Rock [BEB92, FMK^{+}99, SBWB94, UBBS94]. Rocky [MDWS97, WLCG96, HQ90, Pai96, Set92, RKA96]. Role [AA91b, HP95, Jør90, Lon91, MZ97, MHSW99, RHH97, RB94, AS91, BW90, CRKS93, Cra95, De 90, Fro91, Hay96, Hes93, JC95, KHBL98, KHRS98, KSH^{+}99, KHP91, KGFP91, LAC^{+}96b, MB98a, MT90, MZW^{+}95, MH97, PM90b, RHT92, RMT96, RS99, RB97, SJPGH97, SMV98, SGP92, VORJ93, WMTB93b, WPBS99, WWSJ98, Wil95, WHOL99, ZMP98, ZFdSY96]. Roles [Mul92, AML93, PPK95, PFB96, WJRB97]. roots [LKD99, SVC^{+}98]. rosette [HGFC96]. Ross [CDHS98]. rotifer [Rot90b, Rot95]. rotifers [Gil94, Gil96, PV94, Rot90a, Sch90]. rotunda [SB93]. roughness [Eck90, TA97]. routes [RF99]. ROVERSE [GWMB91]. Rowe [Thi94]. rRNA [LK94a, RKG97, SMB^{+}97]. rRNA-targeted [SMB^{+}97, LK94a]. Ru [HRD^{+}95]. rubens [Rot90a]. Rubisco [Hay96, OP92, RC95]. Rueter [Cur94]. running [RW95]. runoff [RM96]. Rutledge [Liv90].

S [Alt94, Ano93g, Bur94, Har95, Joh91, Man95, Mil95, RKG97, Spi96, SW93, War90, LK94a, LBL^{+}98, Mil91, NWS94]. Saccoglossus [KM94]. Sado [Roc98]. safe [RC93]. salamanders [BW99a]. saline [EP96, EP97, GS91, JCM^{+}99, MHSP90, MKBN99, ND91, RM96, WR95, WFG99, WPAA95]. salinity [Fri90, GRG98, JAMH96, KKS^{+}99, MCCD92, SM99b]. salt [BFG94, BFS91, BdBC99, CLMH93, HKvS^{+}92, Hin94, IPP97, KLB^{+}99, LKM98, MM91, OBH^{+}91, OBKH96, RRK94, SD95, SVC^{+}98, TH96, WH94b]. salt-marsh [BFS91, BdBC99, Hin94, LKM98, SD95, TH96]. salt-wedge [IPP97]. saltans [DŠ98]. saltatrix [HC96]. Saltmarsh [Man94]. same [MBKG95]. Sammamish [BMP92a, BMP92b]. sampler [CSPG94, DMF^{+}98, KDZD94, LSB92, SSE90, WF90]. sampler-incubator [LSB92]. samplers [CSPG94, Fil91]. samples [Kap94, KMWK92, KC92, Lig92, LDC99, LHS90a, MNAH96, NJLW97, Sos99, SRPG92]. Sampling [KJS98, BR91, KDZD94, Liv90, WW98, YGB91]. San [HCP90, LvGLC98, CCR^{+}95, LL98, MHSW99, PD90a, WBKF95]. Sanaga [BGC94, BGN98]. sand [App94, DMF^{+}98, JMMH98]. sand-bottom [DMF^{+}98]. sandflat [AOTR97]. sandflats [EG91b]. Sandgren [Leh95a]. sandy [KJ94, SSE90]. Santa [Cot99, ALB90, OC97, RLTB90]. sapidissima [TJC^{+}98]. sapropel [BCM92]. Sargasso [BS90b, CDHS98, CPLD99, Goe98, NB97, UC98]. Sargassum [Lap95]. Sas [Smi90a]. satellite [BF97b, GSG94, MMTJ94, WMC^{+}96, Abb90]. satellite-based [BF97b]. satellite-derived [WMC^{+}96]. Satilla [CW98]. saturation [GA96]. sauce [XB97]. savanna [BGC94]. Savin [GG90, TGG94]. saxitoxin [HSI95]. scalar [Ban90b, Ban92, LPJ92, LM98a]. Scale [Man95, ASC94, ALB90, BMC^{+}96, Cur90, GJC92, Guy91, HCJ96, JGRR99, KSH^{+}99, LFGM96, MCED94, Old94, SB94, WLD91, Har95]. scaled [Siv93]. Scaling [HBH97, WVP97]. scanning [HC97]. scattered [BC90, VdHH^{+}98]. Scattering [VBK98, BSMK95, BSMK98, CW93b, GCP90, Gor91, GWY^{+}92, GWMB91, JX94, Kir91, LM98b, RSK97, WWT^{+}93b, WKZK98]. Scavenging [VJM96, Cur93, JHM96, PRS94, WM92]. scenario [JRT96]. scenarios [MQ96]. Scenedesmus [Jan93, LRvE94, LV97]. Schaefer [Sch92]. Scheiner [Mil95]. Scheldt [FBW96]. Schimel [Alt94]. Scholarly [Kir98]. Science [Thr91, HCPR99, Tho90, Rav95, Lew95]. sciences [Hai90]. scintillans [SW96a, TK98]. scintillation [BW94a]. scleractinian [GKM94]. SCO [HA92, LB95]. Scotia [NCD98, TLvB^{+}91, BdBNvB91, UGU^{+}90]. Scottish [BFM^{+}97]. screens [Lam94, LB96, Log93, LPA94]. sculpin [AMP98]. scyphomedusa [KCKC90]. Se [GHF^{+}97]. Sea [WDGS92, Ano93n, AS94, AMP98, Ban98, Ber91, BR91, CT93, Con90, CMPD92, Cul91, Dad93, Eic94, Fee99b, Fro91, GREGG97, KZM90, KDB91, MaB98, MVM98, MH95, MHP91, MKD90, PD90a, RBMH91, SBPV99, Sma97a, SGP92, SGP96, SHM96, Sti91, Str95, VF94, WWSJ98, dABL91, ASS^{+}98, AS91, BC97, BL94a, BDY^{+}98, BR92, BS90b, CR96, CDHS98, CPLD99, CSM94, DM98, Goe98, GW90, HB95b, JLMV98, KKS^{+}99, KKM^{+}96, KHK^{+}98, KFW94, KKBG91, LP95, MHIHH99, MMSKH94, NB97, OCP92, Ree90, SH96a, SMV98, SN90, UA97, UC98, VDG^{+}97, VPN^{+}90, VRHK99, Sel94, Sut93b, Tho90, Thi94]. sea-ice [AS94, Eic94]. sea/air [Con90]. Seagrass [TNG^{+}92, BLK95, FZP92, JMMH98, LD99, MD97]. seagrasses [DMM95, MCED94]. Seamount [MB90b]. SEAREX [Con90]. seas [HSM^{+}95, Sti91, BdBNvB91, GCD90, TLvB^{+}91]. season [SR95, WCW^{+}91]. Seasonal [BUW96, CR96, DD92a, FP99, GBC95, GGA91, HQCG92, Kli92, KFW94, LKM98, LVM^{+}90, MC95, PFB96, PJ91, PCC^{+}97, PRS94, RLTB90, SPB90, SBGBC96b, Ste94, VT90, VBD^{+}97, AE91, BGN98, CNSR90, CL96, Gae92, Hay95, KGFP91, KS93, MR92a, MR99, MHSW99, Nür98, SBWB94, SHM96, SBB^{+}93b, TGG94, VVIK99]. Seasonality [ZEG^{+}94, CMT93, MCC^{+}96, Ven93]. seasonally [BMP92a, CL91, Goe98, HTDM96, JUSI91, JSL92, MMS93, REL97]. seawater [Ano93t, DSV^{+}99, DHSS95, DHWW97, FN95, GG92, GCD90, HGFC96, HSL^{+}97, JAS95, KHS97, KW98a, Kli92, KNM96, LVvdB99, LD96, LF93, MS93b, Muc91, PL94, PW93, PD90b, PM90b, SBB^{+}93a, SH92, SIO98, Tay95, Tra97, TK90, ZMP98]. seawater-filled [PD90b]. seaweed [HDP^{+}90]. Seaweeds [Man91b, TR98]. Sebastolobus [WS90]. Secchi [KE91, SH96a]. second [CIM92, MWI92, RSZ97]. second-derivative [CIM92]. secondary [CP95a, Kim90, KV99, NK99, Pol90]. secret [LH91]. secrete [LGI94]. sectioning [Ver93]. sector [OAF^{+}99, RGR^{+}98]. Sediment [BK99, CWSX96, HL92, MJC^{+}98, SBWB94, WMM^{+}94, AK92, BH96, BRPR98, BK95, BPS94, BFW92, Bur92, CR93a, CDP96, CR93b, Coa90, CA91, CGSD93, DCML96, Ege95, EG91b, FNWK96, GREGG97, HY97, HSSA94, HHR94, HAAM99, JMA^{+}95, JMMH98, Jön91b, KG95, KAD95, KW93, LHWZ92, Liv90, LHS90b, Mar94, NCRS90, Nod97, PBL99a, PD90b, PHT^{+}93, PK91, REL97, RC93, SCK93, SMV98, SNB92, SGB^{+}97, SGG^{+}99, SGA98, SBGCC91, SKR^{+}91, Ver93, WTH98, WFG99, WWW98, Wri90, WGS97, dJvB95, HTSDZ99]. sediment-water [BH96, SMV98]. Sedimentary [BM96, Edm91a, Meh99, MTLR98, VFW^{+}92, DHP^{+}96, ES90, Guy91, KMC90, PSS^{+}96, SH95a]. Sedimentation [SS99c, VRHK99, Ant99, BG97b, Cra95, FNWK96, GG99, HS98, LC90b, LJA91, San92, Ver99, ZPHH98]. sedimenting [BBS97, LO97]. sediments [And90, Ano93l, AZS^{+}98, ALB90, AR94, Arn98, BEB92, Bal96, BFS91, BGC94, BLDP96, BBH^{+}92, BdBC99, Bou98, BFM^{+}97, BZ98, CFW97, CL91, CSPG94, CV99, CC91, CHC97, CA91, CK92b, CGSD93, DM98, DMM95, FFNS99, Fie97, GPMF96, GLWH90, GAA93, GAA98, GLM92, Han96, Hil90, HRD^{+}95, Hin94, HZF96, HGR95, JAMH96, JA92, JD90, Jør90, KSC^{+}90, Ker95, KHK97, KJ94, KSCG98, LPJ92, LKSA97, LLB^{+}96, MW98a, MSS^{+}95, MB98b, MB99, MBSW98, OBL^{+}91, PL99, PWM^{+}98, RPRNR94, PJ91, PHX92, RGR^{+}98, RH93, RNR98, RA98, RMT96, Roc98, RT92, RW96, Rut92, RRPP^{+}94, SSE90, SH91, SA94, SDJK98, SAK91, SDB97, SMV98, SS99c, SHM96, SBGBC96a, SBGBC96b, Sta91, SSM^{+}95, SK90, SLA93, SM99b, SGSA92, SAG98, SK95, SKR^{+}91, TCCA93, TC96, UPP95]. sediments [UBBS94, Ver97, VBvG91, WHK^{+}93, WF90, WH94b, YK95, YH92, dBGEK97, dHK98]. sediments-Comparison [BBH^{+}92]. see [GSG94]. seep [PY99]. Seepage [BM92a, Les95, Boy94, CCC^{+}96, HK98b, MKBN99]. seeps [BKB97]. segregating [HN94]. Seiche [OYWK96, LJH91, MWI92, RSZ97]. Seiche-induced [OYWK96]. Seine [GBC95]. selected [CDP96, GAAG96a]. selection [App94, CŠ90, DW99, HM99, JD95, MC95, MLB99, SGG^{+}99, TYL92]. Selective [AMP98, MHIHH99, MC96, MBSW98]. Selectivity [Som97, BM99]. Self [AK97, ADC92, GD92, GP94]. Self-regulation [ADC92]. Self-shading [AK97, GD92, GP94]. semiautomatic [RS91]. Seneca [WPAA95]. sensed [BM92b, ZKM93]. sensing [Bar99, JJGS96, SCF99]. sensitive [KT92, KJ92, LWS^{+}96, MMG94a]. Sensitivity [BMC^{+}96, RM96, VV98, WB96, DF94, FBCS90, HRD^{+}95, NCD98, Sta90a, VWF96, LV98]. sensor [FNWK96, KHK97, Old94, SS99d]. Separating [Gla91, KSYK96]. Separation [SAK91, FK98, PPK^{+}99]. separations [LPA94]. sequence [TMLS99]. sequential [Rut92]. series [AS94, CH94, NTH^{+}91, SGB^{+}97, TBM90]. Service [Wet93]. seston [EM93, HKvS^{+}92, KG95]. sestonic [CM91, UW93]. Settlement [PB93, TZFP^{+}94, WLBQ95, AWL94, Eck90, JB94, Joh94, LJA91, Wal92, ZFT94]. settlers [RERB91]. settling [BL94a, HTSDZ99, KG95, LSS^{+}97, NSJ90, SGG^{+}99, WHM97]. sex [Cra95, Sve97]. sex-specific [Sve97]. sexual [Den97]. Seymour [Wri90]. shad [SVW^{+}97, TJC^{+}98]. shading [AK97, GP94, GD92]. shallow [Bac99, BPS94, CAS93, FZD93, FSGG93, JA92, JM98, KV99, LHS90b, Mac93b, MP92b, MBT^{+}98, MMG94b, PASC95, Pra90, RHH97, RLH95, Som89, UNN^{+}98b, VMH^{+}97, WKK^{+}95, Kim92a]. Shallow-water [Kim92a]. shape [PRWG97]. shaped [MMCL94]. shear [KBJ98, Ken93, LCG94, SJL95]. sheath [CBFC98]. shelf [AD92, BOB94, BG97b, FBR91, Fur91, HLL98, MJC^{+}98, MPS^{+}91, PBL99b, RKG97, Sam91, MR99, RS98]. shelfbreak [RYC99]. shell [EG91b, FRLS90, GAA98, UMS^{+}99]. shell-bearing [GAA98]. shell-bound [UMS^{+}99]. shellfish [KDB91]. sheltered [HSL^{+}97]. Sherr [Duc94]. Shield [FHKC96, RKR91, SSM^{+}95, SCEG95, WMC^{+}96]. Shift [Gar91b]. Shift-up [Gar91b]. Shifts [HKvS^{+}92, LJH91, ŠKN^{+}99, MCB90, PS92]. shikimate [SRLRL^{+}99]. Shimizu [Sel94]. shipboard [CBFC98]. shoreline [LS93]. shores [Pai96]. Short [CCL90, DY99, FNWK96, Nod97, BLDP96, DD92b, HPTC97, OR94, Pla98]. Short-range [CCL90]. Short-term [DY99, FNWK96, Nod97, BLDP96, DD92b, HPTC97, Pla98]. shrimp [BW99a, BLK95, CBS99, FH92, FR99, PRCV98]. Si [BP97, LLB^{+}96, TLvB^{+}91]. Siberia [Liv99, SS98]. sic [Wet98]. sidearm [MIM90]. Siderophore [HRF91, RB91, WT94, WMT96]. siderophores [GP99, WMT98]. Sierra [WBM93, WB96]. Siesser [Paa95]. signal [DBZF98, GRG98, JD96a, RA98]. signals [Ege95, PJ91, RS98, SM98b]. signatures [CWSX96, Cla94, GSC^{+}96, MVM97, MSK^{+}97, TJC^{+}98]. Significance [BH96, KCKC90, MDC95, AW93, BV95, GSL97, MPT99, PFDH93, YK93]. significant [CCL90]. Silica [BAO97, BPC^{+}97, BFG94, BP97, Con97, KFEC91, SBPV99, TLvB^{+}91, VL91]. silicate [Mar94, Som94]. Siliceous [SWSC90, SLC^{+}99]. Sillaginodes [JB94]. silt [VdHH^{+}98]. similar [FN95, LVvdB99]. similarities [Har96]. Similarity [SKR^{+}91]. Simple [McB92, ADS94, DK93a, Fil91, GC90, KHK97, KC92, LCG94, SSE90, THK^{+}99, WHB95]. Simplicity [Kim90]. Simulated [SHF^{+}96, SZA92, dABL91]. Simulating [BSC93, ZCGP98]. Simulation [SFW^{+}95, MSH96a, RLH95, RGH^{+}93, Ver91]. Simulations [PJSW96, BAH97, Gor91, MS97, SO91]. simulator [HHL98]. Simultaneous [DHWW97, Lap99, LBL^{+}98, KV99, LSB92]. simultaneously [SGC^{+}97]. Single [GW90, LK94a, DOC99, FMK^{+}99, SM97]. Single-cell [LK94a, DOC99]. single-particle [SM97]. Single-species [GW90]. Sink [VORJ93]. sinking [BPK94, CBC^{+}91, Gra91, Han96, HSM^{+}98, WTH92]. sinks [LAC^{+}96a]. siphonal [MKT^{+}90]. sires [HAL90]. Site [TV99, CJVD97, HM99, Kli92, MKBN99, PRCV98, WMHB95]. Site-specific [TV99, WMHB95]. Sites [App94, dARB91, HSL^{+}97, WPBS99]. situ [Ano93x, AD92, BRZ95, BPA96, CH94, CFT92, DHSS95, FYG98, Fur91, GJ90, HTSDZ99, HJ94, JCH92, KW98b, KF93, LSB92, LS92, LV97, ML90, PSO91, PCFS97, PZ93, SHK95, SGG^{+}99, TT97a, VBvG91, VHS^{+}95, WTH98, WF90, WLD91, dABL91]. six [CT93, KLB^{+}99, MHZ98]. Size [DAC90, FVL98, FVL99, HNS95, MV96, Pro90, WL94, AB96, Ber99b, Bre91, CACM98, CDPR98, CŠ90, CPCS93, DDC92, DD92a, Ebe92, ECA99, FHKC96, FL90, Gae92, Gae93, GGA91, GU94, GL94, HBH94, HBH97, HPJW94, HVHH91, JF96, JF99, LV98, LPA91, Li90, LDL91, Log93, LPA94, MC95, MT94, OCZA90, PW93, RV99, Rot90b, RG92b, Sha94, ŠKN^{+}99, SB97, SSS93, TYL92, TZGS98, TL94, TKL^{+}97, VPBD97, Whe92, YFF^{+}98]. Size-ascent [MV96]. Size-dependent [FVL98, FVL99, Rot90b, Whe92]. Size-fractionated [HNS95, Pro90, WL94, TL94]. sizes [LB96, Rot90a]. Skagerrak [GA96, JD96a, JD96b]. skeletal [VSBZ^{+}94]. Skeletonema [NS93, PVL^{+}96, SZA92]. slightly [BVSBB93]. slope [AD92, BLDP96, MCC^{+}96, PRT^{+}93, WS90]. Small [ASC94, And90, Ano93j, BMT^{+}96, CCK98, CS97, DHK^{+}96, Guy91, HA92, LAYE95, MIM90, PSH^{+}96, RC93, RSZ97, SJDJ94, SBGCC91, Ver99, VCT^{+}99, WL92, WLD91]. Small-scale [ASC94, Guy91, WLD91]. small-volume [HA92]. Smayda [Sel94]. Smearing [CC91]. Smith [Cha96, Edm96c, War90, HR90, Rot91]. Smol [Leh95a]. snow [AGGD90, BAO97, FCC99, GS93, GSL97, HC97, LW90, PW93, Rie92, SE90, SW96a]. Society [Ano90a, Ano90i, Ano91f, Ano94]. soda [Ver99, VCT^{+}99]. soft [CC91, CHC97, EG91b, FGB95, FYG98, HTW90, RHT92, SGG^{+}99, Ver93]. soft-bottom [Ver93]. soft-sediment [SGG^{+}99]. soft-shell [EG91b]. soft-water [HTW90, RHT92]. Sohn [Arc93]. soil [ND91, RC98]. soils [Mat93, ND91]. Solar [De 93, MJC93, Ban90b, GMM96, GC90, HV98, HHL98, MZW^{+}95, ORH99, VRL^{+}98, VGB94]. solar-stimulated [GMM96, HV98, VGB94]. solid [LBL^{+}98]. solid-state [LBL^{+}98]. solidungula [HD97]. solubility [GG92, KNM96]. soluble [Cam94, TZFP^{+}94]. Solute [MW98a, AA92, DMF^{+}98, LMH98, PRT^{+}93, Web92]. solutes [QWR^{+}92]. solution [Swa99]. solutions [AWL94, Arm94, BFG94, SWH92]. Some [VPBD97, BSO92, IMB^{+}99, NML95, WCT90]. Sommer [Pra90]. sorbed [Tay95]. sorption [Bal96, SM99b, TBM95]. sorted [UC98]. sorting [Li94, RC93]. sound [GWY^{+}92, GWMB91, JX94, GAA93, MZ90, PL99, PK94, SLA93, WLY93, WGS97]. sound-scattering [GWY^{+}92, GWMB91]. source [AECP95, Ano93v, BW90, CCL90, DK96, FCG94, SBPV99, SBZD97, TR98, TC94, WBG^{+}91, BW94b]. Sources [CH92a, HKPM99, LTPB98, SCEG95, CW98, CCR^{+}95, DK93b, EF95, Els98, GNL97, HTW90, KCC98, KM95, KKS^{+}99, LAC^{+}96a, MV98, MTLR98, MHT99, Pae97, PH97, SR92, VV98, WC91]. south [CPLD99, PD90a, WBKF95, HC96, HCP90, LvGLC98, MJC^{+}98, WPBS99]. Southeast [HKPM99, FFNS99, Lap97, MBK^{+}99]. southeastern [BFW92, CPMW99, CM91, Eli98, MCB90, MPS^{+}91, SW93]. southern [Ber91, Liv99, MGN92, Mei90, RV99, VWF96, CM94, LBL91, MBHH^{+}91, NS91, RGR^{+}98, SO91, Smi96, WG94]. southwest [Nod97]. southwestern [HCLD95]. sp [DO98, HHB93, OS91, PSEJ99, PSJ99, TV99]. space [PHMH96]. Spanish [KM95, MD97]. Spartina [BFS91, LKD99, WH94c]. Spatial [Fra97b, KW93, LF91, LMH98, MCC^{+}96, MCD^{+}96, MKWD97, Sti92, FHF99, HPTC97, KJLA98, KJS98, MOF90, MBSW98, PPK^{+}99, RG92b, SA94, SW96b, SBWB94, SBGBC96b, Ste94, VV98]. Spatially [BAH97]. Spatio [HWP95]. Spatio-temporal [HWP95]. spawning [HC96]. Specialist [HDP^{+}90]. Speciation [DBT91, MBCB97, PM90b, SGWS96, SH90, XS93]. species [And97, BH95, BZR99, CWW98, ČH99, CT93, CM95, DZC91, DHB93, Dod92b, Dua92, FL90, GW90, GAAG96b, LC93, LB91, MS93a, MBOCK95, NML95, Rot90b, RFHG99, ŠBM^{+}95, SH97, TH92, Ven97, VAH99, WMHB95, WBG94, YK94]. species-specific [CM95, MBOCK95]. specific [BTLC93, BPCL97, CM95, CW95, DW95a, DW91, FR91, GW93, Goe98, GG90, LLLB97, MFLR96, MBOCK95, PPK^{+}99, Sam91, SD95, SD94, Siv93, SW91, Sve97, TV99, WW95, WMHB95, dGF96, DW95a]. specificactivity [RLL93]. specifying [Ano93p]. spectra [GSG94, HS92, JPJ97, Str90, SCKF99, VPBD97]. Spectral [BS90b, KJ92, ABB^{+}95, BM92b, GCP90, GC90, LPJ92, MUGK92, PWM^{+}98, Vod92, Vos92]. spectrally [And93, NS93]. Spectrofluorometric [Hen98]. spectrometry [KLB^{+}99, KSCG98]. spectrophotometric [PA95a]. spectrophotometry [RCCdG99]. spectroscopy [CIM92, HGR95]. spectrum [Ano93y]. speed [CW93a, GW99, Gli98, LPA96]. Sphagnum [NWS94]. sphere [KZ92]. spherically [Dod92a]. spicules [KFEC91, VL91]. spillover [RLTB90]. spination [Ste90]. Spinrad [Str94]. spiny [AB99]. sponge [SJP94, KFEC91, PCFS97, PPS^{+}97, VL91, PPS^{+}97]. sponges [Paw98]. Spongilla [SJP94]. spp [MHMD99, SCKF99, SCF99, Wym92]. spread [MVM98]. Spring [SP90b, CDHS98, DK93a, DDC92, DD92a, GJ96, HHR94, KOWH99, LVM^{+}90, MS93a, MMTJ94, MTM^{+}98, MHIHH99, NTH^{+}91, NHJ91, NCD98, RB94, RYC99, SGZ93, SN90, WMPC^{+}90, WBP99]. springs [FCM^{+}97]. spurious [Ber97]. Squeeze [BBH^{+}92]. Squeeze-water [BBH^{+}92]. St [BTL^{+}95, HL92, KBMSP99, OR94, SS99c, WMD90]. St. [MF99, PLTL96]. Stability [SL98, Ano93l, HPTC97, Mol98, PPK^{+}96, WJRB97]. Stable [Ben94, BW94b, BdBC99, HB94, KCC98, LKSA97, MB90a, TMLS99, Ura90, YMT^{+}99, Alt90, Ano95e, BLK95, CWSX96, Fra95, KSH96, MVM97, MA96, MMH^{+}95, RC95, WKK91, YKM95, ZEG^{+}94, Rav95]. stage [Rom91]. stages [HHR94, MHMD99, TGG94]. stagnant [HCS95]. stain [dGBPP96]. stained [LJD95]. staining [XB97]. stand [RJC^{+}97]. standard [CdG92]. standards [TG94]. Standing [VTCB96]. Starvation [Duc98, CMPD92, Pla98]. Starvation-survival [Duc98]. starved [Jan93]. state [Arm94, BH95, CDPR98, DM98, Gei90, LBL^{+}98, Nür99, SSM^{+}95]. Statement [Ano90a]. States [SW93, ES96, SHF^{+}96, SHKP96]. station [FBS99, DK96, Ven97]. stationary [LP94]. Statistical [Sch92, JJGS96]. statistics [Ber97]. Status [Cul95, Ano93m, Cla94, EG91a, JEP^{+}99, Pra90, Som89, WMD97]. Stauffer [GLH91, NKMO91]. Steady [Arm94, BH95, Gei90, KBJ98]. Steinberg [Web95]. step [BM99]. steps [TvE94]. sterols [EMH95]. Stewart [TMB96]. Stilgoe [Kim92a]. stimulate [SS99b]. stimulated [GMM96, HV98, VGB94]. stimulates [SRLRL^{+}99]. Stimulation [BVSBB93, LK94b, Rot92]. stock [VTCB96]. stocking [MSP^{+}99]. stoichiometry [HCSE97, HCH93, HS99, MDE97]. Stokes [GD96]. stone [MPT99]. Storage [MMPS98, Sos99, GLWH90, GHT^{+}96, Sta91, WH92]. storm [RG92a]. strain [Gil94, Gil96]. Strait [MTM^{+}98, FAE99]. Straits [Ohm96]. stranded [MJG91]. strategic [RGH^{+}93]. Strategies [LB96, AB99, DL96, FGN93, HRF91, Mal98]. strategy [But95, NT93b, VGP^{+}92, WW98]. stratification [DK93a, HGS94, MJOC93, WKK91, JM93b]. stratified [PAS90, KSSP91, LJH91, MFH^{+}93, MPA94, Nür98, REL97, SI98]. Stream [PTJ96, BDV99, Ber99b, BH91, Fin95, FHS98, FDG99, HPM96, HCJ96, HB91, Hur96, JM98, MPT99, Mul92, NCRS90, NML99, SC90, TBM95, TMB96, Wal96, WB96, WMML98, UC98, RW95]. streambed [SA94]. streams [CPT^{+}94, CNSR90, CMN93, DMF^{+}98, ES96, MT95, MMW^{+}97, MMW^{+}99, RS99, SBP^{+}96, TSD97, VMDC96, WCW^{+}91]. streamwater [FSGG93, GK97, VVK97]. strength [AGGD90]. stress [BF98, CACM98, DBS99, LK99, Les96, MB96, Wet95]. stresses [FMK^{+}99]. stressor [BSG^{+}99, MLLR99]. stressors [FCMB99, HC99, LMSP99, PLP99, RFHG99]. Strike [TV99]. Strombidium [OS91, DŠ97]. strong [HFBB99, MB96, XKS95]. structural [GHT^{+}96, PPK95]. Structure [Eic94, RW95, AS91, CNV94, DAC90, DAC92, Dum98, GGA91, Goe98, GWY^{+}92, GWMB91, GW95, HC97, LBH^{+}93, MCC^{+}96, MT94, MDWS97, NSBB98, PMF90, PG97, PCC^{+}97, Tho99, TZGS98, TKL^{+}97, Ura90, Ven97, VFW^{+}92, WDGS92, WBG94, WHOL99, ZKM93, dGGV^{+}96]. structures [HBI95]. structuring [CRKS93]. Student [Ano92e, Ano93a, Ano95c]. Students [Ano90h]. studied [KJLA98, LF91, NCRS90, Pol95, SR92]. Studies [LBL91, BLK95, BPA96, CBS99, FPC99, KYM97, LTPB98, LC90b, LKSA97, LDC99, Liv90, LAC^{+}96a, LAC^{+}96b, MQ96, NS93, Pse92, PC97, SS98, SK96, Wet91d]. study [Ant99, AR94, BW94a, CHC97, DG99, DHP^{+}96, DF94, EW99, Gae93, GGJ97, HCHD94, KY97a, KOWH99, KLR94, Kir94b, KWF90, LJBTB99, LHM^{+}93, MBT^{+}98, MR92a, MSK^{+}97, MKT^{+}90, MWW97, OMk93, OGH^{+}96, Pse90, QLBAP96, RGH94, SJDJ94, ŠBM^{+}95, SS93, TJK^{+}97, TTS93, UA97, VFW^{+}92, VVIK99, WH91, WWW98, ZR98, dHK98]. studying [Ant99, CR93a, FBCS90, WO98]. stylifera [IP93]. Stylophora [FPAG^{+}98, SRLRL^{+}99]. subalpine [WT93]. subantarctic [MR92b]. subaqueous [MCED94]. subarctic [Ban90d, Coa91, KKW90, MP92a, MFW^{+}91, RM92, SW91, WGSP91, WK90, WWMI98]. subcontinent [Wet91a]. subestuary [GJC92]. subglacial [CPT^{+}94]. subitaneous [LHS90a]. subject [MBCB97, Swa99]. sublittoral [SAG98]. submarine [HLL98, Tho99]. submarine-canyon [HLL98]. submerged [BK99]. submersed [CK92b, EW99, NK99, RJC^{+}97]. Submicrometer [YFF^{+}98, Ano93j]. subsequent [Bal96, SH98a]. Subsidence [PD90a]. substance [FT91, Mac93a]. substances [CL91, De 93, KZM90, KHS97, MH94b, NE93, Sha94]. substitution [YM96]. substrata [AML93, Wal92]. substrate [FFF98, FL95, NML99, YD99]. substrates [Ber99b, BL98, WHB95]. substratum [CW93a, HQ90]. subsurface [Fin95, GW90, JM98, KHK^{+}98, WSMS96]. subthermocline [LWMD96]. subtidal [Liv90, Wal92]. subtropical [CNP94, DK93a, EK90, HGS94, KL96, McK96, PASC95, PBL99b, YSSO91]. success [CF96, GH95, SWS^{+}96]. Successes [HCPR99]. succession [GBC95, KKK^{+}93, LP95, SLD97, SWSC90, TGG94]. sugars [FR99, NVH^{+}94]. suggests [Hen98]. sulcatum [DŠ97]. Sulfate [FFNS99, UBBS94, GLS90, HB95a, LB94, MHSP90, Muc91, PPB91, RRK94, SW96b, WR95]. sulfate-reducing [GLS90, HB95a]. Sulfide [RT92, DGW92, HCHD94, KMM90, LKD99, LB95, UA97, VBvG91, WCD^{+}94, WS96]. sulfide-induced [KMM90]. Sulfur [GLWH90, NWS94, SM90, Sta91, PAS90, BEB92, CTF98, HCLD95, Jør90, OBH^{+}91, OCP92, OBKH96, RKR91, RRK94, SFRB98, WTB98, ZFdSY96]. sulfur-disproportionating [CTF98]. sulfuric [RKST90]. sulfurous [GGA91]. Sulphur [Kim92b, KHS^{+}98]. sum [JF96, JF99]. Summer [LCR99, Ano93x, CCL92, Han91, IPPD92, MLR^{+}94, ŠBM^{+}95, Ven93, YK94]. Summerhayes [Cha96]. Sun [HM99]. sunlight [WWSJ98]. sunscreens [GP94]. super [WMM^{+}94]. super-oxygenated [WMM^{+}94]. superba [Dal90]. Superior [KSH96, KLR94, MKWD97, TH98, VFW^{+}92, WH95]. superiority [GKM94]. supersaturation [PDM96]. supplementation [RBMH91]. supply [BC97, BFGS94, DA92, FL95, FSSI97, HD97, Lev96, OR94, PAD^{+}99]. support [Hai96a]. supported [BPS94]. supporting [SCEG95]. supports [CACM98, MLR^{+}94]. suppression [RW96]. surface [BB98, BW94b, CFW97, Car98, CPLD99, DOC99, FFNS99, Fin95, FGN93, GSG94, HG99, IM98, KM94, Ken93, LM98a, Li98, MT90, MPT99, MS99, MH94a, MZ90, MLB99, NSJ90, OHMK91, PL94, PRT^{+}93, Rie92, Sti91, SGWS96, TS97, THR97, TA97, TCA96, VMDC96, VPN^{+}90, VHS^{+}95, WCD^{+}94, WMHB95, WBM93, WGS97, YK94, ZMP98, Ben94, HB94]. surface-active [ZMP98]. surface-deposit [KM94]. surface-subsurface [Fin95]. surface-water [BW94b, Ben94, HB94]. surface/volume [NSJ90]. Surfaces [dARB91, JAS95]. surficial [CA91, CGSD93, PK91, SDJK98, SK90]. surrogate [MO91, Sch92]. surrounding [FCM^{+}97]. survey [BLD99, HWP95, WPBS99]. Survival [MC96, Bur92, DZC91, Duc98, GAA98, GH95, Har96, KSSP91, MC97, NT93b, PK94, RKW94]. susceptibilities [BOFW93]. Susceptibility [Gil94, HDP^{+}90, LMSP99, WCT90]. suspended [BW94a, Bur92, CGB90, CH92b, HSM^{+}98, KW98b, LHS90b, PSS^{+}96, TG92]. Suspension [TZF96, Ano93j, CH94, GW99, Joh90, Jør93, KK91, KHB96, LPA96, NT94, PF90, SJL95, WMTB93a, WSNM98, WMML98]. suspension- [WSNM98]. suspension-feeders [LPA96, WMML98]. suspension-feeding [CH94, KK91, KHB96, PF90]. Sustainable [Eli98]. Sustained [OR94, Rie92]. Sverdrup [NS91, SP90b]. swamp [AB94]. swamps [HCS95]. swarming [BPA96]. swept [Men90]. swimmer [Coa90, HN94]. swimmers [LHWZ92, PD90b]. swimming [DL92a, GD96, MK95, Sai94, vLM91]. Switzerland [AHKH^{+}96, SFRB98, KS93, TMLS99]. symbiont [UMS^{+}99]. symbionts [AMP98]. symbioses [Hay96, RC95]. symbiosis [Ano93u, FRY95]. symbiotic [CMPD92, SJP94, LGI94, Les96, PY99, RCG98, SJPGH97, SRLRL^{+}99, VDG^{+}97]. synchronous [PC97]. Synechoccus [Wym92]. Synechococcus [CJD^{+}99, DŠ98, DŠ99, FCI91, Li98, Mor97, OCZA90, PBL99b, WPB94, WMT96]. Synergism [FCMB99]. synergisms [PPP97]. Synergistic [HD95]. Synoptic [OGH^{+}96]. synthase [Ano93m]. syntheses [Wet98]. Synthesis [Cot99, GW93, LBB^{+}97, TB91]. system [Ant99, AR94, BOCL^{+}95, BW94b, BPC^{+}97, CTP^{+}95, HPTC97, HB94, HGFC96, KV99, LO97, RV99, SBGBC96a, UMS^{+}99, Wet95, WBRL98]. systematics [SS98]. systems [BP91, BGC99, BG94b, EW99, FDG99, Gar91b, GV93, KHK97, MB98a, PL95, SMH94, SB90]. SYTO [dGBPP96]. Szekielda [Abb90].

T [Ano93f, Fen94, HL94, Jum94, Sel94, SW93, Thi94, War90, Won91]. table [AO96]. Tagging [AMF98, LHM^{+}93]. Tahoe [JGRR99]. Taiwan [KL96]. Tanganyika [ESC^{+}93]. tangential [ECA99]. tangential-flow [ECA99]. tank [Ant99, HW94]. tannin [HSRR96]. taphonomy [Whe91]. targeted [LK94a, SMB^{+}97]. tarsalis [BK90]. taste [KK91]. Taxon [WW95, Goe98]. Taxon-specific [WW95, Goe98]. Taxonomic [HB95b, LV98, Str97, WMD97]. taxonomy [GB93]. teaching [Kal91, Wet91c]. technique [CLP^{+}92, CG91, EF95, HC97, Jön91b, Roe98, SHK95, WDGS92]. techniques [KSH96, NT94, NCRS90, Ano93f]. technology [GWMB91]. Temora [Ano93m, IP93, PK94]. Temperate [TH98, BP96, DHK^{+}96, DK93a, HK98b, KOWH99, KBB98, Lev96, Lew90, MSM96, Mul92, NK94, OW96, PMF90, SF95, SD95, SAG98, TN94, WMD97]. Temperature [Ano93y, BN90, HQ90, LG99, SD94, CKR93, CK95, DD92b, ESG91, GMK98, GAAG96b, GRG98, Han91, JA92, Kap92, KOWH99, KLB^{+}99, KAB92, MVT99, MM93, MCCD92, ML90, Rah90, SBB^{+}93a, SHF^{+}96, YD99]. Temperature-dependent [Ano93y, MCCD92]. temperature-resolved [KLB^{+}99]. temperatures [AR95, FSSI97, Les96, Liv99, MF99]. Temporal [FCC99, JB94, LBH^{+}93, SA94, WM92, Ano93u, GAA93, HWP95, JGRR99, LF91, LMH98, MTM^{+}98, MCD^{+}96, PPK^{+}99, REL97, RG92b, SW96b, Sti92, VV98, WMC^{+}96]. temporary [HJ94]. tension [FL90]. tentacles [RBMH91]. tenue [FSSI97]. TEP [PA95a]. terebellid [Mar94]. term [BSY^{+}95, BLDP96, DG99, DY99, DD92b, FNWK96, FHN^{+}96, HGS94, HPTC97, HC99, JCH92, MR98, Nod97, Pla98, PJSW96, Pse90, SH96a, SHF^{+}96, SB94, VCT^{+}99, WH94b, Won91]. terminal [CCK98]. Terminos [RMT96]. Terrestrial [VBD^{+}97, CE96, Kim92b]. terrestrially [MPS^{+}91]. terrigenous [OBA99, PHX92]. tertiolecta [SGC^{+}97, SM91]. Test [Fri90, Bre93, CL90, FS99, LPA96, Lou90, MDE97, MDdM^{+}99, PLP99, Som93, Som95, UW92]. Testable [Sar99]. tested [PC96]. Testing [YD99]. Tests [AS94, FSS94, Whe92]. testudinum [FZP92, LD99]. Tethys [Ver90]. Texas [BMAW97, Ste94, SGWS96]. Th [LSS^{+}97, WM92]. Thalassia [FZP92, LD99]. Thalassiosira [CH96, LTPB98, LRM95, RSK97, RM90]. Thau [VTCB96]. thecate [SB93]. Their [Edm96a, Man91b, Sma97a, AML93, BKT96, Ban90e, BGF92, CPLD99, Eic94, HBH94, HDP^{+}90, LHWZ92, ME98b, MLB99, Paw98, RJ91, SC93b, SE90, SLC^{+}99, Str97, VBK98]. Theoretical [CP97b, Log93, MaB98, Roe98]. theories [Joh91]. Theory [MVM98, Ano93c, DO97, KTF96, Lou90, Pol90, WSMS96, Alt94]. theperiphyton [EM93]. there [GA96]. Thermal [MT94, DMA^{+}96, ES96, MMTJ94, SGZ93, Tho99]. thermo [BMT^{+}96]. thermo- [BMT^{+}96]. thermohaline [AS91]. thiebautii [Kan93, VC90]. thiols [LVvdB99]. thiosulfate [Jør90]. Thornton [Thr90]. thraustochytrids [Ano93o]. threats [Mal98]. Three [CP94, SO91, TGG94, BH95, CK95, DZC91, DG99, KZ92, LB91, MA94, MH97, SGWS96, VDG^{+}97, VAH99]. Three-dimensional [SO91]. three-layered [KZ92]. threshold [Ebe92]. Thresholds [SR94, Lap97]. thymidine [Bel90, GLS90, KV99, TB91]. Thysanmssa [HK98a]. tidal [Ano93v, CCP92, LWMD96, LL95, Pin99, RJC^{+}97, SD95, dJvB95]. Tidally [KBB98, CP99, KMU95]. tide [AWKH98, BMAW97, HGP97, JPJ97, LR99, MSK^{+}97, TSV^{+}91, TS97, dJvB95]. tide-induced [dJvB95]. tides [BL97]. Time [RLL93, AS94, CH94, DC96a, NTH^{+}91, SGB^{+}97, TBM90]. Time-course [RLL93]. time-series [AS94, CH94, SGB^{+}97, TBM90]. times [MM91]. Timing [GJC92]. tissue [HDR^{+}98, TLB99]. titration [Muc91]. TO-PRO [LJD95]. tolerance [Han91]. Tomas [Boo94]. tomography [VSBZ^{+}94]. tonsa [Ano93m, EMH95, KBH98, PWR94, STJ^{+}93, Sai94, Tis92, TJK^{+}97]. tool [HGR95, KMK95, KW98b, SCFB98, VSBZ^{+}94]. top [ADC92, Lap99, Leh91, MMC94, PS92]. top-down [ADC92, Lap99, Leh91, MMC94, PS92]. topography [CW93a]. total [ES90, GH95, LF91, MMH^{+}95, VFAA98]. TOTO [LJD95]. Toxic [Jun92, TT97b, And97, DZC91, Gil94, Gil96, RKW94, Sel97, SH98c, Sel94]. toxicity [GDG99, HCHD94, SH96b]. Toxin [TOKA97]. toxins [DZC91, KDB91, Plu97]. Trace [AMM97, TJC^{+}98, AMF98, BMP92a, BW94a, BSG^{+}99, BDH91, DL96, FCM^{+}97, HTW90, HTSDZ99, HT98, HGFC96, MDC95, MLLR99, MHP91, RF99, TC95b, TCA96, WRLF96, WF96, WF98, WBKF95, Whe91]. traced [AHKH^{+}96]. tracemetal [RS99]. tracer [ASMB98, BW94a, BDY^{+}98, BP97, LAC^{+}96a, LAC^{+}96b, SK96, SB94, Swa99, TMB96]. tracers [BL94a, CCR^{+}95, LSB92, MCED94, Alt94]. tracing [KKS^{+}99]. tract [Bre97]. transect [RGR^{+}98]. transecting [SGB^{+}97]. transfer [ADS94, BA97, BA92, BA95, BS90a, Dad93, EMH95, ESG91, Fri90, HCS95, JHM96, MS97, OMK95, Sha94, SFC90, TA97, TCA96]. transformation [LGT95]. transformations [BB98, HA90, dHJS90]. transient [EKT99]. transition [RYC99]. Transitions [SWJO98]. transmission [BSG^{+}99, MLLR99, SBW^{+}98]. Transparency [VRL^{+}98, JGRR99, MH97]. transparent [GSL97, PA95a, SFW^{+}95, ZMP98]. transplant [Sam91]. Transport [CMN93, HC96, TSV^{+}91, AB99, AA92, BFW92, CCL90, DT91, EG91b, FBR91, FPZFW99, HPM96, HM90, HCP90, Log95, MMS93, MW98a, Pal92, PJ90, PPK^{+}96, RS99, SSFP90, SG93, TS97, WS90, WLBQ95, Won91, Wri90]. transversal [RSZ97]. trap [Coa90, HTSDZ99, HN94, HSSA94, LHWZ92, Nod97, PHT^{+}93, Set92]. Trapped [FLH96]. trapping [Kit96]. traps [FNWK96, HHR94, PD90b, YGB91]. travel [MK95]. trays [SGP92]. treatise [Jón95]. treatments [LHWZ92]. trematodes [SC93b]. Trends [Mal98, GA96, MF99, SH96a]. tributaries [FCM^{+}97]. trichloroethylene [AECP95]. Trichodesmium [Cur94, Kan93, SCKF99, SCF99, VC90]. Trichormus [vEJ97]. tricornutum [LBP97]. triggered [LP95]. triglycerides [KHS97]. tripton [PASC95]. triterpenoid [EMH95]. tritium [MA96]. trivariate [Li90]. trohic [JPG90]. Trophic [AC92, CK92a, EMH95, FDG99, KHP91, LL90, PCFS97, PPS^{+}97, PRCV98, Wic95, CL99, CPLD99, CFSW91, Cla94, DL93, DC99, EG91a, FFF98, Ham94, JEP^{+}99, ME98b, MPG95, NB97, Nür99, PFDH93, RB94, Sch92, SR95, SSM^{+}95, TCA96]. trophy [dGP94]. Tropical [Bie99, BUW96, BDY^{+}98, DHB93, DCM90, FPC91, Fur91, GCB^{+}98, Kit96, Lew90, LDL91, LDV^{+}92, MA94, PPP97, TGSB96, Ver99, VCT^{+}99]. trout [Hur96, MHM96]. Truckee [Ben94, BW94b, HB94]. tube [PB93, YGB91]. tubes [Liv90]. tundra [PFDH93]. tunicate [ADM96, GCD90]. tunnel [HW94]. turbid [CP94, CCP92, ES90, GCP90, Kir94b, LDL91, LVFD93, MaB98, SWJO98]. turbidity [AEL^{+}99, Ant99, KE91, LDL91, LDV^{+}92, RG92a, RCCdG99]. Turbulence [HSM^{+}95, HBL^{+}93, ASC94, Dad93, HvOW99, MMCL94, MH95, SJL95]. turbulent [BSC93, FPZFW99, LR99, MK95, Sai94]. turf [CW93a]. Turkana [RA98]. turnover [BS99, Fie97, KGFP91, LB94]. Twentieth [Fri90]. Twentieth-century [Fri90]. Two [HvOW99, Ano93i, AMP98, Ber91, CJD^{+}99, CMN93, Den97, HNS95, HC94, Har96, HDB93, JPJ97, Jun92, LL98, MAS91, MGWB97, RWH^{+}98, RKST90, SW96b, SM98c, Two94, TCB^{+}98, WPBS99, WS96, YAF^{+}92]. Type [CGB90, CH92b, MEM99, WMT98]. typhoon [RWH^{+}98]. typhoon-induced [RWH^{+}98]. tyrannus [CCK98, KDKS90].

U.S. [And97, BL97, CPMW99, DFN97, HGP97, MPS^{+}91]. U.S.A.

[SLC^{+}99]. ultrafiltration [ECA99, SIO98]. Ultraphytoplankton [LP95, Li94, Li97]. Ultraviolet [SRLRL^{+}99, WSM^{+}96, DC99, GP94, Gle93, GHT^{+}96, HHL98, KL90, KBMSP99, Les96, MJC93, MDdM^{+}99, NCD98, RHH97, SML96, VL96, VL98, WHB95, ZCGP98]. Ultraviolet-B [SRLRL^{+}99, MDdM^{+}99, HHL98, RHH97]. Unbalanced [DCML96]. unbiased [SGP92]. under-ice [Eic94]. underestimates [SSS93]. underlying [RT92]. understanding [Pal92]. underwater [Ban92, BSMK95, Kir91, MKBN99]. undulatus [TJC97]. unguiculata [KCKC90]. unicellular [WKZK98]. Unintended [JRM98]. United [SW93, ES96, SHF^{+}96, SHKP96]. universa [UMS^{+}99]. University [Ano94, Cot99]. unknown [MBKG95]. Unusual [SM92]. unvegetated [RW96]. upland [Mul92]. upon [Mor97]. upper [HP95, PLTL96, SBB99]. upstate [HPB^{+}99]. Uptake [CW92, Jan93, Ano93v, Arm99, AB92, AR96, BW94a, BTM92, BW99b, CHD91, DW95a, DW91, FVL98, FVL99, GGFR91, HLGA98, HFK92, HK95, Hof90, HZF96, IH95, KW98a, KKW90, Leb90, LG99, LvGLC98, MMW^{+}97, MMW^{+}99, NVH^{+}94, PBJD93, PVL^{+}96, PAD^{+}99, PSJ99, PSS91, Pro90, RDK96, RF99, SJDJ94, SCK93, SRG98, SS99c, SH95c, TBM95, TMB96, TA97, Ver91, Véz94, WH92, WHN^{+}98, WKK91]. upward [AK97]. upwelled [Pro90]. upwelling [BS90b, BPC^{+}97, DW95a, DWM90, FPC91, FGN93, Gar91b, HDZB98, KTMI^{+}98, MKWD97, RV99, SZA92, WLBQ95, WBRL98, Cha96]. upwelling-induced [KTMI^{+}98]. Urabe [Bre93]. Uranium [Ano93z]. urban [Kna99]. urbanization [HLQ^{+}99, MVM97]. urea [WHN^{+}98]. USA [WLCG96]. Use [CSPG94, CHC97, KF93, KHB96, MKBN99, RS93, SRG00, SPMW97, WHM94, BM92b, BC90, DC91b, DDG94, ES90, EK90, LL96, MO91, Old94, PH97, PH92b, RJ91, Set92, SLC^{+}99, SJHS98, VHS^{+}95, WF90, WK90]. used [CDP96, CCR^{+}95, GCD90]. usefulness [GP94]. uses [Thr92]. Using [GDG99, SH95a, TMB96, AS94, BW94a, BLK95, CCK98, Fra95, HY97, KTF96, KRB98, Leh93, MMTJ94, MSK^{+}97, NJLW97, QLBAP96, SMB^{+}97, Som95, SK96, SIO98, THK^{+}99, Tho99, VWF96, YD99]. usually [RSZ97]. Utilization [TG92, WSMS96, AB96, CGD90, CNL90, DM90, HHI96, JPJ97, KK93, MBHH^{+}91, NZHB95, OAF^{+}99, PM90a, SBB99, SH98a, TK90, VHH96]. UV [AOTR97, De 93, HDMB97, HM99, JLO97, KZM90, LV98, MZW^{+}95, MH97, NE93, VRL^{+}98, VHS^{+}95, Wil95, WHOL99]. UV-absorbing [JLO97]. UV-B [NE93, Wil95]. UV-driven [AOTR97]. UV-induced [KZM90]. UV-light [De 93]. UV-sensitivity [LV98]. UVB [SLD97].

V [Cha96, Edm96c, Jón95, Thi94]. vacuolate [PAS90]. vacuole [DŠ99]. validated [WHM97]. Validation [BM92b, CG91]. values [RC95]. valved [PHT^{+}93]. vanadium [SMC98]. vanhoeffeni [BDR99]. vanhoefleni [ADM96]. vapor [BW94b]. var [Gen91]. Variability [And90, BSG^{+}99, GSG94, LS93, MLLR99, Bel90, CM94, Cur90, FHKC96, Gae92, GL94, GKS^{+}94, JGRR99, JB94, JW92, KWF90, LBH^{+}93, SBWB94, SD95, TOKA97, TH96, TTS93, Véz94]. Variable [LPA94, NCD98, JEP^{+}99, MP92b, PCC^{+}97, SZA92, Wil92]. variables [LPA91]. variation [BUW96, CNSR90, DB93, Hay95, KM95, LF91, LKM98, MTM^{+}98, MW98b, RPRNR94, PSS91, REL97, Ura90, Vod92, VBD^{+}97]. Variations [BTLC93, CE96, KMC90, LID93, Rot91, ARM96, BBS97, GGA91, Guy91, Kli92, MC95, OGH^{+}96, PC97, RM96, SW96b, SHM96, SBGBC96b, SR93, TJC97, TV99, VWF96, WH99, WM92, WHM97]. varied [IKT99]. various [AM95, BdBNvB91, Cla94, FSSI97, Har96, LHWZ92, SB97]. varve [RLTB90]. varves [KLR94]. varying [Goe98, MT94]. vascular [MH94b]. Vechten [SBGCC91]. vegetated [Ano93p, RW96, WH94b]. vegetation [BGC94, NSZ97, Wet91a]. vehicle [SGB^{+}97]. velocimeter [FHF99]. velocities [Gli98]. velocity [AB92, CCK98, FPC99, PB93, TA97]. Veluwe [Wet95]. Venezuela [GCB^{+}98]. Venezuelan [Two94]. Venice [MDC95]. vent [CJVD97, PRCV98, TV99, dABL91, dARB91]. ventilated [PDM96]. ventilation [Mar94, SGZ93]. vents [VF94]. Vermaat [Wet95]. versatile [SJPGH97]. versus [CM94, CFT92, KSC^{+}90, LV98, LK94b, MHM^{+}99, MN95]. vertebrate [BOFW93]. Vertical [BH96, BOFW93, DFN97, DPC90, Mac93b, MMS93, MTM^{+}98, NK99, ZKM93, AO96, PAS90, BSMK95, DL92a, Dod90, DOC99, FH92, FR99, HCKW90, HPJW94, Hay95, HWP95, Jac93, JM93b, KY97a, KBB98, KM95, KTMI^{+}98, LBL91, Lei90, LW93, LJH91, Loo93, MLBC94, ML95, MWI92, NRR96, RF94c, RS98, RSZ97, RM96, Sar99, SE90, Tjo90, WS90, WKK91, WSMS96, YW93, ZCGP98]. vertically [KC92, LP94]. vesicomyid [BKB97]. vessel [SS99d]. via [Ano93r, HW91, KV99, LSB92, MM91, SAK91]. viability [IP93, MLBC94, VKB96]. Vibrio [HHB93]. Victoria [HBO^{+}94, LB94]. Vierssen [Wet95]. vinelandii [WH95]. Vineyard [MZ90]. viral [GHF^{+}97, MWB^{+}98]. virginica [CNL90, TZFP^{+}94]. Virus [XB97, CS95, PW93]. virus-rich [PW93]. Viruses [FN95, KWS98, HS95, TH98, WWSJ98, XB97]. visibility [EABK99]. visible [CSS96]. visual [THH98, TSH99]. Visualization [HSL^{+}97]. visualizing [LJA91]. vitamin [Hof90]. vitro [CFT92]. vivo [BTL^{+}95, MC97, Wym92, YM96]. Volatile [MHMD99, HCHD94, RKR91, RRK94]. Voltammetric [WTB98, LBL^{+}98]. voltinism [Hur90]. Volume [Ano90f, Ano91d, Kir91, BM99, FLH96, HA92, MBHT94, NSJ90, VRT^{+}92]. vortex [SC90]. vs [PAS90, DFA93, FZD93, Lap95, PD90b]. vulgaris [WWT^{+}93b]. vulnerability [GU94, THH98, TSH99].

W [CT96, Edm96a, Har95, Jum90, Men90, Paa95, Pri94, Str94, SW93, Thr90, Wal96, Web95, Wet95, CLKJ95, RDK96]. Wadden [BC97]. wall [VLHL97]. Ward [Jør93, SRG00]. warm [HGS94, Pin99]. warming [ARM96, CF96, ES96, MHM96, Nic98, PSH^{+}96, RKA96, SBP^{+}96]. washing [BLK95]. Washington [AD92, BMP92a, BMP92b, Edm91a, Ker95, PHX92, RJC^{+}97, Thr92]. Watanabe [Bre93]. Water [ESG91, Kit96, New99, PAS90, Ano93y, AMLH97, AB92, BMP92a, BMP92b, BKB97, BH96, BW94b, BRPR98, BDR99, BBH^{+}92, CR93a, CL91, CNP94, CRKS93, CLMH93, CM94, De 93, DMF^{+}98, ESC^{+}93, Edm96c, EBH92, FHKC96, FPC99, Fri90, GD92, GCD90, HLQ^{+}99, HTW90, HTSDZ99, Han91, HCS95, HBO^{+}94, HRD^{+}95, HB95b, HKP^{+}97, HG99, JRM98, JL94, JLO97, Kap94, KG95, KYM97, KCWM96, Kim92b, KJLA98, KDZD94, LCR99, LJBTB99, LWMD96, LDC99, LG99, LVM^{+}90, MAG95, MMG94b, MT94, MF99, MH94a, MMH^{+}95, Mol98, MKBN99, MB90b, NJLW97, NML99, OHMK91, PG97, PL95, RKA96, Ree90, REL97, RHT92, RERB91, RT92, SSE90, Sai94, SJDJ94, Sar92, SHK95, Sha94, SBWB94, SMH94, Siv93, SBB99, SMV98, SJHS98, SL98, SB94, SBGCC91, SRPG92, TA97, TG94, TvE94, Tra94]. water [Tra97, TZFP^{+}94, UA97, WTH98, WF90, WHM94, WFG99, WM92, WBM93, WL94, WGS97, YK94, ZCGP98, dABL91, dHJS90, Ben94, HB94, Kim92a]. water-filled [SL98]. water-level [Fri90]. water-soluble [TZFP^{+}94]. waterborne [ZFT94]. waterfowl [PL91, Sau92]. waters [AK97, AMM97, AA91a, AA91b, BN90, BPCL97, BRZ95, BL97, BDH91, BP97, CR93a, CW98, Cam94, Car98, CPLD99, CS97, CI94, CTSB98, CK95, CPCS93, CP99, ES90, FSS94, FAE99, FGN93, FAC^{+}99, GMM96, GS91, GB94, HS95, HK95, HS92, HV98, HFBB99, JLMV98, Jon91a, JRT96, KOWH99, KZM90, Kir94b, KBMSP99, Lap95, Li98, LM98b, LBL^{+}98, MaB98, Mal98, MGN92, Mar94, MMG94b, MM93, MPA94, MSSM96, Mei90, MUGK92, MZ90, Mor97, Mor90, Muc91, OB98, PH92a, PRT^{+}93, Pro90, Rie92, RGRW93, SdBdJL98, SML96, SSKW91, SMC98, SGZ93, SGWS96, SH90, TIM96, THR97, TM95b, TCA96, VPN^{+}90, VHS^{+}95, WTB98, WK96, WMHB95, WPB94, Wel99b, RW95]. watershed [BMC^{+}96, GJC92, JCMW91, KPSB95, WB96]. watersheds [BM96, MVM97, MA94]. Wave [Web92, DG90, HSL^{+}97, Men90, SI98, WLD93]. wave-induced [WLD93]. wave-sheltered [HSL^{+}97]. wave-swept [Men90]. waveguide [DSV^{+}99]. waves [TJ98]. way [LJA91, TIOS99]. Weak [FPL^{+}91]. weather [SI98]. weathering [CPT^{+}94, Sta90a]. web [BVSBB93, BKR^{+}90, BAH97, BSG^{+}99, BLK95, CL99, CFP97, CFT92, ES93, FCI91, FPC99, KSH96, LC93, MR98, MLLR99, MSP^{+}99, PCC^{+}97, SS93, TCA96, VL98, YMT^{+}99, ZEG^{+}94]. webs [CCR^{+}95, Fra95, MVM97, MF90, Pro90, TT97b, Wic95]. Weddell [BdBNvB91, NCD98, SN90, TLvB^{+}91]. Weddell-Scotia [BdBNvB91]. wedge [IPP97]. weight [BLS^{+}95, DD92b, GSC^{+}96, NML95, PL99, Str97]. weir [SWJO98]. weisflogii [CH96]. weissflogii [LRM95, RM90]. well [RM90]. well-defined [RM90]. west [HGP97, MMC94]. western [HSM^{+}98, HS92, Lap95, MKWD97, SM92, WL94]. wet [RS91]. wet-oxidation [RS91]. wetland [CTSB98, HSCM97, HC94, KMM90, MB99, RW96, SM90, SLT99, Wet91d]. wetland-derived [SLT99]. Wetlands [Kim92b, CDP96, LWS^{+}96, PJSW96]. Wetzel [McC97]. Where [VRHK99, RSZ97]. wherefore [Bou98]. Which [KAD95]. Whigham [Wet91d]. White [KMC90]. Whiting [SLSLBD97]. Whole [AR96, DC91a, LFHS99, LD99, NCRS90, PC96, RKST90, SKR^{+}91]. whole-core [NCRS90]. Whole-lake [AR96, DC91a, LFHS99, PC96, RKST90]. whole-plant [LD99]. whole-sediment [SKR^{+}91]. Wildlife [KSH^{+}99]. Wind [CAS93, HW94, MC92, dJvB95, Ano93k, CC98, CTP^{+}95, LHS90b, Ver94, Web90, WH94a]. Wind- [dJvB95]. wind-driven [CTP^{+}95, Ver94]. Wind-induced [MC92]. Wind-tunnel [HW94]. Winter [Paa95, VPN^{+}90, AR95, BMT^{+}96, DDC92, DD92a, KOWH99, LP95, UNN^{+}98a, WBRJ91]. winter-spring [DDC92, DD92a, KOWH99]. wintering [CF96]. Wisconsin [Cot99, ARM96, BEB92, BFW92, FMK^{+}99, HA90, SBWB94, UBBS94, VWF96, WMHB95, WKB^{+}96]. within [BAO97, FGN93, GGJ97, HBH97, LS93, NSZ97, PCFS97, PPS^{+}97, Web92, WHM97, WMD90, ZEG^{+}94, dHK98]. within- [dHK98]. without [SRG98]. Wollast [May92]. Wood [Edm96c]. wooded [HCS95]. Woodhead [Sut93b]. world [Edm96a, EBH92, LH91, Mal98, Sal92a]. worm [PB93, WSB98]. Wright [Web95, CA93]. WT [SB94]. Wyoming [Rah90].

X [CHC97, DGW92]. X-ray [CHC97, DGW92].

Yanaco [YK95]. year [Ree90]. year-long [Ree90]. years [LKSA97, MTLR98, Wet99]. yellow [FT91, KE91]. Yellowstone [IKT99, KTF96]. Yes [MBFG90]. yield [BTL^{+}95, GKS^{+}94, PH92b, SM91]. yields [FK98, LP94]. Yo [XB97]. Yo-Pro [XB97]. York [HPB^{+}99, Kna99, WPAA95]. Young [Str95].

Z [Sal92a]. Zagami [Ohm96]. Zanzibar [HDR^{+}98]. Zealand [Hur96, MPT99, MJG91, MT95]. zebra [RF99]. Zeitzschel [Cha96]. zenith [Ban90b]. zinc [Coa91, LBLO91, SH92, SH95b, SH96b, SH98c, XKS95, YM96]. Zn [HTDM96, LL98, RHT92, SH98b]. zone [AEL^{+}99, BDV99, BW99b, Dal90, DK96, DMF^{+}98, Eli98, FSGG93, Fin95, HQ90, KFW94, MD97, Mat93, Meh99, PCFS97, PK91, PPS^{+}97, RG92a, SM92, SCPS98, SN90, SBGBC96b, TGSB96, Was90, WBRL98, WMD90]. zones [Boy94, Gar91a, JB91a, MMW^{+}97, MMW^{+}99]. zoobenthos [Jón95]. zooplankter [HH93, SP90a]. Zooplankton [EG91a, Hai96b, HBH97, PH92a, Sar99, TTS93, AW93, AO96, ASC94, AH91, Bra98, Bre93, BP96, CFSW91, CLP^{+}92, CG95, CF96, CMT93, CT93, CRKS93, CH96, CH92b, De 90, Dod90, Dod92b, Fil91, GSL97, Hai90, HCSE97, HAL90, Hes93, Jac93, KBB98, KSSP91, KC92, Lei90, LV97, MAN91a, MDE97, MLBC94, MKWD97, Nie91, NK94, Ohm96, PWR94, QLBAP96, RF94b, RGRW93, RBH99, Rot91, SWS^{+}96, SZW95, Set92, SSP^{+}94, SHKP96, Str97, TN94, TT97b, Ura90, UW93, VBAO95, VVIK99, VAH99, WB92, WP94, WBG94, WCT90, WWT93a, WSMS96, WHOL99, WC91, dAL94, dGF96, LAC^{+}96a]. zooplankton-phytoplankton [HCSE97]. zooplankton-protozoan [CFSW91]. zooxanthella [FRY95, RCG98]. zooxanthella-giant [FRY95]. zooxanthellae [CMPD92, FGB95]. zooxanthellaefree [FGB95]. zooxanthellate [FPAG^{+}98, GJ90]. Zostera [SBK95].

References

Ammerman:1991:BNAa

[AA91a] James W. Ammerman and Farooq Azam. Bacterial 5’- nucleotidase activity in estuarine and coastal marine waters: Characterization of enzyme activity. Limnology and Oceanography, 36(7):1427–1436, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Ammerman:1991:BNAb

[AA91b] James W. Ammerman and Farooq Azam. Bacterial 5’- nucleotidase activity in estuarine and coastal marine waters: Role in phosphorus regeneration. Limnology and Oceanography, 36(7): 1437–1446, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Aller:1992:MST

[AA92] Robert C. Aller and Josephine Y. Aller. Meiofauna and solute transport in marine muds. Limnology and Oceanography, 37(5): 1018–1033, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Atkinson:1992:EWV

[AB92] M. J. Atkinson and R. W. Bilger. Effects of water velocity on phosphate uptake in coral reef-hat communities. Limnology and Oceanography, 37(2):273–279, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1994:GRC

[AB94] David H. Anderson and Arthur C. Benke. Growth and reproduction of the cladoceran Ceriodaphnia dubia from a forested floodplain swamp. Limnology and Oceanography, 39(7):1517–1527, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Amon:1996:BUD

[AB96] Rainer M. W. Amon and Ronald Benner. Bacterial utilization of different size classes of dissolved organic matter. Limnology and Oceanography, 41(1):41–51, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Acosta:1999:ASR

[AB99] Charles A. Acosta and Mark J. Butler I.V. Adaptive strategies that reduce predation on Caribbean spiny lobster postlarvae during onshore transport. Limnology and Oceanography, 44(3):494– 501, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Abbott:1990:BRS

[Abb90] Mark R. Abbott. Book review: Satellite monitoring of the earth (K.-H. Szekielda). Limnology and Oceanography, 35(2):546–547, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Allali:1995:NMM

[ABB^{+}95] Karima Allali, Annick Bricaud, Marcel Babin, André Morel, and Patrick Chang. A new method for measuring spectral absorption coefficients of marine particles. Limnology and Oceanography, 40 (8):1526–1532, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Alpine:1992:TID

[AC92] Andrea E. Alpine and James E. Cloern. Trophic interactions and direct physical effects control phytoplankton biomass and production in an estuary. Limnology and Oceanography, 37(5):946–955, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Archer:1992:BOF

[AD92] David Archer and Allan Devol. Benthic oxygen fluxes on the Washington shelf and slope: a comparison of in situ microelectrode and chamber flux measurements. Limnology and Oceanography, 37(3):614–629, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Agusti:1990:PAF

[ADC90] Susana Agusti, Carlos M. Duarte, and Daniel E. Canjield, Jr. Phytoplankton abundance in Florida lakes: Evidence for the frequent lack of nutrient limitation. Limnology and Oceanography, 35 (1):181–187, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Agusti:1992:SRB

[ADC92] Susana Agustí, Carlos M. Duarte, and Daniel E. Canfield, Jr. Selfregulation, bottom-up, and top-down control of phytoplankton communities: a reply to the comment by Kamenir. Limnology and Oceanography, 37(3):683–687, May 1992. CODEN LIOCAH. ISSN 0024-3590. See [Kam92].

Acuna:1996:PCM

[ADM96] José Luis Acuña, Don Deibel, and Claude C. Morris. Particle capture mechanism of the pelagic tunicate Oikopleura vanhoefleni. Limnology and Oceanography, 41(8):1800–1814, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Allen:1993:IPL

[ADR93] Yvonne C. Allen, Burt T. De Stasio, and Charles W. Ramcharan. Individual and population level consequences of an algal epibiont on Daphnia. Limnology and Oceanography, 38(3):592–601, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Acuna:1994:SDT

[ADS94] José Luis Acuña, Don Deibel, and Steve Sooley. A simple device to transfer large and delicate planktonic organisms. Limnology and Oceanography, 39(8):2001–2003, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Arts:1991:ODM

[AE91] Michael T. Arts and Marlene S. Evans. Optical-digital measurements of energy reserves in Calanoid copepods: Intersegmental distribution and seasonal patterns. Limnology and Oceanography, 36(2):289–298, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Abrahamsson:1995:MAS

[AECP95] K. Abrahamsson, A. Ekdahl, J. Collén, and M. Pedersén. Marine algae — a source of trichloroethylene and perchloroethylene. Limnology and Oceanography, 40(7):1321–1326, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Abril:1999:OAO

[AEL^{+}99] Gwenaël Abril, Henri Etcheber, Pierre Le Hir, Philippe Bassoullet, Bernard Boutier, and Michel Frankignoulle. Oxic/anoxic oscillations and organic carbon mineralization in an estuarine maximum turbidity zone (The Gironde, France). Limnology and Oceanography, 44(5):1304–1315, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Alldredge:1990:PSM

[AGGD90] Alice L. Alldredge, Timothy C. Granata, Chris C. Gotschalk, and Tommy D. Dickey. The physical strength of marine snow and its implications for particle disaggregation in the ocean. Limnology and Oceanography, 35(7):1415–1428, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Albrecht:1999:ILP

[AGZ99] Achim Albrecht, Gerrit Goudsmit, and Markus Zeh. Importance of lacustrine physical factors for the distribution of anthropogenic 60Co in Lake Biel. Limnology and Oceanography, 44(1):196–206, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Andersen:1991:CNP

[AH91] Tom Andersen and Dag O. Hessen. Carbon, nitrogen, and phosphorus content of freshwater zooplankton. Limnology and Oceanography, 36(4):807–814, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Aeschbach-Hertig:1996:DDE

[AHKH^{+}96] W. Aeschbach-Hertig, R. Kipfer, M. Hofer, D. M. Imboden, and H. Baur. Density-driven exchange between the basins of Lake Lucerne (Switzerland) traced with the^{3}H–^{3}He method. Limnology and Oceanography, 41(4):707–721, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Andersen:1992:IBM

[AK92] Frede Ostergaard Andersen and Erik Kristensen. The importance of benthic macrofauna in decomposition of microalgae in a coastal marine sediment. Limnology and Oceanography, 37(7):1392–1403, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Aas:1997:SSE

[AK97] Eyvind Aas and Brith Korsbø. Self-shading effect by radiance meters on upward radiance observed in coastal waters. Limnology and Oceanography, 42(5):968–974, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Ahner:1995:PPMa

[AKM95] Beth A. Ahner, Shing Kong, and Fraçois M. M. Morel. Phytochelatin production in marine algae. 1. An interspecies comparison. Limnology and Oceanography, 40(4):649–657, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Arndt:1990:CCM

[ALB90] Schimmelmann Arndt, Carina B. Lange, and Wolf H. Berger. Climatically controlled marker layers in Santa Barbara Basin sediments and fine-scale core-to-core correlation. Limnology and Oceanography, 35(1):165–173, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Altabet:1990:OCS

[Alt90] Mark A. Altabet. Organic C, N, and stable isotopic composition of particulate matter collected on glass-fiber and aluminum oxide filters. Limnology and Oceanography, 35(4):902–909, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Altabet:1994:BRT

[Alt94] Mark A. Altabet. Book review: Theory and application of tracers (D. S. Schimel). Limnology and Oceanography, 39(8):2017–2018, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Ahner:1995:PPMb

[AM95] Beth A. Ahner and Françios M. M. Morel. Phytochelatin production in marine algae. 2. Induction by various metals. Limnology and Oceanography, 40(4):658–665, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Anastasia:1998:TCL

[AMF98] Jean R. Anastasia, Steven G. Morgan, and Nicholas S. Fisher. Tagging crustacean larvae: Assimilation and retention of trace elements. Limnology and Oceanography, 43(2):362–368, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Aminot:1990:PDO

[Ami90] Alain Aminot. Precision of dissolved oxygen measurements. A fair comparison of methods. Limnology and Oceanography, 35(1): 237–238, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Abelson:1993:FPI

[AML93] Avigdor Abelson, Touvia Miloh, and Yossi Loya. Flow patterns induced by substrata and body morphologies of benthic organisms, and their roles in determining availability of food particles. Limnology and Oceanography, 38(6):1116–1124, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Arst:1997:CIP

[AMLH97] Helgi Arst, Sirje Mäekivi, Toomas Lukk, and Antti Herlevi. Calculating irradiance penetration into water bodies from the measured beam attenuation coefficient. Limnology and Oceanography, 42(2):379–385, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Ammerman:1996:BRA

[Amm96] James W. Ammerman. Book review: Antarctic microbiology (E. I.Friedmann [ed.]). Limnology and Oceanography, 41(1):193–194, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Abner:1997:TMC

[AMM97] B. A. Abner, F. M. M. Morel, and J. W. Moffett. Trace metal control of phytochelatin production in coastal waters. Limnology and Oceanography, 42(3):601–608, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Augustine:1998:SPM

[AMP98] Leon Augustine and Giséle Muller-Parker. Selective predation by the mosshead sculpin Clinocottus globiceps on the sea anemone Anthopleura elegantissima and its two algal symbionts. Limnology and Oceanography, 43(4):711–715, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1990:VDC

[And90] N. J. Anderson. Variability of diatom concentrations and accumulation rates in sediments of a small lake basin. Limnology and Oceanography, 35(2):497–508, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1993:SAM

[And93] Thomas R. Anderson. A spectrally averaged model of light penetration and photosynthesis. Limnology and Oceanography, 38(7): 1403–1419, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1997:BDT

[And97] Donald M. Anderson. Bloom dynamics of toxic Alexandrium species in the northeastern U.S. Limnology and Oceanography, 42(5 (part 2)):1009–1022, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Andersen:1999:MPG

[And99] Robert A. Andersen. Marine phytoplankton: a guide to naked flagellates and coccolithophorids. Limnology and Oceanography, 44(1):232–233, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1990:ASL

[Ano90a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 35(1):241, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1990:Aa

[Ano90b] Anonymous. Announcement. Limnology and Oceanography, 35 (1):238, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1990:Ab

[Ano90c] Anonymous. Announcement. Limnology and Oceanography, 35 (2):545, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1990:Ac

[Ano90d] Anonymous. Announcement. Limnology and Oceanography, 35 (2):548–458, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1990:Ad

[Ano90e] Anonymous. Announcement. Limnology and Oceanography, 35 (7):1644, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1990:ALO

[Ano90f] Anonymous. Authors in Limnology and Oceanography, volume 35, 1990. Limnology and Oceanography, 35(8):1867–1871, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1990:GEH

[Ano90g] Anonymous. G. Evelyn Hutchinson Medal Award. Limnology and Oceanography, 35(4):985–988, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1990:NCO

[Ano90h] Anonymous. National Council of Oceanography Graduate Students. Limnology and Oceanography, 35(1):244, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1990:PAS

[Ano90i] Anonymous. Proceedings of the American society of limnology and oceanography 53rd annual meeting. Limnology and Oceanography, 35(1):243, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1991:Aa

[Ano91a] Anonymous. Announcement. Limnology and Oceanography, 36 (4):750, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1991:Ab

[Ano91b] Anonymous. Announcement. Limnology and Oceanography, 36 (5):845, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1991:Ac

[Ano91c] Anonymous. Announcement. Limnology and Oceanography, 36 (6):1275, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1991:ALO

[Ano91d] Anonymous. Authors in Limnology and Oceanography, volume 36, 1991. Limnology and Oceanography, 36(8):1966–1970, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1991:GEH

[Ano91e] Anonymous. G. Evelyn Hutchinson Medal Award. Limnology and Oceanography, 36(7):1503–1505, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1991:PAS

[Ano91f] Anonymous. Proceedings of the American society of limnology and oceanography. Limnology and Oceanography, 36(4):828, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1992:GEH

[Ano92a] Anonymous. 1992 G. Evelyn Hutchinson Medal Award. Limnology and Oceanography, 37(8):1834–1836, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1992:LAa

[Ano92b] Anonymous. 1992 Lindeman Award. Limnology and Oceanography, 37(8):1836, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1992:LAb

[Ano92c] Anonymous. 1992 Lindeman Award. Limnology and Oceanography, 37(8):1837, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1992:EPD

[Ano92d] Anonymous. Estimate of phytoplankton division rates by the mitotic index method: The fmax approach revisited. Limnology and Oceanography, 37(3):644–649, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1992:SAA

[Ano92e] Anonymous. Student awards for the 1992 annual meeting. Limnology and Oceanography, 37(5):1126, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:AOS

[Ano93a] Anonymous. 1993 ASLO/ONR Student Poster Awards. Limnology and Oceanography, 38(7):1592, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:GEH

[Ano93b] Anonymous. 1993 G. Evelyn Hutchinson Medal Award. Limnology and Oceanography, 38(7):1590–1592, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:ARR

[Ano93c] Anonymous. Applicability of resource-ratio theory to microbial ecology. Limnology and Oceanography, 38(1):239–249, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:BPK

[Ano93d] Anonymous. Bibliography of Peter Kilham. Limnology and Oceanography, 38(4):707–708, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:BRC

[Ano93e] Anonymous. Book review: Copepod evolution (R. Huys and G. A. Boxshall). Limnology and Oceanography, 38(2):478, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:BRN

[Ano93f] Anonymous. Book review: Nitrogen isotope techniques (R. Knowles and T. H. Blackburn [eds.]). Limnology and Oceanography, 38(2):478–479, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:BRP

[Ano93g] Anonymous. Book review: Plant-animal interactions in the marine benthos (D. M. John, S. J. Hawkins, and J. H. Price [eds.]). Limnology and Oceanography, 38(2):479–480, March 1993. CO-DEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:BRB

[Ano93h] Anonymous. Book review: The biology of the chaetognaths (Q. Bone, H. Kapp, and A. C. Pierrot-Bults [eds.]). Limnology and Oceanography, 38(2):480–481, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:CTL

[Ano93i] Anonymous. Characteristics of two landslide-dammed lakes in a glaciated alpine environment. Limnology and Oceanography, 38 (2):441–445, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:DES

[Ano93j] Anonymous. Diffusional encounter of submicrometer particles and small cells by suspension feeders. Limnology and Oceanography, 38(2):456–465, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:DDB

[Ano93k] Anonymous. Distribution and dynamics of bacterial production in relation to wind perturbations in a Mediterranean lagoon. Limnology and Oceanography, 38(1):193–201, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:ECE

[Ano93l] Anonymous. An ecological cascade effect: Migratory birds affect stability of intertidal sediments. Limnology and Oceanography, 38 (1):225–231, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:ENS

[Ano93m] Anonymous. Effect of nutritional status on citrate synthase activity in Acartia tonsa and Temora longicornis. Limnology and Oceanography, 38(2):414–418, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:EDS

[Ano93n] Anonymous. Embryos of the deep-sea echinoid Echinus affinis require high pressure for development. Limnology and Oceanography, 38(1):178–181, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:EMM

[Ano93o] Anonymous. An epifluorescence microscopy method for direct detection and enumeration of the fungilike marine protists, the thraustochytrids. Limnology and Oceanography, 38(1):182–187, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:GAV

[Ano93p] Anonymous. Gastropod abundance in vegetated habitats: The importance of specifying null models. Limnology and Oceanography, 38(1):217–225, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:HMD

[Ano93q] Anonymous. Hydroacoustic measurement of the density of the Baikal macrozooplankter Macrohectopus branickii. Limnology and Oceanography, 38(2):425–434, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:IGE

[Ano93r] Anonymous. The influence of grazing on the estimation of phytoplankton growth rate via cell cycle analysis: Modeling and experimental evidence. Limnology and Oceanography, 38(1):202–212, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:ILI

[Ano93s] Anonymous. The influence of light intensity on di-methylsulfide production by a marine diatom. Limnology and Oceanography, 38 (2):419–425, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:MBS

[Ano93t] Anonymous. Methanogenic bacteria in seawater. Limnology and Oceanography, 38(2):452–456, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:MTM

[Ano93u] Anonymous. A method for temporal measurement of hemolymph composition in the giant clam symbiosis and its application to glucose and glycerol levels during a diel cycle. Limnology and Oceanography, 38(1):213–217, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:NSU

[Ano93v] Anonymous. Nitrogen source for uptake by gyrodinium cf. aureolum in a tidal front. Limnology and Oceanography, 38(2):446– 451, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:NOP

[Ano93w] Anonymous. Nitrous oxide production by estuarine epiphyton. Limnology and Oceanography, 38(2):435–441, March 1993. CO-DEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:PBS

[Ano93x] Anonymous. Proportionality between in situ carbon assimilation and bio-optical measures of primary production in the Gulf of Maine in summer. Limnology and Oceanography, 38(1):232–238, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:TDA

[Ano93y] Anonymous. Temperature-dependent absorption of water in the red and near-infrared portions of the spectrum. Limnology and Oceanography, 38(1):188–192, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1993:UBP

[Ano93z] Anonymous. Uranium behavior in a permanently anoxic fjord: Microbial control? Limnology and Oceanography, 38(2):408–414, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1994:PAS

[Ano94] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc., 57th Annual Meeting, University of Alberta, Edmonton, 3 June 1994. Limnology and Oceanography, 39 (2):477–478, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1995:ALA

[Ano95a] Anonymous. 1994 ASLO lifetime achievement award. Limnology and Oceanography, 40(1):208–210, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1995:GEH

[Ano95b] Anonymous. 1994 G. Evelyn Hutchinson Medal. Limnology and Oceanography, 40(1):205–207, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1995:AOS

[Ano95c] Anonymous. ASLO/ONR Student Poster Awards. Limnology and Oceanography, 40(1):210, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1995:LA

[Ano95d] Anonymous. Lindeman Award. Limnology and Oceanography, 40 (1):210, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1995:NMG

[Ano95e] Anonymous. New manuscript guidelines for the reporting of stable hydrogen, carbon, and oxygen isotope-ratio data. Limnology and Oceanography, 40(6):1182, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1996:A

[Ano96a] Anonymous. Addenda. Limnology and Oceanography, 41(2):242,

March 1996. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1996:C

[Ano96b] Anonymous. Correction. Limnology and Oceanography, 41(8): 1814, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1996:Ea

[Ano96c] Anonymous. Errata. Limnology and Oceanography, 41(8):809,

June 1996. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1996:Eb

[Ano96d] Anonymous. Errata. Limnology and Oceanography, 41(7):1520,

November 1996. CODEN LIOCAH. ISSN 0024-3590.

Anthony:1999:TSS

[Ant99] Kenneth R. N. Anthony. A tank system for studying benthic aquatic organisms at predictable levels of turbidity and sedimentation: Case study examining coral growth. Limnology and Oceanography, 44(6):1415–1422, September 1999. CODEN LI-OCAH. ISSN 0024-3590.

Aksnes:1996:VLT

[AO96] Dag L. Aksnes and Mark D. Ohman. A vertical life table approach to zooplankton mortality estimation. Limnology and Oceanography, 41(7):1461–1469, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Abele-Oeschger:1997:DUD

[AOTR97] Doris Abele-Oeschger, Helmut Tüg, and Rüdiger Röttgers. Dynamics of UV-driven hydrogen peroxide formation on an intertidal sandflat. Limnology and Oceanography, 42(6):1406–1415, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Agusti:1992:LAC

[AP92] Susana Agusti and Edward J. Phlips. Light absorption by cyanobacteria: Implications of the colonial growth form. Limnology and Oceanography, 37(2):434–441, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Appelmans:1994:SPS

[App94] Nicholas Appelmans. Sites of particle selection determined from observations of individual feeding larvae of the sand dollar Dendraster excentricus. Limnology and Oceanography, 39(2):404–411, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Arnosti:1994:ODA

[AR94] C. Arnosti and D. J. Repeta. Oligosaccharide degradation by anaerobic marine bacteria: Characterization of an experimental system to study polymer degradation in sediments. Limnology and Oceanography, 39(8):1865–1877, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Assel:1995:CWA

[AR95] Raymond A. Assel and Dale M. Robertson. Changes in winter air temperatures near Lake Michigan, 1851–1993, as determined from regional lake-ice records. Limnology and Oceanography, 40 (1):165–176, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Axler:1996:NUP

[AR96] Richard P. Axler and John E. Reuter. Nitrate uptake by phytoplankton and periphyton: Whole-lake enrichments and mesocosm-15N experiments in an oligotrophic lake. Limnology and Oceanography, 41(4):659–671, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Archer:1993:BRC

[Arc93] David Archer. Book review: Chemical oceanography (F. J. Millero and M. L. Sohn). Limnology and Oceanography, 38(8):1843–1844, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Armstrong:1994:GLN

[Arm94] Robert A. Armstrong. Grazing limitation and nutrient limitation in marine ecosystems: Steady state solutions of an ecosystem model with multiple food chains. Limnology and Oceanography, 39(3):597–608, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1996:ERW

[ARM96] Wendy L. Anderson, Dale M. Robertson, and John J. Magnuson. Evidence of recent warming and El Niño-related variations in ice breakup of Wisconsin lakes. Limnology and Oceanography, 41(5): 815–821, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Armstrong:1999:OBM

[Arm99] Robert A. Armstrong. An optimization-based model of iron–light– ammonium colimitation of nitrate uptake and phytoplankton growth. Limnology and Oceanography, 44(6):1436–1446, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Arnosti:1998:RPR

[Arn98] C. Arnosti. Rapid potential rates of extracellular enzymatic hydrolysis in Arctic sediments. Limnology and Oceanography, 43(2): 315–324, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Anati:1991:PTS

[AS91] D. A. Anati and M. Stiller. The post- 1979 thermohaline structure of the Dead Sea and the role of double-diffusive mixing. Limnology and Oceanography, 36(2):342–353, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Arrigo:1994:HRB

[AS94] Kevin R. Arrigo and Cornelius W. Sullivan. A high resolution bio-optical model of microalgal growth: Tests using sea-ice algal community time-series data. Limnology and Oceanography, 39(3): 609–631, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Alcaraz:1994:SST

[ASC94] Miguel Alcaraz, Enric Saiz, and Albert Calbet. Small-scale turbulence and zooplankton metabolism: Effects of turbulence on heartbeat rates of planktonic crustaceans. Limnology and Oceanography, 39(6):1465–1470, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Alstad:1999:ECC

[ASH99] Nina E. W. Alstad, Liz Skardal, and Dag O. Hessen. The effect of calcium concentration on the calcification of Daphnia magna. Limnology and Oceanography, 44(8):2011–2017, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Agusti:1998:DEA

[ASMB98] Susana Agustí, Maria Paola Satta, Maria Paola Mura, and Esther Benavent. Dissolved esterase activity as a tracer of phytoplankton lysis: Evidence of high phytoplankton lysis rates in the northwestern Mediterranean. Limnology and Oceanography, 43(8): 1836–1849, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Allen:1998:EEI

[ASS^{+}98] J. R. Allen, D. J. Slinn, T. M. Shummon, R. G. Hurtnoll, and S. J. Hawkins. Evidence for eutrophication of the Irish Sea over four decades. Limnology and Oceanography, 43(8):1970–1974, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1999:ELA

[ASS^{+}99] Dean E. Anderson, Robert G. Striegl, David I. Stannard, Catherine M. Michmerhuizen, Ted A. McConnaughey, and James W. LaBaugh. Estimating lake-atmosphere CO_{2} exchange. Limnology and Oceanography, 44(4):988–1001, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Alber:1994:BCO

[AV94a] Merryl Alber and Ivan Valiela. Biochemical composition of organic aggregates produced from marine macrophyte-derived dissolved organic matter. Limnology and Oceanography, 39(3):717– 723, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Alber:1994:PMO

[AV94b] Merryl Alber and Ivan Valiela. Production of microbial organic aggregates from macrophyte-derived dissolved organic material. Limnology and Oceanography, 39(1):37–50, January 1994. CO-DEN LIOCAH. ISSN 0024-3590.

Aksnes:1993:MSZ

[AW93] Dug L. Aksnes and Paul Wassmann. Modeling the significance of zooplankton grazing for export production. Limnology and Oceanography, 38(5):978–985, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Amano:1998:CNC

[AWKH98] Kunihiko Amano, Masataka Watanabe, Kunio Kohata, and Shigeki Harada. Conditions necessary for Chattonella antiqua red tide outbreaks. Limnology and Oceanography, 43(1):117–128, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Abelson:1994:HIS

[AWL94] Avigdor Abelson, Daniel Weihs, and Yossi Loya. Hydrodynamic impediments to settlement of marine propagules. and adhesivefilament solutions. Limnology and Oceanography, 39(1):164–169, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Anschutz:1998:BEP

[AZS^{+}98] Pierre Anschutz, Shaojun Zhong, Bjørn Sundby, Alfonso Mucci, and Charles Gobeil. Burial efficiency of phosphorus and the geochemistry of iron in continental margin sediments. Limnology and Oceanography, 43(1):53–64, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Bilger:1992:AMT

[BA92] R. W. Bilger and M. J. Atkinson. Anomalous mass transfer of phosphate on coral reef flats. Limnology and Oceanography, 37 (2):261–272, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Bilger:1995:ENL

[BA95] R. W. Bilger and M. J. Atkinson. Effects of nutrient loading on mass-transfer rates to a coral-reef community. Limnology and Oceanography, 40(2):279–289, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Baird:1997:MPM

[BA97] M. E. Baird and M. J. Atkinson. Measurement and prediction of mass transfer to experimental coral reef communities. Limnology and Oceanography, 42(8):1685–1693, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Bachmann:1999:GRN

[Bac99] Roger W. Bachmann. A guide to the restoration of nutrientenriched shallow lakes. Limnology and Oceanography, 44(3):737, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Blackburn:1997:SES

[BAH97] Nicholas Blackburn, Farooq Azam, and Åke Hagström. Spatially explicit simulations of a microbial food web. Limnology and Oceanography, 42(4):613–622, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Baldwin:1996:EEA

[Bal96] D. S. Baldwin. Effects of exposure to air and subsequent drying on the phosphate sorption characteristics of sediments from a eutrophic reservoir. Limnology and Oceanography, 41(8):1725–1732, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Bannister:1990:CKM

[Ban90a] T. T. Bannister. Comparison of Kiefer-Mitchell and bannisterlaws algal models. Limnology and Oceanography, 35(4):972–979, June 1990. CODEN LIOCAH. ISSN 0024-3590. See reply [Kie90].

Bannister:1990:EER

[Ban90b] T. T. Bannister. Empirical equations relating scalar irradiance to a, b/a, and solar zenith angle. Limnology and Oceanography, 35 (1):173–176, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Banse:1990:BRB

[Ban90c] Karl Banse. Book review: Biological oceanography. An early history. 187-196 (E. L. Mills). Limnology and Oceanography, 35(6): 1410–1411, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Banse:1990:DIR

[Ban90d] Karl Banse. Does iron really limit phytoplankton production in the offshore subarctic Pacific? Limnology and Oceanography, 35 (3):772–775, May 1990. CODEN LIOCAH. ISSN 0024-3590. See reply [Ban90d].

Banse:1990:MTB

[Ban90e] Karl Banse. Mermaids- their biology, culture. and demise. Limnology and Oceanography, 35(1):148–153, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Banse:1991:RPC

[Ban91] Karl Banse. Rates of phytoplankton cell division in the field and in iron enrichment experiments. Limnology and Oceanography, 36 (8):1886–1898, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Bannister:1992:MMC

[Ban92] T. T. Bannister. Model of the mean cosine of underwater radiance and estimation of underwater scalar irradiance. Limnology and Oceanography, 37(4):773–780, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Banse:1998:EGS

[Ban98] Karl Banse. Ecological geography of the sea. Limnology and Oceanography, 43(7):1763, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Brzezinski:1997:SCW

[BAO97] Mark A. Brzezinski, Alice L. Alldredge, and Linda M. O’Bryan. Silica cycling within marine snow. Limnology and Oceanography, 42(8):1706–1713, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Barber:1999:PMR

[Bar99] David G. Barber. Passive microwave remote sensing of oceans. Limnology and Oceanography, 44(3):738, May 1999. CODEN LI-OCAH. ISSN 0024-3590.

Biddanda:1997:CNC

[BB97] Bopaiah Biddanda and Ronald Benner. Carbon, nitrogen, and carbohydrate fluxes during the production of particulate and dissolved organic matter by marine phytoplankton. Limnology and Oceanography, 42(3):506–518, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Benner:1998:PTS

[BB98] Ronald Benner and Bopaiah Biddanda. Photochemical transformations of surface and deep marine dissolved organic matter: Effects on bacterial growth. Limnology and Oceanography, 43(6): 1373–1378, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Bolliger:1992:SWA

[BBH^{+}92] Reinhard Bolliger, Helmut Brandl, Patrick Höhener, Kurt W. Hanselmann, and Reinhard Bachofen. Squeeze-water analysis for the determination of microbial metabolites in lake sedimentscomparison of methods. Limnology and Oceanography, 37(2):448– 455, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Bernasconi:1997:CNI

[BBS97] Stefano M. Bernasconi, Alberto Barbieri, and Marco Simona. Carbon and nitrogen isotope variations in sedimenting organic matter in Lake Lugano. Limnology and Oceanography, 42(8):1755–1765, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Buonaccorsi:1990:AIF

[BC90] Giovanni Buonaccorsi and Alexander Cunningham. Azimuthal inhomogeneity in the forward light scattered from microalgal colonies, and its use as a morphological indicator in flow cytometry. Limnology and Oceanography, 35(5):1170–1175, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Barber:1991:RPP

[BC91] R. T. Barber and F. P. Chavez. Regulation of primary productivity rate in the equatorial Pacific. Limnology and Oceanography, 36(8):1803–1815, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Beukema:1997:LDM

[BC97] J. J. Beukema and G. C. Cadée. Local differences in macrozoobenthic response to enhanced food supply caused by mild eutrophication in a Wadden Sea area. Food is only locally a limiting factor. Limnology and Oceanography, 42(6):1424–1435, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Beal:1993:BRC

[BCBC93] Brian F. Beal, Merrie Cartwright, Desiree Bethune, and Sara A. Chaves. Book review: A critique for ecology (R. H. Peters). Limnology and Oceanography, 38(6):1344–1346, September 1993. CO-DEN LIOCAH. ISSN 0024-3590.

Boudreau:1992:EDM

[BCM92] Bernard P. Boudreau, Donald E. Canfeld, and Alfonso Mucci. Early diagenesis in a marine sapropel, Mangrove Lake, Bermuda. Limnology and Oceanography, 37(8):1738–1753, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Boschker:1999:CMD

[BdBC99] H. T. S. Boschker, J. F. C. de Brouwer, and T. E. Cappenberg. The contribution of macrophyte-derived organic matter to microbial biomass in salt-marsh sediments: Stable carbon isotope analysis of microbial biomarkers. Limnology and Oceanography, 44(2): 309–319, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Buma:1991:MEE

[BdBNvB91] Anita G. J. Buma, Hein J. W. de Baar, Rob F. Nolting, and Aleido J. van Bennekom. Metal enrichment experiments in the Weddell-Scotia Seas: Effects of iron and manganese on various plankton communities. Limnology and Oceanography, 36(8):1865– 1878, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Bruland:1991:IIB

[BDH91] Kenneth W. Bruland, John R. Donat, and David A. Hutchins. Interactive influences of bioactive trace metals on biological production in oceanic waters. Limnology and Oceanography, 36(8): 1555–1577, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Bochdansky:1999:AEB

[BDR99] Alexander B. Bochdansky, Don Deibel, and Richard B. Rivkin. Absorption efficiencies and biochemical fractionation of assimilated compounds in the cold water appendicularian Oikopleura vanhoeffeni. Limnology and Oceanography, 44(2):415–424, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Baker:1999:ARM

[BDV99] Michelle A. Baker, Clifford N. Dahm, and H. Maurice Valett. Acetate retention and metabolism in the hyporheic zone of a mountain stream. Limnology and Oceanography, 44(6):1530–1539, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Brandes:1998:ICN

[BDY^{+}98] Jay A. Brandes, Allan H. Devol, T. Yoshinari, D. A. Jayakumar, and S. W. A. Naqvi. Isotopic composition of nitrate in the central Arabian Sea and eastern tropical North Pacific: a tracer for mixing and nitrogen cycles. Limnology and Oceanography, 43(7): 1680–1689, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Brooks:1994:BCP

[BE94] Arthur S. Brooks and David N. Edgington. Biogeochemical control of phosphorus cycling and primary production in Lake Michigan. Limnology and Oceanography, 39(4):961–968, June 1994. CO-DEN LIOCAH. ISSN 0024-3590.

Baker:1992:RSE

[BEB92] Lawrence A. Baker, Daniel R. Engstrom, and Patrick L. Brezonik. Recent sulfur enrichment in the sediments of Little Rock Lake, Wisconsin. Limnology and Oceanography, 37(4):689–702, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Bell:1990:EVC

[Bel90] Russell T. Bell. An explanation for the variability in the conversion factor deriving bacterial cell production from incorporation of [^{3}H]thymidine. Limnology and Oceanography, 35(4):910–915, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Benson:1994:SIOa

[Ben94] L. V. Benson. Stable isotopes of oxygen and hydrogen in the Truckee River–Pyramid Lake surface-water column. 1. Data analysis and extraction of paleoclimatic information. Limnology and Oceanography, 39(2):344–355, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Berelson:1991:FTD

[Ber91] William M. Berelson. The flushing of two deep-sea basins, southern California borderland. Limnology and Oceanography, 36(6): 1150–1166, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Berges:1997:RRS

[Ber97] John A. Berges. Ratios, regression statistics, and “spurious” correlations. Limnology and Oceanography, 42(5):1006–1007, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Berendonk:1999:IFK

[Ber99a] Thomas U. Berendonk. Influence of fish kairomones on the ovipositing behavior of Chaoborus imagines. Limnology and Oceanography, 44(2):454–458, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Bergey:1999:CRS

[Ber99b] Elizabeth A. Bergey. Crevices as refugia for stream diatoms: Effect of crevice size on abraded substrates. Limnology and Oceanography, 44(6):1522–1529, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Behrenfeld:1997:CGP

[BF97a] Michael J. Behrenfeld and Paul G. Falkowski. A consumer’s guide to phytoplankton primary productivity models. Limnology and Oceanography, 42(7):1479–1491, November 1997. CODEN LI-OCAH. ISSN 0024-3590.

Behrenfeld:1997:PRD

[BF97b] Michael J. Behrenfeld and Paul G. Falkowski. Photosynthetic rates derived from satellite-based chlorophyll concentration. Limnology and Oceanography, 42(1):1–20, January 1997. CODEN LI-OCAH. ISSN 0024-3590.

Berges:1998:PSC

[BF98] John A. Berges and Paul G. Falkowski. Physiological stress and cell death in marine phytoplankton: Induction of proteases in response to nitrogen or light limitation. Limnology and Oceanography, 43(1):129–135, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Berman-Frank:1996:ICP

[BFE96] Ilana Berman-Frank and Jonathan Erez. Inorganic carbon pools in the bloom-forming dinoflagellate Peridinium gatunense. Limnology and Oceanography, 41(8):1780–1789, December 1996. CO-DEN LIOCAH. ISSN 0024-3590.

Barker:1994:EDD

[BFG94] Philip Barker, Jean-Charles Fontes, and Françoise Gasse. Experimental dissolution of diatom silica in concentrated salt solutions and implications for paleoenvironmental reconstruction. Limnology and Oceanography, 39(1):99–110, January 1994. CODEN LI-OCAH. ISSN 0024-3590.

Butman:1994:FEF

[BFGS94] Cheryl Ann Butman, Marcel Fréchette, W. Rockwell Geyer, and Victoria R. Starczak. Flume experiments on food supply to the blue mussel Mytilus edulis L. as a function of boundary-layerflow. Limnology and Oceanography, 39(7):1755–1768, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Berninger:1991:PCB

[BFKL91] Ulrike-G. Berninger, Bland J. Finlay, and Pirjo Kuuppo-Leinikki. Protozoan control of bacterial abundances in freshwater. Limnology and Oceanography, 36(1):139–147, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Bryant:1997:MBS

[BFM^{+}97] C. L. Bryant, J. G. Farmer, A. B. MacKenzie, A. E. Bailey-Watts, and A. Kirika. Manganese behavior in the sediments of diverse Scottish freshwater lochs. Limnology and Oceanography, 42(5): 918–929, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Benner:1991:DBB

[BFS91] Ronald Benner, Marilyn L. Fogel, and E. Kent Sprague. Diagenesis of belowground biomass of Spartina alterniflora in saltmarsh sediments. Limnology and Oceanography, 36(7):1358–1374, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Brown:1992:CPS

[BFW92] Bruce E. Brown, Judy L. Fassbender, and Ronald Winkler. Carbonate production and sediment transport in a marl lake of southeastern Wisconsin. Limnology and Oceanography, 37(1):184–191, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Berman-Frank:1994:CAC

[BFZED94] Ilana Berman-Frank, Tamar Zohary, Jonathan Erez, and Zvy Dubinsky. CO_{2} availability, carbonic anhydrase, and the annual dinoflagellate bloom in Lake Kinneret. Limnology and Oceanography, 39(8):1822–1834, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Berg:1994:GDN

[BG94a] David J. Berg and David W. Garton. Genetic differentiation in North American and European populations of the cladoceran Bythotrephes. Limnology and Oceanography, 39(7):1503–1516, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Bronk:1994:FMD

[BG94b] Deborah A. Bronk and Patricia M. Glibert. The fate of the missing 15N differs among marine systems. Limnology and Oceanography, 39(1):189–195, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Burkholder:1997:PPO

[BG97a] JoAnn M. Burkholder and Howard B. Glasgow, Jr. Pfiesteria piscicida and other Pfiesreria-like dinoflagellates: Behavior, impacts, and environmental controls. Limnology and Oceanography, 42(5 (part 2)):1052–1075, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Buscail:1997:PDO

[BG97b] Roselyne Buscail and Claire Germain. Present-day organic matter sedimentation on the NW Mediterranean margin: Importance of off-shelf export. Limnology and Oceanography, 42(2):217–229, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Bird:1994:EFS

[BGC94] M. I. Bird, P. Giresse, and A. R. Chivas. Effect of forest and savanna vegetation on the carbon-isotope composition of sediments from the Sanaga River, Cameroon. Limnology and Oceanography, 39(8):1845–1854, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Berg:1999:DEE

[BGC99] Gry Mine Berg, Patricia M. Glibert, and Chung-Chi Chen. Dimension effects of enclosures on ecological processes in pelagic systems. Limnology and Oceanography, 44(5):1331–1340, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Byeong-Gweon:1992:DER

[BGF92] Lee Byeong-Gweon and Nicholas S. Fisher. Degradation and elemental release rates from phytoplankton debris and their geochemical implications. Limnology and Oceanography, 37(7):1345– 1360, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Bird:1998:SCC

[BGN98] Michael I. Bird, Pierre Giresse, and Simon Ngos. A seasonal cycle in the carbon-isotope composition of organic carbon in the Sanaga River, Cameroon. Limnology and Oceanography, 43(1):143–146, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Boston:1991:PLR

[BH91] H. L. Boston and W. R. Hill. Photosynthesis–light relations of stream periphyton communities. Limnology and Oceanography, 36(4):644–656, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Berges:1995:NRA

[BH95] John A. Berges and P. J. Harrison. Nitrate reductase activity quantitatively predicts the rate of nitrate incorporation under steady state light limitation: a revised assay and characterization of the enzyme in three species of marine phytoplankton. Limnology and Oceanography, 40(1):82–93, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Benoit:1996:VED

[BH96] Gaboury Benoit and Harold F. Hemond. Vertical eddy diffusion calculated by the flux gradient method: Significance of sedimentwater heat exchange. Limnology and Oceanography, 41(1):157– 168, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Beaton:1997:CBD

[BH97] Margaret J. Beaton and Paul D. N. Hebert. The cellular basis of divergent head morphologies in Daphnia. Limnology and Oceanography, 42(2):346–356, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Balch:1991:BOP

[BHAV91] William M. Balch, Patrick M. Holligan, Steven G. Ackleson, and Kenneth J. Voss. Biological and optical properties of mesoscale coccolithophore blooms in the Gulf of Maine. Limnology and Oceanography, 36(4):629–643, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Bienfang:1999:TM

[Bie99] Paul K. Bienfang. Tropical mariculture. Limnology and Oceanography, 46(4):1183–1184, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Blaustein:1990:IEM

[BK90] Leon Blaustein and Richard Karban. Indirect effects of the mosquitofish Gambusia affinis on the mosquito Culex tarsalis. Limnology and Oceanography, 35(3):767–771, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Blais:1995:ILM

[BK95] Jules M. Blais and Jacob Kalff. The influence of lake morphometry on sediment focusing. Limnology and Oceanography, 40(3):582– 588, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Benoy:1999:SAP

[BK99] Glenn A. Benoy and Jacob Kalff. Sediment accumulation and Pb burdens in submerged macrophyte beds. Limnology and Oceanography, 46(4):1081–1090, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Barry:1997:IPW

[BKB97] J. P. Barry, R. E. Kochevar, and C. H. Baxter. The influence of pore-water chemistry and physiology on the distribution of vesicomyid clams at cold seeps in Monterey Bay: Implications for patterns of chemosynthetic community organization. Limnology and Oceanography, 42(2):318–328, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Balch:1996:CBCb

[BKH^{+}96] William M. Balch, Katherine A. Kilpatrick, Patrick Holligan, Derek Harbour, and Emilio Fernandez. The 1991 coccolithophore bloom in the central North Atlantic. 2. Relating optics to coccolith concentration. Limnology and Oceanography, 41(8):1684–1696, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Belviso:1990:PDP

[BKR^{+}90] S. Belviso, S.-K. Kim, F. Rassoulzadegan, B. Krajka, B. C. Nguyen, N. Mihalopoulos, and P. Buat-Menard. Production of dimethylsulfonium propionate (DMSP) and dimethylsulfide (DMS) by a microbial food web. Limnology and Oceanography, 35(8):1810–1821, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Balch:1996:CBCa

[BKT96] William M. Balch, Katherine A. Kilpatrick, and Charles C. Trees. The 1991 coccolithophore bloom in the central North Atlantic. 1. Optical properties and factors affecting their distribution. Limnology and Oceanography, 41(8):1669–1683, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Branstrator:1991:IPL

[BL91] Donn K. Branstrator and John T. Lehman. Invertebrate predation in Lake Michigan: Regulation of Bosmina longirostris by Leptodora kindtii. Limnology and Oceanography, 36(3):483–495, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Blomqvist:1994:DBE

[BL94a] Sven Blomqvist and Ulf Larsson. Detrital bedrock elements as tracers of settling resuspended particulate matter in a coastal area of the Baltic Sea. Limnology and Oceanography, 39(4):880–896, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Burkhardt:1994:PCP

[BL94b] Stefen Burkhardt and John T. Lehman. Prey consumption and predatory effects of an invertebrate predator (bythotrephes: Cladocera, cercopagidae) based on phosphorus budgets. Limnology and Oceanography, 39(5):1007–1019, July 1994. CODEN LI-OCAH. ISSN 0024-3590.

Bricelj:1997:AAC

[BL97] V. Monica Bricelj and Darcy J. Lonsdale. Aureococcus anophagefferens: Causes and ecological consequences of brown tides in U.S. mid-Atlantic coastal waters. Limnology and Oceanography, 42(5 (part 2)):1023–1038, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Bertilsson:1998:PPC

[BL98] Stefan Bertilsson and Stefan J. Lam. Photochemically produced carboxylic acids as substrates for freshwater bacterioplankton.

Limnology and Oceanography, 43(5):885–895, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Black:1993:PIP

[Bla93] A. Ross Black. Predator-induced phenotypic plasticity in Daphnia pulex: Life history and morphological responses to notonecta and chaoborus. Limnology and Oceanography, 38(5):986–996, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Brosse:1999:PFD

[BLD99] Sébastien Brosse, Sovan Lek, and Francis Dauba. Predicting fish distribution in a mesotrophic lake by hydroacoustic survey and artificial neural networks. Limnology and Oceanography, 44(5): 1293–1303, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Blair:1996:STF

[BLDP96] Neal E. Blair, Lisa A. Levin, David J. DeMaster, and Gayle Plaia. The short-term fate of fresh algal carbon in continental slope sediments. Limnology and Oceanography, 41(6):1208–1219, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Bunn:1995:EAW

[BLK95] S. E. Bunn, N. R. Loneragan, and M. A. Kempster. Effects of acid washing on stable isotope ratios of C and N in penaeid shrimp and seagrass: Implications for food-web studies using multiple stable isotopes. Limnology and Oceanography, 40(3):622–625, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Bruning:1992:EIF

[BLR92] Kees Bruning, Rob Lingeman, and Joop Ringelberg. Estimating the impact of fungal parasites on phytoplankton populations. Limnology and Oceanography, 37(2):252–260, March 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Bianchi:1995:PPB

[BLS^{+}95] Thomas S. Bianchi, Corey Lambert, Peter H. Santschi, M. Baskaran, and Laodong Guo. Plant pigments as biomarkers of highmolecular-weight dissolved organic carbon. Limnology and Oceanography, 40(2):422–428, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Belanger:1992:SME

[BM92a] T. V. Belanger and M. T. Montgomery. Seepage meter errors. Limnology and Oceanography, 37(8):1787–1795, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Berthon:1992:VSL

[BM92b] Jean-François Berthon and André Morel. Validation of a spectral light-photosynthesis model and use of the model in conjunction with remotely sensed pigment observations. Limnology and Oceanography, 37(4):781–796, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Bourbonniere:1996:SGR

[BM96] Richard A. Bourbonniere and Philip A. Meyers. Sedimentary geolipid records of historical changes in the watersheds and productivities of lakes Ontario and erie. Limnology and Oceanography, 41(2):352–359, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Bock:1999:PSG

[BM99] Michael J. Bock and Douglas C. Miller. Particle selectivity, gut volume, and the response to a step change in diet for depositfeeding polychaetes. Limnology and Oceanography, 46(4):1132– 1138, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Buskey:1997:DGP

[BMAW97] Edward J. Buskey, Paul A. Montagna, Anthony F. Amos, and Terry E. Whitledge. Disruption of grazer populations as a contributing factor to the initiation of the Texas brown tide algal bloom. Limnology and Oceanography, 42(5 (part 2)):1215–1222, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Band:1996:EPW

[BMC^{+}96] Lawrence E. Band, D. Scott Mackay, Irena F. Creed, Ray Semkin, and Dean Jeffries. Ecosystem processes at the watershed scale: Sensitivity to potential climate change. Limnology and Oceanography, 41(5):928–938, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Balistrieri:1992:BCT

[BMP92a] Laurie S. Balistrieri, James W. Murray, and Barbara Paul. The biogeochemical cycling of trace metals in the water column of Lake Sammamish, Washington: Response to seasonally anoxic conditions. Limnology and Oceanography, 37(3):529–548, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Balistrieri:1992:CIM

[BMP92b] Laurie S. Balistrieri, James W. Murray, and Barbara Paul. The cycling of iron and manganese in the water column of Lake Sammamish, Washington. Limnology and Oceanography, 37(3):510– 528, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Bengtsson:1996:FIW

[BMT^{+}96] Lars Bengtsson, Joakim Malm, Arkady Terzhevik, Michail Petrov, Pjotr Boyarinov, Alexander Glinsky, and Nikolqj Palshin. Field investigation of winter thermo- and hydrodynamics in a small Karelian lake. Limnology and Oceanography, 41(7):1502–1513, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Berounsky:1990:TAC

[BN90] Veronica M. Berounsky and Scott W. Nixon. Temperature and annual cycle of nitrification in waters of Narragansett Bay. Limnology and Oceanography, 35(7):1610–1617, November 1990. CO-DEN LIOCAH. ISSN 0024-3590.

Buck:1995:FPF

[BN95] Kurt R. Buck and Jan Newton. Fecal pellet flux in Dabob Bay during a diatom bloom: Contribution of microzooplankton. Limnology and Oceanography, 40(2):306–315, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Berendonk:1996:MRC

[BO96] Thomas U. Berendonk and W. John O’Brien. Movement response of chaoborus to chemicals from a predator and prey. Limnology and Oceanography, 41(4):1829–1832, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Biddanda:1994:PRC

[BOB94] Bopaiah Biddanda, Stephen Opsahl, and Ronald Benner. Plankton respiration and carbon flux through bacterioplankton on the Louisiana shelf. Limnology and Oceanography, 39(6):1259–1275, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Benner:1995:BCM

[BOCL^{+}95] Ronald Benner, Stephen Opsahl, Gerardo Chin-Leo, Jeffrey E. Richey, and Bruce R. Forsberg. Bacterial carbon metabolism in the Amazon River system. Limnology and Oceanography, 40(7): 1262–1270, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Bollens:1993:VDS

[BOFW93] Stephen M. Bollens, Kenric Osgood, Bruce W. Frost, and Sidney D. Watts. Vertical distributions and susceptibilities to vertebrate predation of the marine copepods Metridia lucens and Calanus pacificus. Limnology and Oceanography, 38(8):1827– 1837, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Booth:1994:BRM

[Boo94] Beatrice C. Booth. Book review: Marine phytoplankton. A guide to naked flagellates and coccolithophorids (C. R. Tomas [ed.]). Limnology and Oceanography, 39(4):982–983, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Boudreau:1998:MMD

[Bou98] Bernard P. Boudreau. Mean mixed depth of sediments: The wherefore and the why. Limnology and Oceanography, 43(3):524– 526, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Boyle:1994:DSM

[Boy94] D. R. Boyle. Design of a seepage meter for measuring groundwater fluxes in the nonlittoral zones of lakes-evaluation in a boreal forest lake. Limnology and Oceanography, 39(3):670–681, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Baines:1991:PDO

[BP91] Stephen B. Baines and Michael L. Pace. The production of dissolved organic matter by phytoplankton and its importance to bacteria: Patterns across marine and freshwater systems. Limnology and Oceanography, 36(6):1078–1090, September 1991. CO-DEN LIOCAH. ISSN 0024-3590.

Bum:1996:FRP

[BP96] Ben K. Bum and Frances R. Pick. Factors regulating phytoplankton and zooplankton biomass in temperate rivers. Limnology and Oceanography, 41(7):1572–1577, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Brzezinski:1997:ETM

[BP97] Mark A. Brzezinski and Dennis R. Phillips. Evaluation of ^{32}Si as a tracer for measuring silica production rates in marine waters. Limnology and Oceanography, 42(5):856–865, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Buskey:1996:SBC

[BPA96] Edward J. Buskey, Jay O. Peterson, and Julie W. Ambler. he swarming behavior of the copepod Dioithona oculata: In situ and laboratory studies. Limnology and Oceanography, 41(3):513–521, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Brzezinski:1997:SPM

[BPC^{+}97] Mark A. Brzezinski, Dennis R. Phillips, Francisco P. Chavez, Gernot E. Friederich, and Richard C. Dugdale. Silica production in the Monterey, California, upwelling system. Limnology and Oceanography, 42(8):1694–1705, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Bissett:1997:PPC

[BPCL97] W. Paul Bissett, Jennifer S. Patch, Kendall L. Carder, and Z. Ping Lee. Pigment packaging and chl-a-specific absorption in high-light oceanic waters. Limnology and Oceanography, 42(5):961–968, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Baines:1994:WDR

[BPK94] Stephen B. Baines, Michael L. Pace, and David M. Karl. Why does the relationship between sinking flux and planktonic primary production differ between lakes and oceans? Limnology and Oceanography, 39(2):213–226, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Brenner:1994:DBS

[BPS94] Mark Brenner, Arthur J. Peplow, and Claire L. Schelske. Disequilibrium between ^{226}Ra and supported ^{210}Pb in a sediment core from a shallow Florida lake. Limnology and Oceanography, 39(5): 1222–1227, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Boland:1991:DSB

[BR91] Gregory S. Boland and Gilbert T. Rowe. Deep-sea benthic sampling with the GOMEX box corer. Limnology and Oceanography, 36(5):1015–1020, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Brettar:1992:ICA

[BR92] Ingrid Brettar and Gerhard Rheinheimer. Influence of carbon availability on dentrification in the central Baltic Sea. Limnology and Oceanography, 37(6):1146–1163, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Brand:1991:MIR

[Bra91] Larry E. Brand. Minimum iron requirements of marine phytoplankton and the implications for the biogeochemical control of new production. Limnology and Oceanography, 36(8):1756–1771, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Branstrator:1998:PDC

[Bra98] Donn K. Branstrator. Predicting diet composition from body length in the zooplankton predator Leptodora kindti. Limnology and Oceanography, 43(3):530–535, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Brendelberger:1991:FMS

[Bre91] Heinz Brendelberger. Filter mesh size of cladocerans predicts retention efficiency for bacteria. Limnology and Oceanography, 36 (5):884–894, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Brett:1993:CPP

[Bre93] Michael T. Brett. Comment on “Possibility of N or P limitation for planktonic cladocerans: an experimental test” (Urabe and Watanabe) and “Nutrient element limitation of zooplankton production” (Hessen). Limnology and Oceanography, 38(6):1333– 1337, September 1993. CODEN LIOCAH. ISSN 0024-3590. See [UW92, UW93] and reply [Hes93].

Brendelberger:1997:BDE

[Bre97] Heinz Brendelberger. Bacteria and digestive enzymes in the alimentary tract of Radix peregra (Gastropoda, Lymnaeidae). Limnology and Oceanography, 42(7):1635–1638, November 1997. CO-DEN LIOCAH. ISSN 0024-3590.

Berg:1998:IMC

[BRPR98] Peter Berg, Nils Risgaard-Petersen, and Søren Rysgaard. Interpretation of measured concentration profiles in sediment pore water. Limnology and Oceanography, 43(7):1500–1510, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Bruland:1992:CCN

[Bru92] Kenneth W. Bruland. Complexation of cadmium by natural organic ligands in the central North Pacific. Limnology and Oceanography, 37(5):1008–1017, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Bricaud:1995:SMM

[BRZ95] Annick Bricaud, Collin Roesler, and J. Ronald V. Zaneveld. In situ methods for measuring the inherent optical properties of ocean waters. Limnology and Oceanography, 40(2):393–410, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Blanchard:1990:ADE

[BS90a] Duncan C. Blanchard and Lawrence D. Syzdek. Apparatus to determine the efficiency of transfer of bacteria from a bursting bubble to the jet drops. Limnology and Oceanography, 35(1):136– 143, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Bricaud:1990:SAC

[BS90b] Annick Bricaud and Dariusz Stramski. Spectral absorption coefficients of living phytoplankton and nonalgal biogenous matter: a comparison between the Peru upwelling area and the Sargasso Sea. Limnology and Oceanography, 35(3):562–582, May 1990. CO-DEN LIOCAH. ISSN 0024-3590.

Buskey:1990:EMP

[BS90c] Edward J. Buskey and Elijah Swift. An encounter model to predict natural planktonic bioluminescence. Limnology and Oceanography, 35(7):1469–1485, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Bunte:1999:BTD

[BS99] Christoph Bunte and Meinhard Simon. Bacterioplankton turnover of dissolved free monosaccharides in a mesotrophic lake. Limnology and Oceanography, 44(8):1862–1870, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Bowen:1993:SBC

[BSC93] James D. Bowen, Keith D. Stolzenbach, and Sallie W. Chisholm. Simulating bacterial clustering around phytoplankton cells in a turbulent ocean. Limnology and Oceanography, 38(1):36–51, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Breitburg:1999:VRN

[BSG^{+}99] Denise L. Breitburg, James G. Sanders, Cynthia C. Gilmour, Colleen A. Hatfield, Richard W. Osman, Gerhardt F. Riedel, Sybil P. Seitzinger, and Sybil P. Seitzinger. Variability in responses to nutrients and trace elements, and transmission of stressor effects through an estuarine food web. Limnology and Oceanography, 44(3 (part 2)):837–863, May 1999. CODEN LI-OCAH. ISSN 0024-3590.

Berwald:1995:IAS

[BSMK95] Juli Berwald, Dariusz Stramski, Curtis D. Mobley, and Dale A. Kiefer. Influences of absorption and scattering on vertical changes in the average cosine of the underwater light field. Limnology and Oceanography, 40(8):1347–1357, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Berwald:1998:ERS

[BSMK98] Juli Berwald, Dariusz Stramski, Curtis D. Mobley, and Dale A. Kiefer. Effect of Raman scattering on the average cosine and diffuse attenuation coefficient of irradiance in the ocean. Limnology and Oceanography, 43(4):564–576, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Brix:1992:IPC

[BSO92] Hans Brix, Brian K. Sorrell, and Philip T. Orr. Internal pressurization and convective gas flow in some emergent freshwater macrophytes. Limnology and Oceanography, 37(7):1420–1433, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Berman:1995:PPP

[BSY^{+}95] T. Berman, L. Stone, Y. Z. Yacobi, B. Kaplan, Al. Schlichter, A. Nishri, and U. Pollingher. Primary production and phytoplankton in Lake Kinneret: a long-term record (1972–1993). Limnology and Oceanography, 40(6):1064–1076, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Babin:1995:RBM

[BTL^{+}95] Marcel Babin, Jean-Claude Therriault, Louis Legendre, Barbara Nieke, Rainer Reuter, and Alfonso Condal. Relationship between the maximum quantum yield of carbon fixation and the minimum quantum yield of chlorophyll-a in vivo fluorescence in the Gulf of St. Lawrence. Limnology and Oceanography, 41(4):956–968, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Babin:1993:VSA

[BTLC93] Marcel Babin, Jean-Claude Therriault, Louis Legendre, and Alfonso Condal. Variations in the specific absorption coefficient for natural phytoplankton assemblages: Impact on estimates of primary production. Limnology and Oceanography, 38(1):154–177, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Bentzen:1992:IDO

[BTM92] E. Bentzen, W. D. Taylor, and E. S. Millard. The importance of dissolved organic phosphorus to phosphorus uptake by limnetic plankton. Limnology and Oceanography, 37(2):217–231, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Burkholder:1992:PES

[Bur92] JoAnn M. Burkholder. Phytoplankton and episodic suspended sediment loading: Phosphate partitioning and mechanisms for survival. Limnology and Oceanography, 37(5):974–988, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Burdige:1994:BRO

[Bur94] David J. Burdige. Book review: Organic geochemistry: Principles and applications (M. H. Engel and S. A. Macko [eds.]). Limnology and Oceanography, 39(6):1500–1501, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Butman:1995:BRG

[But95] Cheryl Ann Butman. Book review: Global marine biological diversity: a strategy for building conservation into decision making (E. A. Norse [ed.]). Limnology and Oceanography, 40(2):447–448, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Boltovskoy:1996:SVF

[BUW96] Demetrio Boltovskoy, Eleonora Uliana, and Gerald Wefer. Seasonal variation in the flux of microplankton and radiolarian assemblage compositions in the northeastern tropical Atlantic at 2,195 m. Limnology and Oceanography, 41(4):615–635, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Boersma:1995:SNE

[BV95] Maarten Boersma and Jacobus Vijverberg. The significance of nonviable eggs for Daphnia population dynamics. Limnology and Oceanography, 40(7):1215–1224, November 1995. CODEN LI-OCAH. ISSN 0024-3590.

Bell:1993:SMF

[BVSBB93] Russell T. Bell, Katarina Vrede, Ulrika Stensdotter-Blomberg, and Peter Blomqvist. Stimulation of the microbial food web in an oligotrophic, slightly acidified lake. Limnology and Oceanography, 38(7):1532–1538, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Burkholder:1990:EAP

[BW90] JoAnn M. Burkholder and Robert G. Wetzel. Epiphytic alkaline phosphatase on natural and artificial plants in an oligotrophic lake: Re-evaluation of the role of macrophytes as a phosphorus source for epiphytes. Limnology and Oceanography, 35(3):736– 747, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Berger:1991:PGO

[BW91] W. H. Berger and G. Wefer. Productivity of the glacial ocean: Discussion of the iron hypothesis. Limnology and Oceanography, 36(8):1899–1918, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Barbeau:1994:MCL

[BW94a] Katherine Barbeau and Roland Wollast. Microautoradiography (with combined liquid scintillation) applied to the study of trace metal uptake by suspended particles: Initial results using ^{63}Ni as a tracer. Limnology and Oceanography, 39(5):1211–1222, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Benson:1994:SIOb

[BW94b] L. V. Benson and J. W. C. White. Stable isotopes of oxygen and hydrogen in the Truckee River–Pyramid Lake surface-water system. 3. Source of water vapor overlying Pyramid Lake. Limnology and Oceanography, 39(8):1945–1958, December 1994. CODEN LI-OCAH. ISSN 0024-3590.

Bohonak:1999:DFS

[BW99a] Andrew J. Bohonak and Howard H. Whiteman. Dispersal of the fairy shrimp Branchinecta coloradensis (Anostraca): Effects of hydroperiod and salamanders. Limnology and Oceanography, 44 (3):487–493, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Bronk:1999:GNN

[BW99b] D. A. Bronk and B. B. Ward. Gross and net nitrogen uptake and DON release in the euphotic zone of Monterey Bay, California. Limnology and Oceanography, 44(3):573–585, May 1999. CODEN LIOCAH. ISSN 0024-3590. See reply [SRG00].

Burdige:1998:BCD

[BZ98] David J. Burdige and Shilong Zheng. The biogeochemical cycling of dissolved organic nitrogen in estuarine sediments. Limnology and Oceanography, 43(8):1796–1813, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Burkhardt:1999:ECC

[BZR99] Steffen Burkhardt, Ingrid Zondervan, and Ulf Riebesell. Effect of CO_{2} concentration on C : N : P ratio in marine phytoplankton: a species comparison. Limnology and Oceanography, 44(3):683–690, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Crusius:1991:CCS

[CA91] John Crusius and Robert F. Anderson. Core compression and surficial sediment loss of lake sediments of high porosity caused by gravity coring. Limnology and Oceanography, 36(5):1021–1030, July 1991. CODEN LIOCAH. ISSN 0024-3590. See comment [CGSD93].

Crusius:1993:RCC

[CA93] John Crusius and Robert F. Anderson. Reply to the comments of Cumming et al. and Wright. Limnology and Oceanography, 38 (3):701–703, May 1993. CODEN LIOCAH. ISSN 0024-3590. See [CGSD93, Wri93].

Cattaneo:1998:ORC

[CACM98] Antonella Cattaneo, Alessandra Asioli, Patrizia Comoli, and Marina Manca. Organisms’ response in a chronically polluted lake supports hypothesized link between stress and size. Limnology and Oceanography, 43(8):1938–1943, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Camarero:1994:ASR

[Cam94] Lluís Camarero. Assay of soluble reactive phosphorus at nanomolar levels in nonsaline waters. Limnology and Oceanography, 39 (3):707–711, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Carignan:1998:ADC

[Car98] Richard Carignan. Automated determination of carbon dioxide, oxygen, and nitrogen partial pressures in surface waters. Limnology and Oceanography, 43(5):969–975, July 1998. CODEN LI-OCAH. ISSN 0024-3590.

Carrick:1993:WIP

[CAS93] Hunter J. Carrick, Frederick J. Aldridge, and Claire L. Schelske. Wind influences phytoplankton biomass and composition in a shallow, productive lake. Limnology and Oceanography, 38(6): 1179–1192, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Chavez:1995:CRP

[CBB^{+}95] Francisco P. Chavez, Kurt R. Buck, Robert R. Bidigare, David M. Karl, Dale Hebel, Mikel Latasa, Lisa Campbell, and Jan Newton. On the chlorophyll-a retention properties of glass-fiber GF/F filters. Limnology and Oceanography, 40(2):428–433, March 1995. CODEN LIOCAH. ISSN 0024-3590. See reply [DW95b].

Cable:1996:ACA

[CBBC96] Jaye E. Cable, Glynnis C. Bugna, William C. Burnett, and Jeffrey P. Chanton. Application of ^{222}Rn and CH_{4} for assessment of groundwater discharge to the coastal ocean. Limnology and Oceanography, 41(6):1347–1353, September 1996. CODEN LI-OCAH. ISSN 0024-3590.

Chavez:1991:GRG

[CBC^{+}91] Francisco P. Chavez, Kurt R. Buck, K. H. Coale, J. H. Martin, G. R. DiTullio, N. A. Welschmeyer, Aubrie C. Jacobson, and Richard T. Barber. Growth rates, grazing, sinking, and iron limitation of equatorial Pacific phytoplankton. Limnology and Oceanography, 36(8):1816–1833, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Cherrier:1999:RMB

[CBD^{+}99] Jennifer Cherrier, James E. Bauer, Ellen R. M. Druffel, Richard B. Coffin, and Jeffrey P. Chanton. Radiocarbon in marine bacteria: Evidence for the ages of assimilated carbon. Limnology and Oceanography, 44(3):730–736, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Cavender-Bares:1998:DSF

[CBFC98] Kent K. Cavender-Bares, Sheila L. Frankel, and Sallie W. Chisholm. A dual sheath flow cytometer for shipboard analyses of phytoplankton communities from the oligotrophic oceans. Limnology and Oceanography, 43(6):1383–1388, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Cavender-Bares:1999:DRE

[CBMC^{+}99] Kent K. Cavender-Bares, Elizabeth L. Mann, Sallie W. Chisholm, Michael E. Ondrusek, and Robert R. Bidigare. Differential response of equatorial Pacific phytoplankton to iron fertilization. Limnology and Oceanography, 44(2):237–246, March 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Carvalho:1999:CBS

[CBS99] Roberta A. Carvalho, Mark C. Benfield, and Peter H. Santschi. Comparative bioaccumulation studies of colloidally complexed and free-ionic heavy metals in juvenile brown shrimp Penaeus aztecus (Crustacea: Decapoda: Penaeidae). Limnology and Oceanography, 44(2):403–414, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Chant:1991:SGC

[CC91] L. A. Chant and R. J. Cornett. Smearing of gravity core profiles in soft sediments. Limnology and Oceanography, 36(7):1492–1498, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Crawford:1997:ODM

[CC97] G. B. Crawford and R. W. Collier. Observations of a deep-mixing event in Crater Lake, Oregon. Limnology and Oceanography, 42 (2):299–306, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Cole:1998:AEC

[CC98] Jonathan J. Cole and Nina F. Caraco. Atmospheric exchange of carbon dioxide in a low-wind oligotrophic lake measured by the addition of SF_{6}. Limnology and Oceanography, 43(4):647–656, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Corbett:1999:PGD

[CCB^{+}99] D. Reide Corbett, Jeffrey Chanton, William Burnett, Kevin Dillon, Christine Rutkowski, and James W. Fourqurean. Patterns of groundwater discharge into Florida Bay. Limnology and Oceanography, 46(4):1045–1055, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Christensen:1996:PRC

[CCC^{+}96] David L. Christensen, Stephen R. Carpenter, Kathryn L. Cottingham, Susan E. Knight, Joseph P. LeBouton, Daniel E. Schindler, Nicholas Voichick, Jonathan J. Cole, and Michael L. Pace. Pelagic responses to changes in dissolved organic carbon following division of a seepage lake. Limnology and Oceanography, 41(3):553–559, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Cullen:1997:ODA

[CCDL97] John J. Cullen, Áurea M. Ciotti, Richard F. Davis, and Marlon R. Lewis. Optical detection and assessment of algal blooms. Limnology and Oceanography, 42(5 (part 2)):1223–1239, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Cambalik:1998:NMM

[CCK98] John J. Cambalik, David M. Checkley, Jr., and Daniel Kamykowski. A new method to measure the terminal velocity of small particles: a demonstration using ascending eggs of the Atlantic menhaden (Brevoortia tyrannus). Limnology and Oceanography, 43(7):1722–1727, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1998:IDO

[CCKP98] Stephen R. Carpenter, Jonathan J. Cole, James F. Kitchell, and Michael L. Pace. Impact of dissolved organic carbon, phosphorus, and grazing on phytoplankton biomass and production in experimental lakes. Limnology and Oceanography, 43(1):73–80, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Cole:1990:SRA

[CCL90] Jonathan J. Cole, Nina F. Caraco, and Gene E. Likens. Shortrange atmospheric transport: a significant source of phosphorus to an oligotrophic lake. Limnology and Oceanography, 35(6):1230– 1237, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Caraco:1992:NRP

[CCL92] Nina F. Caraco, Jonathan J. Cole, and Gene E. Likens. New and recycled primary production in an oligotrophic lake: Insights for summer phosphorus dynamics. Limnology and Oceanography, 37 (3):590–602, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Cole:1992:CPM

[CCP92] Jonathan J. Cole, Nina F. Caraco, and Benjamin L. Peierls. Can phytoplankton maintaina positive carbon balance in a turbid, freshwater, tidal estuary? Limnology and Oceanography, 37(8): 1608–1617, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Canuel:1995:MIT

[CCR^{+}95] Elizabeth A. Canuel, James E. Cloern, David B. Ringelberg, James B. Guckert, and Greg H. Rau. Molecular and isotopic tracers used to examine sources of organic matter and its incorporation into the food webs of San Francisco Bay. Limnology and Oceanography, 40(1):67–81, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Couillard:1993:RMC

[CCT93] Yves Couillard, Peter G. C. Campbell, and André Tessier. Response of metallothionein concentrations in a freshwater bivalve (Anodonta grandis) along an environmental cadmium gradient. Limnology and Oceanography, 38(2):299–313, March 1993. CO-DEN LIOCAH. ISSN 0024-3590.

Cuthbert:1992:TSM

[CdG92] Iain D. Cuthbert and Paul del Giorgio. Toward a standard method of measuring color in freshwater. Limnology and Oceanography, 37(6):1319–1326, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Carlson:1998:OCP

[CDHS98] Craig A. Carlson, Hugh W. Ducklow, Dennis A. Hansell, and Walker O. Smith, Jr. Organic carbon partitioning during spring phytoplankton blooms in the Ross Sea polynya and the Sargasso Sea. Limnology and Oceanography, 43(3):375–386, May 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Callaway:1996:CUD

[CDP96] J. C. Callaway, R. D. DeLaune, and W. H. Patrick, Jr. Chernobyl^{137}Cs used to determine sediment accretion rates at selected northern European coastal wetlands. Limnology and Oceanography, 41(3):444–450, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Christaki:1998:CPS

[CDPR98] Urania Christaki, John R. Dolan, Silvia Pelegri, and Fereidoun Rassoulzadegan. Consumption of picoplankton-size particles by marine ciliates: Effects of physiological state of the ciliate and particle quality. Limnology and Oceanography, 43(3):458–464, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Clair:1996:VDD

[CE96] Thomas A. Clair and James M. Ehrman. Variations in discharge and dissolved organic carbon and nitrogen export from terrestrial basins with changes in climate: a neural network approach. Limnology and Oceanography, 41(5):921–927, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Cifuentes:1998:MIB

[CE98] Luis A. Cifuentes and Peter M. Eldridge. A mass- and isotopebalance model of DOC mixing in estuaries. Limnology and Oceanography, 43(8):1872–1882, December 1998. CODEN LI-OCAH. ISSN 0024-3590.

Chen:1996:CFW

[CF96] C. Y. Chen and C. L. Folt. Consequences of fall warming for zooplankton over wintering success. Limnology and Oceanography, 41(5):1077–1086, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Carman:1997:RBF

[CFP97] Kevin R. Carman, John W. Fleeger, and Steven M. Pomarico. Response of a benthic food web to hydrocarbon contamination. Limnology and Oceanography, 42(3):561–571, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Carrick:1991:IZP

[CFSW91] Hunter J. Carrick, Gary L. Fahnenstiel, Eugene F. Stoermer, and Robert G. Wetzel. The importance of zooplankton-protozoan trophic couplings in Lake Michigan. Limnology and Oceanography, 36(7):1335–1345, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Currick:1992:GPP

[CFT92] Hunter J. Currick, Gary L. Fahnenstiel, and William D. Taylor. Growth and production of planktonic protozoa in Lake Michigan: In situ versus in vitro comparisons and importance to food web dynamics. Limnology and Oceanography, 37(6):1221–1235, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Canuel:1997:ICL

[CFW97] Elizabeth A. Canuel, Katherine H. Freeman, and Stuart G. Wakeham. Isotopic compositions of lipid biomarker compounds in estuarine plants and surface sediments. Limnology and Oceanography, 42(7):1570–1583, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Chisholm:1991:VAA

[CG91] John R. M. Chisholm and Jean-Pierre Gattuso. Validation of the alkalinity anomaly technique for investigating calcification of photosynthesis in coral reef communities. Limnology and Oceanography, 36(6):1232–1239, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Caron:1993:PER

[CG93] David A. Caron and Joel C. Goldman. Predicting excretion rates of protozoa: Reply to the comment by Landry. Limnology and Oceanography, 38(2):472–474, March 1993. CODEN LIOCAH. ISSN 0024-3590. See [Lan93].

Chandy:1995:EPI

[CG95] Shonali T. Chandy and Charles H. Greene. Estimating the predatory impact of gelatinous zooplankton. Limnology and Oceanography, 40(5):947–955, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Callender:1997:GMB

[CG97] Edward Callender and Liba Granina. Geochemical mass balances of major elements in Lake Baikal. Limnology and Oceanography, 42(1):148–155, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Cuker:1990:TSC

[CGB90] Benjamin E. Cuker, Phumelele T. Gama, and JoAnn M. Burkholder. Type of suspended clay influences lake productivity and phytoplankton community response to phosphorus loading. Limnology and Oceanography, 35(4):830–839, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Caron:1990:CUO

[CGD90] David A. Caron, Joel C. Goldman, and Mark R. Dennett. Carbon utilization by the omnivorous flagellate Paraphysomonas imperforata. Limnology and Oceanography, 35(1):192–201, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Cleckner:1999:MMP

[CGHK99] Lisa B. Cleckner, Cynthia C. Gilmour, James P. Hurley, and David P. Krabbenhoft. Mercury methylation in periphyton of the Florida Everglades. Limnology and Oceanography, 44(7):1815– 1825, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Cumming:1993:CCC

[CGSD93] Brian F. Cumming, John R. Glew, John P. Smol, and Ronald B. Davis. Comment on “Core compression and surficial sediment loss of lake sediments of high porosity caused by gravity coring” (Crusius and Anderson). Limnology and Oceanography, 38(3): 695–699, May 1993. CODEN LIOCAH. ISSN 0024-3590. See [CA91], comment [Wri93], and reply [Wri93].

Cloern:1995:EMP

[CGVL95] James E. Cloern, Christian Grenz, and Lisa Vidergar-Lucas. An empirical model of the phytoplankton chlorophyll : carbon ratio — the conversion factor between productivity and growth rate. Limnology and Oceanography, 40(7):1313–1321, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Cotner:1990:IRE

[CH90] James B. Cotner, Jr. and Robert T. Heath. Iron redox effects on photosensitive phosphorus release from dissolved humic materials. Limnology and Oceanography, 35(5):1175–1181, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Clements:1991:CFP

[CH91] Lee Ann J. Clements and Thomas J. Hilbish. Comparison of fluorescamine and o-phthaldialdehyde methods for measuring amino acid exchange in marine organisms. Limnology and Oceanography, 36(7):1463–1471, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Cowie:1992:SRA

[CH92a] Gregory L. Cowie and John I. Hedges. Sources and reactivities of amino acids in a coastal marine environment. Limnology and Oceanography, 37(4):703–724, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Cuker:1992:TSC

[CH92b] Benjamin E. Cuker and Leon Hudson, Jr. Type of suspended clay influences zooplankton response to phosphorus loading. Limnology and Oceanography, 37(3):566–576, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Cranford:1994:STS

[CH94] P. J. Cranford and B. T. Hargrave. In situ time-series measurement of ingestion and absorption rates of suspension-feeding bivalves: Placopecten magellanicus. Limnology and Oceanography, 39(3):730–738, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Cowie:1996:DAB

[CH96] Gregory L. Cowie and John I. Hedges. Digestion and alteration of the biochemical constituents of a diatom (Thalassiosira weisflogii) ingested by an herbivorous zooplankton (Calanus pacificus). Limnology and Oceanography, 41(4):581–594, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Cerny:1999:IAD

[ČH99] Martin Černý and Paul D. N. Hebert. Intercontinental allozyme differentiation among four holarctic Daphnia species. Limnology and Oceanography, 44(6):1381–1387, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1991:MPP

[CHA91] Robert C. Carpenter, John M. Hackney, and Walter H. Adey. Measurements of primary productivity and nitrogenase activity of coral reef algae in a chamber incorporating oscillatory flow. Limnology and Oceanography, 36(1):40–49, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Chapman:1996:BRU

[Cha96] Piers Chapman. Book review: Upwelling in the ocean: Modern processes and ancient records (C. P. Summerhayes, K.-C. Emeis, M. V. Angel, R. L. Smith, and B. Zeitzschel [eds.]). Limnology and Oceanography, 41(7):1585–1586, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Charbonneau:1997:UXR

[CHC97] Patrick Charbonneau, Landis Hare, and Richard Carignan. Use of X-ray images and a contrasting agent to study the behavior of animals in soft sediments. Limnology and Oceanography, 42(8): 1823–1828, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Cochlan:1991:DPN

[CHD91] W. P. Cochlan, P. J. Harrison, and K. L. Denman. Diel periodicity of nitrogen uptake by marine phytoplankton in nitrate-rich environments. Limnology and Oceanography, 36(8):1689–1700, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Che:1994:DNM

[CI94] Guangwei Che and Johann Ilmberger. Development of a new method for direct measurement of pCO_{2} in natural waters. Limnology and Oceanography, 39(4):976–981, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Crumpton:1992:NOA

[CIM92] William G. Crumpton, Thomas M. Isenhart, and Paul D. Mitchell. Nitrate and organic N analyses with second-derivative spectroscopy. Limnology and Oceanography, 37(4):907–913, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Christaki:1999:GGP

[CJD^{+}99] Urania Christaki, Stéphan Jacquet, John R. Dolan, Daniel Vaulot, and Fereidoun Rassoulzadegan. Growth and grazing on Prochlorococcus and Synechococcus by two marine ciliates. Limnology and Oceanography, 44(1):52–61, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Chevaldonne:1997:HVA

[CJVD97] Pierre Chevaldonne, Didier Jollivet, Annick Vangriesheim, and Daniel Desbruyères. Hydrothermal-vent alvinellid polychaete dispersal in the eastern Pacific. 1. Influence of vent site distribution, bottom currents, and biological patterns. Limnology and Oceanography, 42(1):67–80, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1992:TCB

[CK92a] Stephen R. Carpenter and James F. Kitchell. Trophic cascade and biomanipulation: Interface of research and management — a reply to the comment by DeMelo et al. Limnology and Oceanography, 37 (1):208–213, January 1992. CODEN LIOCAH. ISSN 0024-3590. See [DFM92].

Cufrey:1992:ISP

[CK92b] Jane M. Cufrey and W. Michael Kemp. Influence of the submersed plant, Potamogeton perfoliatus, on nitrogen cycling in estuarine sediments. Limnology and Oceanography, 37(7):1483– 1495, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Christian:1995:BEM

[CK95] James R. Christian and David M. Karl. Bacterial ectoenzymes in marine waters: Activity ratios and temperature responses in three oceanographic provinces. Limnology and Oceanography, 40 (6):1042–1049, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1999:PRC

[CKCP99] S. R. Carpenter, J. F. Kitchell, J. J. Cole, and M. L. Pace. Predicting responses of chlorophyll and primary production to changes in phosphorus, grazing, and dissolved organic carbon (reply to comment by Nürnberg). Limnology and Oceanography, 46(4):1179– 1182, June 1999. CODEN LIOCAH. ISSN 0024-3590. See [Nür99].

Carlotti:1993:GDC

[CKR93] Fraçois Carlotti, Michael Krause, and Günther Radach. Growth and development of Calanus finmarchicus related to the influence of temperature: Experimental results and conceptual model. Limnology and Oceanography, 38(6):1125–1134, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Chalup:1990:TAP

[CL90] Michael S. Chalup and Edward A. Laws. A test of the assumptions and predictions of recent microalgal growth models with the marine phytoplankter Pavlova lutheri. Limnology and Oceanography, 35(3):583–596, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Carignan:1991:RDS

[CL91] R. Carignan and D. R. S. Lean. Regeneration of dissolved substances in a seasonally anoxic lake: The relative importance of processes occurring in the water column and in the sediments. Limnology and Oceanography, 36(4):683–707, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Clarke:1996:SCP

[CL96] Andrew Clarke and Raymond J. G. Leakey. The seasonal cycle of phytoplankton, macronutrients, and the microbial community in a nearshore Antarctic marine ecosystem. Limnology and Oceanography, 41(6):1281–1294, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Calbet:1999:MIM

[CL99] Albert Calbet and Michael R. Landry. Mesozooplankton influences on the microbial food web: Direct and indirect trophic interactions in the oligotrophic open ocean. Limnology and Oceanography, 44(6):1370–1380, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Claustre:1994:TSV

[Cla94] Hervé Claustre. The trophic status of various oceanic provinces as revealed by phytoplankton pigment signatures. Limnology and Oceanography, 39(5):1206–1210, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Carr:1995:PEN

[CLKJ95] Mary-Elena Carr, Marlon R. Lewis, Dan Kelley, and Burton Jones. A physical estimate of new production in the equatorial Pacific along 150^{o}W. Limnology and Oceanography, 40(1):138–147, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Cole:1993:MAC

[CLMH93] Jonathan J. Cole, Judith M. Lane, Roxanne Marino, and Robert W. Howarth. Molybdenum assimilation by cyanobacteria and phytoplankton in freshwater and salt water. Limnology and Oceanography, 38(1):25–35, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Cloern:1998:ECM

[Clo98] James E. Cloern. Eutrophication in coastal marine ecosystems. Limnology and Oceanography, 43(5):1018–1019, July 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Cary:1992:MTI

[CLP^{+}92] S. C. Cary, J. T. Lovette, P. J. Perl, M. E. Huntley, and M. Vernet. A microencapsulation technique for introducing pure compounds in zooplankton diets. Limnology and Oceanography, 37(2):404– 413, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Carlough:1991:BSP

[CM91] Leslie A. Carlough and Judy L. Meyer. Bacterivory by sestonic protists in a southeastern blackwater river. Limnology and Oceanography, 36(5):873–883, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Conversi:1994:NVH

[CM94] Alessandra Conversi and John A. McGowan. Natural versus human-caused variability of water clarity in the Southern California Bight. Limnology and Oceanography, 39(3):632–646, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Coffroth:1995:IMI

[CM95] Mary Alice Coffroth and John M. Mulawka III. Identification of marine invertebrate larvae by means of PCR-RAPD speciesspecific markers. Limnology and Oceanography, 40(1):181–189, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Cushing:1993:TDF

[CMN93] Colbert E. Cushing, G. Wayne Minshall, and J. Denis Newbold. Transport dynamics of fine particulate organic matter in two Idaho streams. Limnology and Oceanography, 38(6):1101–1115, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Cook:1992:AED

[CMPD92] Clayton B. Cook, Gisèle Muller-Parker, and Christopher F. D’Elia. Ammonium enhancement of dark carbon fixation and nitrogen limitation in symbiotic zooxanthellae: Effects of feeding and starvation of the sea anemone Aiptasia pallida. Limnology and Oceanography, 37(1):131–139, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Chiavelli:1993:HPS

[CMT93] Deborah A. Chiavelli, Edward L. Mills, and Stephen T. Threlkeld. Host preference, seasonality, and community interactions of zooplankton epibionts. Limnology and Oceanography, 38(3):574–583, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Crosby:1990:BMU

[CNL90] Michael P. Crosby, Roger I. E. Newell, and Christopher J. Langdon. Bacterial mediation in the utilization of carbon and nitrogen from detrital complexes by Crassostrea virginica. Limnology and Oceanography, 35(3):625–639, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Carignan:1994:LWH

[CNP94] Richard Carignan, Juan José Neiff, and Dolors Planas. Limitation of water hyacinth by nitrogen in subtropical lakes of the Paraná floodplain (Argentina). Limnology and Oceanography, 39(2):439– 443, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Christensen:1990:DNR

[CNSR90] Peter Bondo Christensen, Lars Peter Nielsen, Jan Sørensen, and Niels Peter Revsbech. Denitrification in nitrate-rich streams: Diurnal and seasonal variation related to benthic oxygen metabolism. Limnology and Oceanography, 35(3):640–651, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Campbell:1994:IPC

[CNV94] Lisa Campbell, H. A. Nolla, and Daniel Vaulot. The importance of Prochlorococcus to community structure in the central North Pacific Ocean. Limnology and Oceanography, 39(4):954–961, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Coale:1990:LDD

[Coa90] Kenneth H. Coale. Labyrinth of doom: a device to minimize the “swimmer” component in sediment trap collections. Limnology and Oceanography, 35(6):1376–1381, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Coale:1991:EIM

[Coa91] Kenneth H. Coale. Effects of iron, manganese, copper, and zinc enrichments on productivity and biomass in the subarctic Pacific.

Limnology and Oceanography, 36(8):1851–1864, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Conrad:1990:BRS

[Con90] Ralf Conrad. Book review: SEAREX: The sea/air exchange program (R. A. Duce [ed.]). Limnology and Oceanography, 35(4):984, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Conley:1997:RCB

[Con97] Daniel J. Conley. Riverine contribution of biogenic silica to the oceanic silica budget. Limnology and Oceanography, 42(4):774– 777, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Cottingham:1999:NCE

[Cot99] Kathryn L. Cottingham. National Center for Ecological Analysis and Synthesis, Santa Barbara, California 93101 and Center for Limnology, University of Wisconsin, Madison, Wisconsin 53706. Limnology and Oceanography, 44(3 (part 2)):810–827, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Carignan:1994:RNL

[CP94] Richard Carignan and Dolors Planas. Recognition of nutrient and light limitation in turbid mixed layers: Three approaches compared in the Parani floodplain (Argentina). Limnology and Oceanography, 39(3):580–596, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Cole:1995:BSP

[CP95a] Jonathan J. Cole and Michael L. Pace. Bacterial secondary production in oxic and anoxic freshwaters. Limnology and Oceanography, 40(6):1019–1027, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Cole:1995:WMB

[CP95b] Jonathan J. Cole and Michael L. Pace. Why measure bacterial production? A reply to the comment by Jahnke and Craven. Limnology and Oceanography, 40(2):441–444, March 1995. CODEN LIOCAH. ISSN 0024-3590. See [JC95].

Chase:1997:MCI

[CP97a] Z. Chase and N. M. Price. Metabolic consequences of iron deficiency in heterotrophic marine protozoa. Limnology and Oceanography, 42(8):1673–1684, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Crawford:1997:IPD

[CP97b] David W. Crawford and Duncan A. Purdie. Increase of P_{CO2} during blooms of Emiliania huxleyi: Theoretical considerations on the asymmetry between acquisition of HCO_{3}^{−} and respiration of free CO_{2}. Limnology and Oceanography, 42(2):365–372, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Culver:1999:RPA

[CP99] Mary E. Culver and Mary Jane Perry. The response of photosynthetic absorption coefficients to irradiance in culture and in tidally mixed estuarine waters. Limnology and Oceanography, 44 (1):24–36, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Cole:1993:BBC

[CPCS93] Jonathan J. Cole, Michael L. Pace, Nina F. Caraco, and Gail S. Steinhart. Bacterial biomass and cell size distributions in lakes: More and larger cells in anoxic waters. Limnology and Oceanography, 38(8):1627–1632, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Caron:1999:PNT

[CPLD99] David A. Caron, Emily R. Peele, Ee Lin Lim, and Mark R. Dennett. Picoplankton and nanoplankton and their trophic coupling in surface waters of the Sargasso Sea south of Bermuda. Limnology and Oceanography, 44(2):259–272, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Cai:1999:OCD

[CPMW99] Wei-Jun Cai, Lawrence R. Pomeroy, Mary Ann Moran, and Yongchen Wang. Oxygen and carbon dioxide mass balance for the estuarine-intertidal marsh complex of five rivers in the southeastern U.S. Limnology and Oceanography, 44(3):639–649, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Caron:1990:CNP

[CPS90] David A. Caron, Karen G. Porter, and Robert W. Sanders. Carbon, nitrogen, and phosphorus budgets for the mixotrophic phytoflagellate Poterioochromonas malhamensis (Chrysophyceae) during bacterial ingestion. Limnology and Oceanography, 35(2): 433–443, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Currin:1990:IDC

[CPSA90] Carolyn A. Currin, Hans W. Paerl, Gregory K. Suba, and Randall S. Alberte. Immunofluorescence detection and characterization of N_{2}-fixing microorganisms from aquatic environments. Limnology and Oceanography, 35(1):59–71, January 1990. CO-DEN LIOCAH. ISSN 0024-3590.

Chillrud:1994:CWP

[CPT^{+}94] S. N. Chillrud, F. L. Pedrozo, P. F. Temporetti, H. F. Planas, and P. N. Froelich. Chemical weathering of phosphate and germanium in glacial meltwater streams: Effects of subglacial pyrite oxidation. Limnology and Oceanography, 39(5):1130–1140, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Cai:1993:DPP

[CR93a] Wei-Jun Cai and Clare E. Reimers. The development of pH and pCO_{2} microelectrodes for studying the carbonate chemistry of pore waters near the sediment–water interface. Limnology and Oceanography, 38(8):1762–1773, December 1993. CODEN LI-OCAH. ISSN 0024-3590.

Casamitjana:1993:RSF

[CR93b] X. Casamitjana and E. Roget. Resuspension of sediment by focused groundwater in Lake Banyoles. Limnology and Oceanography, 38(3):643–656, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Carlotti:1996:SDP

[CR96] François Carlotti and Günther Radach. Seasonal dynamics of phytoplankton and Calanus finmarchicus in the North Sea as revealed by a coupled one-dimensional model. Limnology and Oceanography, 41(3):522–539, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Connolly:1999:IRN

[CR99] Sean R. Connolly and Joan Roughgarden. Increased recruitment of northeast Pacific barnacles during the 1997 El Niño. Limnology and Oceanography, 44(2):466–469, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Crawford:1995:RSS

[Cra95] Richard M. Crawford. The role of sex in the sedimentation of a marine diatom bloom. Limnology and Oceanography, 40(1):200– 204, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Christoffersen:1993:PRF

[CRKS93] Kirsten Christoffersen, Bo Riemann, Annette Klysner, and Morten Søndergaard. Potential role of fish predation and natural populations of zooplankton in structuring a plankton community in eutrophic lake water. Limnology and Oceanography, 38 (3):561–573, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Chrzanowski:1990:PSS

[CŠ90] Thomas H. Chrzanowski and Karel Šimek. Prey-size selection by freshwater flagellated protozoa. Limnology and Oceanography, 35 (7):1429–1436, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Cottrell:1995:DLV

[CS95] Matthew T. Cottrell and Curtis A. Suttle. Dynamics of lytic virus infecting the photosynthetic marine picoflagellate Micromonas pusilla. Limnology and Oceanography, 40(4):730–739, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Carrick:1997:WOI

[CS97] Hunter J. Carrick and Claire L. Schelske. Have we overlooked the importance of small phytoplankton in productive waters? Limnology and Oceanography, 42(7):1613–1621, November 1997. CO-DEN LIOCAH. ISSN 0024-3590.

Cota:1994:PPG

[CSM94] Glenn F. Cota, Walker O. Smith, and B. Greg Mitchell. Photosynthesis of Phaeocystis in the Greenland Sea. Limnology and Oceanography, 39(4):948–953, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Carignan:1994:UDS

[CSPG94] Richard Carignan, Sylvie St-Pierre, and Rene Gachter. Use of diffusion samplers in oligotrophic lake sediments: Effects of free oxygen in sampler material. Limnology and Oceanography, 39(2): 468–474, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Choi:1996:RBP

[CSS96] Joon W. Choi, Evelyn B. Sherr, and Barry F. Sherr. Relation between presence-absence of a visible nucleoid and metabolic activity in bacterioplankton cells. Limnology and Oceanography, 41 (6):1161–1168, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Childress:1993:EHP

[CT93] J. J. Childress and E. V. Thuesen. Effects of hydrostatic pressure on metabolic rates of six species of deep-sea gelatinous zooplankton. Limnology and Oceanography, 38(3):665–670, May 1993. CO-DEN LIOCAH. ISSN 0024-3590.

Chapman:1996:BRC

[CT96] Piers Chapman and David W. Townsend. Book review: From Cape Cod to the Bay of Fundy (P. W. Conkling [ed.]). Limnology and Oceanography, 41(7):1586–1587, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Canfield:1998:IFS

[CTF98] D. E. Canfield, B. Thamdrup, and S. Fleischer. Isotope fractionation and sulfur metabolism by pure and enrichment cultures of elemental sulfur-disproportionating bacteria. Limnology and Oceanography, 43(2):253–264, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Commito:1995:DDW

[CTP^{+}95] J. A. Commito, S. F. Thrush, R. D. Pridmore, J. E. Hewitt, and V. J. Cummings. Dispersal dynamics in a wind-driven benthic system. Limnology and Oceanography, 40(8):1513–1518, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Chin:1998:APD

[CTSB98] Yu-Ping Chin, Samuel J. Traina, Cathleen R. Swank, and Debera Backhus. Abundance and properties of dissolved organic matter in pore waters of a freshwater wetland. Limnology and Oceanography, 43(6):1287–1296, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Cullen:1991:HEH

[Cul91] John J. Cullen. Hypotheses to explain high-nutrient conditions in the open sea. Limnology and Oceanography, 36(8):1578–1599, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Cullen:1995:SIH

[Cul95] John J. Cullen. Status of the iron hypothesis after the openocean enrichment experiment. Limnology and Oceanography, 40 (7):1336–1343, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Currie:1990:LSV

[Cur90] David J. Currie. Large-scale variability and interactions among phytoplankton, bacterioplankton, and phosphorus. Limnology and Oceanography, 35(7):1437–1455, November 1990. CODEN LI-OCAH. ISSN 0024-3590.

Curtis:1993:EDO

[Cur93] P. Jeferson Curtis. Effect of dissolved organic carbon on 59Fe scavenging. Limnology and Oceanography, 38(7):1554–1561, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Curr:1994:BRM

[Cur94] N. G. Curr. Book review: Marine pelagic cyanobacteria: Trichodesmium and other diazotrophs (E. J. Carpenter, D. G. Capone, and J. G. Rueter [eds.]). Limnology and Oceanography, 39(6):1498–1499, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Carignan:1999:PAP

[CV99] R. Carignan and P. Vaithiyanathan. Phosphorus availability in the Paraná floodplain lakes (Argentina): Influence of pH and phosphate buffering by fluvial sediments. Limnology and Oceanography, 44(6):1540–1548, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Cotner:1992:UDI

[CW92] James B. Cotner, Jr. and Robert G. Wetzel. Uptake of dissolved inorganic and organic bphosphorus compounds by phytoplankton and bacterioplankton. Limnology and Oceanography, 37(2):232– 243, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1993:EAT

[CW93a] Robert C. Carpenter and Susan L. Williams. Effects of algal turf canopy height and microscale substratum topography on profiles of flow speed in a coral forereef environment. Limnology and Oceanography, 38(3):687–694, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Cleveland:1993:QAA

[CW93b] J. S. Cleveland and A. D. Weidemann. Quantifying absorption by aquatic particles: a multiple scattering correction for glass-fiber filters. Limnology and Oceanography, 38(6):1321–1327, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Coveney:1995:BPS

[CW95] Michael F. Coveney and Robert G. Wetzel. Biomass, production, and specific growth rate of bacterioplankton and coupling to phytoplankton in an oligotrophic lake. Limnology and Oceanography, 40(7):1187–1200, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Cai:1998:CFS

[CW98] Wei-Jun Cai and Yongchen Wang. The chemistry, fluxes, and sources of carbon dioxide in the estuarine waters of the satilla and altamaha rivers, Georgia. Limnology and Oceanography, 43 (4):657–668, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Chandler:1996:SME

[CWSX96] G. Thomas Chandler, Douglas F. Williams, Howard J. Spero, and Gao Xiaodong. Sediment microhabitat effects on carbon stable isotopic signatures of microcosm-cultured benthic Foraminifera. Limnology and Oceanography, 41(4):680–688, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Cao:1998:HIR

[CWW98] Yong Cao, Dudley D. Williams, and Nancy E. Williams. How important are rare species in aquatic community ecology and bioassessment? Limnology and Oceanography, 43(7):1403–1409, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Davis:1992:ECI

[DA92] Cubell S. Davis and Philip Alatalo. Effects of constant and intermittent food supply on life-history parameters in a marine copepod. Limnology and Oceanography, 37(8):1618–1639, December 1992. CODEN LIOCAH. ISSN 0024-3590.

deAngelis:1991:EMM

[dABL91] Marie A. de Angelis, John A. Baross, and Marvin D. Lilley. Enhanced microbial methane oxidation in water from a deep-sea hydrothermal vent field at simulated in situ hydrostatic pressures. Limnology and Oceanography, 36(3):565–570, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Duarte:1990:SPF

[DAC90] Carlos M. Duarte, Susana Agusti, and Daniel E. Canfield, Jr. Size plasticity of freshwater phytoplankton: Implications for community structure. Limnology and Oceanography, 35(8):1846–1851, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Duarte:1992:PPC

[DAC92] Carlos M. Duarte, Susana Agustí, and Daniel E. Canjield, Jr. Patterns in phytoplankton community structure in Florida lakes. Limnology and Oceanography, 37(1):155–161, January 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Dade:1993:NBT

[Dad93] W. Brian Dade. Near-bed turbulence and hydrodynamic control of diffusional mass transfer at the sea floor. Limnology and Oceanography, 38(1):52–69, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Daly:1990:ODG

[Dal90] Kendra L. Daly. Overwintering development, growth, and feeding of larval Euphausia superba in the Antarctic marginal ice zone. Limnology and Oceanography, 35(7):1564–1576, November 1990. CODEN LIOCAH. ISSN 0024-3590.

deAngelis:1994:MPD

[dAL94] Marie A. de Angelis and Cindy Lee. Methane production during zooplankton grazing on marine phytoplankton. Limnology and Oceanography, 39(6):1298–1308, September 1994. CODEN LIOCAH. ISSN 0024-3590.

deAngelis:1991:SHV

[dARB91] Marie A. de Angelis, Anna-Louise Reysenbach, and John A. Baross. Surfaces of hydrothermal vent invertebrates: Sites of elevated microbial CH_{4} oxidation activity. Limnology and Oceanography, 36(3):570–577, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Duarte:1998:PPL

[DASV98] Carlos M. Duarte, Susana Agustí, Maria Paola Satta, and Dolors Vaqué. Partitioning particulate light absorption: a budget for a Mediterranean bay. Limnology and Oceanography, 43(2):236–244, March 1998. CODEN LIOCAH. ISSN 0024-3590.

DeMeester:1993:HVC

[DB93] Luc De Meester and Natalie Beenaerts. Heritable variation in carotenoid content in Daphnia magna. Limnology and Oceanography, 38(6):1193–1199, September 1993. CODEN LIOCAH. ISSN 0024-3590.

deBeer:1997:FRC

[dBGEK97] Dirk de Beer, Anni Glud, Eric Epping, and Michael Kûhl. A fastresponding CO_{2} microelectrode for profiling sediments, microbial mats, and biofilms. Limnology and Oceanography, 42(7):1590– 1600, November 1997. CODEN LIOCAH. ISSN 0024-3590.

DeMeester:1999:ESL

[DBS99] Luc De Meester, Maarten Boersma, and Piet Spaak. Environmental stress and local adaptation in Daphnia magna. Limnology and Oceanography, 44(2):393–402, March 1999. CODEN LIOCAH. ISSN 0024-3590.

DeVitre:1991:SAA

[DBT91] Richard De Vitre, Nelson Belzile, and André Tessier. Speciation and adsorption of arsenic on diagenetic iron oxyhydroxides. Limnology and Oceanography, 36(7):1480–1485, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Decho:1998:CCW

[DBZF98] Alan W. Decho, Kenneth A. Browne, and Richard K. Zimmer-Faust. Chemical cues: Why basic peptides are signal molecules in marine environments. Limnology and Oceanography, 43(7):1410– 1417, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Dini:1991:EWL

[DC91a] Michael L. Dini and Stephen R. Carpenter. he effect of wholelake fish community manipulations on Daphnia migratory behavior. Limnology and Oceanography, 36(2):370–377, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Dolan:1991:PPD

[DC91b] John R. Dolan and D. Wayne Coats. Preliminary prey digestion in a predacious estuarine ciliate and the use of digestion data to estimate ingestion. Limnology and Oceanography, 36(3):558–565, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Davidson:1996:ANP

[DC96a] K. Davidson and A. Cunningham. Accounting for nutrient processing time in mathematical models of phytoplankton growth. Limnology and Oceanography, 41(8):779–783, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Duarte:1996:FMA

[DC96b] Carlos M. Duarte and Just Cebrián. The fate of marine autotrophic production. Limnology and Oceanography, 41(8):1758– 1766, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Donahue:1999:EUR

[DC99] William F. Donahue and John J. Clare. Effects of ultraviolet radiation on trophic interactions not detected? Limnology and Oceanography, 44(2):470–473, March 1999. CODEN LIOCAH. ISSN 0024-3590. See reply [HM99].

Diaz:1990:NAM

[DCM90] Mara R. Diaz, Jorge E. Corredor, and Julio M. Morell. Nitrogenase activity of Microcoleus lyngbyaceus mat communities in a eutrophic, tropical marine environment. Limnology and Oceanography, 35(8):1788–1795, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Dymond:1996:UPF

[DCML96] Jack Dymond, Robert Collier, James McManus, and Gary L. Larson. Unbalanced particle flux budgets in Crater Lake, Oregon: Implications for edge effects and sediment focusing in lakes. Limnology and Oceanography, 41(4):732–743, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Durbin:1992:SCS

[DD92a] Ann G. Durbin and Edward G. Durbin. Seasonal changes in size frequency distribution and estimated age in the marine copepod Acartia hudsonica during a winter-spring diatom bloom in Narragansett Bay. Limnology and Oceanography, 37(2):379–392, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Durbin:1992:ETF

[DD92b] Edward G. Durbin and Ann G. Durbin. Effects of temperature and food abundance on grazing and short-term weight change in the marine copepod Acartia hudsonica. Limnology and Oceanography, 37(2):361–378, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Durbin:1992:BSE

[DDC92] Edward G. Durbin, Ann G. Durbin, and Robert G. Campbell. Body size and egg production in the marine copepod Acartia hudsonica during a winter-spring diatom bloom in Narragansett Bay. Limnology and Oceanography, 37(2):342–360, March 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Drapeau:1994:IFD

[DDG94] David T. Drapeau, Hans G. Dam, and Gary Grenie. An improved flocculator design for use in particle aggregation experiments. Limnology and Oceanography, 39(3):723–729, May 1994. CODEN LIOCAH. ISSN 0024-3590.

DeStasio:1990:RDE

[De 90] Burt T. De Stasio, Jr. The role of dormancy and emergence patterns in the dynamics of a freshwater zooplankton community. Limnology and Oceanography, 35(5):1079–1090, July 1990. CO-DEN LIOCAH. ISSN 0024-3590.

DeHaan:1993:SUL

[De 93] H. De Haan. Solar UV-light penetration and photodegradation of humic substances in peaty lake water. Limnology and Oceanography, 38(5):1072–1076, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Deng:1997:PRS

[Den97] Hong-Wen Deng. Photoperiodic response of sexual reproduction in the Daphnia pulex group is reversed in two distinct habitats. Limnology and Oceanography, 42(3):609–611, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Dusenberry:1994:ISF

[DF94] Jeflrey A. Dusenberry and Sheila L. Frankel. Increasing the sensitivity of a FACScan flow cytometer to study oceanic picoplankton. Limnology and Oceanography, 39(1):206–209, January 1994. CO-DEN LIOCAH. ISSN 0024-3590.

DeLong:1993:PDA

[DFA93] Edward F. DeLong, Diana G. Franks, and Alice L. Alldredge. Phylogenetic diversity of aggregate-attached vs. free-living marine bacterial assemblages. Limnology and Oceanography, 38(5):924– 934, July 1993. CODEN LIOCAH. ISSN 0024-3590.

DeMelo:1992:BHM

[DFM92] Rita DeMelo, Robert France, and Donald J. McQueen. Biomanipulation: Hit or myth? Limnology and Oceanography, 37(1): 192–207, January 1992. CODEN LIOCAH. ISSN 0024-3590. See reply [CK92a].

Dagg:1997:VMF

[DFN97] Michael J. Dagg, Bruce W. Frost, and Jan A. Newton. Vertical migration and feeding behavior of Calanus pacificus females during a phytoplankton bloom in Dabob Bay, U.S. Limnology and Oceanography, 42(5):974–980, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Denny:1990:PMI

[DG90] Mark W. Denny and Steven D. Gaines. On the prediction of maximal intertidal wave forces. Limnology and Oceanography, 35 (1):1–15, January 1990. CODEN LIOCAH. ISSN 0024-3590.

DeMott:1999:PLD

[DG99] William R. DeMott and Ramesh D. Gulati. Phosphorus limitation in Daphnia: Evidence from a long term study of three hypereutrophic Dutch lakes. Limnology and Oceanography, 44(6): 1557–1564, September 1999. CODEN LIOCAH. ISSN 0024-3590.

delGiorgio:1996:FCD

[dGBPP96] Paul A. del Giorgio, David F. Bird, Yves T. Prairie, and Dolors Planas. Flow cytometric determination of bacterial abundance in lake plankton with the green nucleic acid stain SYTO 13. Limnology and Oceanography, 41(8):783–789, June 1996. CODEN LIOCAH. ISSN 0024-3590.

delGiorgio:1996:ESP

[dGF96] Paul A. del Giorgio and Robert L. France. Ecosystem-specific patterns in the relationship between zooplankton and POM or microplankton δ ^{13}C. Limnology and Oceanography, 41(2):359– 365, March 1996. CODEN LIOCAH. ISSN 0024-3590.

delGiorgio:1996:BCS

[dGGV^{+}96] Paul A. del Giorgio, Josep M. Gasol, Dolors Vaqué, Paola Mura, Susana Agustí, and Carlos M. Duarte. Bacterioplankton community structure: Protists control net production and the proportion of active bacteria in a coastal marine community. Limnology and Oceanography, 41(6):1169–1179, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Denny:1998:MMD

[DGHD98] Mark Denny, Brian Gaylord, Brian Helmuth, and Tom Daniel. The menace of momentum: Dynamic forces on flexible organisms. Limnology and Oceanography, 43(5):955–968, July 1998. CODEN LIOCAH. ISSN 0024-3590.

delGiorgio:1994:PPE

[dGP94] Paul A. del Giorgio and Robert H. Peters. Patterns in planktonic P : R ratios in lakes: Influence of lake trophy and dissolved organic carbon. Limnology and Oceanography, 39(4):772–787, June 1994. CODEN LIOCAH. ISSN 0024-3590.

DeMott:1998:EPD

[DGS98] William R. DeMott, Ramesh D. Gulati, and Klaas Siewertsen. Effects of phosphorus-deficient diets on the carbon and phosphorus balance of Daphnia magna. Limnology and Oceanography, 43(6): 1147–1161, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Davison:1992:CLI

[DGW92] W. Davison, G. W. Grime, and C. Woof. Characterization of lacustrine iron sulfide particles with proton-induced X-ray emission. Limnology and Oceanography, 37(8):1770–1777, December 1992. CODEN LIOCAH. ISSN 0024-3590.

DiTullio:1993:IIM

[DHB93] Giacomo R. DiTullio, David A. Hutchins, and Kenneth W. Bruland. Interaction of iron and major nutrients controls phytoplankton growth and species composition in the tropical North Pacific Ocean. Limnology and Oceanography, 38(3):495–508, May 1993. CODEN LIOCAH. ISSN 0024-3590.

deHaan:1990:ATI

[dHJS90] Henk de Haan, Roger I. Jones, and Kalevi Salonen. Abiotic transformations of iron and phosphate in humic lake water revealed by double-isotope labeling and gel filtration. Limnology and Oceanography, 35(2):491–497, March 1990. CODEN LIOCAH. ISSN 0024-3590.

DeStasio:1996:PEG

[DHK^{+}96] Bart T. De Stasio, Jr., David K. Hill, Julie M. Kleinhans, Nathan P. Nibbelink, and John J. Magnuson. Potential effects of global climate change on small north-temperate lakes: Physics, fish, and plankton. Limnology and Oceanography, 41(5):1136– 1149, July 1996. CODEN LIOCAH. ISSN 0024-3590.

denHeyer:1998:OMM

[dHK98] Cornelia den Heyer and Jacob Kalff. Organic matter mineralization rates in sediments: a within- and among-lake study. Limnology and Oceanography, 43(4):695–705, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Duan:1996:PSG

[DHP^{+}96] W. M. Duan, D. B. Hedrick, K. Pye, M. L. Coleman, and D. C. White. A preliminary study of the geochemical and microbiological characteristics of modern sedimentary concretions. Limnology and Oceanography, 41(7):1404–1414, November 1996. CODEN LIOCAH. ISSN 0024-3590.

DeGrandpre:1995:SMS

[DHSS95] M. D. DeGrandpre, T. R. Hammar, S. P. Smith, and F. L. Sayles. In situ measurements of seawater pCO_{2}. Limnology and Oceanography, 41(4):969–975, July 1995. CODEN LIOCAH. ISSN 0024-3590.

DeGrandpre:1997:SMB

[DHWW97] M. D. DeGrandpre, T. R. Hammar, D. W. R. Wallace, and C. D. Wirick. Simultaneous mooring-based measurements of seawater CO_{2} and O_{2} off Cape Hatteras, North Carolina. Limnology and Oceanography, 42(1):21–28, January 1997. CODEN LIOCAH. ISSN 0024-3590.

deJorge:1995:WTI

[dJvB95] V. N. de Jorge and J. E. E. van Beusekom. Wind- and tide-induced resuspension of sediment and microphytobenthos from tidal flats in the Ems estuary. Limnology and Oceanography, 40(4):776–778, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Demers:1993:SMP

[DK93a] Eric Demers and Jacob Kalff. A simple model for predicting the date of spring stratification in temperate and subtropical lakes. Limnology and Oceanography, 38(5):1077–1081, July 1993. CO-DEN LIOCAH. ISSN 0024-3590.

Depetris:1993:CDS

[DK93b] Pedro J. Depetris and Stephan Kempre. Carbon dynamics and sources in the Parani River. Limnology and Oceanography, 38(2): 382–395, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Dore:1996:NEZ

[DK96] John E. Dore and David M. Karl. Nitrification in the euphotic zone as a source for nitrite, nitrate, and nitrous oxide at Station ALOHA. Limnology and Oceanography, 41(8):1619–1628, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Dawidowicz:1992:CSD

[DL92a] Piotr Dawidowicz and Carsten J. Loose. Cost of swimming by Daphnia during diel vertical migration. Limnology and Oceanography, 37(3):665–669, May 1992. CODEN LIOCAH. ISSN 0024-

Dawidowicz:1992:MCD

[DL92b] Piotr Dawidowicz and Carsten J. Loose. Metabolic costs during predator-induced dielvertical migration of Daphnia. Limnology and Oceanography, 37(8):1589–1595, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Decho:1993:EMM

[DL93] Alan W. Decho and Glenn R. Lopez. Exopolymer microenvironments of microbial flora: Multiple and interactive effects on trophic relationships. Limnology and Oceanography, 38(8):1633– 1645, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Decho:1996:FDS

[DL96] Alan W. Decho and Samuel N. Luoma. Flexible digestion strategies and trace metal assimilation in marine bivalves. Limnology and Oceanography, 41(3):568–572, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Decho:1990:BEU

[DM90] Alan W. Decho and David J. W. Moriarty. Bacterial exopolymer utilization by a harpacticoid copepod: a methodology and results. Limnology and Oceanography, 35(5):1039–1049, July 1990. CO-DEN LIOCAH. ISSN 0024-3590.

Downing:1992:NPR

[DM92] John A. Downing and Edward McCauley. The nitrogen : phosphorus relationship in lakes. Limnology and Oceanography, 37(5): 936–945, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Dauwe:1998:AAH

[DM98] Birgit Dauwe and Jack J. Middelburg. Amino acids and hexosamines as indicators of organic matter degradation state in North Sea sediments. Limnology and Oceanography, 43(5):782– 798, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Doran:1996:CFT

[DMA^{+}96] Peter T. Doran, Christopher P. McKay, W. Peter Adams, Michael C. English, Robert A. Wharton, and Michael A. Meyer. Climate forcing and thermal feedback of residual lake-ice covers in the high Arctic. Limnology and Oceanography, 41(5):839–848, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Duff:1998:MDS

[DMF^{+}98] John H. Duff, Fred Murphy, Chrsistopher C. Fuller, Frank J. Triska, Judson W. Harvey, and Alan P. Jackman. A mini drivepoint sampler for measuring pore water solute concentrations in the hyporheic zone of sand-bottom streams. Limnology and Oceanography, 43(6):1378–1383, September 1998. CODEN LI-OCAH. ISSN 0024-3590.

Dauwe:1999:LDA

[DMHH99] Birgit Dauwe, Jack J. Middelburg, Peter M. J. Herman, and Carlo H. R. Heip. Linking diagenetic alteration of amino acids and bulk organic matter reactivity. Limnology and Oceanography, 44(7): 1809–1814, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Duarte:1995:EID

[DMM95] Carlos M. Duarte, Merino Martín, and Gallegos Margarita. Evidence of iron deficiency in seagrasses growing above carbonate sediments. Limnology and Oceanography, 40(6):1153–1158, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Donaghay:1997:TTB

[DO97] Percy L. Donaghay and Thomas R. Osborn. Toward a theory of biological-physical control of harmful algal bloom dynamics and impacts. Limnology and Oceanography, 42(5 (part 2)):1283–1296, July 1997. CODEN LIOCAH. ISSN 0024-3590.

DuRand:1998:DPO

[DO98] Michele D. DuRand and Robert J. Olson. Diel patterns in optical properties of the chlorophyte Nannochloris sp.: Relating individual-cell to bulk measurements. Limnology and Oceanography, 43(6):1107–1118, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Dusenberry:1999:FDP

[DOC99] Jeffrey A. Dusenberry, Robert J. Olson, and Sallie W. Chisholm. Frequency distributions of phytoplankton single-cell fluorescence and vertical mixing in the surface ocean. Limnology and Oceanography, 44(2):431–435, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Dodson:1990:PDV

[Dod90] Stanley Dodson. Predicting diel vertical migration of zooplankton. Limnology and Oceanography, 35(5):1195–1200, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Dodds:1992:MFO

[Dod92a] Walter K. Dodds. A modified fiber-optic light microprobe to measure spherically integrated photosynthetic photon flux density: Characterization of periphyton photosynthesis-irradiance patterns. Limnology and Oceanography, 37(4):871–878, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Dodson:1992:PCZ

[Dod92b] Stanley Dodson. Predicting crustacean zooplankton species richness. Limnology and Oceanography, 37(4):848–856, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Dodson:1999:EEI

[Dod99] Stanley I. Dodson. The ecology and evolution of inducible defenses. Limnology and Oceanography, 44(6):1597, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Dolan:1991:BRE

[Dol91] John R. Dolan. Book review: Ecology of marine protozoa (G. M. Capriulo [ed.]). Limnology and Oceanography, 36(1):217–218, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Dawidowicz:1990:VMC

[DPC90] Piotr Dawidowicz, Joanna Pijanowska, and Konrad Ciechomski. Vertical migration of chaoborus larvae is induced by the presence of fish. Limnology and Oceanography, 35(7):1631–1637, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Dolan:1997:PIM

[DŠ97] John R. Dolan and Karel Šimek. Processing of ingested matter in Strombidium sulcatum, a marine ciliate (Oligotrichida). Limnology and Oceanography, 42(2):393–397, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Dolan:1998:IDA

[DŠ98] John R. Dolan and Karel Šimek. Ingestion and digestion of an autotrophic picoplankter, Synechococcus, by a heterotrophic nanoflagellate, Bodo saltans. Limnology and Oceanography, 43(7): 1740–1746, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Dolan:1999:DPS

[DŠ99] John R. Dolan and Karel Šimek. Diel periodicity in Synechococcus populations and grazing by heterotrophic nanoflagellates: Analysis of food vacuole contents. Limnology and Oceanography, 44(6): 1565–1570, September 1999. CODEN LIOCAH. ISSN 0024-3590.

DSa:1999:DOA

[DSV^{+}99] Eurico J. D’Sa, Robert G. Steward, Anthony Vodacek, Neil V. Blough, and Dave Phinney. Determining optical absorption of colored dissolved organic matter in seawater with a liquid capillary waveguide. Limnology and Oceanography, 46(4):1142–1148, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Duce:1991:ATI

[DT91] Robert A. Duce and Neil W. Tindale. Atmospheric transport of iron and its deposition in the ocean. Limnology and Oceanography, 36(8):1715–1726, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Duarte:1992:NCA

[Dua92] Carlos M. Duarte. Nutrient concentration of aquatic plants: Patterns across species. Limnology and Oceanography, 37(4):882–889, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Duarte:1995:BRM

[Dua95] Carlos Duarte. Book review: Microbial ecology of Lake Plußee (J. Overbeck and R. J. Chróst [ed.]). Limnology and Oceanography, 40(4):841–842, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Ducklow:1994:BRH

[Duc94] Hugh W. Ducklow. Book review: Handbook of methods in aquatic microbial ecology (P. F. Kemp, B. F. Sherr, E. B. Sherr, and J. J. Cole [eds.]). Limnology and Oceanography, 39(6):1496–1497, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Ducklow:1998:BOE

[Duc98] Hugh Ducklow. Bacteria in oligotrophic environments: Starvationsurvival lifestyle. Limnology and Oceanography, 43(5):1021–1022, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Dumont:1998:CLH

[Dum98] H. J. Dumont. The Caspian Lake: History, biota, structure, and function. Limnology and Oceanography, 43(1):44–52, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1991:LSN

[DW91] R. C. Dugdale and F. P. Wilkerson. Low specific nitrate uptake rate: a common feature of high-nutrient, low-chlorophyll marine ecosystems. Limnology and Oceanography, 36(8):1678–1688, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Dickson:1993:CCN

[DW93] Mary-Lynn Dickson and Patricia A. Wheeler. Chlorophyll a concentrations in the North Pacific: Does a latitudinal gradient exist? Limnology and Oceanography, 38(8):1813–1818, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Dickson:1995:NUR

[DW95a] Mary-Lynn Dickson and Patricia A. Wheele. Nitrate uptake rates in a coastal upwelling regime: a comparison of PN-specific, absolute, and chl a-specific rates. Limnology and Oceanography, 40(3): 533–543, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Dickson:1995:RNC

[DW95b] Mary-Lynn Dickson and Patricia A. Wheeler. Reply to the note by Chavez et al. Limnology and Oceanography, 40(2):434–436, March 1995. CODEN LIOCAH. ISSN 0024-3590. See [CBB^{+}95].

DeMeester:1999:DSB

[DW99] Luc De Meester and Lawrence J. Weider. Depth selection behavior, fish kairomones, and the life histories of Daphnia hyalina × galeata hybrid clones. Limnology and Oceanography, 44(5):1248– 1258, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1990:RNP

[DWM90] R. C. Dugdale, F. P. Wilkerson, and A. Morel. Realization of new production in coastal upwelling areas: a means to compare relative performance. Limnology and Oceanography, 35(4):822–829, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Dizon:1999:STR

[DY99] Romeo M. Dizon and Helen T. Yap. Short-term responses of coral reef microphytobenthic communities to inorganic nutrient loading. Limnology and Oceanography, 44(5):1259–1267, July 1999. CODEN LIOCAH. ISSN 0024-3590.

DeMott:1991:ETC

[DZC91] William R. DeMott, Qing-Xue Zhang, and Wayne W. Carmichael. Effects of toxic cyanobacteria and purified toxins on the survival and feeding of a copepod and three species of Daphnia. Limnology and Oceanography, 36(7):1346–1357, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Erdner:1999:FFB

[EA99] Deana L. Erdner and Donald M. Anderson. Ferredoxin and flavodoxin as biochemical indicators of iron limitation during openocean iron enrichment. Limnology and Oceanography, 44(7):1609– 1615, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Eiane:1999:FJQ

[EABK99] Ketil Eiane, Dag L. Aksnes, Espen BagØien, and Stein Kaartvedt. Fish or jellies — a question of visibility? Limnology and Oceanography, 44(5):1352–1357, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Ebert:1992:FIM

[Ebe92] Dieter Ebert. A food-independent maturation threshold and size at maturity in Daphnia magna. Limnology and Oceanography, 37 (4):878–881, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Eilser:1992:IGM

[EBH92] J. M. Eilser, D. F. Brakke, and A. Henriksen. The inapplicability of the Gibbs model of world water chemistry for dilute lakes. Limnology and Oceanography, 37(6):1335–1337, September 1992. CODEN LIOCAH. ISSN 0024-3590. See reply [Gib92].

Everett:1999:HPS

[ECA99] Christopher R. Everett, Yu-Ping Chin, and George R. Aiken. High-pressure size exclusion chromatography analysis of dissolved organic matter isolated by tangential-flow ultrafiltration. Limnology and Oceanography, 44(5):1316–1322, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Eckman:1990:MPS

[Eck90] James E. Eckman. A model of passive settlement by planktonic larvae onto bottoms of differing roughness. Limnology and Oceanography, 35(4):887–901, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1991:SRC

[Edm91a] W. T. Edmondson. Sedimentary record of changes in the condition of Lake Washington. Limnology and Oceanography, 36(5):1031– 1043, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1991:MGE

[Edm91b] Yvette H. Edmondson. In memoriam — G. Evelyn Hutchinson (1903–1991). Limnology and Oceanography, 36(3):618, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1996:BRF

[Edm96a] W. T. Edmondson. Book review: Freshwater algae: Their microscopic world explored (H. Canter-Lund and J. W. G.Lund). Limnology and Oceanography, 41(1):194–195, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1996:BRLb

[Edm96b] W. T. Edmondson. Book review: Limnologie generale (R. Pourriot, M. Meybeck, P. Champ, and J. Arcady-Meyer [eds.]). Limnology and Oceanography, 41(7):1583, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1996:BRLa

[Edm96c] W. T. Edmondson. Book review: Lough Neagh: The ecology of a multipurpose water resource (R. B. Wood and R. V. Smith [eds.]). Limnology and Oceanography, 41(6):1373, September 1996. CO-DEN LIOCAH. ISSN 0024-3590.

Elliott:1997:PHL

[EEL97] J. K. Elliott, J. M. Elliott, and W. C. Leggitt. Predation by Hydra on larval fish: Field and laboratory experiments with bluegill (Lepomis macrochirus). Limnology and Oceanography, 42 (6):1416–1423, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Elser:1995:MGS

[EF95] James J. Elser and David L. Frees. Microconsumer grazing and sources of limiting nutrients for phytoplankton growth: Application and complications of a nutrient-depletion/dilution-gradient technique. Limnology and Oceanography, 40(1):1–16, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Elser:1991:ZEP

[EG91a] James J. Elser and Charles R. Goldman. Zooplankton effects on phytoplankton in lakes of contrasting trophic status. Limnology and Oceanography, 36(1):64–90, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Emerson:1991:CSS

[EG91b] Craig W. Emerson and Jonathan Grant. The control of softshell clam (Mya arenaria) recruitment on intertidal sandflats by bedload sediment transport. Limnology and Oceanography, 36(7): 1288–1300, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Egeberg:1995:ASE

[Ege95] Per Kristian Egeberg. Acidification signals in estuarine sediment cores. Limnology and Oceanography, 40(8):1438–1446, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Eicken:1994:SUI

[Eic94] Hajo Eicken. Structure of under-ice melt ponds in the central Arctic and their effect on, the sea-ice cover. Limnology and Oceanography, 39(3):682–693, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Eppley:1990:NUP

[EK90] Richard W. Eppley and Wolfgang Koeve. Nitrate use by plankton in the eastern subtropical North Atlantic, March–April 1989. Limnology and Oceanography, 35(8):1781–1788, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Epping:1999:PDO

[EKT99] Eric H. G. Epping, Arzhang Khalili, and Roland Thar. Photosynthesis and the dynamics of oxygen consumption in a microbial mat as calculated from transient oxygen microprofiles. Limnology and Oceanography, 44(8):1936–1948, December 1999. CODEN LI-OCAH. ISSN 0024-3590.

Elia:1998:SDS

[Eli98] Christopher F. D. Elia. Sustainable development in the southeastern coastal zone. Limnology and Oceanography, 43(7):1762–1763, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Elser:1998:PNC

[Els98] James J. Elser. Pelagic nutrient cycles: Herbivores as sources and links. Limnology and Oceanography, 43(6):1401–1402, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Engle:1993:CRF

[EM93] Diana L. Engle and John M. Melack. Consequences of riverine flooding for seston and theperiphyton of floating meadows in an Amazon floodplain lake. Limnology and Oceanography, 38(7): 1500–1520, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Ederington:1995:TTF

[EMH95] Melissa C. Ederington, George B. McManus, and H. Rodger Harvey. Trophic transfer of fatty acids, sterols, and a triterpenoid alcohol between bacteria, a ciliate, and the copepod Acartia tonsa. Limnology and Oceanography, 40(5):860–867, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Evans:1996:PEC

[EP96] J. Catherine Evans and Ellie E. Prepas. Potential effects of climate change on ion chemistry and phytoplankton communities in prairie saline lakes. Limnology and Oceanography, 41(5):1063– 1076, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Evans:1997:RII

[EP97] J. Catherine Evans and Ellie E. Prepas. Relative importance of iron and molybdenum in restricting phytoplankton biomass in high phosphorus saline lakes. Limnology and Oceanography, 42 (3):461–472, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Epp:1996:GFC

[Epp96] Gregory T. Epp. Grazing on filamentous cyanobacteria by Daphnia pulicaria. Limnology and Oceanography, 41(3):560–567, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Espie:1995:BID

[ER95] Patrick J. Espie and John C. Roff. A biochemical index of duration of the molt cycle for planktonic Crustacea based on the chitindegrading enzyme, chitobiase. Limnology and Oceanography, 40 (6):1028–1034, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Engie:1990:AUS

[ES90] Diana L. Engie and Orlando Sarnelle. Algal use of sedimentary phosphorus from an Amazon floodplain lake: Implications for total phosphorus analysis in turbid waters. Limnology and Oceanography, 35(2):483–490, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Eldridge:1993:BHR

[ES93] Peter M. Eldridge and Michael E. Sieracki. Biological and hydrodynamic regulation of the microbial food web in a periodically mixed estuary. Limnology and Oceanography, 38(8):1666–1679, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Eaton:1996:ECW

[ES96] John G. Eaton and Robert M. Scheller. Effects of climate warming on fish thermal habitat in streams of the United States. Limnology and Oceanography, 41(5):1109–1115, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Edmond:1993:NCW

[ESC^{+}93] J. M. Edmond, R. F. Stallard, H. Craig, V. Craig, R. F. Weiss, and G. W. Coulter. Nutrient chemistry of the water column of Lake Tanganyika. Limnology and Oceanography, 38(4):725–738, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Ellis:1991:WTD

[ESG91] Christopher R. Ellis, Heinz G. Stefan, and Ruochuan Gu. Water temperature dynamics and heat transfer beneath the ice cover of a lake. Limnology and Oceanography, 36(2):324–334, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Eriksson:1999:ESE

[EW99] Peder G. Eriksson and Stefan E. B. Weisner. An experimental study on effects of submersed macrophytes on nitrification and denitrification in ammonium-rich aquatic systems. Limnology and Oceanography, 44(8):1993–1999, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Fogel:1999:BIC

[FAC^{+}99] Marilyn L. Fogel, Carmen Aguilar, Russell Cuhel, David J. Hollander, Joan D. Willey, and Hans W. Paerl. Biological and isotopic changes in coastal waters induced by Hurricane Gordon. Limnology and Oceanography, 44(6):1359–1369, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Figueiras:1999:IBO

[FAE99] Francisco G. Figueiras, Belén Arbones, and Marta Estrada. Implications of bio-optical modeling of phytoplankton photosynthesis in Antarctic waters: Further evidence of no light limitation in the Bransfield Strait. Limnology and Oceanography, 44(7):1599–1608, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Frankel:1990:HSF

[FBCS90] Sheila L. Frankel, Brian J. Binder, Sallie W. Chisholm, and Howard M. Shapiro. High-sensitivity flow cytometer for studying picoplankton. Limnology and Oceanography, 35(5):1164–1169, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Farrell:1991:CST

[FBR91] Terence M. Farrell, David Bracher, and Jonathan Roughgarden. Cross-shelf transport causes recruitment to intertidal populations in central California. Limnology and Oceanography, 36(2):279– 288, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Fasham:1999:MRC

[FBS99] Michael J. R. Fasham, Philip W. Boyd, and Graham Savidge. Modeling the relative contributions of autotrophs and heterotrophs to carbon flow at a Lagrangian JGOFS station in the Northeast Atlantic: The importance of DOC. Limnology and Oceanography, 44(1):80–94, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Frankignoulle:1996:ACF

[FBW96] Michel Frankignoulle, Isabelle Bourge, and Roland Wollast. Atmospheric CO_{2} fluxes in a highly polluted estuary (the Scheldt). Limnology and Oceanography, 41(2):365–369, March 1996. CO-DEN LIOCAH. ISSN 0024-3590.

Fuhrman:1991:PBC

[FC91] Jed A. Fuhrman and Douglas G. Capone. Possible biogeochemical consequences of ocean fertilization. Limnology and Oceanography, 36(8):1951–1959, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Felip:1999:TCM

[FCC99] Marisol Felip, Lluís Camarero, and Jordi Catalan. Temporal changes of microbial assemblages in the ice and snow cover of a high mountain lake. Limnology and Oceanography, 44(4):973–987, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Frankignoulle:1994:MCS

[FCG94] Michel Frankignoulle, Christine Canon, and Jean-Pierre Gattuso. Marine calcification as a source of carbon dioxide: Positive feedback of increasing atmospheric CO_{2}. Limnology and Oceanography, 39(2):458–462, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Fahnenstiel:1991:PCF

[FCI91] Gary L. Fahnenstiel, Hunter J. Carrick, and Rodolfo Iturriaga. Physiological characteristics and food-web dynamics of Synechococcus in Lakes Huron and Michigan. Limnology and Oceanography, 36(2):219–234, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Falkner:1997:MTE

[FCM^{+}97] K. Kenison Falkner, M. Church, C. I. Measures, G. Lebaron, D. Thouron, C. Jeandel, M. C. Stordal, G. A. Gill, R. Mortlock, P. Froelich, and L.-H. Chan. Minor and trace element chemistry of Lake Baikal, its tributaries, and surrounding hot springs. Limnology and Oceanography, 42(2):329–345, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Folt:1999:SAA

[FCMB99] C. L. Folt, C. Y. Chen, M. V. Moore, and J. Burnaford. Synergism and antagonism among multiple stressors. Limnology and Oceanography, 44(3 (part 2)):864–877, May 1999. CODEN LI-OCAH. ISSN 0024-3590.

Forrester:1999:TIO

[FDG99] Graham E. Forrester, Tom L. Dudley, and Nancy B. Grimm. Trophic interactions in open systems: Effects of predators and nutrients on stream food chains. Limnology and Oceanography, 44(5):1187–1197, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Frenette:1993:LAA

[FDLD93] Jean-Jacques Frenette, Serge Demers, Louis Legendre, and Julian Dodson. Lack of agreement among models for estimating the photosynthetic parameters. Limnology and Oceanography, 38(3): 679–687, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Fee:1999:EF

[Fee99a] Everett Fee. Encyclopedia of fishes. Limnology and Oceanography, 44(3):737, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Fee:1999:RS

[Fee99b] Everett Fee. The restless sea. Limnology and Oceanography, 44 (6):1594, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Fenchel:1994:BRA

[Fen94] Tom Fenchel. Book review: Aquatic microbiology. An ecological approach (T. E. Ford [ed.]). Limnology and Oceanography, 39(2): 475, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Foreman:1998:TDR

[FFF98] C. M. Foreman, P. Franchini, and R. L. Foreman. The trophic dynamics of riverine bacterioplankton: Relationships among substrate availability, ectoenzyme kinetics, and growth. Limnology and Oceanography, 43(6):1344–1352, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Fritz:1992:IBM

[FFJ92] Lowell W. Fritz, Gregory Ferrence, and Timothy R. Jacobsen. Induction of barite mineralization in the Asiatic clam Corbicula fluminea. Limnology and Oceanography, 37(2):442–448, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Ferdelman:1999:SRS

[FFNS99] Timothy G. Ferdelman, Henrik Fossing, Kirsten Neumann, and Horst D. Schulz. Sulfate reduction in surface sediments of the southeast Atlantic continental margin between 15^{o}38’S and 27^{o}57’S (Angola and Namibia). Limnology and Oceanography, 44 (3):650–661, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Fitzgerald:1993:ACB

[FG93] Sharon A. Fitzgerald and Wayne S. Gardner. An algal carbon budget for pelagic-benthic coupling in Lake Michigan. Limnology and Oceanography, 38(3):547–560, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Fabricius:1995:FDH

[FGB95] Katharina E. Fabricius, Amatzia Genin, and Yehuda Benayahu. Flow-dependent herbivory and growth in zooxanthellae-free soft corals. Limnology and Oceanography, 40(7):1290–1301, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Flegal:1993:LID

[FGN93] A. Russell Flegal, David L. Garrison, and Sidney Niemeyer. Lead isotopic disequilibria between plankton assemblages and surface waters reflect life cycle strategies of coastal populations within a northeast Pacific upwelling regime. Limnology and Oceanography, 38(3):670–678, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Forward:1992:EFP

[FH92] Richard B. Forward, Jr. and William F. Hettler, Jr. Effects of feeding and predator exposure on photoresponses during diel vertical migration of brine shrimp larvae. Limnology and Oceanography, 37(6):1261–1270, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Finelli:1999:ESR

[FHF99] Christopher M. Finelli, David D. Hart, and Dina M. Fonseca. Evaluating the spatial resolution of an acoustic Doppler velocimeter and the consequences for measuring near-bed flows. Limnology and Oceanography, 44(7):1793–1801, November 1999. CODEN LI-OCAH. ISSN 0024-3590.

Fee:1996:ELS

[FHKC96] E. J. Fee, R. E. Hecky, S. E. M. Kasian, and D. R. Cruikshank. Effects of lake size, water clarity, and climatic variability on mixing depths in Canadian Shield lakes. Limnology and Oceanography, 41(5):912–920, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Fry:1996:LTD

[FHN^{+}96] Brian Fry, Charles S. Hopkinson, Jr., Amy Nolin, Bosse Norrman, and Ulla Li Zweifel. Long-term decomposition of DOC from experimental diatom blooms. Limnology and Oceanography, 41(6): 1344–1347, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Findlay:1998:MSE

[FHS98] Stuart Findlay, Robert O. Hall, Jr., and William V. Sobczak. Methods on stream ecology. Limnology and Oceanography, 43(5): 1020–1021, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Fiebig:1997:MTA

[Fie97] Douglas M. Fiebig. Microbiological turnover of amino acids immobilized from groundwater discharged through hyporheic sediments. Limnology and Oceanography, 42(4):763–768, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Filion:1991:IZS

[Fil91] Jean-Marc Filion. Improving zooplankton samplers with simple electronics. Limnology and Oceanography, 36(1):204–210, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Findlay:1995:ISS

[Fin95] Stuart Findlay. Importance of surface-subsurface exchange in stream ecosystems: The hyporheic zone. Limnology and Oceanography, 40(1):159–164, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Flameling:1998:LDQ

[FK98] Inez A. Flameling and Jacco Kromkamp. Light dependence of quantum yields for PSII charge separation and oxygen evolution in eucaryotic algae. Limnology and Oceanography, 43(2):284–297, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Forbes:1990:EFC

[FL90] Thomas L. Forbes and Glenn R. Lopez. The effect of food concentration, body size, and environmental oxygen tension on the growth of the deposit-feeding polycheate, Capitella species. Limnology and Oceanography, 35(7):1535–1544, November 1990. CO-DEN LIOCAH. ISSN 0024-3590.

Freeman:1995:BPM

[FL95] Christopher Freeman and Maurice A. Lock. The biofilm polysaccharide matrix: a buffer against changing organic substrate supply? Limnology and Oceanography, 40(2):273–278, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Fechner-Levy:1996:TMV

[FLH96] Elizabeth J. Fechner-Levy and Harold F. Hemond. Trapped methane volume and potential effects on methane ebullition in a northern peatland. Limnology and Oceanography, 41(7):1375– 1383, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Francisco:1994:CCE

[FM94] P. W. Francisco and N. J. McCormick. Chlorophyll concentration effects on asymptotic optical attenuation. Limnology and Oceanography, 39(5):1195–1205, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Falkner:1991:MME

[FMH^{+}91] Kelly Kenison Falkner, Chris I. Measures, Sarah E. Herbelin, John M. Edmond, and Ray F. Weiss. The major and minor element geochemistry of Lake Baikal. Limnology and Oceanography, 36(3):413–423, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Frost:1999:MSS

[FMK^{+}99] Thomas M. Frost, Pamela K. Montz, Timothy K. Kratz, Tracy Badillo, Patrick L. Brezonik, Maria J. Gonzalez, Ronald G. Rada, Carl J. Watras, Katherine E. Webster, James G. Wiener, Craig E. Williamson, and Donald P. Morris. Multiple stresses from a single agent: Diverse responses to the experimental acidification of Little Rock Lake, Wisconsin. Limnology and Oceanography, 44(3 (part 2)):784–794, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Fuhrman:1995:VPC

[FN95] Jed A. Fuhrman and Rachel T. Noble. Viruses and protists cause similar bacterial mortality in coastal seawater. Limnology and Oceanography, 40(7):1236–1242, November 1995. CODEN LI-OCAH. ISSN 0024-3590.

Fischer:1996:STS

[FNWK96] Gerhard Fischer, Susanne Neuer, Gerold Wefer, and Gunther Krause. Short-term sedimentation pulses recorded with a fluorescence sensor and sediment traps at 9-m depth in the Canary basin. Limnology and Oceanography, 41(6):1354–1359, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Fritsen:1999:SCO

[FP99] Christian H. Fritsen and John C. Priscu. Seasonal change in the optical properties of the permanent ice cover on Lake Bonney, Antarctica: Consequences for lake productivity and phytoplankton dynamics. Limnology and Oceanography, 44(2):447–454, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Ferrier-Pages:1998:MZC

[FPAG^{+}98] C. Ferrier-Pagès, D. Allemand, J.-P. Gattuso, J. Jaubert, and F. Rassoulzadegan. Microheterotrophy in the zooxanthellate coral Stylophora pistillata: Effects of light and ciliate density. Limnology and Oceanography, 43(7):1639–1648, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Fiedler:1991:OUP

[FPC91] Paul C. Fiedler, Valerie Philbrick, and Francisco P. Chavez. Oceanic upwelling and productivity in the eastern tropical Pacific. Limnology and Oceanography, 36(8):1834–1850, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Finlay:1999:EWV

[FPC99] Jacques C. Finlay, Mary E. Power, and Gilbert Cabana. Effects of water velocity on algal carbon isotope ratios: Implications for river food web studies. Limnology and Oceanography, 44(5):1198– 1203, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Findlay:1991:WCB

[FPL^{+}91] Stuart Findlay, Michael L. Pace, David Lints, Jonathan J. Cole, Nina F. Caraco, and Benjamin Peierls. Weak coupling of bacterial and algal production in a heterotrophic ecosystem: The Hudson River estuary. Limnology and Oceanography, 36(2):268–278, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Ferrier-Pages:1994:RPP

[FPR94] C. Ferrier-Pages and F. Rassoulzadegan. N remineralization in planktonic protozoa. Limnology and Oceanography, 39(2):411– 419, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Finelli:1999:OTT

[FPZFW99] Christopher M. Finelli, N. Dean Pentcheff, Richard K. Zimmer-Faust, and David S. Wethey. Odor transport in turbulent flows: Constraints on animal navigation. Limnology and Oceanography, 46(4):1056–1071, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Ferrier:1991:DDE

[FR91] Christine Ferrier and Fereidoun Rassoulzadegan. Densitydependent effects of protozoans on specific growth rates in picoand nanoplanktonic assemblages. Limnology and Oceanography, 36(4):657–669, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Forward:1999:BSL

[FR99] Richard B. Forward, Jr. and Dan Rittschof. Brine shrimp larval photoresponses involved in diel vertical migration: Activation by fish mucus and modified amino sugars. Limnology and Oceanography, 44(8):1904–1916, December 1999. CODEN LIOCAH. ISSN 0024-3590.

France:1995:DBL

[Fra95] Robert L. France. Differentiation between littoral and pelagic food webs in lakes using stable carbon isotopes. Limnology and Oceanography, 40(7):1310–1313, November 1995. CODEN LI-OCAH. ISSN 0024-3590.

Franks:1997:MHA

[Fra97a] Peter J. S. Franks. Models of harmful algal blooms. Limnology and Oceanography, 42(5 (part 2)):1273–1282, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Franks:1997:SPD

[Fra97b] Peter J. S. Franks. Spatial patterns in dense algal blooms. Limnology and Oceanography, 42(5 (part 2)):1297–1305, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Fritz:1990:TCS

[Fri90] Sherilyn C. Fritz. Twentieth-century salinity and water-level fluctuations in Devils Lake, North Dakota: Test of a diatom-based transfer function. Limnology and Oceanography, 35(8):1771–1781, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Fritz:1996:PRC

[Fri96] Sherilyn C. Fritz. Paleolimnological records of climatic change in North America. Limnology and Oceanography, 41(5):882–889, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Fahnenstiel:1994:IPP

[FRL94] Gary L. Fahnenstiel, Donald G. Redalje, and Steven E. Lohrenz. Has the importance of photoautotrophic picoplankton been overestimated? Limnology and Oceanography, 39(2):432–438, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Fritz:1990:BBS

[FRLS90] Lowell W. Fritz, Lisa M. Ragone, Richard A. Lutz, and Susan Swapp. Biomineralization of barite in the shell of the freshwater Asiatic clam Corbicula fluminea (Molluscs: Bivalvia). Limnology and Oceanography, 35(3):756–762, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Frost:1991:RGN

[Fro91] Bruce W. Frost. The role of grazing in nutrient-rich areas of the open sea. Limnology and Oceanography, 36(8):1616–1630, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Fitt:1995:RBP

[FRY95] W. K. Fitt, T. A. V. Rees, and D. Yellowlees. Relationship between pH and the availability of dissolved inorganic nitrogen in the zooxanthella-giant clam symbiosis. Limnology and Oceanography, 41(4):976–982, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Floder:1999:DPC

[FS99] Sabine Flöder and Ulrich Sommer. Diversity in planktonic communities: an experimental test of the intermediate disturbance hypothesis. Limnology and Oceanography, 46(4):1114–1119, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Findlay:1993:MSD

[FSGG93] Stuart Findlay, David Strayer, Cheikh Goumbala, and Kim Gould. Metabolism of streamwater dissolved organic carbon in the shallow hyporheic zone. Limnology and Oceanography, 38(7): 1493–1499, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Fenaux:1994:FTF

[FSS94] Lucienne Fenaux, Megumi F. Strathmann, and Richard A. Strathmann. Five tests of food-limited growth of larvae in coastal waters by comparisons of rates of development and form of echinoplutei. Limnology and Oceanography, 39(1):84–98, January 1994. CO-DEN LIOCAH. ISSN 0024-3590.

Fujimoto:1997:NLG

[FSSI97] Naoshi Fujimoto, Ryuichi Sudo, Norio Sugiura, and Yuhei Inamori. Nutrient-limited growth of Microcystis aeruginosa and Phormidium tenue and competition under various N : P supply ratios and temperatures. Limnology and Oceanography, 42(2): 250–256, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Ferrari:1991:ADL

[FT91] Giovanni M. Ferrari and Stelvio Tassan. On the accuracy of determining light absorption by “yellow substance” through measurements of induced fluorescence. Limnology and Oceanography, 36(4):777–786, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Furnas:1991:NSG

[Fur91] Miles J. Furnas. Net in situ growth rates of phytoplankton in an oligotrophic, tropical shelf ecosystem. Limnology and Oceanography, 36(1):13–29, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Frenette:1998:SDC

[FVL98] Jean-Jacques Frenette, Warwick F. Vincent, and Louis Legendre. Size-dependent C : N uptake by phytoplankton as a function of irradiance: Ecological implications. Limnology and Oceanography, 43(6):1362–1368, September 1998. CODEN LIOCAH. ISSN 0024-

Frenette:1999:SDC

[FVL99] J-J. Frenette, W. F. Vincent, and L. Legendre. Size-dependent C : N uptake by phytoplankton as a function of irradiance: Ecological implications. Limnology and Oceanography, 44(1):231, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Fabricius:1998:SDP

[FYG98] Katharina Fabricius, Gitai Yahel, and Amatzia Genin. In situ depletion of phytoplankton by an azooxanthellate soft coral. Limnology and Oceanography, 43(2):354–356, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Fang:1993:NVP

[FZD93] Peggy Fang, Joy B. Zedler, and Regina M. Donohoe. Nitrogen vs. phosphorus limitation of algal biomass in shallow coastal lagoons. Limnology and Oceanography, 38(5):906–923, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Fourqurean:1992:PLP

[FZP92] James W. Fourqurean, Joseph C. Zieman, and George V. N. Powell. Phosphorus limitation of primary production in Florida Bay: Evidence from C : N : P ratios of the dominant seagrass Thalassia testudinum. Limnology and Oceanography, 37(1):162–171, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Gray:1996:TNT

[GA96] John S. Gray and Mohammed I. Abdullah. Are there negative trends in oxygen saturation along the Norwegian Skagerrak coast? Limnology and Oceanography, 40(5):810–812, June 1996. CODEN LIOCAH. ISSN 0024-3590. See reply [JD96b].

Green:1993:CDT

[GAA93] Mark A. Green, Robert C. Aller, and Josephine Y. Aller. Carbonate dissolution and temporal abundances of Foraminifera in Long Island Sound sediments. Limnology and Oceanography, 38 (2):331–345, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Green:1998:ICD

[GAA98] Mark A. Green, Robert C. Aller, and Josephine Y. Aller. Influence of carbonate dissolution on survival of shell-bearing meiobenthos in nearshore sediments. Limnology and Oceanography, 43(1):18– 28, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Graham:1996:EPS

[GAAG96a] James M. Graham, Patricia Arancibia-Avila, and Linda E. Graham. Effects of pH and selected metals on growth of the filamentous green alga Mougeotia under acidic conditions. Limnology and Oceanography, 41(2):263–270, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Graham:1996:PES

[GAAG96b] James M. Graham, Patricia Arancibia-Avila, and Linda E. Graham. Physiological ecology of a species of the filamentous green alga Mougeotia under acidic conditions: Light and temperature effects on photosynthesis and respiration. Limnology and Oceanography, 41(2):253–262, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Gaedke:1992:SDP

[Gae92] Ursula Gaedke. The size distribution of plankton biomass in a large lake and its seasonal variability. Limnology and Oceanography, 37(6):1202–1220, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Gaedke:1993:EAB

[Gae93] Ursula Gaedke. Ecosystem analysis based on biomass size distributions: a case study of a plankton community in a large lake. Limnology and Oceanography, 38(1):112–127, January 1993. CO-DEN LIOCAH. ISSN 0024-3590.

Gargett:1991:PPM

[Gar91a] Ann E. Gargett. Physical processes and the maintenance of nutrient-rich euphotic zones. Limnology and Oceanography, 36(8): 1527–1545, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Garside:1991:SNK

[Gar91b] Chris Garside. Shift-up and the nitrate kinetics of phytoplankton in upwelling systems. Limnology and Oceanography, 36(6):1239– 1244, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Gaiser:1993:ETE

[GB93] Evelyn E. Gaiser and Roger W. Bachmann. The ecology and taxonomy of epizoic diatoms on cladocera. Limnology and Oceanography, 38(3):628–637, May 1993. CODEN LIOCAH. ISSN 0024-

Green:1994:OAF

[GB94] Sarah A. Green and Neil V. Blough. Optical absorption and fluorescence properties of chromophoric dissolved organic matter in natural waters. Limnology and Oceanography, 39(8):1903–1916, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Gamier:1995:SSD

[GBC95] Josette Gamier, Gilles Billen, and Michel Coste. Seasonal succession of diatoms and Chlorophyceae in the drainage network of the Seine River: Observation and modeling. Limnology and Oceanography, 40(4):750–765, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Gregg:1990:SSS

[GC90] Watson W. Gregg and K. L. Carder. A simple spectral solar irradiance model for cloudless maritime atmospheres. Limnology and Oceanography, 35(8):1657–1675, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Gessner:1997:GPA

[GC97] Mark O. Gessner and Eric Chauvet. Growth and production of aquatic hyphomycetes in decomposing leaf litter. Limnology and Oceanography, 42(3):496–505, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Gardner:1998:NCR

[GCB^{+}98] Wayne S. Gardner, Joann F. Cavaletto, Harvey A. Bootsma, Peter J. Lavrentyev, and Federico Troncone. Nitrogen cycling rates and light effects in tropical Lake Maracaibo, Venezuela. Limnology and Oceanography, 43(8):1814–1825, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Gardner:1997:ENL

[GCCJ97] Wayne S. Gardner, Joann F. Cavaletto, James B. Cotner, and Jefsrey R. Johnson. Effects of natural light on nitrogen cycling rates in the Mississippi River plume. Limnology and Oceanography, 42 (2):273–281, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Grebmeier:1990:OIC

[GCD90] Jacqueline M. Grebmeier, Lee W. Cooper, and Michael J. DeNiro. Oxygen isotopic composition of bottom seawater and tunicate cellulose used as indicators of water masses in the northern Bering and Chukchi Seas. Limnology and Oceanography, 35(5):1182– 1195, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Gordon:1997:IDE

[GCJ97] R. Michael Gordon, Kenneth H. Coale, and Kenneth S. Johnson. Iron distributions in the equatorial Pacific: Implications for new production. Limnology and Oceanography, 42(3):419–431, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Gallegos:1990:MSD

[GCP90] Charles L. Gallegos, David L. Correll, and J. W. Pierce. Modeling spectral diffuse attenuation, absorption, and scattering coefficients in a turbid estuary. Limnology and Oceanography, 35(7): 1486–1502, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Gordon:1992:SSW

[GD92] Howard R. Gordon and Kuiyuan Ding. Self-shading of in-water optical instruments. Limnology and Oceanography, 37(3):491–500, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Gorski:1996:FSD

[GD96] Patrick R. Gorski and Stanley I. Dodson. Free-swimming Daphnia pulex can avoid following Stokes’ law. Limnology and Oceanography, 41(4):1815–1821, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Gensemer:1999:UCF

[GDG99] Robert W. Gensemer, D. George Dixon, and Bruce M. Greenberg. Using chlorophyll-a fluorescence to detect the onset of anthracene photoinduced toxicity in Lemna gibba, and the mitigating effects of a commercial humic acid. Limnology and Oceanography, 44(3 (part 2)):878–888, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Gasol:1997:BDM

[GdGD97] Josep M. Gasol, Paul A. del Giorgio, and Carlos M. Duarte. Biomass distribution in marine planktonic communities. Limnology and Oceanography, 42(6):1353–1363, September 1997. CO-DEN LIOCAH. ISSN 0024-3590.

Geider:1990:RBS

[Gei90] Richard J. Geider. The relationship between steady state phytoplankton growth and photosynthesis. Limnology and Oceanography, 35(4):971–972, June 1990. CODEN LIOCAH. ISSN 0024-

Gensemer:1991:EPA

[Gen91] Robert W. Gensemer. The effects of pH and aluminum on the growth of the acidophilic diatom Asterionella ralfsii var. americana. Limnology and Oceanography, 36(1):123–131, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Gould:1990:FMS

[GG90] Diane M. Gould and Eugene D. Gallagher. Field measurement of specific growth rate, biomass, and primary production of benthic diatoms of savin Hill cove, Boston. Limnology and Oceanography, 35(8):1757–1770, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Garcia:1992:OSS

[GG92] Herncin E. Garcia and Louis I. Gordon. Oxygen solubility in seawater: Better fitting equations. Limnology and Oceanography, 37(6):1307–1312, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Gustafsson:1997:ACC

[GG97] Cirjan Gustafsson and Philip M. Gschwend. Aquatic colloids: Concepts, definitions, and current challenges. Limnology and Oceanography, 42(3):519–528, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Gries:1999:ENM

[GG99] Thomas Gries and Hans Güde. Estimates of the nonconsumptive mortality of mesozooplankton by measurement of sedimentation losses. Limnology and Oceanography, 44(2):459–465, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Gasol:1991:SVS

[GGA91] Josep M. Gasol, Ricardo Guerrero, and Carlos Pedró Alió. Seasonal variations in size structure and procaryotic dominance in sulfurous Lake Cisó. Limnology and Oceanography, 36(5):860–872, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Greene:1991:EIL

[GGF91] Richard M. Greene, Richard J. Geider, and Paul G. Falkowski. Effect of iron limitation on photosynthesis in a marine diatom. Limnology and Oceanography, 36(8):1772–1782, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Glibert:1991:DCI

[GGFR91] Patricia M. Glibert, Chris Garside, Jed A. Fuhrman, and Michael R. Roman. Dependent coupling of inorganic and organic nitrogen uptake and regeneration in the plume of the Chesapeake Bay estuary and its regulation by large heterotrophs. Limnology and Oceanography, 36(5):895–909, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Gullberg:1997:FDC

[GGJ97] Karl R. Gullberg, Willem Goedkoop, and Richard K. Johnson. The fate of diatom carbon within a freshwater benthic community — a microcosm study. Limnology and Oceanography, 42(3):452– 460, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Glud:1994:EBD

[GGRJ94] Ronnie Nøhr Glud, Jens Kristian Gundersen, Niels Peter Revsbech, and Bo Barker Jørgensen. Effects on the benthic diffusive boundary layer imposed by microelectrodes. Limnology and Oceanography, 39(2):462–467, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Guisande:1995:ETO

[GH95] Castor Guisande and Roger Harris. Effect of total organic content of eggs on hatching success and naupliar survival in the copepod Calanus helgolandicus. Limnology and Oceanography, 40(3):476– 482, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Gobler:1997:RBC

[GHF^{+}97] Christopher J. Gobler, David A. Hutchins, Nicholas S. Fisher, Elizabeth M. Cosper, and Sergio A. Sañudo-Wilhelmy. Release and bioavailability of C, N, P, Se, and Fe following viral lysis of a marine chrysophyte. Limnology and Oceanography, 42(7):1492– 1504, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Goes:1996:MNM

[GHT^{+}96] Joaquim I. Goes, Nobuhiko Handa, Satoru Taguchi, Takeo Hama, and Hiroaki Saito. Metabolism of neutral monosaccharide constituents of storage and structural carbohydrates in natural assemblages of marine phytoplankton exposed to ultraviolet radiation. Limnology and Oceanography, 41(7):1478–1489, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Gibbs:1992:RCE

[Gib92] Ronald J. Gibbs. A reply to the comment of Eilers et al. Limnology and Oceanography, 37(6):1338–1339, September 1992. CODEN LIOCAH. ISSN 0024-3590. See [EBH92].

Gilbert:1994:SPR

[Gil94] John J. Gilbert. Susceptibility of planktonic rotifers to a toxic strain of Anabaena flos-aquae. Limnology and Oceanography, 39 (6):1286–1297, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:1996:EFA

[Gil96] John J. Gilbert. Effect of food availability on the response of planktonic rotifers to a toxic strain of the cyanobacterium Anabaena flos-aquae. Limnology and Oceanography, 41(7):1565– 1572, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Gattuso:1990:ELO

[GJ90] J.-P. Gattuso and J. Jaubert. Effect of light on oxygen and carbon dioxide fluxes and on metabolic quotients measured in situ in a zooxanthellate coral. Limnology and Oceanography, 35(8):1796– 1804, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Goedkoop:1996:PBC

[GJ96] Willem Goedkoop and Richard K. Johnson. Pelagic-benthic coupling: Profundal benthic community response to spring diatom deposition in mesotrophic Lake Erken. Limnology and Oceanography, 41(4):636–647, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Gallegos:1992:ESR

[GJC92] C. L. Gallegos, T. E. Jordan, and D. L. Correll. Event-scale response of phytoplankton to watershed inputs in a subestuary: Timing, magnitude, and location of blooms. Limnology and Oceanography, 37(4):813–828, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Gremm:1997:DCS

[GK97] Thomas J. Gremm and Louis A. Kaplan. Dissolved carbohydrates in streamwater determined by HPLC and pulsed amperometric detection. Limnology and Oceanography, 42(2):385–393, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Genin:1994:EFC

[GKM94] Amatzia Genin, Lee Karp, and Aharon Miroz. Effects of flow on competitive superiority in scleractinian corals. Limnology and Oceanography, 39(4):913–924, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Greene:1994:PLP

[GKS^{+}94] Richard M. Greene, Zbigniew S. Kolber, Dorothy G. Swift, Neil W. Tindale, and Paul G. Falkowski. Physiological limitation of phytoplankton photosynthesis in the eastern equatorial Pacific determined from variability in the quantum yield of fluorescence. Limnology and Oceanography, 39(5):1061–1074, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Gliwicz:1994:CSV

[GL94] Z. Maciej Gliwicz and Winfried Lampert. Clutch-size variability in Daphnia: Body-size related effects of egg predation by cyclopoid copepods. Limnology and Oceanography, 39(3):479–485, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Glazier:1991:SRR

[Gla91] Douglas S. Glazier. Separating the respiration rates of embryos and brooding females of Daphnia magna: Implications for the cost of brooding and the allometry of metabolic rate. Limnology and Oceanography, 36(2):354–361, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Gleason:1993:DEU

[Gle93] Daniel F. Gleason. Differential effects of ultraviolet radiation on green and brown morphs of the Caribbean coral Porites astreoides. Limnology and Oceanography, 38(7):1452–1463, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Giblin:1991:RCS

[GLH91] Anne E. Giblin, Gene E. Likens, and Robert W. Howarth. A reply to the comment by Stauffer. Limnology and Oceanography, 36(6): 1265–1271, September 1991. CODEN LIOCAH. ISSN 0024-3590. See [Sta91].

Glinsky:1998:CMM

[Gli98] Alexander Glinsky. Current meters for measurement of low-speed velocities in ice-covered lakes. Limnology and Oceanography, 43 (7):1661–1668, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Guilizzoni:1992:PPR

[GLM92] Piero Guilizzoni, Andrea Lami, and Aldo Marchetto. Plant pigment ratios from lakes sediments as indicators of recent acidification in alpine lakes. Limnology and Oceanography, 37(7):1565– 1569, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Gilmour:1990:EAI

[GLS90] C. C. Gilmour, M. E. Leavitt, and M. P. Shiaris. Evidence against incorporation of exogenous thymidine by sulfate-reducing bacteria. Limnology and Oceanography, 35(6):1401–1409, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Graneli:1996:POP

[GLT96] Wilhelm Granéli, Måns Lindell, and Lars Tranvik. Photooxidative production of dissolved inorganic carbon in lakes of different humic content. Limnology and Oceanography, 41(4):698– 706, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Giblin:1990:SSA

[GLWH90] Anne E. Giblin, Gene E. Likens, David White, and Robert W. Howarth. Sulfur storage and alkalinity generation in New England lake sediments. Limnology and Oceanography, 35(4):852–869, June 1990. CODEN LIOCAH. ISSN 0024-3590. See comment [Sta91].

Geider:1996:DMP

[GMK96] Richard J. Geider, Hugh L. MacIntyre, and Todd M. Kana. A dynamic model of photoadaptation in phytoplankton. Limnology and Oceanography, 41(1):1–15, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Geider:1998:DRM

[GMK98] Richard J. Geider, Hugh L. Maclntyre, and Todd M. Kana. A dynamic regulatory model of phytoplanktonic acclimation to light, nutrients, and temperature. Limnology and Oceanography, 43(4): 679–694, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Garcia-Mendoza:1996:RSS

[GMM96] Ernesto Garcia-Mendoza and Helmut Maske. The relationship of solar-stimulated natural fluorescence and primary productivity in Mexican Pacific waters. Limnology and Oceanography, 41(8): 1697–1710, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Guisande:1999:HEA

[GMR99] Cástor Guisande, Isabel Maneiro, and Isabel Riveiro. Homeostasis in the essential amino acid composition of the marine copepod Euterpina acutifrons. Limnology and Oceanography, 44(3):691– 696, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Goodwin:1997:PBD

[GNL97] Kelly D. Goodwin, Wheeler J. North, and Mary E. Lidstrom. Production of bromoform and dibromomethane by Giant Kelp: Factors affecting release and comparison to anthropogenic bromine sources. Limnology and Oceanography, 42(8):1725–1734, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Goericke:1998:RPC

[Goe98] Ralf Goericke. Response of phytoplankton community structure and taxon-specific growth rates to seasonally varying physical forcing in the Sargasso Sea off Bermuda. Limnology and Oceanography, 43(5):921–935, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Gordon:1991:ASE

[Gor91] Howard R. Gordon. Absorption and scattering estimates from irradiance measurements: Monte Carlo simulations. Limnology and Oceanography, 36(4):769–777, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Garcia-Pichel:1994:MIS

[GP94] Ferran Garcia-Pichel. A model for internal self-shading in planktonic organisms and its implications for the usefulness of ultraviolet sunscreens. Limnology and Oceanography, 39(7):1704–1717, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Granger:1999:ISI

[GP99] Julie Granger and Neil M. Price. The importance of siderophores in iron nutrition of heterotrophic marine bacteria. Limnology and Oceanography, 44(3):541–555, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Gagnon:1996:DBM

[GPMF96] Christian Gagnon, Emilien Pelletier, Alfonso Mucci, and William F. Fitzgerald. Diagenetic behavior of methylmercury in organic-rich coastal sediments. Limnology and Oceanography, 41(3):428–434, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Granata:1991:DPG

[Gra91] Timothy C. Granata. Diel periodicity in growth and sinking rates of the centric diatom Coscinodiscus concinnus. Limnology and Oceanography, 36(1):132–139, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Granina:1997:CBL

[Gra97] Liba Granina. The chemical budget of Lake Baikal: a review. Limnology and Oceanography, 42(2):373–379, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Gehlen:1997:DCD

[GREGG97] Marion Gehlen, Christophe Rabouille, Ullah Ezat, and Laurence D. Guidi-Guilvard. Drastic changes in deep-sea sediment porewater composition induced by episodic input of organic matter. Limnology and Oceanography, 42(5):980–986, July 1997. CO-DEN LIOCAH. ISSN 0024-3590.

Gundersen:1998:PSM

[GRG98] Jens Kristian Gundersen, Niels Birger Ramsing, and Ronnie Nøhr Glud. Predicting the signal of O_{2} microsensors from physical dimensions, temperature, salinity, and O_{2} concentration. Limnology and Oceanography, 43(8):1932–1937, December 1998. CODEN LI-OCAH. ISSN 0024-3590.

Gal:1999:ACM

[GRG99] Gideon Gal, Lars G. Rudstam, and Charles H. Greene. Acoustic characterization of Mysis relicta. Limnology and Oceanography, 44(2):371–381, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Grunbaum:1997:HMC

[Grü97] Daniel Grünbaum. Hydromechanical mechanisms of colony organization and cost of defense in an encrusting bryozoan, Membranipora membranacea. Limnology and Oceanography, 42(4):741–752, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Gat:1991:GES

[GS91] Joel R. Gat and Michal Shatkay. Gas exchange with saline waters. Limnology and Oceanography, 36(5):988–997, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Grossart:1993:LMO

[GS93] Hans-Peter Grossart and Meinhard Simon. Limnetic macroscopic organic aggregates (lake snow): Occurrence, characteristics, and microbial dynamics in Lake Constance. Limnology and Oceanography, 38(3):532–546, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Guo:1996:CHM

[GSC^{+}96] Laodong Guo, Peter H. Santschi, Luis A. Cifuentes, Susan E. Trumbore, and John Southon. Cycling of high-molecular-weight dissolved organic matter in the Middle Atlantic Bight as revealed by carbon isotopic (^{13}C and ^{14}C) signatures. Limnology and Oceanography, 41(6):1242–1252, September 1996. CODEN LI-OCAH. ISSN 0024-3590.

Garver:1994:VNS

[GSG94] Sara A. Garver, David A. Siegel, and Mitchell B. Greg. Variability in near-surface particulate absorption spectra: What can a satellite ocean color imager see? Limnology and Oceanography, 39(6): 1349–1367, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Grossart:1997:FMO

[GSL97] Hans-Peter Grossart, Meinhard Simon, and Bruce E. Logan. Formation of macroscopic organic aggregates (lake snow) in a large lake: The significance of transparent exopolymer particles, plankton, and zooplankton. Limnology and Oceanography, 42(8):1651– 1659, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Guo:1995:DDO

[GSW95] Laodong Guo, Peter H. Santschi, and Kent W. Warnken. Dynamics of dissolved organic carbon (DOC) in oceanic environments. Limnology and Oceanography, 40(8):1392–1403, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Gliwicz:1994:CBS

[GU94] Z. Maciej Gliwicz and Gerardo Umana. Cladoceran body size and vulnerability to copepod predation. Limnology and Oceanography, 39(2):419–424, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Guyonnet:1991:NME

[Guy91] Dominique A. Guyonnet. Numerical modeling of effects of smallscale sedimentary variations on groundwater discharge into lakes.

Limnology and Oceanography, 36(4):787–796, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Gasol:1993:LCB

[GV93] Josep M. Gasol and Dolors Vaque. Lack of coupling between heterotrophic nanoflagellates and bacteria: a general phenomenon across aquatic systems? Limnology and Oceanography, 38(3):657– 665, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Gordon:1993:ADF

[GVK93] H. R. Gordon, K. J. Voss, and K. A. Kilpatrick. Angular distribution of fluorescence from phytoplankton. Limnology and Oceanography, 38(7):1582–1586, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Gould:1990:SSD

[GW90] Richard W. Gould, Jr. and Denis A. Wiesenburg. Single-species dominance in a subsurface phytoplankton concentration at a Mediterranean Sea front. Limnology and Oceanography, 35(1): 211–219, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Goericke:1993:CLM

[GW93] Ralf Goericke and Nicholas A. Welschmeyer. The chlorophylllabeling method: Measuring specific rates of chlorophyll-a synthesis in cultures and in the open ocean. Limnology and Oceanography, 38(1):80–95, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Guyot:1994:RPR

[GW94] Jean Loup Guyot and Jean Gabriel Wasson. Regional pattern of riverine dissolved organic carbon in the Amazon drainage basin of Bolivia. Limnology and Oceanography, 39(2):452–458, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Grubaugh:1995:FSP

[GW95] Jack W. Grubaugh and J. Bruce Wallace. Functional structure and production of the benthic community in a Piedmont river: 1956–1957 and 1991–1992. Limnology and Oceanography, 40(3): 490–501, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Genovese:1999:IEF

[GW99] Salvatore J. Genovese and Jon D. Witman. Interactive effects of flow speed and particle concentration on growth rates of an active suspension feeder. Limnology and Oceanography, 46(4):1120–1131, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Greene:1991:PFS

[GWMB91] Charles H. Greene, Peter H. Wiebe, Robert T. Miyamoto, and Janusz Burczynski. Probing the fine structure of ocean soundscattering layers with ROVERSE technology. Limnology and Oceanography, 36(1):193–204, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Green:1992:MBF

[GWY^{+}92] Charles H. Green, Edith A. Widder, Marsh J. Youngbluth, Armando Tamse, and Gary E. Johnson. The migration behavior, fine structure, and bioluminescent activity of krill sound-scattering layers. Limnology and Oceanography, 37(3):650–658, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Hurley:1990:FTA

[HA90] James P. Hurley and David E. Armstrong. Fluxes and transformations of aquatic pigments in Lake Mendota, Wisconsin. Limnology and Oceanography, 35(2):384–398, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Hall:1992:RSV

[HA92] Per O. J. Hall and Robert C. Aller. Rapid, small-volume, flow injection analysis for SCO_{2}, and NH_{4}^{+} in marine and freshwaters. Limnology and Oceanography, 37(5):1113–1119, July 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Hoehler:1999:ACA

[HAAM99] Tori M. Hoehler, Daniel B. Albert, Marc J. Alperin, and Christopher S. Martens. Acetogenesis from CO_{2} in an anoxic marine sediment. Limnology and Oceanography, 44(3):662–667, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Hairston:1990:PPZb

[Hai90] Nelson G. Hairston, Jr. Problems with perception of zooplankton research by colleagues outside aquatic sciences. Limnology and Oceanography, 35(5):1214–1216, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Hairston:1996:BRH

[Hai96a] Nelson G. Hairston, Jr. Book review: How many people can the earth support? (J. E. Cohen). Limnology and Oceanography, 41 (3):578–579, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Hairston:1996:ZEB

[Hai96b] Nelson G. Hairston, Jr. Zooplankton egg banks as biotic reservoirs in changing environments. Limnology and Oceanography, 41(5): 1087–1092, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Hessen:1990:CMH

[HAL90] Dug O. Hessen, Tom Andersen, and Anne Lyche. Carbon metabolism in a humic lake: Pool sires and cycling through zooplankton. Limnology and Oceanography, 35(1):84–99, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Hambright:1994:MCP

[Ham94] K. David Hambright. Morphological constraints in the piscivoreplanktivore interaction: Implications for the trophic cascade hypothesis. Limnology and Oceanography, 39(4):897–912, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Hanazato:1991:ECR

[Han91] Takayuki Hanazato. Effects of a chaoborus-released chemical on Daphnia ambigua: Reduction in the tolerance of the Daphnia to summer water temperature. Limnology and Oceanography, 36(1): 165–171, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Hansson:1992:FRP

[Han92] Lars-Anders Hansson. Factors regulating periphytic algal biomass. Limnology and Oceanography, 37(2):322–328, March 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Hansson:1996:ARL

[Han96] Lam-Anders Hansson. Algal recruitment from lake sediments in relation to grazing, sinking, and dominance patterns in the phytoplankton community. Limnology and Oceanography, 41(6):1312– 1323, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Harrison:1995:BRM

[Har95] Paul J. Harrison. Book review: Measurement of primary production from the molecular to the global scale (W. K. W. Li and S. Y. Maestrini [eds.]). Limnology and Oceanography, 40(6):1183–1185, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Hart:1996:NCG

[Har96] Rob C. Hart. Naupliar and copepodite growth and survival of two freshwater calanoids at various food levels: Demographic contrasts, similarities, and food needs. Limnology and Oceanography, 41(4):648–658, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Hays:1995:OSV

[Hay95] Graeme C. Hays. Ontogenetic and seasonal variation in the diel vertical migration of the copepods Metridia lucens and Metridia longa. Limnology and Oceanography, 40(8):1461–1465, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Haygood:1996:PRF

[Hay96] Margo G. Haygood. The potential role of functional differences between Rubisco forms in governing expression in chemoautotrophic symbioses. Limnology and Oceanography, 41(2):370–371, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Hill:1991:CDA

[HB91] Walter R. Hill and Harry L. Boston. Community development alters photosynthesis-irradiance relations in stream periphyton. Limnology and Oceanography, 36(7):1375–1389, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Hostetler:1994:SIO

[HB94] S. W. Hostetler and L. V. Benson. Stable isotopes of oxygen and hydrogen in the Truckee River–Pyramid Lake surface-water system. 2. A predictive model of δ ^{18}O and δ ^{2}H in Pyramid Lake. Limnology and Oceanography, 39(2):356–364, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Hansen:1995:AAD

[HB95a] Lars Stenvang Hansen and Thlomas Henry Blackburn. Amino acid degradation by sulfate-reducing bacteria: Evaluation of four methods. Limnology and Oceanography, 40(3):502–510, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Hofle:1995:TDM

[HB95b] Manfred G. Höfle and Ingrid Brettar. Taxonomic diversity and metabolic activity of microbial communities in the water column of the central Baltic Sea. Limnology and Oceanography, 40(5): 868–874, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Hansen:1994:SRB

[HBH94] Benni Hansen, Peter Koefoed Bjornsen, and Per Juel Hansen. The size ratio between planktonic predators and their prey. Limnology and Oceanography, 39(2):395–403, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Hansen:1997:ZGG

[HBH97] Per Juel Hansen, Peter Koefoed Bjørnsen, and Benni Winding Hansen. Zooplankton grazing and growth: Scaling within the 2-2,-μ m body size range. Limnology and Oceanography, 42(4): 687–704, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Harvey:1995:EEP

[HBI95] Michel Harvey, Edwin Bourget, and R. Grant Ingram. Experimental evidence of passive accumulation of marine bivalve larvae on filamentous epibenthic structures. Limnology and Oceanography, 40(1):94–104, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Howurth:1993:TPN

[HBL^{+}93] Robert W. Howurth, Thomas Butler, Katharine Lunde, Dennis Swaney, and Chia Ren Chu. Turbulence and planktonic nitrogen fixation: a mesocosm experiment. Limnology and Oceanography, 38(8):1696–1711, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Hecky:1994:DDW

[HBO^{+}94] R. E. Hecky, F. W. B. Bugenyi, P. Ochumba, J. F. Talling, R. Mugidde, M. Gophen, and L. Kaufman. Deoxygenation of the deep water of Lake Victoria, East Africa. Limnology and Oceanography, 39(6):1476–1481, September 1994. CODEN LI-OCAH. ISSN 0024-3590.

Harden:1994:LMR

[HC94] Harmon S. Harden and Jefrey P. Chanton. Locus of methane release and mass-dependent fractionation from two wetland macrophytes. Limnology and Oceanography, 39(1):148–154, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Hare:1996:TML

[HC96] Jonathan A. Hare and Robert K. Cowen. Transport mechanisms of larval and pelagic juvenile bluefish (Pomatomus saltatrix) from South Atlantic Bight spawning grounds to Middle Atlantic Bight nursery habitats. Limnology and Oceanography, 41(6):1264–1280, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Holloway:1997:DSC

[HC97] Charles F. Holloway and James P. Cowen. Development of a scanning confocal laser microscopic technique to examine the structure and composition of marine snow. Limnology and Oceanography, 42(6):1340–1352, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Hughes:1999:MSC

[HC99] T. P. Hughes and J. H. Connell. Multiple stressors on coral reefs: a long-term perspective. Limnology and Oceanography, 44(3 (part 2)):932–940, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Hecky:1993:SCN

[HCH93] R. E. Hecky, P. Campbell, and L. L. Hendzel. The stoichiometry of carbon, nitrogen, and phosphorus in particulate matter of lakes and oceans. Limnology and Oceanography, 38(4):709–724, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Hare:1994:FSM

[HCHD94] Landis Hare, Richard Carignan, and Miguel A. Huerta-Diaz. A field study of metal toxicity and accumulation by benthic invertebrates; implications for the acid-volatile sulfide (AVS) model. Limnology and Oceanography, 39(7):1653–1668, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Hart:1996:FSF

[HCJ96] David D. Hart, Brian D. Clark, and Amal Jasentuliyana. Finescale field measurement of benthic flow environments inhabited by stream invertebrates. Limnology and Oceanography, 41(2):297– 308, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Haney:1990:LCE

[HCKW90] James F. Haney, Althea Craggy, Kenneth Kimball, and Frank Weeks. Light control of evening vertical migrations by Chaoborus punctipennis larvae. Limnology and Oceanography, 35(5):1068– 1078, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Houle:1995:DOC

[HCLD95] Daniel Houle, Richard Carignan, Marks Lachance, and Jacques DuPont. Dissolved organic carbon and sulfur in southwestern Quebec lakes: Relationships with catchment and lake properties. Limnology and Oceanography, 40(4):710–717, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Huzzey:1990:ECL

[HCP90] Linda M. Huzzey, James E. Cioern, and Thomas M. Powell. Episodic changes in lateral transport and phytoplankton distribution in South San Francisco Bay. Limnology and Oceanography, 35(2):472–478, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Hairston:1999:SLF

[HCPR99] Nelson G. Hairston, Jr., Francis Chan, Derek M. Pelletier, and Brian J. Roberts. Successes, limitations, and frontiers in ecosystem science. Limnology and Oceanography, 44(6):1595–1596, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Hedges:1994:OPO

[HCR^{+}94] John I. Hedges, Gregory L. Cowie, Jeffrey E. Richey, Paul D. Quay, Ronald Benner, Mike Strom, and Bruce R. Forsberg. Origins and processing of organic matter in the Amazon River as indicated by carbohydrates and amino acids. Limnology and Oceanography, 39(4):743–761, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Happell:1995:MTA

[HCS95] James D. Happell, Jeffrey P. Chanton, and William J. Showers. Methane transfer across the water–air interface in stagnant wooded swamps of Florida: Evaluation of mass-transfer coefficients and isotropic fractionation. Limnology and Oceanography, 40(2):290–298, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Hassett:1997:ESP

[HCSE97] R. P. Hassett, B. Cardinale, L. B. Stabler, and J. J. Elser. Ecological stoichiometry of N and P in pelagic ecosystems: Comparison of lakes and oceans with emphasis on the zooplanktonphytoplankton interaction. Limnology and Oceanography, 42(4): 648–662, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Hanazato:1995:SEL

[HD95] Takayuki Hanazato and Stanley I. Dodson. Synergistic effects of low oxygen concentration, predator kairomone, and a pesticide on the cladoceran Daphnia pulex. Limnology and Oceanography, 40 (4):700–709, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Henley:1997:ENS

[HD97] William J. Henley and Kenneth H. Dunton. Effects of nitrogen supply and continuous darkness on growth and photosynthesis of the Arctic kelp Laminaria solidungula. Limnology and Oceanography, 42(2):209–216, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Hutchins:1993:IRP

[HDB93] David A. Hutchins, Giacomo R. DiTullio, and Kenneth W. Burland. Iron and regenerated production: Evidence for biological iron recycling in two marine environments. Limnology and Oceanography, 38(6):1242–1255, September 1993. CODEN LI-OCAH. ISSN 0024-3590.

Hill:1997:NAU

[HDMB97] Walter R. Hill, Sarah M. Dimick, Amy E. McNamara, and Cheryl A. Branson. No effects of ambient UV radiation detected in periphyton and grazers. Limnology and Oceanography, 42(4): 769–774, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Hay:1990:SHR

[HDP^{+}90] Mark E. Hay, J. Emmett Dufy, Valerie J. Paul, Paul E. Renaud, and William Fenical. Specialist herbivores reduce their susceptibility to predation by feeding on the chemically defended seaweed Avrainvillea longicaulis. Limnology and Oceanography, 35 (8):1734–1743, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Heikoop:1998:RBL

[HDR^{+}98] Jeffrey M. Heikoop, Jennifer J. Dunn, Michael J. Risk, Ian M. Sandeman, Henry P. Schwartz, and Nigel Waltho. Relationship between light and the δ ^{15}N of coral tissue: Examples from Jamaica and Zanzibar. Limnology and Oceanography, 43(5):909–920, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Hutchins:1998:ILM

[HDZB98] D. A. Hutchins, G. R. DiTullio, Y. Zhang, and K. W. Bruland. An iron limitation mosaic in the California upwelling regime. Limnology and Oceanography, 43(6):1037–1054, September 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Hecky:1999:ECR

[Hec99] Robert E. Hecky. Environmental change and response in east African lakes. Limnology and Oceanography, 44(6):1593, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Hentschel:1998:SQN

[Hen98] Brian T. Hentschel. Spectrofluorometric quantification of neutral and polar lipids suggests a food-related recruitment bottleneck for juveniles of a deposit-feeding polychaete population. Limnology and Oceanography, 43(3):543–549, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Hessen:1993:RMN

[Hes93] Dag O. Hessen. The role of mineral nutrients for zooplankton nutrition: Reply to the comment by Brett. Limnology and Oceanography, 38(6):1340–1343, September 1993. CODEN LI-OCAH. ISSN 0024-3590. See [Bre93].

Hutchins:1999:IPI

[HFBB99] D. A. Hutchins, V. M. Franck, M. A. Brzezinski, and K. W. Bruland. Inducing phytoplankton iron limitation in iron-replete coastal waters with a strong chelating ligand. Limnology and Oceanography, 44(4):1009–1018, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Hoch:1992:IFA

[HFK92] Matthew P. Hoch, Marilyn L. Fogel, and David L. Kirchman. Isotope fractionation associated with ammonium uptake by a marine bacterium. Limnology and Oceanography, 37(7):1447–1459, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Hoppema:1999:RBC

[HG99] Mario Hoppema and Leo Goeyens. Redfield behavior of carbon, nitrogen, and phosphorus depletions in Antarctic surface water. Limnology and Oceanography, 44(1):220–224, January 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Hunter:1996:RSC

[HGFC96] Craig N. Hunter, R. Michael Gordon, Steve E. Fitzwater, and Kenneth H. Coale. A rosette system for the collection of trace metal clean seawater. Limnology and Oceanography, 41(6):1367– 1372, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Horner:1997:HAB

[HGP97] Rita A. Horner, David L. Garrison, and F. Gerald Plumley. Harmful algal blooms and red tide problems on the U.S. west coast. Limnology and Oceanography, 42(5 (part 2)):1076–1088, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Hupfer:1995:PLS

[HGR95] Michael Hupfer, Red Gtichter, and RedHeinz R. Ruegger. Polyphosphate in lake sediments: ^{31}P NMR spectroscopy as a tool for its identification. Limnology and Oceanography, 40(3): 610–617, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Hambright:1994:ILT

[HGS94] K. David Hambright, Moshe Gophen, and Salvador Serruya. Influence of long-term climatic changes on the stratification of a subtropical, warm monomictic lake. Limnology and Oceanography, 39 (5):1233–1242, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Havel:1993:DLN

[HH93] John E. Havel and Paul D. N. Hebert. Daphnia lumholtzi in North America: Another exotic zooplankter. Limnology and Oceanography, 38(8):1823–1827, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Hawkins:1992:MED

[HHB^{+}92] L. E. Hawkins, S. Hutchinson, S. D. Batten, G. L. J. Paterson, P. J. D. Lambshead, and A. L. Rice. A method for the enzymic determination of glycogen in chemically preserved material. Limnology and Oceanography, 37(8):1784–1786, December 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Haygood:1993:APP

[HHB93] Margo G. Haygood, Pamela D. Holt, and Alison Butler. Aerobactin production by a planktonic marine Vibrio sp. Limnology and Oceanography, 38(5):1091–1097, July 1993. CODEN LI-OCAH. ISSN 0024-3590.

Harrison:1996:KNU

[HHI96] W. G. Harrison, L. R. Harris, and B. D. Irwin. The kinetics of nitrogen utilization in the oceanic mixed layer: Nitrate and ammonium interactions at nanomolar concentrations. Limnology and Oceanography, 41(1):16–32, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Hood:1998:DGD

[HHJ98] E. M. Hood, B. L. Howes, and W. J. Jenkins. Dissolved gas dynamics in perennially ice-covered Lake Fryxell, Antarctica. Limnology and Oceanography, 43(2):265–272, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Hurtubise:1998:EUB

[HHL98] Robin D. Hurtubise, John E. Havel, and Edward E. Little. The effects of ultraviolet-B radiation on freshwater invertebrates: Experiments with a solar simulator. Limnology and Oceanography, 43(6):1082–1088, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Head:1994:APL

[HHR94] E. J. H. Head, B. T. Hargrave, and D. V. Subba Rao. Accumulation of a pheophorbide a-like pigment in sediment traps during late stages of a spring bloom: a product of dying algae? Limnology and Oceanography, 39(1):176–181, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Hilton:1990:GMP

[Hil90] J. Hilton. Greigite and the magnetic properties of sediments. Limnology and Oceanography, 35(2):509–520, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Hines:1994:ACO

[Hin94] Mark E. Hines. Acetate concentrations and oxidation in saltmarsh sediments. Limnology and Oceanography, 39(1):140–148, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Hentschel:1994:SCI

[HJ94] Brian T. Hentschel and Peter A. Jumars. In situ chemical inhibition of benthic diatom growth affects recruitment of competing, permanent and temporary meiofauna. Limnology and Oceanography, 39(4):816–838, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Hoch:1995:AUH

[HK95] Matthew P. Hoch and David L. Kirchman. Ammonium uptake by heterotrophic bacteria in the Delaware estuary and adjacent coastal waters. Limnology and Oceanography, 40(5):886–897, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Hagen:1998:LMA

[HK98a] Wilhelm Hagen and Gerhard Kattner. Lipid metabolism of the Antarctic euphausiid Thysanmssa macrura and its ecological implications. Limnology and Oceanography, 43(8):1894–1901, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Hagerthey:1998:GFI

[HK98b] Scot E. Hagerthey and W. Charles Kerfoot. Groundwater flow influences the biomass and nutrient ratios of epibenthic algae in a north temperate seepage lake. Limnology and Oceanography, 43 (6):1227–1242, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Hohmann:1997:PDW

[HKP^{+}97] R. Hohmann, R. Kipfer, F. Peeters, G. Piepke, D. M. Imboden, and M. N. Shimaraev. Processes of deep-water renewal in Lake Baikal. Limnology and Oceanography, 42(5):841–855, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Herut:1999:AIN

[HKPM99] Barak Herut, Michael D. Krom, Gang Pan, and Robert Mortimer. Atmospheric input of nitrogen and phosphorus to the southeast Mediterranean: Sources, fluxes, and possible impact. Limnology and Oceanography, 44(7):1683–1692, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Hemminga:1992:SSC

[HKvS^{+}92] M. A. Hemminga, V. A. Klap, J. van Soelen, J. de Leeuw, and J. J. Boon. Shifts in seston characteristics after inundation of a European coastal salt marsh. Limnology and Oceanography, 37(7): 1559–1564, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Hawley:1992:SRL

[HL92] Nathan Hawley and Barry M. Lesht. Sediment resuspension in Lake St. Clair. Limnology and Oceanography, 37(8):1720–1737, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Hofmann:1994:BRT

[HL94] Eileen E. Hofmann and Cathy M. Lascara. Book review: Towards a model of ocean biogeochemical processes (G. T. Evans and M. J. R. Fasham [eds.]). Limnology and Oceanography, 39(7):1780–1781, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Herrin:1998:HMU

[HLGA98] Russell T. Herrin, Richard C. Lathrop, Patrick R. Gorski, and Anders W. Andren. Hypolimnetic methylmercury and its uptake by plankton during fall destratification: a key entry point of mercury into lake food chains? Limnology and Oceanography, 43(7): 1476–1486, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Harrold:1998:OES

[HLL98] Christopher Harrold, Karen Light, and Susan Lisin. Organic enrichment of submarine-canyon and continental-shelf benthic communities by macroalgal drift imported from nearshore kelp forests. Limnology and Oceanography, 43(4):669–678, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Hall:1999:EAU

[HLQ^{+}99] Roland I. Hall, Peter R. Leavitt, Roberto Quinlan, Aruna S. Dixit, and John P. Smol. Effects of agriculture, urbanization, and climate on water quality in the northern Great Plains. Limnology and Oceanography, 44(3 (part 2)):739–756, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Hudson:1990:LTM

[HM90] Robert J. M. Hudson and FraÇois M. M. Morel. lron transport in marine phytoplankton: Kinetics of cellular and medium coordination reactions. Limnology and Oceanography, 35(5):1002–1020, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Hill:1999:SAS

[HM99] Walter R. Hill and Amy E. McNamara. Sun angle, site selection, parameter choice, and generalizing UV effects in freshwater benthic ecosystems (reply to the comment by Donahue and Clare). Limnology and Oceanography, 44(2):474–475, March 1999. CO-DEN LIOCAH. ISSN 0024-3590. See [DC99].

Hansell:1994:DES

[HN94] Dennis A. Hansell and Jan A. Newton. Design and evaluation of a “swimmer”-segregating particle interceptor trap. Limnology and Oceanography, 39(6):1487–1495, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Haga:1995:SFN

[HNS95] Hiroki Haga, Toshi Nagata, and Mitsuru Sakamoto. Sizefractionated NH_{4}^{+} regeneration in the pelagic environments of two mesotrophic lakes. Limnology and Oceanography, 40(6):1091– 1099, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Hoflmann:1990:DPV

[Hof90] James P. Hoflmann. Dependence of photosynthesis and vitamin B_{12} uptake on cellular vitamin B_{12} concentration in the multicellular alga Cladophora glomerata (Chlorophyta). Limnology and Oceanography, 35(1):100–108, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Holliday:1999:FAO

[Hol99] D. V. Holliday. Fundamentals of acoustical oceanography. Limnology and Oceanography, 46(4):1183, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Hongve:1997:CIM

[Hon97] Dag Hongve. Cycling of iron, manganese, and phosphate in a meromictic lake. Limnology and Oceanography, 42(4):635–647, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Huh:1995:RCU

[HP95] Chih-An Huh and Fredrick G. Prahl. Role of colloids in upper ocean biogeochemistry in the northeast Pacific Ocean elucidated from 238U-234Th disequilibria. Limnology and Oceanography, 40 (3):528–532, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Hairston:1999:PBF

[HPB^{+}99] Nelson G. Hairston, Jr., Linda J. Perry, Andrew J. Bohonak, Meghan Q. Fellows, Colleen M. Kearns, and Daniel R. Engstrom. Population biology of a failed invasion: Paleolimnology of Daphnia exilis in upstate New York. Limnology and Oceanography, 44 (3):477–486, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Hays:1994:IDD

[HPJW94] G. C. Hays, C. A. Proctor, A. W. G. John, and A. J. Warner. Interspecific differences in the diel vertical migration of marine copepods: The implications of size, color, and morphology. Limnology and Oceanography, 39(7):1621–1629, November 1994. CO-DEN LIOCAH. ISSN 0024-3590.

Hall:1996:EIC

[HPM96] Robert O. Hall, Jr., Christopher L. Peredney, and Judy L. Meyer. The effect of invertebrate consumption on bacterial transport in a mountain stream. Limnology and Oceanography, 41(6):1180–1187, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Hewitt:1997:AST

[HPTC97] J. E. Hewitt, R. D. Pridmore, S. F. Thrush, and V. J. Cummings. Assessing the short-term stability of spatial patterns of macrobenthos in a dynamic estuarine system. Limnology and Oceanography, 42(2):282–288, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Hayworth:1990:TLR

[HQ90] Anita M. Hayworth and James F. Quinn. Temperature of limpets in the rocky intertidal zone: Effects of caging and substratum. Limnology and Oceanography, 35(4):967–970, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Hill:1992:SCL

[HQCG92] Cathy Hill, Michael A. Quigley, Joann F. Cavaletto, and Wendy Gordon. Seasonal changes in lipid content and composition in the benthic amphipods Monoporeia affinis and Pontoporeia femorata. Limnology and Oceanography, 37(6):1280–1289, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Howarth:1990:NFC

[HR90] Robert W. Howarth and Marino Roxanne. Nitrogen-fixing cyanobacteria in the plankton of lakes and estuaries: a reply to the comment by Smith. Limnology and Oceanography, 35(8):1859– 1863, December 1990. CODEN LIOCAH. ISSN 0024-3590. See [Smi90b].

Hilton:1995:MIE

[HRD^{+}95] J. Hilton, E. Rigg, W. Davison, J. Hamilton-Taylor, M. Kelly, F. R. Livens, and D. L. Singleton. Modeling and interpreting element ratios in water and sediments: a sensitivity analysis of post-Chernobyl Ru : Cs ratios. Limnology and Oceanography, 40 (7):1302–1309, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Hutchins:1991:SPN

[HRF91] Dave A. Hutchins, John G. Rueter, and William Fish. Siderophore production and nitrogenfixation are mutually exclusive strategies in Anabaena 712. Limnology and Oceanography, 36(1):1–12, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Horrigan:1990:OEA

[HS90] S. G. Horrigan and A. L. Springer. Oceanic and estuarine ammonium oxidation: Effects of light. Limnology and Oceanography, 35 (2):479–482, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Hoepffner:1992:BOC

[HS92] Nicolas Hoepffner and Shubha Sathyendranath. Bio-optical characteristics of coastal waters: Absorption spectra of phytoplankton and pigment distribution in the western North Atlantic. Limnology and Oceanography, 37(8):1660–1679, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Hennes:1995:DCV

[HS95] Kilian P. Hennes and Curtis A. Suttle. Direct counts of viruses in natural waters and laboratory cultures by epifluorescence microscopy. Limnology and Oceanography, 40(6):1050–1055, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Hodell:1998:PSI

[HS98] David A. Hodell and Claire L. Schelske. Production, sedimentation, and isotopic composition of organic matter in Lake Ontario. Limnology and Oceanography, 43(2):200–214, March 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Hillebrand:1999:NSB

[HS99] Helmut Hillebrand and Ulrich Sommer. The nutrient stoichiometry of benthic microalgal growth: Redfield proportions are optimal. Limnology and Oceanography, 44(2):440–446, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Hamilton:1997:AEO

[HSCM97] Stephen K. Hamilton, Suzanne J. Sippel, Dóbora F. Calheiros, and John M. Melack. An anoxic event and other biogeochemical effects of the Pantanal wetland on the Paraguay River. Limnology and Oceanography, 42(2):257–272, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Hodell:1998:BIC

[HSFR98] David A. Hodell, Claire L. Schelske, Gary L. Fahnenstiel, and Lisa L. Robbins. Biologically induced calcite and its isotopic composition in Lake Ontario. Limnology and Oceanography, 43(2): 187–199, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Haney:1995:EPR

[HSI95] James F. Haney, John J. Sasner, and Miyoshi Ikawa. Effects of products released by Aphanizomenon flos-aquae and purified saxitoxin on the movements of Daphnia carinata feeding appendages. Limnology and Oceanography, 40(2):263–272, March 1995. CO-DEN LIOCAH. ISSN 0024-3590.

Hurd:1997:VSF

[HSL^{+}97] Catriona L. Hurd, Craig L. Stevens, Bernard E. Lava, G. A. Lawrence, and P. J. Harrison. Visualization of seawater flow around morphologically distinct forms of the giant kelp Macrocystis integrifolia from wave-sheltered and exposed sites. Limnology and Oceanography, 42(1):156–163, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Howarth:1995:TDP

[HSM^{+}95] R. W. Howarth, D. Swaney, R. Marino, T. Butler, and C. R. Chu. Turbulence does not prevent nitrogen fixation by plankton in estuaries and coastal seas (reply to the comment by Paerl et al.). Limnology and Oceanography, 40(3):639–643, May 1995. CODEN LIOCAH. ISSN 0024-3590. See [PPK95].

Hata:1998:CFS

[HSM^{+}98] Hiroshi Hata, Atsushi Suzuki, Tadashi Maruyama, Norihide Kurano, Sigetoh Miyachi, Yutaka Ikeda, and Hajime Kayanne. Carbon flux by suspended and sinking particles around the barrier reef of Palau, western Pacific. Limnology and Oceanography, 43 (8):1883–1893, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Herrera-Silveira:1996:ENP

[HSRR96] Jorge A. Herrera-Silveira and Javier Ramírez-Ramírez. Effects of natural phenolic material (tannin) on phytoplankton growth. Limnology and Oceanography, 41(5):1018–1023, July 1996. CO-DEN LIOCAH. ISSN 0024-3590.

Hughes:1999:ABC

[HSS^{+}99] Terry Hughes, Alina M. Szmant, Robert Steneck, Robert Carpenter, and Steven Miller. Algal blooms on coral reefs: What are the causes? Limnology and Oceanography, 44(6):1583–1586, September 1999. CODEN LIOCAH. ISSN 0024-3590. See reply [Lap99].

Hargrave:1994:CAS

[HSSA94] B. Hargrave, G. Siddall, G. Steeves, and G. Awalt. A currentactivated sediment trap. Limnology and Oceanography, 39(2):383– 390, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Hudson:1996:MRD

[HT96] Jeff J. Hudson and William D. Taylor. Measuring regeneration of dissolved phosphorus in planktonic communities. Limnology and Oceanography, 41(7):1560–1565, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Hare:1998:AIC

[HT98] Landis Hare and André Tessier. The aquatic insect Chaoborus as a biomonitor of trace metals in lakes. Limnology and Oceanography, 43(8):1850–1859, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Hamilton-Taylor:1996:BCZ

[HTDM96] J. Hamilton-Taylor, W. Davison, and K. Morfett. The biogeochemical cycling of Zn, Cu, Fe, Mn, and dissolved organic C in a seasonally anoxic lake. Limnology and Oceanography, 41(3):408– 418, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Hamilton-Taylor:1999:NDS

[HTSDZ99] J. Hamilton-Taylor, E. J. Smith, W. Davison, and H. Zhang. A novel DGT-sediment trap device for the in situ measurement of element remobilization from settling particles in water columns and its application to trace metal release from Mn and Fe oxides. Limnology and Oceanography, 44(7):1772–1780, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Hamilto-Taylor:1990:QAS

[HTW90] J. Hamilto-Taylor and M. Willis. A quantitative assessment of the sources and general dynamics of trace metals in a soft-water lake. Limnology and Oceanography, 35(4):840–851, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Huryn:1990:GVL

[Hur90] Alexander D. Huryn. Growth and voltinism of lotic midge larvae: Patterns across an Appalachian Mountain basin. Limnology and Oceanography, 35(2):339–351, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Huryn:1996:AAP

[Hur96] Alexander D. Huryn. An appraisal of the Allen paradox in a New Zealand trout stream. Limnology and Oceanography, 41(2): 243–252, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Hu:1998:MSS

[HV98] Chuanmin Hu and Kenneth J. Voss. Measurement of solarstimulated fluorescence in natural waters. Limnology and Oceanography, 43(6):1198–1206, September 1998. CODEN LI-OCAH. ISSN 0024-3590.

Hoge:1993:IOP

[HVB93] Frank E. Hoge, Anthony Vodacek, and Neil V. Blough. Inherent optical properties of the ocean: Retrieval of the absorption coefficient of chromophoric dissolved organic matter from fluorescence measurements. Limnology and Oceanography, 38(7):1394–1402, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Helbling:1991:EIP

[HVHH91] E. Walter Helbling, Virginia Villafañe, and Osmund Holm-Hansen. Effect of iron on productivity and size distribution of Antarctic phytoplankton. Limnology and Oceanography, 36(8): 1879–1885, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Huisman:1999:CDC

[HvOW99] Jef Huisman, Paul van Oostveen, and Franz J. Weissing. Critical depth and critical turbulence: Two different mechanisms for the development of phytoplankton blooms. Limnology and Oceanography, 44(7):1781–1787, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Hurley:1991:IBL

[HW91] James P. Hurley and Carl J. Watras. Identification of bacteriochlorophylls in lakes via reverse-phase HPLC. Limnology and Oceanography, 36(2):307–315, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Hutchinson:1994:DBG

[HW94] Paul A. Hutchinson and Ian T. Webster. On the distribution of blue-green algae in lakes: Wind-tunnel tank experiments. Limnology and Oceanography, 39(2):374–382, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Hutchins:1995:CGB

[HWF95] David A. Hutchins, Wen-Xiong Wang, and Nicholas S. Fisher. Copepod grazing and the biogeochemical fate of diatom iron. Limnology and Oceanography, 41(4):989–994, July 1995. CODEN LI-OCAH. ISSN 0024-3590.

Hays:1995:STP

[HWP95] Graeme C. Hays, Andrew J. Warner, and Catherine A. Proctor. Spatio-temporal patterns in the diel vertical migration of the copepod Metridia lucens in the northeast Atlantic derived from the Continuous Plankton Recorder survey. Limnology and Oceanography, 40(3):469–475, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Hall:1997:CAP

[HY97] Roland I. Hall and Norman D. Yan. Comparing annual population growth estimates of the exotic invader bythotrephes by using sediment and plankton records. Limnology and Oceanography, 42 (1):112–120, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Huettel:1996:FIU

[HZF96] M. Huettel, W. Ziebis, and S. Forster. Flow-induced uptake of particulate matter in permeable sediments. Limnology and Oceanography, 41(2):309–322, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Istvanovics:1995:ENU

[IH95] Vera Istvánovics and Sándor Herodek. Estimation of net uptake and leakage rates of orthophosphate from ^{32}P-uptake kinetics by a linear force-flow model. Limnology and Oceanography, 40(1): 17–32, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Interlandi:1999:RPV

[IKT99] Sebastian J. Interlandi, Susan S. Kilham, and Edward C. Theriot. Responses of phytoplankton to varied resource availability in large lakes of the greater Yellowstone ecosystem. Limnology and Oceanography, 44(3):668–682, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Isiorho:1990:GRL

[IM90] Solomon A. Isiorho and G. Matisoff. Groundwater recharge from Lake Chad. Limnology and Oceanography, 35(4):931–938, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Ibelings:1998:PAI

[IM98] Bas W. Ibelings and Stephen C. Maberly. Photoinhibition and the availability of inorganic carbon restrict photosynthesis by surface blooms of cyanobacteria. Limnology and Oceanography, 43(3): 408–419, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Ianora:1999:FES

[IMB^{+}99] A. Ianora, A. Miralto, I. Buttino, G. Romano, and S. A. Poulet. First evidence of some dinoflagellates reducing male copepod fertilization capacity. Limnology and Oceanography, 44(1):147–153, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Ianora:1993:EVC

[IP93] Adrianna Ianora and Serge A. Poulet. Egg viability in the copepod Temora stylifera. Limnology and Oceanography, 38(8):1615–1626, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Ibanez:1997:CER

[IPP97] Carles Ibañez, Didier Pont, and Narcís Prat. Characterization of the Ebre and Rhone estuaries: a basis for defining and classifying salt-wedge estuaries. Limnology and Oceanography, 42(1):89–101, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Istanovics:1992:EPD

[IPPD92] Vera Istánovics, Kurt Pettersson, Don Pierson, and Russell Duarte. Evaluation of phosphorus deficiency indicators for summer phytoplankton in Lake Erken. Limnology and Oceanography, 37(4):890–900, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Iverson:1990:CMF

[Ive90] Richard L. Iverson. Control of marine fish production. Limnology and Oceanography, 35(7):1593–1604, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Jensen:1992:ITN

[JA92] Henning Skovgaard Jensen and Frede Ostergaard Andersen. Importance of temperature, nitrate, and pH for phosphate release from aerobic sediments of four shallow, eutrophic lakes. Limnology and Oceanography, 37(3):577–589, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Jackson:1993:FFM

[Jac93] George A. Jackson. Flux feeding as a mechanism for zooplankton grazing and its implications for vertical particulate flux. Limnology and Oceanography, 38(6):1328–1331, September 1993. CO-DEN LIOCAH. ISSN 0024-3590. See reply [NT93a].

Jellison:1996:OMA

[JAMH96] Robert Jellison, Robert F. Anderson, John M. Melack, and Darla Heil. Organic matter accumulation in sediments of hypersaline Mono Lake during a period of changing salinity. Limnology and Oceanography, 41(7):1539–1544, November 1996. CODEN LI-OCAH. ISSN 0024-3590.

Jansson:1993:UEE

[Jan93] Mats Jansson. Uptake, exchange and excretion of orthophosphate in phosphate-starved Scenedesmus quadricauda and Pseudomonas K7. Limnology and Oceanography, 38(6):1162–1178, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1995:AFC

[JAS95] Bruce D. Johnson and Kumiko Azetsu-Scott. Adhesion force and the character of surfaces immersed in seawater. Limnology and Oceanography, 40(4):802–808, June 1995. CODEN LIOCAH. ISSN 0024-3590.

James:1991:EPE

[JB91a] William F. James and John W. Barko. Estimation of phosphorus exchange between littoral and pelagic zones during nighttime convective circulation. Limnology and Oceanography, 36(1):179–187, January 1991. CODEN LIOCAH. ISSN 0024-3590.

James:1991:LPP

[JB91b] William F. James and John W. Barko. Littoral-pelagic phosphorus dynamics during nighttime convective circulation. Limnology and Oceanography, 36(5):949–960, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Jenkins:1994:TVS

[JB94] Gregory P. Jenkins and Kerry P. Black. Temporal variability in settlement of a coastal fish (Sillaginodes punctata) determined by low-frequency hydrodynamics. Limnology and Oceanography, 39 (7):1744–1754, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Jansson:1996:NLB

[JBJB96] Mats Jansson, Peter Blomqvist, Anders Jonsson, and Ann-Kristin Bergström. Nutrient limitation of bacterioplankton, autotrophic and mixotrophic phytoplankton, and heterotrophic nanoflagellates in Lake Örträsket. Limnology and Oceanography, 41(7): 1552–1559, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Jahnke:1995:QRH

[JC95] Richard A. Jahnke and Deborah B. Craven. Quantifying the role of heterotrophic bacteria in the carbon cycle: a need for respiration rate measurements. Limnology and Oceanography, 40(2):

436–441, March 1995. CODEN LIOCAH. ISSN 0024-3590. See reply [CP95b].

Judge:1992:ELT

[JCH92] Michael L. Judge, Loren D. Coen, and K. L. Heck, Jr. The effect of long-term alteration of in situ currents on the growth of Mercenaria mercenaria in the northern Gulf of Mexico. Limnology and Oceanography, 37(7):1550–1559, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Joye:1999:OAM

[JCM^{+}99] Samantha B. Joye, Tracy L. Connell, Laurence G. Miller, Ronald S. Oremland, and Robert S. Jellison. Oxidation of ammonia and methane in an alkaline, saline lake. Limnology and Oceanography, 44(1):178–188, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Jordan:1991:NCI

[JCMW91] Thomas E. Jordan, David L. Correll, Joseph Miklas, and Donald E. Weller. Nutrients and chlorophyll at the interface of a watershed and an estuary. Limnology and Oceanography, 36(2): 251–267, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:1990:DBL

[JD90] Bo Barker Jørgensen and David J. Des Marais. The diffusive boundary layer of sediments: Oxygen microgradients over a microbial mat. Limnology and Oceanography, 35(6):1343–1355, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Jurgens:1995:BFP

[JD95] Klaus Jürgens and William R. DeMott. Behavioral flexibility in prey selection by bacterivorous nanoflagellates. Limnology and Oceanography, 40(8):1503–1507, December 1995. CODEN LI-OCAH. ISSN 0024-3590.

Johannesse:1996:DOC

[JD96a] Tore Johannesse and Einar Dahl. Declines in oxygen concentrations along the Norwegian Skagerrak coast, 1927–1993: a signal of ecosystem changes due to eutrophication? Limnology and Oceanography, 41(8):766–778, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Johannessen:1996:HCO

[JD96b] Tore Johannessen and Einar Dahl. Historical changes in oxygen concentrations along the Norwegian Skagerrak coast: Reply to the comment by Gray and Abdullah. Limnology and Oceanography, 41(4):1847–1852, December 1996. CODEN LIOCAH. ISSN 0024-3590. See [GA96].

Jeremiason:1999:BCP

[JEP^{+}99] Jeff D. Jeremiason, Steven J. Eisenreich, Michael J. Paterson, Ken G. Beaty, Robert Hecky, and James J. Elser. Biogeochemical cycling of PCBs in lakes of variable trophic status: a paired-lake experiment. Limnology and Oceanography, 44(3 (part 2)):889–902, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Jonasz:1996:ASD

[JF96] Miroslaw Jonasz and Georges Fournier. Approximation of the size distribution of marine particles by a sum of log-normal functions. Limnology and Oceanography, 41(4):744–754, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Jonasz:1999:ASD

[JF99] Miroslaw Jonasz and Georges Fournier. Approximation of the size distribution of marine particles by a sum of log-normal functions (errata: Corrections and additional results). Limnology and Oceanography, 44(5):1358, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Jassby:1999:OSD

[JGRR99] Alan D. Jassby, Charles R. Goldman, John E. Reuter, and Robert C. Richards. Origins and scale dependence of temporal variability in the transparency of Lake Tahoe, California–Nevada. Limnology and Oceanography, 44(2):282–294, March 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Jannasch:1996:MSR

[JHM96] Hans W. Jannasch, Bruce D. Honeyman, and James W. Murray. Marine scavenging: The relative importance of mass transfer and reaction rates. Limnology and Oceanography, 41(1):82–88, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Jumars:1996:ARS

[JJGS96] Peter A. Jumars, Darrell R. Jackson, Thomas F. Gross, and Christopher Sherwood. Acoustic remote sensing of benthic activity: a statistical approach. Limnology and Oceanography, 41(6): 1220–1241, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Jackson:1992:ECN

[JL92] George A. Jackson and Steve E. Lochmann. Effect of coagulation on nutrient and light limitation of an algal bloom. Limnology and Oceanography, 37(1):77–89, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Johannesson:1994:REE

[JL94] Kevin H. Johannesson and W. Berry Lyons. The rare earth element geochemistry of Mono Lake water and the importance of carbonate complexing. Limnology and Oceanography, 39(5):1141– 1154, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Jacquet:1998:PPD

[JLMV98] Stéphan Jacquet, Jean-François Lennon, Dominique Marie, and Daniel Vaulot. Picoplankton population dynamics in coastal waters of the northwestern Mediterranean Sea. Limnology and Oceanography, 43(8):1916–1931, December 1998. CODEN LI-OCAH. ISSN 0024-3590.

Jokiel:1997:UAC

[JLO97] Paul L. Jokiel, Michael P. Lesser, and Michael E. Ondrusek. UVabsorbing compounds in the coral Pocillopora damicornis: Interactive effects of UV radiation, photosynthetically active radiation, and water flow. Limnology and Oceanography, 42(6):1468–1473, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Jellison:1993:APA

[JM93a] Robert Jellison and John M. Melack. Algal photosynthetic activity and its response to meromixis in hypersaline Mono Lake, California. Limnology and Oceanography, 38(4):818–837, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Jellison:1993:MHMa

[JM93b] Robert Jellison and John M. Melack. Meromixis in hypersaline Mono Lake, California. 1. Stratification and vertical mixing during the onset, persistence, and breakdown of meromixis. Limnology and Oceanography, 38(5):1008–1019, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Jones:1998:MIE

[JM98] Jeremy B. Jones, Jr. and Patrick J. Mulholland. Methane input and evasion in a hardwood forest stream: Effects of subsurface flow from shallow and deep pathway. Limnology and Oceanography, 43(6):1243–1250, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Jensen:1995:PCC

[JMA^{+}95] Henning S. Jensen, P. B. Mortensen, F. O. Andersen, E. Rasmussen, and A. Jensen. Phosphorus cycling in a coastal marine sediment, Aarhus Bay, Denmark. Limnology and Oceanography, 40(5):908–917, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Jellison:1993:MHMb

[JMMD93] Robert Jellison, Laurence G. Miller, John M. Melack, and Gayle L. Dana. Meromixis in hypersaline Mono Lake, California. 2. Nitrogen fluxes. Limnology and Oceanography, 38(5):1020– 1039, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Jensen:1998:FAS

[JMMH98] Henning Skovgaard Jensen, Karen J. McGlathery, Roxanne Marino, and Robert W. Howarth. Forms and availability of sediment phosphorus in carbonate sand of Bermuda seagrass beds. Limnology and Oceanography, 43(5):799–810, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Juhl:1996:ACP

[JOG96] Andrew R. Juhl, Mark D. Ohman, and Ralf Goericke. Astaxanthin in Calanus pacificus: Assessment of pigment-based measures of omnivory. Limnology and Oceanography, 41(6):1198–1207, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1990:FAP

[Joh90] Amy S. Johnson. Flow around phoronids: Consequnces of a neighbor to suspension feeders. Limnology and Oceanography, 35(6): 1395–1401, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1991:BRC

[Joh91] Dale W. Johnson. Book review: Comparative analyses of ecosystems. Patterns, mechanisms, and theories (J. Cole, G. Lovett, and S. Findlay [eds.]). Limnology and Oceanography, 36(7):1506, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1994:ESM

[Joh94] Ladd E. Johnson. Enhanced settlement on microtopographical high points by the intertidal red alga halosaccion glandiforme. Limnology and Oceanography, 39(8):1893–1902, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Jones:1991:IFM

[Jon91a] Ronald D. Jones. An improved fluorescence method for the determination of nanomolar concentrations of ammonium in natural waters. Limnology and Oceanography, 36(4):814–819, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Jonsson:1991:ITM

[Jön91b] Benno Jönsson. A ^{14}C-incubation technique for measuring microphytobenthic primary productivity in intact sediment cores. Limnology and Oceanography, 36(7):1485–1492, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Jonasson:1995:BRT

[Jón95] Pétur M. Jónasson. Book review: A treatise on limnology. V. 4. The zoobenthos. (G. E. Hutchinson). Limnology and Oceanography, 40(4):839–841, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:1990:SCF

[Jør90] Bo Barker Jørgensen. The sulfur cycle of freshwater sediments: Role of thiosulfate. Limnology and Oceanography, 35(6):1329– 1342, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:1993:CMS

[Jør93] C. Barker Jørgensen. A comment on “Mechanisms of suspension feeding in bivalves: Resolution of current controversies by means of endoscopy” (Ward et al.). Limnology and Oceanography, 38(2): 466, March 1993. CODEN LIOCAH. ISSN 0024-3590. See reply [WMTB93b].

Jassby:1990:IFP

[JPG90] Alan D. Jassby, Thomas M. Powell, and Charles R. Goldman. Interannual fluctuations in primary production: Direct physical effects and the trohic cascade at Castle Lake, California. Limnology and Oceanography, 35(5):1021–1038, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Johnsen:1997:FES

[JPJ97] Geir Johnsen, Barbara B. Prézelin, and Raffael V. M. Jovine. Fluorescence excitation spectra and light utilization in two red tide dinoflagellates. Limnology and Oceanography, 42(5 (part 2)): 1166–1177, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Johan:1990:PBA

[JRH90] Wikner Johan, Fereidoun Rassoulzadegan, and Åke Hagström. Periodic bacterivore activity balances bacterial growth in the marine environment. Limnology and Oceanography, 35(2):313–324, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Jellison:1998:OMD

[JRM98] Robert Jellison, José Romero, and John M. Melack. The onset of meromixis during restoration of Mono Lake, California: Unintended consequences of reducing water diversions. Limnology and Oceanography, 43(4):706–711, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Justic:1996:ECC

[JRT96] Dubravko Justić, Nancy N. Rabalais, and R. Eugene Turner. Effects of climate change on hypoxia in coastal waters: a doubled CO_{2} scenario for the northern Gulf of Mexico. Limnology and Oceanography, 41(5):992–1003, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1992:CCS

[JSL92] C. Annette Johnson, Laura Sigg, and Ursula Lindauer. The chromium cycle in a seasonally anoxic lake. Limnology and Oceanography, 37(2):315–321, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Jumars:1990:WHI

[Jum90] Peter A. Jumars. W(h)ither limnology? Limnology and Oceanography, 35(5):1216–1218, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Jumars:1994:BRB

[Jum94] Peter A. Jumars. Book review: Biological oceanography: an introduction (C. M. Lalli and T. R. Parsons). Limnology and Oceanography, 39(4):982, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Jungmann:1992:TCI

[Jun92] Dirk Jungmann. Toxic compounds isolated from microcystis PCC786 that are more active against Daphnia than two microcystins. Limnology and Oceanography, 37(8):1777–1783, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1991:MMM

[JUSI91] C. A. Johnson, M. Ulrich, L. Sigg, and D. M. Imboden. A mathematical model of the manganese cycle in a seasonally anoxic lake. Limnology and Oceanography, 36(7):1415–1426, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Janse:1999:MBP

[JvROG99] Ingmar Janse, Marion van Rijssel, Alex Ottema, and Jan C. Gottschal. Microbial breakdown of Phaeocystis mucopolysaccharides. Limnology and Oceanography, 44(6):1447–1457, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1992:PBC

[JW92] Richard K. Johnson and Torgny Wiederholm. Pelagic–benthic coupling — the importance of diatom interannual variability for population oscillations of Monoporeia affinis. Limnology and Oceanography, 37(8):1596–1607, December 1992. CODEN LI-OCAH. ISSN 0024-3590.

Jurgens:1996:FRM

[JWRS96] Klaus Jürgens, Stephen A. Wickham, Karl Otto Rothhaupt, and Barbara Santer. Feeding rates of macro- and microzooplankton on heterotrophic nanoflagellates. Limnology and Oceanography, 41(4):1833–1839, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Jones:1994:SSL

[JX94] Ian S. F. Jones and Jishi Xie. A sound scattering layer in a freshwater reservoir. Limnology and Oceanography, 39(2):443–448, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Kristensen:1992:EBM

[KAB92] Erik Kristensen, Frede Østerguard Andersen, and Thomas Henry Blackburn. Effects of benthic macrofauna and temperature on degradation of macroalgal detritus: The fate of organic carbon. Limnology and Oceanography, 37(7):1404–1419, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Kristensen:1995:AAD

[KAD95] Erik Kristensen, Saiyed I. Ahmed, and Allan H. Devol. Aerobic and anaerobic decomposition of organic matter in marine sediment: Which is fastest? Limnology and Oceanography, 40(8): 1430–1437, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Kalff:1991:TFL

[Kal91] Jacob Kalff. On the teaching and funding of limnology. Limnology and Oceanography, 36(7):1499–1501, November 1991. CODEN LIOCAH. ISSN 0024-3590. See reply [?].

Kamenir:1992:FEN

[Kam92] Yury G. Kamenir. Further evidence of nonnutrient constraints on phytoplankton abundances. Limnology and Oceanography, 37 (3):681–683, May 1992. CODEN LIOCAH. ISSN 0024-3590. See reply [ADC92].

Kana:1993:ROC

[Kan93] Todd M. Kana. Rapid oxygen cycling in Trichodesmium thiebautii. Limnology and Oceanography, 38(1):18–24, January 1993. CO-DEN LIOCAH. ISSN 0024-3590.

Kaplan:1992:CHT

[Kap92] Louis A. Kaplan. Comparison of high-temperature and persulfate oxidation methods for determination of dissolved organic carbon in freshwaters. Limnology and Oceanography, 37(5):1119–1125, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Kaplan:1994:FLP

[Kap94] Louis A. Kaplan. A field and laboratory procedure to collect, process, and preserve fresh-water samples for dissolved organic carbon analysis. Limnology and Oceanography, 39(6):1470–1476, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Kiefer:1992:REM

[KB92] Dale A. Kiefer and Juli Berwald. A random encounter model for the microbial planktonic community. Limnology and Oceanography, 37(3):457–467, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Kimmerer:1998:TOV

[KBB98] W. J. Kimmerer, J. R. Burau, and W. A. Bennett. Tidally oriented vertical migration and position maintenance of zooplankton in a temperate estuary. Limnology and Oceanography, 43(7):1697– 1709, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Kleppel:1998:NRE

[KBH98] G. S. Kleppel, C. A. Burkart, and L. Houchin. Nutrition and the regulation of egg production in the Calanoid copepod Acartia tonsa. Limnology and Oceanography, 43(5):1000–1007, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Karp-Boss:1998:MDC

[KBJ98] Lee Karp-Boss and Peter A. Jumars. Motion of diatom chains in steady shear flow. Limnology and Oceanography, 43(8):1767–1773, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Kuhn:1999:PUR

[KBMSP99] Penny Kuhn, Howard Browman, Bruce McArthur, and Jean-Francois St-Pierre. Penetration of ultraviolet radiation in the waters of the estuary and Gulf of St. Lawrence. Limnology and Oceanography, 44(3):710–716, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Knoechel:1992:SID

[KC92] R. Knoechel and C. E. Campbell. A simple, inexpensive device for obtaining vertically integrated, quantitative samples of pelagic zooplankton. Limnology and Oceanography, 37(3):675–680, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Kelleyl:1998:SIE

[KCC98] Cheryl A. Kelleyl, Richard B. CofJin, and Luis A. Cifuentes. Stable isotope evidence for alternative bacterial carbon sources in the Gulf of Mexico. Limnology and Oceanography, 43(8):1962–1969, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Kremer:1990:SPE

[KCKC90] Patricia Kremer, John Costello, James Kremer, and Michael Canino. Significance of photosynthetic endosymbionts to the carbon budget of the scyphomedusa Linuche unguiculata. Limnology and Oceanography, 35(3):609–624, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Killworth:1996:MDW

[KCWM96] Peter D. Killworth, Eddy C. Carmack, Ray F. Weiss, and Richard Matear. Modeling deep-water renewal in Lake Baikal. Limnology and Oceanography, 41(7):1521–1538, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Kvitek:1991:PSP

[KDB91] Rikk G. Kvitek, Anthony R. DeGunge, and Mark K. Beitler. Paralytic shellfish poisoning toxins mediate feeding behavior of sea otters. Limnology and Oceanography, 36(2):393–404, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Keller:1990:GJA

[KDKS90] Aimee A. Keller, Peter H. Doering, Stephen P. Kelly, and Barbara K. Sullivan. Growth of juvenile Atlantic menhaden, Brevoortia tyrannus (Pisces: Clupeidae) in MERL mesocosms: Effects of eutrophication. Limnology and Oceanography, 35(1):109–122, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Krom:1994:HRP

[KDZD94] M. D. Krom, P. Davison, H. Zhang, and W. Davison. Highresolution pore-water sampling with a gel sampler. Limnology and Oceanography, 39(8):1967–1972, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Koenings:1991:SDP

[KE91] J. P. Koenings and J. A. Edmundson. Secchi disk and photometer estimates of light regimes in Alaskan lakes: Effects of yellow color and turbidity. Limnology and Oceanography, 36(1):91–105, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Kenney:1993:OCB

[Ken93] Bernard C. Kenney. Observations of coherent bands of algae in a surface shear layer. Limnology and Oceanography, 38(5):1059– 1067, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Kerner:1993:CMF

[Ker93] M. Kerner. Coupling of microbial fermentation and respiration processes in an intertidal mudflat of the Elbe estuary. Limnology and Oceanography, 38(2):314–330, March 1993. CODEN LI-OCAH. ISSN 0024-3590.

Kerfoot:1995:BRL

[Ker95] W. Charles Kerfoot. Bosmina remains in Lake Washington sediments: Qualitative heterogeneity of bay environments and quantitative correspondence to production. Limnology and Oceanography, 40(2):211–225, March 1995. CODEN LIOCAH. ISSN 0024-

Kolber:1993:UAF

[KF93] Zbigniew Kolber and Paul G. Falkowski. Use of active fluorescence to estimate phytoplankton photosynthesis in situ. Limnology and Oceanography, 38(8):1646–1665, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Kratz:1991:RRS

[KFEC91] Timothy K. Kratz, Thomas M. Frost, Joan E. Elias, and Robert B. Cook. Reconstruction of a regional, 12-yr silica decline in lakes by means of fossil sponge spicules. Limnology and Oceanography, 36 (6):1244–1249, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Kristiansen:1994:NCB

[KFW94] Svein Kristiansen, Tove Farbrot, and Patricia A. Wheeler. Nitrogen cycling in the Barents Sea — seasonal dynamics of new and regenerated production in the marginal ice zone. Limnology and Oceanography, 39(7):1630–1642, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Kemer:1995:COC

[KG95] M. Kemer and H. Gramm. Changes in oxygen consumption at the sediment–water interface formed by settling seston from the Elbe estuary. Limnology and Oceanography, 40(3):544–555, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Kling:1991:RST

[KGFP91] George W. Kling, Anne E. Giblin, Brian Fry, and Bruce J. Peterson. The role of seasonal turnover in lake alkalinity dynamics. Limnology and Oceanography, 36(1):106–122, January 1991. CO-DEN LIOCAH. ISSN 0024-3590.

Kreeger:1996:UDL

[KHB96] D. A. Kreeger, A. J. S. Hawkins, and B. L. Bayne. Use of dual-labeled microcapsules to discern the physiological fates of assimilated carbohydrate, protein carbon, and protein nitrogen in suspension-feeding organisms. Limnology and Oceanography, 41 (2):208–215, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Karl:1998:RDO

[KHBL98] David M. Karl, Dale V. Hebel, K. Björkman, and R. M. Letelier. The role of dissolved organic matter release in the productivity of the oligotrophic North Pacific Ocean. Limnology and Oceanography, 43(6):1270–1286, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Klimant:1997:SFS

[KHK97] Ingo Klimant, Gerhard Holst, and Michael Kûhl. A simple fiberoptic sensor to detect the penetration of microsensors into sediments and other biogeochemical systems. Limnology and Oceanography, 42(7):1638–1643, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Kononen:1998:DSC

[KHK^{+}98] Kaisa Kononen, Seija Hällfors, Marjaana Kokkonen, Harri Kuosa, Jaan Laanemets, Juss Pavelson, and Riitta Autio. Development of a subsurface chlorophyll maximum at the entrance to the Gulf of Finland, Baltic Sea. Limnology and Oceanography, 43(6):1089– 1106, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Kleppel:1991:TIB

[KHP91] G. S. Kleppel, D. V. Holliday, and R. E. Pieper. Trophic interactions between copepods and microplankton: a question about the role of diatoms. Limnology and Oceanography, 36(1):172–178, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Keough:1998:CPP

[KHRS98] Janet R. Keough, Cynthia A. Hagley, Elaine Ruzycki, and Michael Sierszen. δ ^{13}C composition of primary producers and role of detritus in a freshwater coastal ecosystem. Limnology and Oceanography, 43(4):734–740, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Kieber:1997:PTF

[KHS97] Robert J. Kieber, Linda H. Hydro, and Pamela J. Seaton. Photooxidation of triglycerides and fatty acids in seawater: Implication toward the formation of marine humic substances. Limnology and Oceanography, 42(6):1454–1462, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Kopacek:1998:RAM

[KHS^{+}98] Jiří Kopáček, Josef Hejzlar, Evžien Stuchlík, Jan Fott, and Josef Veselý. Reversibility of acidification of mountain lakes after reduction in nitrogen and sulphur emissions in Central Europe. Limnology and Oceanography, 43(2):357–361, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Kiefer:1990:RCB

[Kie90] Dale A. Kiefer. A reply to the comment by Bannister. Limnology and Oceanography, 35(4):980–982, June 1990. CODEN LIOCAH. ISSN 0024-3590. See [Ban90a].

Kilham:1990:MCC

[Kil90] Peter Kilham. Mechanisms controlling the chemical composition of lakes and rivers: Data from Africa. Limnology and Oceanography, 35(1):80–83, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Kimmerer:1990:SSP

[Kim90] W. J. Kimmerer. Simplicity in secondary production measurements: a reply to the comment by Polishchuk. Limnology and Oceanography, 35(7):1651–1655, November 1990. CODEN LI-OCAH. ISSN 0024-3590. See [Pol90].

Kim:1992:BRS

[Kim92a] Alison Kim. Book review: Shallow-water dictionary: a grounding in estuary English (J. R. Stilgoe). Limnology and Oceanography, 37(6):1340–1341, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Kim:1992:SCC

[Kim92b] Alison Kim. Sulphur cycling on the continents: Wetlands, terrestrial ecosystems, and associated water bodies. Limnology and Oceanography, 37(6):1341, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Kirk:1991:VSF

[Kir91] John T. O. Kirk. Volume scattering function, average cosines, and the underwater light field. Limnology and Oceanography, 36(3): 455–467, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Kirchman:1994:ECN

[Kir94a] David L. Kirchman. Editorial comment: Natural history of a manuscript, revisited. Limnology and Oceanography, 39(3):739– 741, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Kirk:1994:CLF

[Kir94b] John T. O. Kirk. Characteristics of the light field in highly turbid waters: a Monte Carlo study. Limnology and Oceanography, 39 (3):702–706, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Kirchman:1998:SPE

[Kir98] David L. Kirchman. Scholarly publishing: The electronic frontier. Limnology and Oceanography, 43(5):1017–1018, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Kirchman:1999:BBE

[Kir99] David L. Kirchman. Bacterial biogeochemistry: The ecophysiology of mineral cycling. Limnology and Oceanography, 44(1): 233–234, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Kitheka:1996:WCC

[Kit96] Johnson U. Kitheka. Water circulation and coastal trapping of brackish water in a tropical mangrove-dominated bay in Kenya. Limnology and Oceanography, 41(1):169–176, January 1996. CO-DEN LIOCAH. ISSN 0024-3590.

Kuhl:1992:SLM

[KJ92] Michael Kühl and Bo Barker Jørgensen. Spectral light measurements in microbenthic phototrophic communities with a fiberoptic microprobe coupled to a sensitive diode array detector. Limnology and Oceanography, 37(8):1813–1823, December 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Kuhl:1994:LFM

[KJ94] Michael Kühl and Bo Barker Jorgensen. The light field of microbenthic communities: Radiance distribution and microscale optics of sandy coastal sediments. Limnology and Oceanography, 39(6):1368–1398, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Krembs:1998:NPB

[KJLA98] Christopher Krembs, Andrew R. Juhll, Richard A. Long, and Farooq Azam. Nanoscale patchiness of bacteria in lake water studied with the spatial information preservation method. Limnology and Oceanography, 43(2):307–314, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Krembs:1998:SIP

[KJS98] Christopher Krembs, Andrew R. Juhl, and J. Rudi Strickler. The spatial information preservation method: Sampling the nanoscale spatial distribution of microorganisms. Limnology and Oceanography, 43(2):298–306, March 1998. CODEN LIOCAH. ISSN 0024-

Krause-Jensen:1998:LAP

[KJSJ98] Dorte Krause-Jensen and Kaj Sand-Jensen. Light attenuation and photosynthesis of aquatic plant communities. Limnology and Oceanography, 43(3):396–407, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Kerfoot:1991:DTD

[KK91] W. Charles Kerfoot and Kevin L. Kirk. Degree of taste discrimination among suspension-feeding cladocerans and copepods: Implications for detritivory and herbivory. Limnology and Oceanography, 36(6):1107–1123, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Keil:1993:DCA

[KK93] Richard G. Keil and David L. Kirchman. Dissolved combined amino acids: Chemical form and utilization by marine bacteria. Limnology and Oceanography, 38(6):1256–1270, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Krom:1991:PLP

[KKBG91] M. D. Krom, N. Kress, S. Brenner, and L. I. Gordon. Phosphorus limitation of primary productivity in the eastern Mediterranean Sea. Limnology and Oceanography, 36(3):424–432, May 1991. CO-DEN LIOCAH. ISSN 0024-3590.

Kivi:1993:NLG

[KKK^{+}93] K. Kivi, S. Kaitala, H. Kuosa, J. Kuparinen, E. Leskinen, R. Lignell, B. Marcussen, and T. Tamminen. Nutrient limitation and grazing control of the Baltic plankton community during annual succession. Limnology and Oceanography, 38(5):893–905, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Kononen:1996:ICB

[KKM^{+}96] Kaisa Kononen, Jorma Kuparinen, Kalervo Mäkelä, Jaan Laanemets, Juss Pavelson, and Sulev Nõmmann. Initiation of cyanobacterial blooms in a frontal region at the entrance to the Gulf of Finland, Baltic Sea. Limnology and Oceanography, 41(1): 98–112, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Kolodny:1999:CTS

[KKS^{+}99] Yehoshua Kolodny, Amitai Katz, Abraham Starinsky, Tamar Moise, and Ehud Simon. Chemical tracing of salinity sources in Lake Kinneret (Sea of Galilee), Israel. Limnology and Oceanography, 44(4):1035–1044, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Kirchman:1990:CLA

[KKW90] David L. Kirchman, Richard G. Keil, and Patricia A. Wheeler. Carbon limitation of ammonium uptake by heterotrophic bacteria in the subarctic Pacific. Limnology and Oceanography, 35(6):1258– 1266, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Karentz:1990:EOH

[KL90] Deneb Karentz and Louise H. Lutze. Evaluation of biologically harmful ultraviolet radiation in Antarctica with a biological dosimeter designed for aquatic environments. Limnology and Oceanography, 35(3):549–561, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Kao:1996:POC

[KL96] Shuh-Ji Kao and Kon-Kee Liu. Particulate organic carbon export from a subtropical mountainous river (Lanyang Hsi) in Taiwan. Limnology and Oceanography, 41(8):1749–1757, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Klap:1999:DDS

[KLB^{+}99] Vincent A. Klap, Patrick Louchouarn, Jaap J. Boon, Marten A. Hemminga, and Jos van Soelen. Decomposition dynamics of six salt marsh halophytes as determined by cupric oxide oxidation and direct temperature-resolved mass spectrometry. Limnology and Oceanography, 44(6):1458–1476, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Klick:1992:SVB

[Kli92] Silke Klick. Seasonal variations of biogenic and anthropogenic halocarbons in seawater from a coastal site. Limnology and Oceanography, 37(7):1579–1585, November 1992. CODEN LI-OCAH. ISSN 0024-3590.

Kerfoot:1994:PSC

[KLR94] W. Charles Kerfoot, George Lauster, and John A. Robbins. Paleolimnological study of copper mining around Lake Superior: Artificial varves from Portage Lake provide a high resolution record. Limnology and Oceanography, 39(3):649–669, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Karrh:1994:FRS

[KM94] Renee R. Karrh and Douglas C. Miller. Functional response of a surface-deposit feeder, Saccoglossus kowalewkii. Limnology and Oceanography, 39(6):1455–1464, September 1994. CODEN LI-OCAH. ISSN 0024-3590.

King:1995:DVM

[KM95] Charles E. King and Maria R. Miracle. Diel vertical migration by Daphnia longispina in a Spanish lake: Genetic sources of distributional variation. Limnology and Oceanography, 40(2):226–231, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Kelley:1990:VSC

[KMC90] Cheryl A. Kelley, Christopher S. Martens, and Jefrey P. Chanton. Variations in sedimentary carbon remineralization rates in the White Oak River estuary, North Carolina. Limnology and Oceanography, 35(2):372–383, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Klimant:1995:FOO

[KMK95] Ingo Klimant, Volker Meyer, and Michael Kühl. Fiber-optic oxygen microsensors, a new tool in aquatic biology. Limnology and Oceanography, 40(6):1159–1165, September 1995. CODEN LI-OCAH. ISSN 0024-3590.

Koch:1990:MHS

[KMM90] Marguerite S. Koch, Irving A. Mendelssohn, and Karen L. McKee. Mechanism for the hydrogen sulfide-induced growth limitation in wetland macrophytes. Limnology and Oceanography, 35(2):399– 408, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Kelley:1995:MDA

[KMU95] Cheryl A. Kelley, Christopher S. Martens, and William Ussler III. Methane dynamics across a tidally flooded riverbank margin. Limnology and Oceanography, 40(6):1112–1129, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Kator:1992:RCM

[KMWK92] Howard Kator, Lori Jedlica Morris, Richard L. Wetzel, and Eric T. Koepfler. A rapid chromatographic method for recovery of ^{15}NO_{2} and NO_{3}^{−} produced by nitrification in aqueous samples. Limnology and Oceanography, 37(4):900–907, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Knauer:1999:LEA

[Kna99] Douglas Knauer. Limnological and engineering analysis of a polluted urban lake: Prelude to environmental management of Onondaga Lake, New York. Limnology and Oceanography, 44(2): 476, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Kuma:1996:CII

[KNM96] Kenshi Kuma, Jun Nishioka, and Katsuhiko Matsunaga. Controls on iron(III) hydroxide solubility in seawater: The influence of pH and natural organic chelators. Limnology and Oceanography, 41 (3):396–407, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Keller:1999:PIE

[KOWH99] Aimee A. Keller, Candace A. Oviatt, Henry A. Walker, and Jamie D. Hawk. Predicted impacts of elevated temperature on the magnitude of the winter-spring phytoplankton bloom in temperate coastal waters: a mesocosm study. Limnology and Oceanography, 44(2):344–356, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Kopacek:1995:NPR

[KPSB95] Jiří Kopáçek, Lidmila Procházková, Evžen Stuchlík, and Pavel Blažka. The nitrogen phosphorus relationship in mountain lakes: Influence of atmospheric input, watershed, and pH. Limnology and Oceanography, 40(5):930–937, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Komada:1998:DIC

[KRB98] Tomoko Komada, Glare E. Reimers, and Susan E. Boehme. Dissolved inorganic carbon profiles and fluxes determined using pH and. Limnology and Oceanography, 43(5):769–781, July 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Kroer:1993:BGE

[Kro93] Niels Kroer. Bacterial growth efficiency on natural dissolved organic matter. Limnology and Oceanography, 38(6):1282–1290, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Kerfoot:1999:NAH

[KRW99] W. Charles Kerfoot, John A. Robbins, and Lawrence J. Weider. A new approach to historical reconstruction: Combining descriptive and experimental paleolimnology. Limnology and Oceanography, 44(5):1232–1247, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Kuhn:1993:ACE

[KS93] Annette Kuhn and Laura Sigg. Arsenic cycling in eutrophic Lake Greifen, Switzerland: Influence of seasonal redox processes. Limnology and Oceanography, 38(5):1052–1059, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Kemp:1990:ARV

[KSC^{+}90] W. M. Kemp, P. Sampou, J. Caffrey, M. Mayer, K. Henriksen, and W. R. Boynton. Ammonium recycling versus denitrification in Chesapeake Bay sediments. Limnology and Oceanography, 35 (7):1545–1563, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Kunu:1998:DES

[KSCG98] Todd M. Kunu, Matthew B. Sullivan, Jeffrey C. Cornwell, and Kevin M. Groxzkowski. Denitrification in estuarine sediments determined by membrane inlet mass spectrometry. Limnology and Oceanography, 43(2):334–339, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Keough:1996:ALS

[KSH96] Janet R. Keough, Michael E. Sierszen, and Cynthia A. Hagley. Analysis of a Lake Superior coastal food web with stable isotope techniques. Limnology and Oceanography, 41(1):136–146, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Kitchell:1999:NCL

[KSH^{+}99] James F. Kitchell, Daniel E. Schindler, Brian R. Herwig, David M. Post, Mark H. Olson, and Michael Oldham. Nutrient cycling at the landscape scale: The role of diel foraging migrations by geese at the bosque del Apache national wildlife refuge, New Mexico. Limnology and Oceanography, 44(3 (part 2)):828–836, May 1999. CODEN LIOCAH. ISSN 0024-3590.

King:1991:DSB

[KSSP91] C. Harold King, Robert W. Sanders, Emmett B. Shot, Jr., and Karen G. Porter. Differential survival of bacteria ingested by zooplankton from a stratified eutrophic lake. Limnology and Oceanography, 36(5):829–844, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Kraines:1996:SBP

[KSYK96] Steven Kraines, Yoshimi Suzuki, Koichi Yamada, and Hiroshi Komiyama. Separating biological and physical changes in dissolved oxygen concentration in a coral reef. Limnology and Oceanography, 41(8):1790–1799, December 1996. CODEN LI-OCAH. ISSN 0024-3590.

Karl:1992:MSP

[KT92] David M. Karl and Georgia Tien. MAGIC: a sensitive and precise method for measuring dissolved phosphorus in aquatic environments. Limnology and Oceanography, 37(1):105–116, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Kilham:1996:LPD

[KTF96] Susan S. Kilham, Edward C. Theriot, and Sherilyn C. Fritz. Linking planktonic diatoms and climate change in the large lakes of the Yellowstone ecosystem using resource theory. Limnology and Oceanography, 41(5):1052–1062, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Kiorboe:1998:IAF

[KTMI^{+}98] Thomas Kiørboe, Peter Tiselius, Betty Mitchell-Innes, Jørgen L. S. Hansen, Andre W. Visser, and Xavier Mari. Intensive aggregate formation with low vertical flux during an upwelling-induced diatom bloom. Limnology and Oceanography, 43(1):104–116, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Kirschner:1999:BBS

[KV99] A. K. T. Kirschner and B. Velimirov. Benthic bacterial secondary production measured via simultaneous ^{3}H-thymidine and ^{14}Cleucine incorporation, and its implication for the carbon cycle of a shallow macrophyte-dominated backwater system. Limnology and Oceanography, 44(8):1871–1881, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Kruger:1993:SPS

[KW93] Clinton D. Kruger and Sarah Ann Woodin. Spatial persistence and sediment disturbance of an arenicolid polychaete. Limnology and Oceanography, 38(3):509–520, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Kiene:1998:GBU

[KW98a] Ronald P. Kiene and Lynn P. Hoffmann Williams. Glycine betaine uptake, retention, and degradation by microorganisms in seawater. Limnology and Oceanography, 43(7):1592–1603, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Knowles:1998:SAA

[KW98b] S. C. Knowles and J. T. Wells. In situ aggregate analysis camera (ISAAC): a quantitative tool for analyzing fine-grained suspended material. Limnology and Oceanography, 43(8):1954–1962, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Kolber:1990:NVP

[KWF90] Zbigniew Kolber, Kevin V. Wyman, and Paul G. Falkowski. Natural variability in photosynthetic energy conversion efficiency: a field study in the Gulf of Maine. Limnology and Oceanography, 35(1):72–79, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Kepner:1998:VAL

[KWS98] Raymond L. Kepner, Jr., Robert A. Wharton, Jr., and Curtis A. Suttle. Viruses in Antarctic lakes. Limnology and Oceanography, 43(7):1754–1761, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Kamykowski:1997:SMI

[KY97a] Daniel Kamykowski and Hidekatsu Yamazaki. A study of metabolism-influenced orientation in the diel vertical migration of marine dinoflagellates. Limnology and Oceanography, 42(5 (part 2)):1189–1202, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Karl:1997:PCD

[KY97b] David M. Karl and Katsumi Yanagi. Partial characterization of the dissolved organic phosphorus pool in the oligotrophic North Pacific Ocean. Limnology and Oceanography, 42(6):1398–1405, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Kieber:1997:FFD

[KYM97] David J. Kieber, Brian H. Yocis, and Kenneth Mopper. Freefloating drifter for photochemical studies in the water column. Limnology and Oceanography, 42(8):1829–1833, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Kitchen:1992:TLS

[KZ92] James C. Kitchen and J. Ronald V. Zaneveld. A three-layered sphere model of the optical properties of phytoplankton. Limnology and Oceanography, 37(8):1680–1690, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Kieber:1990:FCC

[KZM90] Robert J. Kieber, Xianliang Zhou, and Kenneth Mopper. Formation of carbonyl compounds from UV-induced photodegradation of humic substances in natural waters: Fate of riverine carbon in the sea. Limnology and Oceanography, 35(7):1503–1515, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Lyche:1996:MTSa

[LAC^{+}96a] Anne Lyche, Tom Andersen, Kirsten Christoffersen, Dag O. Hessen, Paul H. Berger Hansen, and Annette Klysner. Mesocosm tracer studies. 1. Zooplankton as sources and sinks in the pelagic phosphorus cycle of a mesotrophic lake. Limnology and Oceanography, 41(3):460–474, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Lyche:1996:MTSb

[LAC^{+}96b] Anne Lyche, Tom Andersen, Kirsten Christoffersen, Dag O. Hessen, Paul H. Berger Hansen, and Annette Klysner. Mesocosm tracer studies. 2. The fate of primary production and the role of consumers in the pelagic carbon cycle of a mesotrophic lake. Limnology and Oceanography, 41(3):475–487, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Lampert:1994:PPF

[Lam94] Winfried Lampert. Phenotypic plasticity of the filter screens in Daphnia: Adaptation to a low-food environment. Limnology and Oceanography, 39(5):997–1006, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Landry:1993:PER

[Lan93] Michael R. Landry. Predicting excretion rates of microzooplankton from carbon metabolism and elemental ratios. Limnology and Oceanography, 38(2):468–472, March 1993. CODEN LIOCAH. ISSN 0024-3590. See reply [CG93].

Lapointe:1995:CNL

[Lap95] Brian E. Lapointe. A comparison of nutrient-limited productivity in Sargassum natans from neritic vs. oceanic waters of the western North Atlantic Ocean. Limnology and Oceanography, 40(3):625– 633, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Lapointe:1997:NTB

[Lap97] Brian E. Lapointe. Nutrient thresholds for bottom-up control of macroalgal blooms on coral reefs in Jamaica and southeast Florida. Limnology and Oceanography, 42(5 (part 2)):1119–1131, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Lapointe:1999:STB

[Lap99] Brian E. Lapointe. Simultaneous top-down and bottom-up forces control macroalgal blooms on coral reefs (reply to the comment by Hughes et al.). Limnology and Oceanography, 44(6):1586– 1592, September 1999. CODEN LIOCAH. ISSN 0024-3590. See [HSS^{+}99].

Lawrence:1995:NDS

[LAYE95] Gregory A. Lawrence, Ken I. Ashley, Noboru Yonemitsu, and Joanne R. Ellis. Natural dispersion in a small lake. Limnology and Oceanography, 40(8):1519–1526, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Lundstedt:1991:DGR

[LB91] Lisa Lundstedt and Michael T. Brett. Differential growth rates of three cladoceran species in response to mono- and mixed-algal cultures. Limnology and Oceanography, 36(1):159–165, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1994:NDT

[LB94] John T. Lehman and Dorm K. Branstrator. Nutrient dynamics and turnover rates of phosphate and sulfate in Lake Victoria, East Africa. Limnology and Oceanography, 39(2):227–233, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Lustwerk:1995:EDS

[LB95] Rigel L. Lustwerk and David J. Burdige. Elimination of dissolved sulfide interference in the flow injection determination of SCO_{2}, by addition of molybdate. Limnology and Oceanography, 41(4): 1011–1012, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Lampert:1996:SPL

[LB96] Winfried Lampert and Heinz Brendelberger. Strategies of phenotypic low-food adaptation in Daphnia: Filter screens, mesh sizes, and appendage beat rates. Limnology and Oceanography, 41(2): 216–223, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Landry:1997:IGC

[LBB^{+}97] Michael R. Landry, Richard T. Barber, Robert R. Bidigare, Fei Chai, Kenneth H. Coale, Hans G. Dam, Marlon R. Lewis, Steven T. Lindley, James J. McCarthy, Michael R. Roman, Diane K. Stoecker, Peter G. Verity, and Jacques R. White. Iron and grazing constraints on primary production in the central equatorial Pacific: an EqPac synthesis. Limnology and Oceanography, 42 (3):405–418, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Letelier:1993:TVP

[LBH^{+}93] R. M. Letelier, R. R. Bidigare, D. V. Hebel, M. Ondrusek, C. D. Winn, and D. M. Karl. Temporal variability of phytoplankton community structure based on pigment analysis. Limnology and Oceanography, 38(7):1420–1437, November 1993. CODEN LI-OCAH. ISSN 0024-3590.

Lee:1991:SVM

[LBL91] T. Lee, E. Barg, and D. Lal. Studies of vertical mixing in the Southern California Bight with cosmogenic radionuclides ^{32}P and 7Be. Limnology and Oceanography, 36(5):1044–1052, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Luther:1998:SMM

[LBL^{+}98] George W. Luther III, Paul J. Brendel, Brent L. Lewis, Bjørn Sundby, Lucie Lefrançois, Norman Silverberg, and Donald B. Nuzzio. Simultaneous measurement of O_{2}, Mn, Fe, I^{−}, and S(-II) in marine pore waters with a solid-state voltammetric microelectrode. Limnology and Oceanography, 43(2):325–333, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Landing:1991:NCB

[LBLO91] William M. Landing, William C. Burnett, W. Berry Lyons, and William H. Orem. Nutrient cycling and the biogeochemistry of manganese, iron. and zinc in Jellyfish Lake, Palau. Limnology and Oceanography, 36(3):515–525, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Laws:1997:EGR

[LBP97] Edward A. Laws, Robert R. Bidigare, and Brian N. Popp. Effect of growth rate and CO_{2} concentration on carbon isotopic fractionation by the marine diatom Phaeodactylum tricornutum. Limnology and Oceanography, 42(7):1552–1560, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Laws:1990:MGM

[LC90a] Edward A. Laws and Michael S. Chalup. A microalgal growth model. Limnology and Oceanography, 35(3):597–608, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Leavitt:1990:APD

[LC90b] P. R. Leavitt and S. R. Carpenter. Aphotic pigment degradation in the hypolimnion: Implications for sedimentation studies and paleolimnology. Limnology and Oceanography, 35(2):520–534, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1993:FWR

[LC93] John T. Lehman and Carla E. Cáceres. Food-web responses to species invasion by a predatory invertebrate: Bythotrephes in Lake Michigan. Limnology and Oceanography, 38(4):879–891, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Lutz:1994:EBL

[LCG94] Michael I. Lutz, James F. Case, and Robert L. Gran. Excitation of bioluminescence by laminar fluid shear associated with simple Couette flow. Limnology and Oceanography, 39(6):1424–1439, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Lathrop:1999:SWC

[LCR99] Richard C. Lathrop, Stephen R. Carpenter, and Dale M. Robertson. Summer water clarity responses to phosphorus, Daphnia grazing, and internal mixing in Lake Mendota. Limnology and Oceanography, 44(1):137–146, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Ledyard:1996:MCD

[LD96] Kathleen M. Ledyard and John W. H. Dacey. Microbial cycling of DMSP and DMS in coastal and oligotrophic seawater. Limnology and Oceanography, 41(1):33–40, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Lee:1999:INA

[LD99] Kun-Seop Lee and Kenneth H. Dunton. Inorganic nitrogen acquisition in the seagrass Thalassia testudinum: Development of a whole-plant nitrogen budget. Limnology and Oceanography, 44 (5):1204–1215, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Lim:1999:EPI

[LDC99] Ee Lin Lim, Mark R. Dennett, and David A. Caron. The ecology of Paraphysomonas imperforata based on studies employing oligonucleotide probe identification in coastal water samples and enrichment cultures. Limnology and Oceanography, 44(1):37–51, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Lind:1991:ATO

[LDL91] Owen T. Lind and Laura Dávalos-Lind. Association of turbidity and organic carbon with bacterial abundance and cell size in a large, turbid, tropical lake. Limnology and Oceanography, 36(6): 1200–1208, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Lind:1992:CTR

[LDV^{+}92] Owen T. Lind, Robert Doyle, Darrell S. Vodopich, Bruce G. Trotter, J. Gualberto Limon, and Laura Dcivalos-Lind. Clay turbidity: Regulation of phytoplankton production in a large, nutrient-rich tropical lake. Limnology and Oceanography, 37(3):549–565, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Lebo:1990:PUA

[Leb90] Martin E. Lebo. Phosphate uptake along a coastal plain estuary. Limnology and Oceanography, 35(6):1279–1289, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1991:IGL

[Leh91] John T. Lehman. Interacting growth and loss rates: The balance of top-down and bottom-up controls in plankton communities. Limnology and Oceanography, 36(8):1546–1554, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1993:EIA

[Leh93] John T. Lehman. Efficiencies of ingestion and assimilation by an invertebrate predator using C and P dual isotope labeling. Limnology and Oceanography, 38(7):1550–1554, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1995:BRC

[Leh95a] John T. Lehman. Book review: Chrysophyte algae: Ecology, phylogeny and development (C. D. Sandgren, J. P. Smol, and J. Kristiansen [eds.]). Limnology and Oceanography, 40(8):1535, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1995:BRL

[Leh95b] John T. Lehman. Book review: Limnology now: a paradigm of planetary problems (R. Margalef [ed.]). Limnology and Oceanography, 40(3):644, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Leibold:1990:RPC

[Lei90] Mathew A. Leibold. Resources and predators can affect the vertical distributions of zooplankton. Limnology and Oceanography, 35(4):938–944, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Lesack:1995:SEA

[Les95] Lance F. W. Lesack. Seepage exchange in an Amazon floodplain lake. Limnology and Oceanography, 40(3):598–609, May 1995. CO-DEN LIOCAH. ISSN 0024-3590.

Lesser:1996:ETU

[Les96] Michael P. Lesser. Elevated temperatures and ultraviolet radiation cause oxidative stress and inhibit photosynthesis in symbiotic dinoflagellates. Limnology and Oceanography, 41(2):271–283, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Levin:1996:RTD

[Lev96] Phillip S. Levin. Recruitment in a temperate demersal fish: Does larval supply matter? Limnology and Oceanography, 41(4):672– 679, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1990:CPB

[Lew90] William M. Lewis, Jr. Comparisons of phytoplankton biomass in temperate and tropical lakes. Limnology and Oceanography, 35(8): 1838–1845, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1995:BRS

[Lew95] William M. Lewis, Jr. Book review: Science and limnology (E. H. Rigler and R. H. Peters). Limnology and Oceanography, 40(7): 1344, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1998:FBP

[Lew98] William M. Lewis, Jr. Freshwater biology, priorities and development in Danish research. Limnology and Oceanography, 43(5): 1019–1020, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Lee:1991:STV

[LF91] SangHoon Lee and Jed A. Fuhrman. Spatial and temporal variation of natural bacterioplankton assemblages studied by total genomic DNA cross-hybridization. Limnology and Oceanography, 36(7):1277–1287, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Lee:1993:MMC

[LF93] Byeong-Gweon Lee and Nicholas S. Fisher. Microbially mediated cobalt oxidation in seawater revealed by radiotracer experiments. Limnology and Oceanography, 38(8):1593–1602, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Laird:1996:CSP

[LFGM96] Kathleen R. Laird, Sherilyn C. Fritz, Eric C. Grimm, and Pietra G. Mueller. Century scale paleoclimatic reconstruction from Moon Lake, a closed-basin lake in the northern Great Plains. Limnology and Oceanography, 41(5):890–902, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Leavitt:1999:ARD

[LFHS99] P. R. Leavitt, D. L. Findlay, R. I. Hall, and J. P. Smol. Algal responses to dissolved organic carbon loss and pH decline during whole-lake acidification: Evidence from paleolimnology. Limnology and Oceanography, 44(3 (part 2)):757–773, May 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Lomas:1999:TRN

[LG99] Michael W. Lomas and Patricia M. Glibert. Temperature regulation of nitrate uptake: a novel hypothesis about nitrate uptake and reduction in cool-water diatoms. Limnology and Oceanography, 44(3):556–572, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Leonard:1994:DSD

[LGI94] Muscatine Leonard, Ruth D. Gates, and LaFontaine Ingrid. Do symbiotic dinoflagellates secrete lipid droplets? Limnology and Oceanography, 39(4):925–929, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Lindell:1995:EBG

[LGT95] Måns J. Lindell, Wilhelm Granéli, and Lars J. Tranvik. Enhanced bacterial growth in response to photochemical transformation of dissolved organic matter. Limnology and Oceanography, 40(1): 195–199, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Lynch:1991:ADR

[LGW91] Michael Lynch, Wilfred Gabriel, and A. Michelle Wood. Adaptive and demographic responses of plankton populations to environmental change. Limnology and Oceanography, 36(7):1301–1312, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Lindquist:1991:SWN

[LH91] Niels Lindquist and Mark E. Hay. A secret world — natural products of marine life. Limnology and Oceanography, 36(5):1068, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Levin:1993:RET

[LHM^{+}93] Lisa A. Levin, Douglas Huggett, Philip Myers, Todd Bridges, and Jack Weaver. Rare earth tagging methods for the study of larval dispersal by marine invertebrates. Limnology and Oceanography, 38(2):346–360, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Lohner:1990:MDS

[LHS90a] Linda M. Lohner, Nelson G. Hairston, Jr., and William R. Schaflner. A method for distinguishing subitaneous and diapausing eggs in preserved samples of the Calanoid copepod genus Diaptomus. Limnology and Oceanography, 35(3):763–767, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Luettich:1990:DBS

[LHS90b] Richard A. Luettich, Jr., Donald R. F. Harleman, and László Somlyódy. Dynamic behavior of suspended sediment concentrations in a shallow lake perturbed by episodic wind events. Limnology and Oceanography, 35(5):1050–1067, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Lee:1992:EVT

[LHWZ92] Cindy Lee, John I. Hedges, Stuart G. Wakeham, and Ningli Zhu. Effectiveness of various treatments in retarding microbial activity in sediment trap material and their effects on the collection of swimmers. Limnology and Oceanography, 37(1):117–130, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Li:1990:BTA

[Li90] W. K. W. Li. Bivariate and trivariate analysis in flow cytometry: Phytoplankton size and fluorescence. Limnology and Oceanography, 35(6):1356–1368, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Li:1994:PPP

[Li94] William K. W. Li. Primary production of prochlorophytes, cyanobacteria, and eucaryotic ultraphytoplankton: Measurements from flow cytometric sorting. Limnology and Oceanography, 39(1): 169–175, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Li:1997:CDM

[Li97] W. K. W. Li. Cytometric diversity in marine ultraphytoplankton. Limnology and Oceanography, 42(5):874–880, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Li:1998:AAA

[Li98] W. K. W. Li. Annual average abundance of heterotrophic bacteria and Synechococcus in surface ocean waters. Limnology and Oceanography, 43(7):1746–1753, November 1998. CODEN LI-OCAH. ISSN 0024-3590.

Li:1993:DFV

[LID93] W. K. W. Li, B. D. Irwin, and P. M. Dickie. Dark fixation of ^{14}C: Variations related to biomass and productivity of phytoplankton and bacteria. Limnology and Oceanography, 38(3):483–494, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Lignell:1992:PFF

[Lig92] Risto Lignell. Problems in filtration fractionation of ^{14}C primary productivity samples. Limnology and Oceanography, 37(1):172– 178, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:1990:DSC

[Liv90] D. A. Livingstone. The design of sediment coring tubes and its relevance to “Laboratory studies on core sampling with application to subtidal meiobenthos collection” (rutledge and fleeger). Limnology and Oceanography, 35(2):543–545, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:1999:IBS

[Liv99] David M. Livingstone. Ice break-up on southern Lake Baikal and its relationship to local and regional air temperatures in Siberia and to the North Atlantic Oscillation. Limnology and Oceanography, 44(6):1486–1497, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Lindegarth:1991:FMN

[LJA91] Mats Lindegarth, Per R. Jonsson, and Carl André. Fluorescent microparticles: a new way of visualizing sedimentation and larval settlement. Limnology and Oceanography, 36(7):1471–1476, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Leclerc:1999:DIW

[LJBTB99] A.-M. Leclerc, P. Jean-Baptiste, D. Texier, and D. Broc. Densityinduced water circulations in atoll coral reefs: a numerical study. Limnology and Oceanography, 44(5):1268–1281, July 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Li:1995:DDP

[LJD95] W. K. W. Li, J. F. Jellett, and P. M. Dickie. DNA distributions in planktonic bacteria stained with TOTO or TO-PRO. Limnology and Oceanography, 40(8):1485–1495, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Levy:1991:SFV

[LJH91] David A. Levy, Robert L. Johnson, and Jeremy M. Hume. Shifts in fish vertical distribution in response to an internal seiche in a stratified lake. Limnology and Oceanography, 36(1):187–192, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Lee:1994:SCR

[LK94a] SangHoon Lee and Paul F. Kemp. Single-cell RNA content of natural marine planktonic bacteria measured by hybridization with multiple ^{16}S rRNA-targeted fluorescent probes. Limnology and Oceanography, 39(4):869–879, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Lewitus:1994:REP

[LK94b] Alan J. Lewitus and Todd M. Kana. Responses of estuarine phytoplankton to exogenous glucose: Stimulation versus inhibition of photosynthesis and respiration. Limnology and Oceanography, 39 (1):182–189, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Lafferty:1999:HES

[LK99] Kevin D. Lafferty and Armand M. Kuris. How environmental stress affects the impacts of parasites. Limnology and Oceanography, 44(3 (part 2)):925–931, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Lee:1999:OSS

[LKD99] Raymond W. Lee, David W. Kraus, and Jeannette E. Doeller. Oxidation of sulfide by Spartina alterniflora roots. Limnology and Oceanography, 46(4):1155–1159, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Lewitus:1998:SVR

[LKM98] Alan J. Lewitus, Eric T. Koepfler, and James T. Morris. Seasonal variation in the regulation of phytoplankton by nitrogen and grazing in a salt-marsh estuary. Limnology and Oceanography, 43(4): 636–646, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Li:1997:SIS

[LKSA97] Hong-Chun Li, Teh-Lung Ku, Lowell D. Stott, and Robert F. Anderson. Stable isotope studies on Mono Lake (California). 1. d18 in lake sediments as proxy for climatic change during the last 150 years. Limnology and Oceanography, 42(2):230–238, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Lunte:1990:TIL

[LL90] Cynthia C. Lunte and Chris Luecke. Trophic interactions of leptodora in Lake Mendota. Limnology and Oceanography, 35(5): 1091–1100, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Lazar:1991:BCR

[LL91] Boaz Lazar and Yossi Loya. Bioerosion of coral reefs — a chemical approach. Limnology and Oceanography, 36(2):377–383, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Leonard:1995:FHT

[LL95] Lynn A. Leonard and Mark E. Luther. Flow hydrodynamics in tidal marsh canopies. Limnology and Oceanography, 40(8):1474– 1484, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Lauridsen:1996:ADM

[LL96] Torben L. Lauridsen and David M. Lodge. Avoidance by Daphnia magna of fish and macrophytes: Chemical cues and predatormediated use of macrophyte habitat. Limnology and Oceanography, 41(8):794–798, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Lee:1998:IMB

[LL98] Byeong-Gweon Lee and Samuel N. Louma. Influence of microalgal biomass on absorption efficiency of Cd, Cr, and Zn by two bivalves from San Francisco Bay. Limnology and Oceanography, 43(7): 1455–1466, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Lyons:1996:JLP

[LLB^{+}96] W. Berry Lyons, Robert M. Lent, William C. Burnett, Philip Chin, William M. Landing, William H. Orem, and John M. McArthur. Jellyfish Lake, Palau: Regeneration of C, N, Si, and P in anoxic marine lake sediments. Limnology and Oceanography, 41(7):1394–1403, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Latasa:1997:PSG

[LLLB97] Mikel Latasa, Michael R. Landry, Schlüter Louise, and Robert R. Bidigare. Pigment specific growth and grazing rates of phytoplankton in the central equatorial Pacific. Limnology and Oceanography, 42(2):289–298, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Lee:1999:CMC

[LLZ99] Eunjung S. Lee, Alan J. Lewitus, and Richard K. Zimmer. Chemoreception in a marine cryptophyte: Behavioral plasticity in response to amino acids and nitrate. Limnology and Oceanography, 44(6):1571–1574, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Leff:1991:BAD

[LM91] Laura G. Leff and Judy L. Meyer. Biological availability of dissolved organic carbon along the Ogeechee River. Limnology and Oceanography, 36(2):315–323, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Leathers:1998:OSI

[LM98a] R. A. Leathers and N. J. McCormick. Ocean scalar irradiance near-surface maxima. Limnology and Oceanography, 43(5):982– 986, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Loisel:1998:LSC

[LM98b] Hubert Loisel and André Morel. Light scattering and chlorophyll concentration in case 1 waters: a reexamination. Limnology and Oceanography, 43(5):847–858, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Lesack:1998:STD

[LMH98] Lance F. W. Lesack, Philip Marsh, and Robert E. Hecky. Spatial and temporal dynamics of major solute chemistry among Mackenzie Delta lakes. Limnology and Oceanography, 43(7):1530–1543, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Lenihan:1999:IME

[LMSP99] Hunter S. Lenihan, Fiorenza Micheli, Stephen W. Shelton, and Charles H. Peterson. The influence of multiple environmental stressors on susceptibility to parasites: an experimental determination with oysters. Limnology and Oceanography, 44(3 (part 2)): 910–924, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Lundsgaard:1997:OSD

[LO97] Claus Lundsgaard and Michael Olesen. The origin of sedimenting detrital matter in a coastal system. Limnology and Oceanography, 42(5):1001–1005, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Logan:1993:TAS

[Log93] Bruce E. Logan. Theoretical analysis of size distributions determined with screens and filters. Limnology and Oceanography, 38 (2):372–381, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Logan:1995:BRC

[Log95] Bruce E. Logan. Book review: Chemical fate and transport in the environment (H. F. Hemond and E. J. Fechner). Limnology and Oceanography, 40(8):1534–1535, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Longhurst:1991:RMB

[Lon91] Alan R. Longhurst. Role of the marine biosphere in the global carbon cycle. Limnology and Oceanography, 36(8):1507–1526, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Loose:1993:LER

[Loo93] Carsten J. Loose. Lack of endogenous rhythmicity in Daphnia diel vertical migration. Limnology and Oceanography, 38(8):1837– 1841, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Loudon:1990:ETF

[Lou90] Catherine Loudon. Empirical test of filtering theory: Particle capture by rectangular-mesh nets. Limnology and Oceanography, 35(1):143–148, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Lizottel:1994:NFQ

[LP94] Michael P. Lizottel and John C. Prism. Natural fluorescence and quantum yields in vertically stationary phytoplankton from perennially ice-covered lakes. Limnology and Oceanography, 39(6):1399– 1410, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Lindell:1995:UST

[LP95] Debbie Lindell and Anton F. Post. Ultraphytoplankton succession is triggered by deep winter mixing in the Gulf of Aqaba (Eilat), Red Sea. Limnology and Oceanography, 40(6):1130–1141, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Letarte:1991:RBB

[LPA91] Yvon Letarte and Bernadette Pine-Alloul. Relationships between bacterioplankton production and limnological variables: Necessity of bacterial size considerations. Limnology and Oceanography, 36 (6):1208–1216, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Logan:1994:VRD

[LPA94] Bruce E. Logan, Uta Passow, and Alice L. Alldredge. Variable retention of diatoms on screens during size separations. Limnology and Oceanography, 39(2):390–395, March 1994. CODEN LI-OCAH. ISSN 0024-3590.

Lenihan:1996:DFS

[LPA96] Hunter S. Lenihan, Charles H. Peterson, and Jennifer M. Allen. Does flow speed also have a direct effect on growth of active suspension-feeders: an experimental test on oysters. Limnology and Oceanography, 41(6):1359–1366, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Lassen:1992:MPS

[LPJ92] Carsten Lassen, Helle Ploug, and Bo Barker Jorgensen. Microalgal photosynthesis and spectral scalar irradiance in coastal marine sediments of limfjorden, Denmark. Limnology and Oceanography, 37(4):760–772, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Laybourn-Parry:1997:MCM

[LPPSR97] Johanna Laybourn-Parry, Stephen J. Perriss, George G. R. Seaton, and Jan Rohozinski. A mixotrophic ciliate as a major contributor to plankton photosynthesis in Australian lakes. Limnology and Oceanography, 42(6):1463–1467, September 1997. CO-DEN LIOCAH. ISSN 0024-3590.

Latz:1999:LRR

[LR99] Michael I. Latz and Jim Rohr. Luminescent response of the red tide dinoflagellate Lingulodinium polyedrum to laminar and turbulent flow. Limnology and Oceanography, 44(6):1423–1435, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Lee:1995:CNM

[LRM95] Jennlfer G. Lee, Samantha B. Roberts, and François M. M. Morel. Cadmium: a nutrient for the marine diatom Thalassiosira weissflogii. Limnology and Oceanography, 40(6):1056–1063, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Lampert:1994:CIC

[LRvE94] Winfried Lampert, Karl Otto Rothhaupt, and Eric von Elert. Chemical induction of colony formation in a green alga (Scenedesmus acutus) by grazers (Daphnia). Limnology and Oceanography, 39(7):1543–1550, November 1994. CODEN LI-OCAH. ISSN 0024-3590.

Laird:1990:DLD

[LS90] Gwenyth A. Laird and Donald Scavia. Distribution of labile dissolved organic carbon in Lake Michigan. Limnology and Oceanography, 35(2):443–447, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Levine:1992:MPR

[LS92] S. N. Levine and D. W. Schindler. Modification of the N : P ratio in lakes by in situ processes. Limnology and Oceanography, 37(5): 917–935, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Lorang:1993:VSE

[LS93] Mark S. Lorang and Jack A. Stanford. Variability of shoreline erosion and accretion within a beach compartment of Flathead Lake, Montana. Limnology and Oceanography, 38(8):1783–1795, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Langdon:1992:SSI

[LSB92] Christopher Langdon, Raymond N. Sambrotto, and Ivars Bitte. In situ sampler-incubator for simultaneous biological rate measurements via tracers and net chemical change. Limnology and Oceanography, 37(8):1823–1830, December 1992. CODEN LI-OCAH. ISSN 0024-3590.

Lauerman:1997:ERI

[LSS^{+}97] L. M. L. Lauerman, J. M. Smoak, T. J. Shaw, W. S. Moore, and K. L. Smith, Jr. ^{234}Th and ^{210}Pb evidence for rapid ingestion of settling particles by mobile epibenthic megafauna in the abyssal NE Pacific. Limnology and Oceanography, 42(3):589–595, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Laws:1998:SIC

[LTPB98] Edward A. Laws, Peter A. Thompson, Brian N. Popp, and Robert R. Bidigure. Sources of inorganic carbon for marine microalgal photosynthesis: a reassessment of δ ^{13}C data from batch culture studies of Thalassiosira pseudonana and Emiliania huxleyi. Limnology and Oceanography, 43(1):136–142, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Lurling:1997:MCS

[LV97] Miquel Lürling and Ellen Van Donk. Morphological changes in Scenedesmus induced by infochemicals released in situ from zooplankton grazers. Limnology and Oceanography, 42(4):783–788, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Laurion:1998:CSV

[LV98] Isabelle Laurion and Warwick F. Vincent. Cell size versus taxonomic composition as determinants of UV-sensitivity in natural phytoplankton communities. Limnology and Oceanography, 43(8): 1774–1779, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Lovejoy:1993:MGT

[LVFD93] C. Lovejoy, W. F. Vincent, J.-J. Frenette, and J. J. Dodson. Microbial gradients in a turbid estuary: Application of a new method for protozoan community analysis. Limnology and Oceanography, 38(6):1295–1303, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Luoma:1998:MUP

[LvGLC98] Samuel N. Luoma, Alexander van Geen, Byeong-Gweon Lee, and James E. Cloern. Metal uptake by phytoplankton during a bloom in South San Francisco Bay: Implications for metal cycling in estuaries. Limnology and Oceanography, 43(5):1007–1016, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Luecke:1990:SRD

[LVM^{+}90] Chris Luecke, Michael J. Vanni, John J. Magnuson, James F. Kitchell, and Paul T. Jacobson. Seasonal regulation of Daphnia populations by planktivorous fish: Implications for the spring clear-water phase. Limnology and Oceanography, 35(8):1718–1733, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Leal:1999:CIR

[LVvdB99] M. Fernanda C. Leal, M. Teresa S. D. Vasconcelos, and Constant M. G. van den Berg. Copper-induced release of complexing ligands similar to thiols by Emiliania huxleyi in seawater cultures. Limnology and Oceanography, 44(7):1750–1762, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Logan:1990:FGM

[LW90] Bruce E. Logan and Daniel B. Wilkinson. Fractal geometry of marine snow and other biological aggregates. Limnology and Oceanography, 35(1):130–136, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Leibold:1993:EMM

[LW93] Mathew A. Leibold and Colin T. West. Experimental methods for measuring the effect of light acclimation on vertical migration by Daphnia in the field. Limnology and Oceanography, 38(3):638– 643, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Leichter:1996:PDS

[LWMD96] James J. Leichter, Stephen R. Wing, Steven L. Miller, and Mark W. Denny. Pulsed delivery of subthermocline water to Conch Reef (Florida Keys) by internal tidal bores. Limnology and Oceanography, 41(7):1490–1501, November 1996. CODEN LIOCAH. ISSN 0024-3590.

LaBaugh:1996:CAC

[LWS^{+}96] J. W. LaBaugh, T. C. Winter, G. A. Swanson, D. O. Rosenberry, R. D. Nelson, and N. H. Euliss, Jr. Changes in atmospheric circulation patterns affect midcontinent wetlands sensitive to climate. Limnology and Oceanography, 41(5):864–870, July 1996. CODEN LIOCAH. ISSN 0024-3590.

McDowell:1994:ECN

[MA94] William H. McDowell and Clyde E. Asbury. Export of carbon, nitrogen, and major ions from three tropical montane watersheds. Limnology and Oceanography, 39(1):111–125, January 1994. CO-DEN LIOCAH. ISSN 0024-3590.

Miller:1996:EGM

[MA96] Laurence G. Miller and George R. Aiken. Effects of glacial meltwater inflows and moat freezing on mixing in an ice-covered Antarctic lake as interpreted from stable isotope and tritium distributions. Limnology and Oceanography, 41(5):966–976, July 1996. CODEN LIOCAH. ISSN 0024-3590.

MaBone:1998:TDO

[MaB98] Robert A. MaBone. Theoretical developments on the optical properties of highly turbid waters and sea ice. Limnology and Oceanography, 43(1):29–33, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Machacek:1993:CRD

[Mac93a] Jiří Macháček. Comparison of the response of Daphnia galeata and Daphnia obtusa to fish-produced chemical substance. Limnology and Oceanography, 38(7):1544–1561, November 1993. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:1993:VMS

[Mac93b] Sally MacIntyre. Vertical mixing in a shallow, eutrophic lake: Possible consequences for the light climate of phytoplankton. Limnology and Oceanography, 38(4):798–817, June 1993. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:1995:AMN

[MAG95] Sally MacIntyre, Alice L. Alldredge, and Chris C. Gotschalk. Accumulation of marines now at density discontinuities in the water column. Limnology and Oceanography, 40(3):449–468, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Malone:1998:CWW

[Mal98] Thomas C. Malone. Coastal waters of the world: Trends, threats and strategies. Limnology and Oceanography, 43(7):1762, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Madhupratap:1991:EHA

[MAN91a] M. Madhupratap, C. T. Achuthankutty, and S. R. Sreekumaran Nair. Estimates of high absolute densities and emergence rates of demersal zooplankton from the Agatti Atoll, Laccadives. Limnology and Oceanography, 36(3):585–588, May 1991. CODEN LI-OCAH. ISSN 0024-3590.

Mann:1991:STE

[Man91b] Kenneth H. Mann. Seaweeds: Their environment, biogeography, and ecophysiology. Limnology and Oceanography, 36(5):1066, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Mann:1994:BRS

[Man94] Kenneth H. Mann. Book review: Saltmarsh ecology (P. Adam). Limnology and Oceanography, 39(2):475–476, March 1994. CO-DEN LIOCAH. ISSN 0024-3590.

Mann:1995:BRA

[Man95] Kenneth H. Mann. Book review: Aquatic ecology: Scale, pattern and process (P. S. Giller, A. G. Hildrew, and D. G. Raffaelli [eds.]). Limnology and Oceanography, 40(2):445–446, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Marinelli:1994:EBV

[Mar94] Roberta L. Marinelli. Effects of burrow ventilation on activities of a terebellid polychaete and silicate removal from sediment pore waters. Limnology and Oceanography, 39(2):303–317, March 1994. CODEN LIOCAH. ISSN 0024-3590.

McKnight:1991:AFA

[MAS91] Diane M. McKnight, George R. Aiken, and Richard L. Smith. Aquatic fulvic acids in microbially based ecosystems: Results from two desert lakes in Antarctica. Limnology and Oceanography, 36 (5):998–1006, July 1991. CODEN LIOCAH. ISSN 0024-3590.

McKnight:1994:AFA

[MASA94] Diane M. McKnight, Edmund D. Andrews, Sarah A. Spaulding, and George R. Aiken. Aquatic fulvic acids in algal-rich Antarctic ponds. Limnology and Oceanography, 39(8):1972–1979, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Matson:1993:NFT

[Mat93] Ernest A. Matson. Nutrient flux through soils and aquifers to the coastal zone of Guam (Mariana Islands). Limnology and Oceanography, 38(2):361–371, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Matrai:1994:BRD

[Mat94] Patricia A. Matrai. Book review: Dimethylsulphide: Oceans, atmosphere and climate (G. Restelli and G. Angeletti [eds.]). Limnology and Oceanography, 39(6):1497, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Mayer:1992:BRO

[May92] Lawrence M. Mayer. Book review: Ocean margin processes in global change (R. F. C. Mantoura. J.-M. Martin, and R. Wollast [eds.]). Limnology and Oceanography, 37(6):1341–1342, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Matson:1990:SCI

[MB90a] Ernest A. Matson and Mark M. Brinson. Stable carbon isotopes and the C : N ratio in the estuaries of the Pamlico and Neuse Rivers, North Carolina. Limnology and Oceanography, 35 (6):1290–1300, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Mullineaux:1990:REB

[MB90b] Lauren S. Mullineaux and Cheryl Ann Butman. Recruitment of encrusting benthic invertebrates in boundary-layer flows: a deepwater experiment on cross seamount. Limnology and Oceanography, 35(2):409–423, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Moffett:1996:PSE

[MB96] James W. Moffett and Larry E. Brand. Production of strong, extracellular Cu chelators by marine cyanobacteria in response to Cu stress. Limnology and Oceanography, 41(3):388–395, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Marcus:1998:MIB

[MB98a] Nancy H. Marcus and Ferdinando Boero. Minireview: The importance of benthic-pelagic coupling and the forgotten role of life cycles in coastal aquatic systems. Limnology and Oceanography, 43(5):763–768, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Mitchell:1998:EDO

[MB98b] A. Mitchell and D. S. Baldwin. Effects of desiccation/oxidation on the potential for bacterially mediated P release from sediments. Limnology and Oceanography, 43(3):481–487, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Morris:1999:ENL

[MB99] James T. Morris and Paul M. Bradley. Effects of nutrient loading on the carbon balance of coastal wetland sediments. Limnology and Oceanography, 44(3):699–702, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Moffett:1997:CSC

[MBCB97] James W. Moffett, Larry E. Brand, Peter L. Croot, and Katherine A. Barbeau. Cu speciation and cyanobacterial distribution in harbors subject to anthropogenic Cu inputs. Limnology and Oceanography, 42(5):789–799, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Martin:1990:YID

[MBFG90] John H. Martin, William W. Broenkow, Steve E. Fitzwater, and R. Maichael Gordon. Yes, it does: a reply to the comment by Banse. Limnology and Oceanography, 35(3):775–777, May 1990. CODEN LIOCAH. ISSN 0024-3590. See [Ban90d].

Mitchell:1991:LLP

[MBHH^{+}91] B. Greg Mitchell, Eric A. Brody, Osmund Holm-Hansen, Charles McClain, and James Bishop. Light limitation of phytoplankton biomass and macronutrient utilization in the Southern Ocean. Limnology and Oceanography, 36(8):1662–1677, December 1991. CODEN LIOCAH. ISSN 0024-3590.

McGowan:1999:ABG

[MBHM99] Suzanne McGowan, George Britton, Elizabeth Haworth, and Brian Moss. Ancient blue-green blooms. Limnology and Oceanography, 44(2):436–439, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Montagnes:1994:ECN

[MBHT94] David J. S. Montagnes, John A. Berges, Paul J. Harrison, and F. J. R. Taylor. Estimating carbon, nitrogen, protein, and chlorophyll-a from volume in marine phytoplankton. Limnology and Oceanography, 39(5):1044–1060, July 1994. CODEN LI-OCAH. ISSN 0024-3590.

Martinelli:1999:LCC

[MBK^{+}99] Luiz A. Martinelli, Maria Victoria Ballester, Alex V. Krusche, Reynaldo L. Victoria, Plinio B. de Camargo, Marcelo Bernardes, and Jean P. H. B. Ometto. Landcover changes and δ^{13}C composition of riverine particulate organic matter in the Piracicaba River Basin (southeast region of Brazil). Limnology and Oceanography, 44(7):1826–1833, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Mullins:1995:GCR

[MBKG95] Thomas D. Mullins, Theresa B. Britschgi, Robin L. Krest, and Stephen J. Giovannoni. Genetic comparisons reveal the same unknown bacterial lineages in Atlantic and Pacific bacterioplankton communities. Limnology and Oceanography, 40(1):148–158, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Medeiros-Bergen:1995:DHL

[MBOCK95] Dorothy E. Medeiros-Bergen, Richard R. Olson, Janet A. Conroy, and Thomas D. Kocher. Distribution of holothurian larvae determined with species-specific genetic probes. Limnology and Oceanography, 40(7):1225–1235, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Mulsow:1998:BPG

[MBS98] Sandor Mulsow, Bernard P. Boudreau, and John A. Smith. Bioturbation and porosity gradients. Limnology and Oceanography, 43(1):1–9, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Muller:1998:HSR

[MBSW98] Beat Müller, Kerst Buis, Ruth Stierli, and Bernhard Wehrli. High spatial resolution measurements in lake sediments with PVC based liquid membrane ion-selective electrodes. Limnology and Oceanography, 43(7):1728–1733, November 1998. CODEN LI-OCAH. ISSN 0024-3590.

Malm:1998:FSC

[MBT^{+}98] Joakim Malm, Lars Bengtsson, Arkady Terzhevik, Pjotr Boyarinov, Alexander Glinsky, Nikolaj Palshin, and Michail Petrov. Field study on currents in a shallow, ice-covered lake. Limnology and Oceanography, 43(7):1669–1679, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Millet:1992:WIH

[MC92] Bertrand Millet and Philippe Cecchi. Wind-induced hydrodynamic control of the phytoplankton biomass in a lagoon ecosystem. Limnology and Oceanography, 37(1):140–146, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Manca:1995:SCS

[MC95] Marina Manca and Patrizia Comoli. Seasonal changes in size of the feeding basket of Leptodora kindtii (Focke) in Lago Maggiore as related to variations in prey size selection. Limnology and Oceanography, 40(4):834–838, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Morgan:1996:SML

[MC96] Steven G. Morgan and John H. Christy. Survival of marine larvae under the countervailing selective pressures of photodamage and predation. Limnology and Oceanography, 41(3):498–504, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Murphy:1997:EDM

[MC97] Alice M. Murphy and Timothy J. Cowles. Effects of darkness on multi-excitation in vivo fluorescence and survival in a marine diatom. Limnology and Oceanography, 42(6):1444–1453, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Moore:1999:PMC

[MC99] Lisa R. Moore and Sallie W. Chisholm. Photophysiology of the marine cyanobacterium Prochlorococcus: Ecotypic differences among cultured isolates. Limnology and Oceanography, 44(3):628– 638, May 1999. CODEN LIOCAH. ISSN 0024-3590.

McDonald:1990:CMP

[MCB90] Michael E. McDonald, Larry B. Crowder, and Stephen B. Brandt. Changes in Mysis and Pontoporeia populations in southeastern Lake Michigan: a response to shifts in the fish community. Limnology and Oceanography, 35(1):220–227, January 1990. CODEN LIOCAH. ISSN 0024-3590.

McBride:1992:SCD

[McB92] Graham B. McBride. Simple calculation of daily photosynthesis by means of five photosynthesis-light equations. Limnology and Oceanography, 37(8):1796–1808, December 1992. CODEN LIOCAH. ISSN 0024-3590.

McConnaughey:1991:CCC

[McC91] Ted McConnaughey. Calcification in Chara corallina: CO_{2} hydroxylation generates protons for bicarbonate assimilation. Limnology and Oceanography, 36(4):619–628, June 1991. CODEN LIOCAH. ISSN 0024-3590.

McCormick:1992:AOA

[McC92] N. J. McCormick. Asymptotic optical attenuation. Limnology and Oceanography, 37(7):1570–1578, November 1992. CODEN LIOCAH. ISSN 0024-3590.

McCormick:1995:MMM

[McC95] N. J. McCormick. Mathematical models for the mean cosine of irradiance and the diffuse attenuation coefficient. Limnology and Oceanography, 41(4):1013–1018, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Maynou:1996:SSS

[MCC^{+}96] Francesc Maynou, Gérard Y. Conan, Joan E. Cartes, Joan B. Company, and Francisco Sardà. Spatial structure and seasonality of decapod crustacean populations on the northwestern Mediterranean slope. Limnology and Oceanography, 41(1):113–125, January 1996. CODEN LIOCAH. ISSN 0024-3590.

McClain:1997:BRL

[McC97] Michael E. McClain. Book review: Limnology in developing countries (B. Gopal and G. Wetzel [eds.]). Limnology and Oceanography, 42(1):206, January 1997. CODEN LIOCAH. ISSN 0024-3590.

McManus:1992:PPC

[MCCD92] James McManus, Robert W. Collier, Chen-Tung A. Chen, and J. Dymond. Physical properties of Crater Lake, Oregon: a method for the determination of a conductivity- and temperaturedependent expression for salinity. Limnology and Oceanography, 37(1):41–53, January 1992. CODEN LIOCAH. ISSN 0024-3590.

McManus:1996:STD

[MCD^{+}96] James McManus, Robert Collier, Jack Dymond, C. Geoff Wheat, and Gary L. Larson. Spatial and temporal distribution of dissolved oxygen in Crater Lake, Oregon. Limnology and Oceanography, 41 (4):722–731, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Marba:1994:MLS

[MCED94] Núria Marbà, Just Cebrián, Susana Enríquez, and Carlos M. Duarte. Migration of large-scale subaqueous bedforms measured with seagrasses (Cymodocea nodosa) as tracers. Limnology and Oceanography, 39(1):126–133, January 1994. CODEN LIOCAH. ISSN 0024-3590.

McKinnon:1996:GDS

[McK96] A. D. McKinnon. Growth and development in the subtropical copepod Acrocalanus gibber. Limnology and Oceanography, 41(7): 1438–1447, November 1996. CODEN LIOCAH. ISSN 0024-3590.

McLuren:1997:MBE

[McL97] Ian A. McLuren. Modeling biases in estimating production from copepod cohorts. Limnology and Oceanography, 42(3):584–589, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Malone:1990:MBC

[MD90] Thomas C. Malone and Hugh W. Ducklow. Microbial biomass in the coastal plume of Chesapeake Bay: Phytoplanktonbacterioplankton relationships. Limnology and Oceanography, 35 (2):296–312, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Marba:1997:ICS

[MD97] Núria Marbà and Carlos M. Duarte. Interannual changes in seagrass (Posidonia oceanica) growth and environmental change in the Spanish Mediterranean littoral zone. Limnology and Oceanography, 42(5):800–810, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Martin:1995:SCB

[MDC95] Jean-Marie Martin, Min-Han Dai, and Gustave Cauwet. Significance of colloids in the biogeochemical cycling of organic carbon and trace metals in the Venice Lagoon (Italy). Limnology and Oceanography, 40(1):119–131, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Mostajir:1999:ETE

[MDdM^{+}99] Behzad Mostajir, Serge Demers, Stephen de Mora, Claude Belzile, Jean-Pierre Chanut, Michel Gosselin, Suzanne Roy, Piedad Zulema Villegas, Juliette Fauchot, Josée Bouchard, David Bird, Patrick Monfort, and Maurice Levasseur. Experimental test of the effect of ultraviolet-b radiation in a planktonic community. Limnology and Oceanography, 44(3):586–596, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Main:1997:PSO

[MDE97] T. M. Main, D. R. Dobberfuhl, and J. J. Elser. N : P stoichiometry and ontogeny of crustacean zooplankton: a test of the growth rate hypothesis. Limnology and Oceanography, 42(6):1474–1478, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Manley:1997:MIP

[MdlC97] Steven L. Manley and John L. de la Cuesta. Methyl iodide production from marine phytoplankton cultures. Limnology and Oceanography, 42(1):142–147, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Menge:1997:RIO

[MDWS97] Bruce A. Menge, Bryon A. Daley, Patricia A. Wheeler, and P. Ted Strub. Rocky intertidal oceanography: an association between community structure and nearshore phytoplankton concentration. Limnology and Oceanography, 42(1):57–66, January 1997. CO-DEN LIOCAH. ISSN 0024-3590.

MacKay:1998:NRD

[ME98a] Neil A. MacKay and James J. Elser. Nutrient recycling by Daphnia reduces N_{2} fixation by cyanobacteria. Limnology and Oceanography, 43(2):347–354, March 1998. CODEN LIOCAH. ISSN 0024-3590.

MacKay:1998:FPP

[ME98b] Neil A. MacKay and Jumes J. Elser. Factors potentially preventing trophic cascades: Food quality, invertebrate predation, and their interaction. Limnology and Oceanography, 43(2):339–347, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Mehta:1999:SPI

[Meh99] Ashish J. Mehta. Sedimentary processes in the intertidal zone. Limnology and Oceanography, 46(4):1184–1185, June 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Meiack:1990:BRI

[Mei90] John M. Meiack. Book review: Inland waters of southern Africa: an ecological perspective (B. R. Allanson, R. C. Hart, J. H. O’Keefe, and R. D. Robarts). Limnology and Oceanography, 35(8): 1864–1865, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Melack:1999:MBH

[Mel99] John M. Melack. Microbiology and biogeochemistry of hypersaline environments. Limnology and Oceanography, 44(6):1597, September 1999. CODEN LIOCAH. ISSN 0024-3590.

McLeroy-Etheridge:1999:FTC

[MEM99] Stacey L. McLeroy-Etheridge and George B. McManus. Food type and concentration affect chlorophyll and carotenoid destruction during copepod feeding. Limnology and Oceanography, 44(8): 2005–2011, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Menge:1990:BRB

[Men90] Bruce Menge. Book review: Biology and the mechanics of the wave-swept environment (M. W. Denny). Limnology and Oceanography, 35(1):239–240, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Michaels:1990:LMP

[MF90] Anthony F. Michaels and A. Russell Flegal. Lead in marine planktonic organisms and pelagic food webs. Limnology and Oceanography, 35(2):287–295, March 1990. CODEN LIOCAH. ISSN 0024-3590.

McCormick:1999:RCT

[MF99] Michael J. McCormick and Gary L. Fahnenstiel. Recent climatic trends in nearshore water temperatures in the St. Lawrence Great Lakes. Limnology and Oceanography, 44(3):530–540, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Mason:1993:MBC

[MFH^{+}93] R. P. Mason, W. F. Fitzgerald, J. Hurley, A. K. Hanson, Jr., P. L. Donaghay, and J. M. Sieburth. Mercury biogeochemical cycling in a stratified estuary. Limnology and Oceanography, 38(6):1227– 1241, September 1993. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:1999:BMN

[MFJR99] Sally MacIntyre, Kevin M. Flynn, Robert Jellison, and José R. Romero. Boundary mixing and nutrient fluxes in Mono Lake, California. Limnology and Oceanography, 44(3):512–529, May 1999. CODEN LIOCAH. ISSN 0024-3590.

McCormick:1996:CCS

[MFLR96] Michael J. McCormick, Gary L. Fahnenstiel, Steven E. Lohrenz, and Donald G. Redalje. Calculation of cell-specific growth rates: a clarification. Limnology and Oceanography, 41(1):182–189, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Miller:1991:EDS

[MFW^{+}91] Charles B. Miller, Bruce W. Frost, Patricia A. Wheeler, Michael R. Landry, Nicholas Welschmeyer, and Thomas M. Powell. Ecological dynamics in the subarctic Pacific, a possibly ironlimited ecosystem. Limnology and Oceanography, 36(8):1600– 1615, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Moore:1999:PCC

[MFY99] Paul A. Moore, David M. Fields, and Jeannette Yen. Physical constraints of chemoreception in foraging copepods. Limnology and Oceanography, 44(1):166–177, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Martin:1991:CI

[MGF91] John H. Martin, Michael Gordon, and Steve E. Fitzwater. The case for iron. Limnology and Oceanography, 36(8):1793–1802, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Manley:1992:LPB

[MGN92] Steven L. Manley, Kelly Goodwin, and Wheeler J. North. Laboratory production of bromoform, methylene bromide, and methyl iodide by macroalgae and distribution in nearshore southern California waters. Limnology and Oceanography, 37(8):1652–1659, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Mengis:1996:NOE

[MGW96] Martin Mengis, René Gächter, and Bernhard Wehrli. Nitrous oxide emissions to the atmosphere from an artificially oxygenated lake. Limnology and Oceanography, 41(3):548–553, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Mengis:1997:NET

[MGWB97] Martin Mengis, René Gächter, Bernhard Wehrli, and Stefano Bernasconi. Nitrogen elimination in two deep eutrophic lakes. Limnology and Oceanography, 42(7):1530–1543, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Moran:1990:BPH

[MH90] Mary Ann Moran and Robert E. Hodson. Bacterial production on humic and nonhumic components of dissolved organic carbon. Limnology and Oceanography, 35(8):1744–1756, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Millham:1994:FFC

[MH94a] Newton P. Millham and Brian L. Howes. Freshwater flow into a coastal embayment: Groundwater and surface water inputs. Limnology and Oceanography, 39(8):1928–1944, December 1994. CO-DEN LIOCAH. ISSN 0024-3590.

Moran:1994:DHS

[MH94b] Mary Ann Moran and Robert E. Hodson. Dissolved humic substances of vascular plant origin in a coastal marine environment. Limnology and Oceanography, 39(4):762–771, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Moeseneder:1995:ITB

[MH95] Markus M. Moeseneder and Gerhard J. Herndl. Influence of turbulence on bacterial production in the sea. Limnology and Oceanography, 40(8):1466–1473, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Morris:1997:RPD

[MH97] Donald P. Morris and Bruce R. Hargreaves. The role of photochemical degradation of dissolved organic carbon in regulating the UV transparency of three lakes on the Pocono Plateau. Limnology and Oceanography, 42(2):239–249, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Meyer-Harms:1999:SFN

[MHIHH99] Bettina Meyer-Harms, Xabier Irigoien, Robert Head, and Roger Harris. Selective feeding on natural phytoplankton by Calanus finmarchicus before, during, and after the 1997 spring bloom in the Norwegian Sea. Limnology and Oceanography, 44(1):154–165, January 1999. CODEN LIOCAH. ISSN 0024-3590.

McDonald:1996:GWI

[MHM96] Michael E. McDonald, Anne E. Hershey, and Michael C. Miller. Global warming impacts on lake trout in Arctic lakes. Limnology and Oceanography, 41(5):1102–1108, July 1996. CODEN LI-OCAH. ISSN 0024-3590.

Miller:1999:ENV

[MHM^{+}99] M. W. Miller, M. E. Hay, S. L. Miller, D. Malone, E. E. Sotka, and A. M. Szmant. Effects of nutrients versus herbivores on reef algae: a new method for manipulating nutrients on coral reefs. Limnology and Oceanography, 44(8):1847–1861, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Marshall:1999:VBP

[MHMD99] Rhoda A. Marshall, David B. Harper, W. Colin McRoberts, and Matthew J. Dring. Volatile bromocarbons produced by Falkenbergia stages of Asparagopsis spp. (Rhodophyta). Limnology and Oceanography, 44(5):1348–1352, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Morel:1991:LPT

[MHP91] François M. M. Morel, Robert J. M. Hudson, and Neil M. Price. Limitation of productivity by trace metals in the sea. Limnology and Oceanography, 36(8):1742–1755, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Marino:1990:MSC

[MHSP90] Roxanne Marino, Robert W. Howarth, Jennifer Shamess, and Ellie Prepas. Molybdenum and sulfate as controls on the abundance of nitrogen-fixing cyanobacteria in saline lakes in Alberta. Limnology and Oceanography, 35(2):245–259, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Murrell:1999:BDN

[MHSW99] M. C. Murrell, J. T. Hollibaugh, M. W. Silver, and P. S. Wong. Bacterioplankton dynamics in northern San Francisco Bay: Role of particle association and seasonal freshwater flow. Limnology and Oceanography, 44(2):295–308, March 1999. CODEN LI-OCAH. ISSN 0024-3590.

Munger:1999:CSE

[MHT99] Catherine Munger, Landis Hare, and André Tessier. Cadmium sources and exchange rates for Chaoborus larvae in nature. Limnology and Oceanography, 44(7):1763–1771, November 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Methe:1998:CBM

[MHZ98] Barbara A. Methé, William D. Hiorns, and Jonathan P. Zehr. Contrasts between marine and freshwater bacterial community composition: Analyses of communities in Lake George and six other Adirondack lakes. Limnology and Oceanography, 43(2):368– 374, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Millero:1991:OHF

[Mil91] Frank J. Millero. The oxidation of H_{2} S in Framvaren Fjord. Limnology and Oceanography, 36(5):1007–1014, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Miller:1995:BRD

[Mil95] Douglas C. Miller. Book review: The design and analysis of ecological experiments (S. M. Scheiner and J. Gurevitch [eds.]). Limnology and Oceanography, 40(8):1533–1534, December 1995. CO-DEN LIOCAH. ISSN 0024-3590.

Monismith:1990:CMS

[MIM90] Stephen G. Monismith, Jörg Imberger, and Michael L. Morison. Convective motions in the sidearm of a small reservoir. Limnology and Oceanography, 35(8):1676–1702, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Michael:1994:RDN

[MJ94] Coates Michael and Ferris John. The radiatively driven natural convection beneath a floating plant layer. Limnology and Oceanography, 39(5):1186–1194, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Morales:1993:SUR

[MJC93] Raúl G. E. Morales, Gregorio P. Java, and Sergio Cabrera. Solar ultraviolet radiation measurements by o-nitrobenzaldehyde actinometry. Limnology and Oceanography, 38(3):703–705, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Marinelli:1998:SND

[MJC^{+}98] Roberta L. Marinelli, Richard A. Jahnke, Deborah B. Craven, James R. Nelson, and James E. Eckman. Sediment nutrient dynamics on the South Atlantic Bight continental shelf. Limnology and Oceanography, 43(6):1305–1320, September 1998. CODEN LIOCAH. ISSN 0024-3590.

McClatchie:1991:OGF

[MJG91] S. McClatchie, J. B. Jillett, and P. Gerring. Observations of gulls foraging on beach-stranded plankton in Otago Harbor, New Zealand. Limnology and Oceanography, 36(6):1195–1200, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Miller:1993:MHM

[MJOC93] Laurence G. Miller, Robert Jellison, Ronald S. Oremland, and Charles W. Culbertson. Meromixis in hypersaline Mono Lake, California. 3. Biogeochemical response to stratification and overturn. Limnology and Oceanography, 38(5):1040–1051, July 1993. CODEN LIOCAH. ISSN 0024-3590.

MacKenzie:1995:ERS

[MK95] Brian R. MacKenzie and Thomas Kiørboe. Encounter rates and swimming behavior of pause-travel and cruise larval fish predators in calm and turbulent laboratory environments. Limnology and Oceanography, 40(7):1278–1289, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Mortimer:1999:UHR

[MKBN99] R. J. G. Mortimer, M. D. Krom, D. R. Boyle, and A. Nishri. Use of a high-resolution pore-water gel profiler to measure groundwater fluxes at an underwater saline seepage site in Lake Kinneret, Israel. Limnology and Oceanography, 44(7):1802–1809, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Morris:1990:DEP

[MKD90] James T. Morris, Björn Kjerfve, and John M. Dean. Dependence of estuarine productivity on anomalies in mean sea level. Limnology and Oceanography, 35(4):926–930, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Monismith:1990:SMB

[MKT^{+}90] Stephen G. Monismith, Jeffrey R. Koseff, Janet K. Thompson, Catherine A. O’Riordan, and Heidi M. Nepf. A study of model bivalve siphonal currents. Limnology and Oceanography, 35(3): 680–696, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Megard:1997:SDZ

[MKWD97] R. O. Megard, M. M. Kuns, M. C. Whiteside, and J. A. Downing. Spatial distributions of zooplankton during coastal upwelling in western Lake Superior. Limnology and Oceanography, 42(5):827– 840, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Mourelatos:1990:SFR

[ML90] Spiros Mourelatos and Gérard Lacroix. In situ filtering rates of cladocera: Effect of body length, temperature, and food concentration. Limnology and Oceanography, 35(5):1101–1111, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Morris:1992:NLB

[ML92] Donald P. Morris, Jr. and William M. Lewis. Nutrient limitation of bacterioplankton growth in Lake Dillon, Colorado. Limnology and Oceanography, 37(6):1179–1192, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Michalski:1995:DVM

[ML95] J. Michalski and U. Lemmin. Dynamics of vertical mixing in the hypolimnion of a deep lake: Lake Geneva. Limnology and Oceanography, 40(4):809–816, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Marsh:1996:HRP

[ML96] Philip Marsh and Lance F. W. Lesack. The hydrologic regime of perched lakes in the Mackenzie Delta: Potential responses to climate change. Limnology and Oceanography, 41(5):849–856, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Monger:1999:FSH

[MLB99] Bruce C. Monger, Michael R. Landry, and Susan L. Brown. Feeding selection of heterotrophic marine nanoflagellates based on the surface hydrophobicity of their picoplankton prey. Limnology and Oceanography, 44(8):1917–1927, December 1999. CODEN LI-OCAH. ISSN 0024-3590.

Marcus:1994:AVV

[MLBC94] Nancy H. Marcus, Robert Lutz, William Burnett, and Peter Cable. Age, viability, and vertical distribution of zooplankton resting eggs from an anoxic basin: Evidence of an egg bank. Limnology and Oceanography, 39(1):154–158, January 1994. CODEN LI-OCAH. ISSN 0024-3590.

McDowell:1999:VRN

[MLLR99] Judith E. McDowell, Bruce A. Lancaster, Dale F. Leavitt, and Pirjo Rantamaki. Variability in responses to nutrients and trace elements, and transmission of stressor effects through an estuarine food web. Limnology and Oceanography, 44(3 (part 2)):903–909, May 1999. CODEN LIOCAH. ISSN 0024-3590.

McConnaughey:1994:CBG

[MLR^{+}94] Ted A. McConnaughey, James W. LaBaugh, Donald O. Rosenberry, Robert G. Striegl, Michael M. Reddy, Paul F. Schuster, and Virginia Carter. Carbon budget for a groundwater-fed lake: Calcification supports summer photosynthesis. Limnology and Oceanography, 39(6):1319–1332, September 1994. CODEN LI-OCAH. ISSN 0024-3590.

Miller:1991:EEF

[MM91] Ronald L. Miller and Benjamin F. McPherson. Estimating estuarine flushing and residence times in Charlotte harbor, Florida. via salt balance and a box model. Limnology and Oceanography, 36(3):602–612, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Martin:1993:DOC

[MM93] William R. Martin and Daniel C. McCorkle. Dissolved organic carbon concentrations in marine pore waters determined by hightemperature oxidation. Limnology and Oceanography, 38(7):1464– 1479, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Marchessault:1997:GNI

[MM97a] Philippe Marchessault and A. Mazumder. Grazer and nutrient impacts on epilimnetic ciliate communities. Limnology and Oceanography, 42(5):893–900, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Miller:1997:IPM

[MM97b] William L. Miller and Mary Ann Moran. Interaction of photochemical and microbial processes in the degradation of refractory dissolved organic matter from a coastal marine environment. Limnology and Oceanography, 42(6):1317–1324, September 1997. CO-DEN LIOCAH. ISSN 0024-3590.

Moisan:1999:PAP

[MM99] Tiffany A. Moisan and B. Greg Mitchell. Photophysiological acclimation of Phaeocystis antarctica Karsten under light limitation. Limnology and Oceanography, 44(2):247–258, March 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Moss:1994:DPC

[MMC94] Brian Moss, Suzanne McGowan, and Laurence Carvalho. Determination of phytoplankton crops by top-down and bottom-up mechanisms in a group of English lakes, the West Midland meres. Limnology and Oceanography, 39(5):1020–1029, July 1994. CO-DEN LIOCAH. ISSN 0024-3590.

MacKenzie:1994:EDS

[MMCL94] Brian R. MacKenzie, Thomas J. Miller, Stéphane Cyr, and William C. Leggett. Evidence for a dome-shaped relationship between turbulence and larval fish ingestion rates. Limnology and Oceanography, 39(8):1790–1799, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Marcel:1994:IDE

[MMG94a] Babin Marcel, Andé Morel, and Real Gagnon. An incubator designed for extensive and sensitive measurements of phytoplankton photosynthetic parameters. Limnology and Oceanography, 39(3): 694–702, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Maritorena:1994:DRO

[MMG94b] Stéphane Maritorena, André Morel, and Bernard Gentili. Diffuse reflectance of oceanic shallow waters: Influence of water depth and bottom albedo. Limnology and Oceanography, 39(7):1689– 1703, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Miyajima:1995:DSI

[MMH^{+}95] Toshihiro Miyajima, Yoshihiro Miyajima, Yuko T. Hanba, Koichi Yoshii, Tadatoshi Koitabashi, and Eitaro Wada. Determining the stable isotope ratio of total dissolved inorganic carbon in lake water by GC/C/IIRMS. Limnology and Oceanography, 41(4):994– 1000, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Miller:1998:SLC

[MMPS98] Charles B. Miller, Cheryl A. Morgan, Fredrick G. Prahl, and Margaret A. Sparrow. Storage lipids of the copepod Calanus finmarchicus from Georges Bank and the Gulf of Maine. Limnology and Oceanography, 43(3):488–497, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Magyar:1993:VDT

[MMS93] B. Magyar, H. C. Moor, and L. Sigg. Vertical distribution and transport of molybdenum in a lake with a seasonally anoxic hypolimnion. Limnology and Oceanography, 38(3):521–531, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Muller-Mniklas:1994:OCB

[MMSKH94] Gerald Müller-Mniklas, Schuster Stefan, Elisabeth Kaltenböok, and Gerhard J. Herndl. Organic content and bacterial metabolism in amorphous aggregations of the northern Adriatic Sea. Limnology and Oceanography, 39(1):58–68, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Malm:1994:IST

[MMTJ94] Joakim Malm, Dmitrii Mironov, Arkadii Terzhevik, and Lennart Jiinsson. Investigation of the spring thermal regime in Lake Ladoga using field and satellite data. Limnology and Oceanography, 39(6):1333–1348, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:1997:EHZ

[MMW^{+}97] Patrick J. Mulholland, Erich R. Marzolf, Jackson R. Webster, Deborah R. Hart, and Susan P. Hendricks. Evidence that hyporheic zones increase heterotrophic metabolism and phosphorus uptake in forest streams. Limnology and Oceanography, 42(3): 443–451, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:1999:EHZ

[MMW^{+}99] Patrick J. Mulholland, Erich R. Marzolf, Jackson R. Webster, Deborah R. Hart, and Susan P. Hendricks. Evidence that hyporheic zones increase heterotrophic metabolism and phosphorus uptake in forest streams. Limnology and Oceanography, 44(1): 230–231, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Muller-Navarra:1995:BVM

[MN95] Dörthe C. Müller-Navarra. Biochemical versus mineral limitation in Daphnia. Limnology and Oceanography, 40(7):1209–1214, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Muller-Niklas:1996:MDB

[MNAH96] Gerald Müller-Niklas, Martin Agis, and Gerhard J. Herndl. Microscale distribution of bacterioplankton in relation to phytoplankton: Results from 1-nl samples. Limnology and Oceanography, 41(7):1577–1582, November 1996. CODEN LIOCAH. ISSN 0024-3590.

McManus:1991:USF

[MO91] George B. McManus and Akira Okubo. On the use of surrogate food particles to measure protistan ingestion. Limnology and Oceanography, 36(3):613–617, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Mitchell:1990:GNM

[MOF90] James G. Mitchell, Akira Okubo, and Jed A. Fuhrman. Gyrotaxis as a new mechanism for generating spatial heterogeneity and migration in microplankton. Limnology and Oceanography, 35(1): 123–130, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Moflett:1994:RBC

[Mof94] James W. Moflett. The relationship between cerium and manganese oxidation in the marine environment. Limnology and Oceanography, 39(6):1309–1318, September 1994. CODEN LI-OCAH. ISSN 0024-3590.

Moline:1998:PRD

[Mol98] Mark A. Moline. Photoadaptive response during the development of a coustal Antarctic diatom bloom and relationship to water column stability. Limnology and Oceanography, 43(1):146–153, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Morse:1990:ACK

[Mor90] John W. Morse. Aquatic chemical kinetics: Reaction rates of processes in natural waters. Limnology and Oceanography, 35(8): 1865, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Morel:1997:CSB

[Mor97] André Morel. Consequences of a Synechococcus bloom upon the optical properties of oceanic (case 1) waters. Limnology and Oceanography, 42(8):1746–1754, December 1997. CODEN LI-OCAH. ISSN 0024-3590.

Madin:1992:FMG

[MP92a] Laurence P. Madin and Jennifer E. Purcell. Feeding, metabolism. and growth of Cyclosalpa bakeri in the subarctic Pacific. Limnology and Oceanography, 37(6):1236–1251, September 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Mallin:1992:EVI

[MP92b] Michael A. Mallin and Hans W. Paerl. Effects of variable irradiance on phytoplankton productivity in shallow estuaries. Limnology and Oceanography, 37(1):54–62, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Massana:1994:MDI

[MPA94] Ramon Massana and Carlos Pedrós-Alió. A method to determine integrated predation in stratified waters. Limnology and Oceanography, 39(2):248–262, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Miller:1995:ITI

[MPG95] Carolyn A. Miller, Deborah L. Penry, and Patricia M. Glibert. The impact of trophic interactions on rates of nitrogen regeneration and grazing in Chesapeake Bay. Limnology and Oceanography, 41(4):1005–1011, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Moran:1991:DTD

[MPS^{+}91] Mary Ann Moran, Lawrence R. Pomeroy, Edward S. Sheppard, Larry P. Atkinson, and Robert E. Hodson. Distribution of terrestrially derived dissolved organic matter on the southeastern U.S. continental shelf. Limnology and Oceanography, 36(6):1134–1149, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Matthaei:1999:PSS

[MPT99] Christoph D. Matthaei, Kathi A. Peacock, and Colin R. Townsend. Patchy surface stone movement during disturbance in a New Zealand stream and its potential significance for the fauna. Limnology and Oceanography, 46(4):1091–1102, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Mortsch:1996:CCS

[MQ96] Linda D. Mortsch and Frank H. Quinn. Climate change scenarios for Great Lakes Basin ecosystem studies. Limnology and Oceanography, 41(5):903–911, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Mayer:1992:RED

[MR92a] Lawrence M. Mayer and Donald L. Rice. RICE: Early diagenesis of protein: a seasonal study. Limnology and Oceanography, 37(2): 280–295, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Mayzaud:1992:DGP

[MR92b] Patrick Mayzaud and Suzanne Razouls. Degradation of gut pigment during feeding by a subantarctic copepod: Importance of feeding history and digestive acclimation. Limnology and Oceanography, 37(2):393–404, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Manca:1998:CPF

[MR98] M. Manca and D. Ruggiu. Consequences of pelagic food-web changes during a long-term lake oligotrophication process. Limnology and Oceanography, 43(6):1368–1373, September 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Monaghan:1999:DOP

[MR99] E. J. Monaghan and K. C. Ruttenberg. Dissolved organic phosphorus in the coastal ocean: Reassessment of available methods and seasonal phosphorus profiles from the Eel River Shelf. Limnology and Oceanography, 44(7):1702–1714, November 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Makulla:1993:RBR

[MS93a] Alexandra Makulla and Ulrich Sommer. Relationships between resource ratios and phytoplankton species composition during spring in five north German lakes. Limnology and Oceanography, 38(4):846–856, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Mallini:1993:DDM

[MS93b] Leanne J. Mallini and Alan M. Shiller. Determination of dissolved manganese in seawater by flow injection analysis with colorimetric detection. Limnology and Oceanography, 38(6):1290– 1295, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Mobley:1997:EMP

[MS97] Curtis D. Mobley and Dariusz Stramski. Effects of microbial particles on oceanic optics: Methodology for radiative transfer modeling and example simulations. Limnology and Oceanography, 42 (3):550–560, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Migon:1999:PRP

[MS99] Christophe Migon and Valérie Sandroni. Phosphorus in rainwater: Partitioning inputs and impact on the surface coastal ocean. Limnology and Oceanography, 46(4):1160–1165, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Middelburg:1996:ENI

[MSH96a] Jack J. Middelburg, Karline Soetaert, and Peter M. J. Herman. Evaluation of the nitrogen isotope-pairing method for measuring benthic denitrification: a simulation analysis. Limnology and Oceanography, 41(4):1839–1844, December 1996. CODEN LI-OCAH. ISSN 0024-3590. See replies [MSH96a, MSH96b].

Middelburg:1996:RCN

[MSH96b] Jack J. Middelburg, Karline Soetaert, and Peter M. J. Herman. Reply to the comment by Nielsen et al. Limnology and Oceanography, 41(4):1846–1847, December 1996. CODEN LIOCAH. ISSN 0024-3590. See [NRPRB96].

Millie:1997:DHA

[MSK^{+}97] David F. Millie, Oscar M. Schofield, Gary J. Kirkpatrick, Geir Johnsen, Patricia A. Tester, and Bryan T. Vinyard. Detection of harmful algal blooms using photopigments and absorption signatures: a case study of the Florida red tide dinoflagellate, Gymnodinium breve. Limnology and Oceanography, 42(5 (part 2)):1240– 1251, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Michmerhuizen:1996:PME

[MSM96] Catherine M. Michmerhuizen, Robert G. Striegl, and Michael E. McDonald. Potential methane emission from north-temperate lakes following ice melt. Limnology and Oceanography, 41(5):985– 991, July 1996. CODEN LIOCAH. ISSN 0024-3590.

McNaught:1999:RFW

[MSP^{+}99] A. S. McNaught, D. W. Schindler, B. R. Parker, A. J. Paul, R. S. Anderson, D. B. Donald, and M. Agbeti. Restoration of the food web of an alpine lake following fish stocking. Limnology and Oceanography, 44(1):127–136, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Mayer:1995:BAA

[MSS^{+}95] Lawrence M. Mayer, Linda L. Schick, Thomas Sawyer, Craig J. Plante, Peter A. Jumars, and Robert L. Sel. Bioavailable amino acids in sediments: a biomimetic, kinetics based approach. Limnology and Oceanography, 40(3):511–520, May 1995. CODEN LI-OCAH. ISSN 0024-3590.

McClintock:1996:FDP

[MSSM96] James B. McClintock, Daniel P. Swenson, Deborah K. Steinberg, and Anthony A. Michaels. Feeding-deterrent properties of common oceanic holoplankton from Bermudian waters. Limnology and Oceanography, 41(8):798–801, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Marsh:1990:RNR

[MT90] Adam G. Marsh and Kenneth R. Tenore. The role of nutrition in regulating the population dynamics of opportunistic, surface deposit feeders in a mesohaline community. Limnology and Oceanography, 35(3):710–724, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Mazumder:1994:TSL

[MT94] Asit Mazumder and William D. Taylor. Thermal structure of lakes varying in size and water clarity. Limnology and Oceanography, 39(4):968–976, June 1994. CODEN LIOCAH. ISSN 0024-3590.

McIntosh:1995:IIP

[MT95] Angus R. McIntosh and Colin R. Townsend. Impacts of an introduced predatory fish on mayfly grazing in New Zealand streams.

Limnology and Oceanography, 40(8):1508–1512, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Marubini:1999:BAP

[MT99] Francesca Marubini and Brenda Thake. Bicarbonate addition promotes coral growth. Limnology and Oceanography, 44(3):716–720, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Meyers:1998:SRS

[MTLR98] Philip A. Meyers, Gabrielle E. Tenser, Martin E. Lebo, and John E. Reuter. Sedimentary record of sources and accumulation of organic matter in Pyramid Lake, Nevada, over the past 1,000 years. Limnology and Oceanography, 43(1):160–169, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Massana:1998:VDT

[MTM^{+}98] Ramon Massana, Lance T. Taylor, Alison E. Murray, Ke Y. Wu, Wade H. Jeffrey, and Edward F. DeLong. Vertical distribution and temporal variation of marine planktonic archaea in the Gerlache Strait, Antarctica, during early spring. Limnology and Oceanography, 43(4):607–617, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Mucci:1991:PBT

[Muc91] Alfonso Mucci. A potentiometric back-titration method for determining sulfate in seawater and marine pore waters. Limnology and Oceanography, 36(2):409–412, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Mittenzwey:1992:DCI

[MUGK92] K.-H. Mittenzwey, S. Ullrich, A. A. Gitelson, and K. Y. Kondratiev. Determination of chlorophyll-a of inland waters on the basis of spectral reflectance. Limnology and Oceanography, 37(1): 147–149, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:1992:RNC

[Mul92] Patrick J. Mulholland. Regulation of nutrient concentrations in a temperate forest stream: Roles of upland, riparian, and instream processes. Limnology and Oceanography, 37(7):1512–1526, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Moore:1996:SAR

[MV96] J. Keith Moore and Tracy A. Villareal. Size-ascent rate relationships in positively buoyant marine diatoms. Limnology and Oceanography, 41(7):1514–1520, November 1996. CODEN LI-OCAH. ISSN 0024-3590.

McClelland:1998:LNE

[MV98] James W. McClelland and Ivan Valiela. Linking nitrogen in estuarine producers to land-derived sources. Limnology and Oceanography, 43(4):577–585, June 1998. CODEN LIOCAH. ISSN 0024-3590.

McClelland:1997:NSI

[MVM97] James W. McClelland, Ivan Valiela, and Robert H. Michener. Nitrogen-stable isotope signatures in estuarine food webs: a record of increasing urbanization in coastal watersheds. Limnology and Oceanography, 42(5):930–937, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Maffione:1998:TMC

[MVM98] Robert A. Maffione, Jeff M. Voss, and Curtis D. Mobley. Theory and measurements of the complete beam spread function of sea ice. Limnology and Oceanography, 43(1):34–43, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Markager:1999:CFP

[MVT99] Stiig Markager, Warwick F. Vincent, and Evonne P. Y. Tang. Carbon fixation by phytoplankton in high Arctic lakes: Implications of low temperature for photosynthesis. Limnology and Oceanography, 44(3):597–607, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Matisoff:1998:STS

[MW98a] Gerald Matisoff and Xiaosong Wang. Solute transport in sediments by freshwater infaunal bioirrigators. Limnology and Oceanography, 43(7):1487–1499, November 1998. CODEN LI-OCAH. ISSN 0024-3590.

McPeek:1998:GVR

[MW98b] Mark A. McPeek and Gay A. Wellborn. Genetic variation and reproductive isolation among phenotypically divergent amphipod populations. Limnology and Oceanography, 43(6):1162–1169, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Malin:1998:EPD

[MWB^{+}98] Gillian Malin, William H. Wilson, Gunnar Bratbak, Peter S. Liss, and Nicholas H. Mann. Elevated production of dimethylsulfide resulting from viral infection of cultures of Phaeocystis pouchetii. Limnology and Oceanography, 43(6):1389–1393, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Munnich:1992:OSV

[MWI92] M. Münnich, A. Wüest, and D. M. Imboden. Observations of the second vertical mode of the internal seiche in an alpine lake. Limnology and Oceanography, 37(8):1705–1719, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Moorhead:1997:ILR

[MWW97] Daryl L. Moorhead, Craig F. Wolf, and Robert A. Wharton, Jr. Impact of light regimes on productivity patterns of benthic microbial mats in an Antarctic lake: a modeling study. Limnology and Oceanography, 42(7):1561–1569, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Moffett:1990:IHP

[MZ90] James W. Moffett and Oliver C. Zajiriou. An investigation of hydrogen peroxide chemistry in surface waters of Vineyard Sound with H_{2}^{1}8O_{2} and ^{1}8O_{2} . Limnology and Oceanography, 35(6): 1221–1229, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Moran:1997:RPF

[MZ97] Mary Ann Moran and Richard G. Zepp. Role of photoreactions in the formation of biologically labile compounds from dissolved organic matter. Limnology and Oceanography, 42(6):1307–1316, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Morris:1995:ASU

[MZW^{+}95] Donald P. Morris, Horatio Zagarese, Craig E. Williamson, Esteban G. Balseiro, Bruce R. Hargreaves, Beatriz Modenutti, Robert Moeller, and Claudia Queimalinos. The attenuation of solar UV radiation in lakes and the role of dissolved organic carbon. Limnology and Oceanography, 40(8):1381–1391, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Nelson:1997:DGP

[NB97] David M. Nelson and Mark A. Brzezinski. Diatom growth and productivity in an oligo-trophic midocean gyre: a 3-yr record from the Sargasso Sea near Bermuda. Limnology and Oceanography, 42 (3):473–486, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Neale:1998:IMP

[NCD98] Patrick J. Neale, John J. Cullen, and Richard F. Davis. Inhibition of marine photosynthesis by ultraviolet radiation: Variable sensitivity of phytoplankton in the Weddell–Scotia Confluence during the austral spring. Limnology and Oceanography, 43(3):433–448, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Nielsen:1990:DPS

[NCRS90] Lars Peter Nielsen, Peter Bondo Christensen, Niels Peter Revsbech, and Jan Sørensen. Denitrification and photosynthesis in stream sediment studied with microsensor and whole-core techniques. Limnology and Oceanography, 35(5):1135–1144, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Nyman:1991:CES

[ND91] J. A. Nyman and R. D. DeLaune. CO_{2} emission and soil Eh responses to different hydrological conditions in fresh, brackish, and saline marsh soils. Limnology and Oceanography, 36(7):1406– 1414, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Nielsen:1993:EUB

[NE93] Tom Nielsen and Nils G. A. Ekelund. Effect of UV-B radiation and humic substances on growth and motility of Gyrodinium aureolum. Limnology and Oceanography, 38(7):1570–1575, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Newbury:1999:FW

[New99] Robert Newbury. Fresh water. Limnology and Oceanography, 44 (2):476, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Flamarique:1993:FFB

[NFGP^{+}93] I. Novales-Flamarique, S. Griesbach, M. Parent, A. Cattaneo, and R. H. Peters. Fish foraging behavior changes plankton-nutrient relations in laboratory microcosms. Limnology and Oceanography, 38(2):290–298, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Neale:1991:RHI

[NHJ91] Patrick J. Neale, S. Ivan Heaney, and George H. M. Jaworski. Responses to high irradiance contribute to the decline of the spring diatom maximum. Limnology and Oceanography, 36(4):761–768, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Nicholls:1998:NNI

[Nic98] K. H. Nicholls. El Niño, ice cover, and Great Lakes phosphorus: Implications for climate warming. Limnology and Oceanography, 43(4):715–719, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Nielsen:1991:CZG

[Nie91] Torkel Gissel Nielsen. Contribution of zooplankton grazing to the decline of a Ceratium bloom. Limnology and Oceanography, 36(6): 1091–1106, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Neill:1997:AHA

[NJLW97] Craig Neill, Kenneth M. Johnson, Ernie Lewis, and Douglas W. R. Wallace. Accurate headspace analysis of fCO_{2} in discrete water samples using batch equilibration. Limnology and Oceanography, 42(8):1774–1783, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Nagata:1991:RDF

[NK91] Toshi Nagata and David L. Kirchman. Release of dissolved free and combined amino acids by bacterivorous marine flagellates. Limnology and Oceanography, 36(3):433–443, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Nielsen:1994:RZB

[NK94] Torkel Gissel Nielsen and Thomas Kiørboe. Regulation of zooplankton biomass and production in a temperate, coastal ecosystem. 2. Ciliates. Limnology and Oceanography, 39(3):508–519, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Nepf:1999:VSF

[NK99] Heidi M. Nepf and Evamaria W. Koch Koch. Vertical secondary flows in submersed plant-like arrays. Limnology and Oceanography, 46(4):1072–1080, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Norton:1991:ARC

[NKMO91] Stephen A. Norton, Jeffrey S. Kahl, Myron J. Mitchell, and Jeflrey S. Owen. Another reply to the comment by Stauffer. Limnology and Oceanography, 36(6):1271–1274, September 1991. CO-DEN LIOCAH. ISSN 0024-3590. See [Sta91].

Nojiri:1993:ECF

[NKT^{+}93] Yukihiro Nojiri, Minoru Kusakabe, Klaus Tietze, Jun ichi Hirabayashi, Hiroaki Sato, Yuji Sano, Hiroshi Shinohara, Thomas Njine, and Greg Tanyileke. An estimate of CO_{2} flux in Lake Nyos, Cameroon. Limnology and Oceanography, 38(4):739–752, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Nelson:1999:LBM

[NLSG99] Yarrow M. Nelson, Leonard W. Lion, Michael L. Shuler, and William C. Ghiorse. Lead binding to metal oxide and organic phases of natural aquatic biofilms. Limnology and Oceanography, 44(7):1715–1729, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Nightingale:1995:PCO

[NML95] P. D. Nightingale, G. Malin, and P. S. Liss. Production of chloroform and other low molecular-weight halocarbons by some species of macroalgae. Limnology and Oceanography, 40(4):680–689, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Niyogi:1999:IWS

[NML99] Dev K. Niyogi, Diane M. McKnight, and William M. Lewis, Jr. Influences of water and substrate quality for periphyton in a montane stream affected by acid mine drainage. Limnology and Oceanography, 44(3 (part 2)):804–809, May 1999. CODEN LI-OCAH. ISSN 0024-3590.

Nodder:1997:STS

[Nod97] S. D. Nodder. Short-term sediment trap fluxes from chatham rise, southwest Pacific Ocean. Limnology and Oceanography, 42 (4):777–783, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Nielsen:1996:RNM

[NRPRB96] Lars Peter Nielsen, Nils Risgaard-Petersen, Søren Rysgaard, and T. Henry Blackburn. Reply to the note by Middelburg et al. Limnology and Oceanography, 41(4):1845–1846, December 1996. CODEN LIOCAH. ISSN 0024-3590. See [MSH96a] and reply [MSH96b].

Nesbitt:1996:OPP

[NRR96] Leslie M. Nesbitt, Howard P. Riessen, and Charles W. Ramcharan. Opposing predation pressures and induced vertical migration responses in Daphnia. Limnology and Oceanography, 41(6):1306– 1311, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Nelson:1991:SRC

[NS91] David M. Nelson and Walker O. Smith, Jr. Sverdrup revisited: Critical depths, maximum chlorophyll levels, and the control of Southern Ocean productivity by the irradiance-mixing regime. Limnology and Oceanography, 36(8):1650–1661, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Nielsen:1993:PSS

[NS93] Marianne V. Nielsen and Egil Sakshaug. Photobiological studies of Skeletonema costatum adapted to spectrally different light regimes. Limnology and Oceanography, 38(7):1576–1581, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Nikora:1998:EMF

[NSBB98] Vladimir I. Nikora, Alastair M. Suren, Stephanie L. R. Brown, and Barry J. F. Biggs. The effects of the moss Fissidens rigidulus (Fissidentaceae: Musci) on near-bed flow structure in an experimental cobble bed flume. Limnology and Oceanography, 43(6): 1321–1331, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Nielsen:1990:ASM

[NSJ90] Smen Laurentius Nielsen and Kaj Sand-Jensen. Allometric settling of maximal photosynthetic growth rate to surface/volume ratio. Limnology and Oceanography, 35(1):177–180, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Nuccio:1995:BPF

[NSK95] Joseph Nuccio, Pamela J. Seaton, and Robert J. Kieber. Biological production of form aldehyde in the marine environment. Limnology and Oceanography, 40(3):521–527, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Nepf:1997:MDW

[NSZ97] H. M. Nepf, J. A. Sullivan, and R. A. Zavistoski. A model for diffusion within emergent vegetation. Limnology and Oceanography, 42(8):1735–1745, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Najjar:1993:RCJ

[NT93a] Raymond G. Najjar and J. R. Toggweiler. Reply to the comment by Jackson. Limnology and Oceanography, 38(6):1331–1332, September 1993. CODEN LIOCAH. ISSN 0024-3590. See [Jac93].

Nygaard:1993:BAS

[NT93b] Kari Nygaard and August Tobiesen. Bacterivory in algae: a survival strategy during nutrient limitation. Limnology and Oceanography, 38(2):273–279, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Navarro:1994:CED

[NT94] J. M. Navarro and R. J. Thompson. Comparison and evaluation of different techniques for measuring absorption efficiency in suspension feeders. Limnology and Oceanography, 39(1):159–164, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Neale:1991:LTS

[NTH^{+}91] Patrick J. Neale, Jack F. Talling, S. Ivan Heaney, Colin S. Reynolds, and John W. G. Lund. Long time series from the English Lake District: Irradiance-dependent phytoplankton dynamics during the spring maximum. Limnology and Oceanography, 36 (4):751–760, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Nurnberg:1995:QAL

[Nür95] Gertrud K. Nürnberg. Quantifying anoxia in lakes. Limnology and Oceanography, 40(6):1100–1111, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Nurnberg:1998:PAS

[Nür98] Gertrud K. Nürnberg. Prediction of annual and seasonal phosphorus concentrations in stratified and polymictic lakes. Limnology and Oceanography, 43(7):1544–1552, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Nurnberg:1999:DTS

[Nür99] Gertrud K. Nürnberg. Determining trophic state in experimental lakes. Limnology and Oceanography, 46(4):1176–1179, June 1999. CODEN LIOCAH. ISSN 0024-3590. See reply [CKCP99].

Nedoma:1994:ADF

[NVH^{+}94] Jiří Nedoma, Jaroslav Vrba, Josef Hejzlar, Karel Šimek, and Věra Straškrabová. N-acetylglucosamine dynamics in freshwater environments: Concentration of amino sugars, extracellular enzyme activities, and microbial uptake. Limnology and Oceanography, 39(5):1088–1100, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Novak:1994:SDE

[NWS94] Martin Novák, R. Kelman Wieder, and William R. Schell. Sulfur during early diagenesis in Sphagnum peat: Insights from δ ^{34}S ratio profiles in ^{21}Pb-dated peat cores. Limnology and Oceanography, 39(5):1172–1185, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Norrman:1995:PUD

[NZHB95] Bo Norrman, Ulla Li Zweifel, Charles S. Hopkinson, Jr., and Fry Brian. Production and utilization of dissolved organic carbon during an experimental diatom bloom. Limnology and Oceanography, 40(5):898–907, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Olsson:1999:CUE

[OAF^{+}99] Kristina Olsson, Leif G. Anderson, Markus Frank, Anna Luchetta, and William Smethie. Carbon utilization in the Eurasian sector of the Arctic Ocean. Limnology and Oceanography, 44(1):95–105, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Ohman:1996:ICF

[OAR96] Mark D. Ohman, Dag L. Aksnes, and Jeffrey A. Runge. The interrelationship of copepod fecundity and mortality. Limnology and Oceanography, 41(7):1470–1477, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Opsahl:1998:PRD

[OB98] Stephen Opsahl and Ronald Benner. Photochemical reactivity of dissolved lignin in river and ocean waters. Limnology and Oceanography, 43(6):1297–1304, September 1998. CODEN LI-OCAH. ISSN 0024-3590.

Opsahl:1999:MFT

[OBA99] Stephen Opsahl, Ronald Benner, and Rainer M. W. Amon. Major flux of terrigenous dissolved organic matter through the Arctic Ocean. Limnology and Oceanography, 44(8):2017–2023, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Overmann:1991:CDP

[OBH^{+}91] Jörg Overmann, J. Thomas Beatty, Ken J. Hall, Norbert Pfennig, and Tom G. Northcote. Characterization of a dense, purple sulfur bacterial layer in a meromictic salt lake. Limnology and Oceanography, 36(5):846–859, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Overmann:1996:SCC

[OBKH96] Jörg Overmann, J. Thomas Beatty, H. Roy Krause, and Ken J. Hall. The sulfur cycle in the chemocline of a meromictic salt lake. Limnology and Oceanography, 41(1):147–156, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Orem:1991:JLP

[OBL^{+}91] W. H. Orem, W. C. Burnett, W. M. Landing, W. B. Lyons, and W. Showers. Jellyfish Lake, Palau: Early diagenesis of organic matter in sediments of an anoxic marine lake. Limnology and Oceanography, 36(3):526–543, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Osgood:1997:ODA

[OC97] Kenric E. Osgood and David M. Checkley, Jr. Observations of a deep aggregation of Calanus pacificus in the Santa Barbara Basin. Limnology and Oceanography, 42(5):997–1001, July 1997. CO-DEN LIOCAH. ISSN 0024-3590.

Overmann:1992:ELL

[OCP92] Jörg Overmann, Heribert Cypionka, and Norbert Pfennig. An extremely low-light adapted phototrophic sulfur bacterium from the Black Sea. Limnology and Oceanography, 37(1):150–155, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Olson:1996:PPC

[OCS96] Robert J. Olson, Alexander M. Chekalyuk, and Heidi M. Sosik. Phytoplankton photosynethetic characteristics from fluorescence induction assays of individual cells. Limnology and Oceanography, 41(6):1253–1263, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Olson:1990:PSD

[OCZA90] Robert J. Olson, Sallie W. Chisholm, Erik R. Zettler, and E. Virginia Armbrust. Pigments, size, and distributions of Synechococcus in the North Atlantic and Pacific Oceans. Limnology and Oceanography, 35(1):45–58, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Olaizola:1996:SSV

[OGH^{+}96] Miguel Olaizola, Richard J. Geider, William G. Harrison, Lisa M. Graziano, Giovanni M. Ferrari, and Peter M. Schlittenhardt. Synoptic study of variations in the fluorescence based maximum quantum efficiency of photosynthesis across the North Atlantic Ocean.

Limnology and Oceanography, 41(8):755–765, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Ohman:1996:BRA

[Ohm96] Mark D. Ohman. Book review: Atlas of marine zooplankton. Straits of Magellan: Copepods (M. G. Mazzocchi, G. Zagami, A. Ianora, L. Guglielmo, N. Crescenti, and J. Hure). Limnology and Oceanography, 41(7):1584–1585, November 1996. CODEN LIOCAH. ISSN 0024-3590.

OSullivan:1991:MFI

[OHMK91] Daniel W. O’Sullivan, Alfred K. Hanson, William L. Miller, and Dana R. Kester. Measurement of Fe(II) in surface water of the equatorial Pacific. Limnology and Oceanography, 36(8):1727–1741, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Orr:1998:RBM

[OJ98] Philip T. Orr and Gary J. Jones. Relationship between microcystin production and cell division rates in nitrogen-limited Microcystis aeruginosa cultures. Limnology and Oceanography, 43 (7):1604–1614, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Oldham:1994:FRO

[Old94] Carolyn Oldham. A fast-response oxygen sensor for use on finescale and microstructure CTD profilers. Limnology and Oceanography, 39(8):1959–1966, December 1994. CODEN LIOCAH. ISSN 0024-3590.

ORiordan:1993:SCB

[OMk93] C. A. O’Riordan, S. G. Monismith, and J. R. koseff. A study of concentration boundary-layer formation over a bed of model bivalves. Limnology and Oceanography, 38(8):1712–1729, December 1993. CODEN LIOCAH. ISSN 0024-3590.

ORiordan:1995:EBE

[OMK95] C. A. O’Riordan, S. G. Monismith, and J. R. Koseff. The effect of bivalve excurrent jet dynamics on mass transfer in a benthic boundary layer. Limnology and Oceanography, 40(2):330–344, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Orellana:1992:IMR

[OP92] Mónica V. Orellana and Mary Jane Perry. An immunoprobe to measure Rubisco concentrations and maximal photosynthetic rates of individual phytoplankton cells. Limnology and Oceanography, 37(3):478–490, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Ohman:1994:SFW

[OR94] Mark D. Ohman and Jefrey A. Runge. Sustained fecundity when phytoplankton resources are in short supply: Omnivory by Calanus finmarchicus in the Gulf of St. Lawrence. Limnology and Oceanography, 39(1):21–36, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Obernosterer:1999:CES

[ORH99] Ingrid Obernosterer, Bettina Reitner, and Gerhard J. Herndl. Contrasting effects of solar radiation on dissolved organic matter and its bioavailability to marine bacterioplankton. Limnology and Oceanography, 44(7):1645–1654, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Ohman:1991:GKO

[OS91] Mark D. Ohman and Richard A. Snyder. Growth kinetics of the omnivorous oligotrich ciliate strombidium sp. Limnology and Oceanography, 36(5):922–935, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Oviatt:1999:DAB

[Ovi99] Candace Oviatt. Dynamic aquaria. Building living ecosystems. Limnology and Oceanography, 44(6):1598, September 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Ohman:1996:MEP

[OW96] Mark D. Ohman and Simon N. Wood. Mortality estimation for planktonic copepods: Pseudocalanus newmani in a temperate fjord. Limnology and Oceanography, 41(1):126–135, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Ostrovsky:1996:SIM

[OYWK96] I. Ostrovsky, Y. Z. Yacobi, P. Walline, and I. Kalikhman. Seicheinduced mixing: Its impact on lake productivity. Limnology and Oceanography, 41(2):323–332, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Olson:1995:PFC

[OZ95] Robert J. Olson and Erik R. Zettler. Potential of flow cytometry for “pump and probe” fluorescence measurements of phytoplankton photosynthetic characteristics. Limnology and Oceanography, 40(4):816–820, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Passow:1995:DBA

[PA95a] U. Passow and A. L. Alldredge. A dye-binding assay for the spectrophotometric measurement of transparent exopolymer particles (TEP). Limnology and Oceanography, 40(7):1326–1335, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Pedros-Alio:1995:BRM

[PA95b] Carlos Pedrós-Alió. Book review: Microbial ecology of Lake Plußee (J. Overbeck and R. J. Chróst [ed.]). Limnology and Oceanography, 40(7):1345, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Paasche:1995:BRC

[Paa95] E. Paasche. Book review: Coccolithophores (A. Winter and W. G. Siesser [eds.]). Limnology and Oceanography, 40(3):645–646, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Planas:1999:NUD

[PAD^{+}99] Dolors Planas, Susana Agustí, Carlos M. Duarte, Tim C. Granata, and Martín Merino. Nitrate uptake and diffusive nitrate supply in the Central Atlantic. Limnology and Oceanography, 44(1):116– 126, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1997:CEH

[Pae97] Hans W. Paerl. Coastal eutrophication and harmful algal blooms: Importance of atmospheric deposition and groundwater as “new” nitrogen and other nutrient sources. Limnology and Oceanography, 42(5 (part 2)):1154–1165, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Paine:1996:BRB

[Pai96] R. T. Paine. Book review: The biology of rocky shores (C. Little and J. A. Kitching). Limnology and Oceanography, 41(7):1583– 1584, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Palmer:1992:ILM

[Pal92] Margaret A. Palmer. Incorporating lotic meiofauna into our understanding of faunal transport processes. Limnology and Oceanography, 37(2):329–341, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Price:1994:EPO

[PAM94] N. M. Price, B. A. Ahner, and F. M. M. Morel. The equatorial Pacific Ocean: Grazer-controlled phytoplankton populations in an iron-limited ecosystem. Limnology and Oceanography, 39(3): 520–534, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Alio:1990:MDV

[PAS90] Carlos Pedrbs-Alió and Ma. Montserrat Sala. Microdistribution and diel vertical migration of flagellated vs. gas-vacuolate purple sulfur bacteria in a stratified water body. Limnology and Oceanography, 35(7):1637–1644, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Phlips:1995:RBL

[PASC95] E. J. Phlips, F. J. Aldridge, C. L. Schelske, and T. L. Crisman. Relationships between light availability, chlorophyll a, and tripton in a large, shallow subtropical lake. Limnology and Oceanography, 40(2):416–421, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Pawlik:1998:CRS

[Paw98] Joseph R. Pawlik. Coral reef sponges: Do predatory fishes affect their distribution? Limnology and Oceanography, 43(6):1396– 1399, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Peng:1991:FLR

[PB91] Tsung-Hung Peng and Wallace S. Broecker. Factors limiting the reduction of atmospheric CO_{2} by iron fertilization. Limnology and Oceanography, 36(8):1919–1927, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Pawlik:1993:SMT

[PB93] Joseph R. Pawlik and Cheryl Ann Butman. Settlement of a marine tube worm as a function of current velocity: Interacting effects of hydrodynamics and behavior. Limnology and Oceanography, 38 (8):1730–1740, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Pakulski:1994:ADC

[PB94] J. Dean Pakulski and Ronald Benner. Abundance and distribution of carbohydrates in the ocean. Limnology and Oceanography, 39 (4):930–940, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1993:MCD

[PBJD93] Hans W. Paerl, Brad M. Bebout, Samantha B. Joye, and David J. Des Marais. Microscale characterization of dissolved organic matter production and uptake in marine microbial mat communities. Limnology and Oceanography, 38(6):1150–1161, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Pedersen:1999:EED

[PBL99a] Anna-Grethe U. Pedersen, Jørgen Berntsen, and Bente Aa. Lomstein. The effect of eelgrass decomposition on sediment carbon and nitrogen cycling: a controlled laboratory experiment. Limnology and Oceanography, 44(8):1978–1992, December 1999. CODEN LI-OCAH. ISSN 0024-3590.

Phlips:1999:BPC

[PBL99b] Edward J. Phlips, Susan Badylak, and Tammy C. Lynch. Blooms of the picoplanktonic cyanobacterium Synechococcus in Florida Bay, a subtropical inner-shelf lagoon. Limnology and Oceanography, 46(4):1166–1175, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Pace:1996:RBR

[PC96] Michael L. Pace and Jonathan J. Cole. Regulation of bacteria by resources and predation tested in whole-lake experiments. Limnology and Oceanography, 41(7):1448–1460, November 1996. CO-DEN LIOCAH. ISSN 0024-3590.

Pullin:1997:PVD

[PC97] Michael J. Pullin and Stephen E. Cabaniss. Physicochemical variations in DOM-synchronous fluorescence: Implications for mixing studies. Limnology and Oceanography, 42(8):1766–1773, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Post:1997:SEV

[PCC^{+}97] David M. Post, Stephen R. Carpenter, David L. Christensen, Kathryn L. Cottingham, James F. Kitchell, Daniel E. Schindler, and James R. Hodgson. Seasonal effects of variable recruitment of a dominant piscivore on pelagic food web structure. Limnology and Oceanography, 42(4):722–729, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Patterson:1997:TES

[PCFS97] Mark R. Patterson, Valeri I. Chernykh, Vladimir A. Fialkov, and Michael Savarese. Trophic effects of sponge feeding within Lake Baikal’s littoral zone. 1. In situ pumping rates. Limnology and Oceanography, 42(1):171–178, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Patrick:1990:SAS

[PD90a] Wm. H. Patrick, Jr. and R. D. DeLaune. Subsidence. accretion. and sea level rise in south San Francisco Bay marshes. Limnology and Oceanography, 35(6):1389–1395, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Peterson:1990:ICS

[PD90b] William Peterson and Hans G. Dam. he influence of copepod “swimmers” on pigment fluxes in brine-filled vs. ambient seawaterfilled sediment traps. Limnology and Oceanography, 35(2):448– 455, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Prisu:1996:ESN

[PDM96] John C. Prisu, Malcolm T. Downes, and Christopher P. McKay. Extreme supersaturation of nitrous oxide in a poorly ventilated Antarctic lake. Limnology and Oceanography, 41(7):1544–1551, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Persson:1997:PRF

[Per97] Anders Persson. Phosphorus release by fish in relation to external and internal load in a eutrophic lake. Limnology and Oceanography, 42(3):577–583, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Peters:1994:PPP

[Pet94] Francesc Peters. Prediction of planktonic protistan grazing rates. Limnology and Oceanography, 39(1):195–206, January 1994. CO-DEN LIOCAH. ISSN 0024-3590.

Penry:1990:REG

[PF90] Deborah L. Penry and Bruce W. Frost. Re-evaluation of the gutfullness (gut fluorescence) method for inferring ingestion rates of suspension-feeding copepods. Limnology and Oceanography, 35 (5):1207–1214, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Penry:1991:CDC

[PF91] Deborah L. Penry and Bruce W. Frost. Chlorophyll a degradation by Calanus pacificus: Dependence on ingestion rate and digestive acclimation to food resources. Limnology and Oceanography, 36 (1):147–158, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1996:SNF

[PFB96] Hans W. Paerl, Matthew Fitzpatrick, and Brad M. Bebout. Seasonal nitrogen fixation dynamics in a marine microbial mat: Potential roles of cyanobacteria and microheterotrophs. Limnology and Oceanography, 41(3):419–427, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Peterson:1993:TSE

[PFDH93] Bruce Peterson, Brian Fry, Linda Deegan, and Anne Hershey. The trophic significance of epilithic algal production in a fertilized tundra river ecosystem. Limnology and Oceanography, 38(4):872– 878, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Pandolfi:1997:PCS

[PG97] John M. Pandolfi and Benjamin J. Greenstein. Preservation of community structure in death assemblages of deep-water Caribbean reef corals. Limnology and Oceanography, 42(7):1505– 1516, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Peduzzi:1992:ZAF

[PH92a] Peter Peduzzi and Gerhard J. Herndl. Zooplankton activity fueling the microbial loop: Differential growth response of bacteria from oligotrophic and eutrophic waters. Limnology and Oceanography, 37(5):1087–1092, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Prasad:1992:QYE

[PH92b] K. S. Prasad and J. T. Hollibaugh. Quantum yield estimates of phytoplankton on the Grand Banks for use in production models. Limnology and Oceanography, 37(6):1271–1279, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Palenik:1997:UAO

[PH97] Brian Palenik and Sarah E. Henson. The use of amides and other organic nitrogen sources by the phytoplankton Emiliania huxleyi. Limnology and Oceanography, 42(7):1544–1551, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Padilla:1996:IAI

[PHMH96] Dianna K. Padilla, C. Drew Harvell, Jessica Marks, and Brian Helmuth. Inducible aggression and intraspecific competition for space in a marine bryozoan, Membranipora membranacea. Limnology and Oceanography, 41(3):505–512, May 1996. CODEN LI-OCAH. ISSN 0024-3590.

Peterson:1993:FEV

[PHT^{+}93] M. L. Peterson, P. J. Hernes, D. S. Thoreson, J. I. Ifedges, C. Lee, and S. G. Wakeham. Field evaluation of a valved sediment trap. Limnology and Oceanography, 38(8):1741–1761, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Prahl:1992:DIT

[PHX92] F. G. Prahl, J. M. Hayes, and T.-M. Xie. Diploptene: an indicator of terrigenous organic carbon in Washington coastal sediments. Limnology and Oceanography, 37(6):1290–1299, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Pick:1991:ACF

[Pic91] F. R. Pick. The abundance and composition of freshwater picocyanobacteria in relation to light penetration. Limnology and Oceanography, 36(7):1457–1462, November 1991. CODEN LI-OCAH. ISSN 0024-3590.

Pineda:1999:CLD

[Pin99] Jesús Pineda. Circulation and larval distribution in internal tidal bore warm fronts. Limnology and Oceanography, 44(6):1400–1414, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Park:1990:MFE

[PJ90] J. K. Park and A. James. Mass flux estimation and mass transport mechanism in estuaries. Limnology and Oceanography, 35(6): 1301–1313, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Pilskaln:1991:SSL

[PJ91] Cynthia H. Pilskaln and Thomas C. Johnson. Seasonal signals in Lake Malawi sediments. Limnology and Oceanography, 36(3): 544–557, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Poiani:1996:CCN

[PJSW96] Karen A. Poiani, W. Carter Johnson, George A. Swanson, and Tom C. Winter. Climate change and northern prairie wetlands: Simulations of long-term dynamics. Limnology and Oceanography, 41(5):871–881, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Petticrew:1991:PSS

[PK91] Ellen L. Petticrew and Jacob Kalff. Predictions of surficial sediment composition in the littoral zone of lakes. Limnology and Oceanography, 36(2):384–392, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Peterson:1994:PCR

[PK94] William T. Peterson and William J. Kimmerer. Processes controlling recruitment of the marine Calanoid copepod Temora longicornis in Long Island Sound: Egg production, egg mortality, and cohort survival rates. Limnology and Oceanography, 39(7):1594– 1605, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Parmenter:1991:NCF

[PL91] Robert R. Parmenter and Vincent A. Lamarra. Nutrient cycling in a freshwater marsh: The decomposition of fish and waterfowl carrion. Limnology and Oceanography, 36(5):976–987, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Pantoja:1994:CSO

[PL94] Silvio Pantoja and Cindy Lee. Cell-surface oxidation of amino acids in seawater. Limnology and Oceanography, 39(7):1718–1726, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Portielje:1995:CDF

[PL95] Robert Portielje and Lambertus LiJklema. Carbon dioxide fluxes across the air–water interface and its impact on carbon availability in aquatic systems. Limnology and Oceanography, 40(4):690–699, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Pantoja:1999:MWD

[PL99] Silvio Pantoja and Cindy Lee. Molecular weight distribution of proteinaceous material in Long Island Sound sediments. Limnology and Oceanography, 44(5):1323–1330, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Platt:1993:BRC

[Pla93] Trevor Platt. Book review: Concepts in biological oceanography: an interdisciplinary primer (P. A. Jumars). Limnology and Oceanography, 38(8):1842–1843, December 1993. CODEN LI-OCAH. ISSN 0024-3590.

Plath:1998:AFB

[Pla98] Klaus Plath. Adaptive feeding behavior of Duphnia magna in response to short-term starvation. Limnology and Oceanography, 43(4):593–599, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Porter:1999:EMS

[PLP99] James W. Porter, Sarah K. Lewis, and Karen G. Porter. The effect of multiple stressors on the Florida Keys coral reef ecosystem: a landscape hypothesis and a physiological test. Limnology and Oceanography, 44(3 (part 2)):941–949, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Painchaud:1996:BDU

[PLTL96] Jean Painchaud, Denis Lefaivre, Jean-Claude Therriault, and Louis Legendre. Bacterial dynamics in the upper St. Lawrence estuary. Limnology and Oceanography, 41(8):1610–1618, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Plumley:1997:MAT

[Plu97] F. Gerald Plumley. Marine algal toxins: Biochemistry, genetics, and molecular biology. Limnology and Oceanography, 42(5 (part 2)):1252–1264, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Palenik:1990:AAU

[PM90a] B. Palenik and F. M. M. Morel. Amino acid utilization by marine phytoplankton: a novel mechanism. Limnology and Oceanography, 35(2):260–269, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Pettine:1990:CSS

[PM90b] M. Pettine and Frank J. Millero. Chromium speciation in seawater: The probable role of hydrogen peroxide. Limnology and Oceanography, 35(3):730–736, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Price:1991:CPG

[PM91] N. M. Price and F. M. M. Morel. Colimitation of phytoplankton growth by nickel and nitrogen. Limnology and Oceanography, 36 (6):1071–1077, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Pace:1990:PCS

[PMF90] Michael L. Pace, George B. McManus, and Stuart E. G. Findlay. Planktonic community structure determines the fate of bacterial production in a temperate lake. Limnology and Oceanography, 35 (4):795–808, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Polishchuk:1990:CTS

[Pol90] Leonard K. Polishchuk. A comment on “The theory of secondary production calculations for continuously reproducing populations” (kimmerer). Limnology and Oceanography, 35(7):1645– 1651, November 1990. CODEN LIOCAH. ISSN 0024-3590. See reply [Kim90].

Polishchuk:1995:DPE

[Pol95] Leonard V. Polishchuk. Direct positive effect of invertebrate predators on birth rate in Daphnia studied with a new method of birth rate analysis. Limnology and Oceanography, 40(3):483–489, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Peierls:1997:BAO

[PP97] Benjamin L. Peierls and Hans W. Paerl. Bioavailability of atmospheric organic nitrogen deposition to coastal phytoplankton. Limnology and Oceanography, 42(8):1819–1823, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Paulsen:1991:EMD

[PPB91] Denise M. Paulsen, Hans W. Paerl, and Paul E. Bishop. Evidence that molybdenum-dependent nitrogen fixation is not limited by high sulfate concentrations in marine environments. Limnology and Oceanography, 36(7):1325–1334, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1995:CSF

[PPK95] Hans W. Paerl, James L. Pinckney, and Scott A. Kucera. Clarification of the structural and functional roles of heterocysts and anoxic microzones in the control of pelagic nitrogen fixation. Limnology and Oceanography, 40(3):634–638, May 1995. CODEN LI-OCAH. ISSN 0024-3590. See reply [HSM^{+}95].

Peeter:1996:DSN

[PPK^{+}96] F. Peeter, G. Piepke, R. Kipfer, R. Hohmann, and D. M. Imboden. Description of stability and neutrally buoyant transport in freshwater lakes. Limnology and Oceanography, 41(8):1711–1724, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Paul:1999:ECS

[PPK^{+}99] John H. Paul, Scott L. Pichard, Jordan B. Kang, Gregory M. F. Watson, and F. Robert Tabita. Evidence for a clade-specific temporal and spatial separation in ribulose bisphosphate carboxylase gene expression in phytoplankton populations off Cape Hatteras and Bermuda. Limnology and Oceanography, 44(1):12–23, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Pennings:1997:CDT

[PPP97] Steven C. Pennings, Sonia R. Pablo, and Valerie J. Paul. Chemical defenses of the tropical, benthic marine cyanobacterium Hormothamnion enteromorphoides: Diverse consumers and synergisms. Limnology and Oceanography, 42(5):911–917, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Pile:1997:TES

[PPS^{+}97] Adele J. Pile, Mark R. Patterson, Michael Savarese, Valeri I. Chernykh, and Vladimir A. Fialkov. Trophic effects of sponge feeding within Lake Baikal’s littoral zone. 2. Sponge abundance, diet, feeding efficiency, and carbon flux. Limnology and Oceanography, 42(1):178–184, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Prairie:1990:CNS

[Pra90] Yves T. Prairie. A comment on “Nutrient status and nutrient competition of phytoplankton in a shallow, hypertrophic lake” (Sommer). Limnology and Oceanography, 35(3):778–779, May 1990. CODEN LIOCAH. ISSN 0024-3590. See [Som89] and reply [Som90].

Polz:1998:TEM

[PRCV98] Martin F. Polz, Jonathan J. Robinson, Colleen M. Cavanaugh, and Cindy L. Van Dover. Trophic ecology of massive shrimp aggregations at a Mid-Atlantic Ridge hydrothermal vent site. Limnology and Oceanography, 43(7):1631–1638, November 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Priscu:1994:BRP

[Pri94] John C. Priscu. Book review: Physical and biogeochemical processes in Antarctic lakes (W. J. Green and E. I. Friedmann ([eds.])). Limnology and Oceanography, 39(6):1499–1500, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Probyn:1990:SFM

[Pro90] Trevor A. Probyn. Size-fractionated measurements of nitrogen uptake in aged upwelled waters: Implications for pelagic food webs. Limnology and Oceanography, 35(1):202–210, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Priede:1994:SCA

[PRS94] I. G. Priede, P. M. Ragley, and K. L. Smith, Jr. Seasonal change in activity of abyssal demersal scavenging grenadiers Coryphaenoides (Nematonums) armatus in the eastern North Pacific Ocean. Limnology and Oceanography, 39(2):279–285, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Pringle:1993:LLB

[PRT^{+}93] Catherine M. Pringle, Gary L. Rowe, Frank J. Triska, Jose F. Fernandez, and John West. Landscape linkages between geothermal activity and solute composition and ecological response in surface waters draining the Atlantic slope of Costa Rica. Limnology and Oceanography, 38(4):753–774, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Pahlow:1997:ICS

[PRWG97] Markus Pahlow, Ulf Riebesell, and Dieter A. Wolf-Gladrow. Impact of cell shape and chain formation on nutrient acquisition by marine diatoms. Limnology and Oceanography, 42(8):1660–1672, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Psenner:1992:RCB

[PS92] Roland Psenner and Ruben Sommaruga. Are rapid changes in bacterial biomass caused by shifts from top-down to bottom-up control? Limnology and Oceanography, 37(5):1092–1100, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Psenner:1990:IAC

[Pse90] Roland Psenner. From image analysis to chemical analysis of bacteria: a long-term study? Limnology and Oceanography, 35 (1):234–237, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Psenner:1992:BRA

[Pse92] Roland Psenner. Book review: Acidic deposition and aquatic ecosystems. Regional case studies (D. F. Charles [ed.]). Limnology and Oceanography, 37(6):1342–1343, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Ploug:1999:DBLa

[PSEJ99] Helle Ploug, Willem Stolte, Eric H. G. Epping, and Bo Barker Jørgensen. Diffusive boundary layers, photosynthesis, and respiration of the colony-forming plankton algae, phaeocystis sp. Limnology and Oceanography, 44(8):1949–1958, December 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Porter:1996:ACA

[PSH^{+}96] Karen G. Porter, Patricia A. Saunders, Kurt A. Haberyan, Alexander E. Macubbin, Timothy R. Jacobsen, and Robert E. Hodson. Annual cycle of autotrophic and heterotrophic production in a small, monomictic Piedmont lake (Lake Oglethorpe): Analog for the effects of climatic warming on dimictic lakes. Limnology and Oceanography, 41(5):1041–1051, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Ploug:1999:DBLb

[PSJ99] Helle Ploug, Willem Stolte, and Bo Barker Jørgensen. Diffusive boundary layers of the colony-forming plankton alga phaeocystis sp. — implications for nutrient uptake and cellular growth. Limnology and Oceanography, 44(8):1959–1967, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Patterson:1991:SMF

[PSO91] Mark R. Patterson, Kenneth P. Sebens, and R. Randolph Olson. In situ measurements of flow effects on primary production and dark respiration in reef corals. Limnology and Oceanography, 36 (5):936–948, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Prakash:1991:GVO

[PSS91] A. Prakash, R. W. Sheldon, and W. H. Sutcliffe, Jr. Geographic variation of oceanic ^{14}C dark uptake. Limnology and Oceanography, 36(1):30–39, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Peulve:1996:MCS

[PSS^{+}96] S. Peulvé, M.-A. Sicre, A. Saliot, J. W. De Leeuw, and M. Baas. Molecular characterization of suspended and sedimentary organic matter in an Arctic delta. Limnology and Oceanography, 41(3): 488–497, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Poff:1996:SHE

[PTJ96] N. LeRoy Poff, Sezin Tokar, and Peggy Johnson. Stream hydrological and ecological responses to climate change assessed with an artificial neural network. Limnology and Oceanography, 41(5): 857–863, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Pace:1994:IDD

[PV94] Michael L. Pace and Dolors Vaqué. The importance of Daphnia in determining mortality rates of protozoans and rotifers in lakes.

Limnology and Oceanography, 39(5):985–996, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Pennock:1996:IFA

[PVL^{+}96] Jonathan R. Pennock, David J. Velinsky, John M. Ludlam, Jonathan H. Sharp, and Marilyn L. Fogel. Isotopic fractionation of ammonium and nitrate during uptake by Skeletonema costatum: Implications for δ ^{15}N dynamics under bloom conditions. Limnology and Oceanography, 41(3):451–459, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Peduzzi:1993:ECV

[PW93] Peter Peduzzi and Markus G. Weinbauer. Effect of concentrating the virus-rich 2-2nm size fraction of seawater on the formation of algal flocs (marine snow). Limnology and Oceanography, 38(7): 1562–1565, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Paterson:1998:MMS

[PWM^{+}98] D. M. Paterson, K. H. Wiltshire, A. Miles, J. Blackburn, I. Davidson, M. G. Yates, S. McGrorty, and J. A. Eastwood. Microbiological mediation of spectral reflectance from intertidal cohesive sediments. Limnology and Oceanography, 43(6):1207–1221, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Purcell:1994:PGZ

[PWR94] Jennifer E. Purcell, Jacques R. White, and Michael R. Roman. Predation by gelatinous zooplankton and resource limitation as potential controls of Acartia tonsa copepod populations in Chesapeake Bay. Limnology and Oceanography, 39(2):263–278, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Pile:1999:PAR

[PY99] Adele J. Pile and Craig M. Young. Plankton availability and retention efficiencies of cold-seep symbiotic mussels. Limnology and Oceanography, 44(7):1833–1839, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Pinckney:1993:PRI

[PZ93] J. Pinckney and R. G. Zingmark. Photophysiological responses of intertidal benthic microalgal communities to in situ light environments: Methodological considerations. Limnology and Oceanography, 38(7):1373–1383, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Quiblier-Lloberas:1996:QSZ

[QLBAP96] Catherine Quiblier-Llobéras, Gilles Bourdier, Christian Amblard, and Denise Pepin. A qualitative study of zooplankton grazing in an oligo-mesotrophic lake using phytoplanktonic pigments as organic markers. Limnology and Oceanography, 41(8):1767–1779, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Quifiones:1991:RBR

[QP91] Renato A. Quifiones and Trevor Platt. The relationship between the f-ratio and the P : R ratio in the pelagic ecosystem. Limnology and Oceanography, 36(1):211–213, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Quay:1992:CCA

[QWR^{+}92] Paul D. Quay, David O. Wilbur, Jefrey E. Richey, John I. Hedges, Allan H. Devol, and Reynaldo Victoria. Carbon cycling in the Amazon River: Implications from the ^{13}C compositions of particles and solutes. Limnology and Oceanography, 37(4):857–871, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Quay:1995:DOR

[QWR^{+}95] P. D. Quay, D. O. Wilbur, J. E. Richey, A. H. Devol, R. Benner, and B. R. Forsberg. The^{18}O :^{16}O of dissolved oxygen in rivers and lakes in the Amazon Basin: Determining the ratio of respiration to photosynthesis rates in freshwaters. Limnology and Oceanography, 40(4):718–729, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Rivkin:1997:INL

[RA97] Richard B. Rivkin and M. Robin Anderson. Inorganic nutrient limitation of oceanic bacterioplankton. Limnology and Oceanography, 42(4):730–740, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Ricketts:1998:DCB

[RA98] Richard D. Ricketts and Robert F. Anderson. A direct comparison between the historical record of lake level and the δ ^{18}O signal in carbonate sediments from Lake Turkana, Kenya. Limnology and Oceanography, 43(5):811–822, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Rahel:1990:ATO

[Rah90] Frank J. Rahel. Anomalous temperature and oxygen gradients under the ice of a high-plains lake in Wyoming. Limnology and Oceanography, 35(3):751–755, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Raven:1995:BRS

[Rav95] John A. Raven. Book review: Stable isotopes in ecology and environmental science (K. Lathja and R. H. Michener [eds.]). Limnology and Oceanography, 40(3):644–645, May 1995. CODEN LI-OCAH. ISSN 0024-3590.

Raven:1997:PP

[Rav97] J. A. Raven. Phagotrophy in phototrophs. Limnology and Oceanography, 42(1):198–205, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Reid:1991:IMI

[RB91] Richard T. Reid and Alison Butler. Investigation of the mechanism of iron acquisition by the marine bacterium Alteromonas luteoviolaceus: Characterization of siderophore production. Limnology and Oceanography, 36(8):1783–1792, December 1991. CO-DEN LIOCAH. ISSN 0024-3590.

Runck:1994:RBB

[RB94] Clay Runck and Dean W. Blinn. Role of Belostoma bakeri (Heteroptera) in the trophic ecology of a fishless desert spring. Limnology and Oceanography, 39(8):1800–1812, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Rue:1997:ROC

[RB97] Eden L. Rue and Kenneth W. Bruland. The role of organic complexation on ambient iron chemistry in the equatorial Pacific Ocean and the response of a mesoscale iron addition experiment. Limnology and Oceanography, 42(5):901–910, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Romare:1999:ILJ

[RBH99] Pia Romare, Eva Bergman, and Lars-Anders Hansson. The impact of larval and juvenile fish on zooplankton and algal dynamics. Limnology and Oceanography, 44(7):1655–1666, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Roberts:1991:PTD

[RBMH91] D. Roberts, D. S. M. Billett, G. McCartney, and G. E. Hayes. Procaryotes on the tentacles of deep-sea holothurians: a novel form of dietary supplementation. Limnology and Oceanography, 36(7):1447–1451, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Robinson:1993:ESM

[RC93] S. M. C. Robinson and R. A. Chandler. An effective and safe method for sorting small molluscs from sediment. Limnology and Oceanography, 38(5):1088–1091, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Robinson:1995:EFF

[RC95] Jonathan J. Robinson and Colleen M. Cavanaugh. Expression of form I and form II Rubisco in chemoautotrophic symbioses: Implications for the interpretation of stable carbon isotope values. Limnology and Oceanography, 40(8):1496–1502, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Rothfuss:1998:EGB

[RC98] Franz Rothfuss and Ralf Conrad. Effect of gas bubbles on the diffusive flux of methane in anoxic paddy soil. Limnology and Oceanography, 43(7):1511–1518, November 1998. CODEN LI-OCAH. ISSN 0024-3590.

Roland:1999:RPD

[RCCdG99] Fabio Roland, Nina F. Caraco, Jonathan J. Cole, and Paul del Giorgio. Rapid and precise determination of dissolved oxygen by spectrophotometry: Evaluation of interference from color and turbidity. Limnology and Oceanography, 46(4):1148–1154, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Ribes:1998:HFG

[RCG98] Marta Ribes, Rafel Coma, and Josep-Maria Gili. Heterotrophic feeding by gorgonian corals with symbiotic zooxanthella. Limnology and Oceanography, 43(6):1170–1179, September 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Rich:1996:CUN

[RDK96] James H. Rich, Hugh W. Ducklow, and David L. Kirchman. Concentrations and uptake of neutral monosaccharides along 14^{o}W in the equatorial Pacific: Contribution of glucose to heterotrophic bacterial activity and the DOM flux. Limnology and Oceanography, 41(4):595–604, June 1996. CODEN LIOCAH. ISSN 0024-

Reed:1990:YLO

[Ree90] R. K. Reed. A year-long observation of water exchange between the North Pacific and the Bering Sea. Limnology and Oceanography, 35(7):1604–1609, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Regnell:1997:FCT

[REL97] Olof Regnell, Göran Ewald, and Elsmari Lord. Factors controlling temporal variation in methyl mercury levels in sediment and water in a seasonally stratified lake. Limnology and Oceanography, 42 (8):1784–1795, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Ricci:1991:AWI

[RERB91] N. Ricci, F. Erra, A. Russo, and R. Banchetti. The air–water interface: a microhabitat for hypotrichous settlers (Protista, Ciliata). Limnology and Oceanography, 36(6):1178–1188, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Reinfelder:1994:AOI

[RF94a] John R. Reinfelder and Nicholas S. Fisher. The assimilation of elements ingested by marine planktonic bivalve larvae. Limnology and Oceanography, 39(1):12–20, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Reinfelder:1994:REA

[RF94b] John R. Reinfelder and Nicholas S. Fisher. Retention of elements absorbed by juvenile fish (Menidia menidia, Menidia beryllina) from zooplankton prey. Limnology and Oceanography, 39(8):1783– 1789, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Ringelberg:1994:IPF

[RF94c] J. Ringelberg and B. J. G. Flik. Increased phototaxis in the field leads to enhanced diel vertical migration. Limnology and Oceanography, 39(8):1855–1864, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Roditi:1999:RRT

[RF99] Hudson A. Roditi and Nicholas S. Fisher. Rates and routes of trace element uptake in zebra mussels. Limnology and Oceanography, 44(7):1730–1749, November 1999. CODEN LIOCAH. ISSN 0024-

Ruiz:1999:NIS

[RFHG99] Gregory M. Ruiz, Paul Fofonoff, Anson H. Hines, and Edwin D. Grosholz. Non-indigenous species as stressors in estuarine and marine communities: Assessing invasion impacts and interactions. Limnology and Oceanography, 44(3 (part 2)):950–972, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Ritzrau:1992:IMB

[RG92a] Will Ritzrau and Gerhard Graf. Increase of microbial biomass in the benthic turbidity zone of Kiel Bight after resuspension by a storm event. Limnology and Oceanography, 37(5):1081–1086, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Rothhaupt:1992:IST

[RG92b] Karl O. Rothhaupt and Hans Güde. The influence of spatial and temporal concentration gradients on phosphate partitioning between different size fractions of plankton: Further evidence and possible causes. Limnology and Oceanography, 37(4):739–749, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Roman:1997:CGE

[RG97] Michael R. Roman and Anne L. Gauzens. Copepod grazing in the equatorial Pacific. Limnology and Oceanography, 42(4):623–634, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Ross:1993:SSM

[RGH^{+}93] Alex H. Ross, William S. C. Gurney, Michael R. Health, Steven J. Hay, and Eric W. Henderson. A strategic simulation model of a fjord ecosystem. Limnology and Oceanography, 38(1):128–153, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Ross:1994:CSE

[RGH94] Alex H. Ross, William S. C. Gurney, and Michael R. Heath. A comparative study of the ecosystem dynamics of four fiords. Limnology and Oceanography, 39(2):318–343, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Rabouille:1998:ROM

[RGR^{+}98] Christophe Rabouille, Jean-Francois Gaillard, Jean-Claude Relexans, Paul Tréguer, and Marie-Anne Vincendeau. Recycling of organic matter in Antarctic sediments: a transect through the polar front in the Southern Ocean (Indian Sector). Limnology and Oceanography, 43(3):420–432, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Roman:1993:ELO

[RGRW93] Michael R. Roman, Anne L. Gauzens, W. Kirk Rhinehart, and Jacques R. White. Effects of low oxygen waters on Chesapeake Bay zooplankton. Limnology and Oceanography, 38(8):1603–1614, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Renberg:1993:PFC

[RH93] Ingemar Renberg and Hans Hansson. A pump freeze corer for recent sediments. Limnology and Oceanography, 38(6):1317–1321, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Richey:1990:BCA

[RHD^{+}90] Jeffrey E. Richey, John I. Hedges, Allan H. Devol, Paul D. Quay, Reynaldo Victoria, Luiz Martinelli, and Bruce R. Forsberg. Biogeochemistry of carbon in the Amazon River. Limnology and Oceanography, 35(2):352–371, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Reitner:1997:RUB

[RHH97] Bettina Reitner, Gerhard J. Herndl, and Alois Herzig. Role of ultraviolet-B radiation on photochemical and microbial oxygen consumption in a humic-rich shallow lake. Limnology and Oceanography, 42(5):950–960, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Reynolds:1992:RPA

[RHT92] G. L. Reynolds and J. Hamilton-Taylor. The role of planktonic algae in the cycling of Zn and Cu in a productive soft-water lake. Limnology and Oceanography, 37(8):1759–1769, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Riebesell:1992:FLM

[Rie92] Ulf Riebesell. The formation of large marine snow and its sustained residence in surface waters. Limnology and Oceanography, 37(1):63–76, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Raven:1991:MIC

[RJ91] John A. Raven and Andrew M. Johnston. Mechanisms of inorganic-carbon acquisition in marine phytoplankton and their implications for the use of other resources. Limnology and Oceanography, 36(8):1701–1714, December 1991. CODEN LI-OCAH. ISSN 0024-3590.

Rybicki:1997:OTF

[RJC^{+}97] Nancy B. Rybicki, Harry L. Jenter, Virginia Carter, Robert A. Baltzer, and Michael Turtora. Observations of tidal flux between a submersed aquatic plant stand and the adjacent channel in the Potomac River near Washington, D.C. Limnology and Oceanography, 42(2):307–317, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Rahel:1996:PHL

[RKA96] Frank J. Rahel, Christopher J. Keleher, and Jeffry L. Anderson. Potential habitat loss and population fragmentation for cold water fish in the North Platte River drainage of the Rocky Mountains: Response to climate warming. Limnology and Oceanography, 41 (5):1116–1123, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Rappe:1997:PDM

[RKG97] Michael S. Rappé, Paul F. Kemp, and Stephen J. Giovannoni. Phylogenetic diversity of marine coastal picoplankton ^{16}S rRNA genes cloned from the continental shelf off Cape Hatteras, North Carolina. Limnology and Oceanography, 42(5):811–826, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Richards:1991:OVS

[RKR91] S. R. Richards, C. A. Kelly, and John W. M. Rudd. Organic volatile sulfur in lakes of the Canadian Shield and its loss to the atmosphere. Limnology and Oceanography, 36(3):468–482, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Rudd:1990:CAE

[RKST90] John W. M. Rudd, C. A. Kelly, D. W. Schindler, and M. A. Turner. A comparison of the acidification efficiencies of nitric and sulfuric acids by two whole-lake addition experiments. Limnology and Oceanography, 35(3):663–679, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Reinikainen:1994:ECT

[RKW94] Marko Reinikainen, Matti Ketola, and Mari Walls. Effects of the concentrations of toxic Microcystis aeruginosa and an alternative food on the survival of Daphnia pulex. Limnology and Oceanography, 39(2):424–432, March 1994. CODEN LIOCAH. ISSN 0024-

Rogers:1995:ONS

[RLH95] Christopher K. Rogers, Gregory A. Lawrence, and Paul F. Hamblin. Observations and numerical simulation of a shallow icecovered midlatitude lake. Limnology and Oceanography, 40(2): 374–385, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Riemann:1993:TCD

[RLL93] Bo Riemann, Risto Lignell, and Edward Laws. Time-course development of ^{14}C specificactivity of chlorophyll-a, carbon, and proteins in algal cultures. Limnology and Oceanography, 38(1): 96–111, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Reimers:1990:SSV

[RLTB90] Clare E. Reimers, Carina B. Lange, Marialana Tabak, and Joan M. Bernhard. Seasonal spillover and varve formation in the Santa Barbara Basin, California. Limnology and Oceanography, 35(7):1577–1585, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Rich:1990:AWD

[RM90] Hans Walter Rich and François M. M. Morel. Availability of welldefined iron colloids to the marine diatom Thalassiosira weissflogii. Limnology and Oceanography, 35(3):652–662, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Ridal:1992:DOP

[RM92] Jefrey J. Ridal and Robert M. Moore. Dissolved organic phosphorus concentrations in the northeast subarctic Pacific Ocean. Limnology and Oceanography, 37(5):1067–1075, July 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Romero:1996:SVM

[RM96] José R. Romero and John M. Melack. Sensitivity of vertical mixing in a large saline lake to variations in runoff. Limnology and Oceanography, 41(5):955–965, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Rivera-Monroy:1996:RRD

[RMT96] Victor H. Rivera-Monroy and Robert R. Twilley. The relative role of denitrification and immobilization in the fate of inorganic nitrogen in mangrove sediments (Terminos Lagoon, Mexico). Limnology and Oceanography, 41(2):284–296, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Revsbech:1998:NMD

[RNR98] Niels Peter Revsbech, Lars Peter Nielsen, and Niels Birger Ramsing. A novel microsensor for determination of apparent diffusivity in sediments. Limnology and Oceanography, 43(5):986–992, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Rocha:1998:RAR

[Roc98] Carlos Rocha. Rhythmic ammonium regeneration and flushing in intertidal sediments of the Sado estuary. Limnology and Oceanography, 43(5):823–831, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Roesler:1998:TEA

[Roe98] Collin S. Roesler. Theoretical and experimental approaches to improve the accuracy of particulate absorption coefficients derived from the quantitative filter technique. Limnology and Oceanography, 43(7):1649–1660, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Roman:1991:PCI

[Rom91] Michael R. Roman. Pathways of carbon incorporation in marine copepods: Effects of developmental stage and food quantity. Limnology and Oceanography, 36(4):796–807, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Rothhaupt:1990:CFR

[Rot90a] Karl O. Rothhaupt. Changes of the functional responses of the rotifers Brachionus rubens and Brachionus calyciflorus with particle sizes. Limnology and Oceanography, 35(1):24–32, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Rothhaupt:1990:DPS

[Rot90b] Karl O. Rothhaupt. Differences in particle size-dependent feeding efficiencies of closely related rotifer species. Limnology and Oceanography, 35(1):16–23, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Rothhaupt:1991:VZM

[Rot91] Karl Otto Rothhaupt. Variations on the zooplankton menu: a reply to the comment by Smith. Limnology and Oceanography, 36 (4):824–827, June 1991. CODEN LIOCAH. ISSN 0024-3590. See [Smi91].

Rothhaupt:1992:SPL

[Rot92] Karl O. Rothhaupt. Stimulation of phosphorus-limited phytoplankton by bacterivorous flagellates in laboratory experiments. Limnology and Oceanography, 37(4):750–759, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Rothhaupt:1995:ANL

[Rot95] Karl O. Rothhaupt. Algal nutrient limitation affects rotifer growth rate but not ingestion rate. Limnology and Oceanography, 40(7): 1201–1208, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Petersen:1997:NDR

[RPJ97] Nils Risgaard-Petersen and Kim Jensen. Nitrification and denitrification in the rhizosphere of the aquatic macrophyte Lobelia dortmanna L. Limnology and Oceanography, 42(3):529–537, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Petersen:1994:DVD

[RPRNR94] Nils Risgaard-Petersen, Suren Rysgaard, Lars Peter Nielsen, and Niels Peter Revsbech. Diurnal variation of denitrification and nitrification in sediments colonized by benthic microphytes. Limnology and Oceanography, 39(3):573–579, May 1994. CODEN LI-OCAH. ISSN 0024-3590.

Ralf:1992:PPM

[RR92] Goericke Ralf and Daniel J. Repeta. The pigments of Prochlorococcus marinus: The presence of divinylchlorophyll-a and b in a marine procaryote. Limnology and Oceanography, 37(2):425–433, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Richards:1994:OVS

[RRK94] S. R. Richards, J. W. M. Rudd, and C. A. Kelly. Organic volatile sulfur in lakes ranging in sulfate and dissolved salt concentration over five orders of magnitude. Limnology and Oceanography, 39 (3):562–572, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Rysgaard:1994:ORN

[RRPP^{+}94] Søren Rysgaard, Nils Risgaard-Petersen, Sloth Niels Peter, Jensen Kim, and Nielsen Lars Peter. Oxygen regulation of nitrification and denitrification in sediments. Limnology and Oceanography, 39 (7):1643–1652, November 1994. CODEN LIOCAH. ISSN 0024-

Raimbault:1991:SWO

[RS91] Patrick Raimbault and Gerd Slawyk. A semiautomatic, wetoxidation method for the determination of particulate organic nitrogen collected on filters. Limnology and Oceanography, 36(2): 405–408, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Rosenstock:1993:UDC

[RS93] Bernd Rosenstock and Meinhard Simon. Use of dissolved combined and free amino acids by planktonic bacteria in Lake Constance. Limnology and Oceanography, 38(7):1521–1531, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Rippeth:1998:DSV

[RS98] Tom P. Rippeth and John H. Simpson. Diurnal signals in vertical motions on the Hebridean Shelf. Limnology and Oceanography, 43 (7):1690–1696, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Ross:1999:RCT

[RS99] James M. Ross and Robert M. Sherrell. The role of colloids in tracemetal transport and adsorption behavior in New Jersey Pinelands streams. Limnology and Oceanography, 44(4):1019– 1034, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Reynolds:1997:ENL

[RSK97] Rick A. Reynolds, Dariusz Stramski, and Dale A. Kiefer. The effect of nitrogen limitation on the absorption and scattering properties of the marine diatom Thalassiosira pseudonana. Limnology and Oceanography, 42(5):881–892, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Roget:1997:ISC

[RSZ97] E. Roget, G. Salvadé, and F. Zamboni. Internal seiche climatology in a small lake where transversal and second vertical modes are usually observed. Limnology and Oceanography, 42(4):663–673, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Roden:1992:SRE

[RT92] Eric E. Roden and Jon H. Tuttle. Sulfide release from estuarine sediments underlying anoxic bottom water. Limnology and Oceanography, 37(4):725–738, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Ruttenberg:1992:DSE

[Rut92] Kathleen C. Ruttenberg. Development of a sequential extraction method for different forms of phosphorus in marine sediments. Limnology and Oceanography, 37(7):1460–1482, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Richardson:1999:GRC

[RV99] Anthony J. Richardson and Hans M. Verheye. Growth rates of copepods in the southern Benguela upwelling system: The interplay between body size and food. Limnology and Oceanography, 44(2):382–392, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Rosemond:1995:BRS

[RW95] Amy D. Rosemond and J. Bruce Wallace. Book review: Stream ecology. Structure and function of running waters (J. D. Allan). Limnology and Oceanography, 40(8):1535–1536, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Roden:1996:OCO

[RW96] Eric E. Roden and Robert G. Wetzel. Organic carbon oxidation and suppression of methane production by microbial Fe(III) oxide reduction in vegetated and unvegetated freshwater wetland sediments. Limnology and Oceanography, 41(8):1733–1748, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Robarts:1998:RPL

[RWH^{+}98] Richard D. Robarts, Marley J. Waiser, Ora Hadas, Tamar Zohary, and Sally Maclntyre. Relaxation of phosphorus limitation due to typhoon-induced mixing in two morphologically distinct basins of Lake Biwa, Japan. Limnology and Oceanography, 43(6):1023– 1036, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Ryan:1999:ECS

[RYC99] John P. Ryan, James A. Yoder, and Peter C. Cornillon. Enhanced chlorophyll at the shelfbreak of the Mid-Atlantic Bight and Georges Bank during the spring transition. Limnology and Oceanography, 44(1):1–11, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Schmid-Araya:1994:TSD

[SA94] J. M. Schmid-Araya. Temporal and spatial distribution of benthic microfauna in sediments of a gravel streambed. Limnology and Oceanography, 39(8):1813–1821, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Sutherland:1998:EBB

[SAG98] T. F. Sutherland, C. L. Amos, and J. Grant. The effect of buoyant biofilms on the erodibility of sublittoral sediments of a temperate microtidal estuary. Limnology and Oceanography, 43(2):225–235, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Saiz:1994:OFS

[Sai94] Enric Saiz. Observations of the free-swimming behavior of Acartia tonsa: Effects of food concentration and turbulent water motion. Limnology and Oceanography, 39(7):1566–1578, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Schwinghamer:1991:SCL

[SAK91] P. Schwinghamer, D. M. Anderson, and D. M. Kulis. Separation and concentration of living dinoflagellate resting cysts from marine sediments via density-gradient centrifugation. Limnology and Oceanography, 36(3):588–592, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Sale:1992:BRC

[Sal92a] Peter F. Sale. Book review: Coral reefs. Ecosystems of the world (Z. Dubinsky [ed.]). Limnology and Oceanography, 37(4):915, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Sale:1992:BRD

[Sal92b] Peter F. Sale. Book review: Dynamics of marine ecosystems. Biological-physical interactions in the oceans (K. H. Mann and J. R. N. Lazier). Limnology and Oceanography, 37(4):914–915, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Sammarco:1991:GSR

[Sam91] Paul W. Sammarco. Geographically specific recruitment and postsettlement mortality as influences on coral communities: The cross-continental shelf transplant experiment. Limnology and Oceanography, 36(3):496–514, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Sanford:1992:NSR

[San92] Lawrence P. Sanford. New sedimentation, resuspension, and burial. Limnology and Oceanography, 37(6):1164–1178, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Sanders:1999:FLF

[San99] Robert W. Sanders. Free-living freshwater protozoa — a colour guide. Limnology and Oceanography, 44(1):234–235, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Sarnelle:1992:CED

[Sar92] Orlando Sarnelle. Contrasting effects of Daphnia on ratios of nitrogen to phosphorus in a eutrophic, hard-water lake. Limnology and Oceanography, 37(7):1527–1542, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Sarnelle:1999:ZEV

[Sar99] Orlando Sarnelle. Zooplankton effects on vertical particulate flux: Testable models and experimental results. Limnology and Oceanography, 44(2):357–370, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Sauer:1992:CNC

[Sau92] Peter E. Sauer. A comment on “Nutrient cycling in a freshwater marsh: The decomposition of fish and waterfowl carrion” (parmenter and lamarra). Limnology and Oceanography, 37(7):1586– 1587, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Sullivan:1990:FLB

[SB90] Barbara K. Sullivan and Patria V. Banzon. Food limitation and benthic regulation of populations of the copepod Acartia hudsonica Pinhey in nutrient-limited and nutrient-enriched systems. Limnology and Oceanography, 35(7):1618–1631, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Strom:1993:FGB

[SB93] Suzanne L. Strom and Edward J. Buskey. Feeding, growth, and behavior of the thecate heterotrophic dinoflagellate Oblea rotunda. Limnology and Oceanography, 38(5):965–977, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Suijlen:1994:PPR

[SB94] J. M. Suijlen and J. J. Buyse. Potentials of photolytic rhodamine WT as a large-scale water tracer assessed in a long-term experiment in the Loosdrecht lakes. Limnology and Oceanography, 39(6): 1411–1423, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Skoog:1997:AVS

[SB97] Annelie Skoog and Ronald Benner. Aldoses in various size fractions of marine organic matter: Implications for carbon cycling. Limnology and Oceanography, 42(8):1803–1813, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Sharp:1993:REH

[SBB^{+}93a] Jonathan H. Sharp, Ronald Benner, Lenore Bennett, Craig A. Carlson, Rachael Dow, and Steve E. Fitzwater. Re-evaluation of high temperature combustion and chemical oxidation measurements of dissolved organic carbon in seawater. Limnology and Oceanography, 38(8):1774–1782, December 1993. CODEN LI-OCAH. ISSN 0024-3590.

Stone:1993:LKS

[SBB^{+}93b] Lewi Stone, Tom Berman, Richard Bonner, Stephen Barry, and Scott W. Weeks. Lake Kinneret: a seasonal mode1 for carbon flux through the planktonic biota. Limnology and Oceanography, 38 (8):1680–1695, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Skoog:1999:BUD

[SBB99] Annelie Skoog, Bopaiah Biddanda, and Ronald Benner. Bacterial utilization of dissolved glucose in the upper water column of the Gulf of Mexico. Limnology and Oceanography, 44(7):1625–1633, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Starink:1996:BHN

[SBGBC96a] Mathieu Starink, Marie-José Bär-Gilissen, Rolf P. M. Bak, and Thomas E. Cappenberg. Bacterivory by heterotrophic nanoflagellates and bacterial production in sediments of a freshwater littoral system. Limnology and Oceanography, 41(1):62–69, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Starink:1996:SSV

[SBGBC96b] Mathieu Starink, Marie-José Bär-Gilissen, Rolf P. M. Bak, and Thomas E. Cappenberg. Seasonal and spatial variations in heterotrophic nanoflagellate and bacteria abundances in sediments of a freshwater littoral zone. Limnology and Oceanography, 41(2): 234–242, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Sweerts:1991:OCP

[SBGCC91] Jean-Pierre R. A. Sweerts, Marie-Jose Bär-Gilissen, Adi A. Cornelese, and Thomas E. Cappenberg. Oxygen-consuming processes at the profundal and littoral sediment–water interface of a small meso-eutrophic lake (Lake Vechten, The Netherlands). Limnology and Oceanography, 36(6):1124–1133, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Short:1995:MEQ

[SBK95] Frederick T. Short, David M. Burdick, and James E. Kaldy III. Mesocosm experiments quantify the effects of eutrophication on eelgrass, Zostera marina. Limnology and Oceanography, 40(4): 740–749, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Simek:1995:CPE

[ŠBM^{+}95] Karel Šimek, Jitka Bobková, Miroslav Macek, Jiři Nedoma, and Roland Psenner. Ciliategrazing on picoplankton in a eutrophic reservoir during the summer phytoplankton maximum: a study at the species and community level. Limnology and Oceanography, 40(6):1077–1090, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Schindler:1996:ECW

[SBP^{+}96] David W. Schindler, Suzanne E. Bayley, Brian R. Parker, Ken G. Beaty, Dana R. Cruikshank, Everett J. Fee, Eva U. Schindler, and Michael P. Stainton. The effects of climatic warming on the properties of boreal lakes and streams at the Experimental Lakes Area, northwestern Ontario. Limnology and Oceanography, 41(5): 1004–1017, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Shipe:1999:RMO

[SBPV99] Rebecca F. Shipe, Mark A. Brzezinski, Cynthia Pilskaln, and Tracy A. Villareal. Rhizosolenia mats: an overlooked source of silica production in the open sea. Limnology and Oceanography, 44(5):1282–1292, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Santschi:1998:FPM

[SBW^{+}98] Peter H. Santschi, Eric Balnois, Kevin J. Wilkinson, Jingwu Zhang, Jacques Buffle, and Laodong Guo. Fibrillar polysaccharides in marine macromolecular organic matter as imaged by atomic force microscopy and transmission electron microscopy. Limnology and Oceanography, 43(5):896–908, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Sherman:1994:SPW

[SBWB94] Leslie A. Sherman, Lawrence A. Baker, Edward P. Weir, and Patrick L. Brezonik. Sediment pore-water dynamics of Little Rock Lake, Wisconsin: Geochemical processes and seasonal and spatial variability. Limnology and Oceanography, 39(5):1155–1171, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Strom:1997:PGP

[SBZD97] Suzanne L. Strom, Ronald Benner, Susan Ziegler, and Michael J. Dagg. Planktonic grazers are a potentially important source of marine dissolved organic carbon. Limnology and Oceanography, 42(6):1364–1374, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Soluk:1990:DVF

[SC90] Daniel A. Soluk and Douglas A. Craig. Digging with a vortex: Flow manipulation facilitates prey capture by a predatory stream mayfly. Limnology and Oceanography, 35(5):1201–1206, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Schmidt:1993:HCM

[SC93a] Ursula Schmidt and Ralf Conrad. Hydrogen, carbon monoxide, and methane dynamics in Lake Constance. Limnology and Oceanography, 38(6):1214–1226, September 1993. CODEN LI-OCAH. ISSN 0024-3590.

Schwartz:1993:HDP

[SC93b] Steven S. Schwartz and Guy N. Cameron. How do parasites cost their hosts? Preliminary answers from trematodes and Daphnia obtusa. Limnology and Oceanography, 38(3):602–612, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Sterner:1995:SNP

[SCEG95] Robert W. Sterner, Thomas H. Chrzanowski, James J. Elser, and Nicolas B. George. Sources of nitrogen and phosphorus supporting the growth of bacteria and phytoplankton in an oligotrophic Canadian Shield lake. Limnology and Oceanography, 40(2):242– 249, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Subramaniam:1999:BOPb

[SCF99] Ajiit Subramaniam, Edward J. Carpenter, and Paul G. Falkowski. Bio-optical properties of the marine diazotrophic cyanobacteria Trichodesmium spp. II. A reflectance model for remote sensing. Limnology and Oceanography, 44(3):618–627, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Saiz:1998:RCC

[SCFB98] Enric Saiz, Albert Calbet, Angela Fara, and Elisa Berdalet. RNA content of copepods as a tool for determining adult growth rates in the field. Limnology and Oceanography, 43(3):465–470, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Schneider:1990:DAI

[Sch90] Daniel W. Schneider. Direct assessment of the independent effects of exploitative and interference competition between Daphnia and rotifers. Limnology and Oceanography, 35(4):916–922, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Schneider:1992:SEM

[Sch92] David C. Schneider. Statistical evaluation of models in oceanography: Analysis of Schaefer’s trophic dynamic equation, rather than a surrogate. Limnology and Oceanography, 37(5):1076–1080, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Schindler:1999:MLR

[Sch99] David W. Schindler. Management of lakes and reservoirs during global climate change. Limnology and Oceanography, 44(6):1594, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Sinke:1993:MDC

[SCK93] Anja J. C. Sinke, Francis H. M. Cottaar, and Peer Keizer. A method to determine the contribution of bacteria to phosphate uptake by aerobic freshwater sediment. Limnology and Oceanography, 38(5):1081–1087, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Subramaniam:1999:BOPa

[SCKF99] Ajiit Subramaniam, Edward J. Carpenter, Deneb Karentz, and Paul G. Falkowski. Bio-optical properties of the marine diazotrophic cyanobacteria Trichodesmium spp. I. Absorption and photosynthetic action spectra. Limnology and Oceanography, 44 (3):608–617, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Silver:1998:GAI

[SCPS98] Mary W. Silver, Susan L. Coale, Cynthia H. Pilskaln, and Deborah R. Steinberg. Giant aggregates: Importance as microbial centers and agents of material flux in the mesopelagic zone. Limnology and Oceanography, 43(3):498–507, May 1998. CODEN LI-OCAH. ISSN 0024-3590.

Shish:1994:TRH

[SD94] Fuh-Kwo Shish and Hugh W. Ducklow. Temperature regulation of heterotrophic bacterioplankton abundance, production, and specific growth rate in Chesapeake Bay. Limnology and Oceanography, 39(6):1243–1258, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Shiah:1995:MVB

[SD95] Fuh-Kwo Shiah and Hugh W. Ducklow. Multiscale variability in bacterioplankton abundance, production, and specific growth rate in a temperate salt-marsh tidal creek. Limnology and Oceanography, 40(1):55–66, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Skrabal:1997:FCC

[SDB97] Stephen A. Skrabal, John R. Donat, and David J. Burdige. Fluxes of copper-complexing ligands from estuarine sediments. Limnology and Oceanography, 42(5):992–996, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Schoemann:1998:EPB

[SdBdJL98] V. Schoemann, H. J. W. de Baar, J. T. M. de Jong, and C. Lancelot. Effects of phytoplankton blooms on the cycling of manganese and iron in coastal waters. Limnology and Oceanography, 43(7):1427–1441, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Schmidt:1998:CBA

[SDJK98] Jill L. Schmidt, Jody W. Deming, Peter A. Jumars, and Richard G. Keil. Constancy of bacterial abundance in surficial marine sediments. Limnology and Oceanography, 43(5):976–982, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Scranton:1995:NMA

[SDS95] Mary I. Scranton, Percy Donaghay, and John McN. Sieburth. Nocturnal methane accumulation in the pycnocline of an anoxic estuarine basin. Limnology and Oceanography, 40(4):666–672, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Shanks:1990:VFM

[SE90] Alan L. Shanks and Elizabeth W. Edmondson. The vertical flux of metazoans (holoplankton, meiofauna, and larval invertebrates) due to their association with marine snow. Limnology and Oceanography, 35(2):455–463, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Seliger:1994:BRT

[Sel94] H. H. Seliger. Book review: Toxic phytoplankton blooms in the sea (T. J. Smayda and Y. Shimizu [eds.]). Limnology and Oceanography, 39(1):210–211, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Sellner:1997:PET

[Sel97] Kevin G. Sellner. Physiology, ecology, and toxic properties of marine cyanobacteria blooms. Limnology and Oceanography, 42(5 (part 2)):1089–1104, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Setran:1992:NPT

[Set92] Aaron Carr Setran. A new plankton trap for use in the collection of rocky intertidal zooplankton. Limnology and Oceanography, 37 (3):669–674, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Schladow:1995:PRT

[SF95] S. G. Schladow and I. H. Fisher. The physical response of temperate lakes to artificial destratification. Limnology and Oceanography, 40(2):359–373, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Suttle:1990:RAC

[SFC90] Curtis A. Suttle, Jed A. Fuhrman, and Douglas G. Capone. Rapid ammonium cycling and concentration-dependent partitioning of ammonium and phosphate: Implications for carbon transfer in planktonic communities. Limnology and Oceanography, 35(2): 424–433, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Sinsabaugh:1997:EAR

[SFFF97] Robert L. Sinsabaugh, Stuart Findlay, Paula Franchini, and David Fischer. Enzymatic analysis of riverine bacterioplankton production. Limnology and Oceanography, 42(1):29–38, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Schanz:1998:PPP

[SFRB98] F. Schanz, C. Fischer-Romero, and R. Bachofen. Photosynthetic production and photoadaptation of phototrophic sulfur bacteria in Lake Cadagno (Switzerland). Limnology and Oceanography, 43 (6):1262–1269, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Stefan:1995:SDO

[SFW^{+}95] Heinz G. Stefan, Xing Fang, David Wright, John G. Eaton, and J. Howard McCormick. Simulation of dissolved oxygen profiles in a transparent, dimictic lake. Limnology and Oceanography, 40(1): 105–118, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Stevenson:1993:EAD

[SG93] R. Jan Stevenson and R. Glover. Effects of algal density and current on ion transport through periphyton communities. Limnology and Oceanography, 38(6):1276–1281, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Sutherland:1998:ECP

[SGA98] T. F. Sutherland, J. Grant, and C. L. Amos. The effect of carbohydrate production by the diatom Nitzschia curvilineata on the erodibility of sediment. Limnology and Oceanography, 43(1):65– 72, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Smith:1997:ABT

[SGB^{+}97] K. L. Smith, Jr., R. C. Glatts, R. J. Baldwin, S. E. Beaulieu, A. H. Uhlman, R. C. Horn, and C. E. Reimers. An autonomous. bottomtransecting vehicle for making long time-series measurements of sediment community oxygen consumption to abyssal depths. Limnology and Oceanography, 42(7):1601–1612, November 1997. CO-DEN LIOCAH. ISSN 0024-3590.

Sciandra:1997:GCP

[SGC^{+}97] Antoine Sciandra, Jacques Gostan, Yves Collos, Chantal Descolas-Gros, Christophe Leboulanger, Véronique Martin-Jézéquel, Michel Denis, Dominique Lefèvre, Claire Copin-Montégut, and Bernard Avril. Growth-compensating phenomena in continuous cultures of Dunaliella tertiolecta limited simultaneously by light and nitrate. Limnology and Oceanography, 42(6): 1325–1339, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Snelgrove:1999:SHS

[SGG^{+}99] Paul V. R. Snelgrove, Judith P. Grassle, J. Frederick Grassle, Rosemarie F. Petrecca, and Hongguang Ma. In situ habitat selection by settling larvae of marine soft-sediment invertebrates. Limnology and Oceanography, 44(5):1341–1347, July 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Snelgrove:1992:RFP

[SGP92] P. V. R. Snelgrove, J. F. Grassle, and R. F. Petrecca. The role of food patches in maintaining high deep-sea diversity: Field experiments with hydrodynamically unbiased colonization trays. Limnology and Oceanography, 37(7):1543–1550, November 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Snelgrove:1996:EEA

[SGP96] Paul V. R. Snelgrove, J. Frederick Grassle, and Rosemarie F. Petrecca. Experimental evidence for aging food patches as a factor contributing to high deep-sea macrofaunal diversity. Limnology and Oceanography, 41(4):605–614, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Sundby:1992:PCC

[SGSA92] Bjørn Sundby, Charles Gobeil, Norman Silverberg, and Mucci Alfonso. The phosphorus cycle in coastal marine sediments. Limnology and Oceanography, 37(6):1129–1145, September 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Stordal:1996:MPS

[SGWS96] M. C. Stordal, G. A. Gill, L.-S. Wen, and P. H. Santschi. Mercury phase speciation in the surface waters of three Texas estuaries: Importance of colloidal forms. Limnology and Oceanography, 41 (1):52–61, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Shimaraev:1993:DVL

[SGZ93] M. N. Shimaraev, N. G. Granin, and A. A. Zhdanov. Deep ventilation of Lake Baikal waters due to spring thermal bars. Limnology and Oceanography, 38(5):1068–1072, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1990:DCM

[SH90] William G. Sunda and Susan A. Huntsman. Diel cycles in microbial manganese oxidation and manganese redox speciation in coastal waters of the Bahama Islands. Limnology and Oceanography, 35(2):325–338, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Schelske:1991:RCP

[SH91] Claire L. Schelske and David A. Hodeli. Recent changes in productivity and climate of Lake Ontario detected by isotopic analysis of sediments. Limnology and Oceanography, 36(5):961–975, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1992:FIB

[SH92] William G. Sunda and Susan A. Huntsman. Feedback interactions between zinc and phytoplankton in seawater. Limnology and Oceanography, 37(1):25–40, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Schelske:1995:UCI

[SH95a] Claire L. Schelske and David A. Hodell. Using carbon isotopes of bulk sedimentary organic matter to reconstruct the history of nutrient loading and eutrophication in Lake Erie. Limnology and Oceanography, 40(5):918–929, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1995:CZI

[SH95b] William G. Sunda and Susan A. Huntsman. Cobalt and zinc interreplacement in marine phytoplankton: Biological and geochemical implications. Limnology and Oceanography, 40(8):1404–1417, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1995:RCC

[SH95c] William G. Sunda and Susan A. Huntsman. Regulation of copper concentration in the oceanic nutricline by phytoplankton uptake and regeneration cycles. Limnology and Oceanography, 40(1):132– 137, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Sanden:1996:LTT

[SH96a] Per Sandén and Bertil Håkansson. Long-term trends in Secchi depth in the Baltic Sea. Limnology and Oceanography, 41(2):346– 351, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1996:ABC

[SH96b] William G. Sunda and Susan A. Huntsman. Antagonisms between cadmium and zinc toxicity and manganese limitation in a coastal diatom. Limnology and Oceanography, 41(3):373–387, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Spaak:1997:FPD

[SH97] Piet Spaak and J. Rob Hoekstra. Fish predation on a Daphnia hybrid species complex: a factor explaining species coexistence? Limnology and Oceanography, 42(4):753–762, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Stoderegger:1998:PRB

[SH98a] Karen Stoderegger and Gerhard J. Herndl. Production and release of bacterial capsular material and its subsequent utilization by marine bacterioplankton. Limnology and Oceanography, 43(5): 877–884, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1998:IAC

[SH98b] William G. Sunda and Susan A. Huntsman. Interactions among Cu^{2+}, Zn^{2+}, and Mn^{2+} in controlling cellular Mn, Zn, and growth rate in the coastal alga Chlamydomonas. Limnology and Oceanography, 43(6):1055–1064, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1998:IEE

[SH98c] William G. Sunda and Susan A. Huntsman. Interactive effects of external manganese, the toxic metals copper and zinc, and light in controlling cellular manganese and growth in a coastal diatom. Limnology and Oceanography, 43(7):1467–1475, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Shaw:1994:EPD

[Sha94] Peter J. Shaw. The effect of pH, dissolved humic substances, and ionic composition on the transfer of iron and phosphate to particulate size fractions in epilimnetic lake water. Limnology and Oceanography, 39(7):1734–1743, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Swaney:1999:NAE

[SHB99] Dennis P. Swaney, Robert W. Howarth, and Thomas J. Butler. A novel approach for estimating ecosystem production and respiration in estuaries: Application to the oligohaline and mesohaline Hudson River. Limnology and Oceanography, 44(6):1509–1521, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Stefan:1996:SLT

[SHF^{+}96] H. G. Stefan, M. Hondzo, X. Fang, J. G. Eaton, and J. H. Mc-Cormick. Simulated long term temperature and dissolved oxygen characteristics of lakes in the north-central United States and associated fish habitat limits. Limnology and Oceanography, 41(5): 1124–1135, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Sellers:1995:CMC

[SHK95] Patricia Sellers, Raymond H. Hesslein, and Carol A. Kelly Kelly. Continuous measurement of CO_{2} for estimation of air-water fluxes in lakes: an in situ technique. Limnology and Oceanography, 40 (3):575–581, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Suzuki:1995:DPP

[SHKI95] Koji Suzuki, Nobuhiko Handa, Hiroshi Kiyosawa, and Joji Ishizaka. Distribution of the prochlorophyte prochlorococcus in the central Pacific Ocean as measured by HPLC. Limnology and Oceanography, 41(4):983–989, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Stemberger:1996:CFZ

[SHKP96] Richard S. Stemberger, Alan T. Herlihy, Danny L. Kugler, and Steven G. Paulsen. Climatic forcing on zooplankton richness in lakes of the northeastern United States. Limnology and Oceanography, 41(5):1093–1101, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Sondergaard:1995:DDO

[SHM95] Morten Søndergaard, Benni Hansen, and Stiig Markager. Dynamics of dissolved organic car-bon lability in a eutrophic lake. Limnology and Oceanography, 40(1):46–54, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Soetaer:1996:DRD

[SHM96] Karline Soetaer, Peter M. J. Herman, and Jack J. Middelburg. Dynamic response of deep-sea sediments to seasonal variations: a model. Limnology and Oceanography, 41(8):1651–1668, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Sterner:1993:PNL

[SHSS93] Robert W. Sterner, Douglas D. Hagemeier, William L. Smith, and Robert F. Smith. Phytoplankton nutrient limitation and food quality for Daphnia. Limnology and Oceanography, 38(4):857–871, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Saggio:1998:IWW

[SI98] Angelo Saggio and Jörg Imberger. Internal wave weather in a stratified lake. Limnology and Oceanography, 43(8):1780–1795, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Siegel:1998:RCD

[Sie98] David A. Siegel. Resource competition in a discrete environment: Why are plankton distributions paradoxical? Limnology and Oceanography, 43(6):1133–1146, September 1998. CODEN LI-OCAH. ISSN 0024-3590.

Suzumura:1998:CDO

[SIO98] Masahiro Suzumura, Kimitoshi Ishikawa, and Hiroshi Ogawa. Characterization of dissolved organic phosphorus in coastal seawater using ultrafiltration and phosphohydrolytic enzymes. Limnology and Oceanography, 43(7):1553–1564, November 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Siver:1993:ISC

[Siv93] Peter A. Siver. Inferring the specific conductivity of lake water with scaled chrysophytes. Limnology and Oceanography, 38(7): 1480–1492, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Salonen:1994:RSO

[SJDJ94] Kalevi Salonen, Roger I. Jones, Henk De Haan, and Mark James. Radiotracer study of phosphorus uptake by plankton and redistribution in the water column of a small humic lake. Limnology and Oceanography, 39(1):69–83, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Sprules:1998:COP

[SJHS98] W. G. Sprules, E. H. Jin, A. W. Herman, and J. D. Stockwell. Calibration of an optical plankton counter for use in fresh water. Limnology and Oceanography, 43(4):726–733, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Shimeta:1995:ITS

[SJL95] Jeff Shimeta, Peter A. Jumars, and Evelyn J. Lessard. Influences of turbulence on suspension feeding by planktonic protozoa; experiments in laminar shear fields. Limnology and Oceanography, 40(5):845–859, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Jensen:1994:PSA

[SJP94] Kaj Sand-Jensen and Marten Foldager Pedersen. Photosynthesis by symbiotic algae in the freshwater sponge, Spongilla lacustris. Limnology and Oceanography, 39(3):551–561, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Sand-Jensen:1997:RRP

[SJPGH97] Kaj Sand-Jensen, Ole Pedersen, and Ole Geertz-Hansen. Regulation and role of photosynthesis in the colonial symbiotic ciliate Ophrydium versatile. Limnology and Oceanography, 42(5):866– 873, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Sumi:1990:EAA

[SK90] Takayuki Sumi and Isao Koike. Estimation of ammonification and ammonium assimilation in surficial coastal and estuarine sediments. Limnology and Oceanography, 35(2):270–286, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Suzumura:1995:ODI

[SK95] Masahiro Suzumura and Akiyoshi Kamatani. Origin and distribution of inositol hexaphosphate in estuarine and coastal sediments. Limnology and Oceanography, 40(7):1254–1261, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Stewart:1996:CUL

[SK96] Arthur J. Stewart and Lynn A. Kszos. Caution on using lithium (Li^{+}) as a conservative tracer in hydrological studies. Limnology and Oceanography, 41(1):190–191, January 1996. CODEN LIOCAH. ISSN 0024-3590. See reply [TMB96].

Smith:1994:CNB

[SKB94] K. L. Smith, Jr., R. S. Kaufmann, and R. J. Baldwin. Coupling of near-bottom pelagic and benthic processes at abyssal depths in the eastern North Pacific Ocean. Limnology and Oceanography, 39(5):1101–1118, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Simek:1999:SBC

[ŠKN^{+}99] Karel Šimek, Petra Kojecká, Jiří Nedoma, Petr Hartman, Jaroslav Vrba, and John R. Dolan. Shifts in bacterial community composition associated with different microzooplankton size fractions in a eutrophic reservoir. Limnology and Oceanography, 44(7):1634– 1644, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Sweerts:1991:SWS

[SKR^{+}91] Jean-Pierre R. A. Sweerts, Carol A. Kelly, John W. M. Rudd, Ray Hesslein, and Thomas E. Cappenberg. Similarity of wholesediment molecular diffusion coefficients in freshwater sediments of low and high porosity. Limnology and Oceanography, 36(2): 335–341, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Stevens:1998:SMW

[SL98] C. L. Stevens and G. A. Lawrence. Stability and meromixis in a water-filled mine pit. Limnology and Oceanography, 43(5):946– 954, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Sun:1993:AOD

[SLA93] Ming-Yi Sun, Cindy Lee, and Robert C. Aller. Anoxic and oxic degradation of ^{14}C-labeled chloropigments and a ^{14}C-labeled diatom in Long Island Sound sediments. Limnology and Oceanography, 38(7):1438–1451, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Siver:1999:CCC

[SLC^{+}99] Peter A. Siver, Anne Marie Lott, Ethan Cash, Jamal Moss, and Laurence J. Marsicano. Century changes in Connecticut, U.S.A., lakes as inferred from siliceous algal remains and their relationships to land-use change. Limnology and Oceanography, 44(8): 1928–1935, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Santas:1997:URD

[SLD97] Regas Santas, Charalumbia Lianou, and Daniel Danielidis. UVB radiation and depth interaction during primary succession of marine diatom assemblages of Greece. Limnology and Oceanography, 42(5):986–991, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Schultze-Lam:1997:WEB

[SLSLBD97] Susanne Schultze-Lam, Susanne Schultze-Lam, Terrance J. Beveridge, and David J. Des Marais. Whiting events: Biogenic origin due to the photosynthetic activity of cyanobacterial picoplankton. Limnology and Oceanography, 42(1):133–141, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Stepanauskas:1999:BWD

[SLT99] Ramūnas Stepanauskas, Lars Leonardson, and Lars J. Tranvik. Bioavailability of wetland-derived DON to freshwater and marine bacterioplankton. Limnology and Oceanography, 44(6):1477–1485, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Spratt:1990:SCC

[SM90] Henry G. Spratt, Jr. and Mark D. Morgan. Sulfur cycling in a cedar-dominated, freshwater wetland. Limnology and Oceanography, 35(7):1586–1593, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Sosik:1991:AFQ

[SM91] Heidi M. Sosik and B. Greg Mitchell. Absorption, fluorescence, and quantum yield for growth in nitrogen-limited Dunaliella tertiolecta. Limnology and Oceanography, 36(5):910–921, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Sabath:1992:ULF

[SM92] Ana Sabath and Mercedes Masó. Unusual larval fish distribution pattern in coastal zone of the western Mediterranean. Limnology and Oceanography, 37(6):1252–1260, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Stramski:1997:EMP

[SM97] Dariusz Stramski and Curtis D. Mobley. Effects of microbial particles on oceanic optics: a database of single-particle optical properties. Limnology and Oceanography, 42(3):538–549, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Soemundsdottir:1998:BPM

[SM98a] Stefanía Sœmundsdóttir and Patricia A. Matrai. Biological production of methyl bromide by cultures of marine phytoplankton. Limnology and Oceanography, 43(1):81–87, January 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Stibor:1998:IFE

[SM98b] Herwig Stibor and Jiří Macháček. The influence of fish-exuded chemical signals on the carbon budget of Daphnia. Limnology and Oceanography, 43(5):997–1000, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Striegl:1998:HIM

[SM98c] Robert G. Striegl and Catherine M. Michmerhuizen. Hydrologic influence on methane and carbon dioxide dynamics at two northcentral Minnesota lakes. Limnology and Oceanography, 43(7): 1519–1529, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Scarratt:1999:PCH

[SM99a] M. G. Scarratt and R. M. Moore. Production of chlorinated hydrocarbons and methyl iodide by the red microalga Porphyridium purpureum. Limnology and Oceanography, 44(3):703–707, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Sundareshwar:1999:PSC

[SM99b] P. V. Sundareshwar and James T. Morris. Phosphorus sorption characteristics of intertidal marsh sediments along an estuarine salinity gradient. Limnology and Oceanography, 44(7):1693–1701, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Smayda:1997:HAB

[Sma97a] Theodore J. Smayda. Harmful algal blooms: Their ecophysiology and general relevance to phytoplankton blooms in the sea. Limnology and Oceanography, 42(5 (part 2)):1137–1153, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Smayda:1997:WBC

[Sma97b] Theodore J. Smayda. What is a bloom? A commentary. Limnology and Oceanography, 42(5 (part 2)):1132–1136, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Scholin:1997:DQP

[SMB^{+}97] Christopher Scholin, Peter Miller, Kurt Buck, Francisco Chavez, Paul Harris, Paul Haydock, Judith Howard, and Gerard Cangelosi. Detection and quantification of Pseudo-nitzschia australis in cultured and natural populations using LSU rRNA-targeted probes. Limnology and Oceanography, 42(5 (part 2)):1265–1272, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Shiller:1998:DDV

[SMC98] Alan M. Shiller, Lunjin Mao, and Jeramé Cramer. Determination of dissolved vanadium in natural waters by flow injection analysis with colorimetric detection. Limnology and Oceanography, 43(3): 526–529, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Shun:1994:DAP

[SMH94] Ying Shun, Ian D. McKelvie, and Barry T. Hart. Determination of alkaline phosphatase-hydrolyzable phosphorus in natural water systems by enzymatic flow injection. Limnology and Oceanography, 39(8):1993–2000, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Smith:1990:BRL

[Smi90a] Val H. Smith. Book review: Lake restoration by reduction of nutrient loading: Expectations, experiences. extrapolations (H. Sas [ed.]). Limnology and Oceanography, 35(6):1412–1413, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Smith:1990:NPN

[Smi90b] Val H. Smith. Nitrogen, phosphorus, and nitrogen fixation in lacustrine and estuarine ecosystems. Limnology and Oceanography, 35(8):1852–1859, December 1990. CODEN LIOCAH. ISSN 0024-3590. See reply [HR90].

Smith:1991:CBC

[Smi91] Val H. Smith. Competition between consumers. Limnology and Oceanography, 36(4):820–823, June 1991. CODEN LIOCAH. ISSN 0024-3590. See reply [Rot91].

Smith:1992:BBL

[Smi92] K. L. Smith, Jr. Benthic boundary layer communities and carbon cycling at abyssal depths in the central North Pacific. Limnology and Oceanography, 37(5):1034–1056, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Smith:1996:BSO

[Smi96] Walker O. Smith, Jr. The biology of the Southern Ocean (G. A. Knox). Limnology and Oceanography, 41(2):372, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Scully:1996:HPF

[SML96] N. M. Scully, D. J. McQueen, and D. R. S. Lean. Hydrogen peroxide formation: The interaction of ultraviolet radiation and dissolved organic carbon in lake waters along a 43–75^{o}N gradient. Limnology and Oceanography, 41(3):540–548, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Slomp:1998:RAS

[SMV98] C. P. Slomp, J. F. P. Malschaert, and W. Van Raaphorst. The role of adsorption in sediment-water exchange of phosphate in North Sea continental margin sediments. Limnology and Oceanography, 43(5):832–846, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Smith:1990:PGN

[SN90] Walker O. Smith, Jr. and David M. Nelson. Phytoplankton growth and new production in the Weddell Sea marginal ice zone in the austral spring and autumn. Limnology and Oceanography, 35(4): 809–821, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Sloth:1992:NSC

[SNB92] Niels P. Sloth, Lars P. Nielsen, and T. Henry Blackburn. Nitrification in sediment cores measured with acetylene inhibition. Limnology and Oceanography, 37(5):1108–1112, July 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Sarmiento:1991:TDS

[SO91] Jorge L. Sarmiento and James C. Orr. Three-dimensional simulations of the impact of Southern Ocean nutrient depletion on atmospheric CO_{2} and ocean chemistry. Limnology and Oceanography, 36(8):1928–1950, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Sommer:1989:NSN

[Som89] Ulrich Sommer. Nutrient status and nutrient competition of phytoplankton in a shallow, hypertrophic lake. Limnology and Oceanography, 34(7):1162–1173, November 1989. CODEN LI-OCAH. ISSN 0024-3590. See comment [Pra90].

Sommer:1990:RRA

[Som90] Ulrich Sommer. Resource ratios or absolute concentrations: a reply to the comment by Prairie. Limnology and Oceanography, 35(3):779–780, May 1990. CODEN LIOCAH. ISSN 0024-3590. See [Pra90].

Sommer:1992:PLD

[Som92] Ulrich Sommer. Phosphorus-limited Daphnia: Intraspecific facilitation instead of competition. Limnology and Oceanography, 37 (5):966–973, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Sommer:1993:PCP

[Som93] Ulrich Sommer. Phytoplankton competition in Plußsee: a field test of the resource-ratio hypothesis. Limnology and Oceanography, 38(4):838–845, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Sommer:1994:ILI

[Som94] Ulrich Sommer. The impact of light intensity and daylength on silicate and nitrate competition among marine phytoplankton. Limnology and Oceanography, 39(7):1680–1688, November 1994. CO-DEN LIOCAH. ISSN 0024-3590.

Sommer:1995:ETI

[Som95] Ulrich Sommer. An experimental test of the intermediate disturbance hypothesis using cultures of marine phytoplankton. Limnology and Oceanography, 40(7):1271–1277, November 1995. CO-DEN LIOCAH. ISSN 0024-3590.

Sommer:1997:SIC

[Som97] Ulrich Sommer. Selectivity of Idothea chelipes (Crustacea: Isopoda) grazing on benthic microalgae. Limnology and Oceanography, 42(7):1622–1628, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Sommer:1998:ACA

[Som98] Ulrich Sommer. From algal competition to animal production: Enhanced ecological efficiency of Brachionus plicatilis with a mixed diet. Limnology and Oceanography, 43(6):1393–1396, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Sosik:1999:SMP

[Sos99] Heidi M. Sosik. Storage of marine particulate samples for lightabsorption measurements. Limnology and Oceanography, 46(4): 1139–1141, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Sanders:1990:BFF

[SP90a] Robert W. Sanders and Karen G. Porter. Bacterivorous flagellates as food resources for the freshwater crustacean zooplankter Daphnia ambigua. Limnology and Oceanography, 35(1):188–191, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Smetacek:1990:SBI

[SP90b] Victor Smetacek and Uta Passow. Spring bloom initiation and Sverdrup’s critical-depth model. Limnology and Oceanography, 35 (1):228–234, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Sanders:1990:HAM

[SPB90] Robert W. Sanders, Karen G. Porter, and Susan J. Bennett. Heterotrophic, autotrophic, and mixotrophic nanoflagellates: Seasonal abundances and bacterivory in a eutrophic lake. Limnology and Oceanography, 35(8):1821–1832, December 1990. CODEN LI-OCAH. ISSN 0024-3590.

Spigel:1996:BRP

[Spi96] Robert H. Spigel. Book review: Physical processes in a large lake: Lake Biwa, Japan (S. Okuda, J. Imberger, and M. Kumagai [eds.]). Limnology and Oceanography, 40(5):813–814, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Sun:1997:UEC

[SPMW97] L. Sun, E. M. Perdue, J. L. Meyer, and J. Weis. Use of elemental composition to predict bioavailability of dissolved organic matter in a Georgia river. Limnology and Oceanography, 42(4):714–721, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Simon:1992:CNS

[SR92] Meinhard Simon and Bernd Rosenstock. Carbon and nitrogen sources of planktonic bacteria in Lake Constance studied by the composition and isotope dilution of intracellular amino acids. Limnology and Oceanography, 37(7):1496–1511, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Stramski:1993:DVO

[SR93] Dariusz Stramski and Rick A. Reynolds. Diel variations in the optical properties of a marine diatom. Limnology and Oceanography, 38(7):1347–1364, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Sterner:1994:TGD

[SR94] Robert W. Sterner and Judy L. Robinson. Thresholds for growth in Daphnia magna with high and low phosphorus diets. Limnology and Oceanography, 39(5):1228–1232, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Seip:1995:PFA

[SR95] K. L. Seip and C. S. Reynolds. Phytoplankton functional attributes along trophic gradient and season. Limnology and Oceanography, 40(3):589–597, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Slawyk:1998:MGU

[SRG98] Gerd Slawyk, Patrick Raimbault, and Nicole Garcia. Measuring gross uptake of ^{15}N-labeled nitrogen by marine phytoplankton without particulate matter collection: Evidence of low ^{15}N losses to the dissolved organic nitrogen pool. Limnology and Oceanography, 43(7):1734–1739, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Slawyk:2000:UMD

[SRG00] Gerd Slawyk, Patrick Raimbault, and Nicole Garcia. Use of ^{15}N to measure dissolved organic nitrogen release by marine phytoplankton (reply to comment by Bronk and Ward). Limnology and Oceanography, 45(8):1884–1886, December 2000. CODEN LIOCAH. ISSN 0024-3590. See [BW99b].

Shick:1999:UBR

[SRLRL^{+}99] J. M. Shick, S. Romaine-Lioud, S. Romaine-Lioud, C. Ferrier-Pagès, and J.-P. Gattuso. Ultraviolet-B radiation stimulates shikimate pathway-dependent accumulation of mycosporine-like amino acids in the coral Stylophora pistillata despite decreases in its population of symbiotic dinoflagellates. Limnology and Oceanography, 44(7):1667–1682, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Szyper:1992:DCP

[SRPG92] James P. Szyper, Joel Z. Rosenfeld, Raul H. Piedrahita, and Philip Giovannini. Diel cycles of planktonic respiration rates in briefly incubated water samples from a fertile earthen pond. Limnology and Oceanography, 37(6):1193–1201, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Stockner:1993:BRF

[SS93] John G. Stockner and Curtis A. Suttle. Book review: Food web management. A case study of Lake Mendota (J. F Kitchell [ed.]). Limnology and Oceanography, 38(7):1587–1588, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Seal:1998:OHI

[SS98] Robert R. Seal II and Wayne C. Shanks III. Oxygen and hydrogen isotope systematics of Lake Baikal, Siberia: Implications for paleoclimate studies. Limnology and Oceanography, 43(6):1251–1261, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Schulz:1999:PPL

[SS99a] Kimberly L. Schulz and Robert W. Sterner. Phytoplankton phosphorus limitation and food quality for bosmina. Limnology and Oceanography, 44(6):1549–1556, September 1999. CODEN LI-OCAH. ISSN 0024-3590.

Seitzinger:1999:AID

[SS99b] S. P. Seitzinger and R. W. Sanders. Atmospheric inputs of dissolved organic nitrogen stimulate estuarine bacteria and phytoplankton. Limnology and Oceanography, 44(3):721–730, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Smith:1999:SBH

[SS99c] J. N. Smith and C. T. Schafer. Sedimentation, bioturbation, and Hg uptake in the sediments of the estuary and Gulf of St. Lawrence. Limnology and Oceanography, 44(1):207–219, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Stokes:1999:OOS

[SS99d] M. D. Stokes and G. N. Somero. An optical oxygen sensor and reaction vessel for high-pressure applications. Limnology and Oceanography, 44(1):189–195, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Saager:1990:SPW

[SSE90] Paul M. Saager, Jean-Pierre Sweerts, and Hans J. Ellermeijer. A simple pore-water sampler for coarse, sandy sediments of low porosity. Limnology and Oceanography, 35(3):747–751, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Shaw:1990:IAQ

[SSFP90] R. D. Shaw, J. F. H. Shaw, H. Fricker, and E. E. Prepas. An integrated approach to quantify groundwater transport of phosphorus to Narrow Lake, Alberta. Limnology and Oceanography, 35(4):870–886, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Sherwood:1991:DOC

[SSKW91] J. E. Sherwood, F. Stagnitti, M. J. Kokkinn, and W. D. Williams. Dissolved oxygen concentrations in hypersaline waters. Limnology and Oceanography, 36(2):235–250, March 1991. CODEN LI-OCAH. ISSN 0024-3590.

Stephenson:1995:FDS

[SSM^{+}95] M. Stephenson, W. J. Schwartz, M. F. Motycka, D. J. Rowan, C. Kelly, and R. D. Roshon. Fate and distribution in sediments of carbon-14 added to Canadian Shield lakes of differing trophic state. Limnology and Oceanography, 40(4):779–790, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Steinberg:1994:MZC

[SSP^{+}94] Deborah K. Steinberg, Mary W. Silver, Cynthia H. Pilskaln, Susan L. Coale, and Jennifer B. Paduan. Midwater zooplankton communities on pelagic detritus (giant larvacean houses) in Monterey Bay, California. Limnology and Oceanography, 39(7):1606– 1620, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Suzuki:1993:DDC

[SSS93] Marcelino T. Suzuki, Evelyn B. Sherr, and Barry F. Sherr. DAPI direct counting underestimates bacterial abundances and average cell size compared to AO direct counting. Limnology and Oceanography, 38(7):1566–1570, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Staufer:1990:GWS

[Sta90a] Robert E. Staufer. Granite weathering and the sensitivity of alpine lakes to acid deposition. Limnology and Oceanography, 35(5): 1112–1134, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Staufferl:1990:AML

[Sta90b] Robert E. Staufferl. Alkalinities of Maine lakes: Are they really changing? Limnology and Oceanography, 35(6):1238–1257, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Stauffer:1991:CSS

[Sta91] Robert E. Stauffer. A comment on “Sulfur storage and alkalinity generation in New England lake sediments” (Giblin et al.). Limnology and Oceanography, 36(6):1263–1265, September 1991. CODEN LIOCAH. ISSN 0024-3590. See [GLWH90] and replies [GLH91, NKMO91].

Stemberger:1990:FLS

[Ste90] Richard S. Stemberger. Food limitation. spination, and reproduction in Brachionus calyciflorus. Limnology and Oceanography, 35 (1):33–44, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Sterner:1994:SSP

[Ste94] Robert W. Sterner. Seasonal and spatial patterns in macro- and micronutrient limitation in Joe Pool Lake, Texas. Limnology and Oceanography, 39(3):535–550, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Stigebrandt:1991:COF

[Sti91] Anders Stigebrandt. Computations of oxygen fluxes through the sea surface and the net production of organic matter with application to the Baltic and adjacent seas. Limnology and Oceanography, 36(3):444–454, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Stich:1992:STH

[Sti92] Hans Bernd Stich. Spatial and temporal heterogeneity of Daphnia in Lake Constance; intra- and interspecific comparisons. Limnology and Oceanography, 37(6):1327–1334, September 1992. CO-DEN LIOCAH. ISSN 0024-3590.

Saiz:1993:ERE

[STJ^{+}93] Enric Saiz, Peter Tiselius, Per R. Jonsson, Peter Verity, and Gustav-Adolf Paffenhöfer. Experimental records of the effects of food patchiness and predation on egg production of Acartia tonsa. Limnology and Oceanography, 38(2):280–289, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Stramski:1990:AMA

[Str90] Dariusz Stramski. Artifacts in measuring absorption spectra of phytoplankton collected on a filter. Limnology and Oceanography, 35(8):1804–1809, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Stramski:1994:BRO

[Str94] Dariusz Stramski. Book review: Ocean optics (R. W. Spinrad, K. L. Carder, and M. J. Perry [eds.]). Limnology and Oceanography, 39(8):2015–2017, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Strathmann:1995:BRR

[Str95] Richard R. Strathmann. Book review: Reproduction, larval biology, and recruitment of the deep-sea benthos (C. M. Young and K. J. Eckelbarger [eds.]). Limnology and Oceanography, 40(2): 446–447, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Straile:1997:GGE

[Str97] Dietmar Straile. Gross growth efficiencies of protozoan and metazoan zooplankton and their dependence on food concentration, predator-prey weight ratio, and taxonomic group. Limnology and Oceanography, 42(6):1375–1385, September 1997. CODEN LI-OCAH. ISSN 0024-3590.

Smith:1998:EPI

[SU98] David J. Smith and Graham J. C. Underwood. Exopolymer production by intertidal epipelic diatoms. Limnology and Oceanography, 43(7):1578–1591, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Suttle:1993:LA

[Sut93a] Curtis A. Suttle. 1993 Lindeman Award. Limnology and Oceanography, 38(7):1592, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Suttle:1993:BRP

[Sut93b] Curtis A. Suttle. Book review: Primary productivity and biogeochemical cycles in the sea (P. G. Falkowski and A. D. Woodhead [eds.]). Limnology and Oceanography, 38(7):1588, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Sundby:1998:MRC

[SVC^{+}98] Bjørn Sundby, Carlos Vale, Zsabel Caçador, Fernando Catarino, Maria-João Madureira, and Miguel Caetano. Metal-rich concretions on the roots of salt marsh plants: Mechanism and rate of formation. Limnology and Oceanography, 43(2):245–252, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Svensson:1997:CPS

[Sve97] Jan-Erik Svensson. Chaoborus predation and sex-specific mortality in a copepod. Limnology and Oceanography, 42(3):572–577, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Sancetta:1991:MFR

[SVF91] Constance Sancetta, Tracy Villareal, and Paul Falkowski. Massive fluxes of rhizosolenid diatoms: a common occurrence? Limnology and Oceanography, 36(7):1452–1457, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Schaus:1997:NPE

[SVW^{+}97] M. H. Schaus, M. J. Vanni, T. E. Wissing, M. T. Bremigan, J. E. Garvey, and R. A. Stein. Nitrogen and phosphorus excretion by detritivorous gizzard shad in a reservoir ecosystem. Limnology and Oceanography, 42(6):1386–1397, September 1997. CODEN LIOCAH. ISSN 0024-3590.

Strom:1991:PSR

[SW91] Suzanne L. Strom and Nicholas A. Welschmeyer. Pigment-specific rates of phytoplankton growth and microzooplankton grazing in the open subarctic Pacific Ocean. Limnology and Oceanography, 36(1):50–63, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Suttle:1993:BRB

[SW93] Curtis A. Suttle and Robert G. Wetzel. Book review: Biodiversity of the southeastern United States: 1. Aquatic communities (C. T. Hackney, S. M. Adams, and W. H. Martin [eds.]). Limnology and Oceanography, 38(7):1588–1589, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Shanks:1996:FHD

[SW96a] Alan Shanks and Keith Walters. Feeding by a heterotrophic dinoflagellate (Noctiluca scintillans) in marine snow. Limnology and Oceanography, 41(1):177–181, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Shannon:1996:EST

[SW96b] Robert D. Shannon and Jeffrey R. White. The effects of spatial and temporal variations in acetate and sulfate on methane cycling in two Michigan peatlands. Limnology and Oceanography, 41(3): 435–443, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Swaney:1999:ASB

[Swa99] Dennis P. Swaney. Analytical solution of Boudreaús equation for a tracer subject to food-feedback bioturbation. Limnology and Oceanography, 44(3):697–698, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Sanderson:1992:CSI

[SWH92] Brian G. Sanderson, Ian T. Webster, and Stella E. Humphries. Chaotic solutions for irradiance-regulated buoyant motion of cyanobacteria. Limnology and Oceanography, 37(8):1691–1704, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Swift:1992:PCF

[Swi92] Michael C. Swift. Prey capture by the four larval instars of Chaoborus crystallinus. Limnology and Oceanography, 37(1):14– 24, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Sherman:1998:TBA

[SWJO98] Bradford S. Sherman, Ian T. Webster, Gary J. Jones, and Rod L. Oliver. Transitions between Auhcoseira and Anabaena dominance in a turbid river weir pool. Limnology and Oceanography, 43(8): 1902–1915, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Stoermer:1993:PCL

[SWS93] E. F. Stoermer, J. A. Wolin, and C. L. Schelske. Paleolimnological comparison of the Laurentian Great Lakes based on diatoms. Limnology and Oceanography, 38(6):1311–1316, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Sanders:1996:RSH

[SWS^{+}96] Robert W. Sanders, Craig E. Williamson, Paul L. Stutzman, Robert E. Moeller, Clyde E. Goulden, and Rika Aoki-Goldsmith. Reproductive success of “herbivorous” zooplankton fed algal and nonalgal food resources. Limnology and Oceanography, 41(6): 1295–1305, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Stoermer:1990:SMS

[SWSC90] E. F. Stoermer, J. A. Wolin, C. L. Schelske, and D. J. Conley. Siliceous microfossil succession in Lake Michigan. Limnology and Oceanography, 35(4):959–967, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Schulz:1995:DCA

[SY95] Kimberly L. Schulz and Peder M. Yurista. Diet composition from allozyme analysis in the predatory cladoceran Bythotrephes cederstroemi. Limnology and Oceanography, 40(4):821–826, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Smith:1992:MPR

[SZA92] G. Jason Smith, Richard C. Zimmerman, and Randall S. Alberte. Molecular and physiological responses of diatoms to variable levels of irradiance and nitrogen availability: Growth of Skeletonema costatum in simulated upwelling conditions. Limnology and Oceanography, 37(5):989–1007, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Schulze:1995:CBC

[SZW95] Peter C. Schulze, Horacio E. Zagarese, and Craig E. Williamson. Competition between crustacean zooplankton in continuous cultures. Limnology and Oceanography, 40(1):33–45, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Thomas:1997:AUC

[TA97] F. I. M. Thomas and M. J. Atkinson. Ammonium uptake by coral reefs: Effects of water velocity and surface roughness on mass transfer. Limnology and Oceanography, 42(1):81–88, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Taylor:1995:MDS

[Tay95] Gordon T. Taylor. Microbial degradation of sorbed and dissolved protein in seawater. Limnology and Oceanography, 40(5):875–885, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Torreton:1991:EBD

[TB91] J. P. Torréton and M. Bouvy. Estimating bacterial DNA synthesis from [^{3}H]thymidine incorporation: Discrepancies among macromolecular extraction procedures. Limnology and Oceanography, 36(2):299–306, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Twombly:1996:EAN

[TB96] Saran Twombly and Carolyn W. Burns. Exuvium analysis: a nondestructive method of analyzing copepod growth and development. Limnology and Oceanography, 41(6):1324–1329, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Thomson-Bulldis:1998:ANM

[TBK98] Angie Thomson-Bulldis and David Karl. Application of a novel method for phosphorus determinations in the oligotrophic North Pacific Ocean. Limnology and Oceanography, 43(7):1565–1577, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Takahashi:1990:OPA

[TBM90] Kozo Takahashi, John D. Billings, and Julia K. Morgan. Oceanic province: Assessment from the time-series diatom fluxes in the northeastern Pacific. Limnology and Oceanography, 35(1):154– 165, January 1990. CODEN LIOCAH. ISSN 0024-3590.

Tate:1995:PDA

[TBM95] Cathy M. Tate, Robert E. Broshears, and Diane M. McKnight. Phosphate dynamics in an acidic mountain stream: Interactions involving algal uptake, sorption by iron oxide, and photoreduction. Limnology and Oceanography, 40(5):938–946, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Thuesen:1993:EAM

[TC93] Erik V. Thuesen and James J. Childress. Enzymatic activities and metabolic rates of pelagic chaetognaths: Lack of depth-related declines. Limnology and Oceanography, 38(5):935–948, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Thompson:1994:CIF

[TC94] Peter A. Thompson and Stephen E. Calvert. Carbon-isotope fractionation by a marine diatom: The influence of irradiance, daylength, pH, and nitrogen source. Limnology and Oceanography, 39(8):1835–1844, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Thompson:1995:CIF

[TC95a] Peter A. Thompson and Stephen E. Calvert. Carbon isotope fractionation by Emiliania huxleyi. Limnology and Oceanography, 40 (4):673–679, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Twiss:1995:RTM

[TC95b] Michael R. Twiss and Peter G. C. Campbell. Regeneration of trace metals from picoplankton by nanoflagellate grazing. Limnology and Oceanography, 40(8):1418–1429, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Thamdrup:1996:PCO

[TC96] Bo Thamdrup and Donald E. Canfield. Pathways of carbon oxidation in continental margin sediments off central Chile. Limnology and Oceanography, 41(8):1629–1650, December 1996. CODEN LI-OCAH. ISSN 0024-3590.

Twiss:1996:RRT

[TCA96] Michael R. Twiss, Peter G. C. Campbell, and Jean-Christian Auclair. Regeneration, recycling, and trophic transfer of trace metals by microbial food-web organisms in the pelagic surface waters of Lake Erie. Limnology and Oceanography, 41(7):1425–1437, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Tyler:1998:EDC

[TCB^{+}98] Scott W. Tyler, Peter G. Cook, Anya Z. Butt, James M. Thomas, Peter T. Doran, and W. Berry Lyons. Evidence of deep circulation in two perennially ice-covered Antarctic lakes. Limnology and Oceanography, 43(4):625–635, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Tessier:1993:MCP

[TCCA93] A. Tessier, Y. Couillard, P. G. C. Campbell, and J. C. Auclair. Modeling Cd partitioning in oxic lake sediments and Cd concentrations in the freshwater bivalve Anodonta grandis. Limnology and Oceanography, 38(1):1–17, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Tassan:1995:AAA

[TF95] Stelvio Tassan and Giovanni M. Ferrari. An alternative approach to absorption measurements of aquatic particles retained on filters. Limnology and Oceanography, 40(8):1358–1368, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Taghon:1992:UDS

[TG92] Gary L. Taghon and Ron R. Greene. Utilization of deposited and suspended particulate matter by benthic “interface” feeders. Limnology and Oceanography, 37(7):1370–1391, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Thompson:1994:PSM

[TG94] T. Lewis Thompson and Edward P. Glenn. Plaster standards to measure water motion. Limnology and Oceanography, 39(7): 1768–1779, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Trueblood:1994:TSS

[TGG94] Dwight D. Trueblood, Eugene D. Gallagher, and Diane M. Gould. Three stages of seasonal succession on the Savin Hill Cove mudflat, Boston Harbor. Limnology and Oceanography, 39(6):1440–1454, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Turner:1996:GLB

[TGSB96] Benjamin F. Turner, L. R. Gardner, W. E. Sharp, and E. R. Blood. The geochemistry of Lake Bosumtwi, a hydrologically closed basin in the humid zone of tropical Ghana. Limnology and Oceanography, 41(7):1415–1424, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Taylor:1992:DGM

[TH92] Derek J. Taylor and Paul D. N. Hebert. Daphnia galeata mendotae as a cryptic species complex with interspecific hybrids. Limnology and Oceanography, 37(3):658–665, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Teal:1996:IVS

[TH96] John M. Teal and Brian L. Howes. Interannual variability of a salt-marsh ecosystem. Limnology and Oceanography, 41(8):802– 809, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Tapper:1998:TVL

[TH98] Mark A. Tapper and Randall E. Hicks. Temperate viruses and lysogeny in Lake Superior bacterioplankton. Limnology and Oceanography, 43(1):95–103, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Tsuda:1998:EGC

[THH98] Atsushi Tsuda, Saito Hiroaki, and Taro Hirose. Effect of gut content on the vulnerability of copepods to visual predation. Limnology and Oceanography, 43(8):1944–1947, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Thiel:1994:BRD

[Thi94] Hjalmar Thiel. Book review: Deep-sea food chains and the global carbon cycle (G. T. Rowe and V. Pariente [eds.]). Limnology and Oceanography, 39(6):1497–1498, September 1994. CODEN LI-OCAH. ISSN 0024-3590.

Thingstad:1999:BPI

[THK^{+}99] T. F. Thingstad, H. Havskum, H. Kaas, T. G. Nielsen, B. Riemann, D. Lefevre, and P. J. le B. Williams. Bacteria–protist interactions and organic matter degradation under P-limited conditions: Analysis of an enclosure experiment using a simple model. Limnology and Oceanography, 44(1):62–79, January 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Thorpe:1990:BRS

[Tho90] S. A. Thorpe. Book review: The sea. Ocean engineering science (B. LeMahaute and D. M. Hanes [eds.]). Limnology and Oceanography, 35(7):1656, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Thorpe:1999:OTS

[Tho99] S. A. Thorpe. Observations of the thermal structure of a lake using a submarine. Limnology and Oceanography, 44(6):1575– 1582, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Threlkeld:1990:BRR

[Thr90] Stephen T. Threlkeld. Book review: Reservoir limnology; Ecological perspectives (K. W. Thornton. B. L. Kimmel, and F. E. Payne [eds.]). Limnology and Oceanography, 35(6):1411–1412, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Threlkeld:1991:FIO

[Thr91] Stephen T. Threlkeld. The freshwater initiative: an opportunity to enhance limnology at the national science foundation. Limnology and Oceanography, 36(5):1062–1065, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Threlkeld:1992:UEL

[Thr92] Stephen T. Threlkeld. The uses of ecology: Lake Washington and beyond. Limnology and Oceanography, 37(6):1340, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Thingstad:1997:ADD

[THR97] T. Frede Thingstad, Åke Hagström, and Fereidoun Rassoulzadegan. Accumulation of degradable DOC in surface waters: Is it caused by a malfunctioning microbialloop? Limnology and Oceanography, 42(2):398–404, March 1997. CODEN LIOCAH. ISSN 0024-3590.

Turner:1991:CEE

[THS^{+}91] Michael A. Turner, E. Todd Howell, Martha Summerby, Raymond H. Hesslein, David L. Findlay, and Michael B. Jackson. Changes in epilithon and epiphyton associated with experimental acidification of a lake to pH 5. Limnology and Oceanography, 36 (7):1390–1405, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Tanoue:1996:DDP

[TIM96] Eiichiro Tanoue, Masao Ishii, and Takashi Midorikawa. Discrete dissolved and particulate proteins in oceanic waters. Limnology and Oceanography, 41(6):1334–1343, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Thomas:1999:PRN

[TIOS99] Helmuth Thomas, Venugopalan Ittekkot, Christoph Osterroht, and Bernd Schneider. Preferential recycling of nutrients — the ocean’s way to increase new production and to pass nutrient limitation? Limnology and Oceanography, 44(8):1999–2004, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Tiselius:1992:BAT

[Tis92] Peter Tiselius. Behavior of Acartia tonsa in patchy food environments. Limnology and Oceanography, 37(8):1640–1651, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Thorpe:1998:EIW

[TJ98] S. A. Thorpe and R. Jiang. Estimating internal waves and diapycnal mixing from conventional mooring data in a lake. Limnology and Oceanography, 43(5):936–945, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Thorrold:1997:ROM

[TJC97] Simon R. Thorrold, Cynthia M. Jones, and Steven E. Campana. Response of otolith microchemistry to environmental variations experienced by larval and juvenile Atlantic croaker (Micropogonias undulatus). Limnology and Oceanography, 42(1):102–111, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Thorrold:1998:TES

[TJC^{+}98] Simon R. Thorrold, Cynthia M. Jones, Steven E. Campana, James W. McLaren, and Joseph W. H. Lam. Trace element signatures in otoliths record natal river of juvenile American shad (Alosa sapidissima). Limnology and Oceanography, 43(8):1826– 1835, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Tiselius:1997:ECF

[TJK^{+}97] Peter Tiselius, Per R. Jonsson, Stein Kuurtvedt, Esben Moland Olsen, and Tom Jørstud. Effects of copepod foraging behavior on predation risk: an experimental study of the predatory copepod Pareuchaeta norvegica feeding on Acartia clausi and A. tonsa (Copepoda). Limnology and Oceanography, 42(1):164–170, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Tjossem:1990:EFC

[Tjo90] Sara F. Tjossem. Effects of fish chemical cues on vertical migration behavior of Chaoborus. Limnology and Oceanography, 35(7):1456– 1468, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Tupas:1990:AAA

[TK90] Luis Tupas and Isao Koike. Amino acid and ammonium utilization by heterotrophic marine bacteria grown in enriched seawater. Limnology and Oceanography, 35(5):1145–1155, July 1990. CO-DEN LIOCAH. ISSN 0024-3590.

Tiselius:1998:CDA

[TK98] Peter Tiselius and Thomas Kiørboe. Colonization of diatom aggregates by the dinoflagellate Noctiluca scintillans. Limnology and Oceanography, 43(1):154–159, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Thieberger:1995:NNB

[TKAD95] Y. Thieberger, Z. Kizner, Y. Achituv, and Z. Dubinsky. A novel, nondestructive bioassay for assessing areal chlorophyll-a in hermatypic cnidarians. Limnology and Oceanography, 40(6):1166– 1173, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Tremblay:1997:ERO

[TKL^{+}97] Jean-Eric Tremblay, Bert Klein, Louis Legendre, Richard B. Rivkin, and Jean-Claude Therriault. Estimation of f-ratios in oceans based on phytoplankton size structure. Limnology and Oceanography, 42(3):595–601, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Tremblay:1994:MSF

[TL94] Jean-Eric Tremblay and Louis Legendre. A model for the sizefractionated biomass and production of marine phytoplankton. Limnology and Oceanography, 39(8):2004–2014, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Tanner:1999:MDM

[TLB99] Danny K. Tanner, Edward N. Leonard, and John C. Brazner. Microwave digestion method for phosphorus determination of fish tissue. Limnology and Oceanography, 44(3):708–709, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Treuer:1991:PBS

[TLvB^{+}91] Paul Tréuer, Louis Lindner, Aleido Johannes van Bennekom, Aude Leynaert, Michel Panouse, and Guy Jacques. Production of biogenic silica in the Weddell–Scotia Seas measured with ^{32}Si. Limnology and Oceanography, 36(6):1217–1227, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Tait:1995:MCC

[TM95a] V. K. Tait and R. M. Moore. Methyl chloride (CH_{3}Cl) production in phytoplankton cultures. Limnology and Oceanography, 40(1): 189–195, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Thorrold:1995:RLF

[TM95b] Simon R. Thorrold and A. David McKinnon. Response of larval fish assemblages to a riverine plume in coastal waters of the central Great Barrier Reef lagoon. Limnology and Oceanography, 40(1): 177–181, January 1995. CODEN LIOCAH. ISSN 0024-3590.

Tate:1996:ULL

[TMB96] C. M. Tate, D. M. McKnight, and R. E. Broshears. Using lithium (Li+) as a conservative tracer does not prevent algal uptake of phosphate in an acid mine drainage stream (reply to the comment by Stewart and kszos). Limnology and Oceanography, 41(1):191– 192, January 1996. CODEN LIOCAH. ISSN 0024-3590. See [SK96].

Teranes:1999:SIR

[TMLS99] Jane L. Teranes, Judith A. McKenzie, Andre F. Lotter, and Michael Sturm. Stable isotope response to lake eutrophication: Calibration of a high-resolution lacustrine sequence from baldeggersee, Switzerland. Limnology and Oceanography, 44(2):320–333, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Thomas:1994:RZB

[TN94] KiØboe Thomas and Torkel Gissel Nielsen. Regulation of zooplankton biomass and production in a temperate, coastal ecosystem. 1. Copepods. Limnology and Oceanography, 39(3):493–507, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Thresher:1992:SDB

[TNG^{+}92] Ronald E. Thresher, Peter D. Nichols, John S. Gunn, Barry D. Bruce, and Dianne M. Furlani. Seagrass detritus as the basis of a coastal planktonic food chain. Limnology and Oceanography, 37 (8):1754–1758, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Taroncher-Oldenburg:1997:TVD

[TOKA97] Gaspar Taroncher-Oldenburg, David M. Kulis, and Donald M. Anderson. Toxin variability during the cell cycle of the dinoflagellate Alexandrium fundyense. Limnology and Oceanography, 42(5 (part 2)):1178–1188, July 1997. CODEN LIOCAH. ISSN 0024-

Taylor:1998:EPM

[TR98] Richard Taylor and T. Alwyn V. Rees. Excretory products of mobile epifauna as a nitrogen source for seaweeds. Limnology and Oceanography, 43(4):600–606, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Tranvik:1994:ECO

[Tra94] Lars Tranvik. Effects of colloidal organic matter on the growth of bacteria and protists in lake water. Limnology and Oceanography, 39(6):1276–1285, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Tranvik:1997:RFA

[Tra97] Lars J. Tranvik. Rapid fluorometric assay of bacterial density in lake water and seawater. Limnology and Oceanography, 42(7): 1629–1634, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Tester:1997:GBR

[TS97] Patricia A. Tester and Karen A. Steidinger. Gymnodinium breve red tide blooms: Initiation, transport, and consequences of surface circulation. Limnology and Oceanography, 42(5 (part 2)):1039– 1051, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Townsend:1997:IDH

[TSD97] Colin R. Townsend, Mike R. Scarsbrook, and Sylvain Dolédec. The intermediate disturbance hypothesis, refugia, and biodiversity in streams. Limnology and Oceanography, 42(5):938–949, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Tsuda:1999:EGC

[TSH99] Atsushi Tsuda, Hiroaki Saito, and Taro Hirose. Effect of gut content on the vulnerability of copepods to visual predation. Limnology and Oceanography, 46(4):1175, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Tester:1991:ERT

[TSV^{+}91] Patricia A. Tester, Richard P. Stumpf, Fred M. Vukovich, Patricia K. Fowler, and Jeflerson T. Turner. An expatriate red tide bloom: Transport, distribution, and persistence. Limnology and Oceanography, 36(5):1053–1061, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Trevorrow:1997:ASM

[TT97a] Mark V. Trevorrow and Yuji Tanaka. Acoustic and in situ measurements of freshwater amphipods (Jesogammarus annandalei) in Lake Biwa. Japan. Limnology and Oceanography, 42(1):121– 132, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Turner:1997:TMP

[TT97b] Jefferson T. Turner and Patricia A. Tester. Toxic marine phytoplankton, zooplankton grazers, and pelagic food webs. Limnology and Oceanography, 42(5 (part 2)):1203–1213, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Turner:1993:ZFE

[TTS93] Jeferson T. Turner, Patricia A. Tester, and J. Rudi Strickler. Zooplankton feeding ecology: a cinematographic study of animal-to-animal variability in the feeding behavior of Calanus finmarchicus. Limnology and Oceanography, 38(2):255–264, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Trask:1999:SSO

[TV99] Jennifer L. Trask and Cindy Lee Van Dover. Site-specific and ontogenetic variations in nutrition of mussels (Bathymodiolus sp.) from the Lucky Strike hydrothermal vent field, Mid-Atlantic Ridge. Limnology and Oceanography, 44(2):334–343, March 1999. CODEN LIOCAH. ISSN 0024-3590.

Tollrian:1994:EPC

[TvE94] Ralph Tollrian and Eric von Elert. Enrichment and purification of Chaoborus kairomone from water: Further steps toward its chemical characterization. Limnology and Oceanography, 39(4): 788–796, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Threlkeld:1993:CIO

[TW93] Stephen T. Threlkeld and Ruth L. Willey. Colonization, interaction, and organization of cladoceran epibiont communities. Limnology and Oceanography, 38(3):584–591, May 1993. CODEN LI-OCAH. ISSN 0024-3590.

Twombly:1994:CDP

[Two94] Saran Twombly. Comparative demography and population dynamics of two coexisting copepods in a Venezuelan floodplain lake. Limnology and Oceanography, 39(2):234–247, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Tessier:1992:PDB

[TYL92] Alan J. Tessier, Andrew Young, and Mathew Leibold. Population dynamics and body-size selection in Daphnia. Limnology and Oceanography, 37(1):1–13, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Tamburri:1996:SFB

[TZF96] Mario N. Tamburri and Richard K. Zimmer-Faust. Suspension feeding: Basic mechanisms controlling recognition and ingestion of larvae. Limnology and Oceanography, 41(6):1188–1197, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Turner:1994:SOC

[TZFP^{+}94] E. J. Turner, R. K. Zimmer-Faust, M. A. Palmer, M. Luckenbach, and N. D. Pentcheff. Settlement of oyster (Crassostrea virginica) larvae: Effects of water flow and a water-soluble chemical cue. Limnology and Oceanography, 39(7):1579–1593, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Tittel:1998:RBP

[TZGS98] Jörg Tittel, Barbara Zippel, Walter Geller, and Juliane Seeger. Relationships between plankton community structure and plankton size distribution in lakes of northern Germany. Limnology and Oceanography, 43(6):1119–1132, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Thingstad:1998:PLH

[TZR98] T. Frede Thingstad, Ulla Li Zweifel, and Fereidoun Rassoulzadegan. P limitation of heterotrophic bacteria and phytoplankton in the northwest Mediterranean. Limnology and Oceanography, 43 (1):88–94, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Uher:1997:PPC

[UA97] Günther Uher and Meinrat O. Andreae. Photochemical production of carbonyl sulfide in North Sea water: a process study. Limnology and Oceanography, 42(3):432–442, May 1997. CODEN LI-OCAH. ISSN 0024-3590.

Urban:1994:SRD

[UBBS94] N. R. Urban, P. L. Brezonik, L. A. Baker, and L. A. Sherman. Sulfate reduction and diffusion in sediments of Little Rock Lake, Wisconsin. Limnology and Oceanography, 39(4):797–815, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Urbach:1998:GDP

[UC98] Ena Urbach and Sallie W. Chisholm. Genetic diversity in Prochlorococcus populations flow cytometrically sorted from the Sargasso Sea and Gulf Stream. Limnology and Oceanography, 43 (7):1615–1630, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Urabe:1997:PLD

[UCS97] Jotaro Urabe, Jessica Clasen, and Robert W. Sterner. Phosphorus limitation of Daphnia growth: Is it real? Limnology and Oceanography, 42(6):1436–1443, September 1997. CODEN LI-OCAH. ISSN 0024-3590.

Upstill-Goddard:1998:PCL

[UG98] Robert C. Upstill-Goddard. Physics and chemistry of lakes. Limnology and Oceanography, 43(6):1400–1401, September 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Urban:1990:GPC

[UGU^{+}90] Noel R. Urban, Eville Gorham, John K. Underwood, Frank B. Martin, and J. Gordon Ogden III. Geochemical processes controlling concentrations of Al, Fe, and Mn in Nova Scotia lakes. Limnology and Oceanography, 35(7):1516–1534, November 1990. CODEN LIOCAH. ISSN 0024-3590.

Uhle:1999:FNO

[UMS^{+}99] Maria E. Uhle, Stephen A. Macko, Howard J. Spero, David W. Lea, William F. Ruddiman, and Michael H. Engel. The fate of nitrogen in the Orbulina universa Foraminifera-symbiont system determined by nitrogen isotope analyses of shell-bound organic matter. Limnology and Oceanography, 44(8):1968–1977, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Urabe:1995:CMP

[UNK95] Jotaro Urabe, Masami Nakanishi, and Keiichi Kawabata. Contribution of metazoan plankton to the cycling of nitrogen and phosphorus in Lake Biwa. Limnology and Oceanography, 40(2): 232–241, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Utsumi:1998:DDM

[UNN^{+}98a] Motoo Utsumi, Yukihiro Nojiri, Takeshi Nakamura, Takeshi Nozawa, Akira Otsuki, Noriko Takamura, Makoto Watanabe, and Humitake Seki. Dynamics of dissolved methane and methane oxidation in dimictic Lake Nojiri during winter. Limnology and Oceanography, 43(1):10–17, January 1998. CODEN LIOCAH. ISSN 0024-3590.

Utsumi:1998:ODM

[UNN^{+}98b] Motoo Utsumi, Yukihiro Nojiri, Takeshi Nakamura, Nozawa Takeshi, Akira Otsuki, and Humitake Seki. Oxidation of dissolved methane in a eutrophic, shallow lake: Lake Kasumigaura, Japan. Limnology and Oceanography, 43(3):471–480, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Underwood:1995:MMC

[UPP95] G. J. C. Underwood, D. M. Paterson, and R. J. Parkes. The measurement of microbial carbohydrate exopolymers from intertidal sediments. Limnology and Oceanography, 40(7):1243–1253, November 1995. CODEN LIOCAH. ISSN 0024-3590.

Urabe:1990:SHV

[Ura90] Jotaro Urabe. Stable horizontal variation in the zooplankton community structure of a reservoir maintained by predation and competition. Limnology and Oceanography, 35(8):1703–1717, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Urabe:1992:PPL

[UW92] Jotaro Urabe and Yasunori Watanabe. Possibility of N or P limitation for planktonic cladocerans: an experimental test. Limnology and Oceanography, 37(2):244–251, March 1992. CODEN LIOCAH. ISSN 0024-3590. See comment [Bre93].

Urabe:1993:ISE

[UW93] Jotaro Urabe and Yasunori Watanabe. Implications of sestonic elemental ratio in zooplankton ecology: Reply to the comment by Brett. Limnology and Oceanography, 38(6):1337–1340, September 1993. CODEN LIOCAH. ISSN 0024-3590. See [Bre93].

Vrede:1999:PDT

[VAH99] Tobias Vrede, Tom Andersen, and Dag O. Hessen. Phosphorus distribution in three crustacean zooplankton species. Limnology and Oceanography, 44(1):225–229, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Vadstein:1995:EPR

[VBAO95] Olav Vadstein, Oddmund Brekke, Tom Andersen, and Yngvar Olsen. Estimation of phosphorus release rates from natural zooplankton communities feeding on planktonic algae and bacteria. Limnology and Oceanography, 40(2):250–262, March 1995. CO-DEN LIOCAH. ISSN 0024-3590.

Vodacek:1997:SVC

[VBD^{+}97] Anthony Vodacek, Neil V. Blough, Michael D. DeGrandpre, Michael D. DeGrandpre, and Robert K. Nelson. Seasonal variation of CDOM and DOC in the Middle Atlantic Bight: Terrestrial inputs and photooxidation. Limnology and Oceanography, 42(4): 674–686, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Voss:1998:SAP

[VBK98] Kenneth J. Voss, William M. Balch, and Katherine A. Kilpatrick. Scattering and attenuation properties of Emiliania huxleyi cells and their detached coccoliths. Limnology and Oceanography, 43 (5):870–876, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Visscher:1991:SCS

[VBvG91] Pieter T. Visscher, Jan Beukema, and Hans van Gemerden. In situ characterization of sediments: Measurements of oxygen and sulfide profiles with a novel combined needle electrode. Limnology and Oceanography, 36(7):1476–1480, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Villareal:1990:DBR

[VC90] Tracy A. Villareal and Edward J. Carpenter. Diel buoyancy regulation in the marine diazotrophic cyanobacterium Trichodesmium thiebautii. Limnology and Oceanography, 35(8):1832–1837, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Verschuren:1999:LTD

[VCT^{+}99] Dirk Verschuren, Christine Cocquyt, John Tibby, C. Neil Roberts, and Peter R. Leavitt. Long-term dynamics of algal and invertebrate communities in a small, fluctuating tropical soda lake. Limnology and Oceanography, 44(5):1216–1231, July 1999. CODEN LIOCAH. ISSN 0024-3590.

Vidal:1999:DON

[VDA99] Montserrat Vidal, Carlos M. Duarte, and Susana Agusti. Dissolved organic nitrogen and phosphorus pools and fluxes in the central Atlantic Ocean. Limnology and Oceanography, 44(1):106– 115, January 1999. CODEN LIOCAH. ISSN 0024-3590.

vandenBosch:1994:MGD

[vdBG94] Frank van den Bosch and Wilfried Gabriel. A model of growth and development in copepods. Limnology and Oceanography, 39 (7):1528–1542, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Vago:1997:GRT

[VDG^{+}97] R. Vago, Z. Dubinsky, A. Genin, M. Ben-Zion, and Z. Kizner. Growth rates of three symbiotic corals in the Red Sea. Limnology and Oceanography, 42(8):1814–1819, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Volten:1998:LMA

[VdHH^{+}98] H. Volten, J. F. de Haan, J. W. Hovenier, R. Schreurs, W. Vassen, A. G. Dekker, H. J. Hoogenboom, F. Charlton, and R. Wouts. Laboratory measurements of angular distributions of light scattered by phytoplankton and silt. Limnology and Oceanography, 43 (6):1180–1197, September 1998. CODEN LIOCAH. ISSN 0024-3590.

vonElert:1997:PLL

[vEJ97] Eric von Elert and Friedrich Jüttner. Phosphorus limitation and not light controls the extracellular release of allelopathic compounds by Trichormus doliolum (Cyanobacteria). Limnology and Oceanography, 42(8):1796–1802, December 1997. CODEN LI-OCAH. ISSN 0024-3590.

Venrick:1993:PSC

[Ven93] E. L. Venrick. Phytoplankton seasonality in the central North Pacific: The endless summer reconsidered. Limnology and Oceanography, 38(6):1135–1149, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Venrick:1997:CPS

[Ven97] E. L. Venrick. Comparison of the phytoplankton species composition and structure in the Climax area (1973–1985) with that of Station ALOHA (1994). Limnology and Oceanography, 42(7): 1643–1648, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Vermeij:1990:BRL

[Ver90] Geerat J. Vermeij. Book review: The legacy of Tethys (F. D. Por). Limnology and Oceanography, 35(4):983, June 1990. CODEN LI-OCAH. ISSN 0024-3590.

Verity:1991:MSP

[Ver91] Peter G. Verity. Measurement and simulation of prey uptake by marine planktonic ciliates fed plastidic and aplastidic nanoplankton. Limnology and Oceanography, 36(4):729–750, June 1991. CO-DEN LIOCAH. ISSN 0024-3590.

Verschuren:1993:LEA

[Ver93] Dirk Verschuren. A lightweight extruder for accurate sectioning of soft-bottom lake sediment cores in the field. Limnology and Oceanography, 38(8):1796–1802, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Verhagen:1994:MPP

[Ver94] Jan H. G. Verhagen. Modeling phytoplankton patchiness under the influence of wind-driven currents inlakes. Limnology and Oceanography, 39(7):1551–1565, November 1994. CODEN LI-OCAH. ISSN 0024-3590.

Verardo:1997:CAM

[Ver97] David J. Verardo. Charcoal analysis in marine sediments. Limnology and Oceanography, 42(1):192–197, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Verschuren:1999:IDM

[Ver99] Dirk Verschuren. Influence of depth and mixing regime on sedimentation in a small, fluctuating tropical soda lake. Limnology and Oceanography, 46(4):1103–1113, June 1999. CODEN LI-OCAH. ISSN 0024-3590.

Vezina:1994:MVN

[Véz94] Main F. Vézina. Mesoscale variability in nitrogen uptake rates and the f-ratio during a coastal phytoplankton bloom. Limnology and Oceanography, 39(4):854–868, June 1994. CODEN LIOCAH. ISSN 0024-3590.

VanDover:1994:MFR

[VF94] Cindy Lee Van Dover and Brian Fry. Microorganisms as food resources at deep-sea hydrothermal vents. Limnology and Oceanography, 39(1):51–57, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Varela:1998:DCP

[VFAA98] Ramiro A. Varela, Francisco G. Figueras, Belén Arbones, and Susana Agustí. Determining the contribution of pigments and the nonalgal fractions to total absorption: Toward an improved algorithm. Limnology and Oceanography, 43(3):449–457, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Viekman:1992:SFO

[VFW^{+}92] Bruce E. Viekman, Roger D. Flood, Mark Wimbush, Mohammad Faghri, Yutaka Asako, and John C. Van Leer. Sedimentary furrows and organized flow structure: a study in Lake Superior. Limnology and Oceanography, 37(4):797–812, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Vodacek:1994:EMS

[VGB94] Anthony Vodacek, Sarah A. Green, and Neil V. Blough. An experimental model of the solar-stimulated fluorescence of chromophoric dissolved organic matter. Limnology and Oceanography, 39(1):1–11, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Vanderploeg:1992:LLC

[VGP^{+}92] Henry A. Vanderploeg, Wayne S. Gardner, Christopher C. Parrish, James R. Liebig, and Joann F. Cavaletto. Lipids and lifecycle strategy of a hypolimnetic copepod in Lake Michigan. Limnology and Oceanography, 37(2):413–424, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Vallino:1996:MBU

[VHH96] J. J. Vallino, C. S. Hopkinson, and J. E. Hobbie. Modeling bacterial utilization of dissolved organic matter: Optimization replaces Monod growth kinetics. Limnology and Oceanography, 41 (8):1591–1609, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Vodacek:1995:USA

[VHS^{+}95] Anthony Vodacek, Frank E. Hogel, Robert N. Swift, James K. Yungel, Edward T. Peltzer, and Neil V. Blough. The use of in situ and airborne fluorescence measurements to determine UV absorption coefficients and DOC concentrations in surface waters. Limnology and Oceanography, 40(2):411–415, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Vogler:1996:SLC

[VJM96] S. Vogler, M. Jung, and A. Mangini. Scavenging of 234Th and 7be in Lake Constance. Limnology and Oceanography, 41(7):1384– 1393, November 1996. CODEN LIOCAH. ISSN 0024-3590.

Vijverberg:1996:DDE

[VKB96] Jacobus Vijverberg, Dennis F. Kalf, and Maarten Boersma. Decrease in Daphnia egg viability at elevated pH. Limnology and Oceanography, 41(8):789–794, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Vanek:1991:RRS

[VL91] Vladimir Vanek and David Robert Lee. Reconstruction of a regional, 12-yr silica decline in lakes by means of fossil sponge spicules. Limnology and Oceanography, 36(6):1250–1262, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Vinebrooke:1996:EUR

[VL96] Rolf D. Vinebrooke and Peter R. Leavitt. Effects of ultraviolet radiation on periphyton in an alpine lake. Limnology and Oceanography, 41(5):1035–1040, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Vinebrooke:1998:DIE

[VL98] Rolf D. Vinebrooke and Peter R. Leavitt. Direct and interactive effects of allochthonous dissolved organic matter, inorganic nutrients, and ultraviolet radiation on an alpine littoral food web. Limnology and Oceanography, 43(6):1065–1081, September 1998. CODEN LIOCAH. ISSN 0024-3590.

VanDonk:1997:ACW

[VLHL97] E. Van Donk, M. Lürling, D. O. Hessen, and G. M. Lokhorst. Altered cell wall morphology in nutrient-deficient phytoplankton and its impact on grazers. Limnology and Oceanography, 42(2): 357–364, March 1997. CODEN LIOCAH. ISSN 0024-3590.

vanLeeuwen:1991:CSB

[vLM91] H. C. van Leeuwen and E. J. Maly. Changes in swimming behavior of male Diaptomus leptopus (Copepoda: Calanoida) in response to gravid females. Limnology and Oceanography, 36(6):1188–1195, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Vrba:1994:RDE

[VM94] Jaroslav Vrba and Jiŕí Macháéek. Release of dissolved extracellular β-N-acetylglucosaminidase during crustacean moulting. Limnology and Oceanography, 39(3):712–716, May 1994. CODEN LI-OCAH. ISSN 0024-3590.

Valett:1996:PLS

[VMDC96] H. Maurice Valett, John A. Morrice, Clifford N. Dahm, and Michael E. Campana. Parent lithology, surface–groundwater exchange, and nitrate retention in headwater streams. Limnology and Oceanography, 41(2):333–345, March 1996. CODEN LI-OCAH. ISSN 0024-3590.

Valiela:1997:MBS

[VMH^{+}97] Ivan Valiela, James McClelland, Jennifer Hauxwell, Peter J. Behr, Douglas Hersh, and Kenneth Foreman. Macroalgal blooms in shallow estuaries: Controls and ecophysiological and ecosystem consequences. Limnology and Oceanography, 42(5 (part 2)):1105–1118, July 1997. CODEN LIOCAH. ISSN 0024-3590.

Vodacek:1992:ESV

[Vod92] Anthony Vodacek. An explanation of the spectral variation in freshwater CDOM fluorescence. Limnology and Oceanography, 37 (8):1808–1813, December 1992. CODEN LIOCAH. ISSN 0024-3590.

Vadstein:1993:RPB

[VORJ93] Olav Vadstein, Yngvar Olsen, Helge Reinertsen, and Arne Jensen. The role of planktonic bacteria in phosphorus cycling in lakes — sink and link. Limnology and Oceanography, 38(7):1539–1544, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Voss:1992:SMB

[Vos92] Kenneth J. Voss. A spectral model of the beam attenuation coefficient in the ocean and coastal areas. Limnology and Oceanography, 37(3):501–509, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Vidondo:1997:SAA

[VPBD97] Beatriz Vidondo, Yves T. Prairie, Jose M. Blanco, and Carlos M. Duarte. Some aspects of the analysis of size spectra in aquatic ecology. Limnology and Oceanography, 42(1):184–192, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Vaulot:1990:WPP

[VPN^{+}90] Daniel Vaulot, Frédéric Partensky, Jacques Neveux, R. Fauzi C. Mantoura, and Carole A. Llewellyn. Winter presence of prochlorophytes in surface waters of the northwestern Mediterranean Sea. Limnology and Oceanography, 35(5):1156–1164, July 1990. CO-DEN LIOCAH. ISSN 0024-3590.

Viitasalo:1999:SCF

[VRHK99] Markku Viitasalo, Mirja Rosenberg, Anna-Stiina Heiskanen, and Marja Koski. Sedimentation of copepod fecal material in the coastal northern Baltic Sea: Where did all the pellets go? Limnology and Oceanography, 44(6):1388–1399, September 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Vincent:1998:TAI

[VRL^{+}98] Warwick F. Vincent, Rowena Rae, Isabelle Laurion, Clive Howard-Williams, and John C. Priscu. Transparency of Antarctic ice-covered lakes to solar UV radiation. Limnology and Oceanography, 43(4):618–624, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Verity:1992:RBC

[VRT^{+}92] Peter G. Verity, Charles Y. Robertson, Craig R. Tronzo, Melinda G. Andrews, James R. Nelson, and Michael E. Sieracki. Relationships between cell volume and the carbon and nitrogen content of marine photosynthetic nanoplankton. Limnology and Oceanography, 37(7):1434–1446, November 1992. CODEN LI-OCAH. ISSN 0024-3590.

Vago:1994:CTI

[VSBZ^{+}94] R. Vago, Y. Shai, M. Ben-Zion, Z. Duhinsky, and Y. Achituv. Computerized tomography and image analysis: a tool for examining the skeletal characteristics of reef-building organisms. Limnology and Oceanography, 39(2):448–452, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Vanni:1990:SPG

[VT90] Michael J. Vanni and Jo Temte. Seasonal patterns of grazing and nutrient limitation of phytoplankton in a eutrophic lake. Limnology and Oceanography, 35(3):697–709, May 1990. CODEN LI-OCAH. ISSN 0024-3590.

Vaquer:1996:SSD

[VTCB96] André Vaquer, Marc Troussellier, Claude Courties, and Bertrand Bibent. Standing stock and dynamics of picophytoplankton in the Thau Lagoon (northwest Mediterranean coast). Limnology and Oceanography, 41(4):1821–1828, December 1996. CODEN LIOCAH. ISSN 0024-3590.

Vadeboncoeur:1998:DIC

[VV98] Yvonne Vadeboncoeur and David M. Vadeboncoeur. Dissolved inorganic carbon sources for epipelic algal production: Sensitivity of primary production estimates to spatial and temporal distribution of ^{14}C. Limnology and Oceanography, 43(6):1222–1226, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Vrede:1999:ENP

[VVIK99] Katarina Vrede, Tobias Vrede, Anneli Isaksson, and Anna Karlsson. Effects of nutrients (phosphorous, nitrogen, and carbon) and zooplankton on bacterioplankton and phytoplankton — a seasonal study. Limnology and Oceanography, 44(7):1616–1624, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Volk:1997:CCB

[VVK97] Christian J. Volk, Catherine B. Volk, and Louis A. Kaplan. Chemical composition of biodegradable dissolved organic matter in streamwater. Limnology and Oceanography, 42(1):39–44, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Vavrus:1996:MSS

[VWF96] Stephen J. Vavrus, Randolph H. Wynne, and Jonathan A. Foley. Measuring the sensitivity of southern Wisconsin lake ice to climate variations and lake depth using a numerical model. Limnology and Oceanography, 41(5):822–831, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Wigand:1998:EIA

[WAC^{+}98] C. Wigand, F. O. Andersen, K. K. Christensen, M. Holmer, and H. S. Jensen. Endomycorrhizae of isoetids along a biogeochemical gradient. Limnology and Oceanography, 43(3):508–515, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Walters:1992:FSL

[Wal92] Linda J. Walters. Field settlement locations on subtidal marine hard substrata: Is active larval exploration involved? Limnology and Oceanography, 37(5):1101–1107, July 1992. CODEN LI-OCAH. ISSN 0024-3590.

Walker:1996:BRR

[Wal96] K. F. Walker. Book review: River and stream ecosystems (C. E. Cushing, K. W. Cummins, and G. W. Minshall [eds]). Limnology and Oceanography, 41(3):579–580, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Warren:1990:BRP

[War90] Bruce A. Warren. Book review: Poleward flows along eastern ocean boundaries (S. J. Neshyba, C. N. K. Mooers, R. L. Smith, and R. T. Barber [eds.]). Limnology and Oceanography, 35(5): 1219–1220, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Wassmann:1990:RBP

[Was90] Paul Wassmann. Relationship between primary and export production in the boreal coastal zone of the North Atlantic. Limnology and Oceanography, 35(2):464–471, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Watras:1992:MMI

[WB92] Carl J. Watras and Nicolas S. Bloom. Mercury and methylmercury, in individual zooplankton: Implications for bioaccumulation. Limnology and Oceanography, 37(6):1313–1318, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Wolford:1996:HME

[WB96] Ross A. Wolford and Roger C. Bales. Hydrochemical modeling of Emerald Lake watershed, Sierra Nevada, California: Sensitivity of stream chemistry to changes in fluxes and model parameters. Limnology and Oceanography, 41(5):947–954, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1991:CDM

[WBG^{+}91] Patrick J. Walsh, Patricia Blackwelder, Kenneth A. Gill, Eva Danulat, and Thomas P. Mommsen. Carbonate deposits in marine fish intestines: a new source of biomineralization. Limnology and Oceanography, 36(6):1227–1232, September 1991. CODEN LI-OCAH. ISSN 0024-3590.

Wiackowski:1994:DEZ

[WBG94] Krzysztof Wiackowski, Michael T. Brett, and Charles R. Goldman. Differential effects of zooplankton species on ciliate community structure. Limnology and Oceanography, 39(3):486–492, May 1994. CODEN LIOCAH. ISSN 0024-3590.

Wood:1995:DMT

[WBKF95] T. W. Wood, A. M. Baptista, J. S. Kuwabara, and A. R. Flegal. Diagnostic modeling of trace metal partitioning in south San Francisco Bay. Limnology and Oceanography, 40(2):345–358, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Williams:1993:GHC

[WBM93] Mark W. Williams, Aaron D. Brown, and John M. Melack. Geochemical and hydrologic controls on the composition of surface water in a high-elevation basin. Sierra Nevada, California. Limnology and Oceanography, 38(4):775–797, June 1993. CODEN LIOCAH. ISSN 0024-3590.

Weyhenmeyer:1999:CPS

[WBP99] Gesa A. Weyhenmeyer, Thorsten Blenckner, and Kurt Pettersson. Changes of the plankton spring outburst related to the North Atlantic Oscillation. Limnology and Oceanography, 44(7):1788– 1792, November 1999. CODEN LIOCAH. ISSN 0024-3590.

Whitmore:1991:DFL

[WBRJ91] Thomas J. Whitmore, Mark Brenner, Brian E. Rood, and Kate E. Japy. Deoxygenation of a Florida lake during winter mixing. Limnology and Oceanography, 36(3):577–585, May 1991. CODEN LI-OCAH. ISSN 0024-3590.

Wing:1998:MDC

[WBRL98] Stephen R. Wing, Louis W. Botsford, Stephen V. Ralston, and John L. Largier. Meroplanktonic distribution and circulation in a coastal retention zone of the northern California upwelling system. Limnology and Oceanography, 43(7):1710–1721, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Wylie:1991:RIB

[WC91] John L. Wylie and David J. Currie. The relative importance of bacteria and algae as food sources for crustacean zooplankton. Limnology and Oceanography, 36(4):708–728, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1994:BHS

[WCD^{+}94] Russell S. Walsh, Gregory A. Cutter, William M. Dunstan, Joël Radford-Knoery, and Jennifer T. Elder. The biogeochemistry of hydrogen sulfide: Phytoplankton production in the surface ocean. Limnology and Oceanography, 39(4):941–948, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Willey:1990:EEF

[WCT90] Ruth L. Willey, Patti A. Cantrell, and Stephen T. Threlkeld. Epibiotic euglenoid flagellates increase the susceptibility of some zooplankton to fish predation. Limnology and Oceanography, 35 (4):952–959, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Wallace:1991:EFO

[WCW^{+}91] J. Bruce Wallace, T. F. Cuffney, J. R. Webster, G. J. Lugthart, K. Chung, and B. S. Goldowitz. Export of fine organic particles from headwater streams: Effects of season, extreme discharges, and invertebrate manipulation. Limnology and Oceanography, 36 (4):670–682, June 1991. CODEN LIOCAH. ISSN 0024-3590.

Weissenberger:1992:SIC

[WDGS92] Jiirgen Weissenberger, Gerhard Dieckmann, Rolf Gradinger, and Michael Spindler. Sea ice: a cast technique to examine and analyze brine pockets and channel structure. Limnology and Oceanography, 37(1):179–183, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Webster:1990:EWD

[Web90] I. T. Webster. Effect of wind on the distribution of phytoplankton cells in lakes. Limnology and Oceanography, 35(5):989–1001, July 1990. CODEN LIOCAH. ISSN 0024-3590.

Webster:1992:WES

[Web92] Ian T. Webster. Wave enhancement of solute exchange within empty burrows. Limnology and Oceanography, 37(3):630–643, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Webster:1995:BRA

[Web95] Katherine E. Webster. Book review: Acidification of freshwater ecosystems: Implications for the future (C. E. W. Steinberg and R. F. Wright [eds.]). Limnology and Oceanography, 40(4):842–843, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Welschmeyer:1994:FAC

[Wel94] Nicholas A. Welschmeyer. Fluorometric analysis of chlorophylla in the presence of chlorophyll-b and pheopigments. Limnology and Oceanography, 39(8):1985–1992, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Welch:1999:FER

[Wel99a] Harold E. Welch. Freshwater ecosystems: Revitalizing educational programs in limnology. Limnology and Oceanography, 44(1):235– 236, January 1999. CODEN LIOCAH. ISSN 0024-3590.

Wells:1999:MIA

[Wel99b] Mark L. Wells. Manipulating iron availability in nearshore waters. Limnology and Oceanography, 44(4):1002–1008, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1991:BRE

[Wet91a] Robert G. Wetzel. Book review: Ecology and management of aquatic vegetation in the Indian subcontinent (B. Gopal [ed.]). Limnology and Oceanography, 36(5):1066–1068, July 1991. CO-DEN LIOCAH. ISSN 0024-3590.

Wetzel:1991:LER

[Wet91b] Robert G. Wetzel. Limnological education — reply to the comment by Kalff. Limnology and Oceanography, 36(7):1502, November 1991. CODEN LIOCAH. ISSN 0024-3590. See [Kal91].

Wetzel:1991:TLN

[Wet91c] Robert G. Wetzel. On the teaching of limnology: Need for a national initiative. Limnology and Oceanography, 36(1):213–215, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Wetzgel:1991:BRW

[Wet91d] Robert B. Wetzgel. Book review: Wetland ecology and management: Case studies (Whigham, D. F., R. E. Good, and J. Květ [eds.]). Limnology and Oceanography, 36(5):1066–1068, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1993:ADS

[Wet93] Robert G. Wetzel. 1993 ASLO distinguished service award. Limnology and Oceanography, 38(7):1592, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1995:BRL

[Wet95] Robert G. Wetzel. Book review: Lake Veluwe, a macrophytedominated system under eutrophication stress (W. Van Vierssen, M. Hootsmans, and J. Vermaat [eds.]). Limnology and Oceanography, 40(3):646–647, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1998:FSA

[Wet98] Robert G. Wetzel. Freshwaters [sic] of Alaska: Ecological syntheses. Limnology and Oceanography, 43(5):1021, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1999:LEY

[Wet99] Robert G. Wetzel. Lake Erken: 50 years of limnological research. Limnology and Oceanography, 46(4):1184, June 1999. CODEN LIOCAH. ISSN 0024-3590.

Watson:1990:MSP

[WF90] Peter G. Watson and Trish E. Frickers. A multilevel in situ pore-water sampler for use in intertidal sediments and laboratory microcosms. Limnology and Oceanography, 35(6):1381–1389, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Wang:1996:ATE

[WF96] Wen-Xiong Wang and Nicholas S. Fisher. Assimilation of trace elements and carbon by the mussel Mytilus edulis: Effects of food composition. Limnology and Oceanography, 41(2):197–207, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Wang:1998:ATE

[WF98] Wen-Xiong Wang and Nicholas S. Fisher. Accumulation of trace elements in a marine copepod. Limnology and Oceanography, 43 (2):273–283, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Webster:1999:PWC

[WFG99] Ian T. Webster, Phillip W. Ford, and Nicola J. Grigg. Pore-water convection induced by peeper emplacement in saline sediment. Limnology and Oceanography, 44(2):425–430, March 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Wells:1994:DCN

[WG94] Mark L. Wells and Edward D. Goldberg. The distribution of colloids in the North Atlantic and Southern Oceans. Limnology and Oceanography, 39(2):286–302, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Weers:1997:GRD

[WG97] Paul M. M. Weers and Ramesh D. Gulati. Growth and reproduction of Daphnia galeata in response to changes in fatty acids, phosphorus, and nitrogen in Chlamydomonas reinhardtii. Limnology and Oceanography, 42(7):1584–1589, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Wu:1997:ACW

[WGS97] Hanguo Wu, Mark Green, and Mary I. Scranton. Acetate cycling in the water column and surface sediment of Long Island Sound following a bloom. Limnology and Oceanography, 42(4):705–713, June 1997. CODEN LIOCAH. ISSN 0024-3590.

Welschmeyer:1991:PGH

[WGSP91] Nicholas Welschmeyer, Ralf Goericke, Suzanne Strom, and Willis Peterson. Phytoplankton growth and herbivory in the subarctic Pacific: a chemotaxonomic analysis. Limnology and Oceanography, 36(8):1631–1649, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Wikner:1991:ASB

[WH91] J. Wikner and Å. Hagström. Annual study of bacterioplankton community dynamics. Limnology and Oceanography, 36(7):1313– 1324, November 1991. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1992:IPS

[WH92] Russell S. Walsh and Keith A. Hunter. Influence of phosphorus storage on the uptake of cadmium by the marine alga Macrocystis pyrifera. Limnology and Oceanography, 37(7):1361–1369, November 1992. CODEN LIOCAH. ISSN 0024-3590.

Werner:1993:PAC

[WH93a] Ingeborg Werner and James T. Hollibaugh. Potamocorbula amurensis: Comparison of clearance rates and assimilation efficiencies for phytoplankton and bacterioplankton. Limnology and Oceanography, 38(5):949–964, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Wilson:1993:ICP

[WH93b] Chris C. Wilson and Paul D. N. Hebert. Impact of copepod predation on distribution patterns of Daphnia pulex clones. Limnology and Oceanography, 38(6):1304–1310, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Webster:1994:EWD

[WH94a] Ian T. Webster and Paul A. Hutchinson. Effect of wind on the distribution of phytoplankton cells in lakes revisited. Limnology and Oceanography, 39(2):365–373, March 1994. CODEN LIOCAH. ISSN 0024-3590.

White:1994:LTN

[WH94b] David S. White and Brian L. Howes. Long-term ^{15}N-nitrogen retention in the vegetated sediments of a New England salt marsh. Limnology and Oceanography, 39(8):1878–1892, December 1994. CODEN LIOCAH. ISSN 0024-3590.

White:1994:NID

[WH94c] David S. White and Brian L. Howes. Nitrogen incorporation into decomposing litter of Spartina alterniflora. Limnology and Oceanography, 39(1):133–140, January 1994. CODEN LIOCAH. ISSN 0024-3590.

Weaver:1995:BAV

[WH95] Daniel T. Weaver and Randall E. Hicks. Biodegradation of Azotobacter vinelandii exopolymer by Lake Superior microbes. Limnology and Oceanography, 40(6):1035–1041, September 1995. CO-DEN LIOCAH. ISSN 0024-3590.

Wallace:1999:EVI

[WH99] Brett B. Wallace and David P. Hamilton. The effect of variations in irradiance on buoyancy regulation in Microcystis aeruginosa. Limnology and Oceanography, 44(2):273–281, March 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Wetzel:1995:NPU

[WHB95] Robert G. Wetzel, Patrick G. Hatcher, and Thomas S. Bianchi. Natural photolysis by ultraviolet irradiance of recalcitrant dissolved organic matter to simple substrates for rapidbacterial metabolism. Limnology and Oceanography, 40(8):1369–1380, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Wheatcroft:1991:BRT

[Whe91] Robert A. Wheatcroft. Book review: Trace fossils: Biology and taphonomy (R. G. Bromley). Limnology and Oceanography, 36(1): 216–217, January 1991. CODEN LIOCAH. ISSN 0024-3590.

Wheatcroft:1992:ETP

[Whe92] Robert A. Wheatcroft. Experimental tests for particle sizedependent bioturbation in the deep ocean. Limnology and Oceanography, 37(1):90–104, January 1992. CODEN LIOCAH. ISSN 0024-3590.

Wallmann:1993:NPD

[WHK^{+}93] Klaus Wallmann, Kristof Hennies, Iris König, Wilhelm Petersen, and Hans-Diethard Knauth. New procedure for determining reactive Fe(III) and Fe(II) minerals in sediments. Limnology and Oceanography, 38(8):1803–1812, December 1993. CODEN LI-OCAH. ISSN 0024-3590.

Webster:1994:URI

[WHM94] Ian T. Webster, Gary J. Hancock, and Andrew S. Murray. Use of radium isotopes to examine pore-water exchange in an estuary. Limnology and Oceanography, 39(8):1917–1927, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Weyhenmeyer:1997:VMD

[WHM97] Gesa A. Weyhenmeyer, Lars Håkanson, and Markus Meili. A validated model for daily variations in the flux, origin, and distribution of settling particles within lakes. Limnology and Oceanography, 42(7):1517–1529, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Waser:1998:NIF

[WHN^{+}98] N. A. D. Waser, P. J. Harrison, B. Nielsen, S. E. Calvert, and D. H. Turpin. Nitrogen isotope fractionation during the uptake and assimilation of nitrate, nitrite, ammonium, and urea by a marine diatom. Limnology and Oceanography, 43(2):215–224, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Williamson:1999:DUP

[WHOL99] Craig E. Williamson, Bruce R. Hargreaves, Pamela S. Orr, and Patricia A. Lovera. Does UV play a role in changes in predation and zooplankton community structure in acidified lakes? Limnology and Oceanography, 44(3 (part 2)):774–783, May 1999. CO-DEN LIOCAH. ISSN 0024-3590.

Wickham:1995:TRB

[Wic95] Stephen A. Wickham. Trophic relations between cyclopoid copepods and ciliated protists: Complex interactions link the microbial and classic food webs. Limnology and Oceanography, 40(6): 1173–1181, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Wilzbach:1990:NDB

[Wil90] Margaret A. Wilzbach. Nonconcordance of drift and benthic activity in Baetis. Limnology and Oceanography, 35(4):945–952, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Willows:1992:ODI

[Wil92] Robert I. Willows. Optimal digestive investment: a model for filter feeders experiencing variable diets. Limnology and Oceanography, 37(4):829–847, June 1992. CODEN LIOCAH. ISSN 0024-3590.

Williamson:1995:WRD

[Wil95] Craig E. Williamson. What role does UV-B radiation play in freshwater ecosystems? Limnology and Oceanography, 40(2):386– 392, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Wildish:1998:EMB

[Wil98] D. J. Wildish. Ecology of marine bivalves: an ecosystem approach. Limnology and Oceanography, 43(7):1764, November 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Wilkinson:1997:DRP

[WJRB97] Kevin J. Wilkinson, Anthony Joz-Roland, and Jacques Buffle. Different roles of pedogenic fulvic acids and aquagenic biopolymers on colloid aggregation and stability in freshwaters. Limnology and Oceanography, 42(8):1714–1724, December 1997. CODEN LIOCAH. ISSN 0024-3590.

Wheeler:1990:ARL

[WK90] Patricia A. Wheeler and Steven A. Kokkinakis. Ammonium recycling limits nitrate use in the oceanic subarctic Pacific. Limnology and Oceanography, 35(6):1267–1278, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Wanninkhof:1996:CEC

[WK96] Rik Wanninkhof and Molly Knox. Chemical enhancement of CO_{2} exchange in natural waters. Limnology and Oceanography, 41(4): 689–697, June 1996. CODEN LIOCAH. ISSN 0024-3590.

Webster:1996:ILP

[WKB^{+}96] Katherine E. Webster, Timothy K. Kratz, Carl J. Bowser, John J. Magnuson, and William J. Rose. The influence of landscape position on lake chemical responses to drought in northern Wisconsin. Limnology and Oceanography, 41(5):977–984, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Watanabe:1991:DVM

[WKK91] Masataka Watanabe, Kunio Kohata, and Toshihiko Kimura. Diel vertical migration and nocturnal uptake of nutrients by Chattonella antiqua under stable stratification. Limnology and Oceanography, 36(3):593–602, May 1991. CODEN LIOCAH. ISSN 0024-3590.

Watanabe:1995:GCA

[WKK^{+}95] Masataka Watanabe, Kunio Kohata, Toshihiko Kimura, Takejiro Takamatsu, Shin ichiro Yamaguchi, and Teru Ioriya. Generation of a Chattonella antiqua bloom by imposing a shallow nutricline in a mesocosm. Limnology and Oceanography, 40(8):1447–1460, December 1995. CODEN LIOCAH. ISSN 0024-3590.

Witkowski:1998:LSM

[WKZK98] Konrad Witkowski, Tadeusz Król, Andrzej Zielirińki, and Edward Kuteń. A light-scattering matrix for unicellular marine phytoplankton. Limnology and Oceanography, 43(5):859–869, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Wainman:1992:CFL

[WL92] Bruce C. Wainman and David R. S. Lean. Carbon fixation into lipid in small freshwater lakes. Limnology and Oceanography, 37 (5):956–965, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Wu:1994:SFI

[WL94] Jingfeng Wu and George W. Luther III. Size-fractionated iron concentrations in the water column of the western North Atlantic Ocean. Limnology and Oceanography, 39(5):1119–1129, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Wiltshire:1999:MPI

[WL99] Karen Helen Wiltshire and Winfried Lampert. Max Planck Institute for Limnology, P.O. Box 165, 24302 Plön, Germany. Limnology and Oceanography, 44(8):1894–1903, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Wing:1995:STB

[WLBQ95] Stephen R. Wing, John L. Largier, Louis W. Botsford, and James F. Quinn. Settlement and transport of benthic invertebrates in an intermittent upwelling region. Limnology and Oceanography, 40(2):316–329, March 1995. CODEN LIOCAH. ISSN 0024-3590.

Williams:1996:CCH

[WLCG96] Mark W. Williams, Mark Losleben, Nel Caine, and David Greenland. Changes in climate and hydrochemical responses in a highelevation catchment in the rocky mountains, USA. Limnology and Oceanography, 41(5):939–946, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Watson:1991:DSS

[WLD91] Andrew Watson, Peter Liss, and Robert Duce. Design of a smallscale in situ iron fertilization experiment. Limnology and Oceanography, 36(8):1960–1965, December 1991. CODEN LIOCAH. ISSN 0024-3590.

Wing:1993:DMW

[WLD93] Stephen R. Wing, James J. Leichter, and Mark W. Denny. A dynamic model for wave-induced light fluctuations in a kelp forest. Limnology and Oceanography, 38(2):396–407, March 1993. CODEN LIOCAH. ISSN 0024-3590.

Weissman:1993:EPC

[WLY93] Penny Weissman, Darcy J. Lonsdale, and Jeannette Yen. The effect of peritrich ciliates on the production of Acartia hudsonica in Long Island Sound. Limnology and Oceanography, 38(3):613– 622, May 1993. CODEN LIOCAH. ISSN 0024-3590.

Wei:1992:TVA

[WM92] Ching-Ling Wei and James W. Murray. Temporal variations of 234Th activity in the water column of Dabob Bay: Particle scavenging. Limnology and Oceanography, 37(2):296–314, March 1992. CODEN LIOCAH. ISSN 0024-3590.

Wynne:1996:DTC

[WMC^{+}96] Randolph H. Wynne, John J. Magnuson, Murray K. Clayton, Thomas M. Lillesand, and Daniel C. Rodman. Determinants of temporal coherence in the satellite-derived 1987–1994 ice breakup dates of lakes on the Laurentian Shield. Limnology and Oceanography, 41(5):832–838, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Wissmar:1990:COA

[WMD90] Robert C. Wissmar, Diane M. Mcknight, and Clifford N. Dahm. Contribution of organic acids to alkalinity in lakes within the Mount St. Helens blast zone. Limnology and Oceanography, 35 (2):535–542, March 1990. CODEN LIOCAH. ISSN 0024-3590.

Watson:1997:PPT

[WMD97] Susan B. Watson, Edward McCauley, and John A. Downing. Patterns in phytoplankton taxonomic composition across temperate lakes of differing nutrient status. Limnology and Oceanography, 42(3):487–495, May 1997. CODEN LIOCAH. ISSN 0024-3590.

Watras:1995:CMS

[WMHB95] Carl J. Watras, Kenneth A. Morrison, Jodi S. Host, and Nicolas S. Bloom. Concentration of mercury species in relationship to other site-specific factors in the surface waters of northern Wisconsin lakes. Limnology and Oceanography, 40(3):556–565, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Wharton:1994:SOP

[WMM^{+}94] Robert A. Wharton, Jr., Michael A. Meyer, Christopher P. McKay, Rocco L. Mancinelli, and George M. Simmons, Jr. Sediment oxygen profiles in a super-oxygenated Antarctic lake. Limnology and Oceanography, 39(4):839–853, June 1994. CODEN LIOCAH. ISSN 0024-3590.

Wotton:1998:FPD

[WMML98] Roger S. Wotton, Björn Malmqvist, Timo Muotka, and Kristina Larsson. Fecal pellets from a dense aggregation of suspensionfeeders in a stream: an example of ecosystem engineering. Limnology and Oceanography, 43(4):719–725, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Weisse:1990:RML

[WMPC^{+}90] Thomas Weisse, Helga Müller, Ricardo M. Pinto-Coelho, Anette Schweizer, Doris Springmann, and Gisela Baldringer. Response of the microbial loop to the phytoplankton spring bloom in a large prealpine lake. Limnology and Oceanography, 35(4):781–794, June 1990. CODEN LIOCAH. ISSN 0024-3590.

Williamson:1999:DOC

[WMPO99] Craig E. Williamson, Donald P. Morris, Michael L. Pace, and Olaf G. Olson. Dissolved organic carbon and nutrients as regulators of lake ecosystems: Resurrection of a more integrated paradigm. Limnology and Oceanography, 44(3 (part 2)):795–803, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Wilhelm:1996:GIR

[WMT96] Steven W. Wilhelm, Denis P. Maxwell, and Charles G. Trick. Growth, iron requirements, and siderophore production in ironlimited Synechococcus PCC 72. Limnology and Oceanography, 41 (1):89–97, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Wilhelm:1998:EIC

[WMT98] Steven W. Wilhelm, Kevin MacAuley, and Charles G. Trick. Evidence for the importance of catechol-type siderophores in the iron-limited growth of a cyanobacterium. Limnology and Oceanography, 43(5):992–997, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Ward:1993:MSF

[WMTB93a] J. E. Ward, B. A. MacDonald, R. J. Thompson, and P. G. Beninger. Mechanisms of suspension feeding in bivalves: Resolution of current controversies by means of endoscopy. Limnology and Oceanography, 38(2):265–272, March 1993. CODEN LI-OCAH. ISSN 0024-3590.

Ward:1993:RMB

[WMTB93b] J. Evan Ward, Bruce A. MacDonald, Raymond J. Thompson, and Peter G. Beninger. The role of mucus in bivalve feeding — a reply to the comment by Jorgensen. Limnology and Oceanography, 38 (2):467, March 1993. CODEN LIOCAH. ISSN 0024-3590. See [Jør93].

Whitlatch:1998:NDS

[WO98] Robert B. Whitlatch and Richard W. Osman. A new device for studying benthic invertebrate recruitment. Limnology and Oceanography, 43(3):516–523, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Wong:1991:BRR

[Won91] Kuo-Chuin Wong. Book review: Residual currents and long-term transport (R. T. Cheng [ed.]). Limnology and Oceanography, 36 (5):1068–1070, July 1991. CODEN LIOCAH. ISSN 0024-3590.

Wen:1994:EMP

[WP94] Yuan Hua Wen and Robert Henry Peters. Empirical models of phosphorus and nitrogen excretion rates by zooplankton. Limnology and Oceanography, 39(7):1669–1679, November 1994. CODEN LIOCAH. ISSN 0024-3590.

Wing:1995:ISG

[WPAA95] Michael R. Wing, Amy Preston, Nadine Acquisto, and William F. Ahrnsbrak. Intrusion of saline groundwater into Seneca and Cayuga Lakes, New York. Limnology and Oceanography, 40(4): 791–801, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Wells:1994:ILC

[WPB94] Mark L. Wells, Neil M. Price, and Kenneth W. Bruland. Iron limitation and the cyanobacterium Synechococcus in equatorial Pacific waters. Limnology and Oceanography, 39(6):1481–1486, September 1994. CODEN LIOCAH. ISSN 0024-3590.

Whitehouse:1999:CCN

[WPBS99] M. J. Whitehouse, J. Priddle, M. A. Brandon, and C. Swanson. A comparison of chlorophyll/nutrient dynamics at two survey sites near South Georgia, and the potential role of planktonic nitrogen recycled by land-based predators. Limnology and Oceanography, 44(6):1498–1508, September 1999. CODEN LIOCAH. ISSN 0024-3590.

Waker:1995:MNL

[WR95] Marley J. Waker and Richard D. Robarts. Microbial nutrient limitation in prairie saline lakes with high sulfate concentration. Limnology and Oceanography, 40(3):566–574, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Wright:1990:BRN

[Wri90] L. D. Wright. Book review: Nearshore sediment transport (R. J. Seymour [ed.]). Limnology and Oceanography, 35(8):1865–1866, December 1990. CODEN LIOCAH. ISSN 0024-3590.

Wright:1993:CC

[Wri93] H. E. Wright, Jr. Core compression. Limnology and Oceanography, 38(3):699–701, May 1993. CODEN LIOCAH. ISSN 0024-3590. See [CGSD93] and reply [CA93].

Wang:1996:ART

[WRLF96] Wen-Xiong Wang, John R. Reinfelder, Byeong-Gweon Lee, and Nicholas S. Fisher. Assimilation and regeneration of trace elements by marine copepods. Limnology and Oceanography, 41(1): 70–81, January 1996. CODEN LIOCAH. ISSN 0024-3590.

Wakefield:1990:OVM

[WS90] W. W. Wakefield and K. L. Smith, Jr. Ontogcnetic vertical migration in Sebastolobus altivelis as a mechanism for transport of particulate organic matter at continental slope depths. Limnology and Oceanography, 35(6):1314–1328, September 1990. CODEN LIOCAH. ISSN 0024-3590.

Wolfe:1996:GAP

[WS96] Gordon V. Wolfe and Michael Steinke. Grazing-activated production of dimethyl sulfide (DMS) by two clones of Emiliania huxleyi. Limnology and Oceanography, 41(6):1151–1160, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Wheatcroft:1998:IPA

[WSB98] Robert A. Wheatcroft, Victoria R. Starczak, and Cheryl Ann Butman. The impact of population abundance on the deposit-feeding rate of a cosmopolitan polychaete worm. Limnology and Oceanography, 43(8):1948–1953, December 1998. CODEN LIOCAH. ISSN 0024-3590.

Williamson:1996:URN

[WSM^{+}96] Craig E. Williamson, Richard S. Stemberger, Donald P. Morris, Thomas M. Frost, and Steven G. Paulsen. Ultraviolet radiation in North American lakes: Attenuation estimates from DOC measurements and implications for plankton communities. Limnology and Oceanography, 41(5):1024–1034, July 1996. CODEN LIOCAH. ISSN 0024-3590.

Williamson:1996:USF

[WSMS96] Craig E. Williamson, Robert W. Sanders, Robert E. Moeller, and Paul L. Stutzman. Utilization of subsurface food resources for zooplankton reproduction: Implications for diel vertical migration theory. Limnology and Oceanography, 41(2):224–233, March 1996. CODEN LIOCAH. ISSN 0024-3590.

Ward:1998:NEP

[WSNM98] J. E. Ward, L. P. Sanford, R. I. E. Newell, and B. A. MacDonald. A new explanation of particle capture in suspension- feeding bivalve molluscs. Limnology and Oceanography, 43(5):741–752, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Willey:1993:OCE

[WT93] Ruth L. Willey and Stephen T. Threlkeld. Organization of crustacean epizoan communities in a chain of subalpine ponds. Limnology and Oceanography, 38(3):623–627, May 1993. CODEN LI-OCAH. ISSN 0024-3590.

Wilhelm:1994:ILG

[WT94] Steven W. Wilhelm and Charles G. Trick. Iron-limited growth of cyanobacteria: Multiple siderophore production is a common response. Limnology and Oceanography, 39(8):1979–1984, December 1994. CODEN LIOCAH. ISSN 0024-3590.

Wang:1998:VDE

[WTB98] Feiyue Wang, André Tessier, and Jacques Buffle. Voltammetric determination of elemental sulfur in pore waters. Limnology and Oceanography, 43(6):1353–1361, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Waite:1992:DEC

[WTH92] Anya M. Waite, Peter A. Thompson, and Paul J. Harrison. Does energy control the sinking rates of marine diatoms? Limnology and Oceanography, 37(3):468–477, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Warren:1998:MCA

[WTH98] Lesley A. Warren, André Tessier, and Landis Hare. Modelling cadmium accumulation by benthic invertebrates in situ: The relative contributions of sediment and overlying water reservoirs to organism cadmium concentrations. Limnology and Oceanography, 43(7):1442–1454, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Wen:1997:ASC

[WVP97] Yuan Hua Wen, Alain Vézina, and Robert Henry Peters. Allometric scaling of compartmental fluxes of phosphorus in freshwater algae. Limnology and Oceanography, 42(1):45–56, January 1997. CODEN LIOCAH. ISSN 0024-3590.

Waterhouse:1995:TSA

[WW95] Tye Y. Waterhouse and Nicholas A. Welschmeyer. Taxon-specific analysis of microzooplankton grazing rates and phytoplankton growth rates. Limnology and Oceanography, 40(4):827–834, June 1995. CODEN LIOCAH. ISSN 0024-3590.

Welker:1998:CMC

[WW98] Martin Welker and Norbert Walz. Can mussels control the plankton in rivers? — A planktological approach applying a Lagrangian sampling strategy. Limnology and Oceanography, 43(5):753–762, July 1998. CODEN LIOCAH. ISSN 0024-3590.

Wafar:1990:NRC

[WWD90] Mohideen Wafar, Sayeeda Wafar, and Jason J. David. Nitrification in reef corals. Limnology and Oceanography, 35(3):725–730, May 1990. CODEN LIOCAH. ISSN 0024-3590.

Wong:1998:ENP

[WWMI98] C. S. Wong, F. A. Whitney, R. J. Matear, and K. Iseki. Enhancement of new production in the northeast subarctic Pacific Ocean during negative North Pacific index events. Limnology and Oceanography, 43(7):1418–1426, November 1998. CODEN LIOCAH. ISSN 0024-3590.

Wilhelm:1998:RSR

[WWSJ98] Steven W. Wilhelm, Markus G. Weinbauer, Curtis A. Suttle, and Wade H. Jeffrey. The role of sunlight in the removal and repair of viruses in the sea. Limnology and Oceanography, 43(4):586–592, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Willey:1993:PEZ

[WWT93a] Ruth L. Willey, Robert B. Willey, and Stephen T. Threlkeld. Planktivore effects on zooplankton epibiont communities: Epibiont pigmentation effects. Limnology and Oceanography, 38(8):1818– 1822, December 1993. CODEN LIOCAH. ISSN 0024-3590.

Witkowski:1993:IKG

[WWT^{+}93b] Konrad Witkowski, Leszlek Woliński, Zbigniew Turzyński, Danuta Gedziorowska, and Andrzej Zielifiski. The investigation of kinetic growth of Chlorella vulgaris cells by the method of integral and dynamic light scattering. Limnology and Oceanography, 38(7):1365–1372, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Willows:1998:IIB

[WWW98] Robert I. Willows, J. Widdows, and R. G. Wood. Influence of an infaunal bivalve on the erosion of an intertidal cohesive sediment: a flume and modeling study. Limnology and Oceanography, 43(6): 1332–1343, September 1998. CODEN LIOCAH. ISSN 0024-3590.

Wyman:1992:VME

[Wym92] Michael Wyman. An in vivo method for the estimation of phycoerythrin concentrations in marine cyanobacteria (synechoccus spp.). Limnology and Oceanography, 37(6):1300–1306, September 1992. CODEN LIOCAH. ISSN 0024-3590.

Xenopoulos:1997:VSY

[XB97] Marguerite A. Xenopoulos and David F. Bird. Virus à la sauce Yo-Pro: Microwave enhanced staining for counting viruses by epifluorescence microscopy. Limnology and Oceanography, 42(7):1648– 1650, November 1997. CODEN LIOCAH. ISSN 0024-3590.

Xue:1995:CCZ

[XKS95] HanBin Xue, David Kistler, and Laura Sigg. Competition of copper and zinc for strong ligands in a eutrophic lake. Limnology and Oceanography, 40(6):1142–1152, September 1995. CODEN LIOCAH. ISSN 0024-3590.

Xue:1993:FCI

[XS93] HanBin Xue and Laura Sigg. Free cupric ion concentration and Cu(II) speciation in a eutrophic lake. Limnology and Oceanography, 38(6):1200–1213, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Yavitt:1992:MFC

[YAF^{+}92] Joseph B. Yavitt, Lisa L. Angell, Timothy J. Fahey, Christopher P. Cirmo, and Charles T. Driscoll. Methane fluxes, concentrations, and production in two Adirondack beaver impoundments. Limnology and Oceanography, 37(5):1057–1066, July 1992. CODEN LIOCAH. ISSN 0024-3590.

Yager:1999:PMA

[YD99] Patricia L. Yager and Jody W. Deming. Pelagic microbial activity in an Arctic polynya: Testing for temperature and substrate interactions using a kinetic approach. Limnology and Oceanography, 44(8):1882–1893, December 1999. CODEN LIOCAH. ISSN 0024-3590.

Yamasaki:1998:SPN

[YFF^{+}98] Akiko Yamasaki, Hideki Fukuda, Rumi Fukuda, Toshihiro Miyajima, Toshi Nagata, Hiroshi Ogawa, and Isao Koike. Submicrometer particles in northwest Pacific coastal environments: Abundance, size distribution, and biological origins. Limnology and Oceanography, 43(3):536–542, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Yund:1991:CTT

[YGB91] Philip O. Yund, Steven D. Gaines, and Mark D. Bertness. Cylindrical tube traps for larval sampling. Limnology and Oceanography, 36(6):1167–1177, September 1991. CODEN LIOCAH. ISSN 0024-3590.

Young:1992:GMF

[YH92] Leah Bendell Young and Harold H. Harvey. Geochemistry of Mn and Fe in lake sediments in relation to lake acidity. Limnology and Oceanography, 37(3):603–613, May 1992. CODEN LIOCAH. ISSN 0024-3590.

Yamamuro:1993:NMF

[YK93] Masumi Yamamuro and Isao Koike. Nitrogen metabolism of the filter-feeding bivalve Corbicula japonica and its significance in primary production of a brackish lake in Japan. Limnology and Oceanography, 38(5):997–1007, July 1993. CODEN LIOCAH. ISSN 0024-3590.

Yamamuro:1994:DCN

[YK94] Masumi Yamamuro and Isao Koike. Diel changes of nitrogen species in surface and overlying water of an estuarine lake in summer: Evidence for benthic-pelagic coupling. Limnology and Oceanography, 39(7):1726–1733, November 1994. CODEN LI-OCAH. ISSN 0024-3590.

Yamamuro:1995:RDD

[YK95] Masumi Yamamuro and Hajime Kayanne. Rapid direct determination of organic carbon and nitrogen in carbonate-bearing sediments with a yanaco MT-5 CHN analyzer. Limnology and Oceanography, 41(4):1001–1005, July 1995. CODEN LIOCAH. ISSN 0024-3590.

Yamamuro:1995:CNS

[YKM95] Masumi Yamamuro, Hajime Kayanne, and Masao Minagawao. Carbon and nitrogen stable isotopes of primary producers in coral reef ecosystems. Limnology and Oceanography, 40(3):617–621, May 1995. CODEN LIOCAH. ISSN 0024-3590.

Yee:1996:VSZ

[YM96] Donald Yee and François M. M. Morel. In vivo substitution of zinc by cobalt in carbonic anhydrase of a marine diatom. Limnology and Oceanography, 41(3):573–577, May 1996. CODEN LIOCAH. ISSN 0024-3590.

Yoshii:1999:SIA

[YMT^{+}99] Koichi Yoshii, Natalia G. Melnik, Oleg A. Timoshkin, Nina A. Bondarenko, Pavel N. Anoshko, Takahito Yoshioka, and Eitaro Wada. Stable isotope analyses of the pelagic food web in Lake Baikal. Limnology and Oceanography, 44(3):502–511, May 1999. CODEN LIOCAH. ISSN 0024-3590.

Yahel:1998:PDG

[YPF^{+}98] Gitai Yahel, Anton F. Post, Katharina Fabricius, Dominique Marie, Daniel Vaulot, and Amatzias Genin. Phytoplankton distribution and grazing near coral reefs. Limnology and Oceanography, 43(4):551–563, June 1998. CODEN LIOCAH. ISSN 0024-3590.

Yen:1991:FCE

[YSSO91] Jeannette Yen, Brian Sanderson, J. Rudi Strickler, and Akira Okubo. Feeding currents and energy dissipation by Euchaeta rimana, a subtropical pelagic copepod. Limnology and Oceanography, 36(2):362–369, March 1991. CODEN LIOCAH. ISSN 0024-3590.

Yeats:1990:CFF

[YSW90] P. A. Yeats, P. M. Strain, and B. G. Whitehouse. Cross-flow filtration of colloids from aquatic environments. Limnology and Oceanography, 35(6):1368–1375, September 1990. CODEN LI-OCAH. ISSN 0024-3590.

Young:1993:BMC

[YW93] Stephen Young and Penelope Watt. Behavioral mechanisms controlling vertical migration in Daphnia. Limnology and Oceanography, 38(1):70–79, January 1993. CODEN LIOCAH. ISSN 0024-3590.

Zagarese:1998:CMI

[ZCGP98] Horatio E. Zagarese, Walter Cravero, Pablo Gonzalez, and Fernando Pedrozo. Copepod mortality induced by fluctuating levels of natural ultraviolet radiation simulating vertical water mixing. Limnology and Oceanography, 43(1):169–174, January 1998. CO-DEN LIOCAH. ISSN 0024-3590.

Zohary:1994:SSC

[ZEG^{+}94] Tamar Zohary, Jonathan Erez, Moshe Gophen, Ilana Berman-Frank, and Mariana Stiller. Seasonality of stable carbon isotopes within the pelagic food web of Lake Kinneret. Limnology and Oceanography, 39(5):1030–1043, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Zimmer-Faust:1993:APP

[ZF93] Richard K. Zimmer-Faust. ATP: a potential prey attractant evoking carnivory. Limnology and Oceanography, 38(6):1271–1275, September 1993. CODEN LIOCAH. ISSN 0024-3590.

Zimmer-Faust:1996:BCP

[ZFdSY96] Richard K. Zimmer-Faust, Mark P. de Souza, and Duane C. Yoch. Bacterial chemotaxis and its potential role in marine dimethylsulfide production and biogeochemical sulfur cycling. Limnology and Oceanography, 41(6):1330–1334, September 1996. CODEN LIOCAH. ISSN 0024-3590.

Zimme-Faust:1994:CIE

[ZFT94] Richard K. Zimme-Faust and Mario N. Tamburri. Chemical identity and ecological implications of a waterborne, larval settlement cue. Limnology and Oceanography, 39(5):1075–1087, July 1994. CODEN LIOCAH. ISSN 0024-3590.

Zaneveld:1993:VSP

[ZKM93] J. Ronald V. Zaneveld, James C. Kitchen, and James L. Mueller. Vertical structure of productivity and its vertical integration as derived from remotely sensed observations. Limnology and Oceanography, 38(7):1384–1393, November 1993. CODEN LIOCAH. ISSN 0024-3590.

Zhou:1998:RSA

[ZMP98] Jian Zhou, Kenneth Mopper, and Uta Passow. The role of surfaceactive carbohydrates in the formation of transparent exopolymer particles by bubble adsorption of seawater. Limnology and Oceanography, 43(8):1860–1871, December 1998. CODEN LI-OCAH. ISSN 0024-3590.

Zohary:1998:BDS

[ZPHH98] Tamar Zohary, Utsa Pollingher, Ora Hadas, and David Hambright. Bloom dynamics and sedimentation of Peridinium gatunense in Lake Kinneret. Limnology and Oceanography, 43 (2):175–186, March 1998. CODEN LIOCAH. ISSN 0024-3590.

Zohary:1998:ESM

[ZR98] Tamar Zohary and Richard D. Robarts. Experimental study of microbial P limitation in the eastern Mediterranean. Limnology and Oceanography, 43(3):387–395, May 1998. CODEN LIOCAH. ISSN 0024-3590.

Zweifel:1995:DDO

[ZWH^{+}95] Ulla Li Zweifel, Johan Wikner, Åke Hagström, Erik Lundberg, and Bosse Norrman. Dynamics of dissolved organic carbon in a coastal ecosystem. Limnology and Oceanography, 40(2):299–305, March 1995. CODEN LIOCAH. ISSN 0024-3590.